A Bibliography of Publications about SGML, the Standard Generalized Markup Language: 1981–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org, beebe@ieee.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 April 2024 Version 2.34

Title word cross-reference

+ [ML97]. 21/2 [PSTW97]. 3 [And96b, And96c, BG95a, BG95b, CLPR96, CJ98, Dan98b, DLSR98, DM97, HMRL95, ID99, KS98a, LMM96, LCM96, MVKJ99, Mau98, MD98a, MD98b, MB95b, Nic97, RJN95b, RJN95a, Ros97, SS95d, Sun97, Vac96b, VT98, Van98, VB95c, WSoSS98, WS99, War98]. F [AS93a, AS93b]. $GML [Ens97].

-an [Roe87]. -Distribution [AS93b]. -Percentiles [AS93a].

0-470-21126-1 [BC90].

1 [VRM97]. 1-2 [Via98b]. 1.0 [BLFF96]. 1.1 [FGM^{+}97, LO^{+}97]. 1.2 [LO^{+}97, LO^{+}99b]. 1.5 [Oly96]. 1/2 [Win97a]. 101-1983 [Gra83]. 10th [ACM92, Ano89c, Mei97, MR99b, RS97]. 11th [ACM93a, BFH98, CL97, FS99, Rii97, Vez96]. 12th [Ano96a, PE95]. 13th [ACM95b, BCG99, IEE97a, Lit96]. 14772 [VRM97]. 14772-1 [VRM97]. 14th [ACM96d, Rai90, Wil93a]. 15445 [Intxx]. 15th [ACM97a, Ano99a, K^{+}98, Wil94b]. 16th [HW94, IEE96i, Rai92]. 17th [Ano90d, RJ93]. 18 [Ivl95]. 1866 [BLC95]. 1867 [NM95c]. 1874 [Lev95]. 18th [Ano99f, Bee97b, HW96b, IDK98, RJ94, Wil97b]. 193rd [Ano98b]. 1942 [Rag96b]. 1945 [BLFF96]. 1980 [Sei96]. 1986 [Int86c, Int86d]. 1986-10-15 [Int86b]. 1988 [Int88h]. 1990 [HV92]. 1991 [HV92]. 1994 [Ano94a, Ano94n]. 1997 [Int97]. 19th [Ano90d, Ano98a, Ano98m, RJ95]. 1st [Cai94a, Cai94b].

2 [LO99a]. 2-4 [ACM95b]. 2.0 [Ass87b, Ass87c, Ass89, Ano96-44, CB97, CMP97, Gra96a, HW97, STM97, Tan98, Van98, WNB97]. 2000 [Blu97]. 2068 [FGM^{+}97]. 2070 [YNAD97]. 20th [Ano95a, Ano98u, Wil99a]. 2110 [PH97a]. 21st [Ano96n, Ano97-27, Ano97-46, Cze98, Jac98]. 22.50 [BC90]. 22nd [Ano97-31, Ano99h, RMG98]. 23rd [Ano98t]. 24-26 [ACM99b]. 25th [Ano96c, Ano96-39]. 26th [Ano95h, Isk96]. 29th [Ano94n, GSG98]. 2D [ETT97]. 2HTML [Dra94]. 2nd [Ano95v, Joh95b, Rol98, Wil97a].

3 [Ano95-43, Ano95b, Aro96, Bry98, CNH99, Hen99, Nag99, Pel97a, Pel97b, SBS95a]. 3-D [Ano95-43, Ano95b, CNH99, Hen99, Nag99, Pel97a, Pel97b, SBS95a]. 3-space [AP96]. 3.0 [DM96, How96]. 3.2 [Pat97]. 3/4/1991 [Ame91]. 30th [Ano97-43, ERPNS97a, ERPNS97b]. 31st [Rol98]. 3305-19 [WP98]. 3648-15 [Mor99a]. 3654-22 [Mau98]. 3654-25 [VMM98]. 36th [Ano97-42]. 37th [Ano96x]. 3rd [Ano90g, Ano99d, BP96, Gup97, RM95b, Ano92a].

4 [Bro98b, BH98b, Bry99, Cas98, CH99a, Cea98, Cla98, Dar97, Dar98, D^{+}99, Dre97, Dre98, Hol99b, Hol99a, HUW98a, LO99a, LO^{+}99b, LD97a, LD97b, LD98, LT98, LT99, Mul98, OH97a, OH97b, OH97c, Oli98a, Oli98b, Pat98, PS98b, Pfa97a, Pfa98a, PG98, Phi98a, Phi98b, PW99, RLAK98, RR98b, RR98c, RR99a, RR99b, RR99c, San99, SSB98, Ste98, SMiPM98, SM98b, Tay98, TJ98, TSP98, TJD98, TPM99, Via98a, Via98b, WM98, ZFK97]. 4.0 [Eva98a, Eva98b, Gra98, LO^{+}97, LO98, Mac98, MP98b, MP98c, MP99, Ram98, RR97b, SM98a, WF98b, MP97, MP98a, MM99b, Nat98, Nat99a, Ram99, Rob98]. 4.x [MK98d]. 4.X. [MK98e, MS98b, MK98f, MK98g, MK98h, MK98i]. 400 [Ano96y]. 40th [Ano96s]. 41st [Ano94r]. 42nd [Ano95-37, Ano98r]. 43rd [Ano96b]. 44th [Ano97b]. 45th [Ano98u]. 46th [BB94]. 48th [Ano95k]. 4th [Ano93a, Bow97, COW98, HAB98, HB98, MNW98, SC98, Ass96, ZBW94, Ano88a, Gau96b].

5-8 [MDK98]. 5-9 [Ano98r]. 5-star [H^{+}96]. 5.0 [Bar98a]. 500-178 [MBB90]. 53rd [Hen90]. 56th [Bon93]. 57th [Ano97a]. 5th [Ano88f, Ano96-33, ABG96, IEE96j, KAKZ96, LLY97, PK94b, Rol98, RSV99, USE97a].

6 [Web95b]. 6-10 [Ano96-34]. 6-8 [MM97]. 6-in-1 [Sta97]. 63rd [Ano97a]. 6th [Ano93j, Ano95e, Ano98n, Ano99b, BBKF95, CL99, GJTW97, RT95, RM95b].

7 [New90]. 7-th [Ska99]. 7.0 [CN98]. 724 [AS93a]. 77th [Ano96-40]. 7th [Cha96b, CBK98, HT97, HAB98, Ten99].

8 [TD95]. ’88 [Ano88d, Ano88f]. 8879 [Rie94, BC90]. ’89 [Ano89c]. 8e [Bos97]. 8th [Bos97, Bou97b, BDV95, Ibr95, Nad90].

9 [Tow95]. 9-11 [Jac98]. ’90

[Ano90g, Ano90h, Hen90, VD90]. 9000 [DGG^{+}92]. ’91 [Lic91, Mau91]. ’92 [ACM92]. ’93 [Ano93c, Ano93j, Bon93, Sch93]. ’94 [ACM94a, Ano93e, Ano94f, Ano94l, Ano94o, BGG^{+}94, IEE94a, IEE94b, Kho94, PK94a, ZBW94, Wid94, LR94]. ’95 [ACM99a, ABHY96, Ano95e, Bea95b, BH95b, BDV95, IEE95b, KR95, LM95b, NM95b, RT95, Rut96a, Ste95a, Gan96, NF96, WF98b]. 95/Java [Sha95a]. 95/Windows [Mac98]. 9573-11 [Int92a, Int92b]. 9573-13 [Int91c]. 95th [Ano98d]. ’96 [ACM96b, ACM96c, ACM96d, Ano96v, ABG96, BBJR96, GSV96, HK95b, IEE96g, IEE96k, IEE96m, IEE96j, KAKZ96, RM96a, Ass96, VC96, Win97b]. ’97 [ACM99b, AR97, Ano97-37, Ano97-42, AA97, DG97, HF97, Küc97, M^{+}97, PMS97, RS97, SSK97, USE97a, Wae97, YH97, IEE97c, Car97, Ven97]. ’97s [IEE97e]. ’98 [ACM98, A^{+}98, Ano98a, Ano98m, Ano99b, COW98, EAZ98, HAB98, Heu98, HS98, IEE98c, IEE98e, LB98a, LMT98, MNW98, NS98, NDM98, QSBC98, Ska98, SC98]. ’99 [FS99, Ska99, Ten99, ABB99]. 9th [BH95b, FJ92, IEE98e, QSBC98, RM96a, Ste96d].

= [Bos97, CBO93, HiKgK98, HUW98a, Int86b, Int88c, Int88d, Int88i, Int88e, ML97, Mos96].

A-18 [Ivl95]. A/E/C [Kho94, VC96]. AAAI [Ano98q]. AAAI-98 [Ano98q]. AACR2 [Duk94]. ABCD [Als95a]. ABC’s [Per96]. Abductive [MK96a]. Abstract [EAG^{+}95, BR96b, RC95]. Abstracts [IEE95a]. academe [Cop96]. Academic [ACM96d, FG95, KH95b, Rid95, Cox96, KLM94, MM97]. Academy [PKM99]. Accelerated [HLCG97]. Access [Ano93g, BMK95, CFMG95, CHH^{+}95, Dav98, EAG^{+}95, GB98, GB95a, GFR98, HW96b, JMB95, Law98, PR98a, PR95b, Pri96c, Pro95, RP98, RWBS97, Rot97, SBD^{+}95, SZM^{+}95, SMM95a, ZY95, BEE^{+}94a, BEE^{+}94b, BEE96a, BDH^{+}95, CH98, CW97, Dar95, Hoc93a, Jam97, Kah96, KPD96, LG94a, MSIM95, MS95, New92, PD96, PK95a, SDK^{+}96, Sea96b, SMM95b, WE92, WT93]. access/reference [LG94a]. Accessibility [Ano97q, CV98, BEE95]. Accessible [GB98, RW98, Sha97, For96a]. Accessing [KCFP95, NS96, OA98]. Accessories [GNSP95]. accompany [Wil97a]. Achievement [Cea96b]. ACM [Ano93s, Sol88, AR97, DGBH93, HF97, JM96]. ACM/SIGAPP [DGBH93]. Acquisition [DR95d, Ano95v, Col96a, Dar95, FS99]. Acrobat [Ano93u, Gra95d, Kar96]. Across [KCFP95, Ano95-43, BDK95, Sha97, Soa98]. ACSM [Ano97a]. ACSM/ASPRS [Ano97a]. act [Ano97e]. Actes [Bos97, CBO93]. action [Mor96, PS98a, Par99, Pet98a, SP99a, Wel95b]. Active [Ano96-27, Cov99, HJM95, Lil95, MFVG99, Wid94, Wit96b, TB96b]. ActiveX [Kro99, Man97, Vin98]. activities [App89, Bar97, Bar98b, BT99]. actualizada [Alv97b]. Acumen [Kro99]. ad [COW98]. Ada [ACM96c, NF96]. ADAPT [HW90]. Adaptability [PKT95]. Adaptative [MGVS97]. Adapting [GMR95, STM97, STMM98]. Adaptive [SN95, EH96]. ADB [LR94]. ADB-94 [LR94]. ADC [SBGO99]. ADC’96 [Ano96d]. add [Ano95b, KH95a, Men95a]. add-on [KH95a]. add-ons [Men95a]. Adding [Ano94b, Ano94g, BS95a, BE95, DM97, DM96, MYBP98, MB95a, SMG95a, SMG95b, SMG96]. Adds [Whi96e]. ADL [ABHY96, IEE96m]. ADL’99 [IEE99a]. Admin [Bun96]. Administration [Mue95, USE95, VS95]. Administrators [Gel95]. Adobe [Sys97, Ano96-32, Boh96, DM96, Gra96e, KS95a]. Adopting [Faw89, Rid95]. adoption [SR93]. Adopts [Cla98]. ads [HH96, HB96, Wid94]. Advance [Ano88d, Ano95k]. Advanced [ADE93, CLPR96, CJ95, Dav92, Gar99, LM95b, Mic97, Nat99a, Uns96, Via97c, Via97d, ABH93, CL99, Gar98a, HT97, Mul96b, NS98, Ram98, Ram99, RW96a, SR97, SRSC98, TCLV92]. Advances [ABG96, IEE96m, IEE99a, MB97b, SC98, ABHY96]. Advancing [Wor97]. advantage [Dun95e, Ret96]. advantageous [Bat96d]. Advantages [Lun95a]. Adventures [Pow94]. advice [Ros96]. Aerospace [Far96b, IEE93a, Ano90d]. AF [Spi96a]. AFC [Wat99]. Affect [Fun99, Ano99l]. afraid [How99]. After [Ano96-46, PSTW97, Win97a, And97, GMR95]. AGARD [Ano91a]. AGARD-CP-487 [Ano91a]. Age [WD95, Kik96]. agencies [Ano94t]. Agent [Ano96-31, GTM99, GMR96, GL99]. agent-based [GTM99]. Agents [Ano96-31, Cod98, Eic95b, WT95, Eic95a]. Aggregate [PH97a]. aggregation [SN95]. agility [Jac98]. AGORA [Rot97]. agricultural [WB96]. Agriculture [Ano95f]. AHA [Jam97]. ahead [Ano97-55]. AI [RR96c]. Aid [Ple95]. aided [LLY97]. aiding [CR97]. aids [Eft90]. AIM [Fer93a, Fer93b]. AIP [JJ95b]. Air [Ano99b]. aircraft [Zim95]. Ajaib [Ano97c]. AL [Ano95v, Gau96a]. alarm [HKM^{+}96]. Albuquerque [BBJR96, Bor96, CL97]. Algebraic [Küc97]. Algorithm [BW97, PR95b, BCLT95, AS93a]. Algorithms [Len93, SN95]. Alice [MNA91a, MNA91b]. ALLC [Ano88b]. Alliance [Law98]. Alloys [INIK96]. Almost [Pau95]. along [Mat98]. Already [CK97]. alter [Tub95]. Alternatives [Bee96c, Ano96-32, Rob96]. Alto [BBKF95, NW97]. Amalgamating [YIU96]. Ambiguity [MGH93, Mat93]. amendement [Int88e]. amendment [Int88e]. American [AC93, Ame91, Ano98b, Ano98c, Ano98d, BCG99, CL97, Chu98, Org91, Whi96f]. Americas [Gup97, HB98]. Ameritech [Hen96b]. Ames [IEE96n]. AMIA [Cim96, Mas97]. Amiga [FMW95]. Among [BSF98]. AMPHORE [Sul96]. AMS [FLSG95]. Amsterdam [ACM93c, Van93a, WV89]. Anaheim [Ano98u, VC96]. Analysis [Ano93p, BDV95, COW98, CV98, CGW96, FH97, Fer93b, FK96, IEE93b, IEE94b, IEE95f, IEE97f, KN95b, KW97c, PC97, ZMP96, AHB98, BAC93, Bur99a, COVS95, Gra95b, GL93, HW90, HG96a, IS98, HG96b, KN95a, Kir91, LDW^{+}96, Min96, MB97b, Nag95, TC96]. analysts [Ano96-42]. analytical [Ano97q]. Analyzer [Con96]. Anatomical [WF98a]. anesthesia [IDK98]. Angeles [ACM94a, Ano96-39, BT97]. angina [CNH99]. anglais [BKP97]. anglais-francais [BKP97]. animal [Ano97-52, Gol98]. Animating [Lem97]. Animation [OMY97, BCLT95, Fil94]. Animations [TRMR97]. animés [Dil97]. animierter [Amm97]. Anniversary [MR99b]. Annotated [Ano99c, CB97, Lap99, DuC99]. Annotation [Uns96]. Announce [Wat95]. Announces [Lot95]. Announcing [van90a]. Annual [ACM92, ACM93a, ACM94a, ACM95b, ACM96d, Ano88f, Ano88g, Ano89c, Ano93j, Ano93p, Ano93s, Ano94r, Ano95a, Ano95h, Ano95k, Ano95m, Ano95-37, Ano96c, Ano96x, Ano96b, Ano96-36, Ano96-39, Ano97a, Ano97b, Ano97-43, Ano97-42, Ano97-31, Ano98a, Ano98d, Ano98o, Ano98u, Ano98t, Bee97b, Bon93, BBJR96, Bor96, BCG99, CL97, Chu98, FJ92, Gra96d, Hen90, IEE96c, IEE97g, IEE98c, Isk96, Len93, Mas97, Mau93, Mur92b, PE95, PK94a, PK94b, RPWW96, Tea97b, Tea98b, Ass96, USE97a, ACM96a, ACM97a, Ano96s, Ano98m, Ano98r, Cim96, Gar95, GSG98, IEE96d, USE96, USE98]. Ansatz [Rie94]. ANSI [BBG^{+}95b]. Answers [Kle94, FMSDW93, WM92]. Ant [Ano96e]. Anti [Arn95]. Anti-net [Arn95]. Antibes [Ano97h]. Antipiracy [Cla98]. Anwenderpraxis [KS96]. Anwendungen [Ano99j, ER97]. Aonix [Kro99]. aout [CBO93]. API [SB95]. app [Ano95g]. apparatus [Sea94]. Applet [Mee96, SMDO98, Gol96]. Applet-Based [SMDO98]. Applets [ER97, AS96, Got98, HL97]. appliance [APB^{+}99]. appliance-independent [APB^{+}99]. Application [ABO98a, ABO98b, Ano94n, Ano95p, Ano96-31, FR97, Fre98, Gre98b, Int92a, Int92b, KW97c, KH95b, Lam96, OSE97, Ost94b, SBGK98, Thr95, TTS97, Zha95, Zim95, Ass86a, Ass86b, Ass87a, Ass87b, Ass87c, Ass89, Ano94p, BMY95, BK99, Dan98b, Dra96, Hon95, KPIE99, Mau98, NTI91, OSEMV95, OSEM^{+}96, PS88, Roe87, RM95b, SRB96, Smi87a, Smi87b, YT96]. Application-Oriented [ABO98a, ABO98b]. Applications [Ada95, Als97, Ano91c, Ano97g, Ano97z, Ano98q, Ano99d, Ano99j, BP96, BR95, Bro97, BKS96, Coy97, EDLS98, Gar95, GRA92, IEE95g, IEE96d, IEE96c, IEE97g, IEE98c, Ivl95, Lan96, LMT98, LM95b, LR94, Mal98, McG98b, New95b, Nic97, Nie99, QSBC98, RW90, RW92, SV98, SD98, SGJ^{+}95, Ste95a, Thr95, TR92, TW98, UL98, VT99, Whi96e, WHC95, vES95, Ano93q, Ano97t, Ano98e, BD96, BdRD^{+}98, BRR94, CLPR96, CG86, CG87, CL99, CB96, DM95, FMFVN98, Fon93, Fra99, Fuc96b, Gol96, Gol98, HS98, IEE97e, IEE98b, Jah99, Lib98, LLF98, LG97, LC98, MTU99, Mic99, Moh99, MM97, Neg92, RW96a, RT95, RR96c, SGP96, Spi96a, SC99, To98, Tra97, Vig89, Wel95b, Xu97, Ano90b, Ano94a]. applications-technology [Gol96]. Applied [CH99b, DGBH93, Ano98n, Ano97d]. Applying [AS93b, EBR^{+}96]. Approach [ABO98a, ABO98b, Ano97q, Ano97-41, Bee97a, DeV96, Fre98, Kes95, PC97, BMY95, Bus88, CG95, Cha93, CB95a, Chr95, Con93, FX93, Gol81, GB95c, GL99, Hue98, MM90a, NR96c, VMM98, Wal90, Won90, ZR93c]. Approaches [Res95, Ano96-43, Fra99, GW94]. Approaching [MKBB95, Won92]. Approval [Cla98, Ano95l]. Aprende [Fer97]. April [ACM98, Ano95a, Ano95-37, Ano96-33, Ano97a, Ano97h, Ano97x, Ano97-43, Ano98s, Ano99h, CL99, Coy97, DE95, GSV96, HAB98, Lav97, NP96, RM95b, SSB97, Ten99, Van93a, van88]. aquatic [CBK98]. Arbitrary [DP95a]. Archaeology [Mos96]. archeologia [Mos96]. Architectural [Cam98]. Architecture [APP^{+}98, Boh95, Coy97, LHMM97, MKNS87, SOUH97, Ube94, WMGL95a, Dan98b, Jah99, MG94a, MD98b, Mol91, Wal90, WMGL95b, WH98, Yam94, JGKR96]. Architectures [Heb96, Kam90b, Mas98, SP99b]. Archive [RB95a, SBGO99, Sea96a]. Archives [Can95]. Archiving [SBGK98, Bla90]. Area [FK96, MFSF95, SBS95b]. argumentational [Kir91]. Arguments [Whi96c]. Ariadne [TSY^{+}95]. Arizona [DG97, USE87, YH97]. Arles [Van95b]. Armidale [Ano98n]. Array [SMM95a, SZM^{+}95, SMM95b]. arrival [Ano96-42]. Art [DGBH93, Gol95, Hum95, Res97, IDK98, Kam90a, MM90b, Wil98, W^{+}98b, Whi96a]. Article [TA99]. articles [Kir91, Van89]. Artificial [ACM93c, ACM97b, Ano90b, Ano97z, Ano98q, GRA92, IEE97d, Ste96d, Ste96e]. Artist [Fis95]. artistic [HAB98]. artistry [IZ98]. Arts [Van98, BE99]. ASCII [Bar94]. ASCILITE [PE95]. Asia [HT95]. ASIS [Bon93, Hen90, LG97]. asked [Bow95]. askew [Ros96]. ASME [Min96]. ASP [Moh99, Spe99]. Aspects [DAOD95, FJ92, KAKZ96, SSH87, She95, EBR^{+}96, Smi96a]. ASPRS [Ano97a]. Assembly [Buc97]. Assessment [FW96, WD91, BC95, CB95b, WD90]. Assets [ACM98, Ens97, ACM94a]. assignment [COVS95]. ASSIST [Spi96a]. Assistance [Pro95, Hel96, PB88]. Assistant [Ano95-44, Roa95, Har95a, PAV96, PD96, RS96, Ros96, GF93, KH95a]. Assisted [BBJR96, Bor96, Bon92, WZ91]. Assistive [ACM94a, ACM98, EAZ98]. Association [Ano96f, Ano98c, Ano98m, Ano99e, Ano99h, CBK98, Chu98]. Associative [DWV95]. assurance [Bow96]. Astronomical [AHB98, Ano98b]. Astronomy [Ano95x, EH95b, Str95, EH95a, VN95]. Astrophysics [BSB^{+}95, RW95b, MTH99, EAG^{+}95]. AstroWeb [Jac95]. AT&T [Hol95]. ATA [HL95a]. Atelier [Isa98a, Bos97]. Athina [CBK98]. Atlanta [Ano93j, Ano96a, GE95, Joy98, LG94b, Min96]. ATM [NN97]. Atmospheric [IEE94b, JW96, WAO^{+}95]. Atria [IEE97e]. Attribute [BS90]. Auburn [Ano95v]. audience [Ano94t]. Audio [Jam97, Jam98, Uhl95, Oze96]. Audrey [Pri96b]. August [ACM96a, Ano90c, Ano98m, CBO93, EAZ98, Gup97, HB98, IEE95f, IEE96d, IEE96f, IEE97f, IEE97g, IEE98c, IS98, Kum96, PKM99, QSBC98, Rii97, SSK97, TW98, Van95b]. Austin [Ano98b, Bou97b]. Australasian [Ano96d]. Australia [ACM97b, Ano98n, Ano98s, DE95, K^{+}98, Lav97, OW95, SV98, HT95]. Australian [DE95, DE96, OW95, PE95, BS94]. Austria [EAZ98, HK94, IEE98c, KAKZ96, Mau95, QSBC98, TW98]. AusWeb95 [DE95]. AusWeb96 [DE96]. Authenticity [Gol95]. Author [Bry88, Ano93n]. Authoring [Ano95p, Ano97p, Ano97-51, Dav93, EDLS98, Hla96, KS95a, Kes95, LG94a, Mar97a, AMC95, Ano92c, Ano97s, AOS91, BVN95a, BVN95b, Car95b, Col95b, Gar98a, Gar98b, Lew95, MG94a, MR93a, Pao96, QRV95, Wea97, Web95b, WF98b, Whi96a, ZR93a]. Authority [Pet95]. Authorization [LZ95]. Authors [Fly95, Wil96a, Ano96-45, Smi87d, Smi87e]. Auto [Ano93b, Ano97-34, Whi95b, Whi95a]. Auto-FAQ [Whi95b, Whi95a]. Auto-Graphics [Ano93b]. Autobahn [Oly96]. automata [BK93c, Neu97b]. Automated [Cha95a, DeV96, DC98, Sta98b, WP98, CL90, HW90, LS96]. Automatic [AMC95, Aho95, Fin95b, McG92, ACM96e, BRR94, Bus88, COVS95, Cha93]. Automatically [JMB95, O’K95]. Automating [Fly96, Jon95, GW94, LHW96, MBO93]. Automation [Hig96, MM97, SSB97, YKT90]. Automotive [Rol98]. AUTOTESTCON [IEE94a]. Availability [Eng94, Sch88, Cel92, HW94]. Available [Ano94c, Vac96b]. Average [Mue95]. Avignon [ABG96]. Avilon [Kro99]. Aviv [Ano99d, BP96, IEE98e]. Aware [Ano94n, Ano94p].

Baan [Kro99]. Babel [Kor98]. Back [Ano96-29, Mag97, GS95a, LLL^{+}91]. Background [Ada95, NN92]. Backstage [Ano96g, Ano96g]. Bad [LM98, Len93]. Balancing [Kro99]. balisage [Int86b, Int88e]. Ballina [DE95]. Baltimore [HB98, IEE99a]. band [SM93a]. bandwagon [Ano95-27]. Bandwidth [RMR97]. Bankers [PSTW97, Win97a]. Banking [Dra95b, Dra95a, Lun96, Pri95]. bar [Mee96]. bar-chart [Mee96]. Barbara [IEE98a]. Barcelona [Ano97-37]. barrier [PKM99]. Base [Can95, Fer93a, San96, Smi91, Yok93]. Based [Ano95p, Ano97w, Cha95b, Coh95, DCG^{+}95, Deb97, Div98, Dud97, ETT97, Fab95, Fer93a, FFKO99, GFR98, HTL95, HY97, HH99b, Isk96, KHW97, Kes95, KK98, LK95, MB97a, Neg95, OSYM97, PR98a, PR98c, PR98b, PR95b, Rak95, Ric98, SS89, Sch97b, Shi98, SMDO98, Sun98a, SN96, USO97, VT99, Var95, W^{+}98a, WR99, WF98a, WHC95, ABB99, BA90, BUP99, CLPR96, Coh99, CB96, DK97, Dra95b, Dra95a, EJ95, FW93, FMFVN98, GF93, GTM99, GL99, Kah96, Kel98, KCCJ95, KPIE99, Lew97, LG94a, LD95b, LHW96, LC98, MMV95, MSW97, Mar94, ME99, MCPD95, Mau98, MBH95a, MBB95, Mil93, MBM93, NM95c, NKN91a, NKN91b, NR96c, RR94, SP99b, SC98, Wan99, Wit86, WKD96, WH98, Yam93, Yam94, ZR93c, ZMP96, ABB99, KSJ95, AT88, BHK^{+}88, BG96, CT96, FG95]. based [Int97, INIK96, Jan96, JRK95, JRS96, Lig95, MW92, MBH95b, MG94b, Sch95b, VGGB95]. Bases [Mar95a, V^{+}96]. Basic [Con96, KPIE99, Law98, BH98a, HLL93, Buc97, MRR99, Ubo96]. Basis [Ano94c, Rat95, RH97, Wie94, Hay93]. Basque [IEE95c]. Bath [BD97]. Baton [NP96]. Battin [HW96b]. Battle [Ano96h, Bal96b]. BBEdit [Bar98a, Sof97]. BC [IEE98b]. be [Wil96b]. Beach [IEE97d]. Become [Cea96a]. Becoming [Kel94a, Kel94b, Vac96b]. Beginner [OR96, Que97, Que98, Ano97o]. Beginners [TRMW99, WCM99]. beginnt [Bat96c]. Behavior [MB95a, MN95b, NM95a, Ric96]. Behaviour [Ric98]. behind [Ano95l]. Beijing [C^{+}98]. Being [Fis95, Ano92c, Wel95b]. Belgium [Ano90d]. Benefits [MBO93, Ano97e, Glu96]. Bergen [Ano96f]. Berlin [BFH98, Cov99, HS98, IEE96h]. best [Dun95e]. Beta [AS93b]. bets [BS98]. Better [And96d, Law98, LB96, Mue95, Ste95c, Ano93r, Ano96-32, Dob97a]. Between [Ano96h, CAD^{+}95, Tes96a, CM90, Kum96, PW94, STA98a]. Beyond [Ano94a, Ano98e, Cim96, Del95a, GMG^{+}95b, GMG^{+}95a, Kar96, Lib98, Mil95, Nag99, RC95, RC96, Rub96, RM96b, Spe95, Ven97, Jac96, KSGJ96, KDH96, Res97, RM97b]. BeyondPress [Ano95d]. bible [Har99, PG98, T^{+}96]. biblical [Bot92]. bibliographical [Cou95]. Bibliography [CDB91b, CDB91a]. biblioteca [JSY98]. Bibliothek [Neu96, RH97]. Bibliotheken [Neu96]. Bicycling [CCEE98]. Bielefeld [Neu96]. Bielefelder [Neu96]. bien [VB96]. Biennial [CBK98, IEE96g]. Big [Ano95-28, Gar99]. BIIF [Car97]. Billboards [Blu95c]. billion [Ens97, GSE^{+}94]. billion-record [GSE^{+}94]. Binary [Gar97a]. binding [VP95]. Biocomputing [A^{+}98, HK95b]. biological [Tay95a, Tay95b, Wu98]. Biology [EKS95, PF95, PWHK98, SNS95]. Biomedicine [LMT98]. Bioscience [MKEB98]. birth [Adl92]. bis [Neu94]. black [PMSC98, PMK98]. blind [Kah96, KPD96, KK94, PK95a]. Blinded [MM99a]. Block [Gen97, Spi96a]. Blue [Gar99]. Blueprint [TW95]. Blueprints [Jac96, KDH96]. Blur [Wei94]. BNC [Bur99a]. Board [Ano97w, HKH95]. bodies [RFM95]. BOF [Sha95a]. Boiling [Ano96l, Ano96r]. Bolt [Gra96a]. Bolts [Lun95a, Lun95c]. Bombay [V^{+}96]. Bonestell [Hum95]. Book [Ano98f, BC90, BB94, Daw98a, Far96a, HV92, LSW95, Pie97, PH92, Pop92, Pri96a, Swa96, Tho97, USO97, CG95, DeR97, FS98, Glu96, PSG92, PMSC98, PMK98, SP96]. bookmarks [Ano95-46]. Books [Ano96l, Lut98, Aok99]. Boost [Par95]. Borealis [MM96]. Boston [Ano95h, CM96, Coh95, USE97a]. Botanic [FMC95]. Both [Sun98a, Mac96a]. Bottle [Ano97r]. Boulder [BCG99]. Box [CMP97]. braces [Ano96-42]. Brain [PKM99, CNH99, RSV99]. Brazil [MZZ96, MB97b, dN97]. Brazilian [MB97b, dN97]. Breakdowns [Hen95]. breeding [Gol98]. Brick [Fer94]. Bridging [Ano91a, Lav97, ZR93b]. Brief [SM93b, Bar99, Red99]. Briefs [Cla98, Gar99, Law98]. Bring [Gar97a]. Bringing [Ano98c, BE95, Vac96a]. brings [Ano96-42]. Brisbane [Ano98s]. Britannica [Val96]. British [DP98b, MG94c, Neu96, Tea97b, Tea98b]. broad [SM93a]. broad-band [SM93a]. Broadway [Mio96]. brochure [Huc97]. Browse [Ano96-27, RS96]. Browser [APP^{+}98, Ano96h, BG95b, EH97, HS99, JOA98, Law98, OSE97, POA^{+}95, RW96b, SM95, Sun98b, Ano92i, Ano95-29, BG95a, BHS^{+}96, CB96, KCCJ95, Lev96c, Wil99c, DO96]. browser-based [CB96]. Browsers [GNSP95, Lin96, Tus97, GK95b, GK95a, GG96, Ret96]. Browsing [AVFC98, Gra95b, Koh95, Law98, NN97, Spe95, Whi96e, CLS96, Dob97a, HG96a, INIK96, HG96b, Liu97a, Pes95, VMM98]. Brussels [Ano90d, Ano97y]. Bryant [Wil98]. BSDIA’97 [MB97b]. Buchhandler [Neu96]. Buchhandler-Vereinigung [Neu96]. Budapest [EAZ98, SSB97]. Build [NS95a, Pau95, Pel97a, Ano95b, BC96, HN96b, Mee96]. builder [NTI91]. builders [Bal96b]. Building [ACM96d, Ano95c, Bai93, Bea95a, Blu95a, Can95, CF99, DMH98, For96c, Hei96, Hob98, HH99b, Key94, KMS99, Lau94, Lew97, Mar95a, MGBW96, MD98a, McG98b, PCL^{+}95, Pel97b, PGRC97, SC99, TRMR97, Tho97, Thr95, TSWS96, T^{+}97, Tra97, Ano99d, BP96, DPB91, DM95, HW96a, HW97, MD98b, Pes95, WB96, Yok93]. buildings [For99]. Bulgaria [Dob98]. Bulletin [HKH95]. Buoys [SMM95a, SZM^{+}95, SMM95b]. bureaucratique [Int88c, Int88d]. bureautiques [Int86b, Int88e]. Business [Ano97e, BHK^{+}95a, Cha96a, LRG96, SF96, Wel95a, Win97b, Blu95a, Rub92]. buy [Ric96]. buyer [GPE98]. byte [Fic98b].

C [Ano97w, Kho94, Pri96b, Buc97, KMK95, KK95, Mat97, Ret96, VC96, Xu97, McG92]. C-ODA [KMK95]. C/C [Ret96]. CA [ACM99b, Ano96-39, Ano97x, Ano97-31, Ano98a, BT97, Bee97b, BBKF95, MWEK97, NDM98, Ten99, USE95, Ass96, Whi96f]. cable [Ubo96]. Caching [DHR97, PR95b, BC95, CB95b, GS95c]. CAD [Ano96i, PR98b, Ros97, Ano96-29]. CAD/CAM [Ano96i]. CAE [Ano96-29]. Caere [Gan96]. Café [Ano97-51]. CAI [MSMY97, NANK96]. CAI-SYSTEM [MSMY97]. Cal [Van98]. Calculate [AS93a]. Calculus [Gog90]. CALIBER [MM97]. CALIBER-97 [MM97]. CALICO [BBJR96]. California [ACM94a, ACM98, Ano98u, Bea95b, BE99, BFM96, BUP99, CEG96, CC99, FJV96, IEE95b, IEE96e, IEE96j, IEE97d, IEE98a, KJR98, LB98a, LZ98, NM95b, NW97, SSK97, SB97, SR98, USE89, USE96, USE97b, USE98, VC96, W^{+}98b, Goo94]. Caligari [Fis96b]. Call [Ano90c, SR93, Sra96]. calls [Gol96]. CALS [Ano93c, Ano93j, Ano95e, CH90, Hen91, IN96, Kik96, Lev93a, Orl93, Per96, Smi91, Sum89, SIF^{+}96, Zim95]. CALS/CE [Orl93]. CAM [Ano96i]. Cambridge [Ano96-33, Tea97b, Tea98b]. Campos [dN97]. Campus [LK95, RL95a, Sea95, She95, VD98]. Campus-Wide [RL95a, She95]. can [Ano95-35, Ano97k, Tit98]. Canada [Ano97b, Ano99i, Bos97, DP98b, FW96, IEE95f, IEE98b, JM96, MG94c, WNC95, KCFP95]. Canadian [Bos97, VS95]. canadien [Bos97]. cannot [BEE95]. CANSIM [KCFP95]. Capabilities [Col95d, Ano96-32, Dav92, FMFVN98, PS96, Sum89]. capability [Dun95e]. CAPS [AW93, Eng93, WT93]. capture [VY88]. Car [KK98]. Card [AB95]. Care [Gar95, KD98, Kis95, MMRS98, SGJ^{+}95, Tea97b, Tea98b, Chu98, ECvR93, GJTW97, IDK98, Mas97]. careers [Mat97]. Caribbean [Wil98]. Carlo [GSV96]. Carolina [RS97]. carrefour [Ano96t]. cas [BK99]. Cascading [LB97, RR98a, BO98, JT97, JT98, LB99a]. CASCON [BGG^{+}94]. Case [ABB99, Ano97c, Ano97f, Boa95, Cam98, Can95, Dra95b, EGH95, FK96, GK95c, HY96, Per95, Shi98, Wil93c, BIL95, BMN94, Bur99a, Dra95a, GB95a, GS95c, Kir91, SC98, VN95, TTN93, ABB99]. Case-Based [Shi98, ABB99, SC98, ABB99]. Cases [ZJ96]. Casting [Mil95]. Castle [FS99]. Catalog [SS95a, FGGT96, HB95, LMO^{+}96]. Cataloging [Gay94, Pit95]. catalogs [HH96, HB96]. Catalogues [Tau96]. cataloguing [Dra96]. Catamaran [LB98a]. Catch [Not96]. Cathedral [BUP99]. Caught [Pri95]. CAVEview [MCPD95]. CBLIS [Cha95b]. CBR [HC99]. CD [Ano97g, Daw98b, Dod94, DuC97, DuC98, Fee98, GB94, Gre93, GG96, Nak94, NR96b, NR96a, Neo96a, Neo96b, Neo97, NR97, Neo98, Neo99, Ros88, Sad97, She88, Val96, Via98b, Weg96, Wil97a]. CD-ROM [Dod94, Fee98, GB94, Gre93, GG96, Nak94, Neo97, Ros88, Sad97, Via98b, Wil97a, Weg96]. CDF [Pet98c]. CE [Ano93c, Orl93]. CEFC [IEE96g]. CELEX [MNA91a, MNA91b]. Censorship [RB95a]. Center [Ano94r, Ano95m, Ano96b, Ano98a, Col95d, IEE97e, MTH99, Sea94, Ano96-42, Sea96a, Van98]. centered [Ano96s, Cha93]. Central [Int92a, Int92b, SSB97, Ska98, Ska99, TS96]. Centre [Ano88g, Ano97b, Bos97, CBK98, IEE98b, IEE97a]. Centres [CBK98]. Century [Ano97-46, IEE94a, Ano96n, Ano97-27, Jac98]. CERN [DR95d, Nic92, NL95, Van95b]. Cernobbio [M^{+}97]. Certificate [Cea96b]. Certification [Ano91b]. CGI [Gar97b, Coh96, DG95, DG96, Fra97a, GJ96, Gol98, LO96, LOA^{+}96, Lad96, Liu98, SFO97, SD96b, Tit95a, T^{+}96, TGHE96, T^{+}xx]. CGI/Perl [Coh96]. CGM [HL95a]. CGMs [SF90]. CGs [ME99]. Chairman [Win97b]. challenge [Rat98]. Challenges [IEE95b]. Chameleon [Wal90]. chance [Rat98]. chances [Sch95b]. Chandigarh [MDK98]. Changes [Bro94a, Bro94b, DB96, DBCK96]. Changing [Wei96b, ZY95, IEE95b, Wei96a]. channels [Pet98c, Pur98]. Chaos [ACM95b]. Chapter [Web95b]. Character [Ano97-36, Gay95]. characterization [BC95, CB95b]. characters [Bir96b].

Charles [HV92]. Charlotte [Ano98d, Ano99h, RS97]. Chart [CRBB98, Mee96]. Chatting [Bir96a]. cheap [Rud96]. Check [Ano96-30, Ber95]. Checker [Kro99]. Cheese [SK98]. Chelgraph [Cad87]. Chemical [Col95c, MRR99, STM^{+}94, WR95, Col98, Hic89, LG94b]. Chemistry [Bac96, PF95, RL95b, Ano94f, CLPR96]. Cheshire [LMO^{+}96]. Chesley [Hum95]. Chicago [Ano95-34, Ano98r, Gra96d]. Chichester [BC90]. chih [Neu98]. Chile [IEE97c]. China [C^{+}98, IS98, LLY97]. Ching [mWfC97]. Chip [Law98, KS96]. Chiron [VN95]. Chisel [Bry98]. Choboja [OuP97]. Choosing [Ano96j, GPE98]. Christmas [Bjo96c]. chrono [Fra97b, WM98]. chulpan [HiKgK98]. CIESIN [ZY95]. CIKM [ABY94, BFY93]. CIMBLE [TTG98]. cinema [Dan98b]. Circuit [ID99]. CISST [AA97]. Citation [Ebe98]. City [ACM97a, Ano99f, IEE97b, Isk96, Mei97, BD97, IDK98, Sav95]. Civil [Coy97, LH96, Kho94, Kum96, VC96]. claim [BVN95a]. claims [BVN95b]. Clara [Ano97x, IEE96e]. Clarifications [Bjo96a]. Claris [SDSS95]. Class [HTL95, GI96, Rii97]. Classical [HY96]. Classroom [WCP^{+}95b, KDC95, Mey99, PBB^{+}96, WCP^{+}95a]. Classrooms [LaR95]. Clean [Rag98]. Clearinghouse [Mac95, SS95c]. Clearly [Vac98]. CLI [VP95]. Clickable [Dav95a]. Client [CH98, Sco96a, Sei95, Sei96, WHC95, Ber95, CLS96, Dob97a, HL97]. Client-Side [Sco96a, Sei95, Sei96, Dob97a, HL97]. Client/server [CH98]. Clients [Sla95b, Lev96c, Sla95a]. Climate [BB95a, Col95d, HD95]. Clinical [DC98, JD96, OSE97, ZJ96, Gar95]. Clockwork [Mun95]. closer [Ano95-43]. Closing [Wid94, Win97b]. Clothing [Lam96]. clouds [BCG99]. CML [Deb97]. Co [Min97, Die96, WD91, WD90].

Co-existence [WD91, WD90]. Coast [DE96]. CoBase [CMB93]. Code [CGLV94, Fry95b, IZ98, NL95, Ret96, Mol95b]. CODE-consistent [Mol95b]. codebook [HG96a, HG96b]. Coding [Dav95b, Ano93d, Hol87, O’C97, Tau97]. Cognitive [Ano93a, Ran98]. coherence [Kah95, SL97]. Cold [Oly96]. Collaboration [BLG^{+}95, FLF95b, RW95a, Bon92, FLF95a, Fuc96a]. Collaborative [CPR98, IEE96j, LCLL98, MBH95a, MG94a, MBH95b, MR93a, MR93b, ZR93a]. Collaboratory [WAO^{+}95]. Collaborwriter [MG94a]. collar [KTK^{+}95]. collection [FMSDW93, Wil98]. Collections [O’K95, MS95, SMC^{+}96]. Collector [LSW95, PD95]. College [Cha96b, JRV96, She95, Str95, Ano97-43, GB94]. Colleges [Mei97]. COLLIDE [HLCG97]. Collision [HLCG97, Tow95]. Colloquium [Ano91e, Ano93g, Ano93h, Ano94b, Ano94g, Ano94h, Ano95p, Ano96u, Ano97-50]. Color [Ano98g, Ano98j, Ano98h, Ano98k, Ano98i, BE99]. Colorado [Ano97d, BCG99]. ColorSync [Ano95l]. CoLoS [CT96]. Columbia [DP98b, MG94c]. Columbus [Bon93]. COM [Cha99, FFKO99]. COM/MTS [Cha99]. combination [Mac96a]. Combinators [WR99]. combines [Wag97]. Combining [DWV95, DL95c, Fer93a]. Combo [Cas96c]. Comes [Ano96-30, Bat96d]. Coming [Ano94d]. Command [Whi96c, Ano97-27, Mac90a]. Command-Line [Whi96c]. Commands [Har95b]. Commentary [Sch96b, Smi96a]. Commerce [Lut98, McG98b, BB99a, GTM99]. Commercial [HAD^{+}95, TB96a, Dan98b]. Commodity [FFKO99]. Common [Fry95b, BCG99, NL95, Ano96-32, FJ96]. communicate [Mor99a]. Communicating [Ano93a, SNS95, Ste95c]. Communication [ASWS98, Ano91a, Ano94r, Ano95-37, Ano96k, Ano96b, Ano98u, Dav93, DP95b, GS87, Ano93k, IEE96f, IEE97b, KCFP95, Ran98, Wil94a, ACM97a, Chu92, GL99, WT93, MR95]. Communications [Ano96-30, LSC96, Pol95, PS95, Ano96-42, Jac98, Lav97, Ano96t, Ano90g]. Communicator [Dim96]. Community [Bos99, HRS95, Kee95, LHMM97, Mag94, VS95, GW94, GSG98]. Comp.specification.z [Bow95]. companies [Ano92c, Ens97, Mat97]. Companion [Bra96b, Mud98, Ven96b, Bra98, Bra99, GR99a, SR97, SRSC98]. company [d’H95a]. comparaison [NB91]. Comparative [ZMP96, Xu97]. Compares [Glu96]. Comparison [Aok99, GFR98, Huc97, Lev96c, Wil92]. compatibility [WZ91]. COMPCON [IEE92, IEE96e]. Compelling [CMP97, MGBW96]. competition [M^{+}97]. Competitive [Ano97h]. Complete [Daw98b, GR96, Gra97b, Gra98, Hol98, McF98, BO98, C^{+}99, Gra96c, KH97, McF96, McF97a, Pow98a, Pow99, RV99, SWC99]. Complex [CHH^{+}95, Pri97, Rad98, Ano94v, Gol97]. Complexity [ACM95b]. Compliance [Ber96c, Ber95, Ber96b]. Compliant [OISEM97]. Complicated [Tan98]. Component [CK95, Kro99, Sug96]. Components [NF96, NF97]. Composants [Cha99]. composed [War98]. Composing [Ano88c]. Composition [ETT97, VT98]. Compound [Wil93b, Cro92]. Comprehension [CGLV94]. Comprehensive [CDB91a, Car98a, Car99, Man97]. compression [Gil92]. COMPSAC [IEE96d, IEE98c]. COMPSAC’97 [IEE97g]. Computation [IEE96g, Küc97]. Computational [Bac96, CW95, Gar99, PF95, SV98, ZJEP95]. Computations [LFS97]. Computer [ACM93c, ACM96a, ACM97a, Ano89c, Ano92h, Ano93f, Ano93a, Ano93h, Ano96l, Ano98l, Ano98m, Ano99f, BKS98, BBJR96, Bor96, Bot93, Cha95b, Che97, CJ95, Coy97, DCG^{+}95, EAZ98, FJ92, Gar95, Hah91, IEE92, IEE96d, IEE96e, IEE97g, IEE98c, IEE98a, Joy98, Kle96, LLY97, MZZ96, MR95, Mel99, MWEK97, Owe95, Reb95, SSH87, SS89, SSK97, Ska98, Ska99, Tea97b, Tea98b, TS96, Van95c, VD90, WZ91, dN97, FWDV93, FA94, KPD96, Mar94, MR93b, NANK96, Soa98, Sun97, Wil92]. computer-aided [LLY97]. Computer-assisted [WZ91]. Computer-Mediated [MR95]. computer-readable [FWDV93]. Computerized [LG94b, MSW97]. Computers [Ano96f, Ano96c, Ano96-39, Ano99e, EAZ98, KAKZ96, PE95, SSH87, ZBW94, OW95]. Computing [Ano89a, Ano96f, Ano96-33, Ano98a, Ano99e, DBE^{+}95, DGBH93, FJV96, HMMSS95, IEE95g, IEE96l, IEE96i, IDK98, Mei97, Mos96, Nad90, OMY97, Uns96, Van95b, VC96, Ano94e, Ano97v, Ano97-43, CH90, Gar95, GB95c, KJR98, SSK97, Ten99, VB95b, Kho94]. ComSec [Ano90g]. con [McF97b]. Concept [DWV95, Per96, Shr95, Ano99d, BP96]. Conception [Pet90]. Concepts [FB95b, ABH93, BT99, Bar99, Bra92, Ell94, NB91, SWC99, SCW99, SVC99, VY88]. Conceptual [Cov96, Gre98b]. Concerns [Mar95c]. Concise [Bra96b]. Concurrent [Deb97, SMH99, FWDV93]. Condition [Ric95c]. Conference [ACM92, ACM93a, ACM93b, ACM93c, ACM94a, ACM94b, ACM95b, ACM95c, ACM96b, ACM96d, ACM97b, ACM98, ABY94, AR97, ABB99, Ano88b, Ano88e, Ano88d, Ano88f, Ano88g, Ano89a, Ano89c, Ano90a, Ano90h, Ano91b, Ano91c, Ano91d, Ano91f, Ano93e, Ano93f, Ano93a, Ano93i, Ano93s, Ano94e, Ano94m, Ano94o, Ano94r, Ano95e, Ano95k, Ano95m, Ano95-34, Ano95-37, Ano96d, Ano96a, Ano96b, Ano96-34, Ano96-35, Ano96-33, Ano96-31, Ano96-36, Ano97b, Ano97z, Ano97x, Ano97-42, Ano98a, Ano98m, Ano98q, Ano98s, Ano98u, Ano99e, Ano99h, Ano99a, ABG96, AA97, Bac96, BKS98, Bea95b, BFY93, BB94, Bow97, BH95b, BFH98, BDV95, BKS96, BUP99, BCG99, COW98, CM96, Cha96b, Cha95b, CL99, CL97, Col95c, Coy97, DG97, DE95, DE96, Dés86, DP98b, FD91, Fur90, Gra96d, HAB98, IEE92]. Conference [IEE93a, IEE93b, IEE94a, IEE95a, IEE95f, IEE96d, IEE96e, IEE96a, IEE96f, IEE96b, IEE96g, IEE96k, IEE96l, IEE96i, IEE96c, IEE97a, IEE97c, IEE97b, IEE97f, IEE97e, IEE97d, IEE97g, IEE98c, IEE98a, IEE98d, IEE98b, IEE98e, IEE98f, Ibr95, IR94, Isk96, Jac98, JM96, K^{+}98, KAKZ96, Kum96, LB98a, Lav97, Lic91, LM95b, Mau93, Mau95, MTH99, MG94c, M^{+}97, Mur92a, Mur92b, MDK98, NDM98, WWW95, PK94b, PS95, QSBC98, Rii97, RPWW96, RM95b, SSH87, SS89, SSK97, SSB97, SV98, Ska98, Ska99, Sol88, Su96, Ten99, TS96, To98, TR92, USE87, USE89, USE95, USE98, VD90, VD98, Vez96, V^{+}96, Wae97, W^{+}98a, WNC95, ZBW94, van88, ACM96a, ACM97a, Ano90c, Ano93j, Ano95a, Ano95j, Ano96f, Ano96n, Ano97d, Ano97h, Ano97v, Ano97-31, Ano98d, Ano98o, Ano98t, Ano99b, Ano99f]. conference [Ano99i, BBKF95, Cai94b, Col98, Dud97, ERPNS97a, ERPNS97b, Gau96a, GSG98, Gup97, Heu98, HB98, HS98, IEE95b, JRV96, LLY97, Lit96, LR94, Mei97, MR99b, Ner97, NS98, OW95, PE95, PK94a, RT95, RSV99, Tea97b, Tea98b, USE96, Ano88g, Ano92a, Cai94a]. conferences [Ano96-40]. Conferencing [DL95c, MKBB95]. Conferenece [LMT98]. configuration [Gol96]. Conformance [Ano91b, AC93, Ame91, Lew91]. conforming [Ass86a, Ass86b, Ass87a, Ass87b, Ass87c, Ass89]. Congrès [CBO93].

Congress [CBO93, EAZ98, Kho94, Min96, Sch93, Ano96v, Ano98c, Ano98n, GJTW97, VC96]. conjuction [PK94b]. conjunction [Kho94, VC96]. connected [SDK^{+}96]. Connecting [GSG98, LaR95]. Connection [Ano96i, LDWN96]. Connections [RMR97]. connectivity [SBM96]. conquer [And96b]. Consequences [Kli95]. Considerations [MB91, Dun95g]. consistency [Bot93, Mol95b]. consistent [Mol95b]. Consortia [SS95c]. Consortium [BBJR96, Bor96, Mei97, Ano96-44]. constraint [MBM93, NR96c]. constraint-based [MBM93, NR96c]. constraints [Erf93]. Construct [Tho97]. Constructing [HPS95, CB96]. Construction [Cam98, Coy97, FD95, FD96, Har94, Jon95, LHL97, OBO97, TCLV92, AMC95, ACM96e, Ano97n, FS96, KMA^{+}96, Mac98]. Construire [BK99]. consultants [EG90, EG91, Wil95]. Consumable [Ano97-56]. Consumer [IEE96a, IEE98d]. containment [DSDT96a, DSDT96b]. Content [Aho95, Aho97, Ano97q, BKW93, BKW97, CMP97, Eyz95, Fry95a, Fun99, PR98a, PR98c, PR98b, PC97, SD98, Ano99l, FMFVN98, GL99, Kil99, Neu97b, OM96]. Content-analytical [Ano97q]. Content-Based [PR98a, PR98c, PR98b, FMFVN98]. Contents [Ano99c, Lap99]. Context [EtR94, EDLS98, Shr95, JI95]. Context-free [EtR94]. Contexts [RC96]. Contextual [BR95]. Contrasting [Ano97-34]. contributed [LG97]. Contribution [For96b]. Control [ASWS98, Mac93, SHW95, Ano97-27, BBKL94, HT97, NTI91, Sea96b, Wol91]. Controlled [Sla94, AOS91]. Convection [BB95b]. Convection-Diffusion [BB95b]. Convegno [Mos96]. Convenience [Ano92h]. Convention [Ano94r, Ano95m, Ano96b, Ano97a, Ano98a, MM97]. Conventional [DWV95, LG95]. convergence [Wes99]. conversational [HG96a, HG96b]. Conversion [Ano97-56, Fly96, Hig96, Hun94, LM98, Nic97, Yam93, Yam94, ZR93d, COMG96, Rah95, SG93, SHO^{+}96, WP98, ZR93c]. Conversions [Gro97]. Converter [Ret96, Ano95d]. Converting [Ano97-30, For95b, JK98, Thr95, KKK97]. converts [Ano95g, Ano95r]. Cookbook [Ano99c, Jel98, H^{+}96]. Cool [Hei96, SS95d, Tay95e, Tay96, Tay97a, Mar99, Mor96, Tay95c, Tay95d, Tay97b, Tay98, MD97]. cooled [DGG^{+}92]. cooling [DGG^{+}92]. Coollists [Lim95a, Lim95b, Lim96]. Cooperation [Bon92, TS96]. Cooperative [DBE^{+}95, Pao95, M^{+}97]. Coordinate [KW97c]. copies [Ano90d]. Copy [Fer93a, Hen95]. copyright [WT93]. Copyrighted [Mar95c]. CORAL [RRS^{+}93]. Corba [Cha99, Dos99b, FFKO99, HK95a]. Corba/OTS [Cha99]. Core [Gre98a]. Corner [Ano94o, Sch95c]. Cornerstone [CFMG95]. coronary [MWVE99]. Corporate [ACM96d, Ano93d, DP95b, HPS95, PD95, Qui95, RDC99, Ano96-42, CS96, Gro95]. corpus [Dun95g]. Corrections [Bjo96a]. correctness [Mol95b]. Cosmo [Ano96-45]. COSSC [JCDA95]. Cost [Ano97i, IEE94a, Gar95]. cost-effective [Gar95]. Costly [LFS97]. Costs [Gro97]. Could [Gar97a]. Council [Neu96]. Courageously [Win96]. Course [Cea96b, CJ95, KK95, LG95, OSE97, Bar99, Cea98, Lin99, ZJEP95]. Courses [MHP98, Owe95]. Courseware [Div98, HMMSS95, MHMS^{+}94, SS95b, Tew95, Ben97, Mar95b]. Courtz [GMR96]. cover [New90]. Covers [Ano96-29, Kar96, Gra97a]. CoWeb [JGKR96]. CoWeb-architecture [JGKR96]. CP [Ano91a]. creacion [Alv97a, Alv97b, HC97]. Creando [McF97b]. Create [Pau95, SBS95a, Tes96b, Ano95-42, For96a, Gol97, HH96, HB96, HH98, H^{+}96, Mor95a, Ric94, Ubo96, Vac96b]. Creating [Ano97p, BH98a, CMP97, CPW95, DP98a, Fis97, Fry95a, GN95, JMS98, KN95b, KN95a, Loc97, Lun95b, McC97b, PR95a, Sco96a, SD98, Ste96b, SS95d, Tay95c, Tay95d, Tay95e, Tay96, Tay97a, Tay98, TS99, WS95a, BB99a, Car98b, Car98a, Car99, Gut98, McF96, McF97a, McF98, Pow94, Pur98, Tay97b, WF98b]. Creation [BL98, Wei94, Coh96, CJ90, Dun95f, HC97, How96, RMF93]. Creations [HV92]. Creative [WWEI97, WW98]. créez [San99]. Crescent [DG97]. Crete [NS98]. Critical [Pet95]. criticism [CR95b]. critics [Mac96a]. critique [BBG^{+}95b]. Croner [Ano97j]. Cross [CAD^{+}95, MOTA95, Ski94, ACD^{+}95, Wil99c, HS99]. cross-browser [Wil99c]. Cross-Browser-Lösungen [HS99]. Cross-Institutional [MOTA95]. Cross-platform [Ski94]. cross-references [ACD^{+}95]. Crossroads [ACM97a, IEE97b]. Crucible [Cov96]. CS [FB95a]. CSEE [Ibr95]. CSRs [Sra96]. CSS [Kor98, LS99, LB99b]. CSS2 [Gla99]. ctive [Int94]. CU [Ano96l, Ano96r]. CU-SeeMe [Ano96l, Ano96r]. Cues [Ran98, Kah95]. Cultivating [Ano95f]. Cultural [Bus97, Del95a, Fis95, RB95a]. Curitabia [MB97b]. current [BB91, Not96]. Curriculum [Kee95, JRK95, JRS96]. Custom [WHC95, ZJ96, WS95a]. customer [Sra96]. Customers [FNN97, HK95a, Sra96]. Customised [vSSW93]. customized [CJ90]. cutting [Ano97-40]. cutting-edge [Ano97-40]. Cyber [Gib98]. CyberCareers [MM98]. cyberculture [BKP97]. Cyberleaf [Ano95r, Ano95-44]. Cybermouse [SVOM95]. cybernetics [COW98]. CyberPassage [Ano96-27]. cyberserials [CL97]. Cyberspace [Pes96b, Flo96, Kau96, LJ96, Pes95, Whi95b, Whi95a, Ano96-42]. cyberspace-3-D [Ano96-42]. cybertheatre [Sha98]. Cycle [OBO97, Cha95a]. Cyclone [Ano97w]. Cyprus [NS98]. Cyrillic [Bir96b]. Czech [GSV96, PK94a, Ska98, Ska99]. Czechoslovakia [Nad90].

D [BG95a, SS95d, And96b, And96c, Ano95-43, Ano95b, Ano96-42, Anoxx, BG95b, CLPR96, CNH99, CJ98, Dan98b, DLSR98, DM97, Hen99, HMRL95, ID99, KS98a, LMM96, LCM96, MVKJ99, Mau98, MD98a, MD98b, MB95b, Nag99, Nic97, Pel97a, Pel97b, PSG92, RJN95b, RJN95a, Ros97, SBS95a, Sun97, UL98, Vac96b, VT98, Van98, VB95c, WSoSS98, WS99, War98]. D-CAD-Formats [Ros97]. Dagstuhl [FS99]. Dallas [Ano99a]. Dance [Tub95]. Danish [PKM99]. DAPHNE [SF90]. Data [ABS99, Adk88, Ano91e, Ano95q, Ano96-37, Ano97-30, Ano97-56, Bar95, BS95a, CHH^{+}95, CJ98, DR95d, EJ95, Fer93a, FS95, FB95b, Gel95, Gil92, GK95c, Gro93, HD95, HB95, IEE94b, IEE96o, JMB95, JM96, K^{+}98, KMS99, KFHU98, LB98b, Nag99, NK99, Ple95, PD95, RP98, RWBS97, Sha97, SBGO99, SK98, SS95a, SMM95a, STM^{+}94, Str95, Su96, Sun98a, Tes96a, Tow95, VH95, V^{+}96, ZY95, AHB98, Ano96-42, Ano97h, AMFL^{+}95, BB93, CH98, CNH99, Dun95f, Fra99, Ges91, Gre93, Hay93, LG94b, MBO93, MSIM95, Ore88, Orl93, RTW96, Rei93, San96, SG93, SZM^{+}95, SMM95b, Spi96a, Spr96b, Sta98b, Vau94, WB96, Yee99]. data-intensive [Fra99]. data-sets [CNH99]. Database [ACM94c, Ano91c, Ano92b, Ano96d, Ano97-41, Bla94, CP93, CL99, Fin95b, FB95b, Fra95, GB95a, GFR98, Gra95a, HLY^{+}94, KCFP95, LM95b, Mal98, MP94, NS95a, OISEM97, OSEM^{+}96, Pri96c, QSBC98, RT95, Ric98, SDAMZ95, SD96a, SD98, TABA^{+}94, TR92, TW98, TA99, Toy98, Wid94, Wil93c, Zha95, ABH93, Ano94v, Ano97s, Aok99, ABG96, BCK^{+}94, DJ96, IEE96o, Ish94b, IN96, Jad97, Kim94, LLL^{+}91, Neg92, New96, NTR89, OKT92, OM96, OSEMV95, RRS^{+}93, Sak94, Uza92, VY88, Whe96, WF89, WF90, INIK96]. Databases [Ano96-37, Ano97k, Ano97u, DWV95, MYBP98, ML97, Neg95, OL97, PR98c, PR98b, AMFL^{+}95, BMN94, DPB91, DB95, Gon94, GSE^{+}94, HKM^{+}96, Hic94, Ish94a, LR94, Loe94, LZSD95, LG94b, MNA91a, MNA91b, Mil93, NTI91, Ner97, NS96, O’C92b, O’C92c, Rod92, Sch95b, Wei93, Whi96b, WSM93, YIU96]. DataBasics [O’C92a, O’C92b]. DataBlade [Ube94]. DataViz [Ano95g]. Datenbanksystem [RH97]. DAV [Coh99]. Davenport [Dou92a]. Days [Cea96a, Lem96a, MC97a, Lem95b, LD97a, LT98, LT99, NH99]. DBMS [DAE^{+}94, DAOD95, Sto93]. DC [Kho94, ACM96b, Ano95k, Ano95-37, Ano97-27, BFY93, Cim96, IEE97g, LMT98, Oke93, Oke95]. DCE [LZ95]. Debrecen [Ano99e]. Debugging [OMY97]. débuter [VB96]. Dec [PE95]. Decades [PSTW97, Win97a]. December [ACM96c, ABY94, Ano97y, Ano99d, BP96, LLY97, NM95b, Oke93, RJ95, RMG98, Sol88, USE97b]. Decentering [Pet95]. d’échange [Int88c, Int88d]. Decision [Ano95h, Eng94, KD98, KS95b]. Declarations [MGH93]. declarative [LSS96]. Decrease [HLS97]. DECwrite [CM90]. Dedicated [Ano89c]. deductive [RRS^{+}93]. defence [Ano90d]. Defining [Ano97k, CHP88]. definitions [MB90, MB91, PB88]. Definitive [Ano98f, Goo98, KFZ96, MK96b, MK97, MK98j, RLA96, WM99, Pes95, ZFK97]. del [ACM94a, ACM98, sta89, JSY98]. Delaware [FMC95]. Delay [Cla98]. Delegating [Tau96]. Deliver [Lot95, Tow95]. Delivering [Ano96m]. Delivery [BB95c, SGJ^{+}95, TB96a, CJ90, Cze98, De 94, Dou92b, FA94, Hen91, Lig95, Wal92].

DELPHI [DR95d]. Demand

[DP98c, Ano96-32, OSEMV95, OSEM^{+}96]. Demands [For96c, For99]. Demise [Mun95]. demo [CNH99]. demonstrated [Wel95b]. d’entités [Int91d]. Denver [Ano97d]. Department [Str95, KLM94, Sti94]. departments [Sch95b]. dependent [MP95a]. deploying [Hen95]. deployment [Ano97h]. depository [MN95a]. depth [Ano95r]. describe [LT86]. describing [Hei95, LSI98, Rol91, VGGB95]. Description [ASWS98, LR95, Smi92a, Smi92b, Wei96b, BB93, Intxx, Kom91, MK88, New90, Wei96a]. descriptions [Tay95a, Tay95b]. descriptive [CJ90]. Design [ACM96e, Bly96, BR96b, Coy97, DLSR98, Dav98, DiN95, Div98, FNN97, Gra96a, HSCZ99, HTL95, JD96, KCCJ95, Lut98, MD94, MH96, NS95c, OSYM97, PWHK98, Que98, RW96b, Rem97, RKL^{+}98, SSK97, Tea97a, Tho97, WDM97, Wan99, WWEI97, Ano96i, BIL95, CMB93, Com97a, Com97b, DGG^{+}92, Gre98b, HK96b, H^{+}96, KH97, KGK94, Kum96, Lam97, LLY97, YF96, NS95b, O’C97, PMTJ98, PTJ98, Rii97, Rol98, SRB96, Sko98, Spe99, WWN96, WW98, WKV^{+}96]. designed [Sha98]. Designer [Ano96g, Ano96-32, Kau96, Wil92]. designers [Mar99]. Designing [LLF98, LB97, Moh99, SS89, TD95, Wei97, BMY95, Hef95, LMO^{+}96, LB99a]. Desk [Yuz89]. Desktop [Fla96, Tay90, YKT90, Gro95]. desperate [Bal98b]. Destabilizing [Mat97]. Detail [PCL^{+}95, Sch97b]. details [Pri96a]. detecting [Mat93]. Detection [HLCG97]. determining [GPE98]. Deterministic [BKW92a, BKW92b]. Deutschen [Sch90a]. Deutscher [Neu94]. deux [BK99]. Develop [SS98]. Developer [Ada97b, MLD^{+}99, O’R97, VRM98, Cow99, Dos99b, HK98, HL97, Rul98, Rul99]. developer.com [Hol97, Pur98].

Developers [McG98a, Ano96-41, Mul96b]. Developing [Fee98, Hoc93a, Ito98, Ivl95, Jam98, Lap99, LG95, MA96, Mar95b, ML97, She95, VT99, Fra99, MTU99, Reh95]. Development [Ano93s, Ano97c, Bee97a, Boa95, Fer93b, IEE96h, KLM94, KIM^{+}97b, Kro99, LHKC96, MSMY97, NANK96, Nak94, OST94a, Rem97, TSY^{+}95, Tew95, You96, ABB99, Ano90d, Ano97-33, Ano99g, Boo96, BB90, Buc97, Gol96, Kom93b, Lin94, Neg92, Rud96, TF95, YT96, Kro99]. Developments [Ano96u, BS94, BB91, OKT92, vSSW93]. device [BE99, Mor99a]. device-independent [BE99]. DEXA [Ano91c, QSBC98, RT95, TR92]. d’experts [LO99a]. Dexter [KCCJ95, MR93a, ZR93a]. Dexter-Groupware [MR93a, ZR93a]. DF [And97]. DFN [Sch90a, Sch90b]. DHTML [LB99b, Mud98, Ram98, Ram99, Tea98a]. Diagnostics [Col95d]. Dialects [Ano95i]. dico [Dre98]. DICOM [GFH96, Pri96c]. dictionaries [AT88, FWDV93, IV95]. Dictionary [Ano88g, Bra89, Faw93, Zha95]. Diego [Ano97-31, Ano98a, Bea95b, CC99, IEE95b, LB98a, NM95b, Ten99, USE89, USE96, USE98, W^{+}98b, Whi96f]. Diff [Ano97-53]. Difference [Ano96p, BD96]. Different [MH96]. Difficulties [HW88]. Diffie [Lut98]. Diffusion [BB95b]. Digest [Ano91e, Ano93g, Ano93h, Ano94g, Ano94h, Ano95p, Ano96u, IEE92, IEE96e, IEE96a, IEE98d]. Digital [ACM95a, Ano93v, Ano96n, BMK95, Ber96c, CK95, ERPNS97b, FA94, FB95a, Heb96, IEE96m, IEE97a, IEE99a, Lut98, Nor96, Ska99, Whi96f, AR97, AYE^{+}96, BEE96a, CW97, Del95b, FGGT96, Hen91, HAB98, JI95, Jac98, KMA^{+}96, MNW98, NS98, PAV96, RS93, Sug96, Tab96a, WB96, ABHY96, Bry98, CM90]. digitally [Wal94]. Digitization [Cam94]. Dimension [Whi96e, Ano95b, Ano96-27, Gau96b]. dimensional [AP96]. Dimensions [Ame97, Ano92b, Lik95]. d’Internet [BKP97]. DIP [Whi91]. Directed [DeV96]. directions [Kum96]. directives [LT86]. Direji [Gau96b]. DIS [Intxx]. Disabilitie [Mur92b]. Disabilities [Mur92a, RW98, Huc97]. Disability [Ano93g]. disabled [BEE^{+}94a, BEE^{+}94b, BEE96a]. Disambiguation [Aho97]. disaster [Fly93]. Discovering [BR95, Cro95b, Pfa97a, Pfa98a]. Discovery [Boy95, HPS95, SWC^{+}95, SBD^{+}95, BDH^{+}95, HKM^{+}96, Wei95]. discussed [Tit98]. Disjunctions [GL94]. Disparate [BSF98]. Dispenser [PD95]. Display [GLTH98, LG94a, ORCK95, SNLT96, SD98, Ber96a, CM95, NTR89]. displays [BFM96]. displets [VCB97]. Dissection [RJN95b, RJN95a]. Disseminating [GK95c, NK99, Huc97]. dissemination [Bou90, Rot97]. Dissertation [ES99b]. Distance [Ano96u, BBJR96, DL95c]. Distillation [Boy95]. Distributable [Ano97k]. Distributed [DS96, FMBM95, FFKO99, GS87, HLS97, HK94, IEE96b, IEE96i, KMS99, LZ95, Ple95, Roe95, SHW95, WH98, BG96, BHZ96, Fie94, Fuc96b, Hey94, Hue98, KDC95, LDWN96, LDW^{+}96, Moh99, RR96c]. Distributing [Hen99]. Distribution [AS93b, BS95c, Cel92, Cou97, SS95a, Str95, Ano99b, BS95b]. divergent [YWJ96]. Diverse [Van95c, SMC^{+}96]. diversity [DE95]. DLM [Ano97y]. DLM-Forum [Ano97y]. Do [GHJ95, Ano95-35, Ano97k, Ano97-54, Ber96a, Che97, HS97a, JMS98, PW96, Tit98, Wal94, dN97]. DocBook [WM99]. docs [Ano95g]. DOCUMATION [Ano94l]. Document [ABO98a, ABO98b, Ano91d, Ano92b, Ano93e, Ano93p, Ano94e, Ano97-45, AW93, Bou95a, BK93a, BKW92d, BK93b, Can89, Car92, CPW95, CGW96, Cze98, DHR97, Fur90, Gra83, Gue98b, vH87, HH99b, IEE93b, IEE95f, IEE97f, Int86a, Int88f, Int88g, Int88c, Int88d, Lev96a, Lev93b, LQ97, LZ98, MM90b, Nag95, O’K95, OSYM97, OL97, Pop97, Rag94c, RC95, Rob94a, RB95b, RKL^{+}98, Sch88, Smi92a, Smi92b, Sol88, SMH99, SMDO98, TBS97, Uza92, Wei94, van88, Ano93r, Ano94k, Ano95r, Ano97-54, App89, AMFL^{+}95, Bai93, BMN94, BB90, BB91, Bre96, BBKF95, BK93d, BK93e, CHP88, CM90, Con93, sta89, De 88, DMM90, FX93, Fuc96b, FMSDW93, Gol81, Gor91, HW90, Huc97, HLL93, IN96, Kam90a, Kam90b, Kan93a, Koc94, Kom91, LS96, LG94a, LHW96, Lo95, MSW97]. document [MB90, MB91, MM90a, MR93b, Mol91, Mol95b, MNW98, MB97b, NTI91, NKN91a, NKN91b, NR96c, NW97, Paq94, PB88, Rag94b, Rag94d, Rag94e, Rob94b, Sch92, SIF^{+}96, TKB90, War87, Wei93, Wol91, Won90, WR94, Yam93, Yam94, YHHYK96, Ada92, Intxx]. documentary [DPB91]. Documentation [ACM92, ACM95b, ACM97a, Ano93h, Ano97c, BB94, D’H95b, DP98a, Dés86, Fab95, LSC96, RM97a, Ano97-33, Bot93, Bro94a, Bro94b, Cha95a, CP93, CP94, Dou92a, Dou92b, Ev95, Gio95, Sul96, WS95a, d’H95a]. Documenting [Pri97]. Documents [AVFC98, Ano88c, Ano94g, Ano96-32, Ano97p, Ano97q, Ano97-53, Ano97-56, CLK96, Cam98, CAD^{+}95, CACS94, CMV98, CSM97, Fre95, Hah91, Haj93, HSCZ99, Heb96, HW96b, Hoe97, HS97b, HJM95, HY97, Ito98, Lap99, LR95, LH96, Lim95a, LN98, LN99, Lun95b, ME99, MB97a, MP96, MK96a, Myk95, NS95a, O’K95, PH97a, PKT95, RARH99, Ram97, RG97a, SD96a, Ste95c, SMDO98, TBS97, Vin98, Wil93b, AP96, Ano93d, Ano94b, Ano94u, Ano94v, Ano95r, Ano96m, Aza94, BVN95a, BVN95b, BM91, BEE^{+}94a, BEE^{+}94b, BR94, BR96a, BDKK96, BBB88, BC95, Bro89, CDL97, CCD^{+}99, CR94, CB95b, CM95, CMPd91, CL90, Cro92, CJ90, Dob91, ERPNS97b, Ens97, Erf93, ECvR93, FW93, Gol97, Gut98, Hei95, Hen91, HH96, HB96, Hun94, KZSC96, Lau93, LT86, Lev96b, Lim95b, Lim96, Mac90a]. documents [Mac90b, MBHL91, Meg98, Mol91, MS95, PY98, Pet90, PFH95, Roc93a, Rol91, RR94, Rut96b, SDAMZ95, STA98a, San94, SF90, Shi97, Sta98b, Vig89, WSM96, W^{+}98a, WV91, WP98, WAB^{+}96, YIU96, Daw98a]. DOD [Ano94t]. Does [Bur95b, RARH99, Ros97, Ano96i, Ber96a, Bur95a, Cha96a, Sch92]. dokonalych [Kos98]. Dokumentartag [Neu94]. Dokumenten [Rat95, Wie94]. dollar [Ens97]. domain [Lan91]. domaines [Int91d]. dominated [Gon94]. done [Ano97k]. Don’t [Ano96-29, Ano97-54]. d’or [Dre97]. DotsPlus [GB98]. Doubletree [ACM94a]. Douglass [Pri96b]. DRA [Hen96b]. Draft [New90, Ame91]. Dragons [HT95]. drama [Ano97e]. Drawings [Mol95a]. dream [Shi93]. Drive [Law98]. Driven [Ano97-41]. driving [Ano97-49, Dot95]. Drop [DL95b, DL95a]. Drop-in [DL95b, DL95a]. DSP [IEE97a]. DSS [Fer93b]. DSSSL [Ada92, Ham97a]. DTD [AMN97, BP93, YF96]. DTDs [Cov96, Lap99, MA96, St.99a, Wal90, WS95a]. DTP [Ise90, Mur90]. Dual [IEE96c]. Dual-Use [IEE96c]. Dublin [SC98, Ass87a]. Dueling [Ven96a]. Dummies [Ano99c, Kay96, RR96a, RR98b, Tit95b, Tit95c, TJ96b, Von97, Avi98, Hym97, HH99a, RR96b, RR97a, RR98c, TJ96a, TJ97a, TJ97b, TJ98, TMC98, TPM99]. during [BB94]. Dusseldorf [Ano98l, Rol98, RSV99]. d’utilisation [Att97, Int88i, Int91d]. Duty [KH95b]. Dvůr [Nad90]. DWTP [Bro98a]. Dyna [Nic95]. Dynamic [Ano99g, AAM^{+}96, Bel98, Bou95a, CDCD98, Del98, Del99, Dob97a, Dob97b, DHR97, Eyz95, Goo98, GLTH98, Hym97, HH99a, Isa98a, MK98b, MK98c, MS98a, MK98a, Mac98, MCC98, MBB95, MK96a, ORCK95, PS98a, Pet98a, PKT95, PR95b, PMSC98, Pow98b, RR98a, Rou98, Rul98, Rul99, Sch98, SP99a, St.97a, VT98, Vod99, Wil99c, Ano98e, CD97, CF99, Fis97, FHU99, GR97, HK98, Hol97, HU97a, HU97b, Isa98b, Lib98, LC98, Pet98b, SL97, Yee99, HS99, Isa97]. dynamically [MBK^{+}95]. Dynamicke [ID98]. dynamickym [ID98]. dynamickymi [ID98]. Dynamics [CIF98, CW95]. dynamique [SSB98, SM98b]. DynaText [Smi93].

E-Commerce [Lut98, McG98b, GTM99]. E-mail [DK97, PH97a]. E-Tif [Mel95]. E/F [Ivl95]. Each [Fox95]. EAI [Mar97a]. early [Bir96b]. Earn [Cea96b]. earth [MSIM95, JRK95, Kee95]. Earthquake [HM97]. ease [Ano94k, Ano96-32]. Easier [Ano95-44, Ano97u, BVN95b, Ric94]. Easing [BSF98]. East [Ano97-34, Lav97]. East-West [Ano97-34]. Eastern [SSB97]. Easy [Neu97a, Pfa97b, RJKS96, SDSS95, TAB96b, Pfa98b, Pfa98c, Pfa99, Rus99, Spi96a]. ECDL [NS98]. ECMAST [HS98]. Economic [Kli95]. Economists [PSTW97, Win97a]. ed [Joh95b, CM96, Mau93, Mau95]. ED-MEDIA [Mau93, Mau95, CM96]. EDBT [ABG96]. EDGAR [GK95c]. edge [Ano97-40]. EDI [Ben96, Orl93]. Edinburgh [ACM94b]. edit [Pao95]. Editing [Neg95, BVN95a, BVN95b, Bar98a, BBB88, SF90, TSY^{+}95, Wol91]. Edition [Ano98f, CJD^{+}96, HK96a, KS95a, LO96, LOA^{+}96, Lad96, LO^{+}99b, Mor95c, SBJ96, BJ96, GR97, GR99b, HK96b, Hol99b, HU97a, HU97b, Int86b, Mac98, SWH95, Smi96a, Wil97a, Vig89]. editions [Glu96]. Editor [Bee96b, Fly96, LCLL98, PK95b, Ric98, Ano94u, Ano95-29, Cad87, CM90, CMPd91, CB96, HLL93, YF96, Liu98, MR93c, MBM93, TKB90]. editorial [Ano97l]. Editors [SN96, TDT95, Ano95-39, Ano97-55, Bon92, Kam90a, Smi87f, Pri96b]. Education [Ano96-38, BL98, DG97, DL95c, EMH^{+}95, FB95a, FBDM95, Ibr95, Isk96, Joy98, Kle96, LRG96, MWEK97, MM97, PE95, RPWW96, Lit96, OW95, Owe96, VW96, VD98, JRK95]. Educational [Ano98a, CM96, Mau93, Mau95, Ano97v, FMFVN98, Hef95, Kom93b]. Educator [JOA98]. Educator/Researcher [JOA98]. effect [Ish95]. Effective [IEE94a, Rot97, Gar95]. effectiveness [Huc97, Shi97]. efficace [KG97]. Efficient [CIF98, Nag99, NN97, WNB97, PD96]. Eighth [NP96, Ano99i, MWEK97]. einem [Bat96c]. Einführung [HS99, Szi95]. Einsatz [Rie94, Sch90a]. EJB [Cha99]. EKAW [FS99]. Electro [Ano88e]. Electro/88 [Ano88e]. Electromagnetic [IEE96g, PBB^{+}96]. Electronic [Ada95, And88, Ano88d, Ano90d, Ano91e, Ano93e, Ano95f, Bac96, Blu95b, BB94, BBKF95, BKW92c, CR96, ES99b, Fur90, Goo94, HKH95, Heb96, HW96b, HAB98, JJ95b, KL91, Kru90a, LG97, MKNS87, Org91, Neu96, Orl93, PGRC97, Pit97, RLA96, Res97, Rob96, Rob94b, Rob94a, Run95, SM95, Sea95, See96, SS97, Str95, Wus96, Zha95, ZKM95, van95a, van88, Ass87a, Ano92e, Ano93d, Ano95l, Ano97t, Ano97y, Aza94, Bai93, BS94, BEE^{+}94a, BEE95, BB99a, Bla90, Bou90, Col96a, CS96, CJ90, Duk94, ECvR93, FM99, Gay94, Gio95, Gol97, Hen95, Hoc93a, JRK95, Key94, Mel95, Mol91, Mor95a, Oke93, PSG92, Pul95, RMF93, San94, Sea96a, Sea96b, Smi93, Smi97a, TSY^{+}95, VGGB95, Wae97, Wal92, War87, Sea94, Sea96a]. Electronics [Bed95, IEE93a, IEE96a, IEE98d]. électronique [San94]. Electrotechnical [IEE98e]. Electrotechnology [PS95]. Elektronisches [Neu96]. Element [MGH93]. Elements [Zif96]. ELF [AMFL^{+}95]. eliminate [KL91].

Eliminating [LFS97]. Ellis [BC90]. ELSA [Ada95]. email [McL98]. Embedded [Cov96, Kro99]. Embedding [Liu97b, ME99]. embodying [Yok93]. embrace [Ano93d]. emerge [RFM95]. emergence [Mas97]. Emerging [ACM95b, Mar91, MM95a, Wei95]. Emphasis [TRMW99, WCM99, MM97]. empirical [vSSW93]. Empowering [WD95]. Enables [JJ95b, JJ95c, Ano95-43]. Enabling [IEE96j, BEE96a, New96]. Encapsulation [PH97a]. encoded [BM91]. Encoding [BBSM96, CR95a, CR95b, Hoc93b, IV95, MR99b, BBG^{+}95a, Bir96b, Dra96, GB95c, Joh95a, GP90]. encyclopaedias [Ano97l]. encyclopedia [Hic89]. encyclopedic [Man97]. End [Fer93b, Fis95, Law98, Sjo95, Mee96, Ore88]. endangered [CBO93]. endorse [Dei96]. ends [Eft90]. Energy [Ano89a, For96c, PSTW97, Win97b, For99, Vez96, Win97a, Sti94]. Engine [CRV98, PR98c, PR98b, PF95, KRRK93, LHKC96, BD96]. Engineering [Ano93i, Ano97p, Bal96a, Coy97, CH95, DG97, GRA92, HTL95, IEE96o, IEE99b, Ibr95, Isk96, JD96, K^{+}98, LH96, Mel99, Mol95a, PS95, Rii97, SHW95, Sch93, SGP96, Su96, WNC95, Bro94a, Bro94b, Kho94, Kum96, Min96, Mol95b, Rol98, VC96, VW96, ZJEP95, She95]. Engineers [FZ99, Ano96i, Buc97]. England [BKS98]. English [Ano88g, Bra89, Edd97a, Edd97b, Edd98a, Edd98b, Faw93, Szi95, AT88, Edd99, Sch90a]. enhance [JRS96]. Enhanced [CLS96, HMHK97, PFH95]. Enhancements [HBNL95, Ano96y, Gre93]. Enhancing [CGLV94, Vin98, Ano99l]. enough [Coh99, CK97]. enriched [Dra96]. Ensembles [Int91d]. Enter [Sco96a]. Enterprise [Ano95e, DGG^{+}92, Don97, Jac98, RW96a]. Enterprises [IEE96j, W^{+}98a]. enthusiastic [Ano94t]. Enticing [Ric96]. entities [Bir96b]. entity [Int91c]. Entrepreneurs [PSTW97, Win97a]. Entrez [EKS95]. entry [CL90, MNA91a, MNA91b, ZJEP95]. entry-level [ZJEP95]. Entwurf [RH97]. envelope [Ano97j]. Environment [AVN98, Ano97p, CHH^{+}95, FP95, HLC96, vH87, KN95b, KFHU98, Kro99, Nat99b, Nad90, TTG98, TTS97, ZY95, CJ94, Gau96b, GL93, HMRL95, IEE95b, IEE96f, Kas95a, KN95a, KSM^{+}98, LG94a, MG94b, NANK96, Orl93, PJ95, PS96, QRV95, Sug96, Tex99, VB95b, WKV^{+}96, Wol95, Won94, AW95]. Environmental [Cro95b]. Environments [BW97, Bro98a, CPR98, HK94, PH97b, Res95, SOUH97, SN96, BMZ95, GSV96, Wal94, WH98]. environments-do [Wal94]. Envisioning [Ano97-55, Hil96]. EOSDIS [MSIM95]. EP [Ano93e, HAB98]. EP90 [Fur90]. equivalents [Ano95r]. era [Bou90, Jac98, Wes99]. Ergonomic [SSH87]. Ergonomics [Ano96s, Ano98r]. Eric [Ano98f, Joh95b, Van91b, Pop92, Pop94b]. Erstellung [Wie94]. erweitern [Bat96c]. ESA’93 [Len93]. ESIS [AGM94]. ESPRIT [Ano88f]. Essen [HW94, HW96b]. Essential [Farxx, LB99b, RT98]. essentials [Gar98b, Sch98, Zim97]. established [De 88]. estimation [SM93a]. Etat [Ano90d]. Etat-Major [Ano90d]. Ethernet [Cla98]. Ethical [Eic95b, Eic95a]. études [BK99]. EULOGIA [Dra96]. EURASLIC [CBK98]. Eurographics [Ano95j, GSV96, HK94, JRV96, VD90, Ano98m]. Europe [Ano94m, Ano95e, Ano96v, Ano97-37, CBK98, PMS97, SSB97, Ska98, Ska99, TS96, Wae97, Smi91]. European [ACM94b, Ano88a, Ano91b, Ano92a, Ano93f, Ano97h, Ano98m, CBK98, Dés86, FS99, HS98, Len93, Nad90, NS98, PK94b, SC98, To98, VD90, Gri95, Lew91]. EuropIA [Coy97]. EuropIA’97 [Coy97]. EUSIPCO [To98]. EUSIPCO-98 [To98]. EUV [BSB^{+}95]. Evaluating [Buf96, JI95].

Evaluation [JD96, Bus88]. Event [Tan98, WH98]. event-based [WH98]. every [Sea94]. everyday [Van98]. Everyone [Hah91, Pop92]. Everything [Hin98, Mee96, M^{+}96a, Coh99, HH98]. Evolution [Con94, Ens93, PS88, RTW96, Rod92, W^{+}98b]. Evolving [Sko96, Ano95-39, Yok93]. EWCBR [SC98]. EWCBR-98 [SC98]. examination [Buf96]. Example [Far96a, McG98b, Sta96, FS98]. Examples [VH95]. Excavating [GMG^{+}95b]. Excellent [Got98]. Exceptions [KW97a, KW97b]. Exchange [MKNS87, Sin97, YYOO97, Ano94k, Ano99l, LJ93]. Exchanging [Lev96b]. execution [vB89]. Executions [LFS97]. Executive [Gre98a]. Exhibit [CB95c]. Exhibition [Boh95, DLSR98, VD90]. existence [WD90, WD91]. Existing [Thr95, Dan98b]. expandable [MBK^{+}95]. Expansion [AS93b]. Expectation [FM99]. expensive [HS93]. Experience [AVFC98, Boy95, CMP97, HN96a, Men95b, Buf96, Gay94, Gon94]. Experiences [Dug97, GFR98, KMK95, Pin95, DT95, GP90]. Experiential [TTG98]. Experiment [DR95d, MKBB95, Whi95b, WCP^{+}95b, COMG96, RV94, Whi95a, WCP^{+}95a]. Experimental [SMDO98, HE94]. Experiments [GC95, PCL^{+}95]. Expert [Ano91c, BR95, DPB91, DWV95, IR94, PB88, QSBC98, Ste95b, TR92, TW98, Har94, RT95]. Explained [Bry97, Dob97a, Dob97b, Lut98, Vac98]. explains [Ude95b]. Explanation [LH96]. Explanatory [New90]. Exploitation [RDC99]. Exploiting [PCL^{+}95, PW94, RC96, Cim96]. Exploration [AW95, WKV^{+}96]. Exploratory [ZMP96]. Explore [BB95a, Ano95b]. Explorer [Dil97, Weg96, LD95a, MCC98, Lej96, Tuc99]. Explores [Sch96b]. Exploring [Dil96, DM94, Pel97a, GPV97, GSG98]. Expo [Ano93j, Ano95l, Ano96-29]. Exposition [Ano88d, Ano89c, Ano95m, Ano97a, Min96, Ano93j]. EXPRESS [Sha97]. expression [GSG98]. Expressions [BK93a, BKW92d, CC97, BK93c, BK93d, BK93e]. extend [Bat96d]. Extended [BK93a, Toy98, BK93d, BK93e, Cav94]. extender [LLL^{+}91]. Extending [ABG96, Cod98, FMFVN98, FWDV93, FLF95b, FLF95a, Hus96, Mac90a, Pao96, Rag94a, Sla95b, Sla95a, VCB97, Fuc96a, HMRL95]. Extensibility [FMW95, PKT95]. Extensible [CRV98, DC98, KD98, NK99, Ube94, Har98, Kam90b, Kuc98]. Extension [Ara98, BL98, MGVS97, Sei95, Sei96, SMDO98, Rod92, VY88]. Extensions [GN95, Gra97b, Gra98, Kul94, ZJ96]. Extent [PC97]. External [Mar97a]. Extracting [LN98, LN99, Tay95a]. Extraction [Ano97-41, Fre98, HD95, Sta98b]. Extracts [CDB91a]. Extranet [Ano97l]. Extraordinary [Ano97r]. Extreme [Fic98a]. Eye [Lin96].

F [Ivl95]. F/A [Ivl95]. F/A-18 [Ivl95]. Fabrication [WDM97, Bou97b]. Face [Ano93k]. facil [McF97b]. Facilitate [ZY95]. facilités [Int88i, Int91d]. Facilities [CACS94, Int88c, Int88d, Can89, Can90, Int87, Int88f, Int88g, Int88h, Int88a, Int88i, Int88b, Int91a, Int91b, Int91c, Int91d, Int92a, Int92b]. Facility [LH96, Ano95-27, NTI91]. Factors [Ano96s, Ano97p, Ano98r]. Factory [Kro99]. fails [KS95b]. Fair [BB94, Mar95c]. Fall [Cim96, Mas97]. familiar [Ano95-39]. FAQ [DeR97, Whi95b, Whi95a]. FAQs [MMV95, Var95]. Fast [Ano96k, BBKL95, BW97, RJKS96, RT98]. Faster [Gra96a]. Fe [Sol88]. Fear [LL97]. featured [Ano93q]. Feb [GSV96]. February [ACM99b, Ano96a, Ano98o, BFM96, CC99, DGBH93, IEE96e, IEE96o, Joy98, MWEK97, Neu96, SV98, Ska98, Ska99, SB97, SR98, SB99, Su96, USE89]. Federal [Pro95]. Federated [HLY^{+}94]. Federating [SMC^{+}96]. feet [BCG99]. FERRET [HD95]. Festschrift [HW94, HW96b]. feuilles [Gla99]. few [BVN95a, BVN95b]. FIE’96 [Isk96]. Field [GN95, IEE96g]. Fielded [ZKM95]. Fields [MVKJ99]. Fifteenth [Ano98q]. Fifth [ACM93b, Ano93e, Ano96-34, DP98b, IEE95d, IEE96k, New90]. File [Bor95, NM95c, RB95a, SBS95b]. Files [Lun95b, Mar97b, Mol95a, Sun98a, MW92]. Filling [Ano97-46, JGKR96, Oke95]. filtering [PAV96]. Filters [Smi96b, Smi97b, Smi98]. Final [Ano95k, Rut96a]. Financial [GK95c]. Find [AS93b, Dun95e, Tit98]. FIND2 [KTK^{+}95, OST94a]. Finding [Pin95, Ano94t]. Finds [Pri96a]. Finish [Pal98]. Finite [KW97c, BK93c]. Finland [Ano96-40, HT97]. First [ACM94a, Ano96-35, Ano96-31, Ano96-36, BKS96, Cov99, DE95, Gra96d, HL95b, HL96, IEE96e, IEE97g, Int86b, Kho94, Len93, Luo97, MB97b, Smi96a, TDT95, W^{+}97, Ano97j, Heu98, LR94, Mee96, Wel95b]. Fish [DP95a]. Fish-Search [DP95a]. Five [Ano96-46, Str95]. fix [KS95b]. Fizz [Tub95]. FL [Ano95a, Ano96x, Ano98o, Cha96b, Chu98, Ste96d]. FLAIRS [Ste96e]. FLAIRS-96 [Ste96e]. fledged [AP96]. Flexible [Fer93a, Sch95d, de 98]. Flight [Ano97r]. Floor [Ano96-29]. Florida [COW98, Ste96e]. Flow [Mac93]. Fluid [CW95]. fly [Kru98]. flying [Pes96c]. FMS [FFKO99]. Focal [Ano96-30]. Following [BW97]. fonctionnelle [NB91]. fondo [Isa98b]. font [Kom93a]. fonts [Kom91]. FootPrints [Ano97w]. Forces [ACM96d, Mat97]. Ford [Ano96-29]. Forefront [IEE95e]. Form [GHH97, NM95c, Gut98, STMM98]. Form-based [NM95c]. Formal [BK93b, MGH97, Tau96, BH95b, BFH98, JRS96, LZSD95, Mat93]. Formalism [TBS97]. Formalization [PHL97]. Format [Can89, CRBB98, DABM97, Int86a, Int88f, Int88g, Int88c, Int88d, Rag94c, Ano95-31, CM90, Dun95e, LSI98, MW92, Mel95, Rag94b, Rag94d, Rag94e, Rei93, Ret96, Rot97, WSM96, GC95]. Formation [Str95]. Formats [Bor95, DRBS98, Ros97, Sea95, Ber96a, De 88, Wus96]. Formatted [Sun98b, Ano96-32]. Formatting [Sch95d, de 98, BBB88, LT86, Ski94]. formerly [Cim96, Mas97]. Forms [And96d, DWV95, Gut98, Kas95c, Pro95, Ano96-44, Dav95b, Dun95f, HH96, JGKR96, JMS98, Kri99, Oli97b, Yee99, TB96b]. Formsheets [Kri99]. formulas [Ass86a, Ass87b, vCL89]. Forschungsnetz [Sch90a]. Forty [IEE96e]. Forty-First [IEE96e]. Fortythird [MDK98]. Forum [ABHY96, Ano97y, BB95b, IEE96m, IEE99a, Win97b, Ano96v, Bow95, RS93, SGBO^{+}93]. forums [HB96]. Foundation [Bea95a, Sha95b, SS97, DeR97, Ten95, Wil98]. Foundations [Deb97, RM96a, RS97, Tit95a]. Fountain [Fis96b]. Four [HT95]. Fourteenth [Ano97z, IEE95e]. Fourth [ACM93c, FW96, IEE96f, IEE97f, LM95b, MM97, M^{+}97, WWW95, TS96, Oke95, SB99, Ano88g]. Fractal [HV92]. Frame [Tes96a]. FrameMaker [Sys97, JK97, JK98, RR94]. Frames [Dav98, Mic97, Pri97, Tea97a, Tes96a, JMS98, Rob98]. Framework [OSYM97, Ran98, CW97, GTM99]. Framing [Bat96a]. Franca [GG96]. francais [BKP97]. francaise [Vig89]. France [Ano96-34, Ano97h, ABG96, Col98, Dés86, HAB98, Heu98, IEE97e, MNW98, RM95b, Van95b, van88]. Francisco [Bee97b, BUP99, CEG96, SSK97]. Frankfurt [BB94]. Free [BBKL95, Pau95, STMM98, EtR94]. Free-form [STMM98]. Freedom [Ano93o].

Freeform [WDM97, Bou97b]. Frees [Ano92f]. freeWAIS [PFH95]. freeWAIS-sf [PFH95]. French [CJ95, AVFC98, San94]. frequently [Bow95]. Friedrich [Neu94]. Friedrich-Schiller-Universität [Neu94]. Friends [Bly96]. Frog [PWHK98, RJN95b, RJN95a]. Front [Fer93b, Sjo95, Wei96b, Eft90]. Front-End [Fer93b]. Frontiers [Isk96, CL97]. Frontpage [Ven97, Ano97-51]. FTP [Hig96]. Full [JMB95, JJ95b, JJ95c, LRG96, Neg95, Gol89, INIK96, Ish94b, Ish94a, Koc94, Neg92, NTR89, O’C92b, O’C92c, OKT92]. Full-Service [LRG96]. Full-Text [JMB95, JJ95b, JJ95c, INIK96, Ish94b, Ish94a, O’C92b, O’C92c]. fulltext [Sch95b]. fully [AP96, CL90]. fully-fledged [AP96]. Fun [Mor95b, Mor95c, MR96, MS98c]. function [COVS95, CC99]. Functional [RSV99, SMDO98, Bus88, INIK96]. Functionalities [PC97]. Functionality [ML97, JGKR96, PFH95]. functions [WSoSS98, WS99]. fundamental [IN96]. Fundamentals [Nat98]. Fusion [Oly96]. Future [Ano90e, Boh95, Bou97a, BPBV96, CV98, MR99a, Ano96s, BB90, BB91, CBK98, DT95, Gol95, Hil96, ISM95, Kum96, Mur90, Pri93]. futures [And88, Ano97-50]. FUZZ [IEE96k]. FUZZ-IEEE [IEE96k]. Fuzzy [IEE96k, MP96].

G [AKM95]. GA [Ano93j, Ano95l, LG94b]. Gaithersburg [ABY94]. galaxies [FGGT96]. gallery [Wil98, Hum95]. game [ACM95c]. gaming [Ano95-43]. gamuts [Mor99a]. Gap [Ano91a]. Gardens [FMC95]. Gateway [Boi95, FJ96, PW94]. Gateways [DR95d, Eft90]. gathering [GW94, KS98c]. Gazette [Ano98p]. Gbit [Cla98]. GCA [Gra83]. Gdansk [Cze98]. Geh [HW94]. GEIDE [Ano96v]. GenCode [Gra83]. Genealogical [Chr96b]. Genequiz [COVS95]. General [Fre98, OSE97, TTS97, GI96, Ano90d]. generalisé [Int88e, Int86b]. generalised [Stu87]. Generalized [Ass86a, Ass86b, Ass87a, Ass87b, Ass87c, Ass89, Ame85, Ano88c, Ano93d, Ano94t, Ano97f, Ano97-40, Bry88, Can89, CG86, CDB91b, CDB91a, CH95, Fee88a, Fee88b, Fra96, Gra83, Hei94a, Hei94b, Int86c, Int86d, Int86b, Int88e, Kay89, Kis95, Kru90a, MMRS98, Nat88, NBM93, Pit95, Smi86b, Smi87c, Smi87e, SM93b, Sun98a, Woo95a, AC93, Ame91, Buc93, CG87, EG90, EG91, Gol81, Hay93, Smi87a, Smi87d, Smi87f, Smi87b, Smi89]. Generated [O’K95, Wal94]. Generating [AMN97, ELW95, JMB95, GSE^{+}94, Liu98]. Generation [Aho95, Ano97-41, Bou95a, CR96, Fin95b, Koh95, Mat99, SBGO99, And96c, Ano95-32, Ano97-55, AAM^{+}96, Blu97, BRR94, Bus88, Cha93, Cha95a, DJ96, DBRB96, LMO^{+}96, LHW96, MBRB96, MM95a, Oke93]. Generator [Sch97b, GL91, PS88]. Generic [WR99, Hol87]. Genesis [Wit86]. Genetics [Ano98n]. Geneva [Cai94b]. genome [COVS95]. Gentle [Kas95b]. Geographic [AS98]. geographical [GS95c]. GEOLIS [DO96]. Geologic [FR97, Ano97d]. Geology [UL98]. Geometry [Pel97a, Tau97]. Georgia [ACM95b, Ano96a, GE95, Joy98, Min96]. Geoscience [IEE94b, Ste95a]. Geospatial [AS96, Ano99h]. GeoWeb [Ple95]. German [Ano99j, Die96, Rie94, Sch90a, Sch90b, Smi97a]. Germany [AHB98, ABB99, Ano97-48, Ano98l, BFH98, Cov99, Dud97, FS99, HS98, IEE96h, IEE97f, Len93, Rol98, RSV99, SB99, Vez96, BBKL95]. gestion [San94]. gets [Ano95-27, Lun96, Man95, Mul97c, Ric92]. Getting [Ano97c, Ano97m, BDK95, Gro93, Sun97]. gezamenlijke [VB95a]. GhostSpace [NN97]. Gippsland [SV98]. GIS [Ano95m, KDC95]. GIS/LIS [Ano95m].

Give [Law98]. Gives [Cla98]. Giving [DK97]. Glasgow [Kum96, Rut96a]. Global [DBE^{+}95, LSC96, MKBB95, RBBN95, CJ94, Hey94, IEE95b, Kah95]. Globalization [Jac98]. Glossary [BS94]. glyph [MCP91]. glyphs [Bir96b]. GmbH [Neu96]. Go [RARH99, CT96, LM95a]. goals [ISM95]. Going [ACM92]. Gold [DE96, Mor95c]. Goldfarb [HV92]. Good [LM98, Ano95-27, BH95a]. goodies [Bur99b]. Government [Boy95, Can95, Mac95]. Grab [And96b]. Graduate [Fon93]. Graffiti [Fis95]. Grammar [DeV96, FW93]. grammar-based [FW93]. Grammar-Directed [DeV96]. Grammars [AMN97, BK93a, BKW92d, BK93d, BK93e, Con93, EtR94]. Grammatical [TBS97, Bot92, Wal90]. grand [SM98b]. Graph [Cou95, CLS96, LSI98, Ano95l]. Graphic [CG86, EG90, EG91, BE99]. Graphical [ADE93, Bun96, CH95, Döm95b, BDKK96, CCD^{+}99, Döm95a]. Graphically [Fab95]. Graphically-Based [Fab95]. Graphics [Ano89c, Ano93b, Ano98m, BKS98, CG87, Kro99, LCM96, Lun95b, Owe95, RJN95b, Ska98, Ska99, TS96, VD90, dN97, ACM96a, Ano93q, LLY97, MZZ96, RJN95a, Sun97, Tub95, VB95c, Wei97, GK93]. Grass [Qui95]. Grasshopper [Ano96e]. Graz [HK94, Mau95]. Great [Ano97-46, IEE96n, MM99a, HH96, HB96, Lig95]. great-looking [HH96, HB96]. Greece [CBK98, IEE97a, To98]. Grenoble [IEE97e]. Grif [Pao96]. Grips [Ano94d]. Groningen [Ano90c, Ner97]. Ground [Ano95f, BCG99, Ano96-32]. Group [Ano96-36, Ano97-31, Ano98t, Bee97b, Bow97, BCG99, CL97, Cla98, Gra96d, Str95, Tea97b, Tea98b, TS96, Ano96-44, Dou92a]. Groups [Law98]. Groupware [Lun96, TTN93, MR93a, ZR93a]. Growing [ACM95b, BBKL95]. Growth [HT95].

Grumman [Ivl95]. Grundlagen [Min98]. Guarantee [Kle94]. GUI [Gar99, KK98]. Guide [Als95a, Als95b, Ano95-36, Ano98f, Bar92, Bry88, Cas96b, Cas97, DDM96, Eva96, Fis96a, Fly95, GR96, Gra97b, Gra98, Hah91, Lew97, MD97, MK96b, MK97, MK98j, Neo97, OnW95, Que97, Que98, RLA96, SS88, Smi96b, TW95, WM99, Zac97, Ano97o, Min97, BC90, Car98c, Chi95, Com97a, Com97b, Cou95, Cow99, DD94, Dic99, Dos99b, EG90, EG91, Eva95, Eva98a, Fis97, Gol94, GPE98, Gra96c, GQ99, HK96b, HK98, HE95, Hol97, Keo97, KH95a, KH97, Man97, Mar99, McF96, McF97a, McF98, Mul96b, PMTJ98, PTJ98, Pur98, Roa95, Ros96, Rul98, Rul99, SBM96, Smi97b, Smi98, TS97, TSP98]. Guidelines [DC98, MB90, OCL88, Smi87d, Smi87f, Smi87e, van90a]. guides [Cox96]. GUIs [CF99]. Guru [Cea96a]. Gutenberg [And97].

hacker [Bal98b]. Hahn [Pop92]. Halstead [BC90]. Hamamatsu [IDK98]. Handbook [AP 96, Ada97a, Barxx, Bor95, Fly95, GR90, GP98, Har96, KS95a, She88, HW96a, HW97, Bec93, HV92]. Handicapped [ZBW94]. handle [Mac90a]. Handling [Fly98, Ges91, Lic91, Rob94a, BMN94, Gil92, NR96c, Ric95c, Rob94b]. Hands [Bea95b, Lap98, LL98, RAS99, Car98c]. Hands-on [Lap98, LL98, Car98c]. Handwritten [MB97a]. hang [Sun97]. Hans [HW94]. Hans-Peter [HW94]. Happen [Har95b]. haptic [Liu97b]. Hard [Fer93a, Hen95]. Hardcopy [Hoe97, BE99]. harmonizing [BCK^{+}94, WS95b]. Harvest [BDH^{+}95, SBD^{+}95]. Haskell [WR99]. having [Web95b]. Hawaii [A^{+}98, ERPNS97a, ERPNS97b, HK95b, IEE96f, Küc97]. HAWC [ASWS98]. HCI [SSK97]. HDF [MSIM95]. Head [BCG99, CLS96]. headache [CNH99]. header [Gio95]. headers [Dun95g]. Health [KD98, MMRS98, SGJ^{+}95, Tea97b, Tea98b, Chu98, ECvR93, GJTW97, Har94, Hut96, Mas97, Wae97]. Healthcare [Als97, Ano97-38, Ano98o, BSF98, DABM97, IR94, W^{+}98b, Wes99]. Heights [Wil94a]. held [AHB98, Ano90d, Cze98, Dud97, HAB98, IEE96d, IEE98c, IDK98, Kho94, Kum96, MTH99, Ner97, PKM99, PK94b, VC96]. Hellas [IEE97a]. Help [Bug98, Bur95b, Hou98, Ito98, LL97, NF97, RM97a, Wil96a, Bur95a, Ham97c, WWN96, WF98b, WF98b]. Helpdesk [Ano97w]. helped [Ubo96]. Helpers [Car95a, Car95b]. Helping [KAKZ96]. Heraklion [NS98]. Here [AS94a, Ano95b, Ano96-30]. Heritage [Bus97, Del95a]. Herkunft [Min98]. Herwijnen [Pop92, Pop94b, Joh95b, Van91b]. Heterogeneous [RP98]. Heures [Fra97b, WM98]. HeVeA [Mar98]. Hewlett [Ano93f]. HI [ERPNS97a, ERPNS97b]. Hidden [Tes96a]. Hidden-Frame [Tes96a]. HieNet [Cha93]. hierarchic [Mac91]. Hierarchical [BBG^{+}95a, Cov96]. Hierarchies [Pri97, SMH99]. High [Ano89a, Ara98, BG95b, CGL^{+}95, DP95b, SD98, Ten99, AYE^{+}96, BG95a, Chi95, Win96, Win97b]. High-Energy [Ano89a]. High-Level [Ara98]. High-Speed [BG95b, BG95a]. higher [VD98]. Highlights [BDG^{+}95]. Highway [Ano95f, Ano95q, BB95a, EBR^{+}96, IEE96c]. Hill [BUP99]. Hilton [Gar95]. HIM [KR95]. hip [TS97, TSP98]. Hiroshima [IEE96l]. Historical [KW97c]. History [Blu95b, HRS95, MR99a, GB95c, ISM95]. History-of-Use [HRS95]. HL7 [DRBS98]. HMTL [O’C97]. HMTL/SGML/JAVA/VRML [O’C97]. hoc [Dra94]. hohe [Ano96z]. Hold [Ano96-29]. Home [Dra95b, Lam97, LRG96, Sch95c, BIL95, Dra95a, DM95, HN96b, HH96, HB96, Sko98]. hones [Ano95l]. Hong [DS96, IS98].

Honnef [Len93]. honor [HW94]. Honour [HW96b]. hood [Ude95b]. Hooking [AP96]. hoopla [Ano96-42]. Hopes [Law98]. hormones [Wu98]. Horticultural [Ano98d, BL98]. Horwood [BC90]. Horwood/Halstead [BC90]. Hospital [Dug97, Dud97]. Hot [IEE95d, Mor96]. HotDog [Ano96o]. Hotel [ACM94a, ACM96c, ACM98, Ano95k, Ano95-37, Bos97, BUP99, Cim96, DG97, Gar95, IEE96k, LB98a, LMT98, Mas97, Oke93, Oke95, RM95b]. HoTMetaL [Ano95n, Ano95-44, Ano97-51, How96]. HotOS [IEE95d]. HotOS-V [IEE95d]. Hour [Mee96]. hours [Mee96, OH97b, Oli97a, OH97c, Oli98b, Oli99]. house [Ivl95, Ros88]. houses [Sch95b]. How-To [KFZ96, Sun97, KFZ97, ZFK97]. how-to-do-it [JMS98]. HP [Rag98]. HPC [FFKO99, Ten99]. HPCUA [Ano93f]. HPECU [Ano93f]. HPECU/HPCUA [Ano93f]. HPP [DHR97]. hsin [mWfC97]. Hsinchu [CL99]. hsiu [mWfC97]. hsueh [Neu98, mWfC97]. HTML [Ano96w, Ano97w, Ano98f, Die96, ER97, HS99, KS96, Schxxb, WF98b, Alv97a, Alv97b, And96d, Ano95d, Ano95g, Ano95l, Ano95n, Ano95o, Ano95c, Ano95r, Ano95-31, Ano95-39, Ano95-35, Ano95-40, Ano95-41, Ano95-44, Ano95-46, Ano96l, Ano96p, Ano96q, Ano96r, Ano96-28, Ano96-32, Ano96-45, Ano97m, Ano97n, Ano97o, Ano97-29, Ano97-28, Ano97-30, Ano97-47, Ano97-51, Min97, Ano98e, Ano98p, Ano99g, Anoxx, Aok99, AB95, Arm96, Aro94, Aro96, Arp96, AM96, AM97, Att97, Aug99, Aus95, BMDB98, BVN95a, BVN95b, Bar98a, Bar97, Bar98b, BT99, Bar99, BH98a, Bar94, Barxx, Bau95, Bel98, BE97, Ben97, BKP97, Ber95, Ber96c, Ber96b, Bjo96a, Bjo96c, Blu97, Bly96, Boa95, BR96a, Bou98, Bou95b, BL98, Bri98a, Bri98b, BJ96, BA97, Bro98b, BH98b]. HTML [Bry97, Bry98, Bry99, Buc97, Bug98, Bun96, Bur99b, Cal97, Cal98, CLK96, CD97, CDCD98, Car98b, Car98a, Car99, Car95a, Car95b, Car98c, Cas94, Cas96c, Cas96a, Cas96b, Cas97, Cas98, CH99a, Cea96a, Cea96b, Cea98, CC98, CF99, Che97, CV98, CC96, CRV98, CMV98, Coe96, Coh96, C^{+}99, Col96b, Com97a, Com97b, Con94, Con96, CSM97, Dar97, Dar98, D^{+}99, Dav98, DP98a, Dav95b, Dav95a, DeR97, DG95, DG96, Del98, Del99, DBRB96, Del95a, DiN95, Die96, Dob97a, Dob97b, DK97, DHR97, Dre97, Dre98, Dug97, Dun95d, Dun95e, Dun95f, Edd97a, Edd97b, Edd98a, Eri97, ES99b, Eva95, Eva96, Eva97a, Eva97b, Eva98a, Eva98b, Eyz95, Fal97, Farxx, Far96a, FS98, Fee98, Fer97, Fic98b, Fin95b, Fly95, Fly96, For95b, For96c, For99, FD95, FD96, Fra97a]. HTML [FHU99, Fra97b, Fre98, Gre98a, Gar98a, Gar98b, Gar97b, GRSM97, Gla99, GDRR94, GFR98, Gol96, Gol98, Goo98, GR99a, Got98, Gra97a, Gra95c, GR96, Gra96c, Gra97b, Gra97c, Gra98, Gro97, GG96, GR97, Gut98, Hac97, HN96a, Ham97b, Ham97c, Har95a, HLC96, Har95b, HK96b, HK96a, HK98, Heb96, Hei97, HC97, HiKgK98, HLB98, HS97a, HH96, HB96, Hes97, HH98, Hie97, Hig96, Hoe97, Hol97, Hol99b, Hol99a, HUW97, HU97a, HU97b, HUW98a, HUW98b, HB97a, HB97b, HS97b, Hou98, How99, HJM95, HY97, Huc97, HL97, Hym97, HH99a, Intxx, IZ97, IZ98, Isa97, Isa98a, ID98, Isa98b, Ito98, JK97, JK98, JGKR96, Jac96, Jal98, Jam97, Jam98, JW96, JOA98, Jon95, JMS98, Kar96, Kas95c, Kas95b, KS95a, Ker97, KFZ96, KFZ97, Kes95, KH95a, Kid97, KH97, KPIE99]. HTML [KG97, KDH96, Kor98, Kos98, Kri98, Kro99, Kru98, Lib98, LaR95, LO96, LOA^{+}96, Lad96, LO^{+}97, LO98, LO99a, LO^{+}99b, Lam96, LL97, Lan96, Lap98, LM98, Law98, LM96, LD95a, LH96, Lej96, Lem95a, Lem95c, Lem95b, Lem96a, Lem96b, Lem96c, LD97a, LD97b, LD98, LT98, LT99, LS98, LB96, LS99, Lig95, Lim95a, Lim95b, Lim96, LN98, LN99, LHW96, Lin99, Lin94, LSC96, Liu97b, Liu97a, Liu98, Loc97, dL97, Lot95, LG95, Lyn96, LC98, MK98a, MK98d, MK98e, MS98b, MK98f, MK98g, MK98h, MK98i, MP97, MP98a, Mac98, Mag97, MGVS97, MR99a, Man97, Mar98, MP95a, MD97, McA94, MM99b, McC97b, MOTA95, McF96, McF97a, McF97b, McF98, ML97, McL98, McM96, MBRB96, MCC98, Mee96, Men95a, MHP98, Mey99, Mic97, Mil95, MP96, Mor95b, Mor95c, MR96, Mor96, MS98c]. HTML [MS95, Mul96a, Mul97b, Mul97a, Mul98, Mul96b, MP98b, MP98c, MP99, MK96b, MK97, MK98j, Nat98, Nat99a, NANK96, NM95c, NF96, NF97, NR96b, NR96a, Neo96a, Neo96b, Neo97, NR97, Neo98, Neo99, Neu97a, Neu98, O’K95, OSE97, Okl97, OuP97, OH97a, Oli97b, OH97b, Oli97a, OH97c, Oli98a, Oli98b, Oli99, OA98, Opp97, PY98, PH97a, PS98a, PC97, Pat97, Pat98, PS98b, Pau95, Per96, Pet98a, Pet98b, Pfa97a, Pfa98a, PG98, Phi98a, Phi98b, PB98, PI98, PMSC98, PMTJ98, PTJ98, PGRC97, Pop97, Pow95, Pow97, Pow98a, PW98, Pow99, PW99, Pow98b, Pri97, Pri96c, PK95b, Que97, Que98, Rad97, Rag94b, Rag94d, Rag94c, Rag94e, RLA96, Rag96a, Rag96b, Rag98, RLAK98, Ram98, Ram99, RJKS96, Ran98, RR98a, RC95, RC96, RR96a, RR96b, RR97a, RR97b]. HTML [RR98b, RR98c, RR99a, RR99b, RR99c, RW96b, Red99, RV99, RT98, Reh95, Ret96, Ric94, Rob98, RAS99, Rou98, Rub96, RM96b, RM97b, Rul98, Rul99, Rus98, Rus99, Rut96b, Sad97, San99, SFO97, SWH95, SBJ96, Sch95a, Sch96a, Sch97a, SR97, SRSC98, Sch98, Schxxa, SP99a, Sch95c, SPR^{+}96a, Sea95, Sei95, Sei96, Sha97, She96, SWC99, SCW99, SVC99, SS98, Shi97, SL97, Sig96, SK98, Sin97, Sko98, Soi97, SMG95a, SMG95b, SMG96, Spi96a, St.97a, Sta98b, SD96b, Sta96, SR96, Sta97, Ste95b, SM98a, SSB98, Ste98, SMiPM98, SM98b, SMDO98, SDSS95, Sun98a, Sun98b, Swa96, TAB96b, TBS97, Tay95c, Tay95d, Tay95e, Tay96, Tay97a, Tay97b, Tay98, Tea97a, Ten96, TB96b, Thr95, Tit95a, Tit95b, Tit95c, TJ96a, TJ96b, TGHE96, TJ97a, TS97, TJ97b]. HTML [TJ98, TSP98, TJD98, TPM99, TRMW99, T^{+}xx, Tus97, Uni99, Ude95b, USO97, Ven97, Via97b, Via97a, Via97c, Via97e, Via97d, Via98a, Via98b, VB96, Vin98, VCB97, Vod99, Wil98, Wor97, WBSY98, Wag97, Wal95, Wea97, WWEI97, WW98, WU96, WM98, WF98b, Whi96a, Whi96d, Wil96a, WCM99, Wil99b, Wil99c, Wri96, mWfC97, Xu97, Yee99, Zac97, ZFK97, ZMP96, ZKM95, ZJ96]. HTML-script [Gol96]. HTML. [MK98b, MK98c, MS98a]. HTML3.2 [AL97]. HTMLgen [Ham98]. HTMLParser [Tuc99]. HTTP [BLFF96, Coh99, FGM^{+}97, HKH95, HBNL95, Lot95, PM95, RR96c, Soo95]. HTTP/1.0 [BLFF96]. HTTP/1.1 [FGM^{+}97]. HTTPD [Ivl95]. hui [Neu98]. Human [Ano96s, Ano97p, Ano98r, SSH87, SS89, SSK97, Sha95b, ZY95, CNH99, Ric96, RSV99]. Human-Computer [SSH87, SS89, SSK97]. Human-system [Ano98r]. Humanities [Ano96f, Ano99e, Uns96, Hoc93a, Uns96]. Hungary [Ano99e, EAZ98, SSB97]. Huntsville [Gau96a]. huo [HUW98a]. Hyatt [IEE96k, OW95]. Hybrid [Nat99b, RW96b, KSM^{+}98]. Hydro [Ano97f, Roc93b]. Hydrogen [PSTW97, Vez96, Win97b, Bos97, Win96, Win97a]. hydrogène [Bos97]. Hydrology [Ano96a, Ano99a]. HyOctane [KRRK93]. Hyper [Arm96, MKBB95, CG95, AKM95, Intxx, Lin94]. hyper-book [CG95]. Hyper-G [AKM95]. Hyper-Media [MKBB95]. Hyperactive [RL95b]. Hyperbolic [MB95b]. Hyperbraille [KK94]. HyperCode [BHK^{+}95b]. hyperglossary [CLPR96]. hypergrids [MCG^{+}95]. HyperKids [Ric95a].

Hyperlink [LS98]. Hyperlinks [KW97c]. Hypermedia [ACM94b, AT95, Ano93x, Ano95p, BC96, BKS96, CM96, Cox96, DDM96, FMBM95, HK94, KC92, Mau91, Mau93, Mau95, MBH95a, Rat95, Ric95a, RB95b, Van93a, AMC95, AP96, Bea95b, Bry93, DD94, Dra96, EH96, Gol91, GL93, Hey94, Kan93b, KCCJ95, KZSC96, KSC96, Kom94, Kom93b, Lau94, LHKC96, MBK^{+}95, MCG^{+}95, Mar91, MBH95b, NKN91a, NKN91b, NN92, Pri93, Reh95, RV94, SRB96, ZR93b, Int97]. Hypermedia-Based [Ano95p]. Hypermedia-Dokumenten [Rat95]. hypermedia/time [NKN91a, NKN91b, Int97]. hypermedia/time-based [NKN91a, NKN91b, Int97]. Hypertext [ACM93b, ACM96b, Ano93x, Div98, Döm95b, EKS95, Fee98, Har95b, Jal98, KR95, Lam96, Mar94, Mau91, MBB90, ORCK95, Pri97, RBZ95, SS98, Wil93c, Wil94c, YNAD97, ZR93d, Ano92i, CR94, Col95a, Döm95a, Dra94, ELW95, Fie94, GL91, Kah96, Kas95a, KDC95, KK94, KCCJ95, MP95a, Mar95b, Nic92, O’C92b, O’C92c, PS96, Pow94, Put94, Rob94b, Van93a, ZR93c, Ano95-29, Ano95-42, Bar94, CPW95, Ric95b, BLC95, BLFF96, FGM^{+}97]. Hypertext/Hypermedia [Mau91]. Hypertextove [HS97a]. hypertexts [ACM96e]. Hypocenters [HM97]. HyTime [NKN91a, NKN91b, Win96, OISEM97, PSTW97, Win97a, Win97b, ABO98a, ABO98b, Ano97-54, BRR94, Buf96, BQRT95, CBDH94, DD94, DDM96, Ebe98, Erf93, FB95b, Gol91, Int97, KW97c, KRRK93, LHKC96, Mar91, Mar92, NN92, New95a, New95b, PST97, PBFT97, Pop93a, Pop93b, Rat95, RB95b, RBR95, Rut96b, VB95a, Ano96z, Mac93].

i.e [USE98]. i3D [BG95a, BG95b]. IAAI [Ano98q]. IAAI-98 [Ano98q]. IAPR [IS98].

IBM [BS98, DGG^{+}92, LDWN96]. ICA [RBZ95]. ICADD [Ano96-38]. ICC&IC [Kim97a]. ICCBR [ABB99]. ICCBR-99 [ABB99]. ICCE [IEE98d]. ICCHP [EAZ98, KAKZ96, ZBW94]. ICCIMA [SV98]. ICE [IEE96j]. ICED [Rii97]. ICHIM [Bea95b]. ICIAP [BDV95]. ICOIN [IEE98f]. ICOIN-12 [IEE98f]. ICPS [Ano94f]. ICSEE [DG97]. ICT [Lav97]. Identifiers [Tau96, Int88a, Int88b, Int91a, Int91b]. Identifying [Shr95]. Identity [Gol95]. IDI [Ano94c]. idiot [McF96, McF97a, McF98]. IDT [Ano96t]. IE4 [HU97a]. IE5 [FHU99, Hom99, Moh99, Spe99]. IEC [VRM97, Int88h, Int91a, Int91b, Int91c, Int92a, Int92b, Int97, Intxx]. IEE [Ano91e, Ano93g, Ano93h, Ano94g, Ano94h, Ano95p, Ano95q, Ano96u]. IEEE [Ano90g, BKS98, CEG96, LMT98, PS95, W^{+}97]. IEEE/RSJ [IEE97e, IEE98b]. If [Wil96b]. IFIP [Ano93i, EAZ98, Jac98, TS96, VD98]. IGARSS [IEE94b, Ste95a]. IGES [Mar97b]. II [Ano95x, COVS95, Com97b, Dob97b, Gra95b, LMO^{+}96, LLL^{+}91, New95b, O’C92c, SS89]. III [BFM96, HS97a, Mos96]. IIPS [Ano96a, Ano99a]. Illinois [Ano98r, Gra96d]. illustrated [Gra97a, Red99, RV99, Vod99]. Ilmenau [Ano97-48]. im [Sch90a, Vul97]. IMAC [IEE96f]. Image [Ano91d, Ano96-29, BDV95, IEE96f, KS98a, KS98b, Lic91, MGVS97, MM96, Qui95, Sei95, Sei96, Uns96, dN97, Fil94, MZZ96, MB97b, Nag95, Uza92, Yam93, Yam94]. imagemaps [JMS98]. Images [AS94b, Dav95a, MGVS97, MVKJ99, CC99, DM96, FGGT96, RS96]. Imaging [Ano88d, Ano95k, AA97, Car97, Gol95, JD96, Kro99, Pri96c, Run95, Wei94, Ano94f, Ano96v, BE99, CC99, HAB98, IEE96f, Sko96]. IMC [Ano96v]. Immedia [Bat96a, RKW96]. Immersed [STMM98]. Immersive [EH97, STM97]. immun [SFO97]. Impact [CGL^{+}95, LRG96, OBO97, Ano90d, BC95, CB95b, GJTW97, SF96]. impaired [KPD96]. Imperial [JRV96]. Implementation [Bot98, DAOD95, HKH95, HSCZ99, HJM95, HLL93, MP94, RW95b, Rid95, RKL^{+}98, Ste96a, TW95, ACM96e, Alv99, Ano97f, Bed95, Buf96, BR96b, CMB93, KCCJ95, Pet90, Spe99, Wan99, WV89, Wol95, Xu97]. Implementations [Car94]. Implemented [SM95, Kel98]. Implementierung [RH97]. Implementing [Ano93j, BHK^{+}95a, CGW96, Gla95, HL95b, HL96, LHL97, Luo97, OR96, Pet98c, RM97a, Tes96a]. Implications [CV98, Smi86a, Faw89, LG97, War87]. important [SG93]. imprescindible [Fra97a]. Improve [CSM97]. improvement [KTK^{+}95]. Improvements [BR95, Web95b]. improves [Gan96]. Improving [Ano97q, FB95a, Kah96, MP95b, PM95, Pol95, Kae88, Kae90]. In-depth [Ano95r]. In-house [Ivl95, Ros88]. INA [BS90]. Inc. [Ano92b]. Includes [Kru96, Lot95]. InContext [Ano94u]. Incorporating [Uhl95]. Increased [ML97]. Increasing [BEE^{+}94a, BEE^{+}94b]. incremental [FX93]. incrementally [GMS93]. independent [APB^{+}99, BE99, HBNV97, Rag94a]. Index [CN98, Lim95a, Van95c, Lim95b, Lim96]. Indexing [DSDT96a, MK96a, DSDT96b, Van95c]. India [MM97, MDK98, V^{+}96]. Indianapolis [DGBH93, Gup97]. Indispensable [PMK98]. Individualised [KK95]. Individualizing [OPW95]. Industrial [Ano93i, Don97, EDLS98, Gre98b, Won94]. Industrial-strength [Don97]. industries [YKT90]. Industry [Ano97-34, Dei96, SS95c]. INET [PK94a]. INET’94 [PK94b]. INET’94/JENC5 [PK94b]. Inexpensive [MVKJ99]. Inference [MK96a]. InfoAccess [Ano96w]. InfoHarness [STMS95]. informaci [HS97a]. informacyjne [Dan98a]. informatica [Mos96]. Informatics [Ano98c, Chu98, PMS97, COW98, Cim96, GJTW97]. informatics/ [GJTW97]. Information [Ame85, AS98, ABY94, Als95a, Als95b, Als97, Ano88g, Ano90a, Ano91b, Ano92f, Ano93g, Ano93p, Ano93s, Ano95f, Ano95k, Ano95x, Ano96a, Ano96-35, Ano96-33, Ano97k, Ano97-38, Ano98o, Ano99c, Ano99h, Ano99a, BB95a, BKS98, BLG^{+}95, BB95c, BFY93, Boy95, BHK^{+}95a, Bur93, Bus97, Can89, Can90, COW98, CFMG95, CEG96, CV98, CPW95, Coh95, Col95c, Cro95b, DP95a, Del95a, DWV95, EAG^{+}95, ERPNS97a, FD91, Fis96a, For95a, FK96, Fre98, Fun99, GB98, GE95, Gup97, HSCZ99, HK95a, HW96b, Hen99, HB98, HAD^{+}95, Hou98, HY97, IEE95c, IEE95e, IEE96e, IEE98f, Int86b, Int86a, Int87, Int88f, Int88g, Int88c, Int88d, Int88a, Int88i, Int88b, Int88e, Int91a, Int91b, Int91c, Int91d, Int92a, Int92b, Jel98, JDGP95, JCDA95, JJ95a, KS98c, KC92, KR95, Kum96, LK95, LMT98, LN99, Lit96]. Information [Mac95, Mag94, MFVG99, MKEB98, MP96, MM97, Mue95, Neu94, O’K95, Pol95, PS95, Pro95, Rai90, Rai92, RJ93, RJ94, RJ95, RMG98, RG95, RW95b, RPWW96, RL95a, Rut96a, SD96a, SBD^{+}95, STMS95, Sin97, SS95c, SNS95, W^{+}98a, WD95, Whi96f, WHC95, Wil93c, WCP^{+}95b, YYOO97, ZY95, AC93, Ame91, AMC95, ACM96e, AKM95, Ano90d, Ano96n, Ano96t, Ano96-32, Ano96-43, Ano99l, BHMS98, BEE^{+}94a, BEE^{+}94b, BEE95, BEE96a, BKP97, BMN94, Bot92, BDH^{+}95, CHR^{+}96a, CBK98, CPK95, Chu92, Chu98, Col98, CS96, De 94, Del95b, Dud97, Dun95e, EBR^{+}96, For96a, For96b, Hei95, HW94, Hel96, HE94, Huc97, Hut96, IEE96c, IS98, Ish94b, Ish94a, IN96, Kaz95, Kel94a, Kel94b, KTK^{+}95, Kir91, Kom93a, Kom94, Lau94, LJ93, Lig95, LG94b, Mac91, Mor95a, MDK98, NN92, NTR89, O’C97]. information [OST94a, PH96, PAV96, PD96, Pig96, Pow94, Put94, Rat98, RS93, Rot97, Sch92, SR93, Smi91, TSY^{+}95, VN95, VB95b, Wan99, WE92, WT93, Wha92, WCP^{+}95a, YWJ96, d’H95a, AW95, Ano92b, Int86c, Int86d, Int88h, Int97, Intxx]. Information-Security [Ano96-33]. informational [Wes99]. Informationsmodellierung [Lob99]. InfoSleuth [WT95]. infostructures [Fie94]. Infrastructure [Gar99, HPS95, IEE96j, BMZ95, Har94]. Infrastructures [IEE95e, Yok93]. ingredients [H^{+}96]. Initiative [BBSM96, Law98, MR99b, Hoc93b]. Initiatives [Ano96-38]. Inline [Mol95a]. Innopac [Hen96b]. innovatiions [IEE98b]. Innovation [Par95, Rid95, DE95, Jac98]. Innovative [Ano97z, Ano98q, Tew95, IEE97e]. input [Roe87, TSY^{+}95]. Inquiry [BR95]. ins [HL97]. Inserting [Rag96a]. InSight [GK93]. Inspection [NJDF96]. Installation [Lew97]. Instant [Aus95, HUW97, HU97a, HU97b, HUW98a, HUW98b, Ken96, Wri96]. Institute [ACM93c, ABY94, Ano95h, Ano96x, Ano96-40, Lam96, Uns96]. Institutional [DP95b, MOTA95]. Instructing [KDC95]. Instruction [BBJR96, Bor96]. Instructional [RC96]. instructor [Com97a, Com97b]. Instrument [ASWS98]. Integrate [FFKO99, LG95, WDM97]. Integrated [BGG^{+}94, Div98, DK97, ID99, BP93, DMM90, GB95c, GL93, MM90a, RS93, Smi91]. integrates [Ano95-29]. Integrating [AS98, CHH^{+}95, Col95d, CS96, D’H95b, FS95, Hic94, Lo95, Nic95, Owe95, Owe96, Sha97, SS95b, Vau94, Whe96, d’H95a, BMN94, Col95a, GR99a, SF90]. Integration [Ano95e, KD98, LQ97, Lot95, STMS95, Soa98, TB96a, Ano96-32, Ano97e, Ano97s, DPB91, W^{+}98a]. Intel [Law98]. Intellectual [KTK^{+}95]. Intelligence [ACM93c, ACM97b, Ano90b, Ano97z, Ano98q, GRA92, IEE97d, M^{+}97, SV98, Ste96e, GW94, SMG95a, SMG95b, SMG96, Ste96d, Yok93]. Intelligent [Ano96-31, Cod98, ECvR93, IEE97e, IEE98b, Lic91, WT95, Hay93, OM96, PJ95, RM96a, RS97]. IntelliTeaching [CGL^{+}95]. Intensional [WBSY98]. intensive [Fra99, IDK98]. intentions [Hen95]. Inter [Int94]. Interacting [Kas95c]. Interaction [SSH87, SS89, SSK97, Sha95b, Ano98r]. Interactions [ZY95]. Interactive [Ano96a, Ano97t, Ano99a, Ara98, BB95b, BS95c, BS95b, Cea96b, CW95, CB95c, Cro95a, DCG^{+}95, Fic98b, GJ96, GHH97, Hac97, HKH95, Kap97, KSGJ96, KCFP95, LD95b, MVKJ99, OA98, Pet95, Put94, RBBN95, RJN95b, RJN95a, RB95a, Ska99, TRMR97, Tes96b, Wan97, WSoSS98, WS99, Wea97, AT95, Ano98e, Cea98, Chi95, CJ90, Fis97, Gol97, Hic89, HMRL95, Lib98, Lin94, Liu97b, LG95, Qay96, RKW96, SGBO^{+}93, LG94a]. Interactively [BR96a]. interactives [Coh96]. interactivity [Bea95b, Gut98]. interaktivnich [ID98]. InterCam [Ano97w]. Interchange [Ano91e, BSF98, Can89, DABM97, DRBS98, Fis95, GC95, Int86a, Int88f, Int88g, Int88c, Int88d, New95a, New95b, ZR93d, Car92, CM90, sta89, De 88, Dou92b, Gol91, Kom93a, Kom94, LG94b, MW92, Mel95, Mol91, Orl93, Pri93, RBZ95, Rei93, Rob94b, Spr96b, Van89, WSM96, ZR93c]. interchangeability [NW92]. Intercontinental [Ano97-37]. Interdisciplinary [KAKZ96]. Interest [Bow97, BCG99, CL97]. interesting [Str96]. Interface [APP^{+}98, Ano95p, Ano97w, ADE93, Bun96, CJ98, DWV95, For96c, HLR97, Jam98, KSJ95, Mar97a, ML97, Nad90, NS95c, Tea97a, APB^{+}99, BR96b, CMPd91, EBR^{+}96, EH96, For99, Fuc96b, Gan96, GFH96, HE94, HB95, KMA^{+}96, LDWN96, Mac90a, Nad96, NTI91, NS95b, Paq94, TC96, Wan97, DO96, FJ96]. interface/application [NTI91]. Interfaces [Ano93i, Ano97k, RG95, SS89, Eft90, HT97, KSGJ96, Wil92]. Interleaf [Ano92c, Ano95r]. Intermediate [Ano97-28, Ano97-47, Rob98, Via97a]. Internal [LN99, NS95c, NS95b]. International [ACM92, ACM93a, ACM93c, ACM95b, ACM96d, ACM97b, ACM97a, ACM98, ABY94, AR97, ABB99, Ano88d, Ano89a, Ano90g, Ano91c, Ano93e, Ano93a, Ano93j, Ano93s, Ano95e, Ano95v, Ano95-33, Ano95-34, Ano96n, Ano96a, Ano96-34, Ano96-35, Ano96-33, Ano96-31, Ano96-36, Ano97x, Ano97-42, Ano98s, Ano99b, Ano99d, Ano99f, Ano99a, ABG96, AA97, BKS98, Bea95b, BP96, BFY93, BH95b, BFH98, BDV95, BBKF95, BKS96, BUP99, COW98, Cha95b, CL99, Col95c, Col98, Cov99, Coy97, DG97, DP98b, Dob98, FW96, Fur90, Gra96d, HT97, HW94, HW96b, HAB98, IEE92, IEE93b, IEE94b, IEE95a, IEE95c, IEE95f, IEE96d, IEE96e, IEE96a, IEE96f, IEE96b, IEE96k, IEE96l, IEE96i, IEE96h, IEE96o, IEE97a, IEE97b, IEE97f, IEE97e, IEE97d, IEE97g, IEE98c, IEE98d, IEE98b, IEE98f, IEE99a, IEE99b, Int92a, Int92b, IR94, IDK98, IS98, JM96]. International [K^{+}98, KAKZ96, Kom91, Kom93a, Kom94, Küc97, Kum96, Lav97, LMT98, LM95b, Mel99, Min96, MTH99, Mos96, MG94c, M^{+}97, NW97, WWW95, Pig96, PS95, QSBC98, Rai90, Rai92, RJ93, RJ94, RJ95, RMG98, Rii97, RPWW96, RSV99, RM95b, SSH87, SS89, SSK97, Sch93, SSB97, SV98, Ska98, Ska99, Ste95a, Su96, TS96, TR92, TW98, VRM97, W^{+}98a, ZBW94, van88, Ada92, Ano96f, Ano97d, Ano97-27, Ano97-31, Ano98d, Ano98t, Ano99i, CBO93, ERPNS97a, ERPNS97b, GJTW97, Heu98, Jac98, LLY97, LR94, MNW98, Pri93, RM96a, RS97, RT95, V^{+}96, Ass86a, Ass86b, Ass87a, Ass87b, Ass87c, Ass89]. Internationales [Ano97-48]. Internationalization [Tak96, YNAD97]. internazionale [Mos96]. Internet [Ano95-33, CC96, Cim96, For95a, Hes97, HH98, Hie97, Lil95, Loc97, NDM98, PK94b, USE97b, Vac96a, Yee99, And96c, Ano95t, Ano95c, Ano95-29, Ano95s, Ano95u, Ano96l, Ano96q, Ano96r, Ano96y, Ano96-30, Arn95, BD96, Bat95, Ben96, Bro97, Bro98a, Che97, Chi95, Com96, DM94, FG95, FD95, FD96, Fra95, Fry95a, GW94, GK95c, GPV97, Har95b, HIYO96, HN96b, HM97, Hen99, HT95, HH96, HB96, HPS95, JRK95, KH95a, Kor98, LO99a, Lam96, Lau94, LLF98, Lev96b, LG95, Mag97, Man95, MM99a, MHMS^{+}94, MKBB95, MC97b, NR96b, Pri95, Rak95, Res97, Roa95, Ros96, Rus98, SBM96, SPR^{+}96a, Sko96, Tuc99, Ubo96, Wal95, Wei95, Wei96b, WR95, Wil96b, Wit96a, YT96, van95a]. Internet-based [FG95, JRK95]. Internet-Software [Hie97]. Internet/VB [Ano96-30]. internetable [Mas97]. Internetwork [BMZ95]. interoperability [IEE96o, New92]. Interpretation [IEE94b]. Interpretive [Hui96, TA99]. interrogation [Sam95]. interventions [MWVE99]. Intranet [Ano97c, Dug97, Ano96-41, BJ96, Schxxa, Schxxb]. Intranets [SPR^{+}96a, Ano96-43, Lun96, O’C97]. Introducing [Ano97r, Ano97-56, Gil93, RW96a]. Introduction [Ano94j, Ano97-29, Coe96, DP98a, Ell94, Haa93, Hei94b, Kas95b, LaR95, Ost94b, Pel97b, Pri96a, Rie94, She96, SM93b, Tus97, Uni99, Via97b, Via97e, BH98a, Don97, Fic98b, GS95b, Got98, Hol87, Ker97, Lam97, Pri93, Reh95, Sko98, Szi95, Win96, Hei94a]. Introductory [MHP98, Reb95, Car98b, JRS96, SCW99, SVC99, Vod99]. Intrusive [Pol95]. Intuitive [Ano97s]. Inventaire [AVFC98]. Inverse [AS93b]. investigation [Col95a, WAB^{+}96]. Iowa [IEE96n]. IPCC [Ano96k, Ano93k, IEE97b, Wil94a]. IR [Fil94, JMB95]. Ireland [BH95b, IEE99b, SC98]. IRIS [GK93]. IROS [IEE97e]. IRS [Ano90f]. IS&N [M^{+}97]. ISAS [COW98]. ISATA [Rol98]. ISBN [BC90]. ISD [Fer93b]. ISD/LSAR [Fer93b]. ISI [NPPR95]. Island [IEE96f, PWHK98]. ISMIS [RM96a, RS97]. isn‘t [Lig95]. ISO [Ass86a, Ass86b, Ass87a, Ass87b, Ass87c, Ass89, Bar92, BC90, Int92a, Int92b, Rie94, VRM97, Cla98, Int86c, Int86d, Int88f, Int88g, Int88h, Int91a, Int91b, Int91c, Int92a, Int92b, Int97, Intxx, Rie94, SS88]. ISO/IEC [VRM97, Int88h, Int91a, Int91b, Int91c, Int92a, Int92b, Int97, Intxx]. isolation [BBG^{+}95b]. Ispolzovanie [Pat98]. Israel [Ano99d, BP96, IEE98e]. ISSAC [Küc97]. Issue [Ano96-46]. Issues [Ano93h, BHK^{+}95a, Dav98, Gol95, IEE96o, JD96, Mar95c, OR96, Pri97, Smi86b, Smi87c, TB96a, Wid94, Woo95a, AYE^{+}96, And88, BBKL94, BDK95, CDB91b, RTW96]. ITAB [LMT98]. Italy [Jac98, Mos96, M^{+}97]. IV [BE99, Her96]. IVDS [Qay96]. IVEE [AW95]. IWAN’99 [Cov99]. IX [To98, USE95].

J [Pri96b]. Jacksonville [Cha96b]. Jacques [RM95b]. Jane [Pop92]. January [A^{+}98, Ano96a, Ano98b, Ano98n, Ano99a, BE99, BFM96, BUP99, DG97, ERPNS97a, ERPNS97b, FJV96, HK95b, IEE98f, KJR98, LB98a, LZ98, USE89, USE96, W^{+}98b]. Japan [Ano99f, IEE96b, IEE96l, IEE98f, IDK98, MTH99, BLG^{+}95, Yuz89]. Japanese [HY96, Nak94]. Japanese-Language [Nak94]. Java [Amm97, Bat96c, Die96, Hie97, Ano96p, Ano97k, Div98, HJS98, KG97, O’C97, VB96, AVN98, AS96, And96d, Ano96-45, Ano97-51, Bat96d, Ber96a, BKP97, Boo96, Bot98, Bru98, BB99b, Bry98, Buc97, CH98, Cha99, CIF98, Dei96, Die96, EM99, EBR^{+}96, Eng97, Fla96, FFKO99, Gar97b, GFR98, Gol96, Gol98, Got98, Gre98b, Hug96, Hus96, HL97, Kar96, KS96, Kro99, LO96, LOA^{+}96, Lad96, LO^{+}97, LO98, LO99a, LO^{+}99b, LMM96, LB98b, Lot95, Mac96a, MTU99, Mat97, Mau98, MD98a, MD98b, Mee96, M^{+}96a, M^{+}96b, Par95, Pie97, Rem97, RCRBR97, Rus98, Schxxa, Schxxb, Sco96b, Sha95a, Sha97, SD96b, Sun97, TRMR97, Tub95, T^{+}xx, Ubo96, Wag97, Wat99, Whi96c, Xu97, You96, ZJ96]. Java-applet [Gol96]. Java. [ER97]. Java/EJB [Cha99]. Java/VRML [Eng97]. JavaScript [BT99, Buc97, Coh96, Die96, GJ96, Hus96, HL97, JOA98, KDH96, LO^{+}97, LO98, MHP98, MP99, NR96a, Neo96a, Neo96b, Neo97, NR97, Neo98, Neo99, Sco96a, Sha98, Tea97a, Tes96a, Tes96b, Wea97, CC98, Jac96, MP98c]. JavsScript [Amm97]. jazyka [HS97a]. JCALS [Ano92b]. Jena [Neu94]. JENC5 [PK94b]. Jersey [W^{+}97]. Joan [Bar92, BC90]. Jobs [ACM95b]. Johnson [Mei97]. join [Dav92]. Joint [Ano92a, Ano94f, Ano96f, PK94b, SF90, VD98]. jointly [HAB98]. Jordao [dN97]. Jose [BE99, BFM96, FJV96, KJR98, LZ98, MWEK97]. Jour [Lut98]. Journal [SM95, SBGK98, van90a, VRM98, Key94, SGBO^{+}93, Wus96, Ano96-46, MR95, WWW95]. Journals [CR96, Neg95, Pul95, Van89, Wal92, van95a]. July [ACM97b, AR97, ABB99, Ano96x, Ano97z, Ano97-42, Ano98d, Ano98q, AA97, BKS98, Bee97b, COW98, Cov99, DE96, Heu98, IEE97a, KAKZ96, Küc97, OW95, PS95, Ass96, USE97a, WNC95]. Jun [Ano94p]. June [ACM93c, ACM97b, Ano94b, Ano94n, Ano96f, Ano96v, Ano97-27, Ano97v, Ano97-50, Ano98a, AA97, BBJR96, BCG99, CM96, CL97, Cov99, Cze98, Dés86, DP98b, DS96, HT97, HK94, IEE95a, IEE96a, IEE96b, IEE96l, IEE96j, IEE96c, IEE98a, IEE98d, IEE99b, JM96, Kho94, LR94, Mau95, Mel99, PK94a, RM96a, Rol98, RSV99, USE87, VC96, Vez96, W^{+}97]. junior [Chi95]. Just [Bjo96c, GHJ95, Ram97, RG97a, SK98, Sim99, Kay89, Lig95].

Justifying [Ano97i].

KARP [Ano95v]. KARP-95 [Ano95v]. Keep [Ano96l, Ano96r, Ano97t]. keeping [Not96]. keeps [Ano92c, Goo94]. Kennedy [Aok99]. Kenosha [Ano97-43]. Key [Ste96d, Ano96s, BEE^{+}94b]. keying [OCL88]. Keyword [Law98]. kill [Sed95]. kind [Sea94]. Kingdom [RT95]. Kiosk [Rak95, BDK95]. Kit [FD95, FD96, GMS98, Web95a, Ano97n, Ano99g, FS96, HB97a, Mac98, WF98b]. Kitchen [FNN97]. Know [Che97, Mee96, Tay90]. Knowing [Lau93]. Knowledge [ABY94, Ano93a, Ano95v, BFY93, FS99, Fun99, HKM^{+}96, KHW97, LHW96, Mar95a, ME99, RDC99, SS89, Sch93, Ste95c, Ano98c, Ano99l, CBK98, GJTW97, Kel98, Tay95a, Yok93, DO96]. Knowledge-based [LHW96]. Kodak [Ano95l]. Koganei [IEE98f]. Kolloquium [Neu96, Ano97-48]. Kommunikation [GS87]. Konferencyjne [Dan98a]. Kong [DS96, IS98]. Korea [IEE96d]. kouzla [PB98]. Krakatoa [Ano96-43]. kreativ [Hie97]. Kurdi [Gau96b]. Kurdish [Gau96b]. kwa [HiKgK98].

Lab [Ano96-30]. laboratories [JRS96]. laboratory [PBB^{+}96, FMBM95, TCLV92]. Lacking [AMN97]. Lake [ACM97a, IEE97b, Isk96]. Lakes [IEE96n]. LAN [Lik95]. land [Mat98, Sum94]. Lands [Blu95c]. Landscape [Wei96b, Wei96a]. landscaping [DE96]. langage [Aug99, Fon93, Int86b, Int88e, Mic99]. Langley [NB95]. Language [Ass87b, Ass87c, Ass89, ACM99b, Ame85, Ano88c, Ano94g, Ano95-43, Ano97f, Ano97-40, Arm96, Bar94, BMK95, BHZ96, BS90, BBJR96, Bor96, Buc97, Can89, CDB91b, Dan98b, Deb97, Die98, DWV95, DC98, FH97, Fee98, Fee88a, Fee88b, FR97, Gra83, GJ95, Hei94a, Int86b, Int88e, KD98, KK95, Lam96, Lan96, LR95, LD95b, MGH97, MN95b, MMRS98, NM95b, Nak94, NK99, PR95a, Pit95, SNLT96, SAA^{+}94, SB97, SR98, SB99, SM93b, YNAD97, ZJ96, AC93, Ame91, APB^{+}99, And96b, AMFL^{+}95, BA90, Bot98, Buc93, CCD^{+}99, CG87, CJ90, DT99, DFF^{+}99, EM99, EG90, EG91, Gau96b, GJ96, Gre98b, GL99, Har98, Hay93, Jal98, Kim94, KS98c, Kri99, LSS96, LT86, Mac90a, Mac91, Mat93, New90, NKN91a, NKN91b, NN92, Ric94, Rod92, Smi87a, Smi87d, Smi87f, Smi87b]. language [Smi89, Stu87, VY88, YHHYK96, Ass86a, Ass86b, Ass87a, Ada92, And96b, Ano93d, Ano94t, Ano95-29, Ano95-42, Ano96-42, BLC95, Bru98, Bry88, CG86, CPW95, CDB91a, Flo96, Hei94b, Int86c, Int86d, Int97, Intxx, Kay89, Kis95, Kru90a, Kuc98, Lin94, Nat88, NBM93, Ric95b, Smi86b, Smi87c, Smi87e, TTS97, Vac96b, Woo95a]. Languages [ME99, Smi92a, Smi92b, Ano94b, Ano99f, BKW92a, BKW92b, C^{+}98, CBO93, Die96, Intxx, Kom91, RFM95]. langues [CBO93]. LANs [Ano95z]. Large [Bou97a, Fab95, KH95b, Mar95a, MKEB98, Nic95, O’K95, Pop97, TB96a, V^{+}96, Dun95g, FMSDW93, MCG^{+}95, Roc93a, TS99, Yok93]. Large-Scale [KH95b, MCG^{+}95, Yok93]. last [Bra96a]. Late [Mag97, RCRBR97, Pie97]. Latency [MP95b, PM95]. latest [Ish94b, Ish94a]. Latinae [DeV96]. Launches [Lik95]. Laura [LCM96]. Laval [CBO93]. Law [ACM93c, ACM97b, PHL97, Wil93c, Wil94c]. laws [Goo94]. Layout [BDKK96, FNN97, ID99, LR95, Mic97, BDK95, Con93]. layouts [Ano95l]. lazy [McM96]. LDAP [Moh99]. leaders [Dei96]. leading [Bar98a, Ens97]. Leap [Ano96-29]. Leaping [Dar95]. Learn [Cal97, Cal98, LM96, Ram98, Ram99, RJKS96]. Learned [Bee96b, CMP97, Coh95, Gla95, SMDO98, BBSM96, MSW97]. Learning [Ano96u, BBJR96, Cha96b, Cha95b, CGL^{+}95, Fre98, GJ95, Kea96, Lan96, NF96, OW95, PE95, Per95, Pes96a, Pul95, BG96, CJ94, GJ96, GSG98, JI95, Wol95]. Learns [Gib98]. leben [Bat96c]. L’Edition [Vig89, Att97]. Left [Sch95c]. Legacy [Ano97-56, FS95]. Legal [Ebe98, FK96, Lau93]. Legislature [Ano97-35]. Lemay [LCM96]. Less [Ano97-46]. Lessons [BBSM96, Bee96b, Bra89, CMP97, Coh95, Gla95, Jam98, MSW97, SMDO98, Van95c]. Let [Fuc96a, Gra96a, Kay89, Ano94u]. lets [Ano95b, Dun95e]. Letter [Bee96b, Fly96]. Letters [Ano92d, Ano93m, Ano95w, PKM99, Bed95]. Level [Ara98, Bar94, MK98b, MK98c, MS98a, MK98e, MS98b, MK98f, MK98g, MK98h, MK98i, Sch97b, BAC93, Ben97, ZJEP95]. Levels [PCL^{+}95, BBG^{+}95b]. leveraging [Whi95b, Whi95a]. Levy [Tow95]. Lexical [Con96, WSM93]. lexicographic [Gau96b]. LHC [Bal96a]. Librarian [KH95b, Ste95b, NDM98]. Librarians [Kis95, CP96, Wil94c]. Libraries [ACM95a, Ada95, BMK95, CBK98, IEE96m, IEE99a, Neu96, Nor96, PR98a, Rid95, WP95, ABHY96, AR97, AYE^{+}96, Ano96n, Ano98e, Cox96, Del95b, JI95, Lib98, MM97, NS98, Sug96, WB96]. Library [Ano95x, Ano97w, Ara98, CK95, Cze98, ES99b, FB95a, Fry95b, Gro97, Hen96b, JMB95, KMS99, KH95b, Lev93a, MLD^{+}99, O’R97, SSB97, TTS97, BS94, CB95a, Cox96, DM95, FA94, JSY97, KMA^{+}96, MDK98, NL95, Okl97, Pow94, Sea94, Sea96b, Tab96a, TSY^{+}95, Ten95, Aok99, Gay94]. license [Col96a]. Life [OBO97, Bat96d, Cha95a, Gra95d, Ish95, Ric94]. life-cycle [Cha95a]. Lifecycle [Kro99]. lifesaver [Aza94]. Light [Cov96]. Lighthearted [Fis95]. lightning [Not96]. Lightweight [DB95, Liu97a]. like [KSGJ96]. Limerick [BH95b, IEE99b]. limit [Ano98r]. limits [OW95]. Line [Hir95a, Hir95b, Jon95, MB97a, Whi96c, Ano95l, Ano97-35, Dou92a, RV94, Sea96a]. Lines [Wei94]. Lingua [GG96]. Linguae [DeV96]. lingual [Ano96-33]. linguistes [CBO93]. Linguistic [Ano96f, Ano99e, FK96, Ner97, Bur99a, MBM93]. Linguists [CBO93]. link [Cha93, Jad97]. linked [GJ96]. Linking [Ano96-29, Fer93b, NF96, NF97, DeR99b, MCG^{+}95]. links [CR94, ELW95, Lo95, Sra96]. Linux [Ano98f, Ano98p, Gar99, Kro99, Pau95]. Linuxdoc [Sch95d]. Linuxdoc-SGML [Sch95d]. Linz [KAKZ96]. LIS [Ano95m]. LISA [Ano95x, USE95]. LISA-II [Ano95x]. Lisboa [Ano98m]. LispWeb [RR96c]. list [Ano96y, Hon95, Coh95]. lists [Oli97b]. Literary [Ano96f, Ano99e, Kea96, BHK^{+}88]. literate [MG94b]. literature [SMC^{+}96, Wei95]. little [MMJ96, Tit98]. Liturgical [Rad98]. Live [HD95, Soo95, BVN95a, BVN95b, Lig95, Oze96]. Live3D [FS96]. livestock [Ano98n]. Living [Reg98]. Livre [Dre97, SM98b]. Load [Kro99]. Loading [BB93]. Local [BB95c, Kah95, VW96]. locating [Kea96]. Location [Ple95, SM93a]. LoD [AAM^{+}96]. Lodestar [Sch97b]. Logical [STA98a, BAC93, Buc93, Rol91]. logically [Yam93, Yam94]. London [Ano94b, Ano94e, Ano97-50, BKS98, JRV96, RJ95, RMG98, RT95, Wae97]. long [Ano93r, CM95]. long-document [Ano93r]. Look [Ano96-30, McL95, Ano96-28, Rah95, Whi96d]. Looking [APP^{+}98, Swa96, HH96, HB96, HH98]. loosely [ECvR93, SDK^{+}96]. Lösungen [HS99]. lot [Ric94]. Lots [JDGP95]. Loughborough [Ano95j]. Louis [ACM95c, Lit96]. Louisiana [Gar95, IEE96k, IEE96o, NP96, Su96]. Low [RMR97, Thi91, BAC93]. Low-Bandwidth [RMR97]. low-level [BAC93]. Low-tech [Thi91]. Loyola [EMH^{+}95]. LSAR [Fer93b]. Lufthansa [HL95a]. L’Unione [RV94].

Lurching [Kor98].

M [Bar92, BC90, Ano96-29]. M/cad

[Ano96-29]. m4 [Ano98p]. MA

[Ano95h, USE97a]. Mac

[Anoxx, HK96a, NR96a]. Machine [Ano92h, Fre98, Nad90, HT97, MYN96]. Macintosh [Cas96c, Neo97, Ano96q, Bar98a, HK98, HH96, KS95a, LM96, Run95, Smi95]. Macro [DHR97, Sum89, Sun98a]. Macro-Based [Sun98a]. Macro-Preprocessing [DHR97]. Macworld [Ano95y, Tay96, Tay97b]. Made [Ano95-44, Ano97-32, BVN95b, SDSS95, Lew95, MM99b]. Madison [Ano98q, IEE95a]. Madrid [VD98]. Magazine [Gra96b, Ahe93a, Ahe93b]. Magellan [SS95a]. Magic [IZ97, KMS99, MP94, Ude95a, Ude95b, Col95b, PI98, Rou98]. Magic-sets [MP94]. MagiCaster [Ano97w]. mail [DK97, Lev96b, PH97a]. mainframes [Sal96]. Mainstream [Gar97a, Luo97]. Maintaining [Fie94, GMS93]. Maintenance [BL98, Fer93b, Jon95, MG94c, Bot93]. Major [Ano90d, Dug97, WE92]. make [Ano96-32, Ano96-42, BVN95a, BVN95b, Okl97, Sha97]. Maker [Ubo96]. makes [Lig95, Ric94]. Making [Ano97u, DD94, DDM96, Har95b, Lut98, Rad97, RW96a, RW98, Wal95, RKW96]. Malo [HAB98, MNW98]. Man [Nad90, HT97, Ham97b]. Man-Machine [Nad90, HT97]. manage [DT99, McC97a]. Management [ABY94, ABO98a, ABO98b, Ano96x, Ano97-34, Ano97-45, Ano98o, BFY93, Bla94, DM97, Fun99, IEE96f, JM96, LHL97, Lik95, OISEM97, ORCK95, OL97, RDC99, Rob94a, TABA^{+}94, Wei94, WT95, Ano94v, Ano97-54, Ano99l, AAM^{+}96, Bai93, BCK^{+}94, DJ96, DMM90, FS99, HIYO96, Hay93, Jan96, Kaz95, Lev96a, Lo95, MSW97, MM90a, OSEM^{+}96, Rot97, SIF^{+}96, Vau94].

Manager [OnW95, SD96a, Dic99, LDWN96]. Managing [CAD^{+}95, Rid95, YWJ96, ACD^{+}95]. manifesto [SMG95a, SMG95b, SMG96]. Manipulating [GMG^{+}95a, PR95a, CMPd91, PB88]. Manipulation [Ano93e, ETT97, Fur90, STMM98, van88, BBKF95]. Manual [Ano93h, CB97, Fra97a, GP99, Pet98b, TF95, Aro94, Aro96, AL97, Gor91, JMS98, Lig94]. Manuals [Ano97r, DP98a, Nak94, AT95, KL91]. Manufacturing [Nat99b, RWBS97, Jac98, KIM^{+}97b, KSM^{+}98]. Manuscript [Org91, Ass87a]. Manuscripts [MKNS87, Rad98]. Map [Ano97-52]. Mapedit [DM96]. Mapping [BS95c, Ano95r, BS95b, IEE96c, MCP91, RSV99]. Maps [AVFC98, Ano97g, Sei96, Sei95, EH97]. Mar [Ner97]. March [ACM96b, Ano95j, Ano97-31, Ano98t, Ano99f, ABG96, BT97, Gra96d, HAB98, IEE96h, IEE96n, IDK98, JRV96, Joy98, K^{+}98, MTH99, MM97, MNW98, NP96, Su96]. March/April [HAB98]. Marina [ACM94a, ACM98]. Marine [CBK98]. Mark [HV92, Lin94]. Mark-up [Lin94]. marked [Ano94u, Mac90a, Hen95]. marked-up [Ano94u, Mac90a]. market [Car95b]. Marketing [Pol95, Ano96-42]. marketplace [Ano97-49, Mar91]. MarkItUp [FX93]. Markup [Ass86a, Ass86b, Ass87b, Ass87c, Ass89, Ame85, Ano88c, Ano93d, Ano94g, Ano94t, Ano95-29, Ano95-42, Ano95-43, Ano97f, Ano97-40, Arm96, Bar94, BLC95, Bry88, Can89, CG86, CPW95, CDB91b, CDB91a, Cov96, CH95, DC98, FH97, Fee98, Fee88a, Fee88b, Flo96, FK96, Gra83, HY96, Hei94a, Hei94b, Int86c, Int86d, Int86b, Int88e, Intxx, KD98, Kay89, Kis95, Kru90a, Kuc98, LaT89, Lam96, Lap99, MMRS98, MRR99, Org91, Nat88, NK99, NBM93, O’C92b, O’C92c, Pit95, Ric95b, Smi86b, Smi87c, Smi87e, SM93b, TTS97, Vol97, Woo95a, YNAD97, Ass87a, AC93, Ame91, Ano94b, BHK^{+}88, CG87, CJ90, DBRB96, EG90, EG91, Gol81, Har98, Hay93, Jal98, LS96, MA96, MBRB96, PY98, Smi87a, Smi87d, Smi87f, Smi87b, Smi89, Stu87, Wit86]. Marriage [Ano92h, Sco96b]. Marriott [ACM95c, ACM96c, LMT98]. Marshaling [ACM96d]. Maryland [ABY94]. Mass [CM96, Fry95a, IEE95e, Bat96d]. mass-media [Bat96d]. Massenmedium [Bat96c]. Master [Ada97a, Cea96a, Cea96b, Keo97, M^{+}96a, Sad97, Sch98, Wil99b]. Mastering [RR97b, RR99a, RR99b, RR99c, Wil96a, Wit96b, Rob98, Ros96]. Material [Mar95c, NJDF96, Ass86b, Ass87c, Ass89, BEE95, Kah96, New90]. materialized [GMR95]. Materials [Ano96x, Fer93b, LG95, Col96a]. Math [Blu95b]. Mathematical [Boi95, Woo95a, Ass86a, Ass87b, WSoSS98, WS99, vCL89]. Mathematics [Soi97, Fat97, Int91c, Roe87]. mathématique [Int91d]. MathML [Soi97]. Maturation [Qui95]. Maui [A^{+}98, Küc97]. max [SMG95a, SMG95b, SMG96]. May [ACM94c, ABHY96, Ano89b, Ano94f, Ano94r, Ano95k, Ano96b, Ano96-34, Ano96-40, Ano97b, Ano97-37, Ano98c, Ano98u, Ano99i, BBJR96, Bor96, Bos97, Cai94b, CBK98, DE95, FS99, FW96, HS98, IEE98e, IEE99a, LMT98, LG97, LG94b, M^{+}97, Ste96d, Whi96f, Wil97b, Wil99a, And96b, Mac96a]. Maya [Fox95]. MCC [WT95]. McLean [ABHY96]. MCN [Bea95b]. MD [HB98, IEE99a]. Me [Ano93l, Gra96a, TDT95, Vul97]. Mean [Cha96a, RG97a, Dob97a]. meaningful [Hil96]. Means [LaT89, Huc97]. measurement [HT97]. Measuring [PC97]. Mechanical [Min96]. mechatronics [Rol98]. MECS [SMH99]. Media [Ano96u, BB94, Lev95, MKBB95, RC94, Ska99, TB96a, Wel95a, Bat96d, Ber96a, Bou90, COMG96, Kam90b, Pre96, CM96, Mau93, Mau95]. Mediated [MR95]. Medical [Ano98c, BSF98, CC99, Chu98, EMH^{+}95, FBDM95, Gar95, HJS98, JOA98, JD96, MOTA95, PMS97, Pit97, Pri96c, SNLT96, WHC95, YYOO97, Cim96, GP90, IEE96f]. Medicine [IR94, W^{+}98b, Wes99]. medieval [Dob98]. Mediterranean [IEE98e]. Medium [Fry95a]. MEDLINE [DWV95]. Meet [And96b, Wei94]. Meeting [Ano89b, Ano96c, Ano98b, CBK98, Hen90, Rai90, Rai92, RJ93, RJ94, RJ95, RMG98, Wil93a, Wil90, Wil94b, Wil97b, Wil99a, AHB98, Ano90d, Ano94f, Ano95h, Ano95-35, Ano96s, Ano96x, Ano96-39, Ano98r, Bee97b, Bon93, LG97, Whi96f]. Meets [Lin96, UL98, FM99, W^{+}98b, Wes99]. Melbourne [ACM97b, Lav97]. MELECON [IEE98e]. Memorial [MTH99]. Memory [HPS95]. menacées [CBO93]. MendelWeb [Blu95b]. Mental [Ran98]. menus [Spi96a]. Merits [Cov96]. Meshes [ETT97]. Message [DK97, DABM97, HLS97]. Messages [DRBS98]. messengers [Col95a]. Meta [KMS99, PH96, RTW96, FWDV93, WSM93]. meta-databases [WSM93]. meta-dictionaries [FWDV93]. Meta-information [PH96]. Meta-semantic [RTW96]. MetaCenter [BDG^{+}95]. MetaCreation [Kro99]. Metadata [AS94b, BR94, Fic98a, ME99, AS96, LSI98]. Metalanguage [Gra83, Wit86]. Metamorphoses [Gre95]. Metamorphosizing [Gre95]. metaphor [Nad96, PSG92]. metatheory [Wal94]. Meteorological [NK99]. Meteorology [Ano96a, Ano99a]. METEOSAT [Fil94]. Method [LQ97, OA98, Yam93, Yam94]. méthodes [Att97]. Methodology [KHW97]. Methods [CHH^{+}95, DCG^{+}95, JMB95, Pri96c, JRS96, KKK97, Rii97]. Metody [HS97a]. Metrowerks [Kro99]. METU [DAE^{+}94]. Mexico [CL97]. MHEG [Mar91, Mar92, Pri93]. MHTML [PH97a]. MHVR‘94 [BKS96]. Michigan [Bar95]. Micro [Mat97]. Microcomputer [Bea95a, DM95]. Microcomputing [Ano95z]. Microcomputing/LANs [Ano95z]. Microdocument [Ano97u]. Micromanipulation [AVN98]. Microsoft [Ano93n, Ano94k, Ano96h, Ano96-27, Ano96-30, Ano96-42, CN98, KH95a, MCC98, Ros96, Ven97, Wat99, WF98b]. Microsystem [NS95c, NS95b]. Mid [Whi96f]. Mid-year [Whi96f]. middleware [Ano96-41]. midyear [LG97]. migrating [WF98b]. Migration [TW95, HS93]. Miko [Mat98]. military [CH90, Zim95]. Million [Ano96-29]. MIME [Lev96b, PH97a]. MINAR’98 [IS98]. Mind [TD95]. Miniato [Ano94a]. Mining [Min97, Har95b, HY97]. Minneapolis [ACM94c, Ano94r, LG97]. Minnesota [Ano94r]. Minute [Eva96, Zac97, Eva95, Eva98a, HE95, Roa95]. minutes [Eva98b]. Miro [Rut96a, Rut96a, Sto93]. Mirror [KMS99]. missing [Ges91, Nag95]. Missouri [ACM95c]. MMACE [LSW95]. MMIS [KRRK93]. MmNet [IEE97c]. MN [ACM94c, LG97]. MO [Lit96]. Mobile [IEE95g, PD96, GK95b, GK95a, VB95b]. Mobisaic [VB95b]. Model [Ano97-36, Deb97, Gel95, KMS99, KFHU98, Lap99, LK95, MGH93, MGH97, AMFL^{+}95, BCLT95, BM91, BD97, CNH99, Col96a, KCCJ95, YF96, LZSD95, MCP91, MA96, Mat93, Mil93, MR93a, MR93b, RTW96, Spr96b, ZR93a, ZR93b, HB95]. Model-Based [Deb97]. Modeling [ACM99b, And96b, Ano96-42, Bru98, BB99b, BUP99, CIF98, Dan98b, For96c, For99, FR97, MKEB98, Nat99b, NM95b, RBR95, SB97, SR98, SB99, Vac96b, CPK95, FS99, Gre98b, KSM^{+}98, Wan97].

Modelling [BS95c, BQRT95, FB95b, PHL97, VT99, BS95b, Fra96]. Models [AS98, Aho95, Aho97, Bou97a, BK93b, BKW93, BKW97, BKS96, Cov96, Die98, AP96, CLS96, Kil99, KIM^{+}97b, Loe94, Neu97b, Soa98, WB96]. models-pushdown [Neu97b]. Modify [Kle94]. Module [Ham98, Com97a, Com97b]. Modules [Div98, FFKO99, Ric95c]. Molecular [CR96, EKS95, MKEB98, PF95, CLPR96, VB95c]. Molecules [MKEB98, RL95b]. MOMspider [Fie94]. Monaco [GSV96]. Monastery [ABB99]. mondes [Dil97]. monolingual [AT88]. Monolithic [Cov96]. Montage [Ube94]. Monte [GSV96]. Monterey [ACM99b, NDM98, SB97, SR98, USE95, Ass96, USE97b]. Month [Ano96-29]. Montreal [IEE95f, JM96]. MOO [MBH95a, MBH95b, MBB95]. MOO-based [MBH95b, MBH95a, MBB95]. Moodview [ADE93]. Moored [SMM95a, SZM^{+}95, SMM95b]. Morel [DP98a]. MOSAIC [Lin94, MOTA95, Ano95-34, BHK^{+}95a, BR95, Che95, CHH^{+}95, Chi95, CW95, Col95d, DP95a, DWV95, DM94, DL95c, Fab95, FMW95, FP95, FMC95, GMG^{+}95b, HKH95, HD95, Har95b, HE94, HJM95, HMMSS95, Ivl95, KH95b, Lik95, MCPD95, MSIM95, Pet95, PD95, Rak95, RBBN95, Res95, RBBW95, Sav95, SHW95, Sed95, SB95, SZM^{+}95, SMM95a, SMM95b, SGJ^{+}95, Tew95, VSW94, WP95, WCP^{+}95b, WCP^{+}95a]. Mosaic-Based [Rak95, MCPD95]. MOSAIC/HTML [MOTA95]. Mosaic/WWW [BHK^{+}95a, CHH^{+}95]. MosaicForms [GB95a]. Mostly [Fer94]. mother [Com96]. mountains [KL91]. Move [Rad97]. Movement [EH97]. moves [Ano95-43]. movie [GSS90, Sha98, Sul96]. Moving [Ano95l, Dil97, GB94, Dil96, HW96a, HW97]. MP [EGH95]. MPEG [Cla98, Fil94, Mul97c]. MPEG-4 [Cla98].

MPI [KSJ95]. MS [Thr95]. MS-Windows [Thr95]. MSMQ [Moh99]. mSQL [ML97]. MTS [Cha99]. much [Lig95, Tit98]. Multi [Ano96-33, Ano96-31, Ara98, BW97, BE95, Eng97, New92, PH97b, POA^{+}95, RW96b, Sam95, SOUH97, CLS96, HMRL95, Kam90b]. Multi-Agent [Ano96-31]. multi-client [CLS96]. multi-head [CLS96]. Multi-lingual [Ano96-33]. multi-media [Kam90b]. Multi-platform [Sam95]. Multi-Protocol [POA^{+}95]. Multi-Purpose [BW97]. Multi-Server [SOUH97]. multi-tail [CLS96]. Multi-User [Ara98, BE95, Eng97, PH97b, HMRL95]. Multi-vendor [New92]. Multiconference [DG97, LB98a, COW98, BUP99]. multidimensional [CC99]. Multilingual [SMDO98]. multilingualism [Tak96]. Multimedia [ACM92, ACM93d, Ano90g, Ano97w, Bat96b, BHK^{+}95a, BKS96, Bry98, CM96, CPK95, Chu92, DK97, DS96, DL95c, FD91, Fox95, FJV96, FMC95, Heb96, HK94, HF97, HS98, IEE96l, IEE96h, IEE97c, IS98, KJR98, Kaz95, KAH91, KR95, Law98, LRG96, Mau93, Mau95, MH96, MK96a, New95a, New95b, OISEM97, PBFT97, Rut96a, SV98, Shi93, Soo95, W^{+}97, Wha92, ZMP96, Ano95-27, Ano96-42, Bat96a, BR94, CJ90, De 94, Erf93, Fis97, Ish95, Kap97, KDC95, KTK^{+}95, Kom94, Lin94, Loc97, OSEMV95, RB96, Rus98, WKD96, Ano95-27, SGBO^{+}93, Ano90g, Ish95, RS93]. Multimedia-its [Shi93]. Multimedia/Hypermedia [HK94, Kom94]. multimediaapplicaties [VB95a]. Multimedialne [Dan98a]. multimedialnego [CK98]. Multiple [Bro97, Pre96, LDW^{+}96]. Multipurpose [LS99]. multivariate [TC96]. Mumbai [V^{+}96]. Munich [Dug97]. Museum [Rak95, RBBN95, Dra96]. Museums [BT97, Blu95c, Nic97, Bea95b]. music [New90]. Musical [GHH97, MK88]. must [HW94, LM95a, Wil96b]. Mutual [ZR93d, ZR93c]. my [Sed95]. MYRIAD [HLY^{+}94]. Myth [Gro97]. Mythology [Fox95].

NAECON [IEE93a]. Names [BS90]. NAP98 [MTH99]. Naples [Ano96x]. NASA [NB95, Tow95]. Nashville [Ano95m, Ano98t, Mas97]. Nation [DT95]. Nation-wide [DT95]. National [ABY94, Ano89b, Ano97v, Ano97z, Ano98a, Ano98q, Cha96b, CBK98, IEE93a, MTH99, MM97, Wil93a, Wil90, Wil94b, Wil97b, Wil99a]. Nations [Ano96-33]. Native [Lot95]. Natural [BMK95, DWV95]. Nature [Tex99]. Naughton [Ubo96]. Naval [FMBM95]. navigating [CJ94, DBCK96]. Navigation [Ano97g, Döm95b, DWV95, EH97, Gla95, HRS95, KK98, ORCK95, Pri97, CR97, Döm95a, GF93]. NaviSoft [Ano95-29]. NC [ACM96d, Ano96-30, Ano98d, Ano99h]. NC4 [HU97b]. NCGA [Ano89c, Ano90h]. NCSA [BC95, Che95, CB95b, Har95b]. Near [Tow95]. Near-Real [Tow95]. nearby [FGGT96]. NECC [Ano98a]. Need [Che97, Ano97-54, HH98, H^{+}96, Mee96, M^{+}96a, Tay90, Wal94]. needed [Vac96b]. Needs [KAKZ96, WT93]. nerves [Ano96g]. nested [GW87]. Net [Ano95-28, Kas95c, Kro99, Law98, Mil95, Arn95, Coy97, Hil96, Man95, Bat96d, Not96, Tub95]. net-wise [Man95]. NetBiochem [BD95]. Netherlands [ACM93c, Ano90c]. NetMul [TC96]. Netscape [HU97b, Ano96h, Ber96a, HK96b, HK96a, HK98, HL97, Kid97, KH97, LD95a, Lej96, Par95, Tat96]. NetTerminal [Ano97w]. Network [DM97, Fat97, Sch90a, Ano95-43, BS94, Bat96d, HKM^{+}96, Hel96, Hut96, IEE96f, Jan96, KDC95, Kik96, McC97a, MM97, MM95b, NANK96, Pao95, Pet90, RMF93, Sed95, Wei96a, EMH^{+}95, Sch90b]. Networked [BHK^{+}95a, CPR98, FNN97, Wei96b, PH96].

Networking [Ano92a, Coy97, FJV96, IEE97c, IEE98f, PK94b, ACM95c, Hey94, KJR98]. Networks [IR94, MFVG99, PS95, BA90, Com96, Cov99, De 94, M^{+}97, Oke93, Soa98]. Networks/International [PS95]. Networkshop [Ano88a]. Netz [Bat96c]. Neural [IR94]. neuroimaging [CNH99]. neutrality [Wei93]. Nevada [AA97]. Newport [IEE97d]. News [Ano96-29, BS98, Cla98, Gar97a, Gar99, Hir95a, Hir95b, Hir95c, Law98, McL98, Par95, Wat95, OSEMV95, OSEM^{+}96]. news-on-demand [OSEMV95, OSEM^{+}96]. newsletter [SGM87, SGM91]. newsletters [HH96, HB96]. newspaper [Wal95]. newspapers [RV94]. NEWSTRENDS [Ano96-30]. Next [Ano95-28, IEE94a, Koh95, Pes96b, SBGO99, And97, And96c, Ano95-32, Ano97-55, AAM^{+}96, DJ96, Hoy96, LMO^{+}96, Sch96c]. next-generation [DJ96, LMO^{+}96]. NHS [Hut96]. NHS-wide [Hut96]. Nice [van88]. Night [Pie97, RCRBR97]. Nightmares [Fic98a]. Nimes [Col98]. Nineteenth [Gar95]. nineties [Mac96b]. Ninth [Ano97z, IEE97d, Ste96e, To98, TW98, USE95, Joy98]. NIST [ABY94, MBB90]. NLM [DWV95]. NM [Sol88, BBJR96, Bor96]. No [Ano96-30, DP98a, KH95b, Men95b, Ano98r, BVN95a, BVN95b]. No.103 [Ano93g]. No.106 [Ano91e]. No.1993 [Ano93h]. No.1993/170 [Ano93h]. No.1994 [Ano94h, Ano94g]. no.1994/142 [Ano94g]. No.1994/144 [Ano94h]. No.1995 [Ano95p]. No.1995/202 [Ano95p]. No.1996 [Ano96u]. No.1996/148 [Ano96u]. NOAA [DT95]. Nodes [Ran98]. Nomikos [IEE97a]. Non [Alv99, GB98, Pol95]. Non-Intrusive [Pol95]. Non-Textual [GB98]. Non-validating [Alv99]. Non-Visual [GB98]. Nondestructive [Mel99]. nonlinear [MBHL91]. nonstop [Cel92, Eng94, EGH95]. nontraditional [IEE96o]. Norsk [Ano97f]. North [BCG99, CL97, RS97]. Northrop [Ivl95]. Norway [Ano96f]. not-for-profit [Oke93]. Notation [CMV98, BH95b, BFH98, Hei95, Roe87]. Note [Gog90, MOTA95]. Notes [Ano96-30, Lot95]. Nov [W^{+}98a]. Novel [CACS94, GW94, KK98]. Novell [Man95]. November [ACM94a, ABY94, Ano93j, Ano94e, Ano94o, Ano95h, Ano95m, Ano96-39, Ano97d, Ano98l, Bac96, BFY93, Bow97, CPK95, Chu98, Gar95, HF97, IEE97c, IEE97d, Isk96, Mei97, Min96, MB97b, MDK98, MR99b, NDM98, Oke95, VD98]. Novice [NF97]. NSW [DE95]. NT [Mac98, WF98b, Ano96-30, KPIE99, Neo98, Neo99]. Nuclear [Ano96x, NJDF96, SR93]. nuls [TJD98]. Numerical [MTH99]. Nursing [GJTW97, OSE97]. Nuts [Lun95a, Lun95c]. Nutshell [Ano94i, Fla96]. nutzbringenden [Bat96c]. nutzen [Hie97]. Nutzung [Wie94]. NY [Ano94f, IEE96c].

Oahu [IEE96f]. OB [Lut98]. Object [Ano96-37, Bee97a, Bla94, Boo96, CR96, Die98, DAE^{+}94, DAOD95, ETT97, FFKO99, FB95b, KHW97, Kel98, Kul94, LHL97, OISEM97, OL97, PH97b, PST97, C^{+}98, DJ96, Kim94, KZSC96, KSC96, Mil93, MR93c, NR96c, OSEMV95, Pri93, RMF93, Whi96b, Kro99]. Object-Based [ETT97, KHW97, Kel98, Mil93]. Object-Oriented [Ano96-37, Bee97a, Die98, DAE^{+}94, DAOD95, FB95b, Kul94, LHL97, OISEM97, PH97b, PST97, Boo96, C^{+}98, Kim94, KZSC96, KSC96, MR93c, OSEMV95]. object-relational [DJ96]. Objects [Ano97k, Col96b, Deu95, KC92, PBFT97, Rag96a, VT98, Dob97a, Liu97b, MN95a, Paq94]. objektrelationales [RH97]. obsahem [ID98]. Observations [Kim94, GP90]. Oceanography [Ano96a, Ano99a]. Oceans [IEE95b]. OCR [Gan96]. OCRed [HS97b]. October [ACM94a, ACM95b, ACM95c, ACM96d, ACM97a, Ano90d, Ano95-34, Ano98r, Bea95b, BB94, Bon93, CEG96, Cim96, Col98, Gar95, Gau96a, GE95, HW94, HW96b, IEE95b, IEE97b, IEE98b, LRG96, Lit96, Mas97, RS97, TB96a, Wae97, WS95b, YH97, dN97]. Octree [Sch97b]. Octree-Based [Sch97b]. ODA [BBB88, sta89, KMK95, NB91, Mol91, NW92, Sch92, WSM96, WD90, WD91, Yam93, Yam94]. ODA/ODIF [Sch92]. ODIF [Sch92]. ODIL [LR95]. ODMG [Kim94]. ODMG-93 [Kim94]. Off [Ano97c, Lyn96]. offer [Ano96i]. offers [Ubo96]. Office [Ano90a, Int86b, Int86a, Ame85, AC93, Ame91, Can89, Int86c, Int86d, Int88e, MM90a, YKT90, Gri95]. Official [HK96a, HK98, Kid97, KH97, WF98b]. offs [MD94]. OH [Bon93]. Ohio [Ass87a]. Old [Rad98]. OLE [Gro95]. Olympic [MTH99]. omits [Smi96a]. Omni [Ano95k]. OmniPage [Gan96]. OmniPort [FS95]. On-Demand [DP98c]. On-Line [Hir95a, Hir95b, Jon95, MB97a, Ano97-35, Dou92a, RV94, Sea96a]. OnCenter [IEE96c]. ondersteuning [VB95a]. One [Ano96-29, Cea96a, Mee96, Ano92c, Ano97e, Bus88, DT95, GB95c]. Online [ACM92, Ano89b, Ano95-30, Ano95-31, DLSR98, DR95d, Eft90, Gol89, HN96a, HG96a, Hir95c, Hoe97, HG96b, Mar95c, NPPR95, Rai90, Rai92, RJ93, RJ94, RJ95, RMG98, RM97a, VN95, Ven96b, Wil93a, Wil90, Wil94b, Wil97b, Wil99a, Ano96-43, Bat96b, Dou92b, GW94, GP90, HK96b, HH96, HB96, LMO^{+}96, Rat98, Ric96, Roc93a, SGBO^{+}93, Ano95y]. Only [BS90, Ano94t]. ons [Men95a]. Ontario [WNC95, Roc93b]. ontology [SP99b]. ontwikkelen [VB95a]. OODB [Zha95]. OODBMS [CR94, Sjo95]. OODBMSs [ADE93]. OOHDM [SRB96]. OOHS [KSC96]. OPAC [Hen96b, Hil96]. Open [Ano93r, Bla94, BB90, CHR^{+}96a, Fer94, HK94, ML97, RP98, Tau96, BB91, Mol91, Sch92, Wol95, Yam94]. open-document [Yam94]. opening [Ano97j]. OpenPath [Ano97w]. opens [Mio96]. OpenWorlds [DMH98]. Operating [IEE95d, WF98b]. Operations [FP95]. opinions [Ubo96]. Opportunities [BHK^{+}95a, Oke93]. opportunity [WE92]. Opryland [Mas97]. Optimal [SM93a]. Optimization [GW87, Kro99, GI96]. optimized [TS99]. Optimizing [NAM96, vB89, HS93]. Options [Bjo96b, Bjo96c, Wes99]. Oracle [Hob98]. Orchestration [GNSP95]. Order [Ano96-33, Kle94, Hol99a]. Oregon [Can95]. ORES [TABA^{+}94]. organisations [Lun96]. Organizational [BHK^{+}95a, PC97, And88, ERPNS97a, Kik96]. Organizational-Wide [BHK^{+}95a]. Organizations [Ano93a]. Organize [LS98]. Organizing [WCP^{+}95b, WCP^{+}95a, ECvR93]. oriental [Ano99f]. Oriented [ACM96d, ABO98a, ABO98b, Ano96-37, Bee97a, Die98, DAE^{+}94, DAOD95, FB95b, Kro99, Kul94, LHL97, OISEM97, PH97b, PST97, Boo96, C^{+}98, Kim94, KZSC96, KSC96, MR93c, OSEMV95, WF89, WF90]. Orlando [Ano95a, Ano98o, COW98, Chu98]. Orleans [ACM96a, Gar95, IEE96k, IEE96o, Su96]. OS/400 [Ano96y]. OSI [Smi87a, Smi87b]. OSU [Wal90]. Other [Fox95, LZ95, PHL97, Ano94t, Cav94, DWV95, Eft90]. OTS [Cha99]. Our [ACM95b, Ude95b, IEE95b]. outer [Dav92]. Outerjoins [GL94]. outil [San94]. Outlines [Lot95]. Outlining [HJM95]. Output [Bjo96b, Bjo96c, Fer93a, LS98, Sun98b]. Overhead [Ran98]. Overload [Bur93]. Overview [Ada92, Eng93, Jah99, PCL^{+}95, Roe95, Kom91, MM95a]. Ovid [Gre95]. Own [FD95, FD96, KH95b, Pau95, Ste96b, Col95b, HN96b]. owner [Int88a, Int88b, Int91a, Int91b]. Oxford [Ano88g, Ano94p, Bra89, Faw93].

P.M [IEE97a]. PA [AR97, Ano98c, GSG98]. PAAM [Ano96-31]. Paced [Ten96]. Pacific [A^{+}98, DS96, HK95b]. pack [Ano95l]. packages [Mar94, Uns96]. Packaging [Kro99]. Packard [Ano93f]. PACS [JD96]. Paderborn [SB99]. Page [Ano95-29, Ano96-32, BKT96, Bly96, Boa95, Con93, MH96, Run95, Tes96a, Ano94q, BIL95, Com97a, Com97b, Dun95f, Gro96, HN96b, HC97, Hic94, H^{+}96, How96, KH97, Lam97, Lem97, McF96, McF97a, McF98, Mee96, O’C97, SL97, Sko98, SP96, Zim97, Sra96]. Page-on-demand [Ano96-32]. page/sites [O’C97]. PageMaker [Ano95-32]. PageMill [Gra96e, KS95a, Boh96]. Pages [Aus95, FD95, FD96, Ham97b, McC97b, NJDF96, Nør99, SS98, Ste96b, Tay95e, Tay96, Tay97a, Tes96b, Ano95g, Ano95r, Ano95-42, BH98a, Car98b, Car98a, Car99, Coh96, DM95, DM96, HH96, HB96, HH98, H^{+}96, Jon95, KLM94, Lil95, Okl97, Rag98, Ric94, RKW96, San99, SL97, Tay95c, Tay95d, Tay97b, Tay98]. pagina [McF97b]. paginas [Alv97a, Alv97b, HC97]. pain [CNH99]. Palaeobotanic [GB95a]. Palo [BBKF95, NW97]. pan [mWfC97]. Panacea [GDRR94]. Panel [Ano90d, BSF98, Wid94, Deu95]. Paper [Ano88d, DP98a, For95b, Myk95, RG97a, KL91, RR94, Ano96-27]. Papers [Ano95k, Ano97x, BKS96, IEE92, IEE96e, Ner97, Neu96, Ano90c, Ano90d, Ano97a, Ano97-42, Bea95b, BT97, Cai94b, Cha96b, IEE96a, IEE98d, Kim97a, Kum96, LG97, MM97, MDK98, W^{+}98a, ABHY96]. Paradigm [SF96, Chu98, Hey94]. Paradigms [Swa96]. Parallel [Hil96, HMMSS95, IEE96b, OMY97, SN95, vB89]. Parallelism [EGH95, Cel92]. Paris [Ano96v, Ano96-34, Heu98, RM95b]. Park [ACM96d, GSG98]. ParseMe.1st [McG98a]. Parser [And99, HLL93, WV89]. Parsing [MGH97, Con93, HW88]. Part [Dob97a, Dob97b, EM99, Int91c, Int92a, Int92b, O’C92b, O’C92c, New95a, New95b]. Partie [Int91d]. Partitions [GLTH98]. partners [Ano97-27]. Partnership [Gar99]. parts [BP93, Hug96]. Pascal [Buc97]. Passing [HLS97]. PAT [PW94]. Patent [Gri95, BHMS98, Bre96]. Patiala [MM97]. Patient [PGRC97, RG97b]. Patricia [HW96b]. Patrick [Ubo96]. Pattern [Ber96c, IEE98a, Ber96b, Ber95]. Patterns [Lut98, OSYM97, PR95b, Rem97, Sed95]. Pavement [LHL97]. Pavilion [ACM95c]. paying [Oke95]. PC [Ano96-30, Gra96b, LG94a, Sch95b]. PC-based [Sch95b]. PDAs [GK95b, GK95a]. PDF [Gut98]. PDM [Ano94v]. PDR [ZKM95]. Peloton [CCEE98]. Pennsylvania [ACM96c, LRG96, TB96a]. People [Ano93g, CBK98, Ano93k, KAKZ96, Pin95, RW98, Huc97, PK95a]. Peoples [Reg98]. Percentiles [AS93a, AS93b]. Perception [HN96a, CP96]. Performance [LHMM97, OMY97, SD98, AYE^{+}96, Ano99d, BP96, Eng97, LM95a, LDW^{+}96, NAM96, Ten99]. Performing [Van98]. Period [Mue95]. perl [Xu97, Coh96, MN95b, Pau95, TGHE96, Ude95a, Ude95b]. Persistent [Ric95c]. personal [KMA^{+}96, PAV96]. Persons [Mur92a, Mur92b, ZBW94, BEE95, WT93]. Perspective [Del95a]. perspectives [Car98b, Car98a, Car99, Kam90a, PS96]. Perspektiven [Ano99j]. Perspektives [Ano99j]. perusal [BIL95]. PET [CNH99]. Peter [Cam94, HW94]. Petersburg [Mel99]. Peterson [HV92]. petroglyphs [CL97]. Pflege [Wie94]. PHIGS [Tho93]. Philadelphia [ACM96c, AR97, Ano96s, Ano98c, LRG96, TB96a]. Philosophical [Woo95a]. Philosophy [Blu95c]. Phoenicia [LK95]. Phoenix [DG97, USE87, YH97].

Photoshop [DM96]. PHP [ML97]. physical [Wes99]. Physics [Ano89a, Ano94a, CT96, Nad90, Ano94f, Van89]. Physiology [CC99]. Pictures [BMK95, JCDA95]. Pilot [Gre98a, Sti94]. pimil [HiKgK98]. Pioneering [CL97]. Pipeline [Car97, Oke95]. piper [Oke95]. pitfalls [See96]. Place [LHMM97, PGRC97, Lig95]. Placement [Hel94]. Places [Sha95b]. Plain [Edd97a, Edd97b, Edd98a, Edd98b, Edd99]. Plan [Cla98, Tho97]. Planetary [SS95a, HB95]. Planning [Coy97, Dim96, EH97, Ste96a, IEE96c]. Plans [Lot95, Wat95]. Plant [MSMY97]. Plants [FMC95]. Plasma [IEE95a]. Plate [Ano98g, Ano98h, Ano98i]. Plates [Ano98j, Ano98k]. Platform [STMS95, HBNV97, Rag94a, Sam95, Ski94, Web95b]. Platforms [Fer94, LDW^{+}96, Mul98, Wha92]. Platinum [LO96, LOA^{+}96, Lad96, LO^{+}99b]. Play [Kay89]. plays [Ano95l]. plea [Win96]. Please [Ano96-30]. pleases [GMR96]. PLINTH [Cas94]. plug [HL97]. plug-ins [HL97]. Plugging [Gul96, Wil95]. Plus [KFZ97, Pow97, Ano96g]. Plzen [Ska98, Ska99]. PNG [Car97]. pocket [Eck99, Eva97a, Eva97b, TS97, TSP98]. PODDP [MNW98]. PODP’96 [NW97]. podrobny [Kos98]. PODS [ACM94c]. Point [Ano96-30, PST97, Ano98c, Tit98]. Poison [GDRR94]. POL [Chr95]. Poland [Dan98a, RM96a]. Policy [VS95]. Poll [Ano96-29]. polohovanim [ID98]. Poly [Van98]. Polygonal [GLTH98]. pomoca [CK98]. Pooh [Web95b]. Pop [Tub95]. Populating [Bro97]. PORIDGE [Wal94]. port [Ano96-32]. Portable [Ano96-32, CRBB98]. Portugal [Ano98m]. Possibilities [WD91, RW96a, WD90]. post [Dra94]. post-hoc [Dra94]. Poster [WF98a]. posters [Ska98]. postmodern [Wal94]. Pot [Ano96l, Ano96r]. pour [Int88i, Int91d, San94, TJD98]. Power [Dun95d, Dun95e, Dun95f, Dun95a, Dun95b, Dun95c, Hin98, NR96a, Opp97, Ven96b, JMS98]. Powerful [Lot95]. pp [BC90]. practica [Fer97]. Practical [Ano96-31, Ano98f, Dun95g, Fly98, Glu96, GFR98, Hel94, Phi98a, Pop92, Pop94b, Sch95b, Smi96b, Smi97b, Smi98, Ten96, van90b, van94, Ano97d, Chr95, GS95b, RM95b, Szi95, Glu96, Joh95b, Pri96a, Sch96b, Smi96a, Van91b]. practicality [Sch96b]. Practice [Ano92h, DGBH93, DC98, Gon94, IEE98b, Rie94, WS95b, Wu89]. Practices [Ano91e, Bla90]. Prague [GSV96, PK94a]. praktische [Szi95]. Praxis [Rie94]. pre [Ano95l]. pre-press [Ano95l]. Predicate [HS93, Hel94]. predicates [HS93]. Prelude [BPBV96]. premastering [Adk88]. Premier [Lem96a]. Prentice [Eva97a, Eva97b]. Preparation [vH87, Org91, Smi86a, Ass87a, Adk88, Gre93, MB90, MB91, War87]. prepares [Mar91]. Preprocessing [DHR97]. prerequisite [BEE95]. Prescriptive [Hui96]. Present [JDGP95, Dun95e]. Presentation [BS90, DLSR98, Fre95, FMC95, Fry95b, JCDA95, JJ95a, KW97c, PBFT97, Toy98, NL95, RB96, NB91]. presentations [WKD96]. presented [Ano90d, Kum96, Min96, MM97, Ner97, Neu96, W^{+}98a]. Presenting [Bot92, Lig97, MBK^{+}95, Rut96b]. Preservation [HW96b, CW97]. Press [BC90, Ude95b, Ano95l, Sch95b]. price [EGH95]. Pricing [Var95, MMV95]. Primary [Cam94, Tea97b, Tea98b, Rob94c]. Primer [Ano99c, Dun95d, Pow95, Sof91a, Sof91b, Spi96b, Ano97o, JT97, JT98, St.97a, St.97b, St.98, St.99b]. Princess [Ano97-37]. Princeton [W^{+}97]. principled [VCB97]. Principles [ACM94c, Con97, MNW98, MRR99, NW97, Ost94b, New95b]. Print [Dod94, Ano94u, Ano95l, BEE^{+}94a, BEE^{+}94b, BEE96a]. printed [BEE95]. printers [Bla90]. Printing [Ano88d, Ano95k, Ram97, And97]. Privacy [Lut98]. Pro [Gan96, How96, Har95a, Ano97-51, Ano95n, Ano95-44]. Problem [KFZ96, PMK98, ZFK97]. Problem-Solver [KFZ96, ZFK97]. Problems [BB95b, MKBB95, SS97, Ano97d, BHK^{+}88, BBG^{+}95a, Dou92b, LM95a, WZ91]. Procedures [YYOO97, Eis96, Int88a, Int88b, Int91a, Int91b]. Proceeding [Isk96]. Proceedings [ACM92, ACM93a, ACM93b, ACM93c, ACM93d, ACM95b, ACM96b, ACM96c, ACM96d, Ano88f, Ano88g, Ano89c, Ano90h, Ano91b, Ano91c, Ano91d, Ano91f, Ano93f, Ano93a, Ano93p, Ano94l, Ano94m, Ano95e, Ano95-33, Ano96k, Ano96-34, Ano96-35, Ano96-33, Ano96-31, Ano96-36, Ano97-27, Ano97y, Ano97z, Ano97-52, Ano98q, Ano98r, Ano98s, Ano99b, Ano99i, Bac96, BBJR96, Bos97, BGG^{+}94, BDV95, BUP99, Cai94a, Cha95b, Col95c, CBO93, DGBH93, Dés86, DS96, EAZ98, FD91, FJ92, FW96, Fur90, GJTW97, Hen90, HK94, HF97, Ano93k, IEE93a, IEE93b, IEE94a, IEE95c, IEE95e, IEE95d, IEE95f, IEE95g, IEE96k, IEE96l, IEE96i, IEE96h, IEE96m, IEE96j, IEE97f, IEE97e, IEE97d, IEE97g, Ibr95, IR94, JM96, Joy98, Küc97, KR95, LB98a, Len93, Lic91, LM95b, Mau93, Mau95, MWEK97, MBB90, MG94c, Mur92a, Mur92b]. Proceedings [NP96, NDM98, WWW95, PK94b, PS95, Rai90, Rai92, RJ93, RJ94, RJ95, RPWW96, SSH87, SS89, Sch93, SB97, SR98, SB99, Ste96e, Su96, TS96, TR92, TW98, USE87, USE89, USE95, USE96, USE97b, USE98, VD90, V^{+}96, Wil94a, Wil93a, Wil90, Wil94b, WNC95, ZBW94, van88, ACM94b, ACM94c, ACM95c, ACM96a, ACM97b, ACM97a, ACM99b, ABY94, AHB98, AR97, ABB99, Ano89b, Ano94r, Ano95j, Ano95m, Ano95-37, Ano96s, Ano96b, Ano97d, Ano97b, Ano97-37, Ano97-42, Ano98u, ABG96, BKS98, BFY93, BH95b, BFH98, BCG99, COW98, CM96, CBK98, C^{+}98, CL99, CL97, Cim96, Cov99, Coy97, Cze98, DG97, DE95, DE96, DP98b, Dob98, FS99, Gar95, GE95, GSV96, Gra96d, HAB98, Heu98, HS98, IEE95b, IEE96d, IEE96n, IEE96o, IEE97a, IEE97c, IEE97b, IEE98c, IEE98a, IEE98b]. proceedings [IEE98e, IEE98f, IEE99a, IEE99b, IDK98, IS98, Jac98, JRV96, Kho94, Kim97a, K^{+}98, Lav97, LMT98, LR94, Mas97, MTH99, M^{+}97, MNW98, MB97b, NW97, NS98, Oke93, Oke95, PKM99, QSBC98, RMG98, RM96a, RS97, RT95, Rii97, RM95b, Rut96a, SSK97, Ska98, SC98, To98, Van95b, VC96, Vez96, Wil97b, Wil99a, YH97, dN97, Nad90, Neu94, Sol88]. Process [Kro99, RM97a, Bro94a, Bro94b, CG95, Cha95a, See96, W^{+}98a]. Process-Oriented [Kro99]. Processes [Ano93k, Wu98]. Processing [Ano91d, Ano96a, Ano99a, BDV95, BK93b, DWV95, IEE97a, Int86b, Int86a, Int88c, Int88d, SNLT96, Smi92a, Smi92b, SS97, Sol88, Tan98, W^{+}97, WV91, dN97, Ame85, Ano95g, Ano95r, Ano95v, Ano99f, App89, BB90, BB91, Bre96, Bry93, Can89, Can90, CHP88, sta89, Dob97a, ECvR93, Fil94, Gv92, HW90, Int86c, Int86d, Int87, Int88f, Int88g, Int88h, Int88a, Int88i, Int88b, Int88e, Int92a, Int92b, Intxx, Kan93a, Kaz95, MCP91, MZZ96, MM90a, MNW98, NW97, Pet90, San94, Sch92, To98, VTW94, Won90]. processor [Alv99, Sum94]. processors [DGG^{+}92]. Procurement [NPPR95]. Produce [Sun98a, CP94]. Product [DLSR98, LQ97, Sha97, Ano97h, Cha95a, LJ93, Ore88]. Production [Ano93i, Dod94, Pop97, Ano95-27, Ano98n, CP93, HK96b, KH97, WF89, WF90]. Productive [Ano97p]. productivity [KTK^{+}95]. Products [Ano97w, DP98c, Fee98, Ano93k, Sch88, SS95a, GPE98, Shi93, WS95a]. Professional [Bou98, IEE97b, HH98, LJ96, Ano96o, Ano96w]. professional-looking [HH98]. Professionals [Fun99, LB99b, Ano99l]. professionnelle [Ano96t]. Profit [Mor95b, Mor95c, MR96, MS98c, Oke93]. profound [SF96]. Program [AS93a, Ano95k, Ivl95, OSE97, Tea97a, Win97b, ZJ96, Ano92c, Liu98, Ska98, Thi91]. programador [JSY98]. Programmation [CC98, Fra97b, LO99a, WM98]. Programmer [Hom99, Wri96, FHU99, HUW98a, JSY97, Mul97b, Mul98, PW98]. Programmers [Fla96, NF97, CH99b]. programmeur [CDCD98, PW99]. Programmierung [Amm97]. Programming [Ano97-29, Ano97-28, BL98, Col96b, Dun95d, Dun95e, Dun95f, Dun95a, Dun95b, Dun95c, Far96a, Gra96b, Hug96, KK95, MHP98, Nør99, Swa96, Via97b, Via97a, Via97c, Via98a, Via98b, Bar99, EM99, Gre98b, HL97, MG94b, PD96, SBS95a, Smi96a, Spe99, Tit95a, TGHE96, Ude95b, Wat99]. Programs [Ano96-30, DCG^{+}95, Fly98, Nie99, VH95, GJ96, Ski94]. Progress [CDB91a, Vez96, WF98a]. Progressive [GLTH98]. Project [Ano97i, FM99, GK95c, KMK95, LSC96, Mar99, OBO97, Ple95, Sti94, Ada95, Chr95, Eng97, Eng93, Liu97a, Rot97, Bra89, Faw93, Gre95, LZ95, MD97, Pul95, WT95]. Projects [Eri97, PSG92]. PROLAMAT [Jac98]. Proliferate [Ano95s]. Prolog [RM95b, Tay95b]. Promise [Kay89, BVN95a, BVN95b, Mas98]. Promises [Law98, Ano93r]. Promotion [DP95b]. PROMS [IEE97c]. PROMS-MmNet [IEE97c]. proofing [Ano95l]. Proposal [Mag94, Mar97a, Soi97, Tau96, Kim94]. Proposed [Ame91, Sei96]. Proposes [Cla98, Web95b]. prospects [ABH93, GB95c]. Protecting [Lut98]. Protocol [Bro98a, NM95a, POA^{+}95, BLFF96, FGM^{+}97]. Protocols [IEE97c].

Prototype [HMHK97, HLY^{+}94, CL90]. Prototyping [HLC96, Hue98, MSW97]. ProtoVIew [Kro99]. Provide [For95a]. Providence [Ano97z, MR99b]. provider [Kel94a, Kel94b]. Providers [HAD^{+}95, Rat98]. Provides [Ano92b]. Providing [HK95a, Hel96, JMB95, LZ95, MSIM95, Mue95, VH95, MS95]. provisional [Hei95]. prudently [Win96]. pruvodce [Kos98]. przykladzie [CK98]. Psychology [Ano96c, Ano96-39]. Public [Blu95c, Int91c, Lan91, RG95, Tau96, TRMW99, Int88a, Int88b, Int91a, Int91b]. Publication [Ear87, Str95, Dou92a, Kru90b, See96]. Publications [Anoxx, Smi86a, Zim95]. publikacji [HLB98]. publiques [Int91d]. Publish [DP98a, HN96b]. Publisher [Ano96w, FD95, FD96, FS96, Gra97a, Ise90, Man95]. Publishers [Fly95, Kay89, Smi87f]. Publishing [Ano93e, Ano95s, Ano95-30, Ano95-40, Ano95-41, Ano95-44, Ano96q, Ano97-30, BMDB98, BVN95b, BQRT95, CFMG95, DL95b, Dun95f, FD95, FD96, Fry95a, Fur90, HK96a, Hes97, HH98, JJ95b, JJ95c, Lem96a, Lem96c, LD97a, LT98, LT99, Neu96, Pfa97b, RLA96, Res97, Ric92, Rou95, RM96b, Sea95, SS97, Toy98, Ven96b, WMGL95a, Wei94, Wit96a, van88, ABH93, And97, And88, Ano92e, Ano93r, Ano95-29, Ano95-31, Ano98l, APH98, Bar97, Bar98b, Bat95, Bat96b, BC96, Bri98a, Bri98b, BBKF95, CG95, Chr96b, CC96, Cze98, DL95a, Don97, Dor95a, Dor95b, FM99, Fis96b, Gol97, Goo94, Gra96e, HK96b, Hei97, HiKgK98, HAB98, HH96, HB96, Ish95, Ker97, KH95a, Kid97, KH97, Kru90a, Lem95a, Lem95c, Lem95b, Lem96b, LD97b, LS99, LHW96, LG97, NR96b, Neu97a, Oke93]. publishing [Pal98, Pfa98b, Pfa98c, Pfa99, PJ95, Pop91, Pre96, Rob96, Ros88, Sch95b, Smi93, SD96b, Sta97, Tat96, Tay90, WMGL95b, WZ91, WR95, YKT90, Yuz89, van95a, vSSW93]. publishing-what [Tay90]. publishing-why [Dor95a]. Publizieren [Neu96]. Punjab [MDK98]. Pure [Dos99a]. Purpose [BW97]. Push [Cla98, GS95c]. push-caching [GS95c]. pushdown [Neu97b]. Pushes [Law98]. Put [Ano96-32, Flo96, Mag97, Ano95-31, Sal96]. Putting [Ano88f, Myk95, Roc93a, Pul95]. Python [Ham98].

Q [Ber96a, OST94a]. QoS [AAM^{+}96]. QOTA [BW97]. QSBA [RS96]. Qualität [Neu94]. Qualitative [BAC93]. Quality [RARH99, BBKL94, Bow96]. quantification [Fra91]. Quantitative [Ste95a]. Quark [Bat96a, RKW96]. QuarkXPress [Ano95d]. quasi [RB96]. quasi- [RB96]. Quebec [JM96, CBO93, CBO93]. Queensland [DE96]. Queries [Kle94, DSDT96a, ELW95, GW87, GI96, HS93, DSDT96b, vB89]. Query [BS90, CACS94, AMFL^{+}95, BA90, DFF^{+}99, GI96, KS98c, Mac91, Rod92, RS96, YHHYK96]. Querying [CR94, VT98, VT99, AS96, CCD^{+}99, DBCK96, LSS96]. Question [RM95a]. questionnaire [Liu98]. Questions [GRSM97, WM92, Bow95]. qui [San99]. Quick [Fal97, Fla96, Mul96a, Mul97a, Sch97a, Spi96a, TAB96b, Avi98, Gra97a, RR96b, RR97a, RR98c, Sch95a, Sch96a]. Quickly [GSE^{+}94]. QuickStart [Cas96b, Cas97]. QuickTime [Cla98]. Quiet [Kar93, WZ95]. Quill [Wol91]. qwertz [Gor91, Gor95].

R [LO99a, W^{+}98b, PSG92]. RAF [Ano97r]. Raggett [RLAK98]. Raising [Bar94]. RAMA [Del95a]. Ranges [Gar99]. Rank [Gri95]. rapid [Hue98]. Raster [HAB98]. rationalisation [Dra94]. RAVE [EAU96]. Raymond [Ano98f]. RDBMS [Fer94]. RDBs [KMS99]. RDF [KMS99, Mas98]. Re [Isk96, Kea96, MR93b]. Re-Engineering [Isk96]. Re-locating [Kea96]. re-use [MR93b]. RE’99 [IEE99b]. Read [Fer94, Gib98, TDT95]. Read-Mostly [Fer94]. Readability [CGLV94]. readable [FWDV93, Ret96]. readers [BEE96a]. Readiness [IEE94a]. readings [GSG98]. README.1ST [Pri96b]. ready [BVN95b, BVN95a]. Real [GMG^{+}95b, Gro97, Lut98, Mat97, MWVE99, NJDF96, Tow95, Ano97d, GMG^{+}95a, IEE97e, Mul97c, SZM^{+}95, SMM95b, SHO^{+}96, Uhl95, EAU96]. Real-time [MWVE99, NJDF96, SZM^{+}95, SMM95b, Uhl95, EAU96]. real-world [Ano97d, IEE97e]. Realistic [Ell98]. Reality [ACM99b, And96b, Ano95-28, Ano95-43, Ano96-30, Ano96-42, Bow97, BKS96, Bru98, Dan98b, Die98, FH97, Flo96, FR97, Gro97, MCPD95, MD98b, MN95b, NM95b, NM95a, Roe95, SF96, SB97, SR98, SB99, UL98, Vac96a, Vac96b, BHZ96, BFM96, FM99, Gre98b, HT97, Hen96a, Mas98, Nad96, Rag94a, Rot97, Shi93, SGP96, W^{+}98b, Wes99]. really [Mas97]. Realtime [DL95c, MYN96, SMM95a]. reasoners [BB93]. Reasoning [Shi98, ABB99, CDL97, SC98]. recipe [Fly93, SP96]. Recipes [Ano99c, Jel98]. Recognition [IEE93b, IEE95f, IEE97f, IEE98a, MB97a, CB95a, Con93, FX93, LZ98]. Reconstruction [Nic97]. Record [Ano88e, IEE95a, MOTA95, Pit97, GSE^{+}94, PM99, Wae97]. Records [BSF98, PGRC97, RG97b, Ano97y, BB99a, ECvR93]. Red [Fer94, Kro99]. Redefined [DK97]. redefinitions [GMR95]. Reduce [Mue95]. reduction [Ano95-27]. Redundant [LFS97]. réels [BK99]. Reference [AB95, Arm96, CB97, Eij92, Fal97, Fla96, Goo98, Hah91, Hom99, Kis95, Mul96a, Mul97a, RW96b, Sch97a, Spr96b, Wri96, Ano99g, Avi98, Dre98, Eck99, FHU99, Gol89, GP90, Gor91, Gra97a, HUW98a, KCCJ95, LG94a, Man97, Mul97b, Mul98, Pow98a, PW98, Pow99, RR96b, RR97a, RR98c, RMF93, Sch95a, Sch96a, Wil99b]. References [CAD^{+}95, ACD^{+}95, CB95a, PW99]. Referencing [Ebe98, FG95]. reformatting [LS96]. Regency [IEE96k]. Regenstrief [MOTA95]. Registration [Int91a, Int91b, Int88a, Int88b]. Registry [Lev93a]. Regular [BK93a, BKW92d, BK93c, CC97, BKW92a, BK93d, BKW92b, BK93e]. regulations [DPB91]. Related [EH95b, Smi86b, Smi92a, Smi87c, Smi92b, CDB91b, EH95a, For96b, Ish94b, Ish94a, Smi89]. relation [Bla90]. Relational [LLL^{+}91, Mal98, MP94, BCK^{+}94, DJ96, NS96, VY88, WF89, WF90]. Relations [PBFT97, ABS99, BA90, Deu95]. relationship [CM90, Kaz95]. relationships [DSDT96a, Hil96, DSDT96b]. Releases [Ano96-44]. Relevant [ZY95]. Relieve [Mue95]. Remarks [OSE97, Win97b]. Remote [FR97, IEE94b, Nag99, RMR97, SHW95, Ste95a, Ano97d, Dar95]. Rendering [CRV98, HJS98]. RENDEZVOUS [Rad95]. Reordering [GI96]. Repertory [MCG^{+}95]. Replication [GWD94]. Report [Ano97-41, Boy95, FM99, Kul94, MFSF95, Rob94a, Van95c, Wid94, Sha95a]. Reporting [BS95c, KD98, Sun98a, BS95b, Sam95]. reports [van90a, Int92a, Int92b]. Repositories [FB95b, Nic95, DBCK96, Fra96]. Repository [BDG^{+}95, Boi95, RKL^{+}98, SDS97]. Representation [KHW97, KS98a, MP96, Shr95, Ano95v, Gay95, Kel98, OM96, RTW96]. Representational [Gel95]. Representations [CH95, Boa90]. Representative [Ano96-37]. representatives [Sra96]. Representing [CDL97, HS97b, Soi97, RMF93]. Republic [GSV96, PK94a, Ska98, Ska99]. Required [Men95b]. Requirements [FB95b, Fra96, IEE99b, BMZ95, GPE98]. requires [Rot97]. rescue [Ano96-41]. rescue-middleware [Ano96-41]. Research [ACM96d, Ano93s, Ano94a, CBK98, Col95d, FMBM95, IEE96m, IEE96o, IEE99a, JW96, Law98, MM97, NS98, PSTW97, RG95, Sch90a, Sch90b, Ste96d, Ste96e, TCLV92, TA99, WAO^{+}95, Wid94, Win97a, ABHY96, ABB99, Ano90d, Ano97-27, Bou90, Eng97, Liu97a, Shi93]. Researcher [JOA98]. Researching [Dim96, Ste96a]. Residential [For96c, For99]. resolution [Kri98]. Resolving [Tau96]. Resort [ACM97a, IEE97b, LB98a]. Resource [Blu95b, HW94, MLD^{+}99, MP98b, Wei96b, BP93, HL97, MP98c, MP99, Pes95, Wei95, Wei96a]. Resources [Cro95a, LG95, Reb95, Van95c, Ano95c, EG90, EG91, FG95, GSG98, PH96, VGGB95]. ressources [LO99a]. Restaurant [Coh95]. restructuring [BR96a, CCD^{+}99, LSS96]. resultater [sta89]. Results [CV98, Bou90]. rethinking [W^{+}98a]. retooled [Ano96-32]. Retrieval [Ano93p, Ano94n, Che95, CSM97, DWV95, EJ95, EKS95, FK96, HSCZ99, HH99b, KR95, MP96, O’K95, Rut96a, STM^{+}94, Wil93c, AMC95, ACM96e, Ano94p, BMN94, For96b, Gol89, INIK96, IS98, Kas95a, KTK^{+}95, Kir91, Mac90b, MBHL91, PFH95, Shi97, Wan99, Ano93s, DO96]. retrievals [YHHYK96]. Retrieve [MK96a, CM95, Dun95f]. Retrieving [LN99, Shr95, Mac91]. Return [Ber96c, Ber96b]. Reusable [Ano97k, CP93, CP94]. Reuse [DP98b, Gue98b, NF97, WP95, Lev93b]. revealed [MC97b]. Reverse [CH95, WNC95]. Review [Ano98f, BC90, Coh95, Daw98a, Mat97, Pie97, Rag94c, Tho97, Ano95r, Ano95-44, Eft90, Ind96, NBM93, Rag94b, Rag94d, Rag94e]. Review-Based [Coh95]. Reviewing [RB96]. reviews [HV92, PH92, Pop92]. revisited [GW87, Neu97b]. Revolution [Kar93, WZ95, And97, Duk94, Jan96, Sko96, Whi96f]. Rewriting [CJ95, Cou95]. Rey [ACM94a, ACM98]. RFC [BLC95, BLFF96, FGM^{+}97, Lev95, NM95c, PH97a, Rag96b, Sei96, YNAD97]. Rhetorical [Kir91, Wei93]. rhizomatics [Wal94]. Rhodes [To98]. RI [Ano97z, MR99b]. RIAO [Lic91]. RIDE [Wid94]. RIDE-ADS [Wid94]. RIDT [HAB98]. right [Ano95c, Ano96j, GPE98, PW96]. Risk [HN96a, KH95b, Min96]. Rita [CMPd91]. Riverside [Gar95]. Road [DP95b, Vul97]. Robert [Bar92, BC90, Aok99]. Robinson [Cam94]. Robot [Spe95]. Robotic [CB95c]. robotics [HT97, IEE97e]. Robots [IEE97e, IEE98b]. Robust [ML97, PR95b]. Rochester [Ano94f]. Role [Dim96, CJ90, Del95b, Kas95a]. roles [Glu96]. ROM [Dod94, Fee98, GB94, Gre93, GG96, Nak94, Neo97, Ros88, Sad97, She88, Via98b, Weg96, Wil97a]. Rome [Mos96]. Romp [Fis95]. Ronald [Pri96b]. Room [Nie99, RB95a, Ude95b]. Rooms [Reg98, Ano99b]. ROOMVENT [Ano99b]. Roots [Qui95]. Rosemont [IEE96a]. Rouge [NP96]. Roundup [BVN95b]. Routes [Tan98, WNB97]. routines [EM99]. Royal [PKM99]. Rscan [Sam95]. RSJ [IEE97e, IEE98b]. rul [OuP97]. rule [AOS91, GL99]. rule-based [GL99]. rule-controlled [AOS91]. Rules [DiN95, BDKK96, WF89, WF90]. Russia [Mel99]. rzeczywistosci [CK98].

S [Ano98f, ID98]. Safety [Min96]. said [Web95b]. Saint [MNW98]. Saint-Malo [MNW98]. Salt [ACM97a, IEE97b, Isk96]. Sam [Eva98b, OH97b]. Sampler [Rak95, EU95]. SAMS [NH99, Oli98b, Oli99]. San [Ano94a, Ano97-31, Ano98a, Bea95b, Bee97b, BE99, BFM96, BUP99, CEG96, CC99, FJV96, IEE95b, KJR98, LB98a, LZ98, MWEK97, NM95b, SSK97, Ten99, USE89, USE96, USE98, W^{+}98b, Whi96f]. Santa [Ano97x, IEE96e, IEE98a, Sol88]. Santiago [IEE97c]. Santorini [IEE97a]. SARA [Ano94n]. Saratoga [Ano96k]. Sarda [RV94]. SAS [Ano96-36, Ano97-31, Ano98t, BMDB98, BS95a, Gra96d, Hoy96, LS98, Pop97, Spi96a, Sun98b]. SAS/ [Spi96a]. SATAN [Dot95]. Satellite [AS94b]. Savannah [ACM95b]. Savvy [Kro99]. Say [SK98, Ano96-42]. Saying [CK97]. Scalability [AYE^{+}96, Eng94]. Scale [KH95b, MCG^{+}95, Yok93]. Scaling [BA97, RL95a, Wil94a]. SCAMC [Cim96, Mas97]. scattering [MP95a]. Scenarios [Kom93b]. Scene [LCLL98, War98]. Scenes [RMR97]. Schema [AS94b, Kes95, Rod92]. Schema-Based [Kes95]. Schemas [Fin95b]. Scheme [Nør99]. Schiller [Neu94]. Scholarly [Bos99, ML97, Oke93, WMGL95a, Rob96, WMGL95b, Wei95]. scholars [MBM93]. School [ACM97b, BHK^{+}95a, CJ95, Nad90, TD95, Van95b]. Science [Ano98d, AA97, BDG^{+}95, Blu95b, Cha95b, Che97, CJ95, EAG^{+}95, FJ92, IEE95a, JRK95, Joy98, Kee95, Kle96, Mel99, MWEK97, Reb95, Sav95, SS95a, Ste95a, Van95c, Whi96f, Win97b, FA94, W^{+}98b, ZJEP95, Int91c]. Science/Math/History [Blu95b]. Sciences [Ano93a, Ano95h, CBK98, EH95b, JW96, PKM99, EH95a, ERPNS97a, ERPNS97b, MSIM95]. Scientific [Dés86, Gv92, JJ95b, JJ95c, Nad90, Neg95, Owe95, RG95, Smi86a, Bou90, GSV96, HE94, Kir91, Pop91, SMC^{+}96, St.99a]. scientifique [Int91d]. Scientist [GF93]. scientometric [GL93]. SCiP [Ano96c]. Scotland [Kum96, Rut96a]. script [Keo97, Ano96-30, T^{+}xx, Gol96]. Scripting [BKT96, RFM95, Bot98, GJ96]. Scroll [USO97]. sculpting [And96c]. SD-7 [New90]. SDE [BBB88]. SDIF [Int88f, Int88g, Can89, Int86a, Int88c, Int88d]. SDK [Kro99]. SDOC [GL91, GL93]. SDRC [Ano96-29]. SE [Rod92]. sea [Aza94]. Search [DP95a, PR98c, PR98b, WT95, RS96]. searchable [MS95]. Searches [ZKM95]. Searching [CC97, DP95a, PFH95, Ano99l, Eft90, LHKC96]. Seattle [Ano96b, Ano97a, Ano97v, HF97]. Second [ACM99b, Ano93p, Ano95-34, Ano98f, BFY93, Dés86, IEE93b, IEE98c, SSH87, Sch96b, WNC95, AR97, Bac96, DE96, NS98, NDM98, Oke93, SB97, V^{+}96]. Secret [DiN95, Ens97]. Secretariat [Int92a, Int92b]. secretos [HC97]. Secrets [TGHE96, Hei97, HiKgK98, PW96, HC97]. Section [New90]. Secure [Mag94, NJDF96, WR94]. Security [Ano96-33, HBNL95, LJ96]. Seeing [Lut98]. Seek [Law98]. SeeMe [Ano96l, Ano96r]. seen [Web95b]. Seeon [ABB99]. SEI [Ibr95]. Sekrety [HLB98]. Selected [BKS96, Cai94b, Cha96b, Kim97a, ABHY96, Bea95b, BT97, GSG98]. Selection [Kum96, Sea96b]. Selective [JJ95a]. Self [DCG^{+}95, DP95b, OSE97, Ten96]. Self-Paced [Ten96]. Self-Promotion [DP95b]. Self-Study [DCG^{+}95, OSE97]. Sell [Ano96-30]. Semantic [Che95, Deb97, WT95, DBRB96, MBRB96, RTW96, WB96]. Semantics [Ada92, MYBP98, WF89]. semi [YHHYK96]. semi-structured [YHHYK96]. Seminar [Ano94a, Ano95-35, Cze98, MDK98]. seminars [BB94, EG90, EG91, Ano93q]. semistructured [ABS99]. sense [BCG99]. sensible [Lig95]. Sensing [FR97, IEE94b, Ste95a, Ano97d]. Seoul [Ano96n, IEE96d]. September [ACM94b, AHB98, Ano95v, Ano96k, Ano96n, Ano97-48, Ano98m, Bou97b, BH95b, BFH98, BBKF95, C^{+}98, Dan98a, Dob98, Dud97, EAZ98, IEE96k, IEE97e, Jac98, Len93, MG94c, Neu94, NW97, NS98, RT95, Rut96a, SC98, Tea97b, Tea98b, To98, USE95, USE98, Van95b, V^{+}96, Nad90]. sequence [COVS95, Gra95b]. Sequences [KS98a, KS98b]. SEQUOIA [AS94b]. Serial [Ano97w]. serials [BCG99, CL97]. Series [AS93b, Ass87a, ACM96a, IEE98d]. serious [Ric92]. Server [Ano94c, Bou95a, CN98, CPW95, Gla95, GWD94, HKH95, KPIE99, Kro99, Kru96, Lot95, MM96, SVOM95, SS95c, SOUH97, TD95, UL98, WHC95, BC95, CH98, CB95b, Dar95, LDW^{+}96, RR96c, MFSF95]. Server-based [KPIE99]. Server-Side [Bou95a, Kru96]. Servers [Bea95a, FB95a, Car95b, Fat97]. Service [GNSP95, LRG96, EG90, EG91, Hut96, IEE96f, Lew91, Pet90, RS93, Sra96]. Services [ACM95c, Ano95x, BB99b, Can95, Coh95, Dra95b, LZ95, Rid95, TRMW99, WCM99, Ano96n, CHR^{+}96a, Dra95a, EAU96, GPE98, HS98, LDWN96, MM97, M^{+}97, MDK98, Put94]. Servidores [Gar97b]. Serving [AKM95, CLK96, Cro95a, Fis96a]. ses [Fon93]. Session [Ano97-42, Deu95]. Set [WF90, Pao96, Tho93, WF89]. Set-oriented [WF90, WF89]. Sets [Ple95, CNH99, Int91c, MP94, Val96]. Setting [BIL95, RBBW95, Ano96j]. seventeenth [IEE98d]. Seventh [ACM96b, Ano96d, Ano98s, IEE92, IEE96g, Kle96, Mur92b, RPWW96, SSK97]. Several [Wat95]. Seybold [Ano93q]. SFgate [PFH95]. SGI [Par95]. SGML [Ano93n, Ano94n, HV92, Hei94a, Int88f, Int88g, Int91a, Int91b, Int91c, Int92a, Int92b, Joh95b, NBM93, Rat95, Sch90a, Szi95, Van91b, Ass86a, Ass86b, Ass87a, Ass87b, Ass87c, Ass89, Ame85, AC93, Ame91, AP 96, Ada95, Adl92, Sys97, Ahe93a, Ahe93b, Aho95, Aho97, AMN97, AS94a, Als95a, Als95b, Als97, AVFC98, AT88, And96a, Ano90c, Ano90f, Ano92b, Ano92c, Ano92e, Ano92f, Ano93l, Ano93r, Ano93q, Ano93v, Ano93u, Ano94c, Ano94d, Ano94j, Ano94k, Ano94p, Ano94o, Ano94q, Ano94u, Ano94v, Ano95-36, Ano95-35, Ano96-38, Ano97c, Ano97e, Ano97i, Ano97g, Ano97j, Ano97f, Ano97k, Ano97l, Ano97p, Ano97q, Ano97s, Ano97r, Ano97t, Ano97u, Ano97-34, Ano97-36, Ano97-33, Ano97-37, Ano97-38, Ano97-32, Ano97-35, Ano97-45, Ano97-40, Ano97-44, Ano97-39, Ano97-41, Ano97-49, Ano97-46]. SGML [Ano97-52, Ano97-53, Ano97-55, Ano99c, Ano99j, Aza94, BHK^{+}88, BBG^{+}95a, BBSM96, BS95a, Bar89, Bar92, BEE^{+}94a, BEE96a, BG96, Bec93, Bed95, Bee96c, Bir96b, BCK^{+}94, Boa90, BP93, BH95a, BBB88, Bra92, Bra96b, Bra96a, Bre96, BK93a, BKW92d, BK93d, BKW93, BKW97, BK93e, Bry88, Bry97, BR96b, Bur95b, Bur95a, Bur99a, Bus97, Bus88, Can89, Can90, Cad87, CDL97, Cam94, Car94, Cav94, Cha96a, CG86, CG87, CP94, Chr95, CGW96, CJD^{+}96, CK95, CW97, Con96, CP96, CDB91b, CDB91a, Cov96, CGLV94, CH95, D’H95b, sta89, Dav93, Daw98b, De 94, DeR97, DeV96, Dim96, Dob91, Dod94, DABM97, DRBS98, Don97, Dor95a, Dor95b, Dra96, DuC97, DuC98, Ebe98, EG90, EG91, Ell94, Ens93, Ev95, ES99b, FM99, Faw89, Fee88a, Fee88b, FMFVN98]. SGML [Fer93a, Fer93b, Fly93, Fly98, Fon93, For96a, For96b, FWDV93, FB95b, Fra96, FK96, Fre95, Fuc96b, Gra83, Hon95, Gen97, GC95, Gil93, GK93, Glu96, GR90, Gol94, GPE98, Gol97, GP90, Gor91, Gor95, GHJ95, Gro93, Gro97, Haa93, Haj93, Ham97a, Ham97c, HSCZ99, HY96, Hay93, Hay94, HW88, Heb96, Hei94b, Hen95, Hic89, Hic94, Hla96, Hoc93b, Hol87, HL95b, HL96, Hui96, Hun94, vH87, HLL93, Int86c, Int86d, Int86b, Int86a, Int87, Int88h, Int88c, Int88d, Int88a, Int88i, Int88b, Int88e, Int91d, Ish94b, Ish94a, IN96, Jel98, JJ95b, JJ95c, Kae88, Kae90, Kan93a, Kar93, Kay89, KHW97, Kel98, KW97a, KW97b, Kil99, Kru98, Kru90a, LaT89, Lan91, Lap99, NB91, LT86, LHKC96, LS96, LR95, Lev93a, Lev95, Lev96b, Lew91, LQ97, YF96, Lig94, Lob99, Lun95a]. SGML [Lun95b, Lun95c, Luo97, Mac93, MBHL91, Mac96b, MSW97, Mag94, MA96, MB90, MB91, MNA91a, MNA91b, Mat97, Mat93, MGH93, MGH97, Mat99, MYBP98, McG91, McG98a, McL95, MW92, MB97a, MG94b, MMRS98, MK96a, Nat88, Nak94, NTI91, Neg95, Neu97b, New95a, New95b, NS95a, NW92, Nic95, Nor96, OCL88, O’C92b, O’C92c, O’C97, OISEM97, OSYM97, OL97, OnW95, OR96, Ost94b, Pal98, Pao96, Per96, PGRC97, Pit97, PB88, Pop93a, Pop93b, Pop94a, Pop91, PHL97, Pre96, PW94, Pri87, PS88, Pri96a, Rad98, RARH99, RW96a, RTW96, RM95a, Rie94, RG97b, RW90, RW92, Rob94b, Rob94a, Rob94c, Roc93b, Roe87, Rol91, RM97a, Ros88, Rot97, Rub92, RM96b, RM97b, RKL^{+}98, SD96a, SNLT96, STA98a, San94, Sch90a, SF90, Sch90b, Sch96b, Sch92]. SGML [Sch95d, Sha97, Shr95, Smi86a, Smi87a, Smi87d, Smi87f, Smi87b, SS88, Smi92a, Smi96a, Smi96b, Smi97b, Smi98, Smi87e, Smi92b, SS97, Sof91a, Sof91b, SM93b, SMG95a, SMG95b, SMG96, SMH99, Spi96b, SD96b, Ste95c, Ste96a, Sti94, STM^{+}94, SHO^{+}96, Sum94, SIF^{+}96, Szi95, Tan91, Tau96, Tho93, TKB90, TW95, TTS97, TDT95, Uns96, Vau94, VB95a, Vig89, Vin99, Vol97, Von97, Vul97, WZ95, Wal92, Wal95, WV89, WV91, War87, WD90, WD91, WS95a, Wei94, Wei93, WE92, Wie94, WC92, WB96, Wil93c, Wit96a, WP98, Won92, Woo95c, Woo95b, Wu89, WR94, YYOO97, YHHYK96, YIU96, Zha95, ZR93d, Zif96, d’H95a, van90b, van93b, van94, vES95, vCL89, van90a, van91a, McG92, Pop92, Pop94b, BC90]. SGML-based [AT88, BHK^{+}88, BG96, MW92, MG94b, MSW97]. SGML-encoding [GP90]. SGML-L^{A}TEX [van90a].

SGML-marked [Hen95]. SGML-ODA [NB91]. SGML/ [Ano97c]. SGML/HTML [ES99b, Gro97, Per96]. SGML/HYTIME [OISEM97, FB95b, New95a, New95b, VB95a]. SGML/XML [Ano94q, Smi98]. sgml2tdb [Bra96a]. SGML — [WM92]. SGMLtools [de 98]. SGQL [AMFL^{+}95]. Shake [Ano96-29]. shang [HUW98a]. Shape [KFHU98]. Shaping [HC99, CBK98]. Share [SP99b, Ano94m]. Shared [Bro97, Bro98a, HBNV97, SOUH97, HMRL95, PY98, KGK94]. Sharing [BHK^{+}95a, KFHU98, Mar95a, RC94, Tes96a, CBK98, HW94, MM97, OST94a]. Sheet [MM99a]. Sheets [BKW92c, Fal97, LB96, LB97, Zac97, Bou98, PMTJ98, PTJ98, War87, BO98, JT97, JT98, LB99a]. Shelf [Ano97c]. SHEN [HBNL95]. Shenzhen [LLY97]. Sheraton [Ano95-37, Ano97b, Bos97, Cim96, DG97]. Shift [SF96, Chu98]. Shockwave [Kar96]. Shoemaker [Tow95]. Shoemaker-Levy [Tow95]. Shop [Ano96-29]. shoppers [Ric96]. Shoreham [Ano95k]. Short [Cea96a, Hol99a, Lin99]. shou [HUW98a]. Should [Kis95]. Show [Sha95a, Vul97, LM95a, Ret96]. SHyD [ZR93b]. SICE [Ano97-42]. Side [Bou95a, Kru96, Sco96a, Sei95, Sei96, Dob97a, HL97]. sieciowe [Dan98a]. sieciowej [CK98]. SIGACT [ACM94c]. SIGACT-SIGMOD-SIGART [ACM94c]. sigan [OuP97]. SIGAPP [DGBH93]. SIGART [ACM94c]. SIGCSE [Joy98, Kle96, MWEK97]. SIGDOC [ACM92, ACM95b, ACM96d, ACM97a]. SIGDOC’93 [ACM93a]. SIGGRAPH [ACM96a, Sha95a]. SIGIR [Ano93s]. SIGMOD [ACM94c, JM96]. Signal [IEE97a, To98, W^{+}97]. Signature [Mor95c]. signposts [Wal95]. SIGUCCS [ACM95c]. silence [Mac96a]. Silicon [Ano95-27, GK93]. SIMON [FW93]. Simple [And99, BS90, Law98, PR95b, PK95b, Ano97t, Lew95, MM99b, Ano97-32]. Simplification [GLTH98, War98]. simplifying [Yee99]. Simulation [BUP99, DG97, Fab95, Far96b, LB98a, SVOM95, Wu98, Gau96a]. Simulations [Mel99, SS95b, CT96]. Simulator [CCEE98]. simultaneously [Hen95, JGKR96]. Singapore [CPK95, PS95]. Single [NS95a, Ano95-29]. Single-Source [NS95a]. SIPs [Mat97]. Site [BL98, Gel95, Pau95, Ano95c, Ano95b, Ano96j, Ano96-45, BBKL94, Bal96b, BC96, Dan98b, LSI98, PW96, VW96, RPWW96]. Sites [Ano95b, Tho97, Ano93d, Ano96-45, Blu95a, Dob97a, Fis97, Hen96a, Mor96, O’C97]. Six [Pfa97b, Pfa98b, Pfa98c, Pfa99]. Sixteenth [Ano93s]. Sixth [ACM97b, Ano97x, Coy97, IEE96n, IEE96o, Ska98, IEE96c, USE98]. size [Fic98b]. Skalský [Nad90]. Skills [BL98, BE97]. SkiSoft [Ano96w]. Skripte [ER97]. sky [Ano98r]. Slavic [Rad98]. SlickEdit [Ano97w]. slide [Tho93]. slightly [Ros96]. Slow [Duk94]. Small [Ano97-43, Mei97, MKEB98, RM97b]. smaller [Sch95b]. Smalltalk [BR96b, LP96, Lut98]. Smith [Bar92, BC90]. SMP [Fer94]. SNAFU [Hay94]. Snapshot [SK98]. Snowbird [ACM97a, IEE97b]. sobre [Ste98]. Sociable [DR95c]. Social [SSH87, Sav95, Ish95]. Society [Ano94r, Ano95-37, Ano96c, Ano96s, Ano96b, Ano96-39, Ano98b, Ano98d, Ano98o, Ano98r, Ano98u, HMRL95, IEE92, IEE96e, IEE98a, PE95, PK94b, RPWW96, Tea97b, Tea98b, Whi96f, FM99]. Sociological [She95]. Sofia [Ano97-37, Dob98, ASWS98]. SOFIA/HAWC [ASWS98]. Sofitel [RM95b]. Software [Ano94s, Ano95-38, Ano96-27, Ano96-29, BVN95b, Boi95, BA97, Buc97, Cou97, CH95, DP98b, Fab95, FZ99, FW96, Gue98b, Hie97, IEE96d, IEE96h, IEE97g, IEE98c, Ibr95, Kro99, LH96, Lut98, MKNS87, McG98a, MG94c, Uns96, WP95, You96, ZMP96, AHB98, Ano96j, Ano99g, EG90, EG91, Gol96, Hoy96, Ind96, Liu98, Ric92, Ros88, Rud96, Spi96a, TC96, VW96, Wil96b, Kro99]. software- [Gol96]. Soho [Ano95-27]. Solid [Ano96-29, Bou97b, WDM97]. Solution [Fic98a, RG97b, Ano95-29, Hen91]. Solutions [ACM95b, BGG^{+}94, Ano97d, BHK^{+}88, BBG^{+}95a, Ano95-39]. Solve [Hin98]. Solver [KFZ96, PMK98, ZFK97]. Some [MMV95, NN92, PCL^{+}95, Var95, BHK^{+}88, Ubo96]. Sonthofen [AHB98]. Sony [Ano96-27]. Sophisticated [How96]. Sound [Ell98]. Source [LSW95, NS95a, Ret96, EG90, EG91, SM93a]. Sourcebook [GR96, Gra97b, ANM96, ANM97, Gra95c, Gra96c, Gra97c, Gra98, Wil97a]. SourceCraft [Wag97]. Sources [Cam94, Rob94c, VN95]. Southeastern [Gau96a, Mei97, NP96]. SP [LDWN96, MBB90]. Space [EH95b, MB95b, AP96, EH95a, Flo96, Yok93]. space-technologies [Yok93]. SpaceFusion [SOUH97]. Spaces [KW97c]. Spain [Ano97-37, VD98]. Spanish [HC97]. Spatial [HLS97]. Spatio [HM97]. Spatio-Temporal [HM97]. Spatiotemporal [VT99]. Speaking [GB95c]. Spec [Con94, Ano96-42, Ano96-44]. Special [Bow97, BJ96, CJD^{+}96, GR97, GR99b, Hol99b, KAKZ96, KS96, SWH95, SBJ96, Liu98, WT93]. Special/Anwenderpraxis [KS96]. Specialist [BL98, Tea97b, Ano90d]. Specialists [Tea98b]. specialized [RR96c]. Species [BS95c, BS95b]. Specification [Ada92, Deb97, Erf93, Mon98, PBFT97, SAA^{+}94, VRM97, BH95b, BFH98, DuC99, GQ99, WKD96, RC95]. specification-based [WKD96]. specifications [Ano97o]. Specs [Cla98]. Spectroscopy [Ano90e]. Spectrum [VGGB95]. specyfikacji [CK98]. Speed [BG95b, WNB97, BG95a]. speeds [Mat98]. Spezifikation [Rat95]. Spider [Ano95-44, Ind96]. spills [Ano95n]. Spin [Ano96-32]. Spinning [Fis96a, For95a, GB95b, Pow95]. spoken [Joh95a]. sponsored [Neu96]. Sport [Fic98a]. Sprachen [Die96]. Spring [Ano94m, Ano95-35, Ano98c, CJ95, IEE92]. Springs [Ano96k]. SQL [BA90, BBG^{+}95b, BCK^{+}94, CK97, Cel92, CN98, Dav92, Deu95, Eis96, EM99, Eng94, EGH95, Fin95a, Fra91, GW87, GI96, Gog90, JMB95, Kle94, LFS97, LCW93, MM95a, New92, Pri96c, RP98, Ric95c, Rod92, Toy98, VP95, VY88, vB89]. SQL-based [BA90]. SQL/CLI [VP95]. SQL/MP [EGH95]. SQL/SE [Rod92]. SQL3 [Kul94, MD94]. SQL 2 HTML [Fin95b]. SQLJ [EM99]. SSL [LC98]. SSQL [YHHYK96]. St [ACM95c, HAB98, Lit96, Mel99, RM95b]. ST.32 [Bre96]. stabilization [Vac96b]. STACS [FJ92]. staff [Ten95]. stand [Sch92]. Standard [Ass86a, Ass86b, Ass87b, Ass87c, Ass89, Ame85, AC93, Ame91, Ano93o, Ano97f, Ano97-40, Can89, CDB91b, Fee88a, Fee88b, Gra83, Ham97a, Hay93, Hei94a, HMHK97, Int86b, Int88e, LaT89, MMRS98, Org91, Pit95, RB95b, SS95a, Smi89, SM93b, Ass87a, Ada92, Adl92, AT88, Ano93r, Ano97-29, Ano97-28, Bre96, CH90, CG87, Dor95a, Dor95b, Eis96, EG90, EG91, Gol91, MM95a, Pri93, Rei93, Rol91, Smi87a, Smi87d, Smi87f, Smi87b, Stu87, Via97c, WSM96, ZR93c, Ano93d, Ano94t, Bry88, Bug98, CG86, CDB91a, Hei94b, Int86c, Int86d, Kay89, Kis95, Kru90a, Nat88, NBM93, Smi86b, Smi87c, Smi87e, TTS97, Woo95a]. standard-based [ZR93c]. standardiseringens [sta89]. Standardization [BEE95, Dou92b, MBB90, Tan91, YYOO97, Ano95-43, Chu92, Ev95, KAH91, Kom91, Kom93a, Kom94, MK88]. Standardized [Sea95, CNH99].

Standardizing [Bir96b, KC92, WS95b]. Standards [ABY94, Ano91e, Ano92g, Ano93x, Ano97-44, App89, BB91, Bro89, Bry93, Car97, Cro92, Del95a, Fer93a, Gar99, Gol91, Hoc93b, Int92a, Int92b, Koc94, Law98, MR99a, MM99a, McL95, New95a, New95b, RC94, RFM95, Sch88, Smi92a, Smi92b, VTW94, Whi91, AW93, Cav94, sta89, For96b, GB95c, LG94b, Mar91, Mat98, MM90b, RB96, Sch92, SR93, Smi89, Spr96b, Val96, Van93a, Van89, Wei95, WZ91, RH97]. Standards-Based [Fer93a]. standards-CAPS [AW93]. Standards-History [MR99a]. standards-hypermedia [Mar91]. Stanford [IEE96j]. Star [Str95, H^{+}96]. Starburst [LLL^{+}91]. Start [Pal98]. Starter [Web95a, HB97a]. Starting [BBKL94]. Starwave [Ubo96]. State [Ano96b, Can95, IEE96n, IDK98, Kam90a, Pro95, BB91, Bou95b, MM90b, Tub95]. State-of-the-art [IDK98, Kam90a]. States [DGBH93]. Statistics [KCFP95]. Status [Kul94, Lev93a, BB90]. Stay [AS94a]. STEALTH [PAV96]. Step [Neu98, Pes96b, Win96, Ano97e, Ano97-44, BP93, KKK97, Sha97]. STEP/EXPRESS [Sha97]. STEP/SGML [Ano97-44]. Steps [HL95b, HL96, Luo97, Pfa97b, SBGO99, Pfa98b, Pfa98c, Pfa99, RM97b, Rus99]. Stereoscopic [BFM96]. steroid [Wu98]. still [Ano97-40]. stock [Kum96]. Stockholm [Ano99b, GJTW97]. Stops [Fis95]. Storage [IEE95e, Mac90b, VT98, Koc94, Roe87]. Stored [Ric95c, Eis96, Ges91]. Storing [NK99, Cro92]. Story [GK93, Gra95d]. stranek [ID98, Kos98]. strange [Deu95]. Strasbourg [Dés86]. Strategies [Ano97-34, Hou98, OR96, Kru90b]. Strategy [Hay94, ECvR93, LJ93, RR94]. Strathclyde [Kum96]. streaming [Rus98]. Streamline [Dod94]. streamlining [MNA91a, MNA91b]. strength [Don97]. strengths [Pri96a]. Stress [SSH87]. strikes [LLL^{+}91]. Stripper [Bjo96a, Bjo96c]. strong [Ubo96]. stronger [Wat96]. structural [CB95a, Kum96]. Structure [CSM97, Heb96, JMB95, LR95, MB97a, MB95b, Pit97, STA98a, SNS95, SN96, FX93, GF93, Hei95, Kik96, Kir91, LSI98, Rol91, RMF93, Shi97, TKB90]. Structure-Based [SN96, GF93]. Structured [Als95a, Als95b, Als97, ABO98a, ABO98b, Ano99c, BMN94, CACS94, Fly98, FMSDW93, HSCZ99, Jel98, KD98, LSI98, OSYM97, SD96a, Won90, AMFL^{+}95, Bro89, Cad87, CR94, CM95, CMPd91, CL90, DSDT96a, ECvR93, FW93, Gol91, KZSC96, LZSD95, Mac90b, MN95a, NR96c, PFH95, QRV95, SDAMZ95, Sta98b, DSDT96b, WP98, Yam93, Yam94, YHHYK96]. structurée [Vig89]. Structures [LN98, LN99, Lut98, Shr95, BAC93, Buc93, FWDV93, GL91, Mac91, PH96, New95a]. Structuring [BEE^{+}94b, CGW96, Int97, Lap99, Meg98, FMFVN98, NKN91a, NKN91b, Daw98a]. strukturierten [Rat95]. Student [Fon93]. Students [Che97, WD95, Chi95]. Studies [SMDO98, GL93, Kan93b]. studio [BE97]. Study [Ano97c, Ano97f, Boa95, Cam98, Can95, DCG^{+}95, Dra95b, FK96, GK95c, HN96a, HY96, Mue95, OSE97, Per95, ZMP96, BIL95, Dra95a, EGH95, GB95a, Kah96, MP95a, Ric96, Shi97, Xu97, vSSW93]. Studying [MGH93, Sed95]. Stutely [BC90, Bar92]. Stuttgart [Vez96]. Style [Ada92, BKW92c, BO98, Fal97, JT97, JT98, LB96, LB97, LB99a, MM99a, Wol91, Zac97, Zif96, Aro94, Aro96, AL97, Bou98, PMTJ98, PTJ98, VP95, Wor97, War87, van90a]. Styles [SS97, Ano95r, CHP88, Gla99]. stylesheet [Gra97c]. Stylesheets [RR98a]. styly [ID98]. Subdivision [KKK97]. submitting [Van89]. suboptimal [SM93a]. substance [Wor97]. Substantive [DP95b].

Success [GK93, PW96]. such [PH97a].

SUGI [Ano96-36, Gra96d]. Suite [Kro99, Vol97]. Suited [USO97]. Summaries [Ano88d]. summary [Cav94]. Summer [Ano91f, Nad90, USE87]. Sun [Kro99, NS95c, NS95b, Par95, Ubo96]. Sundsted [Ano98f]. sung [mWfC97]. SunWeb [NS95c, NS95b]. Superhighway [Ano95k, IEE96e, RC96, Bar97, Bar98b, CPK95, Cim96, Del95b]. SuperJANET [Pul95]. SuperJournal [Pul95]. Supersedes [Ros97]. SuperSQL [Toy98]. Suppliers [Ano95-38]. Support [Ano93h, Ano95f, Bal98a, BE95, DHR97, EH97, Eng94, Gar99, GN95, IEE94a, Int88c, Int88d, KD98, Lot95, Mol95a, RL95a, Sei95, Ano96y, BHZ96, Ber96a, Bot93, Can89, Can90, CP93, Fuc96a, HMRL95, Int87, Int88f, Int88g, Int88h, Int88a, Int88i, Int88b, Int91a, Int91b, Int91c, Int91d, Int92a, Int92b, KMA^{+}96, KS95b, KGK94, MM95b, OSEM^{+}96, Rag94a, TF95, Ten95, WR94]. supported [MR93b]. Supporting [Ano97w, Ano97-45, Bro97, MKNS87, VW96, YF96, Rod92]. Supports [SHW95]. Sure [Kle94]. Surface [GLTH98, IEE94b]. Surfaces [GLTH98]. Survey [HN96a, Kan93b, LCW93, Fra99, HG96a, Hen96a, HG96b, Loe94, Sum89]. Surveys [Smi96a, Liu98]. Sustainable [MDK98]. Sweden [GJTW97, LR94]. sweeping [Car95b]. Swim [KH95b]. Switzerland [Cai94b, Dra95b, Dra95a]. Sybase [GWD94]. Sydney [K^{+}98]. Symantec [Ano97-51]. Symbolic [Küc97, Fat97]. Symposia [Pao96]. Symposium [ACM94c, ACM99b, A^{+}98, Ano90e, Ano93p, Ano98l, BBJR96, CEG96, Cim96, DGBH93, FJ92, Fon93, FW96, Gar95, HW94, HW96b, HK94, HK95b, IEE94b, IEE95e, IEE96n, IEE99b, IDK98, Joy98, Kle96, Küc97, Len93, LG94b, MZZ96, Mas97, MWEK97, Mos96, MB97b, NM95b, NP96, SB97, SR98, SB99, Ste95a, Ste96e, USE97b, WS95b, dN97, Ano95v, Ano97-27, Ano97-43, Ano99d, BP96, Bor96, Bou97b, Chu98, HT97, Oke93, Oke95, PKM99, RM96a, RS97, Ste96d]. Synchronization [Rad95, WKD96]. Synchronous [FLF95b, RW95a, FLF95a]. syntactic [Con93]. Syntax [DK97, WF89]. synthesis [COW98, Gor95]. Synthetic [MGVS97, GSE^{+}94]. Syracuse [IEE96c]. Sys [Bun96]. System [Ano97w, BA97, DGG^{+}92, DK97, DR95d, ERPNS97a, ERPNS97b, FH97, FNN97, HLS97, HSCZ99, HB95, HLR97, HH99b, IEE97e, Ito98, JRK95, Kee95, KK98, Kro99, LHL97, MBH95a, Mue95, MP94, Nat99b, NP96, Neg95, Nic97, OISEM97, Pop97, Rad95, RW95b, RL95a, RKL^{+}98, SD96a, SHW95, SWC^{+}95, SBD^{+}95, Shi98, SBS95b, Ste95c, Sun98a, TBS97, TABA^{+}94, Zha95, Ano95l, Ano97l, Ano97-54, Ano98r, AOS91, Bai93, BBB88, BDH^{+}95, DJ96, DMM90, EH96, FW93, Har94, HIYO96, Hay93, INIK96, IN96, KK94, KTK^{+}95, KZSC96, KSC96, KS98c, KSM^{+}98, Lau94, LS96, YF96, MSW97, MG94a, MBH95b, MYN96, NANK96, OST94a, OSEMV95, PY98, Pow94, RRS^{+}93, SF90, Smi93, TF95, VB95b, Wan99, Wha92, Wol91, YT96, d’H95a, MSMY97, Bar95, DGG^{+}92, EAG^{+}95, LG94a, MOTA95, MN95b]. System [NM95a]. system-Collaborwriter [MG94a]. system-technology [OST94a]. System/9000 [DGG^{+}92]. Systematic [SRB96, BMY95]. Systemen [GS87]. systemes [Int86b, Int88e]. systemics [COW98]. Systems [ACM94c, ACM95b, AS94a, Ano90a, Ano91c, Ano93i, Ano96a, Ano96-35, Ano97-38, Ano97-45, Ano98o, Ano99a, AA97, BSF98, Bla94, BKS96, Cze98, DWV95, DS96, Ear87, GS87, HMHK97, IEE94a, IEE95e, IEE95d, IEE95g, IEE96b, IEE96k, IEE96l, IEE96i, IEE97c, IEE98b, Int86b, Int86a, IR94, Kho94, KMS99, LK95, LM95b, MKEB98, MP96, MS95, OMY97, QSBC98, RB95b, SS89, Shr95, Sol88, Ste95b, TB96a, Thr95, TR92, TW98, USE95, USE97b, VC96, Wid94, WHC95, Ame85, AC93, Ame91, AT95, AHB98, Ano94f, BCK^{+}94, BFM96, BDK95, Bot92, Can89, COW98, C^{+}98, CL99, Chu98, DPB91, Dud97, ERPNS97a, EG90, EG91, For96b, Gup97, HE94, HB98, IEE96o, Int86c, Int86d, Int88e, KIM^{+}97b, Kru90b, LHKC96, Lev93b, Lit96, LDWN96, Loe94, LG94b, Mas97, Mor95a, PKM99, RM96a, RS97, RT95]. systems [SDAMZ95, SSK97, Uza92, WF98b, WF89, WF90, WR94, Kro99]. Systems-Multimedia [IEE97c]. systemu [CK98]. systemy [Dan98a, HS97a].

T [Ano95k]. Table

[Ano99c, Lap99, SBGK98, Dun95e]. Tables [Mic97, OA98, Rag96b, JMS98, McG92, Oli97b, Rob98, Tho93]. tabular [Ass86b, Ass87c, Ass89]. Tackling [DiN95]. TACT [Bra96a]. TAG [SGM87, Kay89, Pao96]. Tagged [AMN97]. tagger [Lig94]. Tagging [DeV96, PHL97, Rad98, Shr95, BM91, HW90, LM95a]. Tags [PC97, Dun95f, Pao96]. tail [CLS96]. Tailoring [JCDA95]. Taiwan [CL99]. take [Ano95-46, Dun95e]. takes [PW96]. Taking [DP95b, Reg98, SK98, Kum96]. Talairach [CNH99]. Talk [Roe96, Fuc96a]. Tampere [Ano96-40, HT97, Rii97]. TAO [SZM^{+}95, SMM95a, SMM95b]. Target [Ano96-30, FH97]. Tasks [Bun96]. taste [Lin94]. Tastes [Ano97-46]. Tax [Pro95]. Taxing [Pro95]. TC3 [VD98]. TC3/WG3.3 [VD98]. TC5 [Ano93i]. TC5/WG5.10 [Ano93i]. Tcl [Bal98b, Bal98a, SD98, Ass96, USE97a, USE98, WU96]. Tcl/Tk [Ass96, USE97a, USE98, WU96]. Teach [Ano95-40, Ano95-41, Ano97-47, Bry99, CD97, ES99a, Fox95, Lem95a, Lem95c, Lem95b, Lem96a, Lem96b, Lem96c, LD97a, LD97b, LT98, LT99, MC97a, OH97b, Oli97a, OH97c, Tat96, Via97e, Via97d, Eva98b, NH99, Oli98b, Oli99]. Teacher [RPWW96]. Teaching [BHK^{+}95a, Cha96b, Kea96, Lan96, LD95b, WF98a, JI95, Mar94]. Team [HTL95]. Team-Based [HTL95]. teams [KGK94]. Tech [Kau96, Roe96, Thi91]. Technical [Ano90d, Ano94r, Ano95-37, Ano96b, Ano98u, BBG^{+}95a, BG96, Dav93, Ano93k, Int92a, Int92b, Joy98, Kle96, Kru90b, LaT89, LQ97, MFSF95, MWEK97, She95, TA99, Van95c, WCM99, Wil94a, And88, Ano97a, Ano97o, CP93, CP94, FM99, GF93, Hen91, IEE96a, IEE98d, Rob96, St.99a, USE96, Zim95]. Technique [BR95, Fra91, Tes96a, Gol96]. Techniques [BKT96, Bou95a, Con97, DWV95, HLS97, Int87, Int91c, Int92a, Int92b, LHL97, Mic97, Nad90, RBR95, SS98, War98, Zha95, Can90, HS98, Int88i, Int91d, SWC99, SCW99, SVC99, Int88h]. Technische [Ano97-48]. technischen [Wie94]. Technological [ACM96d]. Technologies [ACM94a, ACM98, Ano96u, CV98, FFKO99, HPS95, IEE94b, IEE96e, IEE96j, IEE96c, Int91d, USE97b, Ano96-32, BKP97, Dud97, HW94, Mor95a, OKT92, Sug96, SIF^{+}96, Tub95, Vau94, Wes99, Yok93]. Technology [ACM94b, ABY94, Ano88f, Ano91b, Ano93u, Ano93t, Ano96-31, Ano96-42, Ano97-49, Ano99h, AA97, C^{+}98, IEE94a, IEE95c, IEE96m, IEE99a, Isk96, Kro99, LMT98, Lot95, Mur92a, Mur92b, NB95, OBO97, Owe95, RPWW96, TCLV92, Uns96, AC93, Ame91, ABHY96, And96c, Ano93d, Ano96s, Ano97h, Ano97-27, Ano97-40, ABG96, BEE96a, CBK98, EAZ98, ERPNS97a, Gol96, Int91a, Int91b, Int91c, Int91d, Int97, Intxx, IDK98, Ish94b, Ish94a, IN96, Kum96, Lie95, MM97, M^{+}97, NS98, OST94a, Owe96, Pes95, Rub92, Rus98, Tit98, W^{+}98b, Whe96]. Technology-Based [Isk96]. Tecnologia [Gar97b]. TeCo3D [Mau98]. TEDI [MR93c]. TEI [Ano96-38, Der99a, Dun95g, Gio95, Haa93, ISM95, MW92, Sea95].

TEI-TERM [MW92]. Tel

[Ano99d, BP96, IEE98e]. Tel-Aviv [IEE98e]. Tele [MKBB95]. Tele-Conferencing [MKBB95]. telecommunication [HKM^{+}96]. Telecommunications [Ano91b, Lav97]. Teleconferencing [Ano94h, Ano97-50, Rus98]. telecooperation [Mau98]. Teleoperated [AVN98]. Teleoperation [HMHK97]. telepathology [DBRB96, MBRB96]. Telepresence [Hue98, Soa98]. Telerobotic [Ano97w]. Telescopes [CB95c]. television [Qay96]. TeleWeb [SDK^{+}96]. Telework [SM94]. Teleworking [Ano94h]. Template [Ano97-41, HY97, Kri98, Sad97]. Template-Based [HY97]. Templates [SD98]. Temporal [Ano97-53, HM97, PBFT97, TABA^{+}94, Erf93]. Ten [Eva96, Zac97, Str96]. Tennessee [Ano95m]. Tenth [Ano98q, Jac98]. Term [CJ95, Shr95, MW92]. Term-Concept [Shr95]. terminographic [Rei93]. terminologie [BKP97, Fon93]. Terminology [Fon93, Sch93, MW92, Mel95, MR93c, WS95b]. Terrain [BW97]. Tertiary [PE95]. Test [Arm96, Ber95, Ber96c, Cov96, Dot95, Kro99, SHW95, Vol97, Ame91, Ber96b]. Testbed [Kee95]. Testing [Ano91b, Mel99, AC93, Ame91, Lew91]. TEX [Ham97c, Won94]. TEX-based [Wit86]. Texas [Ano99a]. TEXnician [Eij92]. Text [Ano88g, BBSM96, Bar94, BCK^{+}94, Boa90, CC97, Cov96, Dob98, Fly98, Gon94, Int86b, Int86a, Intxx, JMB95, JJ95b, JJ95c, LaT89, Lap99, Lic91, Lin94, Loe94, MCP91, MR99b, Neg95, Ore88, Pet95, Sea94, Sea96a, Sun98a, TBS97, Zha95, ZR93d, Ame85, AC93, Ano95-31, Bar98a, BBG^{+}95a, BEE95, Bry93, Can89, DSDT96a, Dra94, EtR94, Gil92, Gol89, Gv92, Int88a, Int88b, Int88e, Int91a, Int91b, INIK96, Ish94b, Ish94a, Kas95a, Koc94, LZSD95, Mac91, MBHL91, MN95a, MA96, MNA91a, MNA91b, Neg92, O’C92b, O’C92c, OKT92, San94, DSDT96b, Van89, ZR93c, ZR93b, Hoc93b, Int86c, Int86d]. Text-Hypertext [ZR93d]. Text/relational [BCK^{+}94]. textbook [Smi97a]. textbooks [GB94]. textes [Ano96t]. Textile [Ano96-40]. Textiles [Lam96]. Texts [AMN97, HY96, Rad98, BHK^{+}88, CR95a, Dob98, GF93, Gay94, Gio95, Hen95, Hoc93a, Joh95a, LM95a, Oli97b, Sea96a, Sea96b]. Textual [Cam94, GB98, Mil93, Pet95, Shi98, Shr95, CR95b, Rob94c]. textured [War98]. th [Ska99]. Thanks [Vac96b]. theater [Dan98b]. their [Ano95r, BVN95a, BVN95b]. them [Ano95-35, Ano96-32]. Theme [PS95]. Theoretical [CJ95, FJ92, Ran98]. theories [To98]. Theory [Ano92h, NP96, Gon94, IEE98b, WS95b, Wu89]. Thesaurus [DeV96]. Thesis [ES99b]. Thing [Ano95-28]. Thinking [Pet95, Ano97-55]. Third [ACM98, ABY94, ABB99, Bea95b, IEE96m, NW97, RM95b, SS89, Sch93, Ano97-27, HS98, IEE95f, MM95a, SR98, VC96]. third-generation [MM95a]. Thirteenth [ACM94c]. Thirty [IEE92]. Thirty-Seventh [IEE92]. those [Ret96]. though [Ano96-42]. Three [Ame97, FB95a, AP96]. three-dimensional [AP96]. Throttle [Ano97r]. TIDE [Eng93]. TIDE-CAPS [Eng93]. tidy [Col95b, Rag98]. tied [Ano92c]. Tif [Mel95]. Time [Tow95, BVN95a, BVN95b, EAU96, MP95a, MMJ96, MWVE99, NJDF96, PM99, SZM^{+}95, SMM95b, Uhl95, Win96, Win97b, Law98]. time-based [NKN91a, NKN91b, Int97]. time-dependent [MP95a]. Times [Pro95]. Timothy [HV92, Pri96b]. Title [IEE98d]. Tk [Ass96, USE97a, USE98, WU96]. TKE [Sch93]. TkWWW [Spe95]. TM [Vin98]. TN [Ano98t, Mas97, Mei97]. Today [Gil93, MM99a, BK96, Mur90, Rob96]. Todd [Ano98f]. Todo [Ste98]. TOGA [SMM95a, SMM95b]. TOGA-TAO [SMM95a, SMM95b]. Together [Ano97-44, BE95, Bra96a]. Tokushima [Ano97-42, Ano99f]. Tokyo [IEE96b, IEE98f, MTH99]. tomorrow [BK96, Wil96b]. Too [Mag97]. Tool [CAD^{+}95, Döm95b, EKS95, Fab95, FH97, Fer93b, Kro99, LH96, LN99, MKBB95, OPW95, Pel97b, RW95a, Sin97, ACM96e, ACD^{+}95, Ano95-29, Bar98a, Döm95a, GL93, KTK^{+}95, Liu98, Ric94, San94, VGGB95, Wag97, LD95b, LSW95]. Toolbox [Sig96]. tooling [Gre98b]. Toolkit [Ada97b, CR96, Daw98b, Far96b, Smi95, CH99b, NR96b, SBS95a]. Tools [Ano95s, Con97, CPR98, Dav93, Fly98, Fra99, FW96, Hla96, IEE97d, Jac95, JW96, KMK95, Kro99, LM98, Mat99, Opp97, SS98, Ven96a, WU96, Ano94k, APH98, BVN95a, BVN95b, CR96, Col95b, Lan91, Mor96, Rud96, Rus98, Vac96b, Web95b, C^{+}98]. toolset [BM91]. Top [Sch95c, Str96, Yuz89]. Topic [Ano97g, Ano97-52, Eij92]. Topical [Ano94a, HT97]. Topics [IEE95d, Nat99a]. tops [Ano96y, Jon95]. Toronto [Ano97b, Ano99i, Bos97, FW96, WNC95, Wil96b]. Tough [Ano96-30]. Touring [CJ94]. tourism [Huc97]. TR [Int88h, Int91c, Int92a, Int92b]. Track [Ano96k, ERPNS97a, ERPNS97b, Goo94]. Tracking [DB96]. Trade [Ano96b, IEE97e, MD94, Pig96]. Trade-offs [MD94]. trading [SP99b]. Traditional [DCG^{+}95]. Traffic [Sed95, BC95, CB95b]. Training [Cas96c, Fer93b, TTG98, SM94, Ten95, VD98, WWN96]. Traitement [Int86b, Int88c, Int88d, Int88i, Int88e]. transaction [BB99a]. transactions [Cha99]. Transcription [Cam94, Rob94c]. Transfer [BLFF96, FGM^{+}97, MKEB98, Nag99, NB95, Ano90d, Yee99]. Transformation [Fre95, Hon95, Pit95, STA98a, FW93, Sum94]. Transit [Ano96w]. Transition [Ano97-39, LL97]. Transitioning [NPPR95]. Translating [Ano92h, Hoe97]. Translation [Ano92h, Gog90, WR99, MBO93, MYN96, Wal90]. Translator [SDSS95, Mar98]. translators [Bus88]. Transmission [GLTH98, MGVS97, NTI91]. transportation [Win96]. trek [Kau96]. Trends [Ano90e, Ano93w, Ano96-42, MK88, And96c, KAH91, Str96, VD98]. Trento [Jac98]. TRI [ACM96c]. TRI-Ada [ACM96c]. Triangle [ACM96d]. triangles [War98]. tricks [Ude95b]. Trigger [LFS97]. truth [Ges91]. Tsen [AM97]. tso [Neu98]. TSQL2 [SAA^{+}94]. tsui [mWfC97]. Tuning [CN98]. Tuple [Gog90]. Turner [Pri96b]. turning [Ens97]. tutor [MP95a]. Tutorial [Als97, Hah91, Ten96, Mee96, Oli97b, Via98b, Wea97, van93b]. tutorials [CT96]. tvorba [Kos98]. Twelfth [IEE98f, Ano97d, Su96]. Twentieth [IEE96d]. Twenty [Ano96-36, Gra96d, IEE97g, IEE98c, Joy98, Kle96, MWEK97, NP96, V^{+}96]. Twenty-Eighth [NP96, MWEK97]. Twenty-First [Ano96-36, Gra96d, IEE97g]. Twenty-ninth [Joy98]. Twenty-Second [IEE98c, V^{+}96]. Twenty-Seventh [Kle96]. twice [Ano95-27]. twist [Ano96-42]. Two [Ano96-43, Ano97r, BHK^{+}88, Col95b, Glu96, Mee96, RB96, van90a]. Tworzenie [CK98]. TX [Ano98b, Bou97b]. Tycho [HLR97]. Type [WR99, AOS91, MB90, MB91, PB88]. type- [AOS91]. Type-Based [WR99]. Types [FB95a, Hou98, Lev95, Gor91]. Typesetting [Dob91, Hah91, Thi91]. Typography [Ano93e, Fur90, van88, BBKF95, HAB98]. Tyson [Ano94o]. tzu [mWfC97].

UARC [WAO^{+}95]. UART [Ano97w]. ubiquitous [Ret96]. Ubiquity [Bos99]. Ugly [LM98]. ui [HiKgK98]. UIML [APB^{+}99]. UIUC [She95]. UK [Ano94b, Ano94e, Ano95j, Ano96-33, Bow97, JRV96, Pop93a, Pop93b]. Ulm [Dud97, IEE97f]. Unambiguity [BK93a, BK93d, BK93e, Neu97b]. Unambiguous [BKW92d]. unde [Ude95b]. undergraduate [Owe96]. underground [Ros96]. Understanding [Fly98, Hef95, LC98, Per96, TBS97, DeR97, WF98b]. understood [Tit98]. Underware [And96b]. Unicode [Uns96]. Unified [Van95c]. Uniform [Fry95b, NR96c, BM91, NL95]. unique [GPE98]. UniSQL [DJ96]. United [RT95]. universal [Fra91, HW94, Rot97]. Universe [Mun95]. universes [Hil96]. Universitat [Ano97-48, Neu96, Neu94]. Université [CBO93]. Universiteit [ACM93c]. universities [Pop93a, Pop93b]. University [ACM97b, Ano95j, Ano97-42, BCG99, CL97, Dug97, EMH^{+}95, GSG98, IEE96n, IEE99b, Kum96, MDK98, Ner97, RBBW95, Sav95, Sea94, Bar95, FMC95, Gay94]. UNIX [Ano96-30, Law98, SZM^{+}95, Web95b]. Unix/X [SZM^{+}95]. Unix/X-Windows [SZM^{+}95]. Unleashed [DR94, DR95a, DR95b, DG96, M^{+}96b, Bel98, Dar97, Dar98, D^{+}99, DG95, M^{+}96a, Mor99b, SD96b]. unmasked [Col95b]. Untangle [Tus97]. Untangling [Bat95, Kas95a]. Update [Sun98b, Pop94a]. Upgrades [Ano96-30]. upgrading [Ano94u]. Upload [NM95c]. Upper [WAO^{+}95]. Urban [Bou97a, BB99b, Coy97]. US-Japan [BLG^{+}95]. USA [ABHY96, ACM96d, Ano93j, Ano94f, Ano94o, Ano95a, Ano97d, Ano97x, AA97, BFY93, Bon93, BBJR96, CM96, DGBH93, GE95, HF97, HK95b, IEE95a, LG94b, NW97, Oke93, Oke95, RS97, SSK97, Sol88, USE87, USE89, USE95, USE96, W^{+}97]. Usability [Bar94, BR95]. Usage [LCW93, MOTA95, Sch90a, Sch90b, Var95, MMV95]. Usage-Based [Var95, MMV95]. USD [Ano96-29]. Use [Als97, Ano88f, Ano95f, Arp96, AM96, BB95a, Fab95, FMBM95, HRS95, IEE96c, JW96, Mar95c, McC97a, NF96, Nor96, PC97, Pop93a, Pop93b, Ret96, Sed95, STM97, STM^{+}94, Tow95, WDM97, ZY95, Ano96-45, Ano98c, AA96, Bar89, BEE95, CBDH94, CR96, CC97, Dor95a, Dor95b, Dun95g, Ev95, HB95, Lin94, MR93b, Rie94, Van89, Van98, Xu97]. used [Ano95-27, Ric94]. Useful [GB95b]. usefulness [Shi97]. USENIX [Ano91f]. User [ACM95c, ACM96d, Als95a, Als95b, Ano93f, Ano95p, Ano97k, Ara98, ADE93, Bar92, BR95, BE95, Bun96, Can95, DDM96, FMW95, HLR97, KMA^{+}96, Mag97, MP98b, Nad96, NS95c, NS95b, Pol95, SS88, APB^{+}99, BC90, Cha93, CMPd91, DT95, DD94, EBR^{+}96, Eng97, Fuc96b, HMRL95, Lig94, MP98c, MP99, Orl93, PH97b, TC96, Wan97, Wil92]. user-centered [Cha93]. User-Interface [Ano95p]. User-Oriented [ACM96d]. Users [Ano96-36, Ano97-31, Ano98t, Bee97b, BH95b, BFH98, Bro97, Fly95, Gra96d, Ano94u, KPD96]. Uses [RBBW95, Sav95, Hef95]. Using [AVN98, ASWS98, Ano95-42, Ano97k, Ano97-53, BMDB98, Ber95, BS95c, BHK^{+}95a, BJ96, BH98b, BQRT95, Bur99a, Cam94, CP94, Chi95, CJD^{+}96, CJ98, CGLV94, CH95, CSM97, Dod94, DC98, Ebe98, ETT97, Eyz95, FJ96, For96a, Fox95, FFKO99, FB95b, Gol97, GHH97, Gro95, HLS97, Hah91, HN96a, HK95a, HM97, Hen99, HJS98, Hic89, HB97b, HTL95, Hou98, ID99, Ivl95, JOA98, KS98a, KS98b, Kru96, LO96, LOA^{+}96, Lad96, LO^{+}97, LO98, LO^{+}99b, LH96, Lew97, LB98b, LS98, Lim95a, Lim95b, Lim96, LSC96, LG95, Lun95a, Mac93, Mal98, MVKJ99, Mee96, MHP98, MN95b, Mor95a, Mor99a, Mue95, MK96a, MM95b, Nat99b, Nak94, NF96, NF97, New95a, New95b, NS95a, Pat97, PS98b, Pel97a, Phi98b, PSG92, Ple95, Pop97, Qay96, RJKS96, Res95, RWBS97, Rob94a, RM97a, SS89, SBJ96]. Using [SHW95, Shi98, SBS95b, SR96, Ste96c, Tay95b, Tes96a, Tes96b, Tuc99, Whi96b, YYOO97, ZKM95, ZJ96, ABH93, AS96, BBSM96, BH98a, BB99a, BS95b, Bot98, Can90, Chr95, CR94, CW95, Dob91, Dun95f, DM96, EM99, Erf93, Gau96b, Ges91, GL99, GR97, GR99b, Hef95, Hol99b, HMMSS95, HL97, HLL93, Int87, Int88h, Int88i, Int91c, Int91d, Int92a, Int92b, KDC95, KH95a, KIM^{+}97b, KSM^{+}98, Lap98, Lev96b, LS99, Loc97, McG92, MWVE99, Nad96, NANK96, NR96c, Pao96, Put94, RBZ95, RBBN95, RW96a, Rob94b, Rob94c, SWH95, SD98, Sha98, Spi96a, TRMR97, Tra97, WSoSS98, WS99, Wat99, Wea97]. Utah [ACM97a, IEE97b, Isk96]. utility [SR93]. Utilizes [Ano90f]. Utilizing [FP95, FMC95]. uvodu [HS97a]. UW [Ano88g]. Uxbridge [Bow97].

V [HLCG97, IEE95d, LZ98]. V-COLLIDE [HLCG97]. VA [Ano94o]. Vadstena [LR94]. validating [Alv99]. Validation [BKW93, BKW97, Liu97a]. Value [Ano94g, BS95a, Ano94b]. values [Ges91]. Variability [BB95a]. variety [Dob98, WSM96]. Various [RC96]. VAT [WF98a]. VB [Ano96-30, Keo97, T^{+}xx]. VBI [Qay96]. VBScript [Man97, Tra97]. vcetne [HS97a]. VCR [KSGJ96]. VCR-like [KSGJ96]. VE [HLS97]. Vector [GN95, Mol95a]. Vegas [AA97]. vehicle [Lig95]. vendent [San99]. vendor [New92]. Vendors [Ano95l, Wat95, Hen96b]. Ventilation [Nie99]. Vereinigung [Neu96]. Verilog [Kro99]. Vermlgen [Cou97]. verse [CR95a]. Version [Ass87a, Ass87b, Ass87c, Ass89, Ano97w, Gar97a, Har95a, Alv97b, Gre98a, Gor91]. versioning [Lo95]. Versus [DCG^{+}95, ME99, McG91, WB96]. versus/and [McG91]. Verteilten [GS87]. Very [Mar95a, V^{+}96, Yok93]. Via [Fer93b, GMG^{+}95b, NJDF96, SNS95, DWV95, FMFVN98, Hel96, RJN95b, RJN95a, SZM^{+}95, WR95, WT95, van95a]. viable [Ano96-32]. Victoria [ACM97b, DP98b, IEE98b, MG94c]. Video [BS90, ETT97, Fer93a, KSJ95, Oze96, FA94, KSGJ96, VMM98]. Videoconference [KS98a, KS98b]. Vienna [EAZ98, IEE98c, QSBC98, TW98]. View [DK97, Gib98, HC99, PST97, Sch95c, USO97, Ise90, Orl93, Vac96b]. viewed [YKT90]. Viewer [SMDO98]. Viewers [USO97, Wel95b]. Viewing [Mar97b, Ano95-43, DB96, PSG92]. views [GMS93, GMR95, Spi96a]. VII [AHB98, GRA92]. VIII [Ano90b]. Village [Ber96c]. Virginia [ABHY96, Gay94, Sea94]. Virtual [ACM99b, Ano95-27, Ano95-28, Ano95-43, BW97, Boi95, Bow97, Bro98a, BKS96, Cod98, Dan98b, DLSR98, Die98, FH97, FR97, GSV96, Hen96a, Heu98, HMRL95, LB98a, MCPD95, NM95b, Nic97, NJDF96, Res95, RJN95b, RJN95a, Roe95, Sco96a, Sha95b, SB97, SR98, SB99, SOUH97, TS96, UL98, Vac96a, VS95, BHZ96, BFM96, BMZ95, GPV97, Gre98b, Gro95, HT97, Jac98, KDC95, KMA^{+}96, Nad96, PBB^{+}96, Rag94a, Sha98, SGP96, Ten95, Van98, VD98, W^{+}98b, Wes99, WH98, And96b, Ano96-42, Bru98, Flo96, MN95b, NM95a, Ude95b, Vac96b]. Virtual-Reality [Die98]. virtues [Hay94]. VirtuFlex [Ano97w]. Virtus [Smi95, SBS95a]. Virus [SNS95]. Vision [IEE98a, MM99a, Mat98, Ano93j, BCG99, Mar94, Rub92]. visions [Oke93]. Vista [Oke93, Oke95]. Visual [Ano96-35, Ano97w, Cas96b, Cas97, GB98, Kah95, KPIE99, Sch95a, Sch96a, Sch97a, GSG98, Buc97, Ubo96]. Visualisation [CR96, GHH97, KS98b]. Visualization [AW95, BKS98, CW95, CPR98, DP98c, FH97, GN95, Gra95a, HD95, HM97, ID99, KN95b, LB98b, MVKJ99, Nag99, Owe95, Ska98, Ska99, TS96, UL98, YH97, AS96, GE95, GSV96, KN95a, MWVE99, Van98, WSoSS98, WS99]. Visualization’96 [CEG96]. Visualizing [CMV98, MB95b]. visually [KPD96]. VLSI [IEE96n, WWN96]. VMRL [HW96a]. Vocabulaire [BKP97]. Voice [Law98, GB95c]. Vol [ERPNS97a, ERPNS97b, SS89, Via98b]. Vol.1 [SSH87]. Volatile [JDGP95]. Volume [Ano96-46, HJS98, Pop97]. Volumetric [EJ95]. voor [VB95a]. Vorteile [Min98]. VP [Ubo96]. VR [FNN97]. VRBS [MN95b, NM95a]. Vrije [ACM93c]. VRML [Bat96c, Die96, Pie97, RH97, ACM99a, ACM99b, AS98, Ada97b, Ada97a, AVN98, ANM96, ANM97, Ame97, Amm97, And96b, And96c, Ano95-27, Ano95b, Ano96l, Ano96i, Ano96r, Ano96-27, Ano96-29, Ano96-30, Ano96-37, Ano96-42, Ano96-44, Ano96-45, VRM98, Ara98, AAM^{+}96, Bal96a, Bat96d, Bee97a, BKP97, Bot98, Bou97a, BD97, BE95, BK96, Bro97, BPBV96, BB99b, Cam98, CB97, Car97, CLPR96, CR96, CR97, CMP97, CNH99, CIF98, CK98, Cod98, CJ98, CPR98, Dan98b, DO96, Dei96, DM97, Die96, Die98, Dil97, Div98, EH97, EDLS98, Ell98, EBR^{+}96, Eng97, ETT97, FH97, Far96b, Fis96b, FS96, FR97, Gar97a, GB98, GN95, GPV97, Gra96a, Gra97a, Gra95a, Gra96b, GHH97, Gre98b, Gue98a, Har96, HW97, HE95, Hei96, HM97, Hen99, HJS98, HMHK97, HLCG97, Hue98, HBNV97, ID99]. VRML [Jad97, Jah99, JSY97, JSY98, Kar96, KSJ95, Kay96, Kea96, Ken96, KKK97, KK98, KFHU98, KS98a, KS98b, KIM^{+}97b, KSM^{+}98, LMM96, LHL97, LCM96, Lew97, LB98b, LCLL98, Mac96a, MVKJ99, MC97a, Mar97b, Mau98, MGBW96, MD98a, MD98b, MB95a, MBB95, MWVE99, MSMY97, MKEB98, MN95b, Mor99a, Mul97c, Nat99b, NM95b, NM95a, NAM96, Nad96, Nag99, Nic97, Nie99, NJDF96, NN97, O’C97, OBO97, PCL^{+}95, PR98a, PR98c, PR98b, PH97b, PR95a, Par95, Pel97a, Pel97b, Pes95, Pes96a, Pes96c, Pes96b, POA^{+}95, PWHK98, RMR97, RWBS97, RW98, Ric98, Roe96, RCRBR97, Ros97, SDS97, SF96, Sch97b, Sco96a, Sco96b, Sha98, SBS95a, Smi95, Soa98, SB97, SR98, SB99, SD96b, Ste95d, Ste96b, Ste96c, STM97, STMM98, SS95d, Sun97, TRMR97, Tan98, Tau97, TTG98, Tho97, TSWS96, T^{+}97]. VRML [TS99, Tub95, Ubo96, UL98, VRM97, Vac96a, Vac96b, Vac98, VT98, VT99, Van98, Ven96b, VMM98, Wan97, WDM97, WSoSS98, WS99, War98, WF98a, Wat95, Weg96, Wel95b, Whi96e, Wil97a, Wit96b, WNB97, WH98, Wu98]. VRML-Based [KSJ95]. VRML-Retrieval [DO96]. VRML/ [CNH99]. VRML/Live3D [FS96]. VRML97 [APP^{+}98]. VRwave [APP^{+}98]. VRweb [POA^{+}95]. vs [Ano93u, Ano96w, CNH99, Gro97, WHC95]. VSPLUS [Ara98]. VToD [Jad97]. vulnerabilities [Hay94]. VVT [FH97]. VW [Heu98]. VWSIM [LB98a]. vytvareni [ID98].

W [Anoxx]. W/D [Anoxx]. W3 [Sla95b, Sla95a, SS95b, Wal94, WHC95]. W3C [Cla98]. W3QL [KS98c]. W3QS [KS98c]. WA [Ano96b, Ano97v]. WAB [KPD96]. Wailea [ERPNS97a, ERPNS97b]. WAIS [Cro95b, Mor95a, Ple95]. Wandering [Hol93]. wang [Neu98]. wansong [OuP97]. Want [Ano95b, Pin95]. Wanted [Bjo96c, Lan91]. Wants [Law98, Ano96-42]. Warehouse [CJ98, Fer94, SK98]. wares [Car95b]. warfare [Ano95-27]. Wars [Law98]. Washington [ACM96b, Ano93c, Ano95k, Ano95-37, Ano96b, Ano97a, Ano97-27, BFY93, Cim96, HF97, IEE97g, Kho94, LMT98, Oke93, Oke95]. Wasserstoff [Ano96z]. Watch [Kor98, Mag97, MM99a]. water [DGG^{+}92]. water-cooled [DGG^{+}92]. WATERS [MFSF95]. wave [Car95b, Wel95b, Ubo96]. Wavelets [Lut98]. WAXweb [MBH95a, MBH95b, MBB95].

Way [RJKS96, Dun95e, Mul97c, VCB97]. Ways [CR97, De 94]. Weak [SS97]. weaknesses [Pri96a]. WeatherDigest [COMG96]. weave [Col95b]. Weaving [RMF93, Wat96, Bat96b]. Web [Ano94a, Ano95-42, Ano96w, Ano96-34, Ano96-46, Ano97w, Ano97x, Ano98s, Ano99i, Aus95, Bat96c, BT97, Bee96b, Blu95a, Boh95, BUP99, Cai94a, Cai94b, Cas94, Cas96b, Cea96a, DE95, For95a, IZ97, Keo97, Lam97, Lem96a, Lem96c, LD97a, LD97b, LZ95, McF97a, WWW95, RM96b, Tay95e, Tay96, Tea98a, ABS99, Ano94q, CB95b, Com97a, Com97b, Dan98b, DE96, Liu97a, O’C97, Pri95, RMF93, Rou98, Sin96, Tay98, Ano97-51, HK95a, Oze96, Pow94, Alv97a, Alv97b, Ame97, And96b, AYE^{+}96, AKM95, Ano95i, Ano95c, Ano95-29, Ano95r, Ano95-31, Ano95-34, Ano95-40, Ano95-41, Ano95-44, Ano95b, Ano96j, Ano96m, Ano96-27, Ano96-32, Ano96-45, Ano97-30, AGM94, Att97, Aug99, AA96, BBKL94, BBKL95, BG95a, BG95b, BVN95a, BVN95b, Bal96b, BIL95, BH98a, Bar94, BHMS98, Bat95, Bat96b]. Web [Bat96d, BG96, Bea95a, BHS^{+}96, BB95c, BSB^{+}95, BLCG92, Bir96a, BC96, BKT96, Blu95c, Bly96, Boi95, BDKK96, BS95c, BS95b, Bow96, BL98, BC95, Bri98a, Bri98b, CFMG95, CLK96, Can95, CLS96, Car98b, Car98a, Car99, Car95b, Cas96a, Cas97, Cas98, Chr96b, CB95b, Coh99, CT96, Col95b, Col95d, CB95c, Cox96, CS96, CB96, Cro95a, Cro95b, CJ94, DT95, Dav98, DT99, DL95b, DL95a, DP95a, DR94, DR95a, DR95b, DiN95, Dil97, DP98c, Dob97a, DR95c, DR95d, DB96, DBCK96, DP95b, Dun95f, DM96, EU95, EH95b, Eic95b, Eic95a, EJ95, EAU96, FMFVN98, Fie94, FLSG95, Fis95, Fis96a, Fis97, FD95, FS96, FD96, FFKO99, Fra99, FBDM95, Fry95b, Fuc96a, Gel95, Gib98, Gla95, GJ95, GMG^{+}95b, GMG^{+}95a, Gol98, GMS98, GR99a, Gra97a, Gra95b]. Web [Gra96e, Gro96, GG96, Gul96, Gut98, HW96a, HW97, HN96b, Hef95, Hei97, HC97, HiKgK98, Hel96, Hen96a, Hen96b, HH98, Hey94, HAD^{+}95, Hol95, Hol93, Hol97, H^{+}96, How96, HB95, HMMSS95, HH99b, Ing95, JMB95, JDGP95, JW96, JOA98, JRS96, JJ95a, JJ95b, JJ95c, Kap97, Kas95a, Kel94a, Kel94b, KPD96, KS95a, KN95b, KN95a, Ker97, KH97, KKK97, KLM94, Kli95, Koh95, KS98c, KCFP95, Kro99, KGK94, LP96, LaR95, LSS96, Lap98, LK95, Law98, Lem95a, Lem95c, Lem95b, LCM96, Lem96b, Lem97, LT98, LT99, Lev96a, Lev96c, Lew97, Lew95, LZ95, Lie95, LB97, LB99a, LS99, LSI98, Lig95, Lil95, Lim95a, Lim95b, Lim96, LHW96, Lin94, Lin96, LL98, LDWN96, LDW^{+}96, LSC96, Liu98, LB99b, Lot95, Lyn96, LC98, MHMS^{+}94, Mar95b, ME99, Mar99, MTU99, Mas98, McA94]. Web [McC97a, McC97b, MMJ96, McF96, McF97b, McF98, ML97, Mey99, Mor95a, Mor96, MH96, MM95b, Mul96b, Mun95, MB95b, MYN96, MRR99, Myk95, NPPR95, NB95, Neu97a, New96, NS96, Nic92, Nic95, NS95c, NS95b, NL95, NJDF96, Nør99, O’C97, O’R97, Okl97, OPW95, Owe95, Owe96, Pao96, PC97, Par95, Pau95, PF95, PK95a, Pes96c, PW96, Pfa97b, Pfa98b, Pfa98c, Pfa99, PI98, Ple95, Pow95, PW94, Pur98, Put94, Que98, Qui95, QRV95, RLA96, Rag98, Ran98, RC95, Reb95, Rem97, Ret96, RW95b, Rid95, RJN95b, RJN95a, RL95a, RS96, Ros96, Rou95, RM97b, Rud96, RL95b, Sal96, SM95, San99, SDK^{+}96, Sch96c, Sch95c, See96, SD98, Sha97, SS98, Shi98, Sin97, Sla94, SS95a, SM94, SBS95a, SZM^{+}95, SP96, SBS95b, SNS95]. Web [SMG95a, SMG95b, SMG96, Sra96, SD96b, Sta97, Ste95d, Ste96b, Str96, SS95d, Sun98b, Tak96, Tat96, Tay95c, Tay95d, Tay97a, Tay97b, Tes96a, Tes96b, TC96, Tit95a, TGHE96, Tow95, Tra97, Tus97, Uhl95, Vac96b, VH95, VW96, VSW94, VGGB95, VMM98, VB95c, Wag97, WWN96, Wat96, Wea97, Web95b, WMGL95a, WMGL95b, Wei95, Wei97, Whe96, Whi96a, Whi96b, Whi96e, Wil95, Wol95, WAB^{+}96, Xu97, Yee99, You96, Zim97, van95a, vES95, Hut96]. Web-Based [Ano97w, BUP99, CT96, JRS96, Lig95, VGGB95, HH99b, Coh99, EJ95, Lew97, LC98]. web-browsing [Liu97a]. Web-ready [BVN95a, BVN95b]. Web-Savvy [Kro99]. Web/Commodity [FFKO99]. web2mush [Cro95a]. WebBASE [YT96]. Webbed [Sap95]. Webbing [GJ96]. WebCompiler [Kro99]. Webcraft [Eyz95]. WebCrawler [Pin95]. Webfest [Her96]. WebFX [Ano96-27]. WebGUIDE [DBCK96]. WebHLA [FFKO99]. Webified [Lun96]. Webland [Fis95]. Weblet [GB95b]. WebLinker [ACD^{+}95, CAD^{+}95]. Weblint [Bow96]. WebMaker [RR94]. WebMap [Döm95b, Döm95a]. webmaster [Fis97, Mul96b, SBM96, Web95a]. Webmasters [Mic97, Del98, Del99]. Webmeister [Ude95b]. webovych [ID98]. Webs [Die96, She95, Die96]. website [JMS98]. Websites [Hob98]. Webspeak [Coe96]. WebTech [Ano97w]. Webtop [Bar97, Bar98b]. webu [PB98]. WebView [LN99]. WebWriter [CB96]. Week [Ano95-40, Ano95-41, Cea96a, Lem96c, CD97, Int94, Lem95a, Lem95c, Lem96b, LD97b, Tat96]. weekend [Cal97, Cal98]. Wegner [HV92]. Welcome [Jan96, Fie94]. Welfare [Ano96-33]. Well [De 88, Ano94t]. Well-established [De 88]. Welten [Amm97]. West [Ano97-34, Ste96d, Lav97]. Western [BUP99, DG97, LB98a]. WET [IEE96j]. WFC [Wat99]. WG3.3 [VD98]. WG3.6 [VD98]. WG5.10 [Ano93i]. WG5.2 [Jac98]. WG5.2/5.3 [Jac98]. whence [Sch96c]. Where [RARH99, Sch92]. Which [USO97, Ano95-43, Vul97, Web95b]. white [KTK^{+}95]. white-collar [KTK^{+}95]. whither [Sch96c]. Who [How99, BEE95, Ubo96]. WI [Ano97-43]. Wide [Ame97, AYE^{+}96, Ano94a, Ano95-31, Ano96m, Ano96-34, Ano96-46, Ano97x, Ano98s, Ano99i, AGM94, AA96, Bat96c, Bat96d, BSB^{+}95, Blu95c, BS95c, BS95b, Bow96, Bri98a, Bri98b, BHK^{+}95a, Cai94b, Can95, Cas94, Cas96a, Cas96b, Cas97, Cas98, Chr96b, Col95d, CB95c, Cox96, DT99, DL95b, DL95a, DP95a, DE95, DE96, EU95, FLSG95, Fis96a, FBDM95, Fry95b, GJ95, Gol98, Gra95b, Gut98, HN96b, Hen96a, Hol95, HMMSS95, Ing95, JDGP95, Kap97, Kel94a, Kel94b, KPD96, KLM94, Kli95, KCFP95, LP96, Lap98, Lev96a, LZ95, Lin94, LL98, LSC96, Liu98, Lot95, MHMS^{+}94, Mar95b, McA94, Mor95a, MM95b, Mun95, MB95b, MYN96, NB95, NS96, NL95, Nør99, WWW95, Owe95, Owe96, Pau95, PK95a, Ple95, Pow95, RLA96, Rid95, RL95a, RM96b, SM95, SDK^{+}96, Sch96c, She95]. Wide [Sin97, SM94, SZM^{+}95, SBS95b, SNS95, SMG95a, Ste95d, Tea98a, Tit95a, Tow95, VSW94, VB95c, WWN96, WMGL95a, WMGL95b, Wei95, Whe96, Whi96a, Wil95, Wol95, WAB^{+}96, Xu97, van95a, DT95, MMJ96, San96, Ano95-34, Bar94, BLCG92, Cai94a, DR94, DR95a, DR95b, DR95d, Hef95, Hey94, Hol93, KS98c, LK95, Lie95, MFSF95, Myk95, Pow94, PW94, Put94, QRV95, RW95b, See96, Shi98, SMG95b, SMG96, Str96, Tak96, TC96, Whi96b, Hut96]. Wide-Area [SBS95b]. widely [Vac96b]. WIDL [Mon98]. wihan [OuP97]. Wiley [Gra95d]. Will [Fun99, Gra96a, Ano99l, Mee96, SF96, Sed95, Tub95]. WILT [LD95b]. Window [Ano97w, BLG^{+}95]. Windows [Gan96, Mac98, Via98b, WF98b, Ano95o, Ano96-42, HK96b, HK98, HB96, Mac98, NR96b, NR97, Neo98, Neo99, SZM^{+}95, Thr95, Wat99, WF98b, ZMP96]. Windows-based [ZMP96]. WinHelp [Blu97]. Winning [ACM95c]. Winter [USE89]. WinWord [Ros96]. WIPO [Bre96]. Wired [Sin96, WNB97]. wireless [VB95b]. wirtualnej [CK98]. Wisconsin [Ano98q, IEE95a]. wise [Man95]. Wishful [Ano97-55]. wissenschaftliches [Ano97-48]. Within [SBS95b, AVN98, Sch90a, Sch90b, Sun98b]. Without [BL98, KS95a, LL97, OW95, SBS95a]. witnesses [Dob98]. Wizard [CN98]. Wonderland [MNA91a, MNA91b]. Word [BMK95, SS97, Ano95g, Ano95r, Sum94, Ano93n, Roa95]. word-processing [Ano95r]. WordPerfect [Ano95-45, Ano95-46]. Words [JCDA95]. Work [DDM96, Gol95, Mac93, SSH87, WF98a, DD94, GP90, HL95a, Mas97, MR93b, Pao95, RW96a, RMF93, Vin99, Ano95t, Ano95u]. workbench [Sul96]. workbook [Gro96]. workers [KTK^{+}95]. workers-multimedia [KTK^{+}95]. Workflow [HIYO96]. Workflows [BQRT95]. Working [Ano93i, Ano97-44, Ano99k, Cop96, RMR97, TS96, VD98, W^{+}98a, WNC95, Mee96]. works [Gol89]. Workshop [And96a, Ano90g, Ano93x, Ano94n, Cov99, Dob98, DS96, FS99, IEE95c, IEE95d, IEE95g, IEE96h, IEE96o, IS98, LCM96, Mel99, MBB90, NW97, Rut96a, TW98, Ass96, USE97a, Van93a, W^{+}97, Ano94p, Bos97, HT97, MNW98, SC98]. Workshops [IEE96j, GSV96]. Workstations [Ear87, MVKJ99]. World [ACM92, Ano93v, Ano94a, Ano95-33, Ano96-34, Ano96-33, Ano96-40, Ano97r, Ano97x, Ano98n, Ano98s, Ano99i, Bat96c, Cai94b, COW98, CM96, Cas94, Cas96b, DE95, DE96, EAZ98, GMG^{+}95b, Lut98, Mau93, Mau95, NL95, PWHK98, Pow95, RM96b, Sco96a, Tea98a, Vez96, ZY95, van95a, And96c, Ano96-42, Ano97d, GMG^{+}95a, IEE97e, Jac96, KDH96, Paq94, Rii97, San96, SHO^{+}96, Ame97, AYE^{+}96, Ano95-31, Ano95-34, Ano96m, Ano96-46, AGM94, AA96, Bar94, Bat96d, BSB^{+}95, BLCG92, Blu95c, BS95c, BS95b, Bow96, Bri98a, Bri98b, Cai94a, Can95, Cas96a, Cas97, Cas98, Chr96b, Col95d, CB95c, Cox96, DT99, DL95b, DL95a, DP95a, DR94, DR95a, DR95b, DR95d, EU95, FLSG95, Fis96a, FBDM95, Fry95b, GJ95, Gol98, Gra95b, Gut98, HN96b, Hef95, Hen96a, Hey94, Hol95]. World [Hol93, HMMSS95, Ing95, JDGP95, Kap97, Kel94a, Kel94b, KPD96, KLM94, Kli95, KS98c, KCFP95, LP96, Lap98, LK95, Lev96a, LZ95, Lie95, Lin94, LL98, LSC96, Liu98, Lot95, MHMS^{+}94, Mar95b, McA94, Mor95a, MM95b, Mun95, MB95b, MYN96, Myk95, NB95, NS96, Nør99, WWW95, Owe95, Owe96, Pau95, PK95a, Ple95, Pow94, PW94, Put94, QRV95, RLA96, RW95b, Rid95, RL95a, SM95, SDK^{+}96, Sch96c, See96, Shi98, Sin97, SM94, SZM^{+}95, SBS95b, SNS95, SMG95a, SMG95b, SMG96, Ste95d, Str96, Tak96, TC96, Tit95a, Tow95, VSW94, VB95c, WWN96, WMGL95a, WMGL95b, Wei95, Whe96, Whi96a, Whi96b, Wil95, Wol95, WAB^{+}96, Xu97]. World-Wide [Cai94b, NL95, Pow95, San96, Ano95-34, Bar94, BLCG92, Cai94a, DR94, DR95a, DR95b, DR95d, Hef95, Hey94, Hol93, KS98c, LK95, Lie95, Myk95, Pow94, PW94, Put94, QRV95, RW95b, See96, Shi98, SMG95b, SMG96, Str96, Tak96, TC96, Whi96b]. World-Wide-Web [FLSG95]. Worlds [Ara98, BE95, Dil97, Hei96, Ken96, LB98a, LMM96, LCLL98, MGBW96, Pel97a, Pel97b, Reg98, SS95d, TS96, Tho97, TSWS96, Vac96b, VT98, CR97, Dil96, GPV97, HW96a, HW97, Heu98, HBNV97, MD98a, MD98b, NAM96, SBS95a, T^{+}97, TS99]. WorldWide [Nic92, DP95b, MRR99]. WorldWideWeb [Fly95]. Worth [BMK95]. Write [LB96]. Writers [LaT89, TDT95, Faw89, Pri96b]. Writing [Ano98p, APH98, Ham97b, KW97c, MOTA95, RR94, Wil99c, ZJ96, BG96, Bir96b, Hen95, Reh95]. Wroclaw [Dan98a]. Wroclawia [CK98]. WSCG [Ska98, Ska99, TS96]. www [Kos98, ACD^{+}95, AS96, Ano92i, BCLT95, BMY95, BLG^{+}95, Ber95, Blu95b, BHK^{+}95a, CAD^{+}95, CHH^{+}95, Coe96, Del95a, DWV95, Dra95b, Dra95a, DL95c, EH95a, EKS95, EH96, FP95, FS95, FB95a, Fre95, FLF95b, FLF95a, Gar97b, GK95b, GK95a, GFH96, Gro95, GNSP95, HLB98, HJS98, HMRL95, HTL95, Hoy96, INIK96, IN96, Jan96, Kah95, KSJ95, KSGJ96, LD95b, Lin94, Mac95, ML97, MBH95a, MBH95b, MP95b, Mol95a, Mue95, Nic95, WWW95, OBO97, Pao95, PD96, Per95, PBFT97, PKT95, Pol95, PBB^{+}96, Pro95, Rad95, Rag94a, RW95a, Reh95, Ric94, RR94, SVOM95, Sjo95, Ste95b, SMDO98, Tau96, Tew95, Wat96, Wel95a, ZY95]. WWW-based [INIK96, Jan96, Dra95b, Dra95a]. WWW/Mosaic [DL95c]. WWWinda [GNSP95]. Wyoming [Wal95]. WYSIWYG [Ano95-39, BBB88, CHP88].

X [Ano97w, HJM95, Mio96, dN97]. X-Mosaic [HJM95]. X-Windows [SZM^{+}95]. X3V1.8M [New90]. X3V1.8M/SD [New90]. X3V1.8M/SD-7 [New90]. Xerox [Ano94u, Gri95]. XFDL [BB99a]. XI [Vez96]. XML [Daw98a, ABS99, APB^{+}99, Alv99, ASWS98, And99, Ano94q, Ano97m, Ano97-56, Ano99c, Ano99j, Ano99k, Ano99l, Aok99, Avi98, BSF98, Bal98b, Bal98a, BHMS98, BK99, BB99a, Bos99, Bou98, BdRD^{+}98, BS98, Bra98, Bra99, CH99b, CCD^{+}99, CH98, Cha99, CRBB98, CRV98, Cla98, Coh99, Con97, Cow99, DT99, DeR97, DeR99b, Der99a, DFF^{+}99, Dic99, DRBS98, Dos99b, Dos99a, DuC99, DC98, Eck99, Edd98b, ES99a, Edd99, Fic98a, FZ99, Fly98, FFKO99, Fun99, Gar99, GTM99, GPE98, GP98, GMS98, GP99, GR99a, GQ99, GL99, Gue98b, GR99b, Har98, Har99, HC99, Hin98, Hol98, Hom99, HH99b, Jel98, KD98, Kil99, KMS99, Kor98, Kri98, Kri99, Kru98, Kuc98, LO99a, LO^{+}99b, Lap99, LLF98, LS99, LSI98, Lig97, Lob99, Lut98, Mal98, MFVG99, ME99, Mar99, MTU99]. XML [Mas98, Mat98, McG98b, Meg98, MLD^{+}99, Mic99, Min98, Moh99, Mor99b, MRR99, NK99, NH99, Par99, Pfa97b, Pfa98b, Pfa98c, Pfa99, PMK98, PM99, RDC99, Rat98, SD98, SBGK98, SBGO99, Shi98, Sim99, Smi98, SP99b, Spe99, St.97b, St.98, SC99, St.99a, St.99b, Tit98, TMC98, TA99, WR99, Wan99, Wat99]. XML-based [ME99, SP99b]. XML-GL [CCD^{+}99]. XML/HTML [Kri98, Kru98]. XML/RDF [KMS99]. XTND [SDSS95]. XV [EAZ98]. XVe [CBO93]. XVth [CBO93]. XXIII [ACM95c]. XXX [Wat96]. Xyvision [Ano94v].

yang [AM97]. year [Whi96f]. Years [Ano96-46, Ano97r, Ano97j]. yeh [Neu98]. yesterday [Wil96b]. ying [AM97]. York [Ano89b, Ano96k, Wil97b, Wil99a, Sav95]. Yorker [Ahe93a, Ahe93b]. Yourself [Ano95-41, Lem96a, Lem96c, MC97a, Ano95-31, Ano95-40, Ano97-47, Bry99, CD97, ES99a, Eva98b, Lem95a, Lem95c, Lem95b, Lem96b, LD97a, LD97b, LT98, LT99, NH99, OH97b, Oli97a, OH97c, Oli98b, Oli99, Tat96, Via97e, Via97d]. Youth [MTH99]. yung [AM97].

Z [Bow95, BH95b, BFH98, CMV98, GC95]. za [CK98]. Zabytki [CK98]. Zakopane [RM96a]. Zeit [Ano96z]. zone [Sha98]. zooming [BHS^{+}96]. zpracovani [HS97a]. ZUM [BH95b]. ZUM’98 [BFH98].

References

Altman:1998:PSB

[A^{+}98] Russ B. Altman et al., editors.

Pacific Symposium on Biocomputing ’98: Maui, Hawaii, 4– 9 January, 1998. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1998. ISBN 981-02-3278-0. LCCN QH506.P324 1998.

Ause:1996:HUW

[AA96] Wayne Ause and Scott Arpajian. How to use the World Wide Web. Ziff-Davis Press, Emeryville, CA, USA, 1996. ISBN 1-56276-392-X (paperback). xiii + 223 pp. LCCN TK5105.888.A87 1996.

Arabnia:1997:ICI

[AA97] Hamid R. Arabnia and Farid Ahmed, editors. International Conference on Imaging Science, Systems, and Technology: CISST ’97: June 30–July 3, 1997, Las Vegas, Nevada, USA. Computer Science Research, Education, and Applications Technology, Las Vegas, NV, USA, 1997. ISBN 0-9648666-9-2. LCCN ????

Arikawa:1996:DLQ

[AAM^{+}96] M. Arikawa, A. Amano, K. Maeda, R. Aibara, S. Shimojo, Y. Nakamura, K. Hiraki, K. Nishimura, and M. Terauchi. Dynamic LoD for QoS management in the next generation VRML. In IEEE [IEE96l], pages 24–27. ISBN 0-8186-7436-9. LCCN QA76.575.I623 1996. IEEE Catalog No. 96TB100057.

Armstrong:1995:HRC

[AB95] James C. Armstrong, Jr. and Liem Bahneman. HTML Reference Card. Specialized Systems Consultants, Seattle, WA, USA, 1995. ISBN 0-916151-79-4. 14 pp. LCCN ???? US$4.50.

Althoff:1999:CBR

[ABB99] Klaus-Dieter Althoff, R. Bergmann, and Karl Branting, editors. Case-based reasoning research and development: Third International Conference on Case-Based Reasoning, ICCBR-99, Seeon Monastery, Germany, July 27–30, 1999: proceedings, volume 1650 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. ISBN 3-540-66237-5 (softcover). LCCN QA76.76.E95 I55 1999 Bar.

Apers:1996:ADT

[ABG96] P. M. G. (Petrus Maria Gerardus) Apers, M. (Mokrane) Bouzeghoub, and G. (Georges) Gardarin, editors. Advances in database technology — EDBT ’96: 5th International Conference on Extending Database Technology, Avignon, France, March 25–29, 1996: proceedings, volume 1057 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61057-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55844 1996.

Aberer:1993:PPU

[ABH93] K. Aberer, K. Bohm, and C. Huser. The prospects of publishing using advanced database concepts. Electronic Publishing—Origination, Dissemination, and Design, 6(4):469– 480, December 1993. CODEN EPODEU. ISSN 0894-3982.

Adam:1996:DLR

[ABHY96] Nabil R. Adam, B. K. Bhargava, M. Halem, and Y. Yesha, editors. Digital libraries: research and technology advances: ADL ’95 Forum, McLean, Virginia, USA, May 15–17, 1995: selected papers, volume 1082 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. ISBN 3-540-61410-9. LCCN Z699 .D532 1996.

Andres:1998:AAH

[ABO98a] F. Andres, J. Buford, and K. Ono. An application-oriented approach for HyTime structured document management. Lecture Notes in Computer Science, ?? (1460):260–269, ???? 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andres:1998:AOA

[ABO98b] F. Andres, J. Buford, and K. Ono. An application-oriented approach for HyTime structured document management. Lecture Notes in Computer Science, 1460:260–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abiteboul:1999:DWR

[ABS99] S. (Serge) Abiteboul, Peter Buneman, and Dan Suciu. Data on the web: from relations to semistructured data and XML. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-622-X. xiii + 257 pp. LCCN QA76.9.D3 A258 2000.

Adam:1994:CPT

[ABY94] Nabil R. Adam, Bharat K.

Bhargava, and Yelena Yesha, editors. CIKM 94: proceedings of the Third International Conference on Information and Knowledge Management, November 29 — December 2, 1994, National Institute of Standards and Technology (NIST), Gaithersburg, Maryland. ACM Press, New York, NY, USA, 1994. ISBN 0-89791-674-3. LCCN QA76.9.D3 I55865 1994. URL http://

www.acm.org/pubs/contents/
proceedings/cikm/191246/.

ANSI:1993:ANS

[AC93] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information technology — text and office systems: conformance testing for standard generalized markup language (SGML) systems. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1993. ISBN ???? v + 28 pp. LCCN ????

Aimar:1995:WTM

[ACD^{+}95] Alberto Aimar, James Casey, Nikos Drakos, Ian Hannell, Arash Khodabandeh, Paolo Palazzi, Bertrand Rousseau, and Mario Ruggier. WebLinker, a tool for managing WWW cross-references. Computer Networks and ISDN Systems, 28(1– 2):99–107, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1995&
                                        volume=28&issue=1-2&aid=1513.

ACM:1992:SAI

[ACM92] ACM, editor. SIGDOC ’92. The 10th Annual International Conference Conference Proceedings. Going Online. The New World of Multimedia Documentation. ACM Press, New York, NY, USA, 1992. ISBN 0-89791-532-1. LCCN QA76.9.D6 S741 1992. URL http://

                                        www.acm.org/pubs/contents/
                                        proceedings/doc/147001/.

ACM:1993:CPS

[ACM93a] ACM, editor. Conference Proceedings. SIGDOC’93. The 11th Annual International Conference. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-630-1. LCCN QA76.9.D6I57 1993.

ACM:1993:FAC

[ACM93b] ACM, editor. Fifth ACM Conference on Hypertext Proceedings. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-624-7. LCCN QA76.76.H94A26 1993.

ACM:1993:FIC

[ACM93c] ACM, editor. Fourth International Conference on Artificial Intelligence and Law. Proceedings of the Conference, June 15–18, 1993, Computer Law Institute, Vrije Universiteit Amsterdam, The Netherlands. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-606-9. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/ai/158976.

ACM:1993:PAM

[ACM93d] ACM, editor. Proceedings ACM Multimedia 93. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-596-8. LCCN ????

ACM:1994:AFA

[ACM94a] ACM, editor. ASSETS ’94:

The First Annual ACM Conference on Assistive Technologies, October 31–November 1, 1994, Doubletree Marina del Rey Hotel, Los Angeles, California. ACM Press, New York, NY, USA, 1994. ISBN 0-89791-649-2. LCCN HV1569.5.A24 1994. ACM order number 444940.

ACM:1994:ECH

[ACM94b] ACM, editor. European Conference on Hypermedia Technology 1994 proceedings / ACM European Conference on Hypermedia Technology, Edinburgh, September 18–23, 1994. ACM Press, New York, NY, USA, 1994. ISBN 0-89791-640-9. LCCN QA76.76.I59A35 1994.

ACM:1994:PPT

[ACM94c] ACM, editor. PODS 1994: proceedings of the Thirteenth ACM SIGACT-SIGMOD-SIGART Symposium on Principles of Database Systems: Minneapolis, MN, May 24–26, 1994. ACM Press, New York, NY, USA, 1994. ISBN 0-89791-642-5. LCCN ????

ACM:1995:DL

[ACM95a] ACM, editor. Digital Libraries, volume 16(2) of SIGOIS Bulletin. ACM Press, New York, NY, USA, December 1995. CO-DEN SIGBEL. ISBN ???? ISSN 0894-0819. LCCN ????

ACM:1995:SAI

[ACM95b] ACM, editor. SIGDOC ‘95. 13th Annual International Conference on Systems Documentation Conference Proceedings. Emerging from Chaos: Solutions for the Growing Complexity of Our Jobs: October 2-4, 1995, Savannah, Georgia. ACM Press, New York, NY, USA, 1995. ISBN 0-89791-713-8. LCCN QA 76.9 D6 I57 1995.

ACM:1995:WNG

[ACM95c] ACM, editor. Winning the networking game: proceedings / ACM SIGUCCS User Services Conference XXIII, October 15– 18, 1995, Marriott Pavilion, St. Louis, Missouri, volume 23 of ACM SIGUCCS User Services Conference. ACM Press, New York, NY, USA, 1995. ISBN 0-89791-704-9. LCCN QA 74 A26 1995.

ACM:1996:CGP

[ACM96a] ACM, editor. Computer graphics: proceedings: annual conference series, 1996: SIGGRAPH 96 Conference proceedings, August 4–9, 1996, New Orleans, LA. ACM Press, New York, NY, USA, 1996. ISBN 0-89791-746-4. ISSN 1069-529X. LCCN T385 .S54 1996. ACM order number: 428960.

ACM:1996:HWD

[ACM96b] ACM, editor. Hypertext ’96, Washington, DC, March 16–20, 1996: the Seventh ACM Conference on Hypertext: Proceedings. ACM Press, New York, NY, USA, 1996. ISBN 0-89791-778-2. LCCN QA76.76.H94 A26 1996. ACM order number: 614960.

ACM:1996:PTA

[ACM96c] ACM, editor. Proceedings / TRI-Ada ’96, Philadelphia Marriott Hotel, Philadelphia, Pennsylvania, December 3–7, 1996. ACM Press, New York, NY, USA, 1996. ISBN 0-89791-808-8. LCCN QA 76.73 A35 T75 1996.

ACM:1996:SAI

[ACM96d] ACM, editor. SIGDOC ’96. The 14th Annual International Conference, Conference Proceedings. Marshaling New Technological Forces: Building a Corporate, Academic, and User-Oriented Triangle, October 20–23, 1996, Research Triangle Park, NC, USA. ACM Press, New York, NY, USA, 1996. ISBN 0-89791-799-5. LCCN QA 76.9 D6 I56 1996.

Agosti:1996:DIT

[ACM96e] M. Agosti, F. Crestani, and M. Melucci. Design and implementation of a tool for the automatic construction of hypertexts for information retrieval. Information Processing and Management, 32(4):459–476, July 1996. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

ACM:1997:CCC

[ACM97a] ACM, editor. Crossroads in communication: conference proceedings / SIGDOC 97, the 15th annual International Conference on Computer Documentation, October 19–22, 1997, Salt Lake City, Utah at Snowbird Resort. ACM Press, New York, NY, USA, 1997. ISBN 0-89791-861-4. LCCN QA 76.9 D6 I56 1997.

ACM:1997:AIL

[ACM97b] ACM, editor. The Sixth International Conference on Artificial Intelligence and Law: proceedings of the conference: June 30–July 3, 1997, The University of Melbourne Law School Melbourne, Victoria, Australia, volume 6 of International Conference on Artificial Intelligence and Law. ACM Press, New York, NY, USA, 1997. ISBN 0-89791-924-6. LCCN ????

URL http://www.acm.org/
pubs/contents/proceedings/

ai/261618/. ACM order number 604972.

ACM:1998:ATI

[ACM98]  ACM, editor.  Assets ’98:

the Third International ACM Conference on Assistive Technologies: April 15–17, 1998, Marina del Rey Hotel, Marina del Rey, California. ACM Press, New York, NY, USA, 1998. ISBN 1-58113-020-1. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/assets/274497/.

ACM order number 444980.

ACM:1999:V

[ACM99a] ACM. VRML ’95. ACM Press, New York, NY, USA, 1999. ???? pp.

ACM:1999:VPS

[ACM99b] ACM. VRML ’97 proceedings of the Second Symposium on Virtual Reality Modeling Language, February 24-26 1997, Monterey, CA. ACM Press, New York, NY, USA, 1999. ???? pp.

Adachi:1992:OIS

[Ada92] J. Adachi. Overview on international standard ‘Document Style Semantics and Specification Language (DSSSL)’. Journal of the Information Processing Society of Japan = Joho Shori, 33(2):176–184, ???? 1992. CODEN JOSHA4. ISSN 0447-8053.

Adams:1995:ELS

[Ada95]  Roy J. Adams.  Electronic

libraries SGML applications:

Background to project ELSA. Program, 29(4):397–406, October 1995. CODEN PRGRDU, PRGMBD. ISSN 0033-0337.

Adams:1997:VMH

[Ada97a] Ellen Adams. VRML 2 Master’s Handbook. Prima Tech Publishing, Rocklin, CA, USA, May 1997. ISBN 0-7615-0774-4. 900 pp. LCCN ???? US$45.00.

Adams:1997:VDT

[Ada97b] Ellen Adams. VRML Developer’s Toolkit. McGraw-Hill Companies, New York, NY, USA, May 1997. ISBN 0-07-913105-0. 304 pp. LCCN ???? US$39.95.

Arpinar:1993:MAG

[ADE93] İsmailcem Budak Arpinar, Asuman Doğaç, and Cem Evrendilek. Moodview: An advanced graphical user interface for OODBMSs. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(4):11–18, December 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Adkins:1988:DPP

[Adk88] Allen L. Adkins. Data preparation and premastering. In Sherman [She88], pages 343– 396. ISBN 0-07-056578-3. LCCN TK7882.C56 C332 198. US$59.95.

Adler:1992:BSS

[Adl92] S. Adler. The birth of a standard (SGML). Journal of the American Society for Information Science, 43(8):556–558, September 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ansari:1994:EWW

[AGM94] S. G. Ansari, P. Giommi, and A. Micol. ESIS on the World Wide Web. International Journal of Modern Physics C [Physics and Computers], 5(5): 805–809, October 1994. CO-DEN IJMPEO. ISSN 0129-1831 (print), 1793-6586 (electronic).

Albrecht:1998:ADA

[AHB98] Rudolf Albrecht, Richard N.

Hook, and Howard A. Bushouse, editors. Astronomical data analysis software and systems VII: proceedings from a meeting held in Sonthofen, Germany, 14–17 September 1997, volume 145 of Astronomical Society of the Pacific conference series. Astronomical Society of the Pacific, San Francisco, CA, USA, 1998. ISBN 1-886733-65-1. ISSN 1080-7926. LCCN QB51.3.E43 A87 1997.

Aheam:1993:SNY

[Ahe93a] Hally Aheam. SGML and The New Yorker Magazine. Technical Communication, 40(2):226– ??, May 1, 1993. CODEN TL-CMBT. ISSN 0049-3155.

Ahearn:1993:SNY

[Ahe93b] Hally Ahearn. SGML and The New Yorker Magazine. Technical Communication, 40(2):226– 229, May 1993. CODEN TL-CMBT. ISSN 0049-3155.

Ahonen:1995:AGS

[Aho95] H. Ahonen. Automatic generation of SGML content models. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3):195–206, June/ September 1995. CODEN EPODEU. ISSN 0894-3982.

Ahonen:1997:DSC

[Aho97] H. Ahonen. Disambiguation of SGML content models. In Nicholas and Wood [NW97], pages 27–38. CODEN LNCSD9. ISBN 3-540-63620-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .P63 1996.

Andrews:1995:SIW

[AKM95] Keith Andrews, Frank Kappe, and Hermann Maurer. Serving information to the Web with Hyper-G. Computer Networks and ISDN Systems, 27(6): 919–926, April 3, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

URL http://www.elsevier.
com/cas/tree/store/comnet/
sub/1995/27/6/1452.pdf.

Aronson:1997:HMS

[AL97] Larry Aronson and Joseph Lowery. HTML3.2 manual of style. Ziff-Davis Press, Emeryville, CA, USA, 1997. ISBN 1-56276-529-9. xviii + 385 pp. LCCN QA76.76.H94 A77 1997. Contents: What is HTML? — The HTML language – Writing HTML documents — Tutorials — Advanced design — Interactivity — HTML examples.

Alschuler:1995:ASA

[Als95a] Liora Alschuler. ABCD … SGML: A User’s Guide to Structured Information. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1995. ISBN 1-85032-197-3. xviii + 414 pp. LCCN QA76.76.H94 A43 1995 Reserves. US$39.95.

Alschuler:1995:SUG

[Als95b] Liora Alschuler. SGML: A User’s Guide to Structured Information. Management Information Systems Series. Van Nostrand Reinhold, New York, NY, USA, 1995. ISBN 0-442-01858-4. 250 pp. LCCN ???? US$39.95.

Alschuler:1997:TUS

[Als97] L. Alschuler. Tutorial on the use of structured information (SGML) in healthcare applications. In Dudeck [Dud97], pages 207–208. ISBN 90-5199-363-3, 4-274-90185-8. ISSN 0926-9630. LCCN ???? AlvarezGarcia:1997:HCPa

[Alv97a] Alonso Alvarez Garcia. HTML, creacion de paginas Web. Guias practicas. Anaya Multimedia, Madrid, Spain, 1997. ISBN 84-7614-858-5. 352 pp. LCCN ????

AlvarezGarcia:1997:HCPb

[Alv97b] Alonso Alvarez Garcia. HTML creacion de paginas Web: (actualizada a la version 3.2). Guias Practicas. Anaya Multimedia, Madrid, Spain, second edition, 1997. ISBN 84-415-0242-0. 333 pp. LCCN ????

Alvarado:1999:NXP

[Alv99] Javier Alvarado. Non-validating XML processor implementation. Thesis (b.s.), California Polytechnic State University, San Luis Obispo, CA, USA, 1999. v + 55 pp.

Arpajian:1996:HUHb

[AM96] Scott Arpajian and Robert Mullen. How to Use HTML 3.2. Que Corporation, Indianapolis, IN, USA, November 1996. ISBN 1-56276-496-9. x + 219 pp. LCCN QA76.76.H94A766 1996. US$24.99.

Arpajian:1997:TYY

[AM97] Scott Arpajian and Robert Mullen. Tsen yang ying yung HTML 3.2. Tsai se tien nao tsung shu; 18. Wan li chi kou, Wan li shu tien, Hsiang-kang, 1997. ISBN 962-14-1220-X. 212 pp. LCCN ????

Agosti:1995:AAC

[AMC95] H. Agosti, M. Melucci, and F. Crestani. Automatic authoring and construction of hypermedia for information retrieval. Multimedia Systems, 3(1):15– 24, February 1995. CODEN MUSYEW. ISSN 0942-4962.

ANSI:1985:IPT

[Ame85] American National Standards Institute and International Organization for Standardization. Information processing: text and office systems: Standard Generalized Markup Language (SGML). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1985. ISBN ???? viii + 158 pp. LCCN ????

ANSIX3:1991:PAN

[Ame91] American National Standards Institute Committee X3. Proposed American National Standard: information technology: test and office systems: conformance testing for standard generalized markup language (SGML) systems: draft 3/4/ 1991. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1991. ISBN ???? 28 pp. LCCN ????

Ames:1997:VTD

[Ame97] A. L. Ames. VRML: Three dimensions for the World Wide Web. In Anonymous [Ano97b], pages 151–?? ISBN 0-914548-94-8, 0-914548-95-6. LCCN ????

ArnoldMoore:1995:EDM

[AMFL^{+}95] T. Arnold-Moore, M. Fuller, B. Lowe, J. Thom, and R. Wilkinson. The ELF data model and SGQL query language for structured document databases. Australian Computer Science Communications, 17(2): 17–26, ???? 1995. CODEN AC-SCDD. ISSN 0157-3055.

Ammann:1997:PAW

[Amm97] Eckhard Ammann. Programmierung animierter Welten: Java, JavsScript und VRML. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1997. ISBN 3-8266-0329-X (??invalid checksum??). 310 pp. LCCN ???? 79 DM. URL

                                        http://www.itp.de/. Includes

CD-ROM.

Ahonen:1997:GGS

[AMN97] H. Ahonen, H. Mannila, and E. Nikunen. Generating grammars for SGML tagged texts lacking DTD. Mathematical and computer modelling, 26(1):1–??, ???? 1997. CODEN MCMOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

Andrews:1988:EPF

[And88] K. Andrews. Electronic publishing futures: organizational and technical issues. In Anonymous [Ano88e], pages 15/1/1–4. ISBN ???? LCCN ????

Andaloussi:1996:SW

[And96a] J. Andaloussi. SGML workshop. In Anonymous [Ano96v], pages 75–78. ISBN ???? LCCN ????

Anderson:1996:GYV

[And96b] Lisa Anderson. Grab your VRML underware! meet Virtual Reality Modeling Language, the 3D language that may conquer the Web. CADalyst: the newsletter of the Auto-CAD Users’ Group, ??(??):76– ??, April 1996. ISSN 0820-5450.

Andoh:1996:VWS

[And96c] Y. Andoh. VRML’s 3D world:

sculpting the next generation Internet — technology and trends in VRML. Joho-Shori (J. Information Processing Soc. Japan), 37(10):899–906, October 1996. CODEN JOSHA4. ISSN 0447-8053.

Andrews:1996:JFB

[And96d] David L. Andrews. Java forms:

Better than HTML. BYTE Magazine, 21(11):38, November 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discuss the Jet-

Form (http://www.jetform.
com/) and Caere (http://www.

caere.com/) products for interactive forms applications.

Andersson:1997:DPP

[And97] Mattias Andersson. DF printing and publishing: the next revolution after Gutenberg. Micro Publishing Press, Torrance, CA, USA, 1997. ISBN 0-941845-22-2. vii + 198 pp. LCCN ????

Andrivet:1999:SXP

[And99] Sebastien Andrivet. A simple XML parser. C/C++ Users Journal, 17(7):22, 24, 26–28, 30, 32, July 1999. CODEN CCU-JEX. ISSN 1075-2838.

Ames:1996:VS

[ANM96] Andrea L. Ames, David R.

Nadeau, and John L. Moreland.

The VRML sourcebook. Wiley, New York, NY, USA, 1996. ISBN 0-471-14159-3. xxi + 650 pp. LCCN QA76.76.H94 A52 1996. US$29.95.

Ames:1997:VS

[ANM97] Andrea L. Ames, David R. Nadeau, and John L. Moreland. The VRML 2.0 sourcebook. Wiley, New York, NY, USA, second edition, 1997. ISBN 0-471-16507-7. xxxiii + 654 pp. LCCN QA76.76.H94A52 1997. US$49.95. URL http:

//www.wiley.com/compbooks/
vrml2sbk/cover/cover.htm.

Anonymous:1988:EN

[Ano88a] Anonymous. 4th European Networkshop. Computer Networks and ISDN Systems, 16(1-2):138– 142, September 1988. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Anonymous:1988:AC

[Ano88b] Anonymous. ALLC conference. Literary and Linguistic Computing, 3(3):162–165, 1988. CO-DEN LLCOEI. ISSN 0268-1145.

Anonymous:1988:CDG

[Ano88c] Anonymous. Composing documents with the generalized markup language. Technical Report S544-3421, IBM Corporation, San Jose, CA, USA, March 1988.

Anonymous:1988:EII

[Ano88d] Anonymous, editor. Electronic Imaging ’88. International Electronic Imaging Exposition and Conference. Advance Printing of Paper Summaries. Inst. Graphic Commun, Boston, MA, USA, 1988. ISBN ???? LCCN ???? 2 vol.

Anonymous:1988:ECR

[Ano88e] Anonymous, editor. Electro/ 88 Conference Record. Electron. Conventions Manage, Los Angeles, CA, USA, 1988. ISBN ???? LCCN ????

Anonymous:1988:EPT

[Ano88f] Anonymous, editor. ESPRIT ’88. Putting the Technology to Use. Proceedings of the 5th Annual ESPRIT Conference. North-Holland Publishing Co., Amsterdam, The Netherlands, 1988. ISBN 0-444-87145-4. LCCN QA75.5.E84 1988. Two volumes.

Anonymous:1988:FAC

[Ano88g] Anonymous, editor. Fourth Annual Conference of the UW Centre for the New Oxford English Dictionary: Information in Text. Proceedings of the Conference. University of Waterloo, Waterloo, Ontario, Canada, ???? 1988. ISBN ???? LCCN ????

Anonymous:1989:ICC

[Ano89a] Anonymous. International conference on computing in highenergy physics. Computer Physics Communications, 57 (1–3):244–250, December 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anonymous:1989:NOM

[Ano89b] Anonymous, editor. National Online Meeting proceedings– 1989, New York, May 9– 11, 1989. Learned Information, Medford, NJ, USA, ???? 1989. ISBN 0-938734-34-2. LCCN QA76.55 .N37 1989.

Anonymous:1989:NCP

[Ano89c] Anonymous, editor. NCGA ’89 Conference Proceedings. 10th Annual Conference and Exposition Dedicated to Computer Graphics. National Computer Graphics Association, Fairfax, VA, USA, 1989. ISBN ???? LCCN ???? 3 vol.

Anonymous:1990:COI

[Ano90a] Anonymous. 1990 conference on office information systems. SIGOIS Bulletin, 11(2-3):118– 122, 1990. CODEN SIGBEL. ISSN 0894-0819.

Anonymous:1990:AAI

[Ano90b] Anonymous. Applications of Artificial Intelligence VIII. Proceedings of the SPIE — The International Society for Optical Engineering, 1293(1):368– 379, 1990. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Anonymous:1990:CPS

[Ano90c] Anonymous. Call for papers: SGML & TEX conference, 31 August 1990, Groningen, The Netherlands. TUGboat, 11(1): 126, April 1990. ISSN 0896-3207.

Anonymous:1990:ETI

[Ano90d] Anonymous, editor. Electronic transfer of information and its impact on aerospace and defence research and development: copies of papers presented at the Technical Information Panel specialists’ meeting, held at the Etat-Major Géneřal, Brussels, Belgium, 17th–19th October 1989, volume 466 of AGARD conference proceedings. AGARD, Neuilly sur Seine, France, 1990. ISBN 92-835-0550-6. LCCN TL500.N6 C6 no.466.

Anonymous:1990:FTS

[Ano90e] Anonymous. Future trends in spectroscopy, symposium. Spectrochimica Acta, Part A [Molecular Spectroscopy], Special issue: 11–45, 1990. CODEN SAM-CAS. ISSN 0584-8539.

Anonymous:1990:IUS

[Ano90f]  Anonymous.  IRS utilizes

SGML. Graphic arts monthly: the magazine of the printing industry, 62(2):88–??, February 1, 1990. CODEN GAMOE4. ISSN 1047-9325.

Anonymous:1990:MIC

[Ano90g] Anonymous, editor. Multimedia ’90. 3rd IEEE ComSec International Workshop on Multimedia Communications. CNET, Lannion, France, 1990. ISBN ???? LCCN ????

Anonymous:1990:NCP

[Ano90h] Anonymous, editor. NCGA ’90.

Conference Proceedings. National Computer Graphics Association, Fairfax, VA, USA, 1990. ISBN ???? LCCN ???? 3 vol.

Anonymous:1991:BCG

[Ano91a] Anonymous, editor. Bridging the Communication Gap (AGARD-CP-487). AGARD, Neuilly sur Seine, France, 1991. ISBN ???? LCCN ????

Anonymous:1991:CTC

[Ano91b] Anonymous, editor. Conformance Testing and Certification in Information Technology and Telecommunications. Proceedings of the European Conference. IOS Press, Amsterdam, The Netherlands, 1991. ISBN ???? LCCN ????

Anonymous:1991:DDE

[Ano91c] Anonymous, editor. DEXA 91. Database and Expert Systems Applications Proceedings of the International Conference. Springer-Verlag Inc., New York, NY, USA, 1991. ISBN ???? LCCN ????

Anonymous:1991:DIP

[Ano91d]  Anonymous, editor.  Doc-

ument Image Processing 91. Proceedings of the Conference. Blenheim Online, London, UK, 1991. ISBN 0-86353-256-X. LCCN ????

Anonymous:1991:ICS

[Ano91e] Anonymous, editor. IEE Colloquium on ‘Standards and Practices in Electronic Data Interchange’ (Digest No.106). IEE, London, UK, 1991. ISBN ???? LCCN ????

Anonymous:1991:PSU

[Ano91f] Anonymous, editor. Proceedings of the Summer 1991 USENIX Conference. USENIX Association, Berkeley, CA, USA, 1991. ISBN ???? LCCN ????

Anonymous:1992:JEN

[Ano92a] Anonymous. 3rd Joint European Networking Conference. Computer Networks and ISDN Systems, 25(4-5):454– 459, November 1992. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Anonymous:1992:IDI

[Ano92b] Anonymous. Information Dimensions, Inc. provides SGML document database for JCALS. Library Hi Tech News, ??(91): 12–??, April 1, 1992. ISSN 0741-9058.

Anonymous:1992:ISK

[Ano92c] Anonymous. Interleaf 5 SGML keeps companies from being tied to one authoring program. InfoWorld, 14(??):21, June 1992. CODEN INWODU. ISSN 0199-6649.

Anonymous:1992:L

[Ano92d] Anonymous. Letters. BYTE Magazine, 17(9):20–??, September 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:SEP

[Ano92e] Anonymous. SGML and electronic publishing. OCLC newsletter, ??(200):13–14, November 1992. CODEN OCLNDT. ISSN 0163-898X.

Anonymous:1992:SFI

[Ano92f] Anonymous. SGML frees information. BYTE Magazine, 17 (6):279–286, June 1992. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:S

[Ano92g] Anonymous. Standards. UNIX/ world, 9(11):99–??, November 1992. ISSN 0739-5922.

Anonymous:1992:TCT

[Ano92h] Anonymous, editor. Translating and the Computer 13. The Theory and Practice of Machine Translation — A Marriage of Convenience? Aslib, London, UK, ???? 1992. ISBN 0-85142-295-0. LCCN ????

Anonymous:1992:WHB

[Ano92i] Anonymous. WWW (hypertext browser). SIGBIO Newsletter, 12(3):37–40, September 1992. CODEN SINWDG. ISSN 0163-5697 (print), 1557-9506 (electronic).

Anonymous:1993:ICC

[Ano93a] Anonymous, editor. 4th International Conference on Cognitive and Computer Sciences for Organizations. Communicating Knowledge in Organizations. Proceedings. Univ. Quebec, Montreal, Que., Canada, 1993. ISBN 2-7624-0555-6. LCCN ????

Anonymous:1993:A

[Ano93b] Anonymous. Auto-graphics. Library Hi Tech, ??(101):23–??, April 1993. CODEN LIHTD2. ISSN 0737-8831.

Anonymous:1993:CCW

[Ano93c] Anonymous, editor. CE and CALS Washington ’93. Fabricators and Manuf. Assoc, Rockford, IL, USA, 1993. ISBN ???? LCCN ????

Anonymous:1993:CSE

[Ano93d] Anonymous. Corporate sites embrace the Standard Generalized Markup Language technology for coding electronic documents. PC Week, 10(8):37–??, March 1993. ISSN 0740-1604.

Anonymous:1993:EFI

[Ano93e] Anonymous. EP ’94. fifth international conference on electronic publishing, document manipulation and typography. Electronic Publishing—Origination, Dissemination, and Design, 6(4), December 1993. CO-DEN EPODEU. ISSN 0894-3982.

Anonymous:1993:HPH

[Ano93f] Anonymous, editor. HPECU/ HPCUA Proceedings of the 1993 Hewlett–Packard Computer Users’ European Conference. HP Comput. Users Assoc, Harrow, UK, 1993. ISBN ???? LCCN ????

Anonymous:1993:ICIa

[Ano93g] Anonymous, editor. IEE Colloquium on ‘Information Access for People with Disability’ (Digest No.103). IEE, London, UK, 1993. ISBN ???? LCCN ????

Anonymous:1993:ICIb

[Ano93h] Anonymous, editor. IEE Colloquium on ‘Issues in Computer Support for Documentation and Manuals’ (Digest No.1993/170). IEE, London, UK, 1993. ISBN ???? LCCN ????

Anonymous:1993:ITW

[Ano93i] Anonymous. IFIP TC5/WG5.10 working conference on interfaces in industrial systems for production and engineering. IFIP Transactions B [Applications in Technology], B-10:175–186, 1993. CODEN ITBTEH. ISSN 0926-5481.

Anonymous:1993:IVC

[Ano93j] Anonymous, editor. Implementing the vision: CALS Expo International ’93: 6th Annual conference and exposition — November 1993, Atlanta, GA, USA. NSIA, Washington DC, USA, 1993. ISBN ???? ISSN ???? LCCN ????

IEEE:1993:IPN

[Ano93k] Anonymous, editor. IPCC 93 Proceedings. The New Face of Technical Communication: People, Processes, Products. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, ???? 1993. ISBN 0-7803-1465-4 (paperback), 0-7803-1466-2. LCCN T 10.5 I54 1993. IEEE catalog no. 93CH3367-0.

Anonymous:1993:ISM

[Ano93l] Anonymous. Is SGML for me?

Publishing and production executive, 7(8):22–??, September 1, 1993. CODEN PPEXEP. ISSN 1048-3055.

Anonymous:1993:L

[Ano93m] Anonymous. Letters. BYTE Magazine, 18(13):18–??, December 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:MSA

[Ano93n] Anonymous. Microsoft SGML Author for Word. Desktop Publishing Commentary, 9(6):14– 15, November 1993. CODEN DPCOEQ. ISSN 0957-3178.

Anonymous:1993:NSF

[Ano93o] Anonymous. A new standard of freedom. Information Week, ?? (418):22–??, March 1993. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:1993:PSA

[Ano93p] Anonymous, editor. Proceedings. Second Annual Symposium on Document Analysis and Information Retrieval. Univ. Nevada, Las Vegas, NV, USA, 1993. ISBN ???? LCCN ????

Anonymous:1993:SSF

[Ano93q] Anonymous. Seybold Seminars 93 featured SGML and graphics applications. InfoWorld, 15(17): 26–??, April 1993. CODEN IN-WODU. ISSN 0199-6649.

Anonymous:1993:SOP

[Ano93r]  Anonymous.  SGML open

promises a better standard for long-document publishing. InfoWorld, 15(9):7–??, March 1993. CODEN INWODU. ISSN 0199-6649.

Anonymous:1993:SAI

[Ano93s] Anonymous. Sixteenth Annual International ACM SIGIR Conference on Research and Development in Information Retrieval. SIGIR forum, 1993. CODEN FASRDV. ISSN 0163-5840.

Anonymous:1993:Ta

[Ano93t] Anonymous. Technology. Publishers weekly, 240(33):36–??, August 1993. ISSN 0000-0019.

Anonymous:1993:TSV

[Ano93u] Anonymous. Technology. SGML vs. Acrobat. Publishers weekly, 240(33):36–??, August 16, 1993. ISSN 0000-0019.

Anonymous:1993:TDS

[Ano93v] Anonymous. Toward a digital, SGML world. Publishing and production executive, 7(6): 16–??, July 1, 1993. CODEN PPEXEP. ISSN 1048-3055.

Anonymous:1993:Tb

[Ano93w] Anonymous. Trends. PC Magazine, 12(7):29–??, April 1993. CODEN PCMGEP. ISSN 0888-8507.

Anonymous:1993:WHH

[Ano93x] Anonymous. Workshop on hypermedia and hypertext standards. Information services and use, 13(2):93–102 (??), ???? 1993. CODEN ISUDX8. ISSN 0167-5265.

Anonymous:1994:SMT

[Ano94a] Anonymous. 1994 San Miniato Topical Seminar on World Wide Web and Beyond in Physics Research and Applications. International Journal of Modern Physics C [Physics and Computers], 5(5):??, October 1994. CO-DEN IJMPEO. ISSN 0129-1831 (print), 1793-6586 (electronic).

Anonymous:1994:AVD

[Ano94b] Anonymous, editor. Adding value to documents with markup languages: Colloquium — June 1994, London, UK, number 142 in Colloquium Digest — IEE 1994. IEE, London, UK, 1994. ISBN ???? LCCN ????

Anonymous:1994:BSS

[Ano94c] Anonymous. Basis SGML server available from IDI. Information Retrieval and Library Automation, 29(10):9–??, March 1994. CODEN IRLAAQ. ISSN 0020-0220.

Anonymous:1994:CGS

[Ano94d] Anonymous. Coming to grips with SGML. Against the Grain, 6(3):12–??, June 1, 1994. ISSN 1043-2094.

Anonymous:1994:DCC

[Ano94e]  Anonymous, editor.  Docu-

ment computing: Conference — November 1994, London, UK. Technology Appraisals, ????, 1994. ISBN 1-871802-32-6. LCCN ????

Anonymous:1994:IPC

[Ano94f] Anonymous, editor. ICPS ’94: the physics and chemistry of imaging systems: Joint meeting — May 1994, Rochester, NY, USA, IS and T Annual Conference 1994; Conf 47//V2. Society for Imaging Science and Technology, Bellingham, WA, USA, 1994. ISBN 0-89208-177-5. LCCN ????

Anonymous:1994:ICA

[Ano94g] Anonymous, editor. IEE Colloquium on ‘ Adding Value to Documents with Markup Languages’ (Digest no.1994/142). IEE, London, UK, ???? 1994. ISBN ???? LCCN ????

Anonymous:1994:ICT

[Ano94h]  Anonymous, editor.  IEE

Colloquium on ‘Teleworking and Teleconferencing’ (Digest No.1994/144). IEE, London, UK, 1994. ISBN ???? LCCN ????

Anonymous:1994:N

[Ano94i] Anonymous. In a nutshell. Network World, 11(50):67–??, December 1994. ISSN 0887-7661 (print), 1944-7655 (electronic).

Anonymous:1994:IS

[Ano94j] Anonymous. Introduction to SGML. Asean proprint: the graphic arts magazine, pages 12–??, February 1, 1994. ISSN 1036-9112.

Anonymous:1994:MST

[Ano94k] Anonymous. Microsoft SGML tools ease document exchange. Information Week, ??(495):58– ??, October 1994. CODEN IN-FWE4. ISSN 8750-6874.

Anonymous:1994:PD

[Ano94l] Anonymous, editor. Proceedings of DOCUMATION ’94. Graphic Commun. Assoc, Alexandria, VA, USA, ???? 1994. ISBN ???? LCCN ????

Anonymous:1994:PSE

[Ano94m] Anonymous, editor. Proceedings. SHARE Europe Spring Conference. SHARE Europe (SEAS), Carouge/Geneva, Switzerland, 1994. ISBN ???? LCCN ????

Anonymous:1994:SSA

[Ano94n] Anonymous. SARA (SGML Aware Retrieval Application) Workshop 29th June 1994. The British Library research and development reports, 6173(??):1– ??, 1994. CODEN BLRRDP. ISSN 0308-2385.

Anonymous:1994:SCN

[Ano94o] Anonymous, editor. SGML ’94 Conference — November 1994, Tyson’s Corner, VA, USA, volume 24(6) of Seybold Report on Publishing Systems 1994. Seybold Publications, San Francisco, CA, USA, 1994. ISBN ???? ISSN 0736-7260. LCCN ????

Anonymous:1994:SAR

[Ano94p] Anonymous, editor. SGML aware retrieval application workshop — 1994 Jun: Oxford, volume 6173 of BLRD Reports 1994. Marc Fresco Consultancy, Kenley, 1994. ISBN ???? ISSN 0308-2385. LCCN ????

Anonymous:1994:SXW

[Ano94q] Anonymous. The SGML/XML web page, 1994.

Anonymous:1994:STC

[Ano94r] Anonymous, editor. Society for Technical Communication / 1994 proceedings / 41st Annual Conference, May 15–18, 1994, Minneapolis Convention Center, Minneapolis, Minnesota, volume 41. Society for Technical Communication, Arlington, VA, USA, 1994. ISBN 0-914548-79-4, 0-914548-80-8. LCCN T 10.5 S63 1994.

Anonymous:1994:S

[Ano94s] Anonymous. Software. Australian Personal Computer, 15 (5):14–??, May 1, 1994. ISSN 0725-4415.

Anonymous:1994:SGM

[Ano94t] Anonymous. The Standard Generalized Markup Language is finding an enthusiastic audience not only in DOD but in other agencies as well. Federal computer week, 8(35):28–??, December 1994. ISSN 0893-052X.

Anonymous:1994:XIU

[Ano94u] Anonymous. Xerox is upgrading its SGML editor, InContext, to let users print out marked-up documents. InfoWorld, 16(44): 32–??, October 1994. CODEN INWODU. ISSN 0199-6649.

Anonymous:1994:XPD

[Ano94v] Anonymous. Xyvision’s PDM:

database management for SGML and complex documents. Seybold Report on Publishing Systems, 23(17):3–29, May 1994. CODEN SRPSD9. ISSN 0736-7260.

Anonymous:1995:ACA

[Ano95a] Anonymous, editor. 20th Annual conference — April 1995, Orlando, FL, USA, SUGI 1995; 20th. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-211-0. LCCN ????

Anonymous:1995:WSW

[Ano95b] Anonymous. 3-D Web sites — want to add a new dimension to your Web site? here’s how VRML lets you build or explore the Web in 3-D. PC/Computing, 8(9):520–??, ???? 1995. CO-DEN PCMPEI. ISSN 0899-1847.

Anonymous:1995:IBW

[Ano95c] Anonymous. (a) Internet — building a Web site with HTML and the right resources. LAN times, 12(24):109–??, ???? 1995. ISSN 1040-5917.

Anonymous:1995:BHC

[Ano95d] Anonymous. BeyondPress 1.0:

HTML converter for QuarkX-Press. Macworld, 12(12):73–??, ???? 1995. CODEN MACWEA. ISSN 0741-8647.

Anonymous:1995:CEP

[Ano95e]  Anonymous, editor.  CALS

Europe ’95. Proceedings of the 6th International Conference on CALS and Enterprise Integration in Europe. World Comput. Graphics Found, Fall Church, VA, USA, 1995. ISBN ???? LCCN ????

Anonymous:1995:CNG

[Ano95f] Anonymous, editor. Cultivating New Ground in Electronic Information: Use of the Information Highway to Support Agriculture, volume 3(2) of Journal of Agricultural and Food Information. Haworth Press, New York, NY, USA, 1995. CODEN JFOIEU. ISBN ???? ISSN 1049-6505. LCCN ????

Anonymous:1995:DAC

[Ano95g] Anonymous. DataViz app converts HTML pages to word processing docs. MacWEEK, 9(28): 4–??, ???? 1995. CODEN MWEEEI. ISSN 0892-8118.

Anonymous:1995:DSI

[Ano95h]  Anonymous, editor.  Deci-

sion Sciences Institute: Annual meeting; 26th — November 1995, Boston, MA, Proceedings of the Annual Meeting— Decision Sciences Institute 1995/V2. ????, ????, 1995. ISBN ???? LCCN ????

Anonymous:1995:DW

[Ano95i] Anonymous. Dialects of the Web. BYTE Magazine, 20 (3):30–??, March 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:EUC

[Ano95j] Anonymous, editor. Eurographics UK: conference proceedings, 1995: Loughborough University, 28–30 March 1995. Eurographics UK, Abingdon, UK, 1995. ISBN 0-9521097-2-7. LCCN T385.E975 1995.

Anonymous:1995:FPA

[Ano95k] Anonymous, editor. Final Program and Advance Printing of Papers, IS and T’s 48th Annual Conference. Imaging on the Information Superhighway: May 7–11, 1995, Omni Shoreham Hotel, Washington, DC. Soc. Imaging Sci. and Technol, Springfield, VA, USA, 1995. ISBN 0-89208-184-8. LCCN TA1632.I83 1995.

Anonymous:1995:GGE

[Ano95l] Anonymous. GA — Graph Expo plays to print, pre-press pack — vendors line up behind ColorSync 2.0 — moving electronic layouts to HTML — Kodak hones Approval proofing system. MacWEEK, 9(41):12–??, ???? 1995. CODEN MWEEEI. ISSN 0892-8118.

Anonymous:1995:GAC

[Ano95m] Anonymous, editor. GIS/LIS Annual Conference and Exposition proceedings: November 14–16, 1995, Nashville Convention Center, Nashville, Tennessee. American Society for Photogrammetry and Remote Sensing and American Congress on Surveying and Mapping, Bethesda, MD, USA, 1995. ISBN 1-57083-026-6 (set), 1-57083-027-4 (vol. 1), 1-57083-028-2 (vol. 2). LCCN G70.2 .G5743 1995. Two volumes.

Anonymous:1995:HPS

[Ano95n] Anonymous. HoTMetaL Pro 2.0 spills out HTML 3. MacWEEK, 9(29):16–??, ???? 1995. CO-DEN MWEEEI. ISSN 0892-8118.

Anonymous:1995:HW

[Ano95o] Anonymous. HTML for Windows. Information Retrieval and Library Automation, 31(6):11– ??, ???? 1995. CODEN IR-LAAQ. ISSN 0020-0220.

Anonymous:1995:ICA

[Ano95p] Anonymous, editor. IEE Colloquium on ‘The Authoring and Application of Hypermedia-Based User-Interfaces’ (Digest No.1995/202). IEE, London, UK, 1995. ISBN ???? LCCN ????

Anonymous:1995:IDH

[Ano95q] Anonymous, editor. IEE Data Highway. IEE Switzerland Section, Bremgarten, Switzerland, 1995. ISBN ???? LCCN ????

Anonymous:1995:IRI

[Ano95r] Anonymous. In-depth review: Interleaf’s Cyberleaf converts word-processing documents to Web pages, mapping document styles to their HTML equivalents. Open Systems Today, ??(168):38–??, February 1995. ISSN 1061-0839.

Anonymous:1995:IPT

[Ano95s] Anonymous. Internet publishing tools proliferate. BYTE Magazine, 20(3):30–??, March 1, 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:I

[Ano95t] Anonymous. Internet@Work.

LAN: the network solutions magazine, 3(11):88–??, May 1, 1995. CODEN LANNER. ISSN 1038-359X (or 1069-5621??).

Anonymous:1995:IW

[Ano95u] Anonymous. Internet@Work.

LAN: the network solutions magazine, 3(11):88–??, May 1, 1995. CODEN LANNER. ISSN 1038-359X (or 1069-5621??).

Anonymous:1995:KAR

[Ano95v] Anonymous, editor. Knowledge acquisition, representation and processing: KARP-95: International symposium; 2nd — September 1995, Auburn, AL, International Symposium on Knowledge Acquisition Representation and Processing 1995. ????, ????, 1995. ISBN ???? LCCN ????

Anonymous:1995:L

[Ano95w] Anonymous. Letters. BYTE Magazine, 20(5):17–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:LIS

[Ano95x] Anonymous, editor. Library and Information Services in Astronomy II (LISA-II), volume 39(pt.2) of Vistas in Astronomy. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1995. CODEN VASTA6. ISBN ???? ISSN 0083-6656. LCCN ????

Anonymous:1995:MO

[Ano95y] Anonymous. Macworld Online. Macworld, 12(10):47–??, October 1, 1995. CODEN MACWEA. ISSN 0741-8647.

Anonymous:1995:ML

[Ano95z] Anonymous. Microcomputing/ LANs. Government computer news, 14(5):37–??, March 1995. ISSN 0738-4300.

Anonymous:1995:MSG

[Ano95-27] Anonymous. MULTIMEDIA — Soho gets multimedia production facility. the VRML bandwagon. virtual warfare. so good they used it twice. silicon reduction. Audio Visual, ??(??):47– ??, September 1995. ISSN 0305-2249.

Anonymous:1995:NNB

[Ano95-28] Anonymous. The net’s next big thing: Virtual reality. BYTE Magazine, 20(4):28–??, April 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:IPN

[Ano95-29] Anonymous. On the Internet page: NaviSoft’s publishing solution for the Web integrates a browser and HyperText Markup Language editor into a single tool. ComputerWorld, 29(29): 57–??, 1995. CODEN CMP-WAB. ISSN 0010-4841.

Anonymous:1995:OP

[Ano95-30] Anonymous. Online publishing.

MacUser, 11(10):117–??, October 1, 1995. CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1995:OPF

[Ano95-31] Anonymous. Online publishing: format your text HTML to put yourself on the World Wide Web. MacUser, 11(10):177–??, ???? 1995. CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1995:PNG

[Ano95-32] Anonymous. PageMaker 6.0 — the next generation? Publishing, pages 22, 24, September 1995. CODEN PUBLB6.

Anonymous:1995:PIW

[Ano95-33] Anonymous, editor. Proceedings of Internet World International 1995. Learned Information, Medford, NJ, USA, ???? 1995. ISBN ???? LCCN ????

Anonymous:1995:SIW

[Ano95-34] Anonymous, editor. Second International World-Wide Web Conference: Mosaic and the Web, Chicago, IL, October 17– 20, 1994. National Center for Supercomputer Applications, University of Illinois at Urbana-Champaign, Urbana, IL 61801, USA, 1995. ISBN ???? LCCN ???? A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Anonymous:1995:SHW

[Ano95-35] Anonymous, editor. SGML and HTML: what are they, what can you do with them? Seminar: Spring meeting — 1995, volume 126 of NEWSIDIC 1995. European Association of Information Services (EUSIDIC), ????, 1995. ISBN ???? ISSN 0141-6243. LCCN ????

Anonymous:1995:SG

[Ano95-36] Anonymous. SGML guide. Information Retrieval and Library Automation, 31(6):11–??, 1995. CODEN IRLAAQ. ISSN 0020-0220.

Anonymous:1995:STC

[Ano95-37] Anonymous, editor. Society for Technical Communication: 1995 proceedings / 42nd Annual Conference, April 23–26, 1995, Sheraton Washington Hotel, Washington, DC, volume 42. Society for Technical Communication, Arlington, VA, USA, 1995. ISBN 0-914548-86-7, 0-914548-87-5. LCCN T 11 S63 1995.

Anonymous:1995:SS

[Ano95-38] Anonymous. Software suppliers. Publishing and production executive, 9(12):77–??, December 1995. CODEN PPEXEP. ISSN 1048-3055.

Anonymous:1995:SHE

[Ano95-39] Anonymous. SOLUTIONS:

HTML editors are evolving into more familiar WYSIWYG editors. MacWEEK, 9(33):53–??, ???? 1995. CODEN MWEEEI. ISSN 0892-8118.

Anonymous:1995:TYWa

[Ano95-40] Anonymous. Teach yourself Web publishing with HTML in a week. Managing Information, 2(6):40–??, 1995. CODEN MAIOEG. ISSN 1352-0229.

Anonymous:1995:TYWb

[Ano95-41] Anonymous. Teach yourself Web publishing with HTML in a week. Whole Earth Review, ?? (86):86–??, Summer 1995. ISSN 0749-5056.

Anonymous:1995:UHM

[Ano95-42] Anonymous. Using HyperText Markup Language to create pages on the web. Network World, 12(13):53–??, March 1995. ISSN 0887-7661 (print), 1944-7655 (electronic).

Anonymous:1995:VRM

[Ano95-43] Anonymous. The virtual reality markup language, which enables 3-D network gaming and viewing across a network, moves closer to standardization. Electronic engineering times, ?? (843):4–??, April 1995. ISSN 0192-1541.

Anonymous:1995:WPM

[Ano95-44] Anonymous. Web publishing made easier — A review of Cyberleaf, HoTMetaL Pro, HTML Assistant Pro, and Spider. BYTE Magazine, 20(12): 170–??, ???? 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:W

[Ano95-45] Anonymous. WordPerfect 3.5. MacUser, 11(12):78–??, December 1, 1995. CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1995:WTB

[Ano95-46] Anonymous. WordPerfect 3.5 to take on bookmarks and HTML. MacWEEK, 9(29):6–??, ???? 1995. CODEN MWEEEI. ISSN 0892-8118.

Anonymous:1996:ICI

[Ano96a] Anonymous, editor. 12th International Conference on Interactive Information and Processing Systems (IIPS) for Meteorology, Oceanography, and Hydrology, January 28–February 2, 1996, Atlanta, Georgia. American Meteorological Society, Boston, MA, USA, 1996. ISBN ???? LCCN QC 851 I57 1996.

Anonymous:1996:PAC

[Ano96b] Anonymous, editor. 1996 proceedings / 43rd Annual Conference, Society for Technical Communication, May 5–9, 1996, Washington State Convention and Trade Center, Seattle, WA, volume 43 of Annual Conference — Society for Technical Communication 1996. Society for Technical Communication, Arlington, VA, USA, 1996. ISBN 0-914548-89-1, 0-914548-90-5. LCCN ????

Anonymous:1996:AMS

[Ano96c] Anonymous, editor. 25th Annual Meeting of the Society for Computers in Psychology (SCiP), volume 28(2) of Behavior research methods, instruments, and computers. Psychonomic Soc, ????, May 1996. CO-DEN BRMCEW. ISBN ???? ISSN 0743-3808 (print), 1532-5970 (electronic). LCCN ????

Anonymous:1996:ASA

[Ano96d] Anonymous, editor. ADC’96.

Seventh Australasian Database Conference, volume 18(2) of Australian Computer Science Communications. James Cook Univ, 1996. CODEN ACSCDD. ISBN ???? ISSN 0157-3055. LCCN ????

Anonymous:1996:AG

[Ano96e] Anonymous. The ant and the grasshopper. MacUser, 12(4): 91–??, April 1, 1996. CODEN MCUSEY. ISSN 0884-0997.

Anonymous:1996:ALL

[Ano96f] Anonymous, editor. Association for Literary and Linguistic Computing and Association for Computers and the Humanities: Joint international conference — June 1996, Bergen, Norway, number 70 in Report Series — Norwegian Computing Centre for the Humanities 1996. Norwegian Computing Centre for the Humanities, Bergen, Norway, 1996. ISBN 82-7283-083-3. ISSN 0804-8800. LCCN ????

Anonymous:1996:BNB

[Ano96g] Anonymous. Backstage nerves [Backstage Designer Plus]. Publishing, ??(??):21–??, August 1996. CODEN PUBLFA.

Anonymous:1996:BBB

[Ano96h]  Anonymous.  The browser

battle between Microsoft and Netscape. PC Magazine, 15(10): 40–??, May 28, 1996. CODEN PCMGEP. ISSN 0888-8507.

Anonymous:1996:CCC

[Ano96i] Anonymous. CAD/CAM connection — what does VRML have to offer design engineers? Computer Graphics World, 19 (12):22–??, December 1996. CO-DEN CGWODH. ISSN 0271-4159.

Anonymous:1996:CRS

[Ano96j] Anonymous. Choosing the right software for setting up a Web site. PC User, 20 March-2 April 1996. CODEN PCUSDW. ISSN 0263-5720.

Anonymous:1996:CFT

[Ano96k] Anonymous, editor. Communication on the Fast Track. IPCC 96 Proceedings: Saratoga Springs, New York, September 18–20, 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3689-5 (softbound), 0-7803-3690-9 (casebound), 0-7803-3691-7 (microfiche). LCCN T10.5.I54 1996. IEEE catalog number 96CH36009.

Anonymous:1996:CBH

[Ano96l] Anonymous. Computer books — HTML, VRML, CU-SeeMe keep the Internet pot boiling. Publishers weekly, 243(2):38–??, ???? 1996. ISSN 0000-0019.

Anonymous:1996:DDW

[Ano96m] Anonymous. Delivering documents on the World Wide Web. Software World, 27(4): 8–12, July 1996. CODEN SOFWBG. ISSN 0038-0652.

Anonymous:1996:DLI

[Ano96n] Anonymous, editor. Digital libraries and information services for the 21st century: International conference — September 1996, Seoul. KOLISS, ????, 1996. ISBN 89-87028-00-3. LCCN ????

Anonymous:1996:HP

[Ano96o] Anonymous. HotDog Professional 2.0. PC/Computing, 9(4): 198–??, April 1, 1996. CODEN PCMPEI. ISSN 0899-1847.

Anonymous:1996:HJW

[Ano96p] Anonymous. HTML and JAVA — what’s the difference? Library Systems, 16(2):15–??, ???? 1996. ISSN 0277-0288.

Anonymous:1996:HPI

[Ano96q] Anonymous. HTML Publishing on the Internet for Macintosh. Ventana Communications Group, Incorporated, Research Triangle Park, NC, USA, second edition, May 1996. ISBN 1-56604-419-7. ???? pp. LCCN ????

Anonymous:1996:HVC

[Ano96r] Anonymous. HTML, VRML, CU-SeeMe keep the Internet pot boiling. Publishers weekly, 243(2):38–??, January 8, 1996. ISSN 0000-0019.

Anonymous:1996:HCT

[Ano96s] Anonymous, editor. Human centered technology, key to the future: proceedings of the Human Factors and Ergonomics Society 40th annual meeting, Philadelphia, 1996, volume 2. Human Factors and Ergonomics Society, Santa Monica, CA, USA, 1996. ISBN 0-945289-06-5. ISSN 1071-1813. LCCN TA166.A2 H8 1996. Two volumes.

Anonymous:1996:ICP

[Ano96t] Anonymous, editor. IDT 96: carrefour de l’information professionnelle, textes des communications. … ????, Paris, France, 1996. ISBN 2-901046-95-9. LCCN ????

Anonymous:1996:ICL

[Ano96u] Anonymous, editor. IEE Colloquium on Learning at a Distance: Developments in Media Technologies (Digest No.1996/ 148). IEE, London, UK, 1996. ISBN ???? LCCN ????

Anonymous:1996:IFG

[Ano96v] Anonymous, editor. IMC forum de la GEIDE ’96: The imaging congress of Europe — June 1996, Paris, number 1–2 in IMC Forum de la GEIDE 1996. IMC, ????, 1996. ISBN ???? LCCN ????

Anonymous:1996:IHT

[Ano96w] Anonymous. InfoAccess HTML Transit vs. SkiSoft Web Publisher Professional. PC/ Computing, 9(5):64–??, May 1, 1996. CODEN PCMPEI. ISSN 0899-1847.

Anonymous:1996:INM

[Ano96x] Anonymous, editor. Institute of Nuclear Materials Management: Annual meeting; 37th — July 1996, Naples, FL, volume 25 of Nuclear Materials Management -Annual Meeting- 1996. Institute of Nuclear Materials Management, ????, 1996. ISBN ???? LCCN ????

Anonymous:1996:IST

[Ano96y]  Anonymous.  Internet sup-

port tops list of OS/400 enhancements for 1996. Computer Economics Report (International Edition), 18(4):4, April 1996. ISSN 0739-0874.

Anonymous:1996:IHH

[Ano96z]  Anonymous.  It’s Hytime

— hohe Zeit für Wasserstoff. Chemie in Labor und Biotechnik: CLB, 47(3):124–??, ???? 1996. ISSN 0722-6764.

Anonymous:1996:MAV

[Ano96-27] Anonymous. Microsoft Active VRML, Sony CyberPassage, Paper Software WebFX — browse the Web in another dimension. PC/Computing, 9(4):82– ??, ???? 1996. CODEN PCM-PEI. ISSN 0899-1847.

Anonymous:1996:NLH

[Ano96-28] Anonymous. A new look for HTML. Online & CDROM review: the international journal of online & optical information systems, 20(2):97–??, April 1, 1996. CODEN ONCDEW. ISSN 0309-314X.

Anonymous:1996:NFS

[Ano96-29] Anonymous. News — Ford and SDRC shake for USD 200 million. the M/cad Expo. CAD++ covers VRML. linking software to the shop floor. one solid leap for CAD. don’t hold back. image of the month. CAE poll. Computer-aided engineering: CAE, 15(2):12–??, ???? 1996. CODEN CCAEDJ. ISSN 0733-3536.

Anonymous:1996:NNM

[Ano96-30] Anonymous. NEWSTRENDS — no more upgrades, please. Internet/VB Script: A tough sell. look out, PC; here comes the NC. reality check for VRML programs. communications: The focal point. Microsoft’s target: The Internet. from NT to UNIX?. notes from the lab. and more. Windows Magazine, 7(5):32–??, ????

1996. CODEN WINMEV. ISSN 1060-1066.

Anonymous:1996:PPF

[Ano96-31] Anonymous, editor. PAAM 96. Proceedings of the First International Conference on the Practical Application of Intelligent Agents and Multi-Agent Technology. Practical Application Company, Blackpool, UK, 1996. ISBN ???? LCCN ????

Anonymous:1996:PDP

[Ano96-32] Anonymous. Portable documents put designer spin on the Web — Adobe’s and Common Ground’s retooled technologies port formatted information to the Web with ease. page-on-demand capabilities and better integration make them viable alternatives to HTML. Computer Shopper, ??(??):613–??, April 1996. ISSN 0886-0556.

Anonymous:1996:PIC

[Ano96-33] Anonymous, editor. Proceedings of 5th International Conference on Multi-lingual Computing (The New World Order and Information-Security and Welfare of Nations): Cambridge, UK, 11–13 April 1996. Cambridge University, Cambridge, UK, 1996. ISBN ???? LCCN ????

Anonymous:1996:PFIa

[Ano96-34] Anonymous, editor. Proceedings of the Fifth International World Wide Web Conference, 6-10 May 1996, Paris, France, volume 28(7-11) of Computer Networks and ISDN Systems. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, May 1996. CODEN CNISE9. ISBN ???? ISSN 0169-7552 (print), 1879-2324 (electronic). LCCN TK 5105.5 C647 v.28 no.7-11 1996.

Anonymous:1996:PFIb

[Ano96-35] Anonymous, editor. Proceedings of The First International Conference on Visual Information Systems. Victoria Univ. Technol, Melbourne, Vic., Australia, 1996. ISBN ???? LCCN ????

Anonymous:1996:PTA

[Ano96-36] Anonymous, editor. Proceedings of the Twenty-First Annual SAS Users Group International Conference, SUGI 21. SAS Inst, Cary, NC, USA, 1996. ISBN ???? LCCN ???? Two volumes. Anonymous:1996:RVD

[Ano96-37] Anonymous. Representative of VRML data in an objectoriented databases. In IEEE [IEE96c], pages 91–96. ISBN ???? LCCN ????

Anonymous:1996:SET

[Ano96-38] Anonymous. SGML in education: The TEI and ICADD initiatives. Computers in Libraries, 16(3):26–28, March 1, 1996. CO-DEN CPLIE8. ISSN 1041-7915. Anonymous:1996:SCP

[Ano96-39] Anonymous, editor. Society for Computers in Psychology: Annual meeting; 25th — November 1995, Los Angeles, CA, volume 28(2) of Behavior Research Methods Instruments and Computers 1996. Psychonomic Society, ????, 1996. ISBN ???? ISSN 0743-3808 (print), 1532-5970 (electronic). LCCN ????

Anonymous:1996:TIW

[Ano96-40] Anonymous, editor. Textile Institute: World conferences; 77th — May 1996, Tampere, Finland. Institute of Fiber, Textile and Clothing Science, Tampere University of Technology, Tampere, Finland, 1996. ISBN 951-722-545-8. LCCN ????

Anonymous:1996:RID

[Ano96-41] Anonymous. To the rescuemiddleware for Intranet developers. Software Economics Letter, 5(8):4–6, August 1996. CO-DEN SECLE3. ISSN 1065-6146.

Anonymous:1996:TTN

[Ano96-42] Anonymous. Trends and technology — with a new spec, Microsoft wants to make Windows 95 your’ communications center, though analysts say it’s all marketing hoopla; the corporate world braces for the arrival of multimedia; Virtual Reality Modeling Language (VRML) brings a new twist to cyberspace-3-D data. Computer Shopper, ??(??):83–??, April 1996. ISSN 0886-0556.

Anonymous:1996:TNA

[Ano96-43] Anonymous. Two new approaches to online information [Krakatoa and intranets]. Online & CDROM review: the international journal of online & optical information systems, 20(4): 206–207, August 1996. CODEN ONCDEW. ISSN 0309-314X.

Anonymous:1996:VGR

[Ano96-44] Anonymous. VRML Group releases VRML 2.0 spec: forms VRML Consortium. Silicon Graphics World, 6(10):7, October 1996. ISSN 1057-7041. See

                                        http://vag.vrml.org for de-

tails.

Anonymous:1996:WAU

[Ano96-45] Anonymous. Web authors use Cosmo site for HTML, Java and VRML sites. Silicon Graphics World, 6(11):25, November 1996. ISSN 1057-7041.

Anonymous:1996:WWW

[Ano96-46] Anonymous, editor. World Wide Web Journal: Volume 1, Issue 3: The Web After Five Years. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-210-7. ISSN 1085-2301. LCCN ???? US$24.95.

Anonymous:1997:AAA

[Ano97a]  Anonymous, editor.  1997

ACSM/ASPRS Annual Convention & Exposition: technical papers: ACSM 57th Annual Convention, ASPRS 63rd Annual Convention, Seattle, Washington, April 7–10, 1997, volume 2 of ACSM ASPRS Annual Convention and Exposition. American Congress on Surveying and Mapping and American Society for Photogrammetry and Remote Sensing, Bethesda, MD, USA, 1997. ISBN 1-57083-043-6 (set), 1-57083-044-4 (vol. 1), 1-57083-045-2 (vol. 2), 1-57083-047-9 (vol. 4), 1-57083-048-7 (vol. 5). LCCN ???? Five volumes.

Anonymous:1997:PAC

[Ano97b] Anonymous, editor. 1997 proceedings: 44th Annual Conference, May 11–14, 1997, Sheraton Centre Toronto, Toronto, Canada, volume 44. Society for Technical Communication, Arlington, VA, USA, 1997. ISBN 0-914548-94-8, 0-914548-95-6. LCCN ????

Anonymous:1997:ACS

[Ano97c] Anonymous. Ajaib — A case study of an SGML/ intranet development: (getting documentation off the shelf). In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 269–272. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:AGR

[Ano97d] Anonymous, editor. Applied geologic remote sensing: practical solutions for real-world problems: proceedings of the twelfth international conference, 17–19 November 1997, Denver, Colorado, USA, volume 2. ERIM International, Inc., Ann Arbor, MI, USA, 1997. ISBN ????

ISSN 1067-0106. LCCN QE 33.2 R4 A67 1997. Two volumes.

Anonymous:1997:BBS

[Ano97e] Anonymous. Business benefits of an SGML and STEP integration: A drama in one act. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 69–72. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:CSS

[Ano97f] Anonymous. “case study”, the SGML (standard generalized markup language) implementation at norsk hydro. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 149–152. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:CSA

[Ano97g] Anonymous. CD 13250: SGML applications — topic navigation maps. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 263–268. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:CDP

[Ano97h] Anonymous, editor. Competitive deployment of product data technology: European conference — April 1997, Antibes, France. Quality Marketing Services, ????, 1997. ISBN 1-901782-01-8. LCCN ????

Anonymous:1997:CJY

[Ano97i] Anonymous. Cost justifying your SGML project. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 61–68. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:CSF

[Ano97j] Anonymous. Croner and SGML — the first 3 years: opening the envelope! In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 165–168. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:DRD

[Ano97k] Anonymous. Defining reusable, distributable information objects using or, how SGML can do for databases what JAVA has done for user interfaces. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 307–310. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:ESE

[Ano97l] Anonymous. Extranet SGML editorial system for encyclopaedias. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 163–164. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:GXH

[Ano97m] Anonymous. Getting to XML from HTML. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 189–194. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:HCK

[Ano97n] Anonymous. HTML construction kit, 1997. ISBN 1-57521-318-4. Includes CD-ROM.

Anonymous:1997:HPB

[Ano97o] Anonymous. HTML primer: a beginner’s guide to HTML; and, HTML 2.0 technical specifications, volume ?? Arizona Testing Center, Administrative Office of the United States Courts, Phoenix, AZ, USA, 1997. vari-

                                        ous pp. URL http://156.131.
                                        8.232/techpub.html.

Anonymous:1997:HFE

[Ano97p] Anonymous. Human factors engineering: Creating a productive environment for authoring SGML documents. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 121–126. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:IAS

[Ano97q] Anonymous. Improving the accessibility of SGML documents: A content-analytical approach. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 321–332. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:ISR

[Ano97r] Anonymous. Introducing SGML into the RAF flight manuals world or throttle to bottle in two extraordinary years. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 83–90. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:ISD

[Ano97s] Anonymous. Intuitive SGML:

database integration in SGML authoring. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 119–120. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:KIS

[Ano97t] Anonymous. Keep it simple — interactive electronic applications with SGML. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 273–276. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:MSE

[Ano97u] Anonymous. Making SGML easier with microdocument databases. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 319–320. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:NEC

[Ano97v] Anonymous, editor. National educational computing conference: — June 1997, Seattle, WA, Proceedings of the National Educational Computing Conference 1997. International Society for Technology in Education, Eugene, OR, USA, 1997. ISBN 1-56484-112-X. LCCN ????

Anonymous:1997:NPC

[Ano97w] Anonymous. New products: Cyclone 6 UART Serial Board; Visual SlickEdit Version 2.0 for the X Window System; Foot-Prints, Web-Based Helpdesk; MagiCaster Multimedia Interface; WebTech C++ Library Supporting HTML 3.2; Inter-Cam Telerobotic System; Open-Path Web 2.0; NetTerminal; VirtuFlex. Linux Journal, 35: ??, March 1997. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1997:PSI

[Ano97x] Anonymous, editor. Papers from the Sixth International World Wide Web Conference, 7–11 April 1997, Santa Clara, CA, USA, volume 29(8–13) of Computer Networks and ISDN Systems. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1997. ISBN ???? ISSN 0169-7552 (print), 1879-2324 (electronic). LCCN TK 5105.5 C647 v.29 no.8-13 1997.

Anonymous:1997:PDF

[Ano97y] Anonymous, editor. Proceedings of the DLM-Forum on electronic records, Brussels, 18–20 December 1996. DLM-Forum. Luxembourg: Office for Official Publications of the European Communities, Luxembourg, Luxembourg, 1997. ISBN 92-828-0111-X. LCCN CD974.4 .D46 1996.

Anonymous:1997:PFN

[Ano97z] Anonymous, editor. Proceedings of the Fourteenth National Conference on Artificial Intelligence and the Ninth Innovative Applications of Artificial Intelligence conference, July 1997, Providence, RI. AAAI Press, Menlo Park, CA, USA, 1997. ISBN 0-262-51095-2. LCCN Q334 .N37 1997.

Anonymous:1997:ICC

[Ano97-27] Anonymous, editor. Proceedings of the third international symposium on command and control research and technology symposium: partners for the 21st century, Washington, DC, June 17–20, 1997. National Defense University, Washington, DC, USA, 1997. ISBN ???? LCCN ????

Anonymous:1997:PHSb

[Ano97-28] Anonymous. Programming in HTML standard 3.2. intermediate, 1997. 1 videocassette.

Anonymous:1997:PHSa

[Ano97-29] Anonymous. Programming in HTML standard 3.2. introduction, 1997. 1 videocassette.

Anonymous:1997:PWM

[Ano97-30] Anonymous. Publishing to the Web is more than converting data into HTML. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 181–184. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:SUG

[Ano97-31] Anonymous, editor. SAS Users Group: Annual international conference; 22nd — March 1997, San Diego, CA, SUGI 1997; Conf 22. SAS, ????, 1997. ISBN 1-55544-961-1. LCCN ????

Anonymous:1997:SMS

[Ano97-32] Anonymous. SGML — made SIMPLE. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 127–132. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:SDD

[Ano97-33] Anonymous. SGML and development documentation. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 159– 162. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:SAI

[Ano97-34] Anonymous. SGML and the auto industry: Contrasting east-west management strategies. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 19–20. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:SOL

[Ano97-35] Anonymous. SGML and the on-line legislature. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 169–170. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:SCM

[Ano97-36] Anonymous. The SGML character model. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 245–250. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:SEP

[Ano97-37] Anonymous, editor. SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain. Inso Corporation, Boston, MA, USA, 1997. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:SHI

[Ano97-38] Anonymous. SGML in healthcare information systems. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 195– 204 (or 205–210??). ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:ST

[Ano97-39] Anonymous. SGML in transition. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 297–302 (or 303–306??). ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:SSG

[Ano97-40] Anonymous. SGML (standard generalized markup language), still a cutting-edge technology? In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 21–26. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:STD

[Ano97-41] Anonymous. SGML template driven database extraction: A new approach to report generation. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 311–318. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:SPS

[Ano97-42] Anonymous, editor. SICE ’97: proceedings of the 36th SICE Annual Conference: International Session papers: University of Tokushima, July 29– 31, 1997. Society of Instrument and Control Engineers, Tokyo, Japan, 1997. ISBN 0-7803-4208-9. LCCN TK7870.K37 1997. SICE catalog number 97 PR 0001-3. IEEE catalog number 97 TH 8323.

Anonymous:1997:SCC

[Ano97-43] Anonymous, editor. Small college computing: Annual symposium; 30th — April 1997, Kenosha, WI, SCCS - Proceedings- 1997; 30th. SCCS, ????, 1997. ISBN ???? LCCN ????

Anonymous:1997:SSS

[Ano97-44] Anonymous. STEP/SGML standards working together. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 39–42. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:SSD

[Ano97-45] Anonymous. Supporting SGML in document management systems. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 95–106. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:TGL

[Ano97-46] Anonymous. Tastes great — less filling: SGML for the 21st century. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 333–334. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:TYH

[Ano97-47] Anonymous. Teach yourself HTML 3.2. intermediate, 1997. 1 videocassette (60 min.).

Anonymous:1997:TUI

[Ano97-48] Anonymous, editor. Technische Universitat Ilmenau: Internationales wissenschaftliches kolloquium; 42 — September 1997, Ilmenau, Germany, volume 42(1) of Internationales Wissenschaftliches Kolloquium-Technische Universitat Ilmenau 1997. Technische Universitat Ilmenau, ????, 1997. ISBN ???? ISSN 0943-7207. LCCN ????

Anonymous:1997:TDS

[Ano97-49] Anonymous. Technology driving the SGML marketplace driving technology. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 295–296. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:TFC

[Ano97-50] Anonymous, editor. Teleconferencing futures: Colloquium — June 1997, London, number 121 in Colloquium Digest — IEE 1997. IEE, London, UK, 1997. ISBN ???? ISSN 0963-3308. LCCN ????

Anonymous:1997:WAH

[Ano97-51] Anonymous. THE WEB authoring: HTML authoring, Front-Page 1.1, HoTMetaL PRO 3.0, Java, Symantec Café. Windows Magazine, ??(??):147–??, ???? 1997. CODEN WINMEV. ISSN 1060-1066.

Anonymous:1997:TMS

[Ano97-52] Anonymous. A topic map for SGML 97 proceedings: A new SGML animal. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 335–338. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:TSD

[Ano97-53] Anonymous. Towards an SGML diff using temporal documents. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 251–262. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:WYD

[Ano97-54] Anonymous. Why you do (or don’t) need HyTime in your document management system. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 211–216. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:WTT

[Ano97-55] Anonymous. Wishful thinking or thinking ahead? envisioning the next generation of SGML editors. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 107–114. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1997:XLD

[Ano97-56] Anonymous. XML and legacy data conversion: Introducing “consumable documents”. In SGML Europe ’97 proceedings 11–16 May 1997, Princess Sofia Intercontinental, Barcelona, Spain [Ano97-37], pages 185– 188. ISBN ???? LCCN ???? One CD-ROM.

Anonymous:1998:ANE

[Ano98a] Anonymous, editor. The 19th Annual National Educational Computing Conference: San Diego Convention Center, San Diego, CA, June 22–24, 1998: NECC ’98. International Society for Technology in Education, Eugene, OR, USA, 1998. ISBN 1-56484-134-0. LCCN ????

Anonymous:1998:AAS

[Ano98b] Anonymous, editor. American Astronomical Society: Meeting; 193rd — January 1999, Austin, TX, volume 30(4) of Bulletin — American Astronomical Society 1998. American Astronomical Society, Washington, DC, USA, 1998. ISBN ???? ISSN 0002-7537. LCCN ????

Anonymous:1998:AMI

[Ano98c] Anonymous, editor. American Medical Informatics Association: Bringing knowledge to the point of use: Spring congress — May 1998, Philadelphia, PA. ????, American Medical Informatics Associations, 1998. ISBN ???? LCCN ????

Anonymous:1998:ASH

[Ano98d] Anonymous, editor. American Society for Horticultural Science: Annual international conference; 95th — July 1998, Charlotte, NC, volume 33(3) of HORTSCIENCE 1998. American Society for Horticultural Science, ????, 1998. ISBN ???? ISSN 0018-5345. LCCN ????

Anonymous:1998:BHD

[Ano98e] Anonymous. Beyond HTML dynamic and interactive applications for libraries, July 1, 1998. Two sound cassettes.

Anonymous:1998:BRP

[Ano98f] Anonymous. Book review: A Practical Guide to Linux by Todd Sundsted; HTML: The Definitive Guide, Second Edition by Eric S. Raymond. Linux Journal, 49:??, May 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1998:CPa

[Ano98g] Anonymous. Color plate. In Spencer and Roehl [SR98], page 131. ISBN 1-58113-022-8. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/271897/.

Anonymous:1998:CPc

[Ano98h] Anonymous. Color plate. In Spencer and Roehl [SR98], page 134. ISBN 1-58113-022-8. LCCN ???? URL http://

                                        www.acm.org/pubs/contents/
                                        proceedings/graph/271897/.

Anonymous:1998:CPe

[Ano98i] Anonymous. Color plate. In Spencer and Roehl [SR98], page 138. ISBN 1-58113-022-8. LCCN ???? URL http://

                                        www.acm.org/pubs/contents/
                                        proceedings/graph/271897/.

Anonymous:1998:CPb

[Ano98j]  Anonymous.  Color plates.

In Spencer and Roehl [SR98], pages 132–133. ISBN 1-58113-022-8. LCCN ????

                                        URL http://www.acm.org/
                                        pubs/contents/proceedings/
                                        graph/271897/.

Anonymous:1998:CPd

[Ano98k]  Anonymous.  Color plates.

In Spencer and Roehl [SR98], pages 135–136. ISBN 1-58113-022-8. LCCN ????

                                        URL http://www.acm.org/
                                        pubs/contents/proceedings/
                                        graph/271897/.

Anonymous:1998:CPS

[Ano98l]  Anonymous, editor.  Com-

puter publishing: Symposium — November 1998, Dusseldorf, Germany, Computer Publishing -Symposium- 1998. CSM, Munchen, Germany, 1998. ISBN ???? LCCN ????

Anonymous:1998:EAC

[Ano98m] Anonymous, editor. EURO-GRAPHICS ’98: the European Association for Computer Graphics, 19th annual Conference, Lisboa, Portugal, August 31–September 4, 1998, volume 17(3) of Computer Graphics Forum. Blackwell, Oxford, UK, 1998. ISBN ???? ISSN 0167-7055 (print), 1467-8659 (electronic). LCCN ????

Anonymous:1998:GAL

[Ano98n] Anonymous, editor. Genetics applied to livestock production: World congress; 6th — January 1998, Armidale, Australia, volume 27 of Proceedings of the World Congress on Genetics Applied to Livestock Production 1998; 6th. ????, ????, 1998. ISBN 1-86389-454-3. LCCN ????

Anonymous:1998:HIM

[Ano98o] Anonymous, editor. Healthcare Information and Management Systems Society: Annual conference — February 1998, Orlando, FL, volume 3 of Proceedings of the Annual HIMSS Conference 1998. HIMSS, ????, 1998. ISBN ???? ISSN 1048-8987. LCCN ????

Anonymous:1998:LGW

[Ano98p] Anonymous. Linux Gazette:

Writing HTML with m4. Linux Journal, 47:??, March 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

URL ftp://ftp.ssc.com/pub/
lj/listings/issue47/2393.
tgz.

Anonymous:1998:PFN

[Ano98q] Anonymous, editor. Proceedings / Fifteenth National Conference on Artificial Intelligence (AAAI-98): Tenth Conference on Innovative Applications of Artificial Intelligence (IAAI-98), July 26–30, 1998, Madison, Wisconsin. AAAI Press, Menlo Park, CA, USA, 1998. ISBN 0-262-51098-7. LCCN Q334 .N36 1998.

Anonymous:1998:PHF

[Ano98r] Anonymous, editor. Proceedings of the Human Factors and Ergonomics Society: 42nd annual meeting, October 5-9, 1998, Chicago, Illinois: Humansystem interaction: the sky’s no limit, volume 2 of Proceedings of the Human Factors and Ergonomics Society Annual Meeting 1999. Human Factors and Ergonomics Society, Santa Monica, CA, USA, 1998. ISBN 0-945289-11-1. ISSN 1071-1813. LCCN TA166.A2 .H8 1998. Two volumes.

Anonymous:1998:PSI

[Ano98s] Anonymous, editor. Proceedings of the Seventh International World Wide Web Conference: 14–18 April 1998, Brisbane, Australia, volume 30(1) of Computer Networks and ISDN Systems. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1998. ISBN ???? ISSN 0169-7552 (print), 1879-2324 (electronic). LCCN ????

Anonymous:1998:SUG

[Ano98t] Anonymous, editor. SAS Users Group: Annual international conference; 23rd — March 1998, Nashville, TN, SUGI 1998; Conf 23. SAS, ????, 1998. ISBN 1-58025-149-8. LCCN ????

Anonymous:1998:STC

[Ano98u] Anonymous, editor. Society for Technical Communication: 1998 proceedings: 45th Annual Conference, May 17–20th, 1998, Anaheim, California, volume 45. Society for Technical Communication, Arlington, VA, USA, 1998. ISBN 1-891709-00-3, 1-891709-01-1. LCCN ????

Anonymous:1999:ICI

[Ano99a] Anonymous, editor. 15th International Conference on Interactive Information and Processing Systems (IIPS) for Meteorology, Oceanography, and Hydrology: 10–15 January 1999, Dallas, Texas. American Meteorological Society, Boston, MA, USA, 1999. ISBN ???? LCCN ????

Anonymous:1999:ADR

[Ano99b] Anonymous, editor. Air distribution in rooms: Proceedings of ROOMVENT ’98: International conference; 6th — 1998: Stockholm, number 95 in Aalborg Universitet Department of Building Technology and Structural Engineering -Report- R 1998. Aalborg University, Aalborg, Denmark, 1999. ISBN ???? ISSN 1395-7953. LCCN ????

Anonymous:1999:ATC

[Ano99c] Anonymous. Annotated table of contents: XML for dummies; the XML & SGML cookbook: Recipes for structured information; XML: A primer. Markup Languages: Theory & Practice, 1(2):95–104, Spring 1999. CO-DEN MLTPFG. ISSN 1099-6621.

Anonymous:1999:APC

[Ano99d] Anonymous, editor. Applications of the performance concept in building: International symposium; 3rd — December 1996, Tel Aviv, Israel, volume 8(4) of Automation in Construction 1999. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1999. ISBN ???? ISSN 0926-5805. LCCN ????

Anonymous:1999:ALL

[Ano99e] Anonymous, editor. Association for Literary and Linguistic Computing—Association for Computers and the Humanities: Conference — 1998: Debrecen, Hungary, volume 14(1) of Literary and Linguistic Computing. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN ???? ISSN 0268-1145. LCCN ????

Anonymous:1999:CPO

[Ano99f]  Anonymous, editor.  Com-

puter processing of oriental languages: International conference; 18th — March 1999, Tokushima City, Japan, volume 1 of International Conference on Computer Processing of Oriental Languages 1999. Oriental Languages Computer Society, ????, 1999. ISBN ???? LCCN ????

Anonymous:1999:DHR

[Ano99g] Anonymous. Dynamic HTML reference and software development kit. Microsoft Press, Bellevue, WA, USA, 1999. ISBN 0-7356-0638-2. viii + 1366 pp. LCCN QA76.76.H94D96 1999.

Anonymous:1999:GIT

[Ano99h] Anonymous, editor. Geospatial Information and Technology Association: Conference; 22nd — April 1999, Charlotte, NC, GITA -Proceedings-1999; 22nd. Geospatial Information and Technology Association, ????, 1999. ISBN ???? LCCN ????

Anonymous:1999:PEI

[Ano99i] Anonymous, editor. Proceedings of the eighth international World Wide Web conference, Toronto, Canada, 11–14 May 1999, volume 31(11). Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1999. ISBN 0-444-50264-5. ISSN 1389-1286 (print), 1872-7069 (electronic). LCCN TK5105.888 .I573 1999.

Anonymous:1999:SXA

[Ano99j] Anonymous. SGML und XML:

Anwendungen und Perspektiven. (German) [SGML and XML:

Applications and Perspektives]. Springer-Verlag Inc., New York, NY, USA, 1999. ISBN 3-540-65543-3. 380 (est.) pp. LCCN ????

Anonymous:1999:WX

[Ano99k]  Anonymous.  Working with

XML, 1999.

Anonymous:1999:XHI

[Ano99l] Anonymous. XML how it will affect content, knowledge management and information professionals: enhancing searching and content management: information and content exchange, 1999.

Anonymous:19xx:HPM

[Anoxx] Anonymous. HTML Publications for Mac W/D. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 19xx. ISBN 1-56830-242-8. ???? pp. LCCN ???? US$35.00.

Aoki:1999:CXH

[Aok99] Steven Aoki. Comparison of XML to HTML for Robert E. Kennedy Library’s “new books” database. Thesis (b.s.), California Polytechnic State University, San Luis Obispo, CA, USA, 1999. 34 pp.

Augenstein:1991:TRA

[AOS91] F. Augenstein, T. Ottmann, and J. Schoning. A type- and rulecontrolled authoring system. In Maurer [Mau91], pages 25–33. ISBN 3-540-54145-4. LCCN ????

APProfessionalStaff:1996:SH

[AP 96] AP Professional Staff. SGML Handbook. Academic Press, New York, USA, August 1996. ISBN 0-12-522450-8. ???? pp. LCCN ???? US$29.95.

Andrews:1996:HTM

[AP96] K. Andrews and M. Pichler.

Hooking up 3-space: threedimensional models as fullyfledged hypermedia documents. In Brusilovsky et al. [BKS96], pages 28–44. ISBN 3-540-61282-3. LCCN QA76.575 .M38 1994.

Abrams:1999:UAI

[APB^{+}99] Marc Abrams, Constantinos Phanouriou, Alan L. Batongbacal, Stephen M. Williams, and Jonathan E. Shuster. UIML: an appliance-independent XML user interface language. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16): 1695–1708, May 17, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.
com/cas/tree/store/comnet/
sub/1999/31/11-16/2170.pdf.

AppleComputer:1998:WPT

[APH98] Apple Computer, Inc, Pierian Spring Software, and Humanities Software. Writing and publishing tools, 1998. Includes seven CD-ROMs.

Appelt:1989:SDP

[App89] W. Appelt. Standards in document processing activities. Informatik Spektrum, 12(6):321– 330, December 1989. CO-DEN INSKDW. ISSN 0170-6012 (print), 1432-122X (electronic).

Andrews:1998:LIV

[APP^{+}98] K. Andrews, A. Pesendorfer, M. Pichler, Heinz, and K. Wagenbrunn. Looking inside VRwave: The architecture and interface of the VRwave VRML97 browser. In Spencer and Roehl [SR98], pages 77–82. ISBN 1-58113-022-8. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/271897/.

Allen:1997:ADL

[AR97] Robert B. Allen and Edie Rasmussen, editors. ACM digital libraries ’97: proceedings of the second ACM International Conference on Digital Libraries, Philadelphia, PA, July 23–26, 1997, volume 2. ACM Press, New York, NY, USA, 1997. ISBN 0-89791-868-1. LCCN Z 699 A1 A27 1997.

Araki:1998:VHM

[Ara98] Y. Araki. VSPLUS: A high-level multi-user extension library for interactive VRML worlds. In Spencer and Roehl [SR98], pages 39–48. ISBN 1-58113-022-8. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/271897/.

Armstrong:1996:HRH

[Arm96] James C. Armstrong. HTML Reference: Hyper Test Markup Language Reference. Specialized Systems Consultants, Incorporated, Seattle, third edition, October 1996. ISBN 0-916151-98-0. 16 pp. LCCN ???? US$4.50.

Arnett:1995:IAN

[Arn95] Nick Arnett. The Internet and the anti-net. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Overviews/arnett/Antinet.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Aronson:1994:HMS

[Aro94] Larry Aronson. HTML manual of style. Ziff-Davis Press, Emeryville, CA, USA, 1994. ISBN 1-56276-300-8. ix + 132 pp. LCCN QA76.76.H94 A76 1994. US$19.95, CDN$27.95.

Aronson:1996:HMS

[Aro96] Larry Aronson. HTML 3 manual of style. Ziff-Davis Press, Emeryville, CA, USA, 1996. ISBN 1-56276-352-0. xiv + 197 pp. LCCN QA 76.76 H94 A763 1995. US$24.95.

Arpajian:1996:HUHa

[Arp96] Scott Arpajian. How to Use HTML 3. Ziff-Davis Press, Emeryville, CA, USA, 1996. ISBN 1-56276-390-3. xi + 223 pp. LCCN QA76.76.H94 A76 1996. US$24.99, CDN$33.99, UK£22.99.

Abernathy:1993:APC

[AS93a] Roger W. Abernathy and Robert P. Smith. Algorithm 724: Program to calculate F -percentiles. ACM Transactions on Mathematical Software, 19(4):481–483, December 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/citations/journals/toms/
1993-19-4/p481-abernathy/.

Abernathy:1993:ASE

[AS93b] Roger W. Abernathy and Robert P. Smith. Applying series expansion to the inverse beta distribution to find percentiles of the F -distribution. ACM Transactions on Mathematical Software, 19(4):474– 480, December 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/citations/journals/toms/
1993-19-4/p474-abernathy/.

Alschuler:1994:SSI

[AS94a] Liora Alschuler and Stephen Schettini. Systems: SGML is here to stay. Publish!: the howto magazine of desktop publishing, 9(6):71–??, June 1, 1994. CODEN PBLIEJ. ISSN 0897-6007.

Anderson:1994:SMS

[AS94b] Jean T. Anderson and Michael Stonebraker. SEQUOIA 2000 metadata schema for satellite images. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(4):42– 48, December 1994. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Alper:1996:GMQ

[AS96] N. Alper and C. Stein. Geospatial metadata querying and visualization on the WWW using Java applets. In Card et al. [CEG96], page ?? ISBN 0-8186-7668-X. LCCN TA1634 .I325 1996.

Abernathy:1998:IGI

[AS98] M. Abernathy and S. Shaw. Integrating geographic information in VRML models. In Spencer and Roehl [SR98], pages 107–114. ISBN 1-58113-022-8. LCCN ????

URL http://www.acm.org/
pubs/contents/proceedings/
graph/271897/.

AAP:1986:MMF

[Ass86a] Association of American Publishers, Electronic Manuscript and Project. Markup of mathematical formulas: an SGML application conforming to International Standard ISO 8879, Standard Generalized Markup Language. Electronic manuscript series / Association of American Publishers. Association of American Publishers, Dublin, OH, USA, 1986. ISBN ???? 79 pp. LCCN Z 286 E43 M36 1989.

AAP:1986:MTM

[Ass86b] Association of American Publishers, Electronic Manuscript and Project. Markup of tabular material: an SGML application conforming to International Standard ISO 8879, Standard Generalized Markup Language. Electronic manuscript series / Association of American Publishers. Association of American Publishers, Dublin, OH, USA, 1986. ISBN 1-55653-085-4. 19 + 4 + 1 pp. LCCN Z 286 E43 M37 1989.

AAP:1987:AAP

[Ass87a] Association of American Publishers. Association of American Publishers electronic manuscript series standard for electronic manuscript preparation and markup: an SGML application conforming to International Standard ISO 8879–Standard Generalized Markup Language. Version 2.0 Dublin, Ohio. Association of American Publishers, Dublin, OH, USA, 1987. ISBN ???? various pp. LCCN Z 286 E43 A87 1987. Available from the Electronic Publishing Special Interest Group.

AAP:1987:MMF

[Ass87b] Association of American Publishers, Electronic Manuscript and Project. Markup of mathematical formulas: an SGML application conforming to International Standard ISO 8879, Standard Generalized Markup Language, Version 2.0. Electronic manuscript series / Association of American Publishers. Association of American Publishers, Dublin, OH, USA, 1987. ISBN ???? 80 pp. LCCN ????

AAP:1987:MTM

[Ass87c] Association of American Publishers, Electronic Manuscript and Project. Markup of tabular material: an SGML application conforming to International Standard ISO 8879, Standard Generalized Markup Language, Version 2.0. Electronic manuscript series / Association of American Publishers. Association of American Publishers, Dublin, OH, USA, 1987. ISBN ???? 19 + 4 + 1 pp. LCCN ????

AAP:1989:MTM

[Ass89] Association of American Publishers, Electronic Manuscript and Project. Markup of tabular material: an SGML application conforming to International Standard ISO 8879, Standard Generalized Markup Language, Version 2.0. Electronic manuscript series / Association of American Publishers. Association of American Publishers, Dublin, OH, USA, second edition, 1989. ISBN 1-55653-085-4. iii + 25 pp. LCCN Z 286 E43 M37 1989.

USENIX:1996:ATT

[Ass96] USENIX Association, editor.

4th Annual Tcl/Tk Workshop ’96, July 10–13, 1996. Monterey, CA. USENIX Association, Berkeley, CA, USA, July 10–13, 1996. ISBN 1-880446-78-2. LCCN QA76.73.T44 T44

Ames:1998:UXI

[ASWS98] T. A. Ames, K. B. Sall, C. E. Warsaw, and R. A. Shafer. Using XML for instrument description, communication and control of the SOFIA/HAWC instrument. In Anonymous [Ano98b], page 13.07. ISBN ???? ISSN 0002-7537. LCCN ????

Amsler:1988:SSE

[AT88] R. A. Amsler and F. W. Tompa. An SGML-based standard for English monolingual dictionaries. In Anonymous [Ano88g], pages 61–79. ISBN ???? LCCN ????

Addison:1995:HMI

[AT95] M. A. Addison and H. W. Thimbleby. Hypermedia manuals for interactive systems. In Anonymous [Ano95p], pages 5/1–5/4. ISBN ???? LCCN ????

Attar:1997:LWM

[Att97] Pierre Attar. L’edition sur le Web: méthodes d’utilisation de HTML. Documentation multimedia. Hermes, Paris, France, deuxième edition, 1997. ISBN 2-86601-565-7. 174 pp. LCCN ????

Augier:1999:HLD

[Aug99] Marc Augier. HTML le langage du Web. Que sais-je ?; 3420. Presses universitaires de France, Paris, France, 1ere edition, 1999. ISBN 2-13-049563-X. 127 pp. LCCN ????

Ause:1995:IHW

[Aus95] Wayne Ause. Instant HTML Web Pages. Ziff-Davis Press, Emeryville, CA, USA, October 1995. ISBN 1-56276-363-6. xv + 124 pp. LCCN QA76.76.H94A97 1995. US$24.95.

Amann:1998:BSD

[AVFC98] B. Amann, D. Vodislav, J. Fernandes, and G. Coste. Browsing SGML documents with maps: The French “Inventaire” experience. Lecture Notes in Computer Science, 1460:250–259, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aviram:1998:XDQ

[Avi98] Mariva H. Aviram. XML for dummies quick reference. I D G Books Worldwide, Indianapolis, IN, USA, 1998. ISBN 0-7645-0383-9 (paperback). xiv + 206 pp. LCCN QA76.76.H94A98 1998.

Alex:1998:TMW

[AVN98] J. Alex, B. Vikramaditya, and B. J. Nelson. Teleoperated micromanipulation within a VRML environment using Java. In IEEE [IEE98b], pages 1747–1752. ISBN 0-7803-4465-0 (softbound), 0-7803-4466-9 (casebound), 0-7803-4467-7 (microfiche). LCCN TJ210.3.I443 1998. Three volumes. IEEE Catalog Number 98CH36190.

Ayres:1993:DS

[AW93] N. Ayres and T. Wesley. Document standards-CAPS. In Anonymous [Ano93g], pages 5/1–3. ISBN ???? LCCN ????

Ahlberg:1995:IIV

[AW95] C. Ahlberg and E. Wistrand. IVEE: an Information Visualization and Exploration Environment. In Gershon and Eick [GE95], pages 66–73, 142– 143. ISBN 0-8186-7201-3 (paperback), 0-8186-7203-X (microfiche). LCCN TA1634.I325 1995. IEEE catalog no. 95TB100000.

Andresen:1996:SIH

[AYE^{+}96] D. Andresen, Tao Yang, O. Egecioglu, O. H. Ibarra, and T. R. Smith. Scalability issues for high performance digital libraries on the World Wide Web. In IEEE [IEE96m], pages 139–148. ISBN 0-8186-7402-4. LCCN ???? IEEE Catalog No. 96TB100053.

Azaria:1994:SLS

[Aza94] Adrienne Azaria. SGML: A lifesaver in a sea of electronic documents. Network World, 11(50):67, December 12, 1994. ISSN 0887-7661 (print), 1944-7655 (electronic).

Basu:1990:SBQ

[BA90] Amit Basu and Rafiul Ahad. An SQL-based query language for networks of relations. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(1):8–17, March 1990. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Brown:1997:HSS

[BA97] C. Brown and C. Aamodt.

HTML software scaling system. In Anonymous [Ano97-43], pages 96–101. ISBN ???? LCCN ????

Belaid:1993:QAL

[BAC93] A. Belaid, J. C. Anigbogu, and Y. Chenevoy. Qualitative analysis of low-level logical structures. Electronic Publishing—Origination, Dissemination, and Design, 6(4):435–446, December 1993. CODEN EPODEU. ISSN 0894-3982.

Bachrach:1996:PSE

[Bac96] S. M. Bachrach, editor. Proceedings of the second Electronic Computational Chemistry Conference: 1–30 November 1995, volume 368 of Journal of Molecular Structure. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1996. ISBN ???? ISSN 0166-1280 (print), 1872-7999 (electronic). LCCN ????

Bailey:1993:BED

[Bai93] N. Bailey. Building an electronic document management system. In Anonymous [Ano93f], pages 2–10. ISBN ???? LCCN ????

Balaguer:1996:VLE

[Bal96a] J.-F. Balaguer. VRML for LHC engineering. In Gobel et al. [GSV96], pages 159–168. ISBN 3-211-82886-9. ISSN 0946-2767. LCCN QA76.9.H85 V546 1996.

Baldazo:1996:BWS

[Bal96b] R. Baldazo. Battle of the Web site builders. BYTE Magazine, 21(7):157–158, July 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ball:1998:XST

[Bal98a] S. Ball. XML support for Tcl. In USENIX [USE98], pages 109–120. ISBN 1-880446-98-7. LCCN QA76.73.T44 T34 1998.

Ball:1998:XDT

[Bal98b] Steve Ball. XML and the desperate Tcl hacker. Computer Networks and ISDN Systems, 30(1– 7):713–715, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/comnet/
                                        sub/1998/30/1-7/1869.pdf.

Barron:1989:WUS

[Bar89] D. Barron. Why use SGML? Electronic Publishing—Origination, Dissemination, and Design, 2(1):3–24, April 1989. CO-DEN EPODEU. ISSN 0894-3982.

Barron:1992:SUG

[Bar92] D. Barron. SGML: The User’s Guide to ISO 8879: Joan M. Smith and Robert Stutely. Electronic Publishing—Origination, Dissemination, and Design, 5 (3):157–158, September 1992. CODEN EPODEU. ISSN 0894-3982.

Barry:1994:HML

[Bar94] Jeff Barry. The HyperText markup language (HTML) and the World-Wide Web: Raising ASCII text to a new level of usability. Public-access Computer Systems Review, 5(5):5–62, ???? 1994. ISSN 1063-164X.

Barber:1995:UMD

[Bar95] David Barber. The University of Michigan Data System. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
SocialSci/barber/umdata.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Barksdale:1997:HAW

[Bar97] Karl Barksdale. HTML activities Webtop publishing on the superhighway. South Western, Cincinnati, OH, USA, 1997. xiii + 178 pp.

BareBonesSoftware:1998:BLT

[Bar98a] Bare Bones Software, Inc. BBEdit 5.0 [the leading text and HTML editing tool for the Macintosh], 1998. Includes CD-ROM.

Barksdale:1998:HAW

[Bar98b] Karl Barksdale. HTML activities: Webtop publishing on the superhighway. South-Western Educational Pub., Cincinnati, OH, USA, 1998. ISBN 0-538-67458-X. xiii + 178 pp. LCCN TK5105.888.H76 1997.

Barksdale:1999:HPC

[Bar99]  Karl Barksdale.  HTML

programming concepts: brief course. International Thomson, South-Western, 1999. ISBN 0-538-69118-2. 112 pp. LCCN ????

Bartok:19xx:HH

[Barxx] ?. Bartok. HTML Handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 19xx. ISBN 1-56592-119-4. ???? pp. LCCN ????

Bath:1995:UWP

[Bat95]  M. Bath.  Untangling the

Web (publishing on the Internet). Publishing, pages 30– 32, September 1995. CODEN PUBLB6.

Bath:1996:FMQ

[Bat96a] M. Bath. Framing multimedia [Quark Immedia]. Publishing, page 29, August 1996. CODEN PUBLFA.

Bath:1996:MWW

[Bat96b] M. Bath. Multimedia Web weaving [online publishing]. Publishing, ??(??):34–36, March 1996. CODEN PUBLFA.

Batt:1996:NBL

[Bat96c]  Paul Batt.  Das Netz be-

ginnt zu leben — Java und VRML erweitern das World Wide Web zu einem nutzbringenden Massenmedium. Bulletin de l’Association Suisse des Electriciens, 87(9):11–16, April 1996. CODEN BUSEAH. ISSN 0004-587X.

Batt:1996:NCL

[Bat96d] Paul Batt. The Net comes to life: Java and VRML extend the World Wide Web into an advantageous mass-media network. Bulletin des Schweizerischen Elektrotechnischen Vereins and des Verbandes Schweizerischer Elektrizitaetswerke, 87(9):11– 16, April 26, 1996. CODEN BEVEDP. ISSN 0036-1321.

Bauerschmidt:1995:WIH

[Bau95] Rebecca Bauerschmidt. What is HTML? Against the Grain, 7(3): 31–??, June 1, 1995. ISSN 1043-2094.

Bormann:1990:ODP

[BB90] U. Bormann and C. Bormann.

Open document processing: status and future development. Informationstechnik — IT, 32(3): 176–185, June 1990. CODEN ICSAES. ISSN 0179-9738.

Bormann:1991:SOD

[BB91] U. Bormann and C. Bormann.

Standards for open document processing: current state and future developments. Computer Networks and ISDN Systems, 21 (3):149–163, May 1991. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Borgida:1993:LDD

[BB93] Alex Borgida and Ronald J. Brachman. Loading data into description reasoners. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):217–226, June 1993. CO-DEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Blunden:1994:DSA

[BB94] Brian Blunden and Margot Blunden, editors. Documentation of the seminars at the Frankfurt Electronic Media Conference during the 46th Frankfurt Book Fair, 5–10 October, 1994. Ausstellungs- und Messe, Frankfurt, Germany, 1994. ISBN 3-923352-25-5. LCCN ????

Baker:1995:UIH

[BB95a] C. Bruce Baker and Danny E. Brinegar. The use of the information highway to explore climate variability. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
EarSci/baker/Baker94.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Beck:1995:IFC

[BB95b] Bradley Carl Beck and Antonio M. Baptista. An interactive forum for convectiondiffusion problems. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
GenSci/beck/beck-baptista.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Bednarcyk:1995:LWI

[BB95c] Lauren A. Bednarcyk and Kevin D. Bond. A local Web for information delivery. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
CorInfSys/bednarcyk/bednarcyk.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Blair:1999:XCE

[BB99a] Barclay Blair and John Boyer.

XFDL: creating electronic commerce transaction records using XML. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1611–1622, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/
11-16/2154.pdf.

Bruzzone:1999:MUS

[BB99b] A. Bruzzone and G. Berrino.

Modeling of urban services by VRML and Java. In Bruzzone et al. [BUP99], pages 34– 40. ISBN 1-56555-156-7. LCCN QA76.9.C65 I574 1999.

Bormann:1988:SWE

[BBB88] U. Bormann, C. Bormann, and C. Bathe. SDE — a WYSIWYG editing and formatting system for ODA and SGML documents. In Anonymous [Ano88f], pages 1075–1095 (vol. 2). ISBN 0-444-87145-4. LCCN QA75.5.E84 1988. 2 vol.

Barnard:1995:HET

[BBG^{+}95a] David Barnard, Lou Burnard, Jean-Pierre Gaspart, Lynne A. Price, C. M. Sperberg-McQueen, and Giovanni Battista Varile. Hierarchical encoding of text: Technical problems and SGML solutions. Computers and the humanities, 29(3):211–231, ???? 1995. CODEN COHUAD. ISSN 0010-4817.

Berenson:1995:CAS

[BBG^{+}95b] Hal Berenson, Phil Bernstein, Jim Gray, Jim Melton, Elizabeth O’Neil, and Patrick O’Neil. A critique of ANSI SQL isolation levels. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (2):1–10, May 1995. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Borchardt:1996:CAS

[BBJR96] F. L. Borchardt, C. L. Bradin, E. Johnson, and L. Rhodes, editors. CALICO ’96 Annual Symposium. Proceedings of the Computer Assisted Language Instruction Consortium 1996 Annual Symposium ‘Distance Learning’, Albuquerque, NM, USA, 27 May–1 June 1996. Duke University, Durham, NC, USA, 1996. ISBN ???? LCCN ????

Brown:1995:EPD

[BBKF95] A. Brown, A. Bruggemann-Klein, and A. Feng, editors. Electronic publishing, document manipulation and typography: International conference; 6th — September 1996, Palo Alto, CA, volume 8(2). Wiley, New York, NY, USA, 1995. ISBN ???? ISSN 0894-3982. LCCN ????

Baker:1994:SWS

[BBKL94] T. Baker, I. Bruning, L. Klein, and M. Lenz. Starting a Web site: issues of quality control. Computer Networks and ISDN Systems, 26(Supplement 2–3): S55–S61, ???? 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Baker:1995:FGF

[BBKL95] Thomas Baker, Inke Bruning, Lothar Klein, and Michael Lenz. Not free, but growing fast: The Web in Germany. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Int/
baker/web-in-germany.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Barnard:1996:LLU

[BBSM96] D. T. Barnard, L. Burnard, and C. M. Sperberg-McQueen. Lessons learned from using SGML in the text encoding initiative. Computer Standards and Interfaces, 18(1):3–10, January 1996. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Benjamin:1990:BRS

[BC90] Michele Benjamin and Francis Cave. Book review: SGML — the user’s guide to ISO 8879, by Smith, Joan M. and Stutely, Robert. Chichester: Ellis Horwood/Halstead Press. 172 pp. ISBN 0-470-21126-1. £22.50. Learned Publishing, 3(2):128, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Braun:1995:WTC

[BC95] H-W Braun and K. C. Claffy. Web traffic characterization: An assessment of the impact of caching documents from NCSA’s Web server. Computer Networks and ISDN Systems, 28(1–2):37–??, December 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Blakstad-Cooke:1996:HPH

[BC96] Katie Blakstad-Cooke. Hypermedia publishing: how to build a Web site. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-237694-6. ???? pp. LCCN QA76.76.I59 B53 1996.

Bullington:1999:HCF

[BCG99] Jeffrey S. Bullington, Beatrice L. Caraway, and Beverley Geer, editors. Head in the clouds, feet on the ground: serials vision and common sense: proceedings of the North American Serials Interest Group, Inc. 13th Annual Conference, June 18–21, 1998, University of Colorado, Boulder, Colorado, volume 36(1/2, 3/4) of Serials Librarian. Haworth Information Press, New York, NY, USA, 1999. ISBN 0-7890-0768-1. ISSN 0361-526X. LCCN Z692.S5N67 1998.

Blake:1994:TRD

[BCK^{+}94] G. E. Blake, M. P. Consens, P. Kilpeläinen, P.-Å. Larson, T. Snider, and F. W. Tompa. Text/relational database management systems: harmonizing SQL and SGML. In Litwin and Risch [LR94], pages 267– 280. ISBN 0-387-58183-9, 3-540-58183-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 A67 1994.

Baker:1995:NMA

[BCLT95] J. E. Baker, I. F. Cruz, G. Liotta, and R. Tamassia. A new model for algorithm animation over the WWW. ACM Computing Surveys, 27(4):568– 572, December 1995. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Baggott:1995:N

[BD95] James Baggott and Sharon E. Dennis. NetBiochem. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        BioChem/baggott/baggott.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Ball:1996:IDE

[BD96] T. Ball and F. Douglis. An Internet Difference Engine and its applications. In IEEE [IEE96e], pages 71–76. ISBN 0-8186-7414-8. LCCN QA76 .I49 1996. IEEE Catalog No. 96CB35911.

Bourdakis:1997:VMC

[BD97] V. Bourdakis and A. Day. The VRML model of the city of bath. In Coyne [Coy97], pages 13– 22. ISBN 2-909285-07-3. LCCN ????

Benton:1995:MSH

[BDG^{+}95] Vivian M. Benton, Dan Dwyer, Cordelia Baron Geiken, Greg McArthur, and Joshua Polterock. MetaCenter science highlights repository. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

                                        http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Meta/
                                        metainfo.paper.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Bowman:1995:HID

[BDH^{+}95] C. M. Bowman, P. B. Danzig, D. R. Hardy, U. Manber, and M. F. Schwartz. The Harvest information discovery and access system. Computer Networks and ISDN Systems, 28(1– 2):119–??, December 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Borchers:1995:GIA

[BDK95] J. Borchers, O. Deussen, and C. Knorzer. Getting it across: layout issues for kiosk systems. ACM SIGCHI Bulletin, 27(4): 68–74, October 1995. CODEN SGBUD4. ISSN 0736-6906.

Borchers:1996:LRG

[BDKK96] J. Borchers, O. Deussen, A. Klingert, and C. Knorzer. Layout rules for graphical Web documents. Computers and Graphics, 20(3):415–426, May– June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Boumphrey:1998:XA

[BdRD^{+}98] Frank Boumphrey, Olivia di Renzo, Jon Duckett, Joe Graf, Dave Hollander, Paul Houle, Trevor Jenkins, Peter Jones, Adrian Kingsley-Hughes, Kathie Kingsley-Hughes, Craig McQueen, and Stephen Mohr. XML applications. Wrox Press, Chicago, IL, USA, 1998. ISBN 1-86100-152-5. xix + 649 pp. LCCN ????

Braccini:1995:IAP

[BDV95] Carlo Braccini, Leila DeFloriani, and Gianni Vernazza, editors. Image Analysis and Processing. 8th International Conference, ICIAP ’95. Proceedings. Springer-Verlag Inc., New York, NY, USA, 1995. ISBN 3-540-60298-4. LCCN TA1632.I552 1995.

Broll:1995:BWT

[BE95] W. Broll and D. England. Bringing worlds together: Adding multi-user support to VRML. In Nadeau and Moreland [NM95b], pages 87–94. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Bell:1997:HSS

[BE97] Ian (Ian Andrew) Bell and Marcus Eby. HTML studio skills. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1997. ISBN 1-56830-355-6. x + 309 pp. LCCN ????

Beretta:1999:CID

[BE99] Giordano B. Beretta and Reiner Eschbach, editors. Color imaging: device-independent color, color hardcopy, and graphic arts IV: 26–29 January, 1999, San Jose, California, volume 3648 of SPIE Proceedings. SPIE Optical Engineering Press, Bellingham, WA, USA, 1999. ISBN 0-8194-3119-2. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.3648.

Bean:1995:BWS

[Bea95a]  Greg Bean.  Building Web

servers for microcomputer foundation. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN

???? URL http://www.ncsa.
uiuc.edu/SDG/IT94/Proceedings/
DDay/bean/bean.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Bearman:1995:HHI

[Bea95b] David Bearman, editor. Hands on hypermedia and interactivity in museums: selected papers from the Third International Conference on Hypermedia and Interactivity in Museums (ICHIM ’95 — MCN ’95), San Diego, California, October 9–13, 1995. Archives and Museum Informatics, Pittsburgh, PA, USA, 1995. ISBN 1-885626-12-6. LCCN AM139 .H36 1995.

Beck:1993:SH

[Bec93] L. Beck. The SGML Handbook. In Anonymous [Ano93j], pages AP–15–AP–16. ISBN ???? ISSN ???? LCCN ????

Bedford:1995:ELS

[Bed95] Jeremy Bedford. Electronics Letters — the SGML implementation. Learned Publishing, 8 (1):39–44, January 1995. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bauwens:1994:IAI

[BEE^{+}94a] B. Bauwens, J. Engelen, F. Evenepoel, C. Tobin, and T. Wesley. Increasing access to information for the print disabled through electronic documents in SGML. In ACM [ACM94a], pages 55– 61. ISBN 0-89791-649-2. LCCN HV1569.5.A24 1994. ACM order number 444940.

Bauwens:1994:SDK

[BEE^{+}94b] B. Bauwens, J. Engelen, F. Evenepoel, C. Tobin, and T. Wesley. Structuring documents: the key to increasing access to information for the print disabled. In Zagler et al. [ZBW94], pages 214–221. ISBN 3-540-58476-5. LCCN HV1569.5.I56 1994.

Bauwens:1995:SPA

[BEE95] B. Bauwens, F. Evenepoel, and J. J. Engelen. Standardization as a prerequisite for accessibility of electronic text information for persons who cannot use printed material. IEEE Transactions on Rehabilitation Engineering, 3 (1):84–89, March 1995. CODEN IEEREN. ISSN 1063-6528.

Bauwens:1996:SET

[BEE96a] B. Bauwens, F. Evenepoel, and J. Engelen. SGML as an enabling technology for access to digital information by print disabled readers. Computer Standards and Interfaces, 18(1): 55–69, January 1996. CO-DEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Beebe:1996:LEW

[Bee96b] Nelson H. F. Beebe. Letter to the editor: Web lessons learned. BYTE Magazine, 21 (9):20, 22, September 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Advocates use of HTML/SGML grammar checkers.

Beech:1996:AS

[Bee96c]  S. Beech.  Alternatives to

SGML. Electrotechnology, 7(4): 33–35, August–September 1996. CODEN ETNYBB. ISSN 0306-8552.

Beeson:1997:OAV

[Bee97a] C. Beeson. An object-oriented approach to VRML development. In Spencer and Brutzman [SB97], pages 17–24. ISBN 0-89791-886-X. LCCN ????

URL http://www.acm.org/
pubs/contents/proceedings/
graph/253437/. ACM order

number 434973.

Beeton:1997:TUG

[Bee97b] B. Beeton, editor. TEX Users Group: 18th Annual meeting, July 1997, San Francisco, CA, volume 18(3) of TUGboat. TEX Users Group, San Francisco, CA, USA, 1997. ISBN ???? ISSN 0896-3207. LCCN ????

Belesis:1998:DHU

[Bel98] Peter Belesis. Dynamic HTML unleashed. Sams.net Pub., Indianapolis, IN, USA, second edition, 1998. ISBN 1-57521-353-2. xxvii + 714 pp. LCCN ????

Benson:1996:EI

[Ben96] P. Benson. EDI and the Internet. DBMS, Supplemental issue: 29–30, 32, May 1996. CODEN DBMSEO. ISSN 1041-5173.

Bennett:1997:HLC

[Ben97] Graeme Bennett. HTML: level I, courseware 7106. MasterTrak. CCI Computer Courseware International, Maple Ridge, BC, Canada, 1997. ISBN 1-55070-606-3. various pp. LCCN ????

Berghel:1995:UWT

[Ber95] H. Berghel. Using the WWW Test Pattern to check HTML client compliance. Computer, 28 (9):63–65, September 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berg:1996:JQH

[Ber96a] C. Berg. Java Q and A: How do I display media formats that Netscape does not support? Dr. Dobb’s Journal of Software Tools, 21(10):111, 114– 115, 129, October 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Berghel:1996:HCR

[Ber96b] H. Berghel. HTML compliance and the return of the test pattern. Communications of the ACM, 39(2):19–22, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berghel:1996:DVH

[Ber96c] Hal Berghel. Digital village:

HTML compliance and the return of the test pattern. Communications of the ACM, 39(2): 19–22, February 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/cacm/230799.
html.

Bowen:1998:ZZF

[BFH98] J. P. Bowen, Andreas Fett, and Michael G. Hinchey, editors. ZUM’98: the Z formal specification notation: 11th International Conference of Z Users, Berlin, Germany, September 24–26, 1998: proceedings, volume 1493 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. ISBN 3-540-65070-9. LCCN QA267.A1 L43 no.1493.

Bolas:1996:SDV

[BFM96] Mark T. Bolas, Scott S. Fisher, and John O. Merritt, editors. Stereoscopic displays and virtual reality systems III: 30 January– 2 February, 1996, San Jose, California, volume 2653 of Proceedings of the SPIE — The International Society for Optical Engineering. SPIE Optical Engineering Press, Bellingham, WA, USA, 1996. CO-DEN PSISDG. ISBN 0-8194-2027-1. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.2653.

Bhargava:1993:CPS

[BFY93] Bharat Bhargava, Timothy Finin, and Yelena Yesha, editors. CIKM 93: proceedings of the Second International Conference on Information and Knowledge Management, November 1–5, 1993, Washington, DC, USA, CIKM - International Conference 1993, 2nd. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-626-3. LCCN QA76.9.D3I5586 1993.

Balaguer:1995:IHS

[BG95a] J.-F. Balaguer and E. Gobbetti. i3D: a high-speed 3D Web browser. In Nadeau and Moreland [NM95b], pages 69– 76. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Balaguer:1995:IHW

[BG95b] J.-F. Balaguer and E. Gobbetti. i3D: A high-speed 3D Web browser. In Nadeau and Moreland [NM95b], pages 69– 76. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Beam:1996:TWW

[BG96] P. Beam and P. Goldsworthy. Technical writing on the Web — distributed SGML-based learning. In ACM [ACM96d], pages 35–41. ISBN 0-89791-799-5. LCCN QA 76.9 D6 I56 1996.

Botsford:1994:PCI

[BGG^{+}94] J. Botsford, A. Gawman, M. Gentleman, E. Kidd, K. Lyons, and J. Slonim, editors. Proceedings. CASCON ’94. Integrated Solutions. Nat. Res. Council Canada, Ottawa, Ont., Canada, 1994. ISBN ???? LCCN ????

Boeri:1995:WGI

[BH95a] Robert J. Boeri and Martin Hensel. What good is SGML? CD-ROM professional, 8(4): 108–110, April 1995. CODEN CRPFEX. ISSN 1049-0833.

Bowen:1995:ZFS

[BH95b] Jonathan P. Bowen and Michael G.

Hinchey, editors. ZUM ’95, the Z formal specification notation: 9th International Conference of Z Users, Limerick, Ireland, September 7–9, 1995: proceedings, number 967 in Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. ISBN 3-540-60271-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.Z2 I58 1995.

Baron:1998:CWP

[BH98a] J. Baron and S. Heath. Creating Web pages using basic HTML: an introduction. In Charou et al. [CBK98], pages 149–162. ISBN 960-85952-4-X. LCCN Z675.M35 E87 1998.

Brown:1998:UH

[BH98b] Mark Robbin Brown and Jerry Honeycutt. Using HTML 4. Que Corporation, Indianapolis, IN, USA, fourth edition, 1998. ISBN 0-7897-1449-3. xx + 1103 pp. LCCN QA76.76.H94B787 1998.

Barnard:1988:SML

[BHK^{+}88] David Barnard, Ron Hayter, Maria Karababa, G. Logan, and J. McFadden. SGML-based markup for literary texts: two problems and some solutions. Computers and the humanities, 22(4):265–276, ???? 1988. CO-DEN COHUAD. ISSN 0010-4817.

Brien:1995:UMW

[BHK^{+}95a] John Brien, Peggy Havukainen, Matthew Krenmayr, Antonio Artese, and Mark Silva. Using Mosaic/WWW in a business school: New opportunities for information sharing and teaching. issues in implementing organizational-wide networked multimedia. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
CorInfSys/brien/Chicago.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Brown:1995:H

[BHK^{+}95b] Jeremy Brown, Jake Harris, Lara Karbiner, Massimiliano Poletto, Andre’ DeHon, Ian Eslick, and Thomas F. Knight, Jr. HyperCode. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
brown/hypercode/hypercode.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Barth:1998:XPI

[BHMS98] A. Barth, L. Helmes, J. Mattes, and T. Strelow. XML, patent information and the Web. In Collier [Col98], pages 92–111. ISBN 1-873699-58-1. LCCN ????

Bederson:1996:ZWB

[BHS^{+}96] B. B. Bederson, J. D. Hollan, J. Stewart, D. Rogers, A. Druin, and D. Vick. A zooming Web browser. In Freeman et al. [FJV96], pages 260–271. CO-DEN PSISDG. ISBN 0-8194-2041-7. ISSN 0277-786X (print), 1996-756X (electronic). LCCN QA76.575 .M784 1996.

Benes:1996:LSD

[BHZ96] B. Benes, A. Holecek, and J. Zara. Language support for distributed virtual reality. In Thalman and Skala [TS96], pages 233–243. ISBN 80-7082-238-4. LCCN ???? 2 vol.

Bar-Ilan:1995:SWH

[BIL95] J. Bar-Ilan and S. S. Lazinger.

Setting up a Web home page: a case study in ‘design by perusal’. In Raitt and Jeapes [RJ95], pages 189–196. ISBN 0-904933-94-6. LCCN ????

Birnbaum:1996:CW

[Bir96a] Bradley Birnbaum. Chatting on the Web. Linux Journal, 21: ??, January 1996. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Birnbaum:1996:SCG

[Bir96b] D. J. Birnbaum. Standardizing characters, glyphs, and SGML entities for encoding early Cyrillic writing. Computer Standards and Interfaces, 18(3):201–252, June 1996. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Brown:1996:SEU

[BJ96] Mark Brown and John Jung. Special edition Using Intranet HTML. Que Corporation, Indianapolis, IN, USA, special edition, July 1996. ISBN 0-7897-0852-3. xli + 1080 pp. LCCN QA76.76.H94S87 1996. US$59.99.

Bjorner:1996:CCH

[Bjo96a] Susanne Bjorner. Corrections and clarifications — an HTML stripper. Online, 20(3):60–??, May 1, 1996. CODEN ONLIDN. ISSN 0146-5422.

Bjorner:1996:OO

[Bjo96b] Susanne Bjorner. Output options. Online, 20(1):57–??, January 1, 1996. CODEN ONLIDN. ISSN 0146-5422.

Bjorner:1996:OOJ

[Bjo96c] Susanne Bjorner. Output options — just what you wanted for Christmas: An HTML stripper! Online, 20(1):57–??, ???? 1996. CODEN ONLIDN. ISSN 0146-5422.

Brueggemann-Klein:1993:UER

[BK93a] A. Brueggemann-Klein. Unambiguity of extended regular expressions in SGML document grammars. In Lengauer [Len93], pages 73–84. ISBN 3-540-57273-2, 0-387-57273-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43E83 1993.

Bruggemann-Klein:1993:FMD

[BK93b] A. Brüggemann-Klein. Formal Models in Document Processing. Habilitation, Institut für Informatik, Universität Freiburg, Freiburg, Germany, ???? 1993.

Bruggemann-Klein:1993:REF

[BK93c] A. Brüggemann-Klein. Regular expressions into finite automata. Theoretical Computer Science, 120(2):197–213, November 1993. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruggemann-Klein:1993:UER

[BK93d] A. Brüggemann-Klein. Unambiguity of extended regular expressions in SGML document grammars. In Lengauer [Len93], pages 73–84. ISBN 3-540-57273-2, 0-387-57273-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43E83

BruggemannKlein:1993:UER

[BK93e] A. Bruggemann-Klein. Unambiguity of extended regular expressions in SGML document grammars. In Lengauer [Len93], pages 73–84. ISBN 3-540-57273-2, 0-387-57273-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43E83 1993.

Broll:1996:VTT

[BK96] W. Broll and T. Koop. VRML: today and tomorrow. Computers and Graphics, 20(3):427– 434, May–June 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bernadac:1999:CAX

[BK99] Jean-Christophe Bernadac and François Knab. Construire une application XML: avec deux études de cas réels. Solutions dévéloppeurs. Eyrolles, Paris, France, 1999. ISBN 2-212-09081-1. xv + 476 pp. LCCN ???? URL http:

//www.eyrolles.com/livres/

bernadac/. Télédéchargez le code source des études de cas, consultez les mises à jour et complements, dialoguez avec les auteurs.

Bergeron:1997:VDH

[BKP97] Marcel Bergeron, Corinne Kempa, and Yolande Perron. Vocabulaire d’Internet: HTML, Java, VRML, cyberculture: terminologie des technologies de l’information: vocabulaire anglais-français. Cahiers de l’Office de la langue française. Gouvernement du Québec, Office de la langue française, Québec, PQ, Canada, deuxième edition, 1997. ISBN 2-551-17286-1. 140 pp. LCCN TK5105.875.I57B47 1997.

Brusilovsky:1996:MHV

[BKS96] Peter Brusilovsky, Piet Kommers, and Norbert Streitz, editors. Multimedia, Hypermedia, and Virtual Reality. Models, Systems, and Applications. First International Conference, MHVR‘94. Selected Papers, volume 1077 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. ISBN 3-540-61282-3. LCCN QA76.575 .M38 1994.

Banissi:1998:ICI

[BKS98] Ebad Banissi, Farzad Khosrowshahi, and Muhammad Sarfraz, editors. 1998 IEEE Conference on Information Visualization: an International Conference on Computer Visualization & Graphics: proceedings, July 29–31, 1998, London, England. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8509-3, 0-8186-8510-7, 0-8186-8511-5. ISSN 1093-9547. LCCN TA1634.I132 1998. IEEE catalog number 97TB1200165.

Bloomberg:1996:WPS

[BKT96] Jason Bloomberg, Jeff Kawski, and Paul Treffers. Web Page Scripting Techniques. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1996. ISBN 1-56830-307-6. xviii + 510 pp. LCCN QA76.625.B556 1996. US$50.00, CDN$70.95, UK£46.95. URL http://www.

                                        mcp.com/hayden;   http://
                                        www.mcp.com/hayden/internet/
                                        scripting/index.html. Dis-

cusses JavaScript, VBScript, and advanced HTML scripting techniques.

Bruggemann-Klein:1992:DRL

[BKW92a] A. Brüggemann-Klein and D. Wood. Deterministic regular languages. In Finkel and Jantzen [FJ92], pages 173–184. ISBN 3-540-55210-3. LCCN QA75.5.S958 1992.

BruggemannKlein:1992:DRL

[BKW92b] A. Bruggemann-Klein and D. Wood. Deterministic regular languages. In Finkel and Jantzen [FJ92], pages 173–184. ISBN 3-540-55210-3. LCCN QA75.5.S958 1992.

Bruggemann-Klein:1992:ESS

[BKW92c] Anne Brüggemann-Klein and Derick Wood. Electronic style sheets. Bericht 45, Institut fü Informatik, Universität Freiburg, Freiburg, Germany, 1992. Also available as University of Western Ontario (London, Ontario, Canada) Computer Science Department Technical Report 350.

Bruggemann-Klein:1992:URE

[BKW92d] Anne Brüggemann-Klein and Derick Wood. Unambiguous regular expressions and SGML document grammars. Technical Report 337, Computer Science Department, University of Western Ontario, London, Ontario, Canada, November 1992. ISBN 0-7714-1454-4.

Bruggemann-Klein:1993:VSC

[BKW93] Anne Brüggemann-Klein and Derick Wood. The validation of SGML content models. Technical Report 355, Computer Science Department, University of Western Ontario, London, Ontario, Canada, March 1993.

Bruggemann-Klein:1997:VSC

[BKW97] A. Brüggemann-Klein and D. Wood. The validation of SGML content models. Mathematical and computer modelling, 25(4):73–84, ???? 1997. CO-DEN MCMOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

Brandenberger:1998:CMH

[BL98] L. Brandenberger and D. Lineberger.

Creation and maintenance of a horticultural education Web site by an extension specialist without HTML programming skills. In Anonymous [Ano98d], pages 355–?? ISBN ???? ISSN 0018-5345. LCCN ????

Blake:1990:PPR

[Bla90] M. Blake. Practices of printers in relation to electronic archiving. Electronic Library, 8(4): 239–243, August 1990. CODEN ELLIDZ. ISSN 0264-0473.

Blakeley:1994:OOD

[Bla94] José A. Blakeley. Open object database management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):520, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Berners-Lee:1995:RHM

[BLC95] T. Berners-Lee and D. Connolly. RFC 1866: Hypertext Markup Language — 2.0, November 1995. URL ftp://ftp.

internic.net/rfc/rfc1866.
txt;   http://www.math.utah.
edu/pub/rfc/rfc1866.txt. Sta-

tus: PROPOSED STANDARD. Berners-Lee:1992:WWW

[BLCG92] T. J. Berners-Lee, R. Cailliau, and J.-F. Groff. The World-Wide Web. Computer Networks and ISDN Systems, 25 (4–5):454–459, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Berners-Lee:1996:RHT

[BLFF96] T. Berners-Lee, R. Fielding, and H. Frystyk. RFC 1945: Hypertext Transfer Protocol — HTTP/1.0, May

1996.  URL ftp://ftp.
internic.net/rfc/rfc1945.
txt;   http://www.math.utah.
edu/pub/rfc/rfc1945.txt. Sta-

tus: INFORMATIONAL.

Bayle:1995:JWU

[BLG^{+}95] Michael L. Bayle, Burton H.

Lee, A. Goto, Y. Sakamoto, J. Hong, V. Lam, T. Takada, M. Washisaka, A. Yamazaki, and K. Kada. Japan window: A US-Japan WWW collaboration for Japan information. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Int/lee/japan_window.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Blum:1995:BBW

[Blu95a] Adam Blum. Building business Web sites. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1995. ISBN 1-55828-431-1. ???? pp. LCCN HD30.37 .B56 1995.

Blumberg:1995:MES

[Blu95b] Roger B. Blumberg. Mendel-Web: An electronic science/math/history resource for the WWW. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/
blumberg.mendelweb/MendelWeb94.
blumberg.html. A selected few

of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Blumberg:1995:MPL

[Blu95c] Roger B. Blumberg. Museums, public lands, and billboards: Toward a philosophy of the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Overviews/blumberg.museums/
                                        WebPhil94.Blumberg.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

BlueSkySoftware:1997:WHG

[Blu97] Blue Sky Software Corp. Win-Help 2000 the HTML generation of WinHelp!, 1997. Includes CD-ROM.

Blyth:1996:WPD

[Bly96] D. S. Blyth. Web page design with HTML and friends. In Anonymous [Ano96k], pages 89–103. ISBN 0-7803-3689-5 (softbound), 0-7803-3690-9 (casebound), 0-7803-3691-7 (microfiche). LCCN T10.5.I54 1996. IEEE catalog number 96CH36009.

Barnes:1991:MTU

[BM91] J. A. Barnes and S. A. Mamrak. A model and toolset for the uniform tagging of encoded documents. Electronic Publishing—Origination, Dissemination, and Design, 4(2):63–85, June 1991. CODEN EPODEU. ISSN 0894-3982.

Bahler:1998:SHH

[BMDB98] C. C. Bahler, S. Muller, D. Doolittle, and A. Barrios. SAS and HTML — HTML publishing using SAS. In Anonymous [Ano98t], pages 229–237. ISBN 1-58025-149-8. LCCN ????

Beck:1995:WWP

[BMK95] Howard Beck, Amir Mobini, and Viswanath Kadambari. A word is worth 1000 pictures: Natural language access to digital libraries. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Searching/beck/beckmain.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Bohm:1994:SDH

[BMN94] K. Bohm, A. Muller, and E. Neuhold. Structured document handling — a case for integrating databases and information retrieval. In Adam et al. [ABY94], pages 147– 154. ISBN 0-89791-674-3. LCCN QA76.9.D3 I55865 1994.

URL http://www.acm.org/
pubs/contents/proceedings/
cikm/191246/.

Balasubramanian:1995:SAD

[BMY95] V. Balasubramanian, Bang Min Ma, and Joonhee Yoo. A systematic approach to designing a WWW application. Communications of the ACM, 38 (8):47–48, August 1995. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brutzman:1995:IIR

[BMZ95] D. P. Brutzman, M. R. Macedonia, and M. J. Zyda. Internetwork infrastructure requirements for virtual environments. In Nadeau and Moreland [NM95b], pages 95–104. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Busch:1998:CSS

[BO98] David D. Busch and J. W. Olsen. Cascading Style Sheets complete. McGraw-Hill, New York, NY, USA, 1998. ISBN 0-07-913703-2 (paperback). xxi + 424 pp. LCCN TK5105.888 .B86 1998 Stacks.

Boar:1990:TRS

[Boa90] K. Boar. Text representations (SGML). In Peter Gillman, editor, Text retrieval, the state of the art. Proceedings of the Institute of Information Scientists text retrieval conferences ‘The User’s Perspective’ (1988) and ‘Text Management’ (1989), pages 97–100. Taylor Graham, London, UK, ???? 1990. ISBN 0-947568-44-1. LCCN Z699.A1 I65 1988.

Boalch:1995:HPD

[Boa95] G. Boalch. HTML page development — A case study. In Debreceny and Ellis [DE95], pages 193–198. ISBN 0-646-23946-5. LCCN ????

Bohrer:1995:AFW

[Boh95] Rand Bohrer. Architecture of the future: — A web exhibition. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Museum/bohrer/bohrer.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Bohush:1996:AP

[Boh96] Peter Bohush. Adobe Pagemill 1.0. Photo Electronic Imaging, 39(4):15–??, April 1, 1996. CO-DEN PELIE6. ISSN 0146-0153.

Boisvert:1995:WGV

[Boi95] Ronald F. Boisvert. A Web gateway to a virtual mathematical software repository. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ???? URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
boisvert/boisvert.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Bonin:1992:CCA

[Bon92] H. E. G. Bonin. Cooperation and collaboration assisted by editors. Wirtschaftsinformatik, 34(6):590–598, December 1992. CODEN WIINE9. ISSN 0937-6429.

Bonzi:1993:AMO

[Bon93] Susan Bonzi, editor. ASIS ’93: 56th Annual meeting — October 1993, Columbus, OH, USA, volume 30 of Proceedings of the Annual Meeting — American Society for Information Science. Learned Information, Medford, NJ, USA, 1993. ISBN 0-938734-78-4. ISSN 0044-7870. LCCN Z699.A1 A5 1993.

Booch:1996:ODJ

[Boo96] G. Booch. Object-oriented development with Java. Report on Object Analysis and Design, 2(6):2–4, March–April 1996. ISSN 1075-2528.

Born:1995:FFH

[Bor95] Günter Born. The File Formats Handbook. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1995. ISBN 1-85032-117-5, 1-85032-128-0. xvii + 1274 pp. LCCN QA76.9.F5 B67 1995. US$59.95.

Borchart:1996:CAL

[Bor96] F. L. Borchart, editor. Computer Assisted Language Instruction Consortium: Annual symposium — May 1996, Albuquerque, NM, CALICO Annual Symposium 1996. Duke University, ????, 1996. ISBN ???? ISSN 1085-3480. LCCN ????

Bose:1997:CHW

[Bos97]  T. K. Bose, editor.  Pro-

ceedings of the 8th Canadian hydrogen workshop (May 28, 29, 30, 1997, Sheraton Centre Hotel, Toronto, Canada) = Actes du 8e atelier canadien sur l’hydrogène. Canadian Hydrogen Association, Toronto, ON, Canada, 1997. ISBN 0-9696869-2-7 (valid ISBN allocated to two books: the 7th and 8th workshop). LCCN QA76.9.S63 I56 1998.

Bosak:1999:XUS

[Bos99] J. Bosak. XML ubiquity and the scholarly community. In Mylonas and Renear [MR99b], pages 199–206. ISBN ???? ISSN 0010-4817. LCCN ????

Bothma:1992:PGI

[Bot92] T. J. D. Bothma. Presenting grammatical information in biblical information systems. In Raitt [Rai92], pages 163–172. ISBN 0-904933-83-0. LCCN ????

Bottaci:1993:CSD

[Bot93] L. Bottaci. Computer support for documentation consistency maintenance. In Anonymous [Ano93h], pages 11/1–3. ISBN ???? LCCN ????

Bottomley:1998:ISL

[Bot98] Gareth Bottomley. Implementation of a scripting language for VRML using Java. University of Leeds, School of Computer Studies, Leeds, 1998.

Boumans:1990:DRS

[Bou90] P. W. J. M. Boumans. The dissemination of the results of scientific research in the era of electronic media. Spectrochimica Acta, Part A [Molecular Spectroscopy], special issue:11– 45, ???? 1990. CODEN SAM-CAS. ISSN 0584-8539.

Boutell:1995:TSS

[Bou95a] Thomas Boutell. Techniques for server-side dynamic document generation. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

                                        http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/
                                        boutell/servergen.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Bouvier:1995:SH

[Bou95b] D. J. Bouvier. The state of HTML. SIGICE Bulletin, 21(2): 8–13, October 1995. CODEN SGBLEC.

Bourdakis:1997:FVL

[Bou97a] V. Bourdakis. The future of VRML on large urban models. In Bowden [Bow97], pages 32– 40. ISBN 1-872166-64-4, 1-872166-69-5. LCCN ????

Bourell:1997:SFF

[Bou97b] D. L. Bourell, editor. Solid freeform fabrication symposium: 8th — September 1997, Austin, TX, Solid Freeform Fabrication Proceedings 1997. University of Texas, ????, 1997. ISBN ???? ISSN 1053-2153. LCCN ????

Boumphrey:1998:PSS

[Bou98]  Frank Boumphrey.  Profes-

sional style sheets for HTML and XML. Wrox Press, Chicago, IL, USA, 1998. ISBN 1-86100-165-7. xxi + 673 pp. LCCN QA76.76.H94 B68 1998.

Bowen:1995:CZF

[Bow95] J. P. Bowen. Comp.specification.z and Z FORUM frequently asked questions. In Bowen and Hinchey [BH95b], pages 561– 569. ISBN 3-540-60271-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.Z2 I58 1995.

Bowers:1996:WQA

[Bow96] Neil Bowers. Weblint: quality assurance for the World Wide Web. Computer Networks and ISDN Systems, 28 (7–11):1283–1290, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=
1591.

Bowden:1997:UVR

[Bow97] R. Bowden, editor. UK Virtual Reality Special Interest Group: Conference; 4th — November 1997, Uxbridge, Proceedings of the UK Virtual Reality Special Interest Group Conference 1997; 4th. Brunel University, ????, 1997. ISBN 1-872166-64-4, 1-872166-69-5. LCCN ????

Boyer:1995:IDD

[Boy95] Paul Boyer. Information discovery and distillation in government: An experience report. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
NationWideSites/scott/index.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Bodarky:1993:SDS

[BP93] Scott Bodarky and Scott W. Paisley. An SGML DTD for the STEP integrated resource parts. Technical report, U.S. Dept. of Commerce, National Institute of Standards and Technology, Gaithersburg, MD, USA, 1993. iv + 31 pp.

Becker:1996:APC

[BP96] R. Becker and M. Paciuk, editors. Applications of the performance concept in building: International symposium; 3rd — December 1996, Tel Aviv, Israel, volume 1 of Applications of the Performance Concept in Building -International Symposium-1996. National Building Research Institute, Haifa, Israel, 1996. ISBN 965-222-752-8. LCCN ????

Brutzman:1996:VPF

[BPBV96] D. Brutzman, M. Pesce, G. Bell, and A. Van Dam. VRML: Prelude and future. In ACM [ACM96a], pages 489–490. ISBN 0-89791-746-4. ISSN 1069-529X. LCCN T385 .S54 1996. ACM order number: 428960.

Burger:1995:UHM

[BQRT95] Franz Burger, Gerald Quirchmayr, Siegfried Reich, and A. Min Tjoa. Using HyTime for modelling publishing workflows. SIGOIS Bulletin, 16(1): 39–??, ???? 1995. CODEN SIG-BEL. ISSN 0894-0819.

Bohm:1994:MMD

[BR94] K. Bohm and T. C. Rakow.

Metadata for multimedia documents. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(4):21– 26, December 1994. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Britton:1995:DUI

[BR95] David R. Britton, Jr. and Arthur A. Reyes. Discovering usability improvements for Mosaic: Applications of the contextual inquiry technique with an expert user. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
HCI/britton/britton_reyes.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Bonhomme:1996:IRH

[BR96a] Stéphane Bonhomme and Cécile Roisin. Interactively restructuring HTML documents. Computer Networks and ISDN Systems, 28(7–11):1075–1084, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1996&
                                        volume=28&issue=7-11&aid=
                                        1595.

Burger:1996:DIA

[BR96b] F. Burger and S. Reich. Design and implementation of an abstract SGML interface in Smalltalk. Computer Standards and Interfaces, 18(1): 71–78, January 1996. CO-DEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Bray:1989:LNO

[Bra89] Tim Bray. Lessons of the New Oxford English Dictionary Project. In USENIX Association [USE89], pages 187–199. ISBN ???? LCCN ???? This dictionary is coded with SGML markup.

Bradley:1992:SC

[Bra92]  Neil Bradley.  SGML con-

cepts. ASLIB Proceedings, 44(7-8):271–274, July–August 1992. CODEN ASLPAO. ISSN 0001-253X (print), 1758-3748 (electronic).

Bradley:1996:TST

[Bra96a] J. Bradley. TACT and SGML together at last: sgml2tdb. In Anonymous [Ano96f], pages 42– 43. ISBN 82-7283-083-3. ISSN 0804-8800. LCCN ????

Bradley:1996:CSC

[Bra96b] Neil Bradley. The Concise SGML Companion. Addison-Wesley Longman, Reading, MA, USA, November 1996. ISBN 0-201-41999-8. xi + 324 pp. LCCN QA76.76.H94 B73 1996. US$29.95. URL http://www.

aw.com/devpress/titles/41999.
html.

Bradley:1998:XC

[Bra98] Neil Bradley. The XML companion. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-34285-5 (paperback). xi + 440 pp. LCCN QA76.76.H94B77 1998.

Bradley:1999:XC

[Bra99] Neil Bradley. The XML companion. Addison-Wesley, Reading, MA, USA, second edition, 1999. ISBN 0-201-67486-6. x + 566 pp. LCCN QA76.76.H94 B77 2000 Bar.

Brewin:1996:SPD

[Bre96]  P. Brewin.  SGML and

patent document processing. I. WIPO standard ST.32. World Patent Information, 18(4):183– 192, December 1996. CODEN WPAID2. ISSN 0172-2190.

Bride:1998:HPWa

[Bri98a] Mac Bride. HTML publishing on the World Wide Web. Teach yourself Teach yourself books. Hodder and Stoughton, London, UK, second edition, 1998. ISBN 0-340-71220-1. vii + 216 pp. LCCN ????

Bride:1998:HPWb

[Bri98b] Mac Bride. HTML publishing on the World Wide Web. Teach yourself books (Lincolnwood, Ill.). NTC/Contemporary Pub., Lincolnwood (Chicago), IL, USA, second edition, 1998. ISBN 0-8442-0218-5 (paperback). vii + 216 pp. LCCN ????

Brown:1989:SSD

[Bro89] H. Brown. Standards for structured documents. The Computer Journal, 32(6):505–514, December 1989. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Brooks:1994:CDE

[Bro94a] G. Brooks. Changes in the documentation engineering process. Electrical Communication, ???? 1994. CODEN ELCMAX. ISSN 0013-4252.

Brooks:1994:DEP

[Bro94b] G. Brooks. Changes in the documentation engineering process. Electrical Communication, pages 173–180, Second Quarter 1994. CODEN ELCMAX. ISSN 0013-4252.

Broll:1997:PIS

[Bro97]  W. Broll.  Populating the

Internet: Supporting multiple users and shared applications with VRML. In Spencer and Brutzman [SB97], pages 33–40. ISBN 0-89791-886-X. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/253437/.

ACM order number 434973.

Broll:1998:DIP

[Bro98a] W. Broll. DWTP — an Internet protocol for shared virtual environments. In Spencer and Roehl [SR98], pages 49– 56. ISBN 1-58113-022-8. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/271897/.

Brown:1998:H

[Bro98b] Mark Brown. HTML 4. Le Macmillan. Simon & Schuster Macmillan (France), Paris, France, 1998. ISBN 2-7440-0310-7. xxvii + 1034 pp. LCCN ????

Buford:1994:TAG

[BRR94] J. F. Buford, L. Rutledge, and J. L. Rutledge. Toward automatic generation of HyTime applications. In Herzner and Kappe [HK94], pages 57–76. ISBN 3-211-82587-8. LCCN QA76.575 .M833 1994.

Brutzman:1998:VRM

[Bru98] Don Brutzman. The Virtual Reality Modeling Language and Java. Communications of the ACM, 41(6):57–64, June 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1998-41-6/p57-
                                        brutzman/.

Bryan:1988:SAG

[Bry88] Martin Bryan. SGML: An Author’s Guide to the Standard Generalized Markup Language. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-17535-5. xvii + 364 pp. LCCN QA76.73.S44 B79 1988. UK£16.95.

Bryan:1993:STH

[Bry93] M. Bryan. Standards for text and hypermedia processing. Information services and use, 13 (2):93–102, ???? 1993. CODEN ISUDX8. ISSN 0167-5265.

Bryan:1997:SHE

[Bry97] Martin Bryan. SGML and HTML Explained. Addison-Wesley Longman, Reading, MA, USA, second edition, February 1997. ISBN 0-201-40394-3. xx + 234 pp. LCCN QA76.76.H94 B79 1997. US$34.95.

Bryan:1998:MJH

[Bry98] S. Bryan. Multimedia, Java, and HTML with Digital Chisel 3. In Anonymous [Ano98a], pages 48– ?? ISBN 1-56484-134-0. LCCN ????

Bryant:1999:TYH

[Bry99] Stephanie Cottrell Bryant. Teach yourself HTML 4. I D G Books Worldwide, Indianapolis, IN, USA, 1999. ???? pp.

Blum:1990:ISQ

[BS90] Bruce I. Blum and Ralph D.

Semmel. The INA: A simple query language with only attribute names: A video presentation. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):390, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Barry:1994:GAE

[BS94] A. Barry and D. Stanton. Glossary: Australian electronic library and network developments. Australian Library Review, 11(1):87–89, February 1994. CODEN ALIRE4. ISSN 1034-8042.

Barnum:1995:AVS

[BS95a] J. Barnum and C. Sampson.

Adding value to SAS data with SGML. In Anonymous [Ano95a], pages 750–759. ISBN 1-55544-211-0. LCCN ????

Boston:1995:ISDb

[BS95b] A. N. Boston and D. R. B.

Stockwell. Interactive species distribution reporting, mapping and modelling using the World Wide Web. Computer Networks and ISDN Systems, 28(1– 2):231–238, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/

tree/store/comnet/cas_sub/
browse/browse.cgi?year=1995&
volume=28&issue=1-2&aid=1537.

Boston:1995:ISDa

[BS95c] Tony Boston and David Stockwell. Interactive species distribution reporting, mapping, and modelling using the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
EarSci/boston/species_paper.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Bowen:1998:NIB

[BS98] Ted Smalley Bowen and Ed Scannell. News: IBM bets on XML. JavaWorld: IDG’s magazine for the Java community, 3(0), January 1998. CO-DEN ???? ISSN 1091-8906.

URL http://www.javaworld.
com/javaworld/jw-00-1998/
jw-00-iw-xml.htm.

Behram:1995:EAW

[BSB^{+}95] Antia Behram, Brett Stroozas, William Boyd, Carol Christian, and Kuei Chen. EUV astrophysics and the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Astronomy/antia/antia.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Ball:1998:PXH

[BSF98] M. J. Ball, A. Sethi, and R. Ferrans. Panel: XML in healthcare: Easing the interchange of medical records among disparate systems. In Anonymous [Ano98c], pages 94–?? ISBN ???? LCCN ????

Bearman:1997:MWM

[BT97] David Bearman and Jennifer Trant, editors. Museums and the Web: Museums and the Web 97: selected papers, Los Angeles, CA, March 1997. Archives and Museum Informatics, Pittsburgh, PA, USA, 1997. ISBN 1-885626-13-4. LCCN AM2.M876 1997.

Barksdale:1999:HJC

[BT99] Karl Barksdale and E. Turner.

HTML and JavaScript: concepts and activities. South-Western Educational Pub., Cincinnati, OH, USA, 1999. ISBN 0-538-68822-X. 192 pp. LCCN ????

Buckley:1993:GLL

[Buc93] Mark. Buckley. A generalized language for logical structures. Bellcore Exchange, 9(1):18–??, January 1, 1993. ISSN 1040-

Buchanan:1997:SDE

[Buc97] William J. Buchanan. Software development for engineers with C, Pascal, C++, Assembly Language, Visual Basic, HTML, JavaScript, and Java. Arnold, London, UK, 1997. ISBN 0-340-70014-9 (paperback). xiv + 674 pp. LCCN QA76.76.D47B82 1997.

Buford:1996:EHE

[Buf96] J. F. Buford. Evaluating Hy-Time: an examination and implementation experience. In ACM [ACM96b], pages 105–115. ISBN 0-89791-778-2. LCCN QA76.76.H94 A26 1996. ACM order number: 614960.

Bugg:1998:NHH

[Bug98] Keith Bugg. The new HTML Help Standard. C/C++ Users Journal, 16(8):??, August 1998. CODEN CCUJEX. ISSN 1075-2838.

Bundy:1996:HGU

[Bun96] Charles Bundy. HTML as a graphical user interface for sys admin tasks. SysAdmin, 5(2): 51–56, 59, 60, 62, 63, February 1996. CODEN SYADE7. ISSN 1061-2688.

Bruzzone:1999:PIC

[BUP99] Agostino G. Bruzzone, Adelinde Uhrmacher, and Ernest H. Page, editors. Proceedings of the 1999 International Conference on Web-based Modeling and Simulation: 1999 Western Multi-Conference, San Francisco, California, January 17–20, 1999, Cathedral Hill Hotel, volume 31(3) of Simulation Series. Society for Computer Simulation International, San Diego, CA, USA, 1999. ISBN 1-56555-156-7. LCCN QA76.9.C65 I574 1999.

Burgard:1993:IO

[Bur93] Mike Burgard. Information overload! UNIX/world, 10(6): 111–??, June 1, 1993. ISSN 0739-5922.

Burnard:1995:WSH

[Bur95a] L. Burnard. What is SGML and how does it help? Computers and the humanities, 29(1): 41–50, ???? 1995. CODEN CO-HUAD. ISSN 0010-4817.

Burnard:1995:WIS

[Bur95b] Lou Burnard. What is SGML and how does it help? Computers and the humanities, 29(1 (or 3??)):41–??, 1995. CODEN CO-HUAD. ISSN 0010-4817.

Burnard:1999:USL

[Bur99a] Lou Burnard. Using SGML for linguistic analysis: the case of the BNC. Markup Languages: Theory & Practice, 1 (2):31–51, Spring 1999. CODEN MLTPFG. ISSN 1099-6621.

Burns:1999:HG

[Bur99b] Joe Burns. HTML goodies. Que Corporation, Indianapolis, IN, USA, 1999. ISBN 0-7897-1823-5. xxvix + 527 pp. LCCN QA76.76.H94 B87 1999.

Bushek:1988:LSF

[Bus88] Joan Bushek. L^{A}TEX and SGML: a functional evaluation of one approach to automatic generation of translators. Thesis (m.s.), Department of Computer and Information Science. Ohio State University, Columbus, OH, USA, 1988. 132–?? pp.

Busch:1997:SCH

[Bus97] J. A. Busch. SGML for cultural heritage information. In Logan and Gluck [LG97], pages 9– 26. ISBN 1-57387-036-6. LCCN Z286.E43 L64 1997.

Baldazo:1995:HAE

[BVN95a] Rex Baldazo and Steven J. Vaughan-Nichols. HTML authoring and editing tools promise to make your documents Web-ready in no time. but few live up to their claim. BYTE Magazine, 20(12):170– ??, December 1, 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Baldazo:1995:SRW

[BVN95b] Rex Baldazo and Steven J. Vaughan-Nichols. Software roundup: Web publishing made easier — HTML authoring and editing tools promise to make your documents Web-ready in no time. but few live up to their claims. BYTE Magazine, 20 (12):170–??, ???? 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barrus:1997:QFM

[BW97] J. W. Barrus and R. C. Waters.

QOTA: A fast, multi-purpose algorithm for terrain following in virtual environments. In Spencer and Brutzman [SB97], pages 59–64. ISBN 0-89791-886-X. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/253437/.

ACM order number 434973.

Chen:1998:TOO

[C^{+}98] Jian Chen et al., editors. Technology of object-oriented languages and systems, TOOLS 24: proceedings: September, 1997, Beijing, China. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8551-4, 0-8186-8553-0. LCCN QA76.64.T66 1997a. IEEE Computer Society Order Number PR08551. IEEE Order Plan Catalog Number 97TB100240.

Coleman:1999:HC

[C^{+}99] Pat Coleman et al., editors.

HTML complete. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1999. ISBN 0-7821-2467-4. xxiv + 1028 pp. LCCN QA76.76.H94H763 1999.

Christophides:1994:SDN

[CACS94] V. Christophides, S. Abiteboul, S. Cluet, and M. Scholl. From structured documents to novel query facilities. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):313–324, June 1994. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cadogan:1987:CSS

[Cad87] Pat H. Cadogan. The Chelgraph SGML structured editor. In Earnshaw [Ear87], pages 190– 195. ISBN 0-387-96527-0, 3-540-96527-0. LCCN Z286.E43 W67 1987. US$29.50.

Casey:1995:WTM

[CAD^{+}95] James Casey, Alberto Aimar, Nikos Drakos, Ian Hannell, Arash Khodabandeh, Paolo Palazzi, Bertrand Rousseau, and Mario Ruggier. WebLinker: A tool for managing cross references between WWW documents. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Searching/casey/WLPaper_1.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Cailliau:1994:PWW

[Cai94a] R. Cailliau. Proceedings of the 1st World-Wide Web Conference. Computer Networks and ISDN Systems, 27(2):135– 341, November 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Cailliau:1994:SPW

[Cai94b] R. Cailliau, editor. Selected papers of the 1st World-Wide Web conference, 25–27 May 1994, Geneva, Switzerland, volume 27(2) of Computer Networks and ISDN Systems. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, November 1994. CODEN CNISE9. ISBN ???? ISSN 0169-7552 (print), 1879-2324 (electronic). LCCN ???? Published in Computer Networks and ISDN Systems, 27, No. 2.

Callihan:1997:LHW

[Cal97]  Steven E. Callihan.  Learn

HTML in a weekend. In a weekend. Prima Tech Publishing, Rocklin, CA, USA, 1997. ISBN 0-7615-1293-4 (paperback). xii + 419 pp. LCCN QA76.76.H94C35 1997.

Callihan:1998:LHW

[Cal98]  Steven E. Callihan.  Learn

HTML in a weekend. In a weekend. Prima Tech Publishing, Rocklin, CA, USA, revised edition, 1998. ISBN 0-7615-1800-2. xii + 409 pp. LCCN QA76.76.H94C35 1998.

Campbell:1994:RPD

[Cam94] J. Campbell. Robinson, Peter. The Digitization of Primary Textual Sources and The Transcription of Primary Textual Sources Using SGML. College and Research Libraries, 55 (6):568–??, 1994. CODEN CR-LIAI. ISSN 0010-0870.

Campbell:1998:VAC

[Cam98] D. A. Campbell. VRML in architectural construction documents: A case study. In Spencer and Roehl [SR98], pages 115–

120. ISBN 1-58113-022-8. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/271897/.

CSA:1989:IPT

[Can89] Canadian Standards Association and Standards Council of Canada. Information processing, text and office systems: Standard Generalized Markup Language (SGML), and SGML support facilities: SGML Document Interchange Format (SDIF). National standard of Canada CAN/CSA-Z243.210-89 [CSA standard], 0317-5669. Canadian Standards Association, Rexdale, ON, Canada, 1989. ISBN ???? viii + 164 pp. LCCN ???? URL http:

//www.iso.ch/cate/d16387.
html;   http://www.iso.ch/
cate/d16388.html; http://
www.iso.ch/cate/d16643.html;
 http://www.iso.ch/cate/d28557.
html; http://www.iso.ch/cate/
d30235.html.

CSA:1990:IPS

[Can90] Canadian Standards Association and Standards Council of Canada. Information processing: SGML support facilities: techniques for using SGML. National Standard of Canada CAN/CSA-Z243.210.1-89 [CSA standard] 0317-5669.

Canadian Standards Association, Rexdale, ON, Canada, 1990. ISBN ???? vii + 124 pp. LCCN ???? URL http:

//www.iso.ch/cate/d17319.
html;   http://www.iso.ch/
cate/d17330.html; http://
www.iso.ch/cate/d17332.html.

Cantrall:1995:OSA

[Can95] Dan Cantrall. The Oregon State Archives on the World Wide Web: A case study in building a new user base for government services. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
LocalCom/cantrall/cantrall.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Carr:1992:DI

[Car92] Richard Carr. Document interchange. DISC, BSi Standards, London, UK, 1992. ISBN 1-85554-182-3. ii + 44 pp. LCCN TK5105.577 .C37 199.

Carr:1994:IS

[Car94] Nick Carr. Implementations of SGML. Asean proprint: the graphic arts magazine, pages 44–??, April 1, 1994. ISSN 1036-9112.

Carl:1995:HH

[Car95a] Jeremy Carl. HTML helpers.

Internet World, 6(9):62–??, September 1, 1995. CODEN IERNE8. ISSN 1064-3923.

Carl:1995:HHW

[Car95b] Jeremy Carl. HTML HELPERS — A wave of new Web authoring wares and servers are sweeping into the market. Internet World, 6(9):62–??, ???? 1995. CODEN IERNE8. ISSN 1064-3923.

Carson:1997:SPP

[Car97] George S. Carson. Standards pipeline: PNG, VRML 97, BIIF, imaging standards. Computer Graphics, 31(3):18–20, August 1997. CODEN CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Carey:1998:NPCb

[Car98a] Patrick Carey. New perspectives on creating Web pages with HTML: comprehensive. New perspectives series. Course Technology, Cambridge, MA, USA, 1998. ISBN 0-7600-5533-5. various pp. LCCN ???? Includes CD-ROM.

Carey:1998:NPCa

[Car98b] Patrick Carey. New perspectives on creating Web pages with HTML: introductory. New perspectives series. Course Technology, Cambridge, MA, USA, 1998. ISBN 0-7600-5532-7. various pp. LCCN ????

Carr:1998:HHG

[Car98c] Heather Carr. HTML: a handson guide. Computer Literacy Press, Gilroy, CA, USA, 1998.

ISBN 1-57426-047-2. vii + 134 pp. LCCN ????

Carey:1999:NPC

[Car99] Patrick Carey. New perspectives on creating Web pages with HTML: comprehensive. New perspectives series. Course Technology, Cambridge, MA, USA, 1999. ISBN 0-7600-6484-9. various pp. LCCN ???? Includes CD-ROM.

Casson:1994:PHW

[Cas94]  Andrew Casson.  PLINTH,

HTML and the world wide web. Technical report, Artificial Intelligence Applications Institute, University of Edinburgh, Edinburgh, Scotland, 1994. 9 pp.

Castro:1996:HWWa

[Cas96a] Elizabeth Castro. HTML for the World Wide Web. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1996. ISBN 0-201-88448-3 (??invalid checksum??). xvi + 176 pp. LCCN QA76.76.H94 C37 1996. US$17.95, CDN$24.00.

Castro:1996:HWWb

[Cas96b] Elizabeth Castro. HTML For the World Wide Web: Visual QuickStart Guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, January 1996. ISBN 0-201-88448-8. xvi + 176 pp. LCCN QA76.76.H94C37 1996. US$17.95.

Castro:1996:HTC

[Cas96c]  Elizabeth Castro.  HTML

Training Combo for Macintosh. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, June 1996. ISBN 0-201-88682-0. ???? pp. LCCN ???? US$49.95.

Castro:1997:HWW

[Cas97] Elizabeth Castro. HTML for the World Wide Web: Visual Quickstart Guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, second edition, February 1997. ISBN 0-201-68862-X. 255 pp. LCCN QA 76.76 H94 C37 1997. US$17.95.

Castro:1998:HWW

[Cas98] Elizabeth Castro. HTML 4 for the World Wide Web. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1998. ISBN 0-201-69696-7. xi + 336 pp. LCCN QA76.76.H94C3698 1998.

Cave:1994:SOS

[Cav94] F. Cave. SGML and other standards — extended summary. In Blunden and Blunden [BB94], pages 145–147. ISBN 3-923352-25-5. LCCN ????

Chenevoy:1995:SAL

[CB95a] Y. Chenevoy and A. Belaid. A structural approach for library references recognition. Traitement du signal, 12(6):663–671, ???? 1995. CODEN TRSIE6. ISSN 0765-0019.

Claffy:1995:WTC

[CB95b] Kimberly C. Claffy and Hans-Werner Braun. Web traffic characterization: an assessment of the impact of caching documents from NCSA’s web server. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
DDay/claffy/main.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Cox:1995:RTI

[CB95c] Mark J. Cox and John E. F.

Baruch. Robotic telescopes: An interactive exhibit on the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Museum/cox/markcox.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Crespo:1996:WBB

[CB96] Arturo Crespo and Eric A. Bier.

WebWriter: A browser-based editor for constructing Web applications. Computer Networks and ISDN Systems, 28(7–11): 1291–1306, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=
1616.

Carey:1997:VAR

[CB97] Rikk Carey and Gavin Bell. The VRML 2.0 Annotated Reference Manual. Addison-Wes-ley, Reading, MA, USA, January 1997. ISBN 0-201-41974-2, 0-201-30945-9. xxi + 501 pp. LCCN QA76.76.H94C368 1997. US$39.99. URL http:

//www.best.com/~rikk/Book/

book.shtml. Includes CD-ROM.

Carr:1994:WUH

[CBDH94] L. A. Carr, D. W. Barron, H. C. Davis, and W. Hall. Why use HyTime? Electronic Publishing—Origination, Dissemination, and Design, 7(3):163– 178, September 1994. CODEN EPODEU. ISSN 0894-3982.

Charou:1998:PTS

[CBK98] E. Charou, J. Baron, and D. Kazepis, editors. People and technology: sharing knowledge and shaping the future of aquatic information in Europe: proceedings of the 7th Biennial Meeting of EURASLIC (European Association of Aquatic Sciences Libraries and Information Centres), 6–8 May 1998, National Centre for Marine Research, Athina, Greece, Proceedings of the Meeting — European Association of Aquatic Sciences Libraries and Information Centres EURASLIC 1998; 7th. National Centre for Marine Research, Athina, Greece, 1998. ISBN 960-85952-4-X. LCCN Z675.M35 E87 1998.

Crochetiere:1993:ADX

[CBO93] André Crochetière, Jean-Claude Boulanger, and Conrad Ouellon, editors. Actes du XVe Congrès international des linguistes: Québec, Université Laval, 9–14 aout 1992: les langues menacées = Proceedings of the XVth International Congress of Linguists: Quèbec, Université Laval, 9–14 August 1992: endangered languages, volume 1 of Proceedings of the International Congress of Linguists 1992. Les Presses de l’Universit’e Laval, Qu’ebec, PQ, Canada, 1993. ISBN 2-7637-7353-2. LCCN P21.I6 1992. Four volumes.

Chu:1996:HPI

[CC96] Kenny Chu and Francis Chin.

HTML publishing on the Internet. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-912100-4 (paperback). xx + 444 pp. LCCN QA76.76.H94 C58 1996.

Clarke:1997:URE

[CC97] Charles L. A. Clarke and Gordon V. Cormack. On the use of regular expressions for searching text. ACM Transactions on Programming Languages and Systems, 19(3): 413–426, May 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        1997-19-3/p413-clarke/.

Chaleat:1998:PHJ

[CC98] Philippe Chaleat and Daniel Charnay. Programmation HTML, Javascript. Eyrolles, Paris, France, 1998. ISBN 2-212-09024-2. 450 pp. LCCN ????

Chen:1999:MIP

[CC99] Chin-Tu Chen and Anne V. Clough, editors. Medical imaging 1999. Physiology and function from multidimensional images: 21–23 February 1999, San Diego, California, number 3660 in SPIE Proceedings. SPIE Optical Engineering Press, Bellingham, WA, USA, 1999. ISBN 0-8194-3132-X. ISSN 0277-786X (print), 1996-756X (electronic). LCCN RC78.7.D53 M45 1999.

Ceri:1999:XGL

[CCD^{+}99] Stefano Ceri, Sara Comai, Ernesto Damiani, Piero Fraternali, Stefano Paraboschi, and Letizia Tanca. XML-GL: a graphical language for querying and restructuring XML documents. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1171– 1187, May 17, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.elsevier.

com/cas/tree/store/comnet/
sub/1999/31/11-16/2140.pdf.

Carraro:1998:PBS

[CCEE98] G. U. Carraro, M. Cortes, J. T. Edmark, and J. R. Ensor. The Peloton bicycling simulator. In Spencer and Roehl [SR98], pages 63–70. ISBN 1-58113-022-8. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/271897/.

Campbell:1997:TYD

[CD97] Bruce Campbell and Rick Darnell. Teach yourself dynamic HTML in a week. Sams.net Pub., Indianapolis, IN, USA, 1997. ISBN 1-57521-335-4. xxvi + 508 pp. LCCN QA76.76.H94 C36 1997.

Cover:1991:PSS

[CDB91a] R. Cover, N. Duncan, and D. T. Barnard. The progress of SGML (Standard Generalized Markup Language): Extracts from a comprehensive bibliography. Literary and Linguistic Computing, 6(3):197–199, ???? 1991. CODEN LLCOEI. ISSN 0268-1145.

Cover:1991:BSS

[CDB91b] Robin Cover, Nicholas Duncan, and David T. Barnard. Bibliography on SGML (standard generalized markup language) and related issues. Technical report, Dept. of Computing and Information Science, Queen’s University at Kingston, Kingston, Ontario, Canada, 1991. 312 pp.

Campbell:1998:PDH

[CDCD98] Bruce Campbell, Rick Darnell, Pierre Couzy, and Jerome Duclos. Le programmeur Dynamic HTML. Le Programmeur. Simon & Schuster Macmillan (France), Paris, France, 1998. ISBN 2-7440-0370-0. viii + 503 pp. LCCN ????

Calvanese:1997:RRS

[CDL97] D. Calvanese, G. De Giacomo, and M. Lenzerini. Representing and reasoning on SGML documents. Lecture Notes in Computer Science, 1325:540– 549, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cearley:1996:DMH

[Cea96a] Kent Cearley. 7 Days to Master HTML: Become a Web Guru in One Short Week. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN 1-57169-066-2. xxxvi + 582 pp. LCCN QA76.76.H94 C43 1996. US$34.95.

Cearley:1996:HIC

[Cea96b] Kent Cearley. HTML 3 Interactive Course: Master HTML and Earn a Certificate of Achievement. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN 1-57169-066-2. xxxvi + 582 pp. LCCN QA76.76.H94C43 1996. US$39.99. URL http:/

                                        /w3.waite.com/waite/waite/
                                        books.new/HTML_3_Interactive_
                                        Course/html/h3iccov.html.

Cearley:1998:HIC

[Cea98] Kent Cearley. HTML 4 interactive course. Waite Group Press, Corte Madera, CA, USA, 1998. ISBN 1-57169-130-8. xxix + 606 pp. LCCN QA76.76.H94C44 1998.

Card:1996:ISI

[CEG96] Stuart Card, Stephen G. Eick, and Nahum Gershon, editors. IEEE Symposium on Information Visualization’96: October 28–29, 1996, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7668-X. LCCN TA1634 .I325 1996.

Celis:1992:DPA

[Cel92] Pedro Celis. Distribution, parallelism, and availability in nonstop SQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 21 (2):225, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Champeon:1999:BDH

[CF99] Steven Champeon and David S.

Fox. Building dynamic HTML GUIs. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1999. ISBN 0-7645-3267-7 (paperback). xxxi + 711 pp. LCCN QA76.9.U83C46 199.

Campbell:1995:IAC

[CFMG95] Paul Campbell, Lawrence Fitzpatrick, David C. Macdonald, and Adrienne Griffith. Information access: A cornerstone of Web publishing. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
WebProd/griffith/witehtml.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Chamberlin:1986:GAS

[CG86] D. D. (Donald Dean) Chamberlin and Charles F. Goldfarb. Graphic applications of the Standard Generalized Markup Language (SGML). Technical Report Research Report RJ 5440., IBM Corporation, Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1986. 33 pp.

Chamberlin:1987:GAS

[CG87] Donald D. Chamberlin and Charles F. Goldfarb. Graphics applications of the standard generalized markup language (SGML). Computers and Graphics, 11(4):343–358, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Catenazzi:1995:PPH

[CG95] N. Catenazzi and F. Gibb. The publishing process: the hyperbook approach. Journal of Information Science, Principles and Practice, 21(3):161–172, ???? 1995. CODEN JIOSED. ISSN 0165-5515 (print), 1741-6485 (electronic).

Cho:1995:IHI

[CGL^{+}95] Pagogh Cho, James Gutierrez, Anne LeMaistre, M. D., Cameron Payne, Russell Pruitt, and Charlotte Wilcox. IntelliTeaching — high impact learning. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
MedEd/cho/ncsa-cho.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Cowan:1994:ECR

[CGLV94] D. D. Cowan, D. M. German, C. J. P. Lucena, and A. Von Staa. Enhancing code for readability and comprehension using SGML. In Muller and Georges [MG94c], pages 181–190. ISBN 0-8186-6332-4, 0-8186-6330-8, 0-8186-6331-6. LCCN QA76.76.S64 I58 1994. IEEE Catalog No. 94CH3485-0.

Coggin:1996:SDS

[CGW96] W. O. Coggin, L. K. Gasser, and B. A. Williams. SGML document structuring: Implementing document analysis. In Anonymous [Ano96b], pages 265–??

ISBN 0-914548-89-1, 0-914548-90-5. LCCN ????

Carson:1990:CSM

[CH90] G. S. Carson and L. Henderson. The CALS standard (military computing). In Anonymous [Ano90h], pages 570–598 (vol. 1). ISBN ???? LCCN ???? 3 vol.

Cross:1995:UGM

[CH95] J. H. Cross and T. D. Hendrix. Using generalized markup and SGML for reverse engineering graphical representations of software. In Wills et al. [WNC95], pages 2–7 (or 2–6??). ISBN 0-8186-7111-4. LCCN TA 168.5 W67 1995. IEEE Catalog No. 95TB8101.

Chang:1998:CSD

[CH98] Daniel Chang and Dan Harkey. Client/server data access with Java and XML. Wiley, New York, NY, USA, 1998. ISBN 0-471-24577-1 (paperback/CD-ROM). xxvii + 606 pp. LCCN QA76.9.C55 C45 1998. US$50.

                                        URL http://www.wiley.com/
                                        compbooks/catalog/24577-1.
                                        htm.

Castro:1999:H

[CH99a] Elizabeth Castro and Servane Heudiard. HTML 4. Guide quick start. Editions First Interactive, Paris, France, 1999. ISBN 2-84427-072-7. 335 pp. LCCN ????

Ceponkus:1999:AXT

[CH99b] Alex Ceponkus and Faraz Hoodbhoy. Applied XML: a toolkit for programmers. Wiley, New York, NY, USA, 1999. ISBN 0-471-34402-8 (paperback / CD-ROM ). xx + 474 pp. LCCN QA76.76.H94 C46 1999.

Chang:1993:HUA

[Cha93] D. T. Chang. HieNet: a user-centered approach for automatic link generation. In ACM [ACM93b], pages 145–158. ISBN 0-89791-624-7. LCCN QA76.76.H94A26 1993.

Chang:1995:APD

[Cha95a] Wey-Tyng Chang. Automated product documentation generation in the product lifecycle process. In Anonymous [Ano95e], pages 85–94. ISBN ???? LCCN ????

Chapman:1995:CCB

[Cha95b] G. M. Chapman, editor. CBLIS 95. Computer Based Learning in Science Proceedings An International Conference. Open Educ. and Sci, Opava, Czech Republic, 1995. ISBN 80-901974-0-X. LCCN ????

Chahuneau:1996:WDS

[Cha96a] F. Chahuneau. What does SGML mean to your business? In Anonymous [Ano96v], pages 145–148. ISBN ???? LCCN ????

Chambers:1996:CTL

[Cha96b] Jack A. Chambers, editor. Selected papers from the 7th National Conference on College Teaching and Learning, Jacksonville, FL. Florida Community College at Jacksonville, Jacksonville, FL, USA, 1996. ISBN ???? LCCN ????

Chauvet:1999:CTC

[Cha99] Jean-Marie Chauvet. Composants et transactions: COM/ MTS, Corba/OTS, Java/EJB, XML. Collection dirigée par Guy Hervier. Eyrolles: Informatiques magazine, Paris, France, 1999. ISBN 2-212-09075-7. v + 274 pp. LCCN ????

Chen:1995:SRN

[Che95] Hsinchun Chen. The semantic retrieval for NCSA Mosaic. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Searching/chen/chenschatz.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Chenoweth:1997:WDC

[Che97] J. D. Chenoweth. What do computer science students need to know about HTML and the Internet? In Meinke [Mei97], pages 10–14. ISBN ???? LCCN ????

Chhabra:1995:ICD

[CHH^{+}95] Bhavna Chhabra, Darren R. Hardy, Allan Hundhausen, David Merkel, John D. Noble, and Michael F. Schwartz. Integrating complex data access methods into the Mosaic/ WWW environment. In Anonymous [Ano95-34], page ????

ISBN ???? LCCN ???? URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
schwartz/schwartz.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Chiu:1995:UMI

[Chi95] Wanda Chiu. Using Mosaic and the Internet: an interactive guide for junior high students. Technical report TR 95-06, University of Arizona, Dept. of Computer Science, Tucson, AZ, USA, July 7, 1995. 74 pp.

Chamberlin:1988:DDS

[CHP88] Donald D. Chamberlin, Helmut F. Hasselmeier, and Dieter P. Paris. Defining document styles for WYSIWYG processing. In van Vliet [van88], pages 121–137. ISBN 0-521-36294-6. LCCN Z286.E43 I571 1988. US$47.50.

Christoffersen:1995:PPP

[Chr95] O. Christoffersen. The POL project: a practical approach using SGML. In Anonymous [Ano95e], pages 173–183. ISBN ???? LCCN ????

Carr:1996:OIS

[CHR^{+}96a] Leslie Carr, Gary Hill, David De Roure, Wendy Hall, and Hugh Davis. Open information services. Computer Networks and ISDN Systems, 28(7–11):1027– 1036, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1996&
                                        volume=28&issue=7-11&aid=
                                        1603.

Christian:1996:GPW

[Chr96b] P. Christian. Genealogical publishing on the World Wide Web.

2. Computers in Genealogy, 5 (10):437–442, June 1996. CO-DEN CGENER. ISSN 0263-3248.

Chung:1992:MIC

[Chu92] Kwang-Soo Chung. Multimedia information communication standardization. Korea Information Science Society Review, 10(5):10–19, ???? 1992. CO-DEN CHKWEN. ISSN 1015-9908.

Chute:1998:AMI

[Chu98] C. G. Chute, editor. American Medical Informatics Association: A paradigm shift in health care information systems: Annual symposium — November 1998, Orlando, FL, AMIA Annual Symposium 1998. Hanley and Belfus, Philadelphia, PA, USA, 1998. ISBN ???? ISSN 1091-8280. LCCN ????

Chenney:1998:EDM

[CIF98] S. Chenney, J. Ichnowski, and D. Forsyth. Efficient dynamics modeling for VRML and Java. In Spencer and Roehl [SR98], pages 15–24. ISBN 1-58113-022-8. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/271897/.

Cimino:1996:BSE

[Cim96] J. J. Cimino, editor. Beyond the superhighway: exploiting the Internet with medical informatics: proceedings, 1996 AMIA annual Fall Symposium, formerly SCAMC, October 26– 30, 1996, Sheraton Washington Hotel, Washington, DC. Hanley and Belfus, Philadelphia, PA, USA, 1996. ISBN 1-56053-208-4. LCCN W 26.5 S978 1996.

Cruz:1990:RDM

[CJ90] G. C. Cruz and T. H. Judd.

The role of a descriptive markup language in the creation of interactive multimedia documents for customized electronic delivery. In Furuta [Fur90], pages 277–290. ISBN 0-521-40246-8. LCCN Z286.E43I58 1990.

Crowcroft:1994:TNG

[CJ94] Jon Crowcroft and Gordon Joly.

Touring and navigating a global learning environment — the Web. Computer Networks and ISDN Systems, 26(Supplement 2–3):S71–S74, ???? 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Comon:1995:TRF

[CJ95] H. Comon and J.-P. Jouannaud, editors. Term Rewriting. French Spring School of Theoretical Computer Science. Advanced Course. Springer-Verlag Inc., New York, NY, USA, 1995. ISBN 3-540-59340-3. LCCN QA75.5.F74 1993.

Coors:1998:UVI

[CJ98] V. Coors and V. Jung. Using VRML as an interface to the 3D data warehouse. In Spencer and Roehl [SR98], pages 121– 127. ISBN 1-58113-022-8. LCCN ???? URL http://

                                        www.acm.org/pubs/contents/
                                        proceedings/graph/271897/.

Colby:1996:SEU

[CJD^{+}96] Martin Colby, David S. Jackson, Steven J. DeRose, et al. Special Edition Using SGML: Special Edition. Que Corporation, Indianapolis, IN, USA, dlx edition, March 1996. ISBN 0-7897-0414-5. xxv + 590 pp. LCCN QA76.73.S44C65 1996¡. US$49.99.

Cole:1995:SCD

[CK95] Timothy W. Cole and Michelle M. Kazmer. SGML as a component of the digital library. Library Hi Tech, 13(4):75–90, ???? 1995. CODEN LIHTD2. ISSN 0737-8831.

Carey:1997:SEA

[CK97] Michael J. Carey and Donald Kossmann. On saying “Enough Already!” in SQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):219–230, June 1997. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Choros:1998:TWR

[CK98] K. Choros and L. Krzywiecki.

Tworzenie wirtualnej rzeczywistosci sieciowej za pomoca specyfikacji VRML na przykladzie multimedialnego systemu “Zabytki wroclawia”. In Danilowicz [Dan98a], pages 101– 108. ISBN 83-7085-371-4. LCCN ????

Crawford:1990:PFA

[CL90] R. G. Crawford and S. Lee. A prototype for fully automated entry of structured documents. Canadian Journal of Information Science, 15(4):39–50, December 1990. CODEN CJISDE. ISSN 0380-9218.

Christiansen:1997:PNS

[CL97] C. Christiansen and C. Leathem, editors. Pioneering new serials frontiers: from petroglyphs to cyberserials: proceedings of the North American Serials Interest Group, Inc., 11th Annual Conference, June 20–23, 1996, University of New Mexico, Albuquerque, volume 30(3/4), 31(1/2) of Serials Librarian. Haworth Press, New York, NY, USA, 1997. ISBN 0-7890-0324-4. ISSN 0361-526X. LCCN Z692.S5N67 1996.

Chen:1999:DSA

[CL99] Arbee L. P. Chen and Frederick H. Lochovsky, editors. Database systems for advanced applications: proceedings 6th International Conference on Database Systems for Advanced Applications: April 19–21, 1999, Hsinchu, Taiwan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN 0-7695-0084-6, 0-7695-0085-4 (casebound), 0-7695-0086-2 (microfiche). LCCN QA76.9.D3 I5565 1999.

Clark:1998:NBW

[Cla98] David Clark. News briefs; W3C approval gives XML a push; ISO adopts QuickTime for MPEG-4; group proposes antipiracy plan; delay for Gbit Ethernet specs. Computer, 31(4):18– 19, April 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1998/pdf/r4018.
                                        pdf.

Campbell:1996:SHD

[CLK96] Scott Campbell, Josh Linder, and Robert J. Kohlhepp. Serving up HTML documents on the Web. Network Computing, 7(1): 92–??, January 1996. CODEN NCOMEV. ISSN 1046-4468.

Casher:1996:AVB

[CLPR96] O. Casher, C. Leach, C. S. Page, and H. S. Rzepa. Advanced VRML based chemistry applications: a 3D molecular hyperglossary. In Bachrach [Bac96], pages 49–56. ISBN ???? ISSN 0166-1280 (print), 1872-7999 (electronic). LCCN ????

Capps:1996:EGM

[CLS96] Michael Capps, Brian Ladd, and David Stotts. Enhanced graph models in the Web: multiclient, multi-head, multi-tail browsing. Computer Networks and ISDN Systems, 28(7–11): 1105–1112, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=
1589.

Cohen:1990:RBD

[CM90] S. S. Cohen and W. E. Morgan.

The relationship between the DECwrite editor and the Digital document interchange format. Digital Technical Journal, 2(1):73–82, Winter 1990. CO-DEN DTJOEL. ISSN 0898-901X.

Corral:1995:HRD

[CM95] M.-L. Corral and J. Mothe.

How to retrieve and display long structured documents? In IEEE [IEE95c], pages 10–19. ISBN 0-8186-7110-6. LCCN QA76.9.D3 B37 1995.

Carlson:1996:EMH

[CM96] Patricia Carlson and Fillia Makedon, editors. Educational Multimedia and Hypermedia, 1994: proceedings of ED-MEDIA 96 — World Conference on Educational Multimedia and Hypermedia, Boston, Mass., USA, June 17–22, 1996. Association for the Advancement of Computing in Education, Charlottesville, VA, USA, 1996. ISBN 1-880094-21-5. LCCN ????

Chu:1993:DIC

[CMB93] Wesley W. Chu, Matthew A. Merzbacher, and Ladislav Berkovich. The design and implementation of CoBase. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2): 517–522, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:1997:BEL

[CMP97] S. Chen, R. Myers, and R. Pasetto. The out of box experience: Lessons learned creating compelling VRML 2.0 content. In Spencer and Brutzman [SB97], pages 83–92. ISBN 0-89791-886-X. LCCN ????

URL http://www.acm.org/
pubs/contents/proceedings/
graph/253437/. ACM order

number 434973.

Cowan:1991:REU

[CMPd91] D. D. Cowan, E. W. Mackie, G. M. Pianosi, and G. de V. Smit. Rita—an editor and user interface for manipulating structured documents. Electronic Publishing—Origination, Dissemination, and Design, 4 (3):125–150, September 1991. CODEN EPODEU. ISSN 0894-3982.

Ciancarini:1998:VZN

[CMV98] P. Ciancarini, C. Mascolo, and F. Vitali. Visualizing Z notation in HTML documents. In Bowen et al. [BFH98], pages 81– 95. ISBN 3-540-65070-9. LCCN QA267.A1 L43 no.1493.

Chaudhuri:1998:MIT

[CN98] Surajit Chaudhuri and Vivek Narasayya. Microsoft index tuning wizard for SQL Server 7.0. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):553–554, June 1998. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chen:1999:BMV

[CNH99] A. C. N. Chen, F. A. Nielsen, and L. K. Hansen. A 3-D brain model of VRML/ talairach for standardized data-sets in neuroimaging of human pain: PET data in headache vs. angina (demo). In Rosen et al. [RSV99], page S183. ISBN ???? ISSN 1053-8119. LCCN ????

Codognet:1998:EVI

[Cod98] P. Codognet. Extending VRML for intelligent virtual agents. In Landauer and Bellman [LB98a], page 241. ISBN 1-56555-137-0. ISSN 0735-9276. LCCN QA76.9.V5 V578 1998.

Coe:1996:WIH

[Coe96] M. A. Coe. Webspeak: An introduction to HTML and the WWW. In Anonymous [Ano96b], pages 406–?? ISBN 0-914548-89-1, 0-914548-90-5. LCCN ????

Cohen:1995:RBI

[Coh95] Ellis S. Cohen. Review-based information services: Lessons learned from The Boston Restaurant List. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
InfOrg/cohen/info-serv.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Cohen:1996:CPJ

[Coh96]  Isaac Cohen.  CGI/Perl et

JavaScript: creation de pages HTML interactives. Eyrolles, Paris, France, 1996. ISBN 2-212-08918-X. vii + 294 pp. LCCN ???? 185 FF. URL http:

//www.estp.fr/eyrolles. In-

cludes floppy disk.

Cohen:1999:WEH

[Coh99] J. Cohen. Web-based everything: are HTTP, DAV and XML enough? In Anonymous [Ano99i], pages 1756– 1757. ISBN 0-444-50264-5. ISSN 1389-1286 (print), 1872-7069 (electronic). LCCN TK5105.888 .I573 1999.

Coleman:1995:IIH

[Col95a] J. P. H. Coleman. An investigation into integrating hypertext and messengers. In Poo and Seumahu [PS95], pages 181–184. ISBN 0-7803-2579-6. LCCN TK5105.2 .S5 1995. IEEE Catalog No. 95TH8061.

Coleman:1995:WAU

[Col95b] J. W. Coleman. Web authoring unmasked: two tidy tools to weave your own Web magic. Info Canada, 20(11):30, 32, November 1995. CODEN IFCAE3. ISSN 1187-7081.

Collier:1995:PIC

[Col95c] H. Collier, editor. Proceedings of the 1995 International Chemical Information Conference. Infonortics, Calne, UK, 1995. ISBN ???? LCCN ????

Collins:1995:CRW

[Col95d] Julia A. Collins. Climate research and the Web: Integrating World Wide Web and Mosaic capabilities into the climate diagnostics center. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
EarSci/collins/collins.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Collier:1996:MLA

[Col96a] M. W. Collier. A model license for acquisition of electronic materials. In Helal and Weiss [HW96b], pages 51–67. ISBN 3-922602-21-5. LCCN Z678.9.A1 I52 1996.

Colner:1996:HPO

[Col96b] Don Colner. HTML programming with objects. C/C++ Users Journal, 14(7):17–18, 21– 24, 26–28, 29–30, July 1996. CODEN CCUJEX. ISSN 1075-2838.

Collier:1998:ICI

[Col98] H. Collier, editor. International chemical information conference: – October 1998, Nimes, France, Proceedings of the International Chemical Information Conference 1998. Infonortics Ltd, ????, 1998. ISBN 1-873699-58-1. LCCN ????

Comes:1996:IMA

[Com96] R. Comes. Is the Internet the mother of all networks? Sistemi and Impresa, 42(3):17–20, April 1996. CODEN SIIMEU.

ComputerPREP:1997:HWPa

[Com97a] ComputerPREP, Inc. HTML and web page design module I: instructor’s guide. Courseware series. ComputerPREP, Phoenix, AZ, USA, 1997. various pp.

ComputerPREP:1997:HWPb

[Com97b] ComputerPREP, Inc. HTML and web page design module II: instructor’s guide. Courseware series. ComputerPREP, Phoenix, AZ, USA, 1997. various pp.

Correia:1996:WEM

[COMG96] N. Correia, I. Oliveira, J. Martins, and N. Guimaraes. WeatherDigest: an experiment on media conversion. In Tescher and Bove, Jr. [TB96a], pages 50–59. CODEN PSISDG. ISBN 0-8194-1979-6. ISSN 0277-786X (print), 1996-756X (electronic). LCCN QA76.575 .I576 1996.

Conway:1993:PGP

[Con93] A. Conway. Page grammars and page parsing. A syntactic approach to document layout recognition. In IEEE [IEE93b], pages 761–764. ISBN 0-8186-4960-7. LCCN TA1640.I57 1992. IEEE catalog no. 93TH0578-5.

Connolly:1994:EHS

[Con94] Dan Connolly. Evolution of the HTML spec. Dr. Dobb’s Journal of Software Tools, 19(15): 20, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Connolly:1996:LAH

[Con96] D. Connolly. A lexical analyzer for HTML and basic SGML. In Anonymous [Ano96-46], pages 43–64. ISBN 1-56592-210-7. ISSN 1085-2301. LCCN ???? US$24.95.

Connolly:1997:XPT

[Con97] Dan Connolly, editor. XML:

Principles, Tools, and Techniques, volume 2(4) of The World Wide Web Journal. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Winter 1997. ISBN 1-56592-349-9. ISSN 1085-2301. ix + 248 pp. US$29.95.

                                        URL http://www.amazon.com/
                                        exec/obidos/ISBN%3D1-56592-
                                        349-9/8788-6902305-556196;
                                         http://www.ora.com/catalog/
                                        wjfall97/.

Copp:1996:WA

[Cop96]  S. Copp.  Working with

academe. In ACM [ACM96d], pages 1–2. ISBN 0-89791-799-5. LCCN QA 76.9 D6 I56 1996.

Courcelle:1995:GRB

[Cou95] B. Courcelle. Graph rewriting: a bibliographical guide. In Comon and Jouannaud [CJ95], page 74. ISBN 3-540-59340-3. LCCN QA75.5.F74 1993.

Couch:1997:VSD

[Cou97]  J. Couch.  Vermlgen soft-

ware distribution. Virtual Light Company, December 1997.

                                        URL http://www.vlc.com.au/
                                        JVerge.

Cover:1996:CTM

[Cov96] R. Cover. Into the crucible: Test the merits of hierarchical models, embedded markup, and monolithic SGML DTDs in light of conceptual models of text. In Anonymous [Ano96f], pages 106–107. ISBN 82-7283-083-3. ISSN 0804-8800. LCCN ????

Covaci:1999:ANF

[Cov99] Stefan Covaci, editor. Active networks: First International Workshop, IWAN’99, Berlin, Germany, June 30–July 2, 1999: proceedings, volume 1653 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. ISBN 3-540-66238-3. LCCN TK5105.5 .I95 1999 Bar.

Casari:1995:GIA

[COVS95] G. Casari, C. Ouzounis, A. Valencia, and C. Sander. Genequiz II: automatic function assignment for genome sequence analysis. In Hunter and Klein [HK95b], pages 707–709. ISBN 981-02-2578-4. LCCN QH323.5.P33 1996.

Callaos:1998:WMS

[COW98] Nagib Callaos, Olu Omolayole, and Lipo Wang, editors. World multiconference on systemics, cybernetics ad informatics: proceedings, ISAS ’98: (4th International Conference on Information systems, Analysis and synthesis). July 12–16, 1998 Orlando, Florida, volume 1. International Institute of Informatics and Systemics, Orlando, FL, USA, 1998. ISBN 980-07-5078-9 (vol. 1), 980-07-5079-7 (vol. 2), 980-07-5080-0 (vol. 3), 980-07-5081-9 (vol. 4). LCCN ???? Four volumes.

Cowan:1999:XDG

[Cow99] John Cowan. XML: a developer’s guide. McGraw-Hill, New York, NY, USA, 1999. ISBN 0-07-135124-8. 600 pp. LCCN ????

Cox:1996:HLG

[Cox96] A. Cox. Hypermedia library guides for academic libraries on the World Wide Web. Program, 30(1):39–50, January 1996. CO-DEN PRGMBD. ISSN 0033-0337.

Coyne:1997:DNP

[Coy97] Richard Coyne, editor. Design and the net: proceedings of the Sixth International EuropIA Conference Applications of Computer Networking in Architecture, Construction, Design, Civil Engineering and Urban Planning: EuropIA’97 2– 3 April 1997. EuropIA Productions, Paris, France, 1997. ISBN 2-909285-07-3. LCCN ????

Chelsom:1993:DSP

[CP93] J. Chelsom and M. Pengelly. Database support for the production of reusable technical documentation. In Anonymous [Ano93h], pages 10/1–2. ISBN ???? LCCN ????

Chelsom:1994:USP

[CP94] J. J. L. Chelsom and M. Pengelly. Using SGML to produce reusable technical documentation. In Anonymous [Ano94b], pages 6/1–6/3. ISBN ???? LCCN ????

Corthouts:1996:SLP

[CP96] J. Corthouts and R. Philips. SGML: a librarian‘s perception. Electronic Library, 14(2):101– 110, April 1996. CODEN EL-LIDZ. ISSN 0264-0473.

Chua:1995:MMT

[CPK95] Tat-Seng Chua, Hung Keng Pung, and Tosiyasu L. Kunii, editors. Multimedia modeling: towards information superhighway: Singapore, 14–17 November 1995. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1995. ISBN 981-02-2502-4. LCCN QA76.575.M834 1995.

Crutcher:1998:VCE

[CPR98] R. M. Crutcher, R. L. Plante, and P. Rajlich. VRML and collaborative environments: New tools for networked visualization. In Albrecht et al. [AHB98], pages 3–6. ISBN 1-886733-65-1. ISSN 1080-7926. LCCN QB51.3.E43 A87 1997.

Chu:1995:CHM

[CPW95] Josey Y. M. Chu, William L.

Palya, and Donald E. Walter. Creating a HyperText Markup Language document for an information server. Behavior research methods, instruments, and computers, 27(2): 200–??, 1995. CODEN BRM-CEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Christophides:1994:QSD

[CR94] V. Christophides and A. Rizk.

Querying structured documents with hypertext links using OODBMS. In ACM [ACM94b], pages 186–197. ISBN 0-89791-640-9. LCCN QA76.76.I59A35

Chisholm:1995:EVT

[CR95a] D. Chisholm and D. Robey. Encoding verse texts. Computers and the humanities, 29(2):99– 111, ???? 1995. CODEN CO-HUAD. ISSN 0010-4817.

Cover:1995:ETC

[CR95b] R. C. Cover and P. M. W. Robinson. Encoding textual criticism. Computers and the humanities, 29(2):123–136, ???? 1995. CODEN COHUAD. ISSN 0010-4817.

Casher:1996:MOT

[CR96] O. Casher and H. S. Rzepa. The molecular object toolkit: A new generation of VRML visualisation tools for use in electronic journals. In Jones et al. [JRV96], pages 173–178. ISBN 0-9521097-3-5 (paperback),0-9521097-4-3 (vol. 1), 0-9521097-5-1 (vol. 2). LCCN ???? Two volumes.

Charitos:1997:WAN

[CR97] D. Charitos and P. Rutherford. Ways of aiding navigation in VRML worlds. In Coyne [Coy97], pages 119–132. ISBN 2-909285-07-3. LCCN ????

Chueh:1998:XPC

[CRBB98] H. C. Chueh, W. F. Raila, D. A. Berkowicz, and G. O. Barnett. An XML portable chart format. In Chute [Chu98], pages 730– 734. ISBN ???? ISSN 1091-8280. LCCN ????

Crossman:1992:SSC

[Cro92] R. Crossman. Standards for storing compound documents.

Info DB, 7(1):17–21, Winter 1992. CODEN IFDBEB. ISSN 0891-6004.

Crocker:1995:WSI

[Cro95a] Glenn A. Crocker. web2mush:

Serving interactive resources to the Web. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
crocker/tech.html. A selected

few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Crossley:1995:WTW

[Cro95b] David Crossley. WAIS through the Web — discovering environmental information. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Searching/crossley/paper.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Ciancarini:1998:ERE

[CRV98] Paolo Ciancarini, Alfredo Rizzi, and Fabio Vitali. An extensible rendering engine for XML and HTML. Computer Networks and ISDN Systems, 30 (1–7):225–237 (or 225–238??), April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

                                        www.elsevier.com/cas/tree/
                                        store/comnet/sub/1998/30/
                                        1-7/1926.pdf.

Crandall:1996:IEI

[CS96] Mike Crandall and Mark C. Swenson. Integrating electronic information through a corporate Web. Computer Networks and ISDN Systems, 28(7–11): 1175–1186, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1996&
                                        volume=28&issue=7-11&aid=
                                        1593.

Cutler:1997:USH

[CSM97] M. Cutler, Y. Shi, and W. Meng. Using the structure of HTML documents to improve retrieval. In USENIX [USE97b], pages 241–252. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

Cole:1996:PGW

[CT96] R. Cole and S. Tooker. Physics to go: Web-based tutorials for CoLoS physics simulations. In Iskander [Isk96], pages 3– ?? ISBN 0-7803-3348-9 (softbound), 0-7803-3349-7 (casebound), 0-7803-3350-0 (microfiche), 0-7803-3720-4 (CD-ROM). LCCN T 62 F76 1996. Three volumes. IEEE catalog number 96CH35946.

Chisholm:1998:RAA

[CV98] W. Chisholm and G. C. Vanderheiden. Results of accessibility analysis of HTML and the implications for future information technologies. In Anonymous [Ano98r], pages 1028– 1032. ISBN 0-945289-11-1. ISSN 1071-1813. LCCN TA166.A2 .H8 1998. Two volumes.

Clucas:1995:IVC

[CW95] Jean Clucas and Velvin (Val)

Watson. Interactive visualization of computational fluid dynamics using Mosaic. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ???? URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
clucas/clucas.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Coleman:1997:SFD

[CW97] James Coleman and Don Willis.

SGML as a framework for digital preservation and access. Report (Commission on Preservation and Access). Commission on Preservation and Access, Washington, DC, 1997. ISBN 1-887334-51-3. 47 pp. LCCN Z681.3.D53 C6 1997. “July 1997.”.

Czerminski:1998:DPD

[Cze98] Jurand B. Czerminski, editor.

Document publishing and delivery: proceedings of the 21st Library Systems Seminar held in Gdansk on 18–20 June 1997. University of Gdansk, Gdansk, Poland, 1998. ISBN 83-86230-35-5. LCCN ????

Darnell:1999:HU

[D^{+}99] Rick Darnell et al. HTML 4 unleashed. SAMS Publishing, Indianapolis, IN, USA, second edition, 1999. ISBN 0-672-31347-2 (paperback). xxxii + 1285 pp. LCCN QA76.76.H94H765 1999.

Dolin:1997:SMI

[DABM97] R. H. Dolin, L. Alschuler, T. Bray, and J. E. Mattison. SGML as a message interchange format in healthcare. In Masys [Mas97], pages 635–639. ISBN ???? ISSN 1091-8280. LCCN W26.5 S978 1997.

Dogac:1994:MOO

[DAE^{+}94] Asuman Dogac, Budak Arpinar, Cem Evrendilek, Cetin Ozkan, Ilker Altintas, Ilker Durusoy, Mehmet Altinel, Tansel Okay, and Yuksel Saygin. METU object-oriented DBMS. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23(2):513, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Danilowicz:1998:MSS

[Dan98a] C. Danilowicz, editor. Multimedialne i sieciowe systemy informacyjne: Konferencyjne — September 1998, Wroclaw, Poland. OWPW, ????, 1998.

ISBN 83-7085-371-4. LCCN ????

Dann:1998:CWS

[Dan98b] Ru-Tsuen Dann. A 3D commercial web site of an existing cinema theater: the application of VRML (Virtual Reality Modeling Language) in architecture. ????, ????, 1998. 53 pp.

Dogac:1995:IAO

[DAOD95] Asuman Dogac, Mehmet Altinel, Cetin Ozkan, and Ilker Durusoy. Implementation aspects of an object-oriented DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):9–14, March 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Darling:1995:LRA

[Dar95] C. B. Darling. Leaping into remote access (server acquisition). Datamation, 41(7):59–66, April 1995. CODEN DTMNAT. ISSN 0011-6963.

Darnell:1997:HU

[Dar97] Rick Darnell. HTML 4 unleashed. Sams.net Pub., Indianapolis, IN, USA, 1997. ISBN 1-57521-299-4 (paperback). xxxvii + 1030 pp. LCCN QA76.76.H94 H764 1997.

Darnell:1998:HU

[Dar98] Rick Darnell. HTML 4 unleashed. Sams.net Pub., Indianapolis, IN, USA, professional reference edition, 1998. ISBN 1-57521-380-X. xlii + 1195 pp. LCCN ????

David:1992:ACO

[Dav92] Michael M. David. Advanced capabilities of the outer join (SQL). SIGMOD Record (ACM Special Interest Group on Management of Data), 21(1):65–70, March 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Davidson:1993:SAT

[Dav93] W. J. Davidson. SGML authoring tools for technical communication. Technical Communication, 40(3):403–409, August 1993. CODEN TLCMBT. ISSN 0049-3155.

Davison:1995:CIH

[Dav95a] Andrew Davison. Clickable images in HTML. Dr. Dobb’s Journal of Software Tools, 20(9):18– 20, 22, 24, 27, 98–100, September 1995. CODEN DDJOEB. ISSN 1044-789X.

Davison:1995:CHF

[Dav95b] Andrew Davison. Coding with HTML forms. Dr. Dobb’s Journal of Software Tools, 20(6):70, 72–75, 106–109, June 1995. CO-DEN DDJOEB. ISSN 1044-789X.

Davidson:1998:HFW

[Dav98] B. H. Davidson. HTML frames: Web design and access issues. In Nixon et al. [NDM98], pages 104–106. ISBN 1-57387-067-6. LCCN Z674.82.I59 I585 1998.

Dawson:1998:BRS

[Daw98a]  Terry Dawson.  Book re-

view: Structuring XML Documents. Linux Journal, 55:94–95, November 1998. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Dawson:1998:SCC

[Daw98b] Terry Dawson. SGML CD: A complete SGML toolkit. Linux Journal, 47:??, March 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Dobson:1995:LD

[DB95] S. A. Dobson and V. A. Burrill.

Lightweight databases. Computer Networks and ISDN Systems, 27(6):1009–1015, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/
store/comnet/sub/1995/27/
6/1436.pdf.

Douglis:1996:TVC

[DB96] F. Douglis and T. Ball. Tracking and viewing changes on the Web. In USENIX Association [USE96], pages 165–176. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

Douglis:1996:WQN

[DBCK96] Fred Douglis, Thomas Ball, Yih-Farn Chen, and Eleftherios Koutsofios. WebGUIDE: querying and navigating changes in Web repositories. Computer Networks and ISDN Systems, 28 (7–11):1335–1344, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=
1612.

DeHon:1995:GCC

[DBE^{+}95] Andre’ DeHon, Jeremy Brown, Ian Eslick, Jake Harris, Lara Karbiner, and Thomas F. Knight, Jr. Global cooperative computing. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
dehon.global/gcc_www94.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

DellaMea:1996:HGS

[DBRB96] V. Della Mea, C. A. Beltrami, V. Roberto, and D. Brunato. HTML generation and semantic markup for telepathology. In Anonymous [Ano96-34], pages 1085–1094. CODEN CNISE9. ISBN ???? ISSN 0169-7552 (print), 1879-2324 (electronic). LCCN TK 5105.5 C647 v.28 no.7-11 1996.

Dubey:1998:UEM

[DC98] A. K. Dubey and H. C. Chueh. Using the extensible markup language (XML) in automated clinical practice guidelines. In Chute [Chu98], pages 735–739. ISBN ???? ISSN 1091-8280. LCCN ????

Davis:1995:ICB

[DCG^{+}95] Nancy B. Davis, Pagogh Cho, James Gutierrez, Cameron Payne, Russell Pruitt, Charlotte Wilcox, and Anne LeMaistre. Interactive computer based programs versus traditional methods for self-study. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
MedEd/davis/nbdavis.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

DeRose:1994:MHW

[DD94] Steven J. DeRose and David G.

Durand. Making hypermedia work: a user’s guide to HyTime. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-9432-1. xxii + 384 pp. LCCN QA76.76.I59 D47 1994. UK£51.00, US$68.00.

Derose:1996:MHW

[DDM96] Steven J. Derose, David G.

Durand, and Gregory Murphy. Making hypermedia work: A user’s guide to HyTime. Computers and the humanities, 30 (1):93–??, ???? 1996. CODEN COHUAD. ISSN 0010-4817.

DeLaBeaujardiere:1988:WDI

[De 88] Jean-M. De La Beaujardiere. Well-established document interchange formats. In van Vliet [van88], pages 83–94. ISBN 0-521-36294-6. LCCN Z286.E43 I571 1988. US$47.50.

DeKemp:1994:NWD

[De 94] A. De Kemp. New ways for the delivery of information: IT, networks, SGML, multimedia and so on …. In Helal and Weiss [HW94], pages 208–?? ISBN 3-922602-18-5. ISSN 0931-7503. LCCN Z672.5 .I58 1994.

deGroot:1998:MFF

[de 98] Cees de Groot. More flexible formatting with SGMLtools. Linux Journal, 55:66, 68–69, November 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Debreceny:1995:AID

[DE95] Roger Debreceny and Allan Ellis, editors. AusWeb95, innovation and diversity: the World Wide Web in Australia: proceedings of AusWeb95, the First Australian World Wide Web Conference, Ballina, NSW, Australia, 30 April to 2 May 1995. Norsearch, Southern Cross University, Lismore, NSW, Australia, 1995. ISBN 0-646-23946-5. LCCN ????

Debreceny:1996:ALW

[DE96] Roger Debreceny and Allan Ellis, editors. AusWeb96: landscaping the web: proceedings of AusWeb96, the second Australian World Wide Web Conference, Gold Coast, Queensland, 7–9 July 1996. Southern Cross University Press, Lismore, NSW, Australia, 1996. ISBN 1-875855-16-5. LCCN ????

Debbabi:1997:MBC

[Deb97] Maurad Debbabi. A modelbased concurrent specification language over CML: Semantic foundations. Parallel Processing Letters, 7(3):329–??, September 1997. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Deitz:1996:ILE

[Dei96] Dan Deitz. Industry leaders endorse Java and VRML. Mechanical Engineering: the journal of the American Society of Mechanical Engineers, 118(??): 28–??, ???? 1996. CODEN MEENAH. ISSN 0025-6501.

Delouis:1995:BWH

[Del95a] D. Delouis. Beyond WWW and HTML — standards for cultural heritage information: The RAMA perspective. In Bearman [Bea95b], pages 281–282. ISBN 1-885626-12-6. LCCN AM139 .H36 1995.

Delp:1995:RDL

[Del95b] E. J. Delp. The role of digital libraries on the information superhighway. In Anonymous [Ano95k], page 78. ISBN 0-89208-184-8. LCCN TA1632.I83

Dell:1998:DHW

[Del98] Tom Dell. Dynamic HTML for WebMasters. AP Professional, Boston, MA, USA, 1998. ISBN 0-12-209065-9 (paperback). ???? pp. LCCN QA76.76.H94D45 1998.

Dell:1999:DHW

[Del99] Tom Dell. Dynamic HTML for WebMasters. AP Professional, Boston, MA, USA, 1999. ISBN 0-12-209065-9 (paper), 0-12-209066-7 (CD-ROM). xxi + 336 pp. LCCN QA76.76.H94 D45 1999.

DeRose:1997:SFB

[DeR97] Steven J. DeRose. The SGML FAQ book: understanding the foundation of HTML and XML. Electronic publishing series. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-9943-9. xxvi + 250 pp. LCCN QA76.76.H94D47 1997.

Derose:1999:XT

[Der99a] S. Derose. XML and the TEI. In Mylonas and Renear [MR99b], pages 11–30. ISBN ???? ISSN 0010-4817. LCCN ????

DeRose:1999:XL

[DeR99b] Steven J. DeRose. XML linking. ACM Computing Surveys, 31(4es):??, December 1999. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

                                        tronic). URL http://www.acm.
                                        org/pubs/citations/journals/

surveys/1999-31-4es/a21-derose/

.

Desarmenien:1986:TSD

[Dés86] Jacques Désarménien, editor.

TEX for Scientific Documentation. Second European Conference. Strasbourg, France, June 19–21, 1986. Proceedings, number 236 in Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1986. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN Z253.4.T47 T49 1986.

Deutsch:1995:OSP

[Deu95] Donald R. Deutsch. Objects and SQL (panel session): strange relations? SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (2):466, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

DeVito:1996:GAA

[DeV96] A. F. DeVito. A grammardirected approach to automated SGML tagging: The Thesaurus Linguae Latinae. In Anonymous [Ano96f], pages 64–66. ISBN 82-7283-083-3. ISSN 0804-8800. LCCN ????

Deutsch:1999:QLX

[DFF^{+}99] Alin Deutsch, Mary Fernandez, Daniela Florescu, Alon Levy, and Dan Suciu. A query language for XML. Computer Networks (Amsterdam, Netherlands: 1999), 31(11–16):1155– 1169, May 17, 1999. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.elsevier.
com/cas/tree/store/comnet/
sub/1999/31/11-16/2146.pdf.

December:1995:HCU

[DG95] John December and Mark Ginsburg. HTML and CGI unleashed. SAMS Publishing, Indianapolis, IN, USA, 1995. ISBN 0-672-30745-6. xxvii + 830 pp. LCCN QA76.76.H94 D43 1995 Reserves. Contents: The World Wide Web as a communication system — A developer’s tour of the Web — Options for Web connections — Web development principles and methodology overview — Web planning — Web analysis – Web design — Web implementation — Web promotion – Web innovation — Design and implementation style and techniques — Basic HTML — Advanced HTML — Forms – Multimedia — Basics of imagemaps — Implementation tools — Development and language environments. Contents: Principles of gateway programming — Gateway programming fundamentals — Gateway programming I: programming libraries and databases — Gateway programming II: text search and retrieval tools – Scripting for the unknown: the control of chaos – Transaction security and security administration – Gateway programming language options and a server modification case study — VRML on the Web — C-based gateway scripting — Writing CGI scripts in REXX — A Web coloring book — A campus-wide information system — A hypertext news interface — A graphical Web page counter. Contents: Appendixes: A. Sources of further information — B. HTML tag and supporting information summary — C. Glossary. PC and Macintosh compatible CD-ROM includes: HTML examples in the book, CGI scripts that you can use for forms, guestbooks, searching, and more, Web development tools such as HTML editors, graphics, and Internet search and retrieval applications.

December:1996:HCU

[DG96] John December and Mark Ginsburg. HTML 3.2 and CGI Unleashed. Sams.net Pub., Indianapolis, IN, USA, second edition, September 1996. ISBN 1-57521-177-7. xliv + 1321 pp. LCCN QA76.76.H94D42 1996. US$49.99.

Davani:1997:ICS

[DG97] Darush Davani and Clark C.

Guest, editors. 1997 International Conference on Simulation in Engineering Education (ICSEE ’97), January 12–15, 1997, Sheraton Crescent Hotel, Phoenix, Arizona: proceedings of the 1997 Western Multiconference, volume 29(2) of Simulation series. Society for Computer Simulation, San Diego, CA, USA, 1997. ISBN 1-56555-104-4. ISSN 0735-9276. LCCN T65.5.C65 I583 1997.

Deaton:1993:ACS

[DGBH93] E. Deaton, K. M. George, H. Bergel, and G. Hedrick, editors. Applied Computing: States of the Art and Practice — 1993 Proceedings of the 1993 ACM/ SIGAPP Symposium on Applied Computing — February 1993, Indianapolis, IN, USA, Applied Computing — Symposium 1993. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-567-4. LCCN QA76.76.A65 S95 1993.

Delia:1992:SCD

[DGG^{+}92] D. J. Delia, T. C. Gilgert, N. H. Graham, U. Hwang, P. W. Ing, J. C. Kan, R. G. Kemink, G. C. Maling, R. F. Martin, K. P. Moran, J. R. Reyes, R. R. Schmidt, and R. A. Steinbrecher. System cooling design for the water-cooled IBM Enterprise System/9000 processors. IBM Journal of Research and Development, 36(4): 791–803, July 1992. CODEN IB-MJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

dHarcourt:1995:IDC

[d’H95a] J.-C. d’Harcourt. Integrating documentation into the company information system with SGML. Managing Information, 2(3):25–27, March 1995. CO-DEN MAIOEG. ISSN 1352-0229.

DHarcourt:1995:IDS

[D’H95b] J.-C. D’Harcourt. Integrating documentation with SGML. Managing Information, 2(3):25– ??, 1995. CODEN MAIOEG. ISSN 1352-0229.

Douglis:1997:HHM

[DHR97] F. Douglis, A. Haro, and M. Rabinovich. HPP: HTML macropreprocessing to support dynamic document caching. In USENIX [USE97b], pages 83– 94. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

Dick:1999:XMG

[Dic99] Kevin Dick. XML: a manager’s guide. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-43335-4. ???? pp. LCCN QA76.76.H94D534 1999.

Diehl:1996:JCS

[Die96] Stephan Diehl. Java & Co:

Die Sprachen des Webs: HTML, VRML, Java, JavaScript. (German) [Java & Co: The languages of the Webs: HTML, VRML, Java, JavaScript]. Addison-Wesley, Reading, MA, USA, 1996. ISBN 3-8273-1124-1. 288 pp. LCCN ????

49,90 DM. URL http://www.
addison-wesley.de/katalog/
item.ppml?textexpr=Java&id=
00099.

Diehl:1998:VLO

[Die98]  S. Diehl.  VRML++: A

language for object-oriented virtual-reality models. In Chen et al. [C^{+}98], pages 141–150.

ISBN 0-8186-8551-4, 0-8186-8553-0. LCCN QA76.64.T66 1997a. IEEE Computer Society Order Number PR08551. IEEE Order Plan Catalog Number 97TB100240.

Dille:1996:EMW

[Dil96] Ed Dille. Exploring moving worlds. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-467-7. ???? pp. LCCN QA76.76.H94 D55 1996.

Dille:1997:EMW

[Dil97] Ed Dille. Explorer Moving Worlds: VRML 2.0 et les mondes animés sur le Web. International Thomson Publishing France, Paris, France, 1997. ISBN 2-84180-146-2. xii + 226 pp. LCCN ???? Translated by Frederic Lefevere-Laoide.

Dimick:1996:RPS

[Dim96] S. A. Dimick. Researching and planning for SGML: The communicator’s role. In Anonymous [Ano96b], pages 238–?? ISBN 0-914548-89-1, 0-914548-90-5. LCCN ????

DiNucci:1995:THS

[DiN95]  Darcy DiNucci.  Tackling

HTML: The secret rules of Web design. U & lc: The International Journal of Typographics, 22(2):44–??, ???? 1995. CO-DEN ULCCDC. ISSN 0362-6245.

Divjak:1998:DHB

[Div98] S. Divjak. Design of hypertext based courseware with the integrated JAVA and VRML modules. In IEEE [IEE98e], pages 178–181. ISBN 0-7803-3879-0 (softbound), 0-7803-3880-4 (casebound), 0-7803-3881-2 (microfiche). LCCN TK5 .M43 1998; TK5101.A1 M38 1998. Two volumes. IEEE catalog number 98CH36056.

DAndrea:1996:UNG

[DJ96] Albert D’Andrea and Phil Janus. UniSQL’s next-generation object-relational database management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(3):70–76, September 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Dokko:1997:MES

[DK97] S.-J. Dokko and T.-Y. Kim.

Multimedia E-mail system giving an integrated message view based on the redefined HTML syntax. In Lavery [Lav97], pages 735–740. ISBN 0-7326-1257-8 (vol. 1), 0-7326-1258-6 (vol. 2), 0-7326-1259-4 (vol. 3), 0-7326-1256-X (set). LCCN ???? Three volumes.

Davis:1995:DPWb

[DL95a] J. Davis and C. Lagoze. “dropin” publishing with the World Wide Web. Computer Networks and ISDN Systems, 28(1– 2):247–??, December 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Davis:1995:DPWa

[DL95b] Jim Davis and Carl Lagoze. Drop-in publishing with the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ???? URL

                                        http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Pub/
                                        davis/davis-lagoze.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Dybvik:1995:CWM

[DL95c] Per Einar Dybvik and Hakon W. Lie. Combining WWW/Mosaic with realtime multimedia conferencing in distance education. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Educ/lie/lie.html. A selected

few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

LorenzoNieto:1997:H

[dL97] Angel de Lorenzo Nieto. HTML. Al Dia en una Hora; 39. Anaya Multimedia, Madrid, Spain, 1997. ISBN 84-415-0067-3. 128 pp. LCCN ????

Dauner:1998:PPO

[DLSR98] J. Dauner, J. Landauer, E. Stimpfig, and D. Reuter. 3D product presentation online: The virtual design exhibition. In Spencer and Roehl [SR98], pages 57–62. ISBN 1-58113-022-8. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/271897/.

Duval:1994:EIM

[DM94] B. K. Duval and L. Main. Exploring the Internet with Mosaic. Library Software Review, 13(4):269–279, Winter 1994. CODEN LSREEA. ISSN 0742-5759.

Duval:1995:MAL

[DM95] B. K. Duval and L. Main. Microcomputer applications in the library: building home pages. Library Software Review, 14(4): 218–227, Winter 1995. CODEN LSREEA. ISSN 0742-5759.

Duval:1996:AIW

[DM96] B. K. Duval and L. Main.

Adding images to Web pages using Adobe Photoshop 3.0 and Mapedit. Library Software Review, 15(2):88–102, Summer 1996. CODEN LSREEA. ISSN 0742-5759.

Deri:1997:VAN

[DM97] L. Deri and D. Manikis. VRML:

Adding 3D to network management. In Mullery et al. [M^{+}97], pages 469–477. CO-DEN LNCSD9. ISBN 3-540-63135-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.A1 I47985 1997.

Diefenbach:1998:BO

[DMH98] P. J. Diefenbach, P. Mahesh, and D. Hunt. Building Open-Worlds. In Spencer and Roehl [SR98], pages 33–38. ISBN 1-58113-022-8. LCCN ????

URL http://www.acm.org/
pubs/contents/proceedings/
graph/271897/.

Dengel:1990:IDM

[DMM90] A. Dengel, N. M. Mattos, and B. Mitschang. An integrated document management system. Proceedings of the SPIE — The International Society for Optical Engineering, 1293(pt.1): 368–379, ???? 1990. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

deFigueiredo:1997:XBS

[dN97] Luiz Henrique de Figueiredo and Marcio Lobo Netto, editors. X Brazilian Symposium on Computer Graphics and Image Processing: proceedings, Campos do Jordao, Brazil, October 14–17, 1997. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-8102-0, 0-8186-8103-9, 0-8186-8104-7. LCCN T385 .S59 1997. IEEE Computer Society order number PR08102. IEEE order plan catalog number 97TB100164.

Dassler:1996:KBV

[DO96] R. Dassler and A. Otto. Knowledge Browser — a VRML-Retrieval Interface for GEO-LIS. Nachrichten für Dokumentation, 47(3):151–158, May– June 1996. CODEN NADOAW. ISSN 0027-7436.

Dobrowolski:1991:TSD

[Dob91] Andrew E. Dobrowolski. Typesetting SGML documents using TEX. TUGboat, 12(34):409–414, November 1991. ISSN 0896-3207.

Dobson:1997:DHEa

[Dob97a] Rick Dobson. Dynamic HTML explained, part 1 — dynamic objects and client-side processing mean better Web sites and better browsing. BYTE Magazine, 22(11):53–54, November 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dobson:1997:DHEb

[Dob97b] Rick Dobson. Dynamic HTML explained, part II. BYTE Magazine, 22(12):61–62, December 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Dobreva:1998:TVW

[Dob98] M. Dobreva, editor. Text variety in the witnesses of medieval texts: proceedings, International Workshop, 21–23 September 1997, Sofia, Bulgaria. Institute of Mathematics and Informatics, Bulgarian Academy of Sciences; 1998, Sofia, Bulgaria, 1998. ISBN 954-9650-02-2. LCCN ????

Dodge:1994:USS

[Dod94] Daniel R. Dodge. Using SGML to streamline print and CD-ROM production. CD-ROM professional, 7(2):77–78, 80– 82, March 1, 1994. CODEN CRPFEX. ISSN 1049-0833.

Domel:1995:WGHb

[Döm95a] P. Dömel. WebMap: A graphical hypertext navigation tool. Computer Networks and ISDN Systems, 28(1–2):85–??, December 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Domel:1995:WGHa

[Döm95b] Peter Dömel. WebMap — A graphical hypertext navigation tool. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Searching/doemel/www-fall94.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Donovan:1997:ISI

[Don97]  Truly Donovan.  Industrial-

strength SGML: an introduction to enterprise publishing. The Charles F. Goldfarb series on open information management. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-216243-1. xx + 201 pp. LCCN QA76.76.H94D66 1997.

Dorward:1995:SPU

[Dor95a]  A. Dorward.  SGML in

publishing-why use the standard? Electronic Library, 13(1): 53–56, February 1995. CODEN ELLIDZ. ISSN 0264-0473.

Dorward:1995:SPW

[Dor95b] Andrew Dorward. SGML in publishing — why use the standard? Electronic Library, 13(1): 53–??, February 1, 1995. CO-DEN ELLIDZ. ISSN 0264-0473.

Doss:1999:PX

[Dos99a] George Doss. Pure XML. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-672-31601-3. 500 pp. LCCN ????

Doss:1999:CDG

[Dos99b] George M. Doss. CORBA developer’s guide with XML. Wordware Publishing, Plano, TX, USA, 1999. ISBN 1-55622-668-3. ???? pp. LCCN QA76.76.H94 D68 1999.

Doty:1995:TDS

[Dot95] T. Doty. Test driving SATAN.

Computer Security Journal, 11 (2):9–14, Fall 1995. CODEN CSJLDR. ISSN 0277-0865.

Dougherty:1992:ODP

[Dou92a] D. Dougherty. On-line documentation and publication: the Davenport Group. EurOpen Newsletter, 12(2):15–18, Summer 1992. CODEN EONLE8. ISSN 1011-4211.

Dougherty:1992:SPI

[Dou92b] D. Dougherty. Standardization: problems of interchange and delivery of documentation online. In ACM [ACM92], pages 145–147. ISBN 0-89791-532-1. LCCN QA76.9.D6 S741 1992. URL http://

www.acm.org/pubs/contents/
proceedings/doc/147001/.

DeBra:1995:SAI

[DP95a] Paul De Bra and R. D. J. Post. Searching for arbitrary information in the World Wide Web: the fish-search for Mosaic. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Searching/debra/article.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Duffy:1995:THR

[DP95b] Robert A. Duffy and Michael J. Palmer. Taking the high road to institutional self-promotion: Substantive corporate communication on the Worldwide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Overviews/duffy/main.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Davis:1998:PPM

[DP98a] M. L. Davis and C. S. Patridge. Publish paper manuals no morel: An introduction to creating HTML documentation. In Anonymous [Ano98t], pages 1508–1516. ISBN 1-58025-149-8. LCCN ????

Devanbu:1998:FIC

[DP98b] Premkumar Devanbu and Jeffrey Poulin, editors. Fifth International Conference on Software Reuse: proceedings: June 2–5, 1998, Victoria, British Columbia, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8377-5, 0-8186-8378-3 (case), 0-8186-8379-1 (microfiche). ISSN 1085-9098. LCCN QA76.76.R47 I587 1998. IEEE Order Plan Catalog Number 98TB100203. IEEE Computer Society Order Number PR08377.

Djurcilov:1998:VPO

[DP98c] S. Djurcilov and A. Pang. Visualization products on-demand through the Web. In Spencer and Roehl [SR98], pages 7– 14. ISBN 1-58113-022-8. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/271897/.

Debras:1991:ESD

[DPB91] P. Debras, P. Poyet, and E. Brisson. Expert systems and documentary databases integration (for building regulations). Microcomputers in Civil Engineering, 6(4):281–289, ???? 1991. CODEN MCENE7. ISSN 0885-9507.

December:1994:WU

[DR94] John December and Neil Randall. World-Wide Web Unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, November 1994. ISBN 0-672-30617-4. 1058 pp. LCCN TK 5105.875 I57 D43 1994. US$35.00, CDN$47.95. Computer Literacy Bookshops says: Not only the largest book on the Web, but also the most complete. 300 pages cover fundamental concepts and individual Wab browsers for different platforms. Covers navigation and searching techniques and list web sites by subject. Written by a team of WWW experts, the best part is the inclusion of details on technical information. Loads of information on designing and creating a Web site, working with HTML editors and filters, creating forms and image mapping.

December:1995:WWUa

[DR95a] John December and Neil Randall. World-Wide Web Unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 1995. ISBN 0-672-30737-5. xxxvii + 1346 pp. LCCN TK5105.875.I57 D32 1995. US$35.00, CDN$47.95. Includes CD-ROM.

December:1995:WWUb

[DR95b] John December and Neil Randall. World-Wide Web Unleashed 1996. Howard W. Sams, Indianapolis, IN 46268, USA, third edition, November 1995. ISBN 1-57521-040-1. xxxix + 1392 pp. LCCN TK5105.888 .D43 1995a. US$49.99. Includes CD-ROM.

Donath:1995:SW

[DR95c] Judith Donath and Niel Robertson. The sociable Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
CSCW/donath/SociableWeb.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Donszelmann:1995:GWW

[DR95d] M. Donszelmann and K. Rodden. Gateways for World-Wide Web in the ‘online’ data acquisition system of the DELPHI experiment at CERN. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ???? URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
donszelmann/paper.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Drakos:1994:THP

[Dra94] Nikos Drakos. From text to hypertext: A post-hoc rationalisation of L^{A}TEX2HTML. Computer Networks and ISDN Systems, 27(2):215–224, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1994&
                                        volume=27&issue=2&aid=1394.

Dratva:1995:WBHb

[Dra95a] R. Dratva. WWW-based home banking services in Switzerland: A case study. Computer Networks and ISDN Systems, 28(1– 2):199–??, December 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Dratva:1995:WBHa

[Dra95b]  Richard Dratva.  WWW-

based home banking services in Switzerland: a case study. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        ComEc/dratva/dratva.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Drandaki:1996:EHA

[Dra96] A. Drandaki. EULOGIA: a hypermedia application for museum cataloguing enriched with SGML encoding. In Moscati [Mos96], pages 1039–1046. ISBN 88-7814-109-7. ISSN 1120-6861. LCCN ???? Two volumes.

Dolin:1998:SXI

[DRBS98] R. H. Dolin, W. Rishel, P. V.

Biron, and J. Spinosa. SGML and XML as interchange formats for HL7 messages. In Chute [Chu98], pages 720–724. ISBN ???? ISSN 1091-8280. LCCN ????

Dreyfus:1997:LDH

[Dre97] Michel Dreyfus. Livre d’or HTML 4. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1997. ISBN 2-7361-2750-1. xxv + 773 pp. LCCN ????

Dreyfus:1998:DRH

[Dre98] Michel Dreyfus. Le dico reference HTML 4. Simon & Schuster Macmillan (France), Paris, France, 1998. ISBN 2-7440-0548-7. v + 382 pp. LCCN ????

Du:1996:PPW

[DS96] David Du and Olivia R. Liu Sheng, editors. Proceedings of the 1996 Pacific Workshop on Distributed Multimedia Systems, June 27–28, 1996, Hong Kong. Knowledge Syst. Inst. and Dept. of Information and Systems Management, Hong Kong University — of Science and Technology, Skokie, IL, USA and Hong Kong, 1996. ISBN 0-9641699-4-0. LCCN ????

Dao:1996:IST

[DSDT96a] Tuong Dao, R. Sacks-Davis, and J. A. Thom. Indexing structured text for queries on containment relationships. In Anonymous [Ano96d], pages 82–91. CODEN ACSCDD. ISBN ???? ISSN 0157-3055. LCCN ????

TuongDao:1996:IST

[DSDT96b] Tuong Dao, R. Sacks-Davis, and J. A. Thom. Indexing structured text for queries on containment relationships. Aust. Comput. Sci. Commun. (Australia), Australian Computer Science Communications, 18(2):82–91, ???? 1996. CODEN ACSCDD. ISSN 0157-3055.

Daddio:1995:NWE

[DT95] Ernest Daddio and William Turnbull. NOAA on the Web — the experiences and future of one nation-wide user. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
NationWideSites/turnbull/

paper.html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Davis-Tanous:1999:XLM

[DT99] Jennifer R. Davis-Tanous. XML, a language to manage the World Wide Web. Syracuse, NY, USA, 1999. Shipping list no. 2001-0191-M.

DuCharme:1997:SC

[DuC97]  Robert DuCharme.  SGML

CD. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, July 1997. ISBN 0-13-475740-8. xix + 353 pp. LCCN QA76.76.H94D83 1998. US$49.95. URL http://www.

snee.com/bob/sgmlfree/. In-

cludes CD-ROM.

DuCharme:1998:SC

[DuC98] Bob DuCharme. SGML CD.

The Charles F. Goldfarb series on open information management. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-475740-8. xix + 353 pp. LCCN QA76.76.H94 D83 1998.

DuCharme:1999:XAS

[DuC99] Bob DuCharme. XML: the annotated specification. The Charles F. Goldfarb series on open information management. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-082676-6. xix + 339 pp. LCCN QA76.76.H94 D835 1999.

Dudeck:1997:NTH

[Dud97] J. (Joachim) Dudeck, editor.

New technologies in hospital information systems:Based on conference held in Ulm, Germany, 20–23 September, 1997, number 45 in Studies in health technology and informatics. IOS Press, Amsterdam, The Netherlands, 1997. ISBN 90-5199-363-3, 4-274-90185-8. ISSN 0926-9630. LCCN ????

Dugas:1997:IHA

[Dug97] M. Dugas. Intranet and HTML at a major university hospital — experiences from Munich. In Pappas et al. [PMS97], pages 119–122. ISBN 90-5199-343-9, 4-274-90169-6. ISSN 0926-9630. LCCN W 26.5 M489-7 1997. Two volumes.

Duke:1994:SRE

[Duk94] J. K. Duke. Slow revolution: the electronic AACR2. Library Resources and Technical Services, 38(2):190–194, April 1994. CO-DEN LRTSAH. ISSN 0024-2527.

Duncan:1995:PPa

[Dun95a] Ray Duncan. Power programming. PC Magazine, 14(11): 261–??, June 13, 1995. CODEN PCMGEP. ISSN 0888-8507.

Duncan:1995:PPb

[Dun95b] Ray Duncan. Power programming. PC Magazine, 14(18): 255–??, October 24, 1995. CO-DEN PCMGEP. ISSN 0888-8507.

Duncan:1995:PPc

[Dun95c] Ray Duncan. Power programming. PC Magazine, 14(21): 391–??, December 5, 1995. CO-DEN PCMGEP. ISSN 0888-8507.

Duncan:1995:PPHa

[Dun95d] Ray Duncan. Power programming — an HTML primer. PC Magazine, 14(11):261–??, ???? 1995. CODEN PCMGEP. ISSN 0888-8507.

Duncan:1995:PPHb

[Dun95e] Ray Duncan. Power programming — HTML 3.0 now lets you present information in table format. find out the best way to take advantage of this new capability. PC Magazine, 14(18): 255–??, ???? 1995. CODEN PCMGEP. ISSN 0888-8507.

Duncan:1995:PPP

[Dun95f]  Ray Duncan.  Power pro-

gramming — publishing HTML forms on the Web: How to retrieve data from your Web page using HTML 2.0’s forms creation tags. PC Magazine, 14 (21):391–??, ???? 1995. CO-DEN PCMGEP. ISSN 0888-8507.

Dunlop:1995:PCU

[Dun95g] D. Dunlop. Practical considerations in the use of TEI headers in a large corpus. Computers and the humanities, 29(1): 85–98, ???? 1995. CODEN CO-HUAD. ISSN 0010-4817.

Doszkocs:1995:ACN

[DWV95] Tamas E. Doszkocs, Seth B.

Widoff, and Bruno M. Vasta. Associative concept navigation in MEDLINE and other NLM databases via a Mosaic — forms — WWW interface combining natural language processing, expert systems and (un)conventional information retrieval techniques.

In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Searching/doszkocs/doszkocs.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Eichhorn:1995:AAS

[EAG^{+}95] G. Eichhorn, A. Accomazzi, C. S. Grant, M. J. Kurtz, and S. S. Murray. Access to the Astrophysics Science Information and Abstract System. In Anonymous [Ano95x], pages 217–225. CODEN VASTA6. ISBN ???? ISSN 0083-6656. LCCN ????

Earnshaw:1987:WPS

[Ear87] Rae A. Earnshaw, editor. Workstations and Publication Systems. Springer-Verlag Inc., New York, NY, USA, 1987. ISBN 0-387-96527-0, 3-540-96527-0. viii + 229 pp. LCCN Z286.E43 W67 1987. US$29.50.

England:1996:RRT

[EAU96] Paul England, Robert Allen, and Ron Underwood. RAVE: Real-time services for the Web. Computer Networks and ISDN Systems, 28(7–11):1547– 1558, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&

volume=28&issue=7-11&aid=
1620.

Edwards:1998:CAT

[EAZ98] Alistair D. N. Edwards, Andras Arato, and Wolfgang L. Zagler, editors. Computers and assistive technology ICCHP ’98: Proceedings of the XV. IFIP World Computer Congress, 31 August– 4 September 1998, Vienna, Austria and Budapest, Hungary, volume 118 of Schriftenreihe der Osterreichischen Computer Gesellschaft. Osterreichische Computer Gesellschaft, Vienna, Austria, 1998. ISBN 3-85403-118-1. LCCN ????

Ebenhoch:1998:LCR

[Ebe98]  P. Ebenhoch.  Legal cita-

tion referencing using SGML and HyTime. In Tjoa and Wagner [TW98], pages 625– 632. ISBN 0-8186-8353-8, 0-8186-8355-4. LCCN QA76.9.D3 D15592 1998. IEEE Computer Society Order number PR08353.

Encarnacao:1996:NUI

[EBR^{+}96] José Encarnação, Klaus M.

Bauer, H. Roland, K. Tanja, and S. Ulrike. New user interface aspects to the information highway — applying Java and VRML. International Journal of Information Technology, 2(1): 79–95, June 1996. ISSN 0218-7957.

Eckstein:1999:XPR

[Eck99] Robert Eckstein. XML pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-709-5. iii + 107 pp. LCCN QA76.76.H94 E25 1999.

Essin:1993:IPL

[ECvR93] D. J. Essin, M. F. Collen, A. M. van Ginneken, and A. L. Rector. Intelligent processing of loosely structured documents as a strategy for organizing electronic health care records. Methods of Information in Medicine, 32 (4):265–273, 341, August 1993. CODEN MIMCAI. ISSN 0026-1270.

Eddy:1997:HPEa

[Edd97a] Sandra E. Eddy. HTML in Plain English. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1997. ISBN 1-55828-511-3. ???? pp. LCCN QA76.76.H94E33 1997. US$16.95.

Eddy:1997:HPEb

[Edd97b] Sandra E. Eddy. HTML in Plain English. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, second edition, 1997. ISBN 1-55828-587-3. vi + 728 pp. LCCN QA76.76.H94E32 1998. US$16.95.

Eddy:1998:HPE

[Edd98a] Sandra E. Eddy. HTML in plain English. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, second edition, 1998. ISBN 1-55828-587-3. vi + 728 pp. LCCN QA76.76.H94E32 1998.

Eddy:1998:XPE

[Edd98b] Sandra E. Eddy. XML in plain English. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1998. ISBN 0-7645-7006-4. xxxi + 559 pp. LCCN QA76.76.H94 E35 1998.

Eddy:1999:XPE

[Edd99] Sandra E. Eddy. XML in plain English. Transworld, IDG, 1999. ISBN 0-7645-7006-4. xxxi + 559 pp. LCCN QA76.76.H94E35 1998.

Elcacho:1998:VAC

[EDLS98] C. Elcacho, R. Doerner, V. Luckas, and A. Schaefer. VRML authoring in the context of industrial applications. In Jacucci [Jac98], pages 523–540. ISBN 0-412-83540-1. LCCN TS155.6 .I569 1998. Includes CD-ROM.

Efthimiadis:1990:OSA

[Eft90] E. Efthimiadis. Online searching aids: a review of front ends, gateways and other interfaces. Journal of Documentation, 46(3):218–262, September 1990. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Elledge:1990:SSG

[EG90] Marion Elledge and Graphic Communications Association (U. S.). The SGML source guide: the Graphic Communications Association’s guide to standard generalized markup language (SGML) systems, software, service, consultants, seminars, and resources. Graphic Communications Association, 100 Daingerfield Road, Alexandria, VA 22314-2804, USA, 1990. ISBN 0-933505-13-2. LCCN ????

Elledge:1991:SSG

[EG91] Marion Elledge and Graphic Communications Association (U. S.). The SGML source guide: the Graphic Communications Association’s guide to standard generalized markup language (SGML) systems, software, service, consultants, seminars, and resources. Graphic Communications Association, 100 Daingerfield Road, Alexandria, VA 22314-2804, USA, revised June 1991 edition, 1991. ISBN 0-933505-13-2. 212 pp. LCCN ???? US$72.00.

Englert:1995:PPC

[EGH95] Susanne Englert, Ray Glasstone, and Waqar Hasan. Parallelism and its price: A case study of NonStop SQL/MP. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(4):61–71, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Egret:1995:WARb

[EH95a] D. Egret and A. Heck. WWW in astronomy and related space sciences. Computer Networks and ISDN Systems, 28(1–2): 161–??, December 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Egret:1995:WARa

[EH95b] Daniel Egret and Andre’ Heck.

The Web in astronomy and related space sciences. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Astronomy/egret/egret.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Espinoza:1996:WIA

[EH96] F. Espinoza and K. Hook. A WWW interface to an adaptive hypermedia system. In Anonymous [Ano96-31], pages 143– 167. ISBN ???? LCCN ????

Edwards:1997:MMP

[EH97] J. D. M. Edwards and C. Hand.

MAPS: Movement and planning support for navigation in an immersive VRML browser. In Spencer and Brutzman [SB97], pages 65–74. ISBN 0-89791-886-X. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/253437/.

ACM order number 434973.

Eichmann:1995:EWAb

[Eic95a] D. Eichmann. Ethical Web agents. Computer Networks and ISDN Systems, 28(1–2): 127–??, December 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Eichmann:1995:EWAa

[Eic95b]  Dave Eichmann.  Ethical

Web agents. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Agents/eichmann.ethical/eichmann.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Eijkhout:1992:TTT

[Eij92] Victor Eijkhout. TEX by Topic, a TEXnician’s Reference. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56882-9. vii + 307 pp. LCCN Z253.4.T47 E38

                                        1992. URL http://eijkhout.
                                        net/texbytopic/. See [Pop92,

PH92].

Eisenberg:1996:NSS

[Eis96] Andrew Eisenberg. New standard for stored procedures in SQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(4):81– 88, December 1996. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Elvins:1995:WBV

[EJ95] T. Todd Elvins and R. Jain. Web-based volumetric data retrieval. In Nadeau and Moreland [NM95b], pages 7–12. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Epstein:1995:WEH

[EKS95] Jonathan A. Epstein, Jonathan A.

Kans, and Gregory D. Schuler. WWW Entrez: A hypertext retrieval tool for molecular biology. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
BioChem/epstein/WWW_entrez.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Ellison:1994:ISC

[Ell94] P. Ellison. Introduction to SGML concepts. In Anonymous [Ano94g], pages 1/1–1/5. ISBN ???? LCCN ????

Ellis:1998:TMR

[Ell98] S. Ellis. Towards more realistic sound in VRML. In Spencer and Roehl [SR98], pages 95– 100. ISBN 1-58113-022-8. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/271897/.

Engel:1995:GHL

[ELW95] A. Engel, D. Langkafel, and W. Weber. Generating hypertext links through queries. In Kuhlen and Rittberger [KR95], pages 219–231. ISBN 3-87940-509-3. LCCN ????

Eisenberg:1999:SPS

[EM99] Andrew Eisenberg and Jim Melton. SQLJ part 1: SQL routines using the Java programming language. SIGMOD Record (ACM Special Interest Group on Management of Data), 28(4):58–63, December 1999. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Espiritu:1995:LUM

[EMH^{+}95] Baltazar Espiritu, John A. Mc-Nulty, James R. Halama, Arcot J. Chandrasekhar, Ronald Price, Jr., and Brain Maggi. The Loyola University Medical Education Network. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        MedEd/espiritu/LUMEN.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Engelen:1993:OTP

[Eng93] J. J. Engelen. Overview of the TIDE-CAPS project. In Anonymous [Ano93g], pages 4/1–8. ISBN ???? LCCN ????

Englert:1994:NSS

[Eng94] Susanne Englert. NonStop SQL: Scalability and availability for decision support. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):491, June 1994. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Engberg:1997:MJV

[Eng97] David A. Engberg. Multi-user Java/VRML performance: research project. Master of science, plan ii, Department of Electrical Engineering and Computer Sciences, University of California, Berkeley, Berkeley, CA, USA, 1997. 10 pp.

Ensign:1993:SE

[Ens93] Chet Ensign. SGML by evolution. Technical Communication, 40(3):387–393, August 1993. CODEN TLCMBT. ISSN 0049-3155.

Ensign:1997:GBD

[Ens97] Chet Ensign. $GML — the billion dollar secret: how leading companies are turning documents into assets. The Charles F. Goldfarb series on open information management. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-226705-5. xxvi + 213 pp. LCCN HF5738.E57 1997.

Erlenkotter:1997:JHS

[ER97] Helmut Erlenkötter and Volker Reher. Java. HTML, Skripte, Applets und Anwendungen. Rowohlt Taschenbuch, Reinbek bei Hamburg, Germany, 1997. ISBN 3-499-19872-X. 320 pp. LCCN ???? 20 DM. URL http:

//www.abc.de/rowohlt.html.

Erfle:1993:STC

[Erf93] R. Erfle. Specification of temporal constraints in multimedia documents using HyTime. Electronic Publishing—Origination, Dissemination, and Design, 6(4):397–411, December 1993. CODEN EPODEU. ISSN 0894-3982.

Ericksen:1997:PH

[Eri97] Linda Ericksen. Projects for HTML. SELECT lab series. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-30421-X. xi + 210 pp. LCCN ????

El-Rewini:1997:SSVa

[ERPNS97a] H. El-Rewini, Y. N. Patt, J. F. Nunamaker, and R. H. Sprague, editors. System sciences Vol 3, Information systems track — organizational systems and technology: Hawaii international conference; 30th — January 1997, Wailea; HI, volume 30 of Proceedings of the Hawaii International Conference on System Sciences 1997; //v3. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7734-1, 0-8186-7749-X. ISSN 1060-3425. LCCN ????

El-Rewini:1997:SSVb

[ERPNS97b] H. El-Rewini, Y. N. Patt, J. F. Nunamaker, and R. H. Sprague, editors. System sciences Vol 6, Digital documents track: Hawaii international conference; 30th — January 1997, Wailea; HI, volume 30 of Proceedings of the Hawaii International Conference on System Sciences 1997; //v6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7734-1, 0-8186-7862-3. ISSN 1060-3425. LCCN ????

Eddy:1999:TYX

[ES99a] Sandra E. Eddy and John E.

Schnyder. Teach yourself XML. IDG Books, San Mateo, CA, USA, 1999. ISBN 0-7645-7513-9. xxii + 522 pp. LCCN QA76.76.H94 E34 1999.

Erickson:1999:SHE

[ES99b] J. Erickson and M. Stoeffler. An SGML/HTML electronic thesis and dissertation library. In Mylonas and Renear [MR99b], pages 175–184. ISBN ???? ISSN 0010-4817. LCCN ????

Ehrenfeucht:1994:CFT

[EtR94] A. Ehrenfeucht, P. ten Pas, and G. Rozenberg. Context-free text grammars. Acta Informatica, 31 (2):161–206, ???? 1994. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). From Joachim Schrod: This paper describes “the basic theory behind text grammars needed for document structure specification”.

Eren:1997:OVM

[ETT97] P. E. Eren, C. Toklu, and A. M. Tekalp. Object-based video manipulation and composition using 2D meshes in VRML. In Wang et al. [W^{+}97], pages 257–261. ISBN 0-7803-3780-8, 0-7803-3781-6, 0-7803-3782-4. LCCN TK5102.9.I338 1997. IEEE catalog number 97TH8256.

EMWAC:1995:WWW

[EU95] European Microsoft Windows NT Academic Centre and University of Edinburgh. Computing Services. World Wide Web sampler, 1995. Includes CD-ROM.

Erfle:1995:USS

[Ev95] R. Erfle and G. von Zadow. The use of SGML for standardization of documentation. Informationstechnik und technische Informatik: IT + TI, 37(6):36– 44, December 1995. CODEN ITINEY. ISSN 0944-2774.

Evans:1995:MGH

[Eva95] Tim Evans. 10 minute guide to HTML. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0541-9. x + 196 pp. LCCN QA76.76.H94 E93 1995.

Evans:1996:TMG

[Eva96] Tim Evans. Ten Minute Guide to HTML. Que Corporation, Indianapolis, IN, USA, second edition, November 1996. ISBN 0-7897-0965-1. x + 228 pp. LCCN QA76.76.H94E93 1996. US$14.99.

Evans:1997:PPHa

[Eva97a] Tim Evans. Prentice pocket HTML. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 84-89660-72-7. xvii + 223 pp. LCCN ????

Evans:1997:PPHb

[Eva97b] Tim Evans. Prentice pocket HTML 3.2. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1997. ISBN 84-89660-43-3. xx + 264 pp. LCCN ???? Translated by Vuelapluma.

Evans:1998:MGH

[Eva98a] Tim Evans. 10 minute guide to HTML 4.0. Que Corporation, Indianapolis, IN, USA, third edition, 1998. ISBN 0-7897-1491-4. xvi + 230 pp. LCCN QA 76.76 H94 E93 1998.

Evans:1998:STY

[Eva98b] Tim Evans. Sams’ teach yourself HTML 4.0 in 10 minutes. Sams Publishing, Indianapolis, IN, USA, 1998. ISBN 0-672-31325-1. xvi + 230 pp. LCCN QA76.76.H94E96 1998.

Eyzaguirre:1995:WDC

[Eyz95] Alan Eyzaguirre. Webcraft: Dynamic Content Using HTML. Ziff-Davis Press, Emeryville, CA, USA, June 1995. ISBN 1-56276-322-9. ???? pp. LCCN ???? US$29.95.

Fox:1994:DVD

[FA94] E. A. Fox and G. Abdulla. Digital video delivery for a digital library in computer science. Proceedings of the SPIE — The International Society for Optical Engineering, 2188:60–66, ???? 1994. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Fabozzi:1995:UMD

[Fab95] Donald J. Fabozzi II. The use of Mosaic as a documentation tool for large, graphically-based simulation software. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
fabozzi/fabozzi.paper.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Falla:1997:HSS

[Fal97] Rob Falla. HTML Style Sheets Quick Reference. Que Corporation, Indianapolis, IN, USA, January 1997. ISBN 0-7897-1035-8. xvii + 203 pp. LCCN QA76.76.H94F35 1997. US$19.99.

Farrar:1996:HPE

[Far96a] Edward Farrar. HTML Programming Example Book. Programmer’s Example Ser. Wordware Publishing, Plano, TX, USA, September 1996. ISBN 1-55622-517-2. 400 pp. LCCN ???? US$34.95.

Farrell:1996:VTA

[Far96b] J. Farrell. VRML toolkit for aerospace simulation. In Gauthier [Gau96a], pages 92–96. ISBN ???? LCCN ????

Farell:19xx:EH

[Farxx] Cecilia Farell. Essential HTML 3.0. SAMS Publishing, Indianapolis, IN, USA, 19xx. ISBN 1-57521-044-4. ???? pp. LCCN ???? US$25.00.

Fateman:1997:NSS

[Fat97] Richard J. Fateman. Network servers for symbolic mathematics. In Küchlin [Küc97], pages 249–256. ISBN ???? LCCN ????

Fawcett:1989:ASI

[Faw89] Heather Fawcett. Adopting SGML: the implications for writers. Technical report, University of Waterloo Centre for the New Oxford Dictionary, Waterloo, ON, Canada, 1989. 10 pp.

Fawcett:1993:NOE

[Faw93] Heather Fawcett. The New Oxford English Dictionary Project. Technical Communication, 40 (3):379–382, August 1993. CO-DEN TLCMBT. ISSN 0049-3155.

Fox:1995:IET

[FB95a] E. Fox and N. Dwight Barnette. Improving education through a CS digital library with three types of WWW servers. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/
barnette/barnette.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Francois:1995:SHR

[FB95b] P. François and P. Bazex. SGML/HyTime repositories: Requirements and data modelling using object-oriented database concepts. Lecture Notes in Computer Science, ?? (978):259–271, ???? 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frisby:1995:MEW

[FBDM95] Anthony J. Frisby, James Baggott, Sharon Dennis, and Rod Murray. Medical education and the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
MedEd/frisby/frisby.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Feeney:1991:MIP

[FD91] Mary Feeney and Shirley Day, editors. Multimedia Information. Proceedings of a Conference. Bowker Saur, London, UK, 1991. ISBN 1-85739-090-3. LCCN QA76.575 .M85 1991.

Fox:1995:HWP

[FD95] David Fox and Troy Downing. HTML Web Publisher’s Construction Kit: Publishing Your Own HTML Pages on the Internet. Waite Group Press, Corte Madera, CA, USA, 1995. ISBN 1-57169-018-2. xxxv + 673 pp. LCCN QA76.76.H94 F69 1995. US$36.95, CDN$53.95. URL

http://w3.waite.com/waite/
waite/books.new//HTML_WebPub_
Construc_Kit/html/hwpckcov.

html. System requirements for accompanying computer disc: 386 DX or better; VGA card and display; 4 MB RAM; hard drive; CD-ROM drive; access to the Internet. Includes index.

Fox:1996:WPCb

[FD96] David Fox and Troy Downing.

Web Publisher’s Construction Kit with HTML 3.2: Publishing Your Own HTML Pages on the Internet. Web Publisher’s Construction Kits Ser. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN 1-57169-079-4. xxix + 898 pp. LCCN QA76.76.H94 F693 1996. US$44.99.

Feeney:1988:SGM

[Fee88a] Mary Feeney. The Standard Generalized Markup Language (SGML), volume 9 of Library & information briefing. British Library Research and Development Dept. and Library & Information Technology Centre, London, UK, 1988. ISSN 0954-1829. 15 pp. LCCN Z286.E43 F44 1988.

Feeney:SGM88

[Fee88b] Mary Feeney. The Standard Generalized Markup Language (SGML), volume 9 of Library & information briefing. British Library Research and Development Dept. and Library & Information Technology Centre, London, UK, 1988. ISSN 0954-1829. 15 pp. LCCN Z286.E43 F44 1988.

Fee:1998:DCP

[Fee98] S. B. Fee. Developing CD-ROM products with hypertext markup language (HTML). In Griffin et al. [GSG98], pages 107–112. ISBN 0-945829-11-6. LCCN ????

Ferrall:1993:ASC

[Fer93a] J. Ferrall. AIM and SGML: Combining data base and SGML for flexible, standards-based video and hard copy output. In Anonymous [Ano93j], pages ST– 144–?? ISBN ???? ISSN ???? LCCN ????

Ferrall:1993:ILD

[Fer93b] J. Ferrall. ISD/LSAR DSS and AIM: Linking a training materials front-end analysis tool with a development and maintenance tool via SGML. In Anonymous [Ano93j], pages LO–93–LO–97. ISBN ???? ISSN ???? LCCN ????

Fernandez:1994:RBW

[Fer94] Phillip M. Fernandez. Red brick warehouse: A read-mostly RDBMS for open SMP platforms. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):492, June 1994. CODEN SRECD8.

ISSN 0163-5808 (print), 1943-5835 (electronic).

FernandezRequejo:1997:APH

[Fer97] Antonio Fernandez Requejo, tr.

Aprende y practica HTML 3.2. Libro PC Magazine. Anaya Multimedia, Madrid, Spain, 1997. ISBN 84-415-0179-3. 289 pp. LCCN ????

Fox:1999:UWI

[FFKO99] G. C. Fox, W. Furmanski, G. Krishnamurthy, and H. T. Ozdemir. Using WebHLA to integrate HPC FMS modules with Web/commodity based distributed object technologies of CORBA, Java, COM and XML. In Tentner [Ten99], pages 273–278. ISBN 1-56555-166-4. LCCN ????

Fletcher:1995:ARI

[FG95] G. Fletcher and A. Greenhill. Academic referencing of Internet-based resources. ASLIB Proceedings, 47(11-12): 245–252, November–December 1995. CODEN ASLPAO. ISSN 0001-253X (print), 1758-3748 (electronic).

Frei:1996:CDI

[FGGT96] Z. Frei, P. Guhathakurta, J. E.

Gunn, and J. A. Tyson. A catalog of digital images of 113 nearby galaxies. Astronomical Journal, 111(1):174–181, January 1996. CODEN ANJOAA. ISSN 0004-6256.

Fielding:1997:RHT

[FGM^{+}97] R. Fielding, J. Gettys, J. Mogul, H. Frystyk, and T. Berners-Lee. RFC 2068: Hypertext Transfer Protocol — HTTP/1.1, January 1997. URL ftp://ftp.

                                        internic.net/rfc/rfc2068.
                                        txt;   http://www.math.utah.
                                        edu/pub/rfc/rfc2068.txt. Sta-

tus: PROPOSED STANDARD.

Fagan:1997:VRM

[FH97] B. Fagan and C. Haritatos. Virtual reality markup language (VRML) visualization tool (VVT) for target system analysis. In Anonymous [Ano97-27], pages 117– 125. ISBN ???? LCCN ????

Francis:1999:IDH

[FHU99] Brian Francis, Alex Homer, and Chris Ullman. IE5 dynamic HTML programmer’s reference. Programmer’s reference. Wrox Press, Chicago, IL, USA, 1999. ISBN 1-86100-174-6. xiii + 551 pp. LCCN ????

Fichter:1998:XNE

[Fic98a] D. Fichter. XML: A new “Extreme” sport or a solution to metadata nightmares? In Nixon et al. [NDM98], pages 124–125. ISBN 1-57387-067-6. LCCN Z674.82.I59 I585 1998.

Fickas:1998:IHB

[Fic98b] Stephen Fickas. Interactive HTML: a byte size introduction. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-919507-6. 50 pp. LCCN QA76.76.H94F54 1998.

Fielding:1994:MDH

[Fie94] Roy T. Fielding. Maintaining distributed hypertext infostructures: welcome to MOMspider’s Web. Computer Networks and ISDN Systems, 27 (2):193–204, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1994&
volume=27&issue=2&aid=1390.

Filippini:1994:MII

[Fil94] L. Filippini. METEOSAT IR image processing and MPEG animation. International Journal of Modern Physics C [Physics and Computers], 5(5): 831–833, October 1994. CO-DEN IJMPEO. ISSN 0129-1831 (print), 1793-6586 (electronic).

Finke:1995:S

[Fin95a] J. Finke. SQL. In USENIX Association [USE95], pages 133– 138. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

Finke:1995:SAG

[Fin95b] Jon Finke. SQL 2 HTML: Automatic generation of HTML database schemas. In USENIX Association [USE95], pages 133– 138. CODEN ???? ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

Fischer:1995:GWC

[Fis95] Elizabeth Fischer. Graffiti on the Web: A cultural interchange A lighthearted romp with an artist in Webland that stops being lighthearted at the end. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Arts/fischer/fischer.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Fisher:1996:SWG

[Fis96a] Yuval Fisher. Spinning the Web: A Guide to Serving Information on the World Wide Web. Springer-Verlag Inc., New York, NY, USA, 1996. ISBN 0-387-94539-3. xiv + 536 pp. LCCN TK5105.888.F58 1996. US$27.95. Discusses Java, JavaScript, and VRML.

Fiske:1996:VPC

[Fis96b] Teresa Fiske. VRML publishing with Caligari Fountain. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-450-2. ???? pp. LCCN QA76.76.H94 F57 1996.

Fisher:1997:CDW

[Fis97] Scott Fisher. Creating dynamic Web sites: a webmaster’s guide to interactive multimedia. Addison-Wesley Developers Press, Reading, MA, USA, 1997. ISBN 0-201-44207-8. ???? pp. LCCN QA76.76.I59F56 1997.

Finkel:1992:SAS

[FJ92] A. Finkel and M. Jantzen, editors. STACS 92. 9th Annual Symposium on Theoretical Aspects of Computer Science. Proceedings. Springer-Verlag Inc., New York, NY, USA, 1992. ISBN 3-540-55210-3. LCCN QA75.5.S958 1992.

Fawcett:1996:UCG

[FJ96] J. W. Fawcett and R. W. Jepson. Using the Common Gateway Interface. UNIX review, 14 (5):39–40, 42–44, 46, May 1996. CODEN UNRED5. ISSN 0742-3136.

Freeman:1996:MCN

[FJV96] Martin Freeman, Paul Jardetzky, and Harrick M. Vin, editors. Multimedia Computing and Networking 1996: 29–31 January 1996, San Jose, California, volume 2667 of Proceedings of the SPIE — The International Society for Optical Engineering. SPIE Optical Engineering Press, 1996. CODEN PSISDG. ISBN 0-8194-2041-7. ISSN 0277-786X (print), 1996-756X (electronic). LCCN QA76.575 .M784 1996.

Franzen:1996:LAS

[FK96] K. Franzen and G. Kaellgren.

Linguistic analysis and SGML markup. A case study from the area of legal information retrieval. In Anonymous [Ano96f], pages 275–?? ISBN 82-7283-083-3. ISSN 0804-8800. LCCN ????

Flanagan:1996:JND

[Fla96] David Flanagan. Java in a Nutshell: A Desktop Quick Reference for Java Programmers. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 1996. ISBN 1-56592-183-6. xix + 438 pp. LCCN QA76.73.J38F553 1996. US$14.95. URL http://www.

ora.com/info/java.

Frivold:1995:EWSb

[FLF95a] T. J. Frivold, R. E. Lang, and M. W. Fong. Extending WWW for synchronous collaboration. Computer Networks and ISDN Systems, 28(1–2):69–??, December 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Frivold:1995:EWSa

[FLF95b] Thane J. Frivold, Ruth E. Lang, and Martin W. Fong. Extending WWW for synchronous collaboration. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/
frivold/frivold.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Flohr:1996:PSC

[Flo96]  U. Flohr.  Put the space

in cyberspace [Virtual Reality Markup Language]. BYTE Magazine, 21(3):61–62, 64, March 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Filiatrault:1995:AWW

[FLSG95] David W. Filiatrault, Craig A.

Lightle, James E. Simmons, and Jerrold M. Grochow. AMS on the World-Wide-Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
CorInfSys/grochow/grochow.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Flynn:1993:TSR

[Fly93] Peter Flynn. TEX and SGML: a recipe for disaster? TUGboat, 14(3):227–230, October 1993. ISSN 0896-3207.

Flynn:1995:WHH

[Fly95]  Peter Flynn.  The World-

WideWeb Handbook: An HTML Guide for Users, Authors and Publishers. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1995. ISBN 1-85032-205-8. xix + 351 pp. LCCN TK5105.888 .F56 1995. Foreword by Tim Berners-Lee. Three sections deal with (1)

Getting connected to the Internet and using Internet software;

(2) Writing HTML (2.0) files for the WorldWideWeb; (3) Running a HTTP server and providing a Web service. Author is a member of the IETF Working Group on HTML. Text includes additional material on SGML; choice of editors, browsers and servers; copyright and intellectual property; and advance details of HTML3.

Flynn:1996:LEA

[Fly96] Peter Flynn. Letter to the editor: Automating HTML conversion. Dr. Dobb’s Journal of Software Tools, 21(10):8, October 1996. CODEN DDJOEB. ISSN 1044-789X.

Flynn:1998:USX

[Fly98] Peter Flynn. Understanding SGML and XML Tools: Practical Programs for Handling Structured Text. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 0-7923-8169-6. xxvi + 432 pp. LCCN QA76.76.H94F59 1998.

Fahrenholz-Mann:1999:PRS

[FM99] Sally Fahrenholz-Mann. Project report: SGML for electronic publishing at a technical society: Expectation meets reality. Markup Languages: Theory & Practice, 1(2):1–30, Spring 1999. CODEN MLTPFG. ISSN 1099-6621.

Fernandez:1995:UDH

[FMBM95] Mark R. Fernandez, Jim Miller, Jeffrey L. Becklehimer, and Bill Maloy. The use of distributed hypermedia at the Naval Research Laboratory. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
NationWideSites/fernandez/
fernandez.html. A selected few

of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Frett:1995:MPP

[FMC95] John J. Frett, Elizabeth O.

Mackenzie, and Dewey Caron. A multimedia presentation of plants in the University of Delaware Botanic Gardens utilizing Mosaic. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/
mackenzie/mackenzie.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Fernandez-Manjon:1998:EWE

[FMFVN98] B. Fernandez-Manjon, A. Fernandez-Valmayor, and A. Navarro. Extending Web educational applications via SGML structuring and content-based capabilities. In Verdejo and Davies [VD98], pages 244–259. ISBN 0-412-83550-9. LCCN LC5803.C65I353 1997.

Fuller:1993:SAL

[FMSDW93] M. Fuller, E. Mackie, R. Sacks-Davis, and R. Wilkinson. Structured answers for a large structured document collection. SI-GIR forum, ???? 1993. CODEN FASRDV. ISSN 0163-5840.

Fischer:1995:UEA

[FMW95] Michael Fischer, Michael Meyer, and Michael Witbrock. User extensibility in Amiga Mosaic. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        DDay/witbrock/amosaic/index.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Fukuda:1997:NVS

[FNN97] T. Fukuda, R. Nagahama, and J. Nomura. Networked VR system: Kitchen layout design for customers. In Spencer and Brutzman [SB97], pages 93–100. ISBN 0-89791-886-X. LCCN ???? URL http://

                                        www.acm.org/pubs/contents/
                                        proceedings/graph/253437/.

ACM order number 434973.

Fontaine:1993:GSS

[Fon93] J. Fontaine. Graduate students’ symposium on terminology: Le langage SGML et ses applications en terminologie. In Crochetière et al. [CBO93], pages 407–408. ISBN 2-7637-7353-2. LCCN P21.I6 1992. Four volumes.

Ford:1995:SWH

[For95a] Andrew Ford. Spinning the Web: How to Provide Information on the Internet. Van Nostrand Reinhold, New York, NY, USA, 1995. ISBN 0-442-01996-3 (Van Nostrand Reinhold), 1-85032-141-8 (ITP UK). xviii + 227 + 8 pp. LCCN TK5105.888 .F67 1995.

Ford:1995:CPH

[For95b] K. S. Ford. Converting from paper to HTML. In ACM [ACM95c], pages 115–118. ISBN 0-89791-704-9. LCCN QA 74 A26 1995.

Forer:1996:USC

[For96a] D. Forer. Using SGML to create accessible information. In Klaus et al. [KAKZ96], pages 219–228. ISBN 3-85403-087-8, 3-486-23797-7 (München), 3-7029-0413-1 (Wien). LCCN HV1569.5.T59 1996. Two volumes.

Forin:1996:CSR

[For96b] Y. Forin. Contribution of the SGML, and related standards to information retrieval systems. In Anonymous [Ano96t], pages 188–191. ISBN 2-901046-95-9. LCCN ????

Forowicz:1996:MED

[For96c] T. Forowicz. Modeling of energy demands for residential building with HTML interface. In Becker and Paciuk [BP96], pages 4– 115–4–126. ISBN 965-222-752-8. LCCN ????

Forowicz:1999:MED

[For99] T. Forowicz. Modeling of energy demands for residential buildings with HTML interface. In Anonymous [Ano99d], pages 481–488. ISBN ???? ISSN 0926-5805. LCCN ????

Fox:1995:MMM

[Fox95] James A. Fox. Maya mythology & multimedia: Using each to teach the other. T H E journal (Technological Horizons in Education), 23(5):64–??, December 1, 1995. CODEN THEJD4. ISSN 0192-592X.

Fischer:1995:UMW

[FP95] Thomas Fischer and Mike Pingleton. Utilizing Mosaic and the WWW in an operations environment. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/http://gto.ncsa.
uiuc.edu/tfischer/realpaper.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Freda:1997:AVR

[FR97] D. Freda and V. J. Realmuto. Application of the virtual reality modeling language (VRML) to geologic remote sensing. In Anonymous [Ano97d], pages II– 338. ISBN ???? ISSN 1067-0106. LCCN QE 33.2 R4 A67 1997. Two volumes.

Fratarcangeli:1991:TUQ

[Fra91] Claudio Fratarcangeli. Technique for universal quantification in SQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(3): 16–24, September 1991. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Frank:1995:DI

[Fra95] Maurice Frank. Database and the Internet. DBMS, 8(13):44– ??, December 1, 1995. CODEN DBMSEO. ISSN 1041-5173.

Francois:1996:GSR

[Fra96] P. François. Generalized SGML repositories: Requirements and modelling. Computer Standards and Interfaces, 18(1): 11–24, January 1996. CO-DEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

FraminanTorres:1997:MIH

[Fra97a] Jose Manuel Framinan Torres. Manual imprescindible de HTML y CGI. Manuales Imprescindibles. Anaya Multimedia, Madrid, Spain, 1997. ISBN 84-415-0057-6. xvi + 350 pp. LCCN ????

Franke:1997:PHH

[Fra97b] Matthias Franke. Programmation HTML: en 16 Heures chrono ! Micro Application, Paris, France, 1997. ISBN 2-7429-1005-0. 492 pp. LCCN ????

Fraternali:1999:TAD

[Fra99] Piero Fraternali. Tools and approaches for developing dataintensive Web applications: a survey. ACM Computing Surveys, 31(3):227–263, September 1999. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL http://

www.acm.org/pubs/articles/
journals/surveys/1999-31-
3/p227-fraternali/p227-fraternali.
pdf;   http://www.acm.org/
pubs/citations/journals/surveys/
1999-31-3/p227-fraternali/
;   http://www.acm.org/pubs/
citations/journals/surveys/
1999-31-3/p227-fraternali/
#indterms.

Freese:1995:TSD

[Fre95] Eric Freese. The transformation of SGML documents for presentation on the WWW. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Autools/freese/freese.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Freitag:1998:IEH

[Fre98] D. Freitag. Information extraction from HTML: Application of a general machine learning approach. In Anonymous [Ano98q], pages 517–523. ISBN 0-262-51098-7. LCCN Q334 .N36 1998.

Fry:1995:PNM

[Fry95a] Andrew L. Fry. Publishing in the new mass medium: Creating content on the Internet. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Overviews/fry/fry.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Frystyk:1995:TUL

[Fry95b] Henrik Frystyk. Towards a uniform library of common code: A presentation of the World Wide Web library. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
frystyk/LibraryPaper.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Ford:1995:OIL

[FS95] Shelley G. Ford and Robert C.

Stern. OmniPort: Integrating legacy data into the WWW. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/

                                        CorInfSys/stern/stern-ford.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Fox:1996:WPCa

[FS96] David Fox and Philip Shaddock. Web publisher’s construction kit with VRML/ Live3D. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN 1-57169-068-9. 600 pp. LCCN T385 .F684 1996. US$39.99. URL http:/

                                        /w3.waite.com/waite/waite/
                                        books.new/VRML_Constuction_
                                        Kit/html/vrmlckcov.html.

Farrar:1998:HEB

[FS98] Edward Farrar and Norman E. (Norman Earl) Smith. The HTML example book. Wordware Publishing, Plano, TX, USA, 1998. ISBN 1-55622-566-0 (paperback) 1-55622-517-2 (paperback). xvi + 492 pp. LCCN QA76.76.H94F37 1998.

Fensel:1999:KAM

[FS99] Dieter Fensel and Rudi Studer, editors. Knowledge acquisition modeling and management: 11th European Workshop, EKAW ’99, Dagstuhl Castle, Germany, May 26–29, 1999: proceedings, volume 1621 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1999. ISBN 3-540-66044-5 (softcover). LCCN QA76.76.E95 E92 1999.

Fuchs:1996:LTE

[Fuc96a] M. Fuchs. Let’s talk: extending the Web to support collaboration. In IEEE [IEE96j], pages 316–321. ISBN 0-8186-7446-6. LCCN TS155.6 .W56 1996.

Fuchs:1996:UID

[Fuc96b] M. Fuchs. The user interface as document: SGML and distributed applications. Computer Standards and Interfaces, 18(1):79–92, January 1996. CO-DEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Funke:1999:XHI

[Fun99] S. A. Funke. XML: How it will affect content, knowledge management, and information professionals. In Williams [Wil99a], pages 123–132. ISBN 1-57387-084-6. LCCN ????

Furuta:1990:EPI

[Fur90] R. Furuta, editor. EP90 Proceedings of the International Conference on Electronic Publishing, Document Manipulation and Typography. Cambridge University Press, New York, NY, USA, 1990. ISBN 0-521-40246-8. LCCN Z286.E43I58 1990.

Feng:1993:SGT

[FW93] A. Feng and T. Wakayama. SI-MON: a grammar-based transformation system for structured documents. Electronic Publishing—Origination, Dissemination, and Design, 6(4):361– 372, December 1993. CODEN EPODEU. ISSN 0894-3982.

Frieder:1996:PFI

[FW96] O. Frieder and J. Wigglesworth, editors. Proceedings of the Fourth International Symposium on Assessment of Software Tools, May 22–24, 1996, Toronto, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7389-3, 0-8186-7390-7 (softbound), 0-8186-7391-5 (microfiche). LCCN QA76.76.D47I595 1996. IEEE Catalog No. 96TB100054.

Fought:1993:ESC

[FWDV93] J. Fought, M. Wesler, H. Davenport, and C. Van Ess-Dykemag. Extending SGML concurrent structures: toward computerreadable meta-dictionaries. Literary and Linguistic Computing, 8(1):33–38, ???? 1993. CODEN LLCOEI. ISSN 0268-1145.

Fankhauser:1993:MIA

[FX93] P. Fankhauser and Y. Xu. Mark-ItUp! an incremental approach to document structure recognition. Electronic Publishing—Origination, Dissemination, and Design, 6(4):447– 456, December 1993. CODEN EPODEU. ISSN 0894-3982.

Finkelstein:1999:XSE

[FZ99] A. Finkelstein and A. Zisman. XML for software engineers. In IEEE [IEE99b], pages 193–194.

ISBN 0-7695-0188-5, 0-7695-0190-7 (microfiche). ISSN 1090-705X. LCCN QA76.751.I36 1999. IEEE Computer Society Order Number PR00188. IEEE Order Plan Catalog Number PR00188.

Gann:1996:CIO

[Gan96] R. Gann. Caere improves its OCR interface [OmniPage Pro for Windows 95]. PC User, 21 August-3 September 1996. CO-DEN PCUSDW. ISSN 0263-5720.

Gardner:1995:NAS

[Gar95] Reed M. Gardner, editor. Nineteenth annual Symposium on Computer Applications in Medical Care: toward cost-effective clinical computing: proceedings. October 28–November 1, 1995, New Orleans Hilton Riverside Hotel, New Orleans, Louisiana, volume 19 of Annual Symposium on Computer Applications in Medical Care. Hanley and Belfus, Philadelphia, PA, USA, 1995. ISBN 1-56053-123-1. ISSN 0195-4210. LCCN W 26.5 S978 1995.

Garber:1997:NBV

[Gar97a] Lee Garber. News: Binary version could bring VRML into mainstream. Computer, 30(4): 25–27, April 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garcia:1997:HCJ

[Gar97b] Alonso Alvarez Garcia. HTML CGI Java Servidores Tecnologia WWW. Anaya Multimedia, Madrid, Spain, 1997. ISBN 84-7614-959-X. ???? pp. LCCN

                                        ???? URL http://www.anaya.
                                        es/Catalogo/CGA?act=L&art=
                                        2321005&ed=2300;   http://
                                        www.anaya.es/multimedia/.

GGL:1998:HAA

[Gar98a] Gartner Group Learning. HTML advanced authoring, 1998. Includes CD-ROM.

GGL:1998:HAE

[Gar98b] Gartner Group Learning. HTML authoring essentials, 1998. Includes CD-ROM.

Garber:1999:NBL

[Gar99]  Lee Garber.  News briefs:

Linux support ranges from GUI to Big Blue; new XML standards; partnership for advanced computational infrastructure. Computer, 32(5):20– 22, May 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1999/pdf/r5020.
                                        pdf.

Gauthier:1996:SSC

[Gau96a] J. S. Gauthier, editor. Southeastern simulation conference:

— October 1996, Huntsville, AL, Proceedings of the Southeastern Simulation Conference 1996. Society for Computer Simulation, San Diego, CA, USA, 1996. ISBN ???? LCCN ????

Gautier:1996:DKL

[Gau96b] G. H. Gautier. Direji Kurdi:

a lexicographic environment for Kurdish language using 4th Dimension. In Anonymous [Ano96-33], pages 9/2/1– 9/2/25. ISBN ???? LCCN ????

Gaynor:1994:CET

[Gay94] E. Gaynor. Cataloging electronic texts: the University of Virginia Library experience. Library Resources and Technical Services, 38(4):403–413, October 1994. CODEN LRTSAH. ISSN 0024-2527.

Gaylord:1995:CR

[Gay95] H. E. Gaylord. Character representation. Computers and the humanities, 29(1):51–73, ???? 1995. CODEN COHUAD. ISSN 0010-4817.

Gossage:1994:MCT

[GB94] L. Gossage and B. Bayer. Moving college textbooks to CD-ROM. CD-ROM professional, 7 (2):90–92, March 1994. CODEN CRPFEX. ISSN 1049-0833.

Gee:1995:MDA

[GB95a] David A. Gee and Mike Boulter.

MosaicForms database access ?: A palaeobotanic case study. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
GenSci/gee/pfrwww.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Globus:1995:SUW

[GB95b] Al Globus and Chris Beaumont. Spinning a useful weblet. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/
globus/RNR-94-017.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Greenstein:1995:SOV

[GB95c] D. Greenstein and L. Burnard. Speaking with one voice: encoding standards and the prospects for an integrated approach to computing in history. Computers and the humanities, 29(2): 137–148, ???? 1995. CODEN COHUAD. ISSN 0010-4817.

Gardner:1998:NAN

[GB98] J. A. Gardner and V. Bulatov. Non-visual access to non-textual information through DotsPlus and accessible VRML. In Edwards et al. [EAZ98], pages 136–140. ISBN 3-85403-118-1. LCCN ????

German:1995:EZI

[GC95] D. M. German and D. D. Cowan. Experiments with the Z Interchange Format and SGML. In Bowen and Hinchey [BH95b], pages 224–236 (or 224–233??).

ISBN 3-540-60271-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.73.Z2 I58 1995.

Glushko:1994:HP

[GDRR94] R. Glushko, D. Dougherty, A. Rizk, and D. Russell. HTML — poison or panacea. In ACM [ACM94b], pages 245–246. ISBN 0-89791-640-9. LCCN QA76.76.I59A35 1994.

Gershon:1995:IVP

[GE95] N. Gershon and S. Eick, editors. Information visualization: proceedings, October 30– 31, 1995, Atlanta, Georgia, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7201-3 (paperback), 0-8186-7203-X (microfiche). LCCN TA1634.I325 1995. IEEE catalog no. 95TB100000.

Gelb:1995:RDM

[Gel95] Laurie Gelb. A representational data model for Web site administrators. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
InfOrg/gelb/Gelb.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Gennusa:1997:SB

[Gen97] Pamela L. Gennusa. SGML on the Block. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, January 1997. ISBN 0-13-190372-1. ???? pp. LCCN ???? US$31.50.

Gessert:1991:HMD

[Ges91] G. H. Gessert. Handling missing data by using stored truth values. SIGMOD Record (ACM Special Interest Group on Management of Data), 20(3):30– 42, September 1991. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gauch:1993:SAS

[GF93] S. Gauch and R. P. Futrelle. The Scientist’s Assistant: structurebased navigation of technical texts. In Anonymous [Ano93p], pages 367–377. ISBN ???? LCCN ????

Grevera:1996:WDI

[GFH96] G. J. Grevera, E. Feingold, and S. C. Horii. A WWW to DICOM interface. In Jost and Dwyer III [JD96], pages 109–117. CO-DEN PSISDG. ISBN 0-8194-2086-7. ISSN 0277-786X (print), 1996-756X (electronic). LCCN R857.P52 M427 1996.

Goeschka:1998:DAH

[GFR98] K. M. Goeschka, J. Falb, and W. Radinger. Database access with HTML and Java — A comparison based on practical experiences. In IEEE [IEE98c], pages 588–595. ISBN 0-8186-8585-9 (softbound), 0-7803-5168-1 (casebound), 0-8186-8587-5 (microfiche). ISSN 0730-3157. LCCN QA76.6 .I546 1998.

Guenette:1996:CWB

[GG96] D. R. Guenette and R. Gustavson. CD-ROM and Web browsers: ¡HTML¿ as the Lingua Franca. CD-ROM professional, 9(8):26–30, 32–34, 36, 38–42, 44, 46–48, 50–51, August 1996. CODEN CRPFEX. ISSN 1049-0833.

Graves:1997:IVM

[GHH97] A. Graves, C. Hand, and A. Hugill. Interactive visualisation of musical form using VRML. In Bowden [Bow97], pages 98–109. ISBN 1-872166-64-4, 1-872166-69-5. LCCN ????

Green:1995:SMT

[GHJ95] C. Green, D. G. High, and J. M. Jordan. SGML: More than “just do it!”. In Anonymous [Ano95-37], pages 273– 276. ISBN 0-914548-86-7, 0-914548-87-5. LCCN T 11 S63 1995.

Goel:1996:SQO

[GI96] Piyush Goel and Bala Iyer. SQL query optimization: Reordering for a general class of queries. SIGMOD Record (ACM Special Interest Group on Management of Data), 25(2):47–56, June 1996. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gibbs:1998:CVW

[Gib98] W. Wayt Gibbs. Cyber view: The Web learns to read. Scientific American, 268(6):40, June 1998. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). Discusses HTML, MathML, MusicML, ChemML, and XML.

Gillman:1992:DHT

[Gil92] P. Gillman. Data handling and text compression. Journal of Information Science, Principles and Practice, 18(2):105– 110, ???? 1992. CODEN JIS-CDI. ISSN 0165-5515 (print), 1741-6485 (electronic).

Gilmore:1993:ITS

[Gil93] Elizabeth Gilmore. Introducing today’s SGML. Technical Communication, 40(2):210–218, May 1993. CODEN TLCMBT. ISSN 0049-3155.

Giordano:1995:THD

[Gio95] R. Giordano. The TEI header and the documentation of electronic texts. Computers and the humanities, 29(1):75–84, ???? 1995. CODEN COHUAD. ISSN 0010-4817.

Godwin-Jones:1995:LLW

[GJ95] Robert Godwin-Jones. Language learning and the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

                                        http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Arts/
                                        godwin-jones/godwin.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Godwin-Jones:1996:IWC

[GJ96]  R. Godwin-Jones.  Interac-

tive Webbing: CGI scripting, JavaScript and linked programs for language learning. In Borchardt et al. [BBJR96], page ?? ISBN ???? LCCN ????

Gerdin-Jelger:1997:NII

[GJTW97] Ulla Gerdin-Jelger, Marianne Tallberg, and Paul Wainwright, editors. Nursing informatics: the impact of nursing knowledge on health care informatics/ Proceedings of Nursing Informatics 1997, the 6th international congress, Stockholm, Sweden, number 46 in Studies in Health Technology and Informatics. IOS Press, Amsterdam, The Netherlands, 1997. ISBN 90-5199-362-5, 4-274-90184-X. ISSN 0926-9630. LCCN ????

Glushko:1993:SGI

[GK93] Robert J. Glushko and Ken Kershner. Silicon Graphics’ IRIS In-Sight: An SGML success story. Technical Communication, 40 (3):394–401, August 1993. CO-DEN TLCMBT. ISSN 0049-3155.

Gessler:1995:PMWb

[GK95a] S. Gessler and A. Kotulla. PDAs as mobile WWW browsers. Computer Networks and ISDN Systems, 28(1–2):53–??, December 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Gessler:1995:PMWa

[GK95b] Stefan Gessler and Andreas Kotulla. PDAs as mobile WWW browsers. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
gessler/www_pda.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Ginsburg:1995:EPC

[GK95c] Mark Ginsburg and Ajit Kambil. The EDGAR project: A case study in disseminating financial data on the Internet. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Fin/ginsburg/edgar.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Grivel:1991:SGH

[GL91] L. Grivel and J.-C. Lamirel. SDOC —a generator of hypertext structures. In Feeney and Day [FD91], pages 67–81. ISBN 1-85739-090-3. LCCN QA76.575 .M85 1991.

Grivel:1993:SAT

[GL93] L. Grivel and J.-C. Lamirel.

SDOC, an analysis tool for scientometric studies integrated in an hypermedia environment. In Anonymous [Ano93a], pages 146–154. ISBN 2-7624-0555-6. LCCN ????

Galindo-Legaria:1994:OD

[GL94] César A. Galindo-Legaria. Outerjoins as disjunctions. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23(2):348–358, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Grosof:1999:AUX

[GL99] Benjamin N. Grosof and Yannis Labrou. An approach to using XML and a rule-based content language with an agent communication language. Research report RC. International Business Machines Corporation. Research Division ; 21491 International Business Machines Corporation. Research Division. Research report ; RC 21491. RC 21491 (96965), IBM T.J. Watson Research Center, Yorktown Heights, NY, USA, 1999. i + 21

pp. URL http://wwwis.win.
tue.nl/ac199/.

Glazer:1995:LLI

[Gla95] David Glazer. Lessons learned implementing a navigation server for the Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        HCI/glazer/glazer.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Glazman:1999:CFS

[Gla99] Daniel Glazman. CSS2: feuilles de styles HTML. Eyrolles, Paris, France, 1999. ISBN 2-212-09051-X. vi + 248 pp. LCCN ????

Gueziec:1998:SPP

[GLTH98] A. Gueziec, F. Lazarus, G. Taubin, and W. Horn. Surface partitions for progressive transmission and display, and dynamic simplification of polygonal surfaces. In Spencer and Roehl [SR98], pages 25–32. ISBN 1-58113-022-8. LCCN ???? URL http://

                                        www.acm.org/pubs/contents/
                                        proceedings/graph/271897/.

Glushko:1996:HPI

[Glu96] Robert J. Glushko. How practical is Practical SGML? compares the roles and benefits of this book’s two editions. Asterisk: the journal of computer documentation, 20(2):39–??, May 1, 1996. CODEN ASTRF7. ISSN 0731-1001.

Goldberg:1995:BWM

[GMG^{+}95a] K. Goldberg, M. Mascha, S. Gentner, J. Rossman, N. Rothenberg, C. Sutter, and J. Wiegley. Beyond the Web: Manipulating the real world. Computer Networks and ISDN Systems, 28(1– 2):209–??, December 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Goldberg:1995:BWE

[GMG^{+}95b] Ken Goldberg, Michael Mascha, Steven Gentner, Nick Rothenberg, Carl Sutter, and Jeff Wiegley. Beyond the Web: Excavating the real world via Mosaic. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/
goldberg/goldberg.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Gupta:1995:AMV

[GMR95] Ashish Gupta, Inderpal S. Mumick, and Kenneth A. Ross. Adapting materialized views after redefinitions. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):211–222, May 1995. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goldman:1996:CAP

[GMR96] C. V. Goldman, Y. Mor, and J. S. Rosenschein. Courtz: an agent that pleases you. In Anonymous [Ano96-31], pages 837–842. ISBN ???? LCCN ????

Gupta:1993:MVI

[GMS93] Ashish Gupta, Inderpal Singh Mumick, and V. S. Subrahmanian. Maintaining views incrementally. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2): 157–166, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goldfarb:1998:XWK

[GMS98] Charles F. Goldfarb, Sean Mc-Grath, and John Simpson. XML Web Kit. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-020354-8. ???? pp. LCCN ???? URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130203548.html.

Ginis:1995:CVE

[GN95] R. Ginis and D. Nadeau. Creating VRML extensions to support vector field visualization. In Nadeau and Moreland [NM95b], pages 13–20. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Gutfreund:1995:WOS

[GNSP95] Yechezkal-Shimon Gutfreund, John Nicol, Russel Sasnett, and Vincent Phuah. WWWinda: An orchestration service for WWW browsers and accessories. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/

Agents/gutfreund/gutfreund.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Gogolla:1990:NTS

[Gog90] Martin Gogolla. A note on the translation of SQL to tuple calculus. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(1):18–22, March 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Goldfarb:1981:GAD

[Gol81] C. F. Goldfarb. A generalized approach to document markup. ACM SIGPLAN Notices, 16(6):68–73, June 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldstein:1989:ORW

[Gol89] C. M. Goldstein. Online reference works and full text retrieval. In Anonymous [Ano89b], pages 171–177. ISBN 0-938734-34-2. LCCN QA76.55 .N37 1989.

Goldfarb:1991:SHS

[Gol91] Charles F. Goldfarb. Standards: HyTime: a standard for structured hypermedia interchange. Computer, 24(8):81– 84, August 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goldfarb:1994:SG

[Gol94] Charles F. Goldfarb. The SGML guide. Interleaf, Inc., Waltham, MA, USA, 1994. ISBN ???? 83 pp. LCCN ????

Gollifer:1995:IFI

[Gol95] Sue Gollifer. ‘imaging the future’: Issues of identity and authenticity of a work of art. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Arts/gollifer/gollifer.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Goldammer:1996:HCJ

[Gol96] G. Goldammer. HTML-script calls Java-applet… a new development technique under a software- and applicationstechnology configuration. IM Information Management, 11 (3):6–14, August 2, 1996. ISSN 0930-5181.

Goldie:1997:USC

[Gol97] P. Goldie. Using SGML to create complex interactive documents for electronic publishing. IEEE Transactions on Professional Communication, 40 (2):129–138, June 1997. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Golden:1998:WWW

[Gol98] B. L. Golden. The World Wide Web: HTML, CGI and Java applications in animal breeding. In Anonymous [Ano98n], pages 409–414. ISBN 1-86389-454-3. LCCN ????

Gonnet:1994:TDD

[Gon94] G. H. Gonnet. Text dominated databases, theory, practice and experience. In ACM [ACM94c], pages 301–302. ISBN 0-89791-642-5. LCCN ????

Goodwin:1994:EPK

[Goo94] R. Goodwin. Electronic publishing keeps California laws on track. Work Process Improvement Today, 17(2):19–21, November–December 1994. CO-DEN WPITE8. ISSN 1073-2233.

Goodman:1998:DHD

[Goo98] Danny Goodman. Dynamic HTML: The Definitive Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1998. ISBN 1-56592-494-0. xii + 1073 pp. LCCN QA76.76.H94 G66 1998. US$39.95. URL http:

//www.oreilly.com/catalog/
dhtmlref.

Gordon:1991:QSD

[Gor91] Thomas Gordon. The qwertz SGML document types: version 1.1 reference manual.

Technical Report Arbeitspapiere der GMD, Nr. 588 0723-0508, Gesellschaft für Mathematik und Datenverarbeitung mbH, Sankt Augustin, Germany, 1991. 59 pp.

Gordon:1995:QSS

[Gor95] Thomas F. Gordon. The qwertz synthesis of SGML and L^{A}TEX. Computer Standards and Interfaces, 17(1):25–33, January 1995. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Gottleber:1998:EHI

[Got98] Timothy T. Gottleber. Excellent HTML with an introduction to Java applets. Irwin/Mc-Graw Hill, Boston, MA, USA, 1998. ISBN 0-07-024772-2, 0-07-913675-3. xx + 428 pp. LCCN QA76.76.H94G68 1998.

Goldstein:1990:SOM

[GP90] C. M. Goldstein and M. F. Prettyman. SGML-encoding of an online medical reference work: experiences and observations. In Williams [Wil90], pages 137–144. ISBN 0-938734-44-X. LCCN ????

Goldfarb:1998:XH

[GP98] Charles F. Goldfarb and Paul Prescod. The XML Handbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-081152-1. xliv + 639 pp. LCCN QA76.76.H92G65 1998. US$44.95. URL http:

//www.phptr.com/ptrbooks/
ptr_0130811521.html.

Goldfarb:1999:MX

[GP99] Charles F. Goldfarb and Paul Prescod. Manual de XML. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 84-8322-105-5. xxix + 524 pp. LCCN ???? Translation to Spanish by Beatriz Paredes Alejandro Ruiz.

Goldfarb:1998:SBG

[GPE98] Charles F. Goldfarb, Steve Pepper, and Chet Ensign. SGML buyer’s guide: a unique guide to determining your requirements and choosing the right SGML and XML products and services. The Charles F. Goldfarb series on open information management. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-681511-1. xxxv + 1145 pp. LCCN QA76.76.H94G65 1997.

Goralski:1997:VEV

[GPV97] Walter Goralski, Matthew Poli, and Peter Vogel. VRML: exploring virtual worlds on the Internet. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-486960-5, 0-614-14500-7. xi + 481 pp. LCCN QA76.76.H94G67 1997. US$39.95. Includes some discussion of Java.

Graham:1999:XSG

[GQ99] Ian S. Graham and Liam Quin.

XML specification guide. Wiley, New York, NY, USA, 1999. ISBN 0-471-32753-0. xxii + 431 pp. LCCN QA76.76.H94 G7335 1999.

Goldfarb:1990:SH

[GR90] Charles F. Goldfarb and Yuri Rubinsky. The SGML Handbook. Clarendon Press, Oxford, UK, 1990. ISBN 0-19-853737-9 (hardcover). xxiv + 663 pp. LCCN Z286.E43 G64 1990. US$75.00.

Graham:1996:HSCa

[GR96] Ian Graham and Eric J. Ray. The HTML sourcebook: A complete guide to HTML. Technical Communication, 43(1):89– ??, ???? 1996. CODEN TL-CMBT. ISSN 0049-3155.

Gulbransen:1997:SEU

[GR97] David Gulbransen and Kenrick Rawlings. Special edition using dynamic HTML. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1482-5. xxiii + 579 pp. LCCN QA76.76.H94 G85 1997.

Goossens:1999:LWC

[GR99a] Michel Goossens and Sebastian Rahtz. The L^{A}TEX Web companion: integrating TEX, HTML, and XML. Tools and Techniques for Computer Typesetting. Addison-Wesley Longman, Reading, MA, USA, 1999. ISBN 0-201-43311-7. xxii + 522 pp. LCCN QA76.76.H94G66 1999. US$36.95. With Eitan M. Gurari and Ross Moore and Robert S. Sutor.

Gulbransen:1999:SEU

[GR99b] David Gulbransen and Kenrick Rawlings. Special edition using XML. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-7897-1996-7. 1000 pp. LCCN ????

GCA:1983:GSD

[Gra83] Graphic Communications Association. GCA Standard 101-1983, Document Markup Metalanguage: GenCode and the Standard Generalized Markup Language (SGML). Graphic Communications Association, 100 Daingerfield Road, Alexandria, VA 22314-2804, USA, 1983. ISBN 0-89740-244-8. xxiv + 264 pp. LCCN QA76.9.T48 G33 1983. US$48.00. Adopted by Department of Defense, 10 August 1983.

Grierson:1992:AAI

[GRA92] D. E. Grierson, G. Rzevski, and R. A. Adey, editors. Applications of Artificial Intelligence in Engineering VII. Comput. Mech. Pubications, Southampton, UK, 1992. ISBN 1-85166-787-3. LCCN TA174.I469 1992.

Graham:1995:DVV

[Gra95a] C. Graham. Database visualization and VRML. In Nadeau and Moreland [NM95b], pages 21– 24. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Graham:1995:BWW

[Gra95b] D. Graham. Browsing the World Wide Web for sequence analysis. II. SciTech Journal, 5(4):12–13, April 1995. CODEN SITJE8. ISSN 1072-0995.

Graham:1995:HS

[Gra95c] Ian S. Graham. The HTML sourcebook. Wiley, New York, NY, USA, 1995. ISBN 0-471-11849-4. xvi + 416 pp. LCCN QA76.76.H94 G73 1995. ????

Granger:1995:ISL

[Gra95d] Geeti Granger. The inside story of life at Wiley with SGML, L^{A}TEX and Acrobat. TUGboat, 16(2):151–154, June 1995. ISSN 0896-3207.

Grabowski:1996:WVL

[Gra96a] Ralph Grabowski. Will VRML 2.0 let me design a bolt faster? Computer Graphics World, 19 (12):22, 24, December 1996. CODEN CGWODH. ISSN 0271-4159.

Graham:1996:PMP

[Gra96b] Clay Graham. PC Magazine Programming VRML. Ziff-Davis Press, Emeryville, CA, USA, March 1996. ISBN 1-56276-380-6. ???? pp. LCCN ???? US$59.99.

Graham:1996:HSCb

[Gra96c] Ian S. Graham. The HTML sourcebook: a complete guide to HTML 3.0. Wiley, New York, NY, USA, second edition, 1996. ISBN 0-471-14242-5. xvii + 688 pp. LCCN QA76.76.H94G73 1996. US$29.95, CDN$39.50.

Grant:1996:SPT

[Gra96d] Paul Grant, editor. SUGI 21:

proceedings of the Twenty-First Annual SAS Users Group International Conference, Chicago, Illinois, March 10–13, 1996, SUGI 1996; CONF 21. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1996. ISBN 1-55544-747-3. LCCN ????

Gray:1996:WPA

[Gra96e] Daniel Gray. Web publishing with Adobe PageMill. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-458-8. ???? pp. LCCN QA76.76.H94 G734 1996.

Grabowski:1997:WPI

[Gra97a] Ralph Grabowski. The Web publisher’s illustrated quick reference: covers HTML 3.2 and VRML 2.0. Springer-Verlag Inc., New York, NY, USA, 1997. ISBN 0-387-94831-7 (paperback). viii + 258 pp. LCCN QA76.76.H94G72 1997.

Graham:1997:HSC

[Gra97b] Ian S. Graham. HTML Sourcebook: A Complete Guide to HTML 3.2 and HTML Extensions. Wiley, New York, NY, USA, third edition, February 1997. ISBN 0-471-17575-7. xix + 620 pp. LCCN QA76.76.H94G73 1997. US$29.95.

Graham:1997:HSS

[Gra97c]  Ian S. Graham.  HTML

stylesheet sourcebook. Wiley Computer Publishers, New York, NY, USA, 1997. ISBN 0-471-19664-9 (paperback). xii + 420 pp. LCCN QA76.76.H94 G733 1997.

Graham:1998:HS

[Gra98] Ian S. Graham. HTML 4.0 sourcebook: A Complete Guide to HTML 4.0 and HTML Extensions. Wiley computer publishing. Wiley, New York, NY, USA, fourth edition, 1998. ISBN 0-471-25724-9. xxiv + 631 pp. LCCN QA76.76.H94G73 1998.

Greenrich:1993:CDP

[Gre93] Edward Greenrich. CD-ROM data preparation enhancements. In Proceedings of the 14th National Online Meeting May 4– 6, 1993, New York, NY, USA, pages 159–163. Learned Information, Medford, NJ, USA, 1993. ISBN ???? LCCN ????

Greenberg:1995:OPM

[Gre95] Hope A. Greenberg. The Ovid Project: Metamorphosizing the metamorphoses. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/
greenberg/greenberg.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

GBPRO:1998:CEP

[Gre98a] Great Britain.Public Record Office. The core executive pilot HTML version, 1998.

URL http://www.pro.gov.uk/
finding/corehtml.htm.

Gregoire:1998:AVR

[Gre98b] David Remi Gregoire. Application of virtual reality modeling language (VRML), and the Java programming language to conceptual design of industrial tooling. ????, ????, ????, 1998. x + 97 pp.

Grimshaw:1995:RXE

[Gri95] A. Grimshaw. Rank Xerox and the European Patent Office. Information Management & Technology, 28(4):163–164, July 1995. CODEN IMTHEM. ISSN 0266-6960.

Gross:1993:GYD

[Gro93] Mark Gross. Getting your data in SGML. Technical Communication, 40(2):219–225, May 1993. CODEN TLCMBT. ISSN 0049-3155.

Gross:1995:UOW

[Gro95] C. Gross. Using OLE and WWW as a virtual corporate desktop. In Anonymous [Ano95q], pages A43/1–27 (or A42/1–4??). ISBN ???? LCCN ????

Groves:1996:WPW

[Gro96] Dawn Groves. The Web page workbook. Franklin, Beedle & Associates, Wilsonville, OR, USA, 1996. ISBN 1-887902-05-8. ???? pp. LCCN QA76.76.H94 G75 1996.

Gross:1997:RCL

[Gro97] M. Gross. The real costs of library conversions to SGML/ HTML: Myth vs reality. In Williams [Wil97b], pages 139– 144. ISBN 1-57387-043-9. LCCN ????

Gerrior:1997:HWQ

[GRSM97] S. Gerrior, M. Rodrigues, M. E. Stein, and F. C. McGill. To HTML or not: What are the questions? In Anonymous [Ano97b], pages 387–390. ISBN 0-914548-94-8, 0-914548-95-6. LCCN ????

Gerner:1987:KVS

[GS87] N. Gerner and O. Spaniol, editors. Kommunikation in Verteilten Systemen (Communication in Distributed Systems). Springer-Verlag Inc., New York, NY, USA, 1987. ISBN 3-540-17213-0. LCCN TK5105.5 .G561 1987.

Goossens:1995:THB

[GS95a] Michel Goossens and Janne Saarela. TEX to HTML and back. TUGboat, 16(2):174–214, June 1995. ISSN 0896-3207.

Goossens:1995:PIS

[GS95b] Michel Goossens and Janne Saarela. A practical introduction to SGML. TUGboat, 16(2): 103–145, June 1995. ISSN 0896-3207. A very clear presentation, with discussion of the HTML grammar, and tools for support of SGML and HTML.

Gwertzman:1995:CGP

[GS95c] J. S. Gwertzman and M. Seltzer.

The case for geographical pushcaching. In IEEE [IEE95d], pages 51–55. ISBN 0-8186-7081-9. LCCN QA76.76.O63.W6666 1995. IEEE catalog no. 95TH8059.

Gray:1994:QGB

[GSE^{+}94] Jim Gray, Prakash Sundaresan, Susanne Englert, Ken Baclawski, and Peter J. Weinberger. Quickly generating billion-record synthetic databases. SIGMOD Record (ACM Special Interest Group on Management of Data), 23 (2):243–252, June 1994. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Griffin:1998:CCE

[GSG98] Robert E. (Robert Edward)

Griffin, Carol B. Schiffman, and William J. Gibbs, editors. Connecting with the community: exploring resources for visual learning & expression: selected readings. 29th annual conference, University Park, PA. International Visual Literacy Association, ????, 1998. ISBN 0-945829-11-6. LCCN ????

GCA:1990:SM

[GSS90] Graphic Communications Association (U. S.), SGML Users’ Group, and SoftQuad Inc. SGML, the movie, 1990. 1 videocassette (18 min.).

Gobel:1996:VES

[GSV96] M. Gobel, P. Slavik, and J. J. Van Wijk, editors. Virtual environments and scientific visualization ’96: proceedings of the Eurographics workshops in Monte Carlo, Monaco, Feb. 19– 20, 1996, and in Prague, Czech Republic, April 23–25, 1996. Springer-Verlag Inc., New York, NY, USA, 1996. ISBN 3-211-82886-9. ISSN 0946-2767. LCCN QA76.9.H85 V546 1996.

Glushko:1999:XFA

[GTM99] Robert J. Glushko, Jay M. Tenenbaum, and Bart Meltzer. An XML framework for agentbased E-commerce. Communications of the ACM, 42(3):106– 109, 111–114, March 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic).  URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1999-42-3/p106-
                                        glushko/.

Guerrero:1998:V

[Gue98a] Javier Guerrero. VRML 2.0. Informatica profesional; 2. Abeto, Alcobendas (Madrid), Spain, 1998. ISBN 84-89973-24-5. 192 pp. LCCN ????

Guerrieri:1998:SDR

[Gue98b] E. Guerrieri. Software document reuse with XML. In Devanbu and Poulin [DP98b], pages 246–254. ISBN 0-8186-8377-5, 0-8186-8378-3 (case), 0-8186-8379-1 (microfiche). ISSN 1085-9098. LCCN QA76.76.R47 I587 1998. IEEE Order Plan Catalog Number 98TB100203. IEEE Computer Society Order Number PR08377.

Gulesian:1996:PW

[Gul96] M. Gulesian. Plugging into the Web. DBMS, 9(13):69–70, 72, December 1996. CODEN DBM-SEO. ISSN 1041-5173.

Gupta:1997:ISA

[Gup97] J. N. D. Gupta, editor. Information systems: Americas conference; 3rd — August 1997, Indianapolis, IN, Proceedings of the Americas Conference on Information Systems — Association for Information Systems 1997; 3rd. AIS, Pittsburgh, PA, USA, 1997. ISBN ???? LCCN ????

Guthrie:1998:FIW

[Gut98] Malcolm Guthrie. Forms: interactivity for the World Wide Web: creating HTML and PDF form documents. Adobe Press, Mountain View, CA, USA, 1998. ISBN 1-56830-455-2. 701 pp. LCCN HF5371.G9247 1998.

Goossens:1992:STP

[Gv92] M. Goossens and E. van Herwijnen. Scientific text processing. International Journal of Modern Physics C [Physics and Computers], 3(3):479–546, June 1992. CODEN IJMPEO. ISSN 0129-1831 (print), 1793-6586 (electronic).

Ganski:1987:ONS

[GW87] Richard A. Ganski and Harry K. T. Wong. Optimization of nested SQL queries revisited. SIGMOD Record (ACM Special Interest Group on Management of Data), 16(3):23–33, December 1987. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Gee:1994:NAA

[GW94] D. Gee and P. Woolliams. Novel approaches to automating the gathering of intelligence from the online community through the Internet. In Raitt and Jeapes [RJ94], pages 501–511. ISBN 0-904933-88-1. LCCN ????

Gorelik:1994:SRS

[GWD94] Alex Gorelik, Yongdong Wang, and Mark Deppe. Sybase replication server. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):469, June 1994. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Horton:1996:WPD

[H^{+}96] William K. (William Kendall) Horton et al. The Web page design cookbook: all the ingredients you need to create 5-star Web pages. Wiley, New York, NY, USA, 1996. ISBN 0-471-13039-7. xx + 649 pp. LCCN QA76.76.H94 W43 1996.

Haas:1993:IST

[Haa93] S. W. Haas. Introduction to SGML and the TEI. In Bonzi [Bon93], pages 285–?? ISBN 0-938734-78-4. ISSN 0044-7870. LCCN Z699.A1 A5 1993.

Hersch:1998:EPA

[HAB98] Roger Hersch, Jacques André, and Heather Brown, editors. Electronic publishing, artistic imaging, and digital typography: 7th International Conference on Electronic Publishing, EP ’98 held jointly with the 4th International Conference on Raster Imaging and Digital Typography, RIDT ’98, St. Malo, France, March/April 1998: proceedings, volume 1375 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-64298-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z250.7.I58 1998.

Hackborn:1997:IH

[Hac97] Dianne Hackborn. Interactive HTML. Thesis (m.s.), Oregon State University, Corvallis, OR, USA, 1997. 359 pp.

Hoffman:1995:CWI

[HAD^{+}95] Paul Hoffman, Randy Adams, Dale Dougherty, Robert Frasca, and Mark Ingalls. Commercial Web information providers. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Pub/hoffman/hoffman.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Hahn:1991:LER

[Hah91] Jane Hahn. L^{A}TEX for Everyone — A Reference Guide and Tutorial for Typesetting Documents Using a Computer. PersonalTEX Inc., 12 Madrona Street, Mill Valley, CA 94941, USA, 1991. ISBN 0-9631044-0-3. xi + 346 pp. LCCN Z253.4.L38 H34 1991. US$19.95. See [Pop92, PH92].

Hajagos:1993:DS

[Haj93] Lani Hajagos. Documents and SGML. UNIX review, 11(3): 38–??, March 1993. CODEN UNRED5. ISSN 0742-3136.

Hamilton:1997:SDS

[Ham97a] Kate Hamilton. SGML and the DSSSL Standard. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, January 1997. ISBN 0-13-652371-4. ???? pp. LCCN ???? US$33.75.

Hamilton:1997:WMP

[Ham97b] Michael Hamilton. Writing man pages in HTML. Linux Journal, 35:??, March 1997. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Hamlin:1997:SHH

[Ham97c] C. B. Hamlin. From SGML to HTML with help from TEX. In Beeton [Bee97b], pages 170–174. ISBN ???? ISSN 0896-3207. LCCN ????

Hamilton:1998:PHM

[Ham98] Michael Hamilton. The Python HTMLgen module. Linux Journal, 55:22, 24–26, November 1998. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL

ftp://ftp.ssc.com/pub/lj/
listings/issue55/2986.tgz.

Hara:1994:CIH

[Har94] S. Hara. Construction of the infrastructure for health expert system. In Ifeachor and Rosen [IR94], pages 97–104. ISBN ???? LCCN ????

Harawitz:1995:HAP

[Har95a] Howard Harawitz. HTML assistant pro, Version 1.5. Brooklyn North Software Works, Bedford, NS, Canada, 1995. ISBN ???? LCCN ????

Harris:1994:MIM

[Har95b] Judi Harris. Mining the Internet. making the hypertext happen: HTML commands in NCSA’s Mosaic. Computing Teacher, 22(4):36–41, December/January 1994–1995. ISSN 0278-9175.

Hartman:1996:VH

[Har96] Jed Hartman. The VRML Handbook. Addison-Wesley Longman, Reading, MA, USA, April 1996. ISBN 0-614-14427-2. ???? pp. LCCN ???? US$24.95.

Harold:1998:XEM

[Har98] Elliotte Rusty Harold. XML: extensible markup language. IDG Books, San Mateo, CA, USA, 1998. ISBN 0-7645-3199-9 (paperback). xxiii + 426 pp. LCCN QA76.76.H94H35 1998.

URL http://www.amazon.com/
exec/obidos/ISBN=0764531999/
cafeaulaitA/.

Harold:1999:XB

[Har99] Elliotte Rusty Harold. XML bible. I D G Books Worldwide, Indianapolis, IN, USA, 1999. ISBN 0-7645-3236-7. xxxiii + 1015 pp. LCCN QA76.76.H94 H34 1999.

Haycox:1993:SGM

[Hay93]  Jamie Haycox.  Standard

generalized markup language (SGML) as a basis for an intelligent data management system. In IEEE [IEE93a], pages 1017– 1020 (vol. 2). ISBN ???? LCCN ???? Two volumes. IEEE catalog no. 93CH3306-8.

Hayhoe:1994:SSV

[Hay94] George F. Hayhoe. Strategy or SNAFU? the virtues and vulnerabilities of SGML. In Wilkes [Wil94a], pages 378–379. ISBN ???? LCCN ???? IEEE cat no. 94CH3442-1.

Hughes:1995:PDS

[HB95] J. Steven Hughes and Ann M. Bernath. Planetary Data System Web catalog interface — another use of the Planetary Data System Data Model. Digest of Papers — IEEE Symposium on Mass Storage Systems, pages 263–273, 1995. CO-DEN DPISDX. ISSN 1051-9173. IEEE catalog number 95CB35860.

Heslop:1996:HPIb

[HB96] Brent D. Heslop and Larry Budnick. HTML publishing on the Internet for Windows: create great-looking documents online: home pages, newsletters, catalogs, ads & forums. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-229-1. xxii + 480 pp. LCCN QA76.76.H94 H483 1995. US$34.95.

Honeycutt:1997:HSK

[HB97a] Jerry Honeycutt and Mark Robbin Brown. HTML 3.2 starter kit. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1249-0. xxvii + 450 pp. LCCN QA 76.76 H94 H658 1997.

Honeycutt:1997:UH

[HB97b] Jerry Honeycutt and Mark Robbin Brown. Using HTML 3.2. Que Corporation, Indianapolis, IN, USA, third edition, 1997. ISBN 0-7897-1097-8 (paperback). xxxvii + 1042 pp. LCCN ????

Hoadley:1998:ISA

[HB98] E. D. Hoadley and I. Benbasat, editors. Information systems: Americas conference; 4th — August 1998, Baltimore, MD, Proceedings of the Americas Conference on Information Systems — Association for Information Systems 1998; 4th. AIS, Pittsburgh, PA, USA, 1998. ISBN ???? LCCN ????

Hallam-Baker:1995:SSE

[HBNL95] Phillip M. Hallam-Baker, Henrik Frystyk Nielsen, and Ari Luotonen. SHEN security enhancements to HTTP. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Security/hallam-baker/Shen/
                                        Documentation/paper.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Husberg:1997:SPI

[HBNV97] F. Husberg, M. Blomqvist, M. Niskanen, and K. Visala. Shared platform independent VRML worlds. In Halttunen and Tuokko [HT97], pages 146– 151. ISBN 952-5183-03-3. LCCN ????

Heid:1997:SHC

[HC97] Jim Heid and David Cruz Langreo. Los secretos de HTML y creacion de paginas Web. (Spanish) [Secrets of HTML and Web page creation]. Anaya Multimedia, Madrid, Spain, 1997. ISBN 84-415-0358-3. 741 pp. LCCN ????

Hayes:1999:SCV

[HC99] C. Hayes and P. Cunningham. Shaping a CBR view with XML. Lecture Notes in Computer Science, 1650:468–481, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hankin:1995:LVE

[HD95] Steve Hankin and Jerry Davison. Live visualization and extraction of climate data with Mosaic and FERRET. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
EarSci/hankin/hankin.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Hinds:1994:EMI

[HE94] N. Hinds and A. W. England.

An experimental Mosaic interface to scientific information systems. In IEEE [IEE94b], pages 863–865 (vol. 2). ISBN 0-7803-1497-2. LCCN QE33.2.R4I57 1994. Four volumes. IEEE catalog no. 94CH3378-7.

Hassinger:1995:MGV

[HE95] Sebastian Hassinger and Mike Erwin. 60 minute guide to VRML. IDG Books, San Mateo, CA, USA, 1995. ISBN 1-56884-710-6. xx + 218 pp. LCCN TK 5105.888 H38 1995.

Heba:1996:DAS

[Heb96] G. M. Heba. Digital architectures: SGML, HTML, multimedia and the structure of electronic documents. In Anonymous [Ano96b], pages 213–216. ISBN 0-914548-89-1, 0-914548-90-5. LCCN ????

Hefley:1995:UUD

[Hef95] W. E. Hefley. Understanding, using and designing for educational uses of the World-Wide Web. In Ibrahim [Ibr95], pages 431–432. ISBN 3-540-58951-1. LCCN QA76.758.S45 1995.

Heimbuerger:1994:ISG

[Hei94a] A. Heimbuerger. Introduction to Standard Generalized Markup Language (SGML). Microcomputers for Information Management, 11(4):239–??, 1994. CO-DEN MIIMEW. ISSN 0742-2342.

Heimburger:1994:ISG

[Hei94b]  Anneli Heimburger.  Intro-

duction to Standard Generalized Markup Language (SGML). Microcomputers for Information Management, 11(4):239– 260, December 1994. CODEN MIIMEW. ISSN 0742-2342.

Heine:1995:PND

[Hei95] M. H. Heine. A provisional notation for describing the information structure of documents. Journal of Documentation, 51(4):339–359, December 1995. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Heinrichs:1996:BCW

[Hei96] L. R. Heinrichs. Building cool worlds with VRML. In Little [Lit96], pages 5–?? ISBN ???? LCCN ????

Heid:1997:HWP

[Hei97] Jim Heid. HTML and Web publishing secrets. I D G Books Worldwide, Indianapolis, IN, USA, 1997. ISBN 0-7645-4003-3 (paperback). xxxii + 626 pp. LCCN QA 76.76 H94 H45 1997. US$49.99.

Hellerstein:1994:PPP

[Hel94] Joseph M. Hellerstein. Practical predicate placement. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):325–335, June 1994. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Held:1996:PNI

[Hel96] G. Held. Providing network information and assistance via the Web. International Journal of Network Management, 6(2):69– 76, March–April 1996. ISSN 1055-7148.

Henderson:1990:APA

[Hen90] D. Henderson, editor. ASIS ’90.

Proceedings of the 53rd ASIS Annual Meeting. American Society for Information Science, Medford, NJ, USA, 1990. ISBN 0-938734-48-2. LCCN Z699.A1 .A506 v.27.

Henderson:1991:CSD

[Hen91] L. R. Henderson. CALS as a solution for digital delivery of technical documents. Computer Aided Design, 23(4):252–256, May 1991. CODEN CAIDA5. ISSN 0010-4485 (print), 1879-2685 (electronic).

Hendry:1995:BWI

[Hen95] D. G. Hendry. Breakdowns in writing intentions when simultaneously deploying SGMLmarked texts in hard copy and electronic copy. Behaviour & information technology, 14(2):80– 92, March–April 1995. CODEN BEITD5. ISSN 0144-929X.

Henderson:1996:VRW

[Hen96a] J. F. Henderson. Virtual reality on the World Wide Web: a survey of Web sites. International Journal of Virtual Reality, 2(1):5–16, Winter 1996. CO-DEN IJVRF8. ISSN 1081-1451.

Henry:1996:LOV

[Hen96b] M. Henry. Library OPAC vendors and the Web: Innopac, Ameritech, DRA. CWIS Campus-Wide Information Services, 13(2):10–12, ???? 1996. CODEN CISYF9. ISSN 1065-0741.

Henderson:1999:DII

[Hen99] D. B. Henderson. Distributing 3-D information over the Internet using VRML. In Anonymous [Ano99h], pages 561–566. ISBN ???? LCCN ????

Herbst:1996:WI

[Her96] Kris Herbst. Webfest IV. Internet World, 7(3):22–??, March 1, 1996. CODEN IERNE8. ISSN 1064-3923.

Heslop:1997:HPI

[Hes97] Heslop. HTML Publishing on the Internet. Ventana Communications Group, Incorporated, Research Triangle Park, NC, USA, February 1997. ISBN 1-56604-625-4. ???? pp. LCCN QA76.76.H94H48 1998. US$49.99.

Heudin:1998:VWF

[Heu98] J. C. Heudin, editor. Virtual worlds: first international conference, VW ’98, Paris, France, July 1–3, 1998: proceedings, volume 1434 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. ISBN 3-540-64780-5 (softcover). LCCN QA76.9.H85 V628 1998 Bar.

Heylighen:1994:WWD

[Hey94]  F. Heylighen.  World-Wide

Web: a distributed hypermedia paradigm for global networking. In Anonymous [Ano94m], pages 355–368. ISBN ???? LCCN ????

Hollan:1997:PAM

[HF97] James D. Hollan and James D.

Foley, editors. Proceedings: ACM Multimedia ’97, November 9–13, 1997, Seattle, Washington, USA. Addison-Wesley and ACM Press, Reading, MA, USA and New York, NY, USA, 1997. ISBN 0-201-32232-3 (Addison-Wesley), 0-89791-991-2 (ACM). LCCN QA76.575.A36 1997.

He:1996:OCB

[HG96a] Jianzhang He and F. Gey. Online codebook browsing and conversational survey analysis. Social science computer review, 14 (2):181–186, Summer 1996. CO-DEN SSCREH. ISSN 0894-4393.

JianzhangHe:1996:OCB

[HG96b] Jianzhang He and F. Gey. Online codebook browsing and conversational survey analysis. Social science computer review, 14 (2):181–186, Summer 1996. CO-DEN SSCREH. ISSN 0894-4393.

Heslop:1996:HPIa

[HH96] Brent D. Heslop and David A. Holzgang. HTML publishing on the Internet for Macintosh: create great-looking documents online: home pages, newsletters, catalogs, ads & forms. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-228-3. xxvi + 539 pp. LCCN QA76.76.H94H48 1995. US$49.95. Includes CD-ROM.

Heslop:1998:HPI

[HH98] Brent D. Heslop and David A. Holzgang. HTML Publishing on the Internet: everything you need to create professionallooking Web pages. Ventana Communications Group, Incorporated, Research Triangle Park, NC, USA, second edition, May 1998. ISBN 1-56604-625-4. xxvi + 721 pp. LCCN QA76.76.H94H48 1998. US$49.95.

Hyman:1999:DHD

[HH99a] Michael I. Hyman and Nate Holdread. Dynamic HTML for dummies. –For dummies. I D G Books Worldwide, Indianapolis, IN, USA, second edition, 1999. ISBN 0-7645-0467-3. xxiv + 405 pp. LCCN ????

Hyunhee:1999:BXW

[HH99b] K. Hyunhee and C. Hyewon.

Building an XML and Webbased document retrieval system. In Williams [Wil99a], pages 251–262. ISBN 1-57387-084-6. LCCN ????

Hickey:1989:UST

[Hic89] T. B. Hickey. Using SGML and TEX for an interactive chemical encyclopedia. In Anonymous [Ano89b], pages 187–195. ISBN 0-938734-34-2. LCCN QA76.55 .N37 1989.

Hickey:1994:ISP

[Hic94] T. Hickey. Integrating SGML and page databases. In Anonymous [Ano94f], pages 690–?? ISBN 0-89208-177-5. LCCN ????

Hierle:1997:HJI

[Hie97] Alexander Hierle. HTML und Java: Internet-Software kreativ nutzen. Sybex Verlag GmbH, Erkrather Straße 345-349, 40231 Düsseldorf, Germany, 1997. ISBN 3-8155-7509-5. ???? pp. LCCN ???? 20 DM. URL http:

//www.sybex.de/. Includes two

CD-ROMs.

Hightower:1996:HCF

[Hig96] Lauren Hightower. HTML conversion and FTP automation. Dr. Dobb’s Journal of Software Tools, 21(4):82, 84, 86, 94, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Heid:1998:HKW

[HiKgK98] Jim Heid, Kyong il Kim, and Chae gak Kim. HTML kwa Web chulpan ui pimil = HTML and Web publishing secrets. Yongjin, Soul, Korea, 1998. ISBN 89-314-0841-2 (paperback). 779 pp. LCCN ????

Hillmann:1996:PUM

[Hil96] D. I. Hillmann. “parallel universes” or meaningful relationships: envisioning a future for the OPAC and the net. Cataloging & classification quarterly, 22(3-4):97–103, ???? 1996. CO-DEN CCQUDB. ISSN 0163-9374.

Hinchley:1998:XPS

[Hin98] A. Hinchley. XML: The power to solve everything. In Teasdale [Tea98b], pages 50–?? ISBN ???? LCCN ????

Hirsch:1995:ONa

[Hir95a] Marina Hirsch. On-line news. Publish!: the how-to magazine of desktop publishing, 10(7):32– ??, July 1, 1995. CODEN PBLIEJ. ISSN 0897-6007.

Hirsch:1995:ONb

[Hir95b] Marina Hirsch. On-line news. Publish!: the how-to magazine of desktop publishing, 10(8):28– ??, August 1, 1995. CODEN PBLIEJ. ISSN 0897-6007.

Hirsch:1995:ONc

[Hir95c] Marina Hirsch. Online news. Publish!: the how-to magazine of desktop publishing, 10(9):30– ??, September 1, 1995. CODEN PBLIEJ. ISSN 0897-6007.

Hayami:1996:WMS

[HIYO96] H. Hayami, T. Iwakiri, K. Yamaguchi, and K. Ohbayashi. Workflow management system on the Internet. NTT R and D, 45(3):291–298, ???? 1996. CODEN NTTDEC. ISSN 0915-2326.

Hsu:1995:AOH

[HJM95] Jui-Chieh Hsu, William Johnston, and John McCarthy. Active outlining for HTML documents: An X-Mosaic implementation. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/HCI/

hsu/hsu.html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Hendin:1998:MVR

[HJS98] O. Hendin, N. W. John, and O. Shochet. Medical volume rendering over the WWW using VRML and JAVA. In Westwood et al. [W^{+}98b], pages 34– 40. ISBN 90-5199-386-2, 4-274-90210-2. ISSN 0926-9630. LCCN W26.5 M45 1998.

Herzner:1994:MOD

[HK94] Wolfgang Herzner and Frank Kappe, editors. Multimedia/ Hypermedia in Open Distributed Environments: Proceedings of the Eurographics Symposium in Graz, Austria, June 6–9, 1994. Springer-Verlag Inc., New York, NY, USA, 1994. ISBN 3-211-82587-8. LCCN QA76.575 .M833 1994.

Hastings:1995:PCI

[HK95a] Edwin Hastings and Dilip H. Kumar. Providing customers information using the WEB and CORBA. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
hastings/hastings.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Hunter:1995:PSB

[HK95b] Lawrence Hunter and Teri E. Klein, editors. Pacific Symposium on Biocomputing ’96: Hawaii, USA, 3–6 January, 1996. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1995. ISBN 981-02-2578-4. LCCN QH323.5.P33 1996.

Harris:1996:OHP

[HK96a] Stuart Harris and Gayle Kidder. Official HTML Publishing for Netscape, Mac: Mac Edition. Ventana Communications Group, Incorporated, Research Triangle Park, NC, USA, September 1996.

ISBN 1-56604-417-0. 600 pp. LCCN QA76.73.H85H37 1996. US$39.99.

Harris:1996:HPN

[HK96b] Stuart (Stuart H.) Harris and Gayle Kidder. HTML publishing for Netscape: your guide to online design & production: Windows edition. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-288-7. xxvi + 564 pp. LCCN QA76.76.H94H372 1995. Includes CD-ROM.

Harris:1998:OND

[HK98] Stuart (Stuart H.) Harris and Gayle Kidder. Official Netscape dynamic HTML developer’s guide: Windows and Macintosh. Ventana Press, Chapel Hill, NC, USA, 1998. ISBN 1-56604-797-8. xxii + 344 pp. LCCN QA76.76.H94H374 1998.

Halama:1995:IEB

[HKH95] James R. Halama, Anne Kreft, and Robert E. Henkin. An interactive electronic bulletin board implementation for Mosaic and HTTP server. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/
halama/halama.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Hatonen:1996:KDT

[HKM^{+}96] K. Hatonen, M. Klemettinen, H. Mannila, P. Ronkainen, and H. Toivonen. Knowledge discovery from telecommunication network alarm databases. In Su [Su96], pages 115–122. ISBN 0-8186-7240-4. LCCN QA76.9.D3 I5582 1996. IEEE Catalog No. 96CB35888.

Herrmann:1995:CAA

[HL95a] T. Herrmann and U. Lasche. CGM and ATA at work at Lufthansa. In Anonymous [Ano95e], pages 410–422. ISBN ???? LCCN ????

Holman:1995:ISF

[HL95b] G. K. Holman and R. N. Luoma. Implementing SGML: The first steps. In Anonymous [Ano95-37], pages 493–?? ISBN 0-914548-86-7, 0-914548-87-5. LCCN T 11 S63 1995.

Holman:1996:ISF

[HL96] G. K. Holman and R. N. Luoma. Implementing SGML: The first steps. In Anonymous [Ano96b], page 569. ISBN 0-914548-89-1, 0-914548-90-5. LCCN ????

Husain:1997:JDR

[HL97] Kamran Husain and Jason Levitt. JavaScript developer’s resource: client-side programming using HTML, Netscape plug-ins and Java applets. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-267923-X. xxiv + 568 pp. LCCN QA76.73.J39 H87 1997. US$39.95.

Hlava:1996:SAT

[Hla96] M. Hlava. SGML authoring tools. In Whitney [Whi96f], pages 365–?? ISBN 1-57387-028-5. LCCN ????

Heid:1998:SHP

[HLB98] Jim Heid, Dorota Lewandowska, and Marzena Baranowska. Sekrety HTML i publikacji WWW. Seria bestsellerow. Wydawn. PLJ, Warszawa, 1998. ISBN 83-7101-382-5. 662 pp. LCCN ????

Harmon:1996:HPE

[HLC96] C. M. Harmon, R. J. Logan, and S. Comer. HTML as a prototyping environment. In Anonymous [Ano96s], pages 862–?? ISBN 0-945289-06-5. ISSN 1071-1813. LCCN TA166.A2 H8 1996. Two volumes.

Hudson:1997:VAC

[HLCG97] T. C. Hudson, M. C. Lin, J. Cohen, and S. Gottschalk. V-COLLIDE: Accelerated collision detection for VRML. In Spencer and Brutzman [SB97], pages 117–124. ISBN 0-89791-886-X. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/253437/.

ACM order number 434973.

Hyun:1993:ISD

[HLL93] Deck Chang Hyun, Kwang Taek Lim, and Soo Youn Lee. Implementation of SGML document editor using SGML basic parser. Journal of the Korea Information Science Society = Chongbo Kwahakhoe nonmunji, 20(4):484–494, April 1993. CO-DEN HJKHDC. ISSN 0258-9125.

Hylands:1997:TUI

[HLR97] Christopher Hylands, Edward A. Lee, and H. John Reekie. The Tycho user interface system. In USENIX [USE97a], pages 149– ?? ISBN 1-880446-87-1. LCCN QA76.73.T44 T44 1997.

Hagsand:1997:UST

[HLS97] O. Hagsand, R. Lea, and M. Stenius. Using spatial techniques to decrease message passing in a distributed VE system. In Spencer and Brutzman [SB97], pages 7–16. ISBN 0-89791-886-X. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/253437/.

ACM order number 434973.

Hwang:1994:MFD

[HLY^{+}94] S.-Y. Hwang, E.-P. Lim, H.-R. Yang, S. Musukula, K. Mediratta, M. Ganesh, D. Clements, J. Stenoien, and J. Srivastava. The MYRIAD federated database prototype. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):518, June 1994. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Henderson:1997:SVE

[HM97] D. B. Henderson and H. L. McKague. Spatio-temporal visualization of earthquake hypocenters over the Internet using VRML.

In Anonymous [Ano97a], pages 369–373. ISBN 1-57083-043-6 (set), 1-57083-044-4 (vol. 1), 1-57083-045-2 (vol. 2), 1-57083-047-9 (vol. 4), 1-57083-048-7 (vol. 5). LCCN ???? Five volumes.

Hirukawa:1997:PST

[HMHK97] H. Hirukawa, T. Matsui, S. Hirai, and K. Konaka. A prototype of standard teleoperation systems on an enhanced VRML. In IEEE [IEE97e], pages 1801– 1806. ISBN 0-7803-4119-8 (softbound), 0-7803-4120-1 (casebound), 0-7803-4121-X (microfiche). LCCN TJ210.3.I443 1997. Three volumes. IEEE catalog number 97CH36108.

Hurley:1995:CPC

[HMMSS95] S. Hurley, A. D. Marshall, S. N.

McIntosh-Smith, and N. M. Stephens. Courseware for parallel computing using Mosaic and the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Educ/hurley/hurley.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Honda:1995:VSE

[HMRL95] Y. Honda, K. Matsuda, J. Rekimoto, and R. Lea. Virtual society: extending the WWW to support a multi-user interactive shared 3D environment. In Nadeau and Moreland [NM95b], pages 109–116. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Halstead-Nussloch:1996:UHO

[HN96a] R. Halstead-Nussloch. Using HTML for an online survey of risk perception: An experience study. In Mintz [Min96], pages 15–16. ISBN 0-7918-1538-2. LCCN T55.A1 S226 1996.

Hayward:1996:HPI

[HN96b] Michael Hayward and Maureen Nicholson. How to publish on the Internet: build your own home page on the World Wide Web. Motion Works Publishing, San Francisco, CA, USA, 1996. ISBN 1-57712-003-5. ???? pp. LCCN QA76.76.H94 H39 1996.

Hobuss:1998:BOW

[Hob98] James J. Hobuss. Building Oracle Websites. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-079841-X. xii + 488 pp. LCCN QA76.625.H64 1998. US$44.95. Discusses Java, JavaScript, and HTML.

Hockey:1993:DAE

[Hoc93a] S. Hockey. Developing access to electronic texts in the humanities. Computers in Libraries, 13 (2):41–43, February 1993. CO-DEN CPLIE8. ISSN 1041-7915.

Hockey:1993:ESS

[Hoc93b] S. Hockey. Encoding standards: SGML and the Text Encoding Initiative: What and why? In Okerson [Oke93], pages 59– 64. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Hoess:1997:THD

[Hoe97] L. Hoess. Translating hardcopy documents to HTML online documents. In Anonymous [Ano97b], pages 378–381. ISBN 0-914548-94-8, 0-914548-95-6. LCCN ????

Holloway:1987:IGC

[Hol87] Henry Holloway. An introduction to generic coding and SGML. British Library research paper 27. British Library Board, London, UK, 1987. ISBN 0-7123-3132-8 (spiral). 20 pp. LCCN ????

Holyer:1993:WWW

[Hol93] A. Holyer. Wandering the World-Wide Web. Artificial Intelligence and Simulation of Behavior quarterly, Winter 1993. CODEN AISBEJ. ISSN 0268-4179.

Holland:1995:AWW

[Hol95] William H. Holland. AT&T and the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
CorInfSys/holland/mospap.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Holzner:1997:WDG

[Hol97]  Steven Holzner.  Web de-

veloper.com guide to dynamic HTML. Wiley, New York, NY, USA, 1997. ISBN 0-471-24102-4 (paperback). xii + 484 pp. LCCN QA76.76.H94H65 1997.

Holzner:1998:XC

[Hol98] Steven Holzner. XML Complete. McGraw-Hill, New York, NY, USA, 1998. ISBN 0-07-913702-4. xiv + 515 pp. LCCN QA76.76.H94H652 1998. US$44.95. URL http://www.

                                        amazon.com/exec/obidos/ISBN=
                                        0-07-913702-4/r/8788-6902305-
                                        556196.

Holzschlag:1999:SOH

[Hol99a] Molly E. Holzschlag. Short order HTML 4. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-7897-2049-3. 300 pp. LCCN ????

Holzschlag:1999:SEU

[Hol99b] Molly E. Holzschlag. Special edition using HTML 4. Que Corporation, Indianapolis, IN, USA, fifth edition, 1999. ISBN 0-7897-1851-0. xxiii + 1168 pp. LCCN QA76.76.H94H653 1999¡.

Homer:1999:XIP

[Hom99] Alex Homer. XML in IE5: Programmer’s Reference. Wrox Press, Chicago, IL, USA, 1999. ISBN 1-86100-157-6. xi + 479 pp. LCCN QA76.76.H94 H6575 1999. US$29.99.

GaoHong:1995:TLS

[Hon95] Gao Hong. Transformation list for SGML application. Journal of Computer Science and Technology [English Language Edition], 10(5):455–462, September 1995. CODEN JCTEEM. ISSN 1000-9000.

Houser:1998:SUI

[Hou98] R. Houser. Strategies for using information types in HTML help. In Anonymous [Ano98u], pages 354–358. ISBN 1-891709-00-3, 1-891709-01-1. LCCN ????

Howlett:1996:SWP

[How96] D. Howlett. Sophisticated Web page creation [HoTMetaL Pro 3.0]. PC User, ??(287):47–??, July 10–23, 1996. CODEN PCUSDW. ISSN 0263-5720.

Howard:1999:WAH

[How99] Todd M. Howard. Who’s afraid of HTML? Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 0-12-356915-X, 0-12-356916-8 (CD-ROM). xiv + 196 pp. LCCN QA76.76.H94 H69 1999.

Hoyle:1996:SSW

[Hoy96] L. Hoyle. SAS software and the WWW — what next? In Anonymous [Ano96-36], pages 867–874. ISBN ???? LCCN ???? Two volumes.

Huynh:1995:CCM

[HPS95] Minh Huynh, Laird Popkin, and Matthew Stecker. Constructing a corporate memory infrastructure from Internet discovery technologies. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
CorInfSys/huynh/cmi.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Hill:1995:CHU

[HRS95] Will Hill, Mark Rosenstein, and Larry Stead. Community and history-of-use navigation. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
HCI/hill/home-page.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Hellerstein:1993:PMO

[HS93] Joseph M. Hellerstein and Michael Stonebraker. Predicate migration: optimizing queries with expensive predicates. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22(2):267–276, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Hendl:1997:MZI

[HS97a] Jan Hendl and Vilem Sklenak. Metody zpracovani informaci. III, Hypertextove systemy (vcetne uvodu do jazyka HTML). Vysoka skola ekonomicka, Praha, Czech Republic, 1997. ISBN 80-7079-896-3. 293 pp. LCCN ????

Hong:1997:ROD

[HS97b] T. Hong and S. Srihari. Representing OCRed documents in HTML. In IEEE [IEE97f], pages 831–834. ISBN 0-8186-7899-2, 0-8186-7898-4, 0-8186-7900-X. LCCN TA1640 .I57 1997. Two volumes. IEEE Order catalog number 97TB100138. IEEE Computer Society order number PR07898.

Hutchison:1998:MAS

[HS98] David Hutchison and Ralf Schafer, editors. Multimedia applications, services, and techniques: ECMAST ’98, third European conference, Berlin, Germany, May 26–28, 1998: proceedings, volume 1425 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. ISBN 3-540-64594-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1425.

Heston:1999:DHE

[HS99] Frank Heston and Klaus Schöning.

Dynamic HTML: Einführung und Cross-Browser-Lösungen. Siemens, Berlin, Germany, 1999. ISBN 3-89578-122-3. 374 pp. LCCN ????

Han:1999:DIS

[HSCZ99] S.-G. Han, J.-H. Son, J.-W. Chang, and Z.-C. Zhoo. Design and implementation of a structured information retrieval system for SGML documents. In Chen and Lochovsky [CL99], pages 81–88. ISBN 0-7695-0084-6, 0-7695-0085-4 (casebound), 0-7695-0086-2 (microfiche). LCCN QA76.9.D3 I5565 1999.

Henri:1995:IGA

[HT95] Philip Tsang James Henri and Sandy Tse. Internet growth in Australia and Asia’s Four Dragons. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Int/
tsang/tsang.html. A selected

few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Halttunen:1997:ISM

[HT97] J. Halttunen and R. Tuokko, editors. International symposium on measurement and control in robotics: topical workshop on virtual reality and advanced man-machine interfaces: International symposium; 7th — June 1997, Tampere, Finland, number 9 in IMEKO -World Congress- 1997; B. Finnish Society of Automation, ????, 1997. ISBN 952-5183-03-3. LCCN ????

Hong:1995:UWT

[HTL95] Jack Hong, George Toye, and Larry Lefier. Using the WWW for a team-based engineering design class. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/
hong/hong.html. A selected few

of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Homer:1997:IID

[HU97a] Alex Homer and Chris Ullman. Instant IE4 dynamic HTML, IE4 edition. Wrox Press, Chicago, IL, USA, 1997. ISBN 1-86100-068-5. xxi + 567 pp. LCCN QA76.76.H94H657 1997.

Homer:1997:IND

[HU97b] Alex Homer and Chris Ullman. Instant Netscape dynamic HTML, NC4 edition. Wrox Press, Chicago, IL, USA, 1997. ISBN 1-86100-119-3. xv + 431 pp. LCCN QA 76.76 H94 H6562 1997.

Huck:1997:CEH

[Huc97] Susan R. Huck. A comparison of the effectiveness of an HTML document to a brochure as a means of disseminating tourism information to people with disabilities. Thesis (m.a.), California State University, Chico, Chico, CA, USA, 1997. x + 77 pp.

Huegel:1998:TDR

[Hue98] Joel C. Huegel. Telepresence for distributed rapid prototyping: a VRML approach. Thesis (m.s.m.e.), University of Washington, Seattle, WA, USA, 1998. iv + 38 pp.

Hughes:1996:PPJ

[Hug96] J. Hughes. Programming with PARTS for Java. Web Techniques, 1(8):71–73, November 1996. CODEN WETEFA. ISSN 1086-556X.

Huitfeldt:1996:WSP

[Hui96]  C. Huitfeldt.  Why SGML

is prescriptive and interpretive. In Anonymous [Ano96f], pages 108–?? ISBN 82-7283-083-3. ISSN 0804-8800. LCCN ????

Humm:1995:CBA

[Hum95] Kathryn M. Humm. Chesley Bonestell Art Gallery. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Arts/humm/bonestell.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Hunter:1994:CDS

[Hun94] B. Hunter. Conversion of documents to and from SGML. In Anonymous [Ano94b], pages 5/1–5/4. ISBN ???? LCCN ????

Husain:1996:EJJ

[Hus96] K. Husain. Extending JavaScript with Java. Web Techniques, 1 (7):58–??, 60–62, October 1996. CODEN WETEFA. ISSN 1086-556X.

Hutchison:1996:NWH

[Hut96] B. Hutchison. The NHS-wide web [health service information network]. BJHC and IM-British Journal of Healthcare Computing and Information Management, 13(2):50–51, March 1996. ISSN 0265-5217.

Homer:1997:IH

[HUW97] Alex Homer, Chris Ullman, and Steve Wright. Instant HTML. Wrox Press, Chicago, IL, USA, second edition, 1997. ISBN 1-86100-156-8. xiii + 433 pp. LCCN ????

Homer:1998:HHS

[HUW98a] Alex Homer, Chris Ullman, and Steve Wright. HTML 4 huo su shang shou = Instant HTML programmer’s reference. Chi feng tzu hsun ku fen yu hsien kung ssu, Tai-pei shih, chu pan. edition, 1998. ISBN 957-566-332-2. xi + 476 pp. LCCN ????

Homer:1998:IH

[HUW98b] Alex Homer, Chris Ullman, and Steve Wright. Instant HTML. Wrox Press, Chicago, IL, USA, 2nd , updated 1998. edition, 1998. xiii + 433 pp.

Hankins:1992:BRS

[HV92] Christopher Hankins and John Vince. Book reviews: The SGML Handbook, by Charles Goldfarb, 1990; Fractal Creations, by Timothy Wegner and Mark Peterson, 1991. The Computer Journal, 35(1):40, February 1992. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//www3.oup.co.uk/computer_
journal/hdb/Volume_35/Issue_
01/tiff/40.tif.

Heath:1988:DPS

[HW88] Jim Heath and Larry Welsch. Difficulties in parsing SGML. In Solem [Sol88], pages 71–77. ISBN 0-89791-291-8. LCCN QA 76.9 T48 A25 1988.

Handley:1990:AAD

[HW90] J. Handley and S. Weibel. ADAPT: automated document analysis processing and tagging. In Furuta [Fur90], pages 183–192. ISBN 0-521-40246-8. LCCN Z286.E43I58 1990.

Helal:1994:RSN

[HW94] Ahmed H. Helal and Joachim W. Weiss, editors. Resource sharing: new technologies as a must for universal availability of information: 16th International Essen Symposium, 18 October– 21 October 1993: Festschrift in honor of Hans-Peter Geh, volume 17. Universitätsbibliothek Essen, Essen, Germany, 1994. ISBN 3-922602-18-5. ISSN 0931-7503. LCCN Z672.5 .I58 1994.

Hartman:1996:VHB

[HW96a] Jed Hartman and Josie Wernecke. The VMRL 2.0 handbook: building moving worlds on the Web. Addison-Wesley Longman, Reading, MA, USA, January 1996. ISBN 0-201-47944-3. 204 pp. LCCN T385.H3475 1996. US$29.95.

Helal:1996:EDI

[HW96b] Ahmed H. Helal and Joachim W.

Weiss, editors. Electronic Documents and Information: From Preservation to Access. 18th International Essen Symposium, 23 October–26 October 1995: Festschrift in Honour of Patricia Battin, volume 20 of Veroffentlichungen der Universitätsbibliothek Essen. Universitätsbibliothek Essen, Essen, Germany, 1996. ISBN 3-922602-21-5. LCCN Z678.9.A1 I52 1996.

Hartman:1997:VHB

[HW97] Jed Hartman and Josie Wernecke. The VRML 2.0 handbook: building moving worlds on the Web. Addison-Wesley, Reading, MA, USA, third edition, 1997. ISBN 0-201-47944-3. xxv + 412 pp. LCCN T385.H3475 1996.

Hara:1996:SMJ

[HY96] S. Hara and H. Yasunaga.

SGML markup of Japanese classical texts: A case study. In Anonymous [Ano96f], pages 131–134. ISBN 82-7283-083-3. ISSN 0804-8800. LCCN ????

Hsu:1997:TBI

[HY97] J. Yung-Jen Hsu and W.-T. Yih. Template-based information mining from HTML documents. In Anonymous [Ano97z], pages 256–262. ISBN 0-262-51095-2. LCCN Q334 .N37 1997.

Hyman:1997:DHD

[Hym97] Michael I. Hyman. Dynamic HTML for dummies. –For dummies. I D G Books Worldwide, Indianapolis, IN, USA, 1997. ISBN 0-7645-0283-2. xviii + 396 pp. LCCN QA76.76.H94H94 1997¡.

Ibrahim:1995:SEE

[Ibr95] R. L. Ibrahim, editor. Software Engineering Education. 8th SEI CSEE Conference Proceedings. Springer-Verlag Inc., New York, NY, USA, 1995. ISBN 3-540-58951-1. LCCN QA76.758.S45 1995.

Isaacs:1998:DHV

[ID98] Scott Isaacs and Jaroslav Dudr. Dynamicke HTML: vytvareni interaktivnich webovych stranek s dynamickym obsahem, dynamickymi styly, dynamickym polohovanim. Internet. Profi. Computer Press, Praha, vyd. 1. edition, 1998. ISBN 80-7226-083-9 (vaz.). 436 pp. LCCN ????

Indrusiak:1999:ICL

[ID99] L. S. Indrusiak and R. A. Da Luz Reis. 3D integrated circuit layout visualization using VRML. In Bruzzone et al. [BUP99], pages 177–181. ISBN 1-56555-156-7. LCCN QA76.9.C65 I574 1999.

Ikeda:1998:SAT

[IDK98] Kazuyuki Ikeda, Matsuyuki Doi, and Tomiei Kazama, editors. State-of-the-art technology in anesthesia and intensive care: proceedings of the 18th International Symposium on Computing in Anesthesia and Intensive Care held in Hamamatsu city, Japan on 18–21 March 1998, volume 1168 of International congress series. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1998. ISBN 0-444-50008-1. LCCN WO 200 I61-2 1998; RD80.95 .I68 1998.

IEEE:1992:DPC

[IEE92] IEEE, editor. Digest of Papers. COMPCON Spring 1992. Thirty-Seventh IEEE Computer Society International Conference. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2655-0 (paperback), 0-8186-2656-9 (microfiche). LCCN QA75.5.C58 1992. IEEE catalog no. 92CH3098-1.

IEEE:1993:PIN

[IEE93a]  IEEE, editor.  Proceedings

of the IEEE 1993 National Aerospace and Electronics Conference. NAECON 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993.

ISBN 0-7803-1295-3. LCCN ???? Two volumes. IEEE catalog no. 93CH3306-8.

IEEE:1993:PSI

[IEE93b] IEEE, editor. Proceedings of the Second International Conference on Document Analysis and Recognition. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4960-7. LCCN TA1640.I57 1992. IEEE catalog no. 93TH0578-5.

IEEE:1994:CPA

[IEE94a] IEEE, editor. Conference Proceedings. AUTOTESTCON ’94. IEEE Systems Readiness Technology Conference. ‘Cost Effective Support Into the Next Century’. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1910-9. LCCN TK 7895 A8 A98 1994. IEEE catalog no. 94CH3436-3.

IEEE:1994:IIG

[IEE94b] IEEE, editor. IGARSS ’94. International Geoscience and Remote Sensing Symposium. Surface and Atmospheric Remote Sensing: Technologies, Data Analysis and Interpretation. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1497-2. LCCN QE33.2.R4I57 1994. Four volumes. IEEE catalog no. 94CH3378-7.

IEEE:1995:ICR

[IEE95a]  IEEE, editor.  IEEE Con-

ference Record — Abstracts. 1995 IEEE International Conference on Plasma Science: 5– 8 June 1995, Madison, Wisconsin, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2669-5. LCCN TA 2020 I612c 1995. IEEE Catalog No. 95CH35796.

IEEE:1995:OOC

[IEE95b]  IEEE, editor.  Oceans ’95:

“Challenges of our changing global environment”: conference proceedings, October 9– 12, 1995, San Diego, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-933957-14-9. LCCN TC 1505 O33197 1995. Three volumes. IEEE Catalog No. 95CH35870.

IEEE:1995:PBI

[IEE95c]  IEEE, editor.  Proceedings.

Basque International Workshop on Information Technology. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7110-6. LCCN QA76.9.D3 B37 1995.

IEEE:1995:PFW

[IEE95d]  IEEE, editor.  Proceedings

Fifth Workshop on Hot Topics in Operating Systems (HotOS-V). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7081-9. LCCN QA76.76.O63.W6666 1995. IEEE catalog no. 95TH8059.

IEEE:1995:PFI

[IEE95e] IEEE, editor. Proceedings of the Fourteenth IEEE Symposium on Mass Storage Systems. Storage - At the Forefront of Information Infrastructures. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7064-9. LCCN TK7895.M4 I5 1995. IEEE catalog no. 95CB35860.

IEEE:1995:PTI

[IEE95f] IEEE, editor. Proceedings of the third International Conference on Document Analysis and Recognition: August 14–16, 1995, Montreal, Canada, volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7128-9. LCCN TA1650 .I58 1995. Two volumes.

IEEE:1995:PWM

[IEE95g]  IEEE, editor.  Proceedings.

Workshop on Mobile Computing Systems and Applications. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-6345-6. LCCN QA76.59.W67 1994. IEEE catalog no. 94TH06734.

IEEE:1996:DTP

[IEE96a]  IEEE, editor.  1996 digest

of technical papers: International Conference on Consumer Electronics: June 5–7, 1996, Rosemont, IL, volume 42(3) of IEEE Transactions on Consumer Electronics. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, August 1996. CODEN ITCEDA. ISBN 0-7803-3029-3 (softbound), 0-7803-3030-7 (casebound), 0-7803-3031-5 (microfiche). ISSN 0098-3063. LCCN TK7869 .I16 1996.

IEEE:1996:ICP

[IEE96b] IEEE, editor. 1996 International Conference on Parallel and Distributed Systems: June 3–6, 1996, Tokyo, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7267-6. LCCN QA76.58 .I5443 1996. IEEE Catalog No. 96TB100045.

IEEE:1996:SAD

[IEE96c] IEEE, editor. 1996 sixth Annual Dual-Use Technologies and Applications Conference: mapping & planning the information highway, June 3–6, 1996, OnCenter, Syracuse, NY. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN ???? LCCN ????

IEEE:1996:CPT

[IEE96d] IEEE, editor. COMPSAC 96: proceedings of the Twentieth annual International Computer Software & Applications Conference held August 21–31, 1996 at Seoul, Korea. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7579-9. LCCN QA 76.6 C6295 1996. IEEE catalog number 96CB35986.

IEEE:1996:DPC

[IEE96e] IEEE, editor. Digest of Papers. COMPCON ‘96. Technologies for the Information Superhighway. Forty-First IEEE Computer Society International Conference: February 25–28, 1996 Santa Clara, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7414-8. LCCN QA76 .I49 1996. IEEE Catalog No. 96CB35911.

IEEE:1996:FIC

[IEE96f] IEEE, editor. Fourth International Conference on Image Management and Communication (IMAC 95): medical imaging service in a network environment, August 20–24, 1995, Oahu Island, Hawaii. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7560-8. LCCN R857.O6 I57 1995. IEEE Computer Society Press order number PR07560.

IEEE:1996:ICS

[IEE96g] IEEE, editor. IEEE CEFC ’96. The Seventh Biennial IEEE Conference on Electromagnetic Field Computation. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN ???? LCCN ????

IEEE:1996:PIW

[IEE96h] IEEE, editor. Proceedings International Workshop on Multimedia Software Development: March 25–26, 1996, Berlin, Germany. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, March 1996. ISBN 0-8186-7511-X. LCCN QA76.575 .M84 1996.

IEEE:1996:PICb

[IEE96i] IEEE, editor. Proceedings of the 16th International Conference on Distributed Computing Systems. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7399-0. LCCN ???? IEEE Catalog No. 96CB35954.

IEEE:1996:PWE

[IEE96j] IEEE, editor. Proceedings of the 5th Workshops on Enabling Technologies: Infrastructure for Collaborative Enterprises (WET ICE ’96): June 19–21, 1996, Stanford, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7446-6. LCCN TS155.6 .W56 1996.

IEEE:1996:PFI

[IEE96k] IEEE, editor. Proceedings of the Fifth IEEE International Conference on Fuzzy Systems: FUZZ-IEEE ’96, September 8– 11, 1996, Hyatt Regency Hotel, New Orleans, Louisiana, volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3646-1 (casebound), 0-7803-3645-3 (softbound), 0-7803-3647-X (microfiche). LCCN TJ212.2.I3249 1996. IEEE catalog number 96CH35998.

IEEE:1996:PICa

[IEE96l]  IEEE, editor.  Proceedings

of the International Conference on Multimedia Computing and Systems: June 17–23, 1996, Hiroshima, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7436-9. LCCN QA76.575.I623 1996. IEEE Catalog No. 96TB100057.

IEEE:1996:PTF

[IEE96m] IEEE, editor. Proceedings of the Third Forum on Research and Technology Advances in Digital Libraries - ADL ’96. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7402-4. LCCN ???? IEEE Catalog No. 96TB100053.

IEEE:1996:SGL

[IEE96n] IEEE, editor. The Sixth Great Lakes Symposium on VLSI, Iowa State University, Ames, Iowa, March 22–23, 1996: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7502-0. LCCN TK 7874.75 G74 1996. IEEE Catalog No. 96TB100041.

IEEE:1996:SIW

[IEE96o] IEEE, editor. Sixth International Workshop on Research Issues in Data Engineering: interoperability of nontraditional database systems: proceedings, February 26–27, 1996, New Orleans, Louisiana. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7289-7. LCCN QA76.9.D3 I593 1996. IEEE Catalog No. 96TB100021.

IEEE:1997:ICD

[IEE97a] IEEE, editor. 1997 13th International Conference on Digital Signal Processing proceedings / DSP 97, July 2–4, 1997, conference centre “P.M. Nomikos”, Santorini, Hellas (Greece), volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-4137-6, 0-7803-4138-4, 0-7803-4139-2. LCCN TK 5102.9 I58 1997.

IEEE:1997:IIP

[IEE97b] IEEE, editor. 1997 IEEE International Professional Communication Conference: crossroads in communication: IPCC 97 proceedings, October 22–25, 1997 in Salt Lake City, Utah at Snowbird Resort. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-4184-8 (softbound), 0-7803-4185-6 (casebound), 0-7803-4186-4 (microfiche). LCCN T 10.5 I54 1997. IEEE catalog number 97CH36123.

IEEE:1997:ICP

[IEE97c] IEEE, editor. IEEE Conference on Protocols for Multimedia Systems-Multimedia Networking: PROMS-MmNet ’97: proceedings, November 24–27, 1997, Santiago, Chile. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7916-6, 0-8186-7918-2 (microfiche). LCCN QA76.575 .I33 1997. IEEE Computer Society order number PR07916. IEEE order plan catalog number 97TB100116.

IEEE:1997:PNI

[IEE97d] IEEE, editor. Proceedings / Ninth IEEE International Conference on Tools with Artificial Intelligence, November 3– 8, 1997, Newport Beach, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-8203-5, 0-8186-8204-3 (case), 0-8186-8205-1 (microfiche). ISSN 1082-3409. LCCN Q334 .I54 1997.

IEEE:1997:PIR

[IEE97e] IEEE, editor. Proceedings of the 1997 IEEE/RSJ International Conference on Intelligent Robots and System, IROS ’97s: innovative robotics for real-world applications, September 7–11, 1997, World Trade Center Atria, Grenoble, France, volume 3. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-4119-8 (softbound), 0-7803-4120-1 (casebound), 0-7803-4121-X (microfiche). LCCN TJ210.3.I443 1997. Three volumes. IEEE catalog number 97CH36108.

IEEE:1997:PFI

[IEE97f] IEEE, editor. Proceedings of the Fourth International Conference on Document Analysis and Recognition, August 18– 20, 1997, Ulm, Germany, volume 2. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7899-2, 0-8186-7898-4, 0-8186-7900-X. LCCN TA1640 .I57 1997. Two volumes. IEEE Order catalog number 97TB100138. IEEE Computer Society order number PR07898.

IEEE:1997:PTF

[IEE97g] IEEE, editor. Proceedings, the Twenty-First Annual International Computer Software & Applications Conference (COMP-SAC’97), August 13–15, 1997, Washington, DC. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-8106-3, 0-8186-8107-1, 0-8186-8105-5. ISSN 0730-3157. LCCN QA76.6 .C62 1997. IEEE Catalog Number 97CB36112. IEEE Computer Society Order Number PR08105.

IEEE:1998:ICS

[IEE98a]  IEEE, editor.  1998 IEEE

Computer Society Conference on Computer Vision and Pattern Recognition: proceedings, June 23–25, 1998, Santa Barbara, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8497-6 (softbound), 0-7803-5063-4 (casebound), 0-8186-8499-2 (microfiche). ISSN 1063-6919. LCCN TA1632 .I36 1998. IEEE Catalog Number 98CB36231. IEEE Computer Society Order Number PR08497.

IEEE:1998:IRI

[IEE98b] IEEE, editor. 1998 IEEE/ RSJ International Conference on Intelligent Robots and Systems: proceedings: innovatiions in theory, practice, and applications: October 13–17, 1998, Victoria Conference Centre, Victoria, BC, Canada, volume 3. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-7803-4465-0 (softbound), 0-7803-4466-9 (casebound), 0-7803-4467-7 (microfiche). LCCN TJ210.3.I443 1998. Three volumes. IEEE Catalog Number 98CH36190.

IEEE:1998:CPT

[IEE98c] IEEE, editor. COMPSAC ’98:

proceedings [of the] Twenty-Second Annual International Computer Software & Applications Conference: held August 19–21, 1998, Vienna, Austria, COMPSAC -New York-1998. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8585-9 (softbound), 0-7803-5168-1 (casebound), 0-8186-8587-5 (microfiche). ISSN 0730-3157. LCCN QA76.6 .I546 1998.

IEEE:1998:IIC

[IEE98d] IEEE, editor. ICCE International Conference on Consumer Electronics. Title: International Conference on Consumer Electronics: 1998 digest of technical papers: seventeenth in the ICCE series: June 2–4, 1998. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-7803-4357-3 (softbound), 0-7803-4358-1 (casebound), 0-7803-4359-X (microfiche). LCCN TK7869 .I16 1998. IEEE Catalog Number 98CH36160.

IEEE:1998:MME

[IEE98e] IEEE, editor. MELECON ’98, 9th Mediterranean Electrotechnical Conference: proceedings, May 18–20. 1998, Tel-Aviv, Israel, volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-7803-3879-0 (softbound), 0-7803-3880-4 (casebound), 0-7803-3881-2 (microfiche). LCCN TK5 .M43 1998; TK5101.A1 M38 1998. Two volumes. IEEE catalog number 98CH36056.

IEEE:1998:TIC

[IEE98f] IEEE, editor. Twelfth International Conference on Information Networking, (ICOIN-12), Koganei, Tokyo, Japan, January 21–23, 1998: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-7225-0. LCCN TK5105.5 .I53 1998.

IEEE:1999:IIF

[IEE99a] IEEE, editor. IEEE International Forum on Research and Technology Advances in Digital Libraries: [ADL’99]: proceedings: May 19–21, 1999, Baltimore, MD. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN 0-7695-0219-9, 0-7695-0221-0 (microfiche). ISSN 1092-9959. LCCN TK5103.5.F678 1999; ZA4080 .F67 1999. IEEE Computer Society Order Number PR00219.

IEEE:1999:RII

[IEE99b] IEEE, editor. RE’99: IEEE International Symposium on Requirements Engineering: proceedings: June 7–11, 1999, University of Limerick, Limerick, Ireland. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN 0-7695-0188-5, 0-7695-0190-7 (microfiche). ISSN 1090-705X. LCCN QA76.751.I36 1999. IEEE Computer Society Order Number PR00188. IEEE Order Plan Catalog Number PR00188.

Ishizuka:1996:SFT

[IN96] H. Ishizuka and M. Negishi.

SGML as a fundamental technology for an information system — document database, WWW and CALS. Joho-Shori (J. Information Processing Soc. Japan), 37(3):207–212, March 1996. CODEN JOSHA4. ISSN 0447-8053.

Indermaur:1996:SSR

[Ind96] K. Indermaur. Spider 1.5 [software review]. DBMS, 9(8):38, 67, July 1996. CODEN DBM-SEO. ISSN 1041-5173.

Ingram:1995:WWW

[Ing95] P. Ingram. The World Wide Web. Computers and Geosciences, 21(6):799–816, July 1995. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Iijima:1996:FAD

[INIK96] K. Iijima, K. Nagahashi, H. Ishiguro, and H. Kaizuka. Functional Alloys Database — WWW-based full-text retrieval and browsing system. Joho kanri = Information & documentation, 39(8):599–608, November 1996. CODEN JOKAAB. ISSN 0021-7298.

ISO:1986:IPTb

[Int86a] International Organization for Standardization. Information Processing: Text and Office Systems: (SGML) Document Interchange Format (SDIF). International Organization for Standardization, Geneva, Switzerland, 1986. ISBN ???? 4 pp. LCCN QA76.73.S44 I56 1986. Draft International Standard ISO/DIS 9069.

ISO:1986:IPTa

[Int86b] International Organization for Standardization. Information Processing, Text and Office Systems, Standard Generalized Markup Language (SGML) = Traitement de l’information, systemes bureautiques, langage standard généralisé de balisage (SGML). First edition, 1986-10-15. International Organization for Standardization, Geneva, Switzerland, October 15, 1986. ISBN ???? xi + 155 pp. LCCN QA76.9.T48

I57 1986. URL http://www.
iso.ch/cate/d16387.html. In-

ternational Standard ISO 8879-1986. Federal information processing standard; FIPS PUB 152.

ISO:1986:IIP

[Int86c] International Organization for Standardization. ISO 8879:1986: Information processing — Text and office systems — Standard Generalized Markup Language (SGML). International Organization for Standardization, Geneva, Switzerland, August 17, 1986. 155 pp. CHF

212. URL http://www.iso.ch/
cate/d16387.html.

ISO:1986:IIPa

[Int86d] International Organization for Standardization. ISO 8879:1986: Information processing — Text and office systems — Standard Generalized Markup Language (SGML). International Organization for Standardization, Geneva, Switzerland, August 17, 1986. ISBN ???? 155 pp. LCCN ???? CHF

212. URL http://www.iso.ch/
cate/d16387.html.

ISO:1987:IPS

[Int87] International Organization for Standardization. Information processing: SGML support facilities: Techniques for using SGML. International Organization for Standardization, Geneva, Switzerland, 1987. ISBN ???? vi + 96 pp. LCCN ????

ISO:1988:IPSc

[Int88a] International Organization for Standardization. Information processing, SGML support facilities: registration procedures for public text owner identifiers. International standard ISO 9070. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? iii + 5 pp. LCCN

???? URL http://www.iso.
ch/cate/d16645.html.

ISO:1988:IPSe

[Int88b] International Organization for Standardization, Geneva, Switzerland. Information processing, SGML support facilities: registration procedures for public text owner identifiers, 1988. iii + 5 pp.

ISO:1988:IPSa

[Int88c] International Organization for Standardization, Geneva, Switzerland. Information Processing, SGML Support Facilities: SGML Document Interchange Format (SDIF) = Traitement de l’information, bureaucratique: format d’échange de document SGML (SDIF), 1988. iii + 6 pp. International Standard ISO/DIS 9069.

ISO:1988:IPSb

[Int88d] International Organization for Standardization. Information Processing, SGML Support Facilities: SGML Document Interchange Format (SDIF) = Traitement de l’information, bureaucratique: format d’échange de document SGML (SDIF). International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? iii + 6 pp. LCCN QA76.73.S44 I55 1988. International Standard ISO/DIS 9069.

ISO:1988:IPT

[Int88e] International Organization for Standardization, Geneva, Switzerland. Information processing: text and office systems: Standard Generalized Markup Language (SGML): amendment 1 = Traitement de l’information: systèmes bureautiques: langage standard generalisé de balisage (SGML): amendement 1, 1988. 15–?? pp.

ISO:1988:IIP

[Int88f] International Organization for Standardization. ISO 9069:1988: Information processing — SGML support facilities — SGML Document Interchange Format (SDIF). International Organization for Standardization, Geneva, Switzerland, 1988. 8 pp. CHF 50. URL http://www.

iso.ch/cate/d16643.html.

ISO:1988:IIPf

[Int88g] International Organization for Standardization. ISO 9069:1988: Information processing — SGML support facilities — SGML Document Interchange Format (SDIF). International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 8 pp. LCCN ????

CHF 50. URL http://www.
iso.ch/cate/d16643.html.

ISO:1988:IIT

[Int88h] International Organization for Standardization. ISO/IEC TR 9573:1988: Information processing — SGML support facilities — Techniques for using SGML. International Organization for Standardization, Geneva, Switzerland, 1988. ISBN ???? 124 pp. LCCN ????

                                        CHF 200. URL http://www.
                                        iso.ch/cate/d17319.html.

ISO:1988:IPSd

[Int88i] International Standards Organization. Information processing, SGML support facilities: Traitement de l’information, facilités de support pour SGML: techniques for using SGML = techniques d’utilisation du SGML. Technical report ISO/IEC TR 9573, International Organization for Standardization, Geneva, Switzerland, 1988. vii + 124 pp. URL http:

                                        //www.iso.ch/cate/d17319.
                                        html;   http://www.iso.ch/
                                        cate/d17330.html; http://
                                        www.iso.ch/cate/d17332.html.

ISO:1991:III

[Int91a] International Organization for Standardization. ISO/IEC 9070:1991 Information technology — SGML support facilities — Registration procedures for public text owner identifiers. International Organization for Standardization, Geneva, Switzerland, 1991. 12 pp. CHF

                                        62. URL http://www.iso.ch/

cate/d16645.html. Available

in English only.

ISO:1991:IIIf

[Int91b] International Organization for Standardization. ISO/IEC 9070:1991 Information technology — SGML support facilities — Registration procedures for public text owner identifiers. International Organization for Standardization, Geneva, Switzerland, 1991. ISBN ???? 12 pp. LCCN ???? CHF

62. URL http://www.iso.ch/
cate/d16645.html. Available

in English only.

ISO:1991:IIT

[Int91c] International Organization for Standardization. ISO/IEC TR 9573-13:1991: Information technology — SGML support facilities — Techniques for using SGML — Part 13: Public entity sets for mathematics and science. International Organization for Standardization, Geneva, Switzerland, 1991. ISBN ???? 82 pp. LCCN ????

CHF 176. URL http://www.
iso.ch/cate/d17332.html.

ISO:1991:ITS

[Int91d] International Organization for Standardization and International Electrotechnical Commission. Information technology– SGML support facilities–techniques for using SGML. technologies de l’information: facilités de support pour SGML: techniques d’utilisation du SGML:

Partie 13, ensembles d’entités publiques pour les domaines mathématique et scientifique. Technical report ISO/IEC TR 9573-13, International Organization for Standardization, Geneva, Switzerland, 1991. iii + 82 pp. URL http:

//www.iso.ch/cate/d17319.
html;   http://www.iso.ch/
cate/d17330.html; http://
www.iso.ch/cate/d17332.html.

ISO:1992:IIT

[Int92a] International Organization for Standardization. ISO/IEC TR 9573-11:1992 Information processing — SGML support facilities — Techniques for using SGML — Part 11: Application at ISO Central Secretariat for International Standards and Technical reports. International Organization for Standardization, Geneva, Switzerland, 1992. 73 pp. CHF

164. URL http://www.iso.ch/
cate/d17330.html.

ISO:1992:IITb

[Int92b] International Organization for Standardization. ISO/IEC TR 9573-11:1992 Information processing — SGML support facilities — Techniques for using SGML — Part 11: Application at ISO Central Secretariat for International Standards and Technical reports. International Organization for Standardization, Geneva, Switzerland, 1992. ISBN ???? 73 pp. LCCN ????

CHF 164. URL http://www.
iso.ch/cate/d17330.html.

IW:1994:IW

[Int94] Inter@ctive week, 1994. ISSN 1080-9406. URL http://www.

interactive-week.ziff.com.

InterActive Enterprises, Garden City, NY.

ISO:1997:IIIb

[Int97] International Organization for Standardization. ISO/IEC 10744:1997: Information technology — Hypermedia/Timebased Structuring Language (HyTime). International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 468 pp. LCCN ????

CHF 320. URL http://www.
iso.ch/cate/d29303.html.

ISO:19xx:IIDi

[Intxx] International Organization for Standardization. ISO/IEC DIS 15445: Information technology — Document description and processing languages — Hyper Text Markup Language (HTML). International Organization for Standardization, Geneva, Switzerland, 19xx. ISBN ???? ???? pp. LCCN

???? URL http://www.iso.
ch/cate/d27688.html.

Ifeachor:1994:PIC

[IR94] E. C. Ifeachor and K. G. Rosen, editors. Proceedings of the International Conference on Neural Networks and Expert Systems in Medicine and Healthcare. Univ. Plymouth, Plymouth, UK, 1994. ISBN ???? LCCN ????

Ip:1998:MIA

[IS98] Horace H. S. Ip and Arnoldus Wilhelmus Maria Smeulders, editors. Multimedia information analysis and retrieval: IAPR International Workshop MI-NAR’98, Hong Kong, China, August 1998, proceedings, volume 1464 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. ISBN 3-540-64826-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .I318 1998.

Isaacs:1997:IDH

[Isa97] Scott Isaacs. Inside Dynamic HTML. Microsoft programming series. Microsoft Press, Bellevue, WA, USA, 1997. ISBN 1-57231-686-1. xvi + 473 pp. LCCN QA76.76.H94 I83 1997. US$39.99.

Isaacs:1998:ADH

[Isa98a] Scott Isaacs. Atelier Dynamic HTML. Microsoft Press, Bellevue, WA, USA, 1998. ISBN 2-84082-284-9. xi + 504 pp. LCCN ????

Isaacs:1998:FDH

[Isa98b] Scott Isaacs. A fondo dynamic HTML. Serie de programacion Microsoft. McGraw-Hill, Madrid, Spain, 1a en español edition, 1998. ISBN 84-481-2024-8. xix + 429 pp. LCCN ????

Iseri:1990:PVD

[Ise90] M. Iseri. A publisher’s view of DTP. Journal of the Information Processing Society of Japan = Joho Shori, 31(11): 1501–1507, ???? 1990. CODEN JOSHA4. ISSN 0447-8053.

Ishizuka:1994:SFD

[Ish94a] H. Ishizuka. SGML and full-text databases. 2. the latest information related technology. Joho kanri = Information & documentation, 37(2):148–159, May 1994. CODEN JOKAAB. ISSN 0021-7298.

Ishizuka:1994:LIR

[Ish94b] Hidehiro Ishizuka. The latest information related technology [2]: SGML and full-text database. Joho kanri = Information & documentation, 37(2): 148–??, May 1, 1994. CODEN JOKAAB. ISSN 0021-7298.

Ishizuka:1995:EMS

[Ish95] H. Ishizuka. The effect of multimedia on social life. 3. Multimedia and publishing. Joho kanri = Information & documentation, 38(4):353–368, July 1995. CODEN JOKAAB. ISSN 0021-7298.

Iskander:1996:TBR

[Isk96] Magdy F. Iskander, editor.

Technology-Based Re-Engineering. Engineering Education. Proceeding of Frontiers in Education FIE’96 26th Annual Conference, November 6–9, 1996, Salt Lake City, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3348-9 (softbound), 0-7803-3349-7 (casebound), 0-7803-3350-0 (microfiche), 0-7803-3720-4 (CD-ROM). LCCN T 62 F76 1996. Three volumes. IEEE catalog number 96CH35946.

Ide:1995:THG

[ISM95] N. M. Ide and C. M. Sperberg-McQueen. The TEI: history, goals, and future. Computers and the humanities, 29(1):5–15, ???? 1995. CODEN COHUAD. ISSN 0010-4817.

Ito:1998:DHD

[Ito98]  S. Ito.  Developing HTML

documents and help system. In Anonymous [Ano98u], pages 307–308. ISBN 1-891709-00-3, 1-891709-01-1. LCCN ????

Ide:1995:ED

[IV95] N. Ide and J. Veronis. Encoding dictionaries. Computers and the humanities, 29(2):167–179, ???? 1995. CODEN COHUAD. ISSN 0010-4817.

Ivler:1995:DHA

[Ivl95] J. M. Ivler. Developing inhouse applications for Northrop Grumman on the F/A-18 E/F program using Mosaic and HTTPD. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
CorInfSys/ivler/edstlk1.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Ibanez:1997:HWM

[IZ97] Ardith Ibanez and Natalie Zee.

HTML Web Magic. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, January 1997. ISBN 1-56830-335-1. x + 230 pp. LCCN QA76.76.H94I23 1997. US$39.99.

Ibanez:1998:HAM

[IZ98] Ardith Ibanez and Natalie Zee. HTML artistry: more than code. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1998. ISBN 1-56830-454-4. 286 pp. LCCN QA76.76.H94I227

1998.  URL http://www.
htmlartistry.com.

Jackson:1995:AT

[Jac95]  Robert E. Jackson.  As-

troWeb tools. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Astronomy/jackson/jackson.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Jacobs:1996:JBW

[Jac96]  Russell Jacobs.  Javascript:

Blueprints to a world beyond HTML. Prima Tech Publishing, Rocklin, CA, USA, July 1996. ISBN 0-7615-0685-3. xii + 337 pp. LCCN QA76.73.J39K66 1996. US$35.00.

Jacucci:1998:GMD

[Jac98] G. Jacucci, editor. Globalization of manufacturing in the digital communications era of the 21st century: innovation, agility, and the virtual enterprise: proceedings of the tenth international IFIP WG5.2/5.3 International Conference PROLAMAT 98, Trento, Italy, September 9-11 & 12, 1998. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 0-412-83540-1. LCCN TS155.6 .I569 1998. Includes CD-ROM.

Jadhav:1997:VVD

[Jad97] Nitin Jadhav. VToD: VRML to database link. Thesis (m.s.), Wayne State University, Detroit, MI, USA, 1997. v + 49 pp.

Jahn:1999:OVA

[Jah99]  G. V. Jahn.  Overview of

VRML applications to architecture. In Melker [Mel99], pages 214–219. ISBN 0-8194-3161-3. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.3687.

Jalil:1998:HML

[Jal98] Adnan Jalil. Hypertext markup language (HTML). ????, ????, 1998. vi + 64 pp.

James:1997:AAH

[Jam97] Frankie James. AHA: Audio HTML access. Computer Networks and ISDN Systems, 29(8– 13):1395–1404, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1997&
volume=29&issue=08-13&aid=
1720.

James:1998:LDA

[Jam98] F. James. Lessons from developing audio HTML interface. In ACM [ACM98], pages 27–34. ISBN 1-58113-020-1. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/assets/274497/.

ACM order number 444980.

Jander:1996:WRW

[Jan96] M. Jander. Welcome to the revolution [WWW-based network management]. Data Communications International, 25(16): 38–42, 44, 46, 48, 50, 52– 53, ???? 21, 1996. CODEN DCINEU. ISSN 0363-6399.

Jordan:1995:TIP

[JCDA95] James M. Jordan, Mark Cresitello-Dittmar, and Jesse Allen. Tailoring information presentation with words and pictures at the COSSC. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/

Astronomy/jordan/jordan.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Jost:1996:MIP

[JD96] R. Gilbert Jost and Samuel J. Dwyer III, editors. Medical Imaging 1996. PACS Design and Evaluation: Engineering and Clinical Issues, volume 2711 of Proceedings of the SPIE — The International Society for Optical Engineering. SPIE Optical Engineering Press, Bellingham, WA, USA, 1996. CO-DEN PSISDG. ISBN 0-8194-2086-7. ISSN 0277-786X (print), 1996-756X (electronic). LCCN R857.P52 M427 1996.

Jennings:1995:HPL

[JDGP95] Donald Jennings, Peter Damon, Maia Good, and Ryszard Pisarski. How to present lots of volatile information on the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Astronomy/jennings/paper.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Jelliffe:1998:XSC

[Jel98] Rick Jelliffe. The XML and SGML Cookbook: Recipes for Structured Information. Charles F. Goldfarb series on open information management. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-614223-0. 656 pp. LCCN QA76.73.X56J45 1998. US$55.00.

Jacobs:1996:FHF

[JGKR96] Stephan Jacobs, Michael Gebhardt, Stefanie Kethers, and Wojtek Rzasa. Filling HTML forms simultaneously: CoWebarchitecture and functionality. Computer Networks and ISDN Systems, 28(7–11):1385– 1395, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=
1607.

Jacobson:1995:EDL

[JI95] F. F. Jacobson and E. Ignacio. Evaluating digital libraries in the context of learning and teaching. In ACM [ACM95a], pages 21–22. CODEN SIGBEL. ISBN ???? ISSN 0894-0819. LCCN ????

Joshi:1995:SPI

[JJ95a] Rajeev Joshi and Hans Joseph.

Selective presentation of information on the Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/

                                        EvolvingInter/joshi/Joseph-

Joshi.html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Judd:1995:EAS

[JJ95b] Peggy Judd and F. Scott Johnson. Electronic AIP: SGML enables full-text scientific publishing on the Web. Computers in Physics, 9(4):369–??, 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Judd:1995:SEF

[JJ95c] Peggy Judd and Scott Johnson. SGML enables full-text scientific publishing on the Web. Computers in Physics, 9(4):369– 370, July–August 1, 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Jackson:1997:FH

[JK97] Ken Jackson and Sonya E. Keene. FrameMaker to HTML. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-31204-2. xii + 327 pp. LCCN QA76.76.H94J33 1997. US$36.53. Includes CD-ROM.

Jackson:1998:CFH

[JK98] K. Jackson and S. Keene. Converting FrameMaker to HTML. In Anonymous [Ano98u], pages 522–?? ISBN 1-891709-00-3, 1-891709-01-1. LCCN ????

Jagadish:1996:PAS

[JM96] H. V. Jagadish and Inderpal Singh Mumick, editors. Proceedings of the 1996 ACM SIG-MOD International Conference on Management of Data, Montreal, Quebec, Canada, June 4– 6, 1996, volume 25(2) of SIG-MOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY, USA, June 1996. CODEN SRECD8. ISBN 0-89791-794-4. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.25 no.2 1996.

Jacobsen:1995:PAD

[JMB95] Lynn Jacobsen, David Millman, and Walter Bourne. Providing access to a data library: SQL and full-text IR methods of automatically generating Web structure. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
SocialSci/jacobsen/jacobsen.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Junion-Metz:1998:CPW

[JMS98] Gail Junion-Metz and Brad Stephens. Creating a power website: HTML, tables, imagemaps, frames, and forms: a how-to-do-it manual. How-to-do-it manuals for librarians; no. 88 Howto-do-it manuals for libraries; no. 88. Neal-Schuman Publishers, New York, NY, USA, 1998. ISBN 1-55570-323-2. xviii + 199 pp. LCCN Z674.75.W67J86 1998.

Johnson:1998:WBM

[JOA98] C. W. Johnson, G. Oser, and A. J. Abedor. Web browser as medical educator/researcher using HTML and JavaScript. In Chute [Chu98], pages 1023–?? ISBN ???? ISSN 1091-8280. LCCN ????

Johansson:1995:EST

[Joh95a] S. Johansson. The encoding of spoken texts. Computers and the humanities, 29(2):149–158, ???? 1995. CODEN COHUAD. ISSN 0010-4817.

Johnson:1995:PSE

[Joh95b] Rick Johnson. Practical SGML, 2nd ed., Eric van Herwijnen. Technical Communication, 42 (1):??, February 1995. CODEN TLCMBT. ISSN 0049-3155.

Jones:1995:TLA

[Jon95]  Kennie Jones.  TOPS on-

line — automating the construction and maintenance of HTML pages. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Autools/jones/kjones.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Joyce:1998:PTN

[Joy98] Daniel T. Joyce, editor. Proceedings of the Twenty-ninth SIGCSE Technical Symposium on Computer Science Education: February 25–March 1, 1998, Atlanta, Georgia, volume 30(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY, USA, 1998. ISBN 0-89791-994-7. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA1 .A86 v.30, no.1.

Johnson:1995:ESS

[JRK95] D. R. Johnson, M. Ruzek, and M. Kalb. Earth System Science Education: an Internetbased electronic curriculum. In Stein [Ste95a], pages 572–574 (vol. 1). ISBN 0-7803-2568-0, 0-7803-2567-2 (paperback). LCCN QE 33.2 R4 I58 1995. Three volumes. IEEE catalog no. 95CH35770.

Jones:1996:WLI

[JRS96] R. P. Jones, F. Ruehr, and R. Salter. Web-based laboratories in the introductory curriculum enhance formal methods. In Klee [Kle96], pages 160– 164. CODEN SIGSD3. ISBN 0-89791-757-X. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA1 .A86. ACM order number: 457960.

Jones:1996:EUC

[JRV96] H. (Huw) Jones, Ralph Raby, and Dion Vicars, editors. Eurographics UK: conference proceedings, 1996: Imperial College, London, 26–28 March 1996, volume 1. Eurographics UK Chapter, Abingdon, UK, 1996. ISBN 0-9521097-3-5 (paperback),0-9521097-4-3 (vol. 1), 0-9521097-5-1 (vol. 2). LCCN ???? Two volumes.

Jamsa:1997:VPL

[JSY97] Kris A. Jamsa, Phil Schmauder, and Nelson Yee. VRML programmer’s library. Jamsa Press, Las Vegas, NV, USA, 1997. ISBN 1-884133-51-7 (paperback). viii + 548 pp. LCCN QA76.76.H94J36 1997.

Jamsa:1998:VBD

[JSY98] Kris A. Jamsa, Phil Schmauder, and Nelson Yee. VRML: biblioteca del programador. Mc-Graw-Hill, New York, NY, USA, 1a en espanol. edition, 1998. ISBN 970-10-1924-5. x + 548 pp. LCCN ????

Jones:1997:CSS

[JT97] Joseph R. Jones and Paul B. Thurrott. Cascading Style Sheets: a primer. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1997. ISBN 1-55828-579-2. xxi + 308 pp. LCCN TK5105.888 .J653 1998.

Jones:1998:CSS

[JT98] Joseph R. Jones and Paul Thurrott. Cascading Style Sheets: a primer. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1998.

ISBN 1-55828-579-2. xxi + 308 pp. LCCN TK5105.888.J653 1998. US$24.95.

Jewett:1996:UHW

[JW96] B. F. Jewett and R. B. Wilhelmson. Use of HTML and Web tools in atmospheric sciences research. In Anonymous [Ano96a], pages 423–426. ISBN ???? LCCN QC 851 I57 1996.

Kitsuregawa:1998:ICD

[K^{+}98] Masaru Kitsuregawa et al., editors. 15th International Conference on Data Engineering: proceedings: March 23–26, 1999, Sydney, Australia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-7695-0071-4 (softbound), 0-7803-5581-4 (casebound), 0-7695-0073-0 (microfiche). ISSN 1063-6382. LCCN QA76.9.D3 I5582 1999. IEEE Catalog Number 99CB36337.

Kaelbling:1988:IS

[Kae88]  Michael J. Kaelbling.  On

improving SGML. Technical research report OSU-CISRC-7/88-TR22, Computer and Information Science Research Center, Ohio State University, Columbus, OH, USA, 1988. 5 pp.

Kaelbling:1990:IS

[Kae90] M. J. Kaelbling. On improving SGML. Electronic Publishing—Origination, Dissemination, and Design, 3(2):93–98, May 1990. CODEN EPODEU. ISSN 0894-3982.

Kim:1991:MST

[KAH91] Hyung-Jun Kim, Byung-Jun Ahn, and Jin-Ho Ham. Multimedia standardization trends. Korea Information Science Society Review, 9(3):33–40, ???? 1991. CODEN CHKWEN. ISSN 1015-9908.

Kahn:1995:VCL

[Kah95] P. Kahn. Visual cues for local and global coherence in the WWW. Communications of the ACM, 38(8):67–69, August 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kahlisch:1996:IAH

[Kah96] T. Kahlisch. Improving access to hypertext based study material for the blind. In Klaus et al. [KAKZ96], pages 229–236. ISBN 3-85403-087-8, 3-486-23797-7 (München), 3-7029-0413-1 (Wien). LCCN HV1569.5.T59 1996. Two volumes.

Klaus:1996:IAC

[KAKZ96] Joachim Klaus, E. Auff, W. Kremser, and W. L. Zagler, editors. Interdisciplinary Aspects on Computers Helping People with Special Needs. 5th International Conference, ICCHP ’96, Linz, Austria, July 1996, volume 792. Osterreichische Computer Gesellschaft, Wien, Austria, 1996. ISBN 3-85403-087-8, 3-486-23797-7 (München), 3-7029-0413-1 (Wien). LCCN HV1569.5.T59 1996. Two volumes.

Kamibayashi:1990:SPD

[Kam90a] N. Kamibayashi. State-of-the-art and perspectives on document editors. Journal of the Information Processing Society of Japan = Joho Shori, 31(11): 1535–1542, ???? 1990. CODEN JOSHA4. ISSN 0447-8053.

Kamper:1990:TEM

[Kam90b] J. Kamper. Towards extensible multi-media document architectures. In Anonymous [Ano90g], page 6. ISBN ???? LCCN ????

Kanazy:1993:SDP

[Kan93a] Y. Kanazy. SGML document processing. Fujitsu, 44(6):521– 526, ???? 1993. CODEN FUJ-TAR. ISSN 0016-2515.

Kaneko:1993:SHS

[Kan93b] A. Kaneko. Survey of hypermedia studies. Journal of the Information Processing Society of Japan = Joho Shori, 34(1): 60–71, January 1993. CODEN JOSHA4. ISSN 0447-8053.

Kaplan:1997:IMW

[Kap97] H. Kaplan. Interactive multimedia and the World Wide Web. Educom review, 32(1): 48–51, January–February 1997. CODEN EDREEW. ISSN 1045-9146.

Karney:1993:SQR

[Kar93] James Karney. SGML: The quiet revolution. PC Magazine, 12(??):246, ???? 1993. CODEN PCMGEP. ISSN 0888-8507.

Karpinski:1996:BHC

[Kar96] Richard Karpinski. Beyond HTML: Covers Acrobat, Java, Shockwave, VRML, and More. Osborne/McGraw-Hill, Berkeley, CA, USA, 1996. ISBN 0-07-882198-3. xxiii + 472 pp. LCCN QA76.76.H94K37 1996. US$27.95.

Kasper:1995:UWR

[Kas95a] Theresa Kasper. Untangling the Web — the role of text retrieval in a hypertext environment. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Autools/kasper/kasper.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Kasten:1995:HGI

[Kas95b] Eric Kasten. HTML: A gentle introduction. Linux Journal, 15:??, July 1995. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic).  URL http://www.
linuxjournal.com/issue15/
html15.html.

Kasten:1995:HFI

[Kas95c] Eric Kasten. HTML forms: Interacting with the net. Linux Journal, 16:??, August 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

tronic).  URL http://www.
linuxjournal.com/issue16/
html16.html.

Kaufman:1996:TTD

[Kau96] S. Kaufman. Tech trek: the designer in cyberspace. CWIS Campus-Wide Information Services, 13(2):17–20, ???? 1996. CODEN CISYF9. ISSN 1065-0741.

Kaysen:1989:LPT

[Kay89] Jesse Kaysen. Let’s play tag:

The promise of SGML: Standard Generalized Markup Language — it’s not just for publishers. OCLC Micro, 5(5):26–28, October 1989. CODEN OCMIEK. ISSN 8756-5196.

Kay:1996:VD

[Kay96] Muder Kay. VRML for Dummies. I D G Books Worldwide, Indianapolis, IN, USA, February 1996. ISBN 1-56884-611-8. ???? pp. LCCN ???? US$29.99.

Kazuhiko:1995:MRI

[Kaz95] T. Kazuhiko. Multimedia and its relationship with information processing and management. Joho kanri = Information & documentation, 38(1):47–63, April 1995. CODEN JOKAAB. ISSN 0021-7298.

Kretz:1992:SHI

[KC92] Francis Kretz and Françoise Colaitis. Standardizing hypermedia information objects. IEEE Communications Magazine, 30 (5):60–70, May 1992. CODEN ICOMD9. ISSN 0163-6804. Subscriptions from IEEE Service Center 445 Hoes Lane Piscataway, NJ 08854-4150 (908)981-0060 (800)678-IEEE.

Kim:1995:DIH

[KCCJ95] Hyunki Kim, Kijin Chang, Jaewoo Chang, and Jinok Jeon. Design and implementation of hypermedia browser based on Dexter hypertext reference model. Journal of KISS[C] [Computing Practices], 1(1):54–64, September 1995.

Kroeker:1995:IDC

[KCFP95] Brian Kroeker, Hilde Colenbrander, Henry Fan, and George Phillips. Interactive database communication across the World Wide Web: Accessing Statistics Canada’s CAN-SIM database. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
SocialSci/kroeker/kroekeru.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Kahn:1998:EML

[KD98] C. E. Kahn and N. B. De la Cruz. Extensible markup language (XML) in health care: Integration of structured reporting and decision support. In Chute [Chu98], pages 725–729. ISBN ???? ISSN 1091-8280. LCCN ????

Keller:1995:IGV

[KDC95] C. P. Keller, T. J. Davis, and R. R. Canessa. Instructing GIS in the virtual classroom using network distributed multimedia and hypertext. In Anonymous [Ano95m], pages 545–553 (vol. 2). ISBN 1-57083-026-6 (set), 1-57083-027-4 (vol. 1), 1-57083-028-2 (vol. 2). LCCN G70.2 .G5743 1995. Two volumes.

Kooros:1996:JBW

[KDH96] Paul Kooros, Michele De Wolfe, and David Heib. JavaScript: Blueprints to a world beyond HTML. Prima Tech Publishing, Rocklin, CA, USA, August 1, 1996. ISBN 0-7615-0685-3. xii + 337 pp. LCCN A76.73.J39K66 1996. US$35.00.

Kearsley:1996:RLT

[Kea96] C. D. Kearsley. Re-locating literary teaching and learning with VRML. In Chambers [Cha96b], pages 137–144. ISBN ???? LCCN ????

Keeler:1995:ESS

[Kee95] Michael J. Keeler. The Earth System Science Community Curriculum Testbed. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Educ/keeler/ESSCCTestbed/
ESSCCTOC.html. A selected few

of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Kelly:1994:BIPa

[Kel94a] B. Kelly. Becoming an information provider on the World Wide Web. In Plattner and Kiers [PK94b], pages 122/1– 122/5 (vol. 1). ISBN ???? LCCN ???? 2 vol.

Kelly:1994:BIPb

[Kel94b] Brian Kelly. Becoming an information provider on the World Wide Web. Computer Networks and ISDN Systems, 27 (3):353–360, December 16, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1994&
                                        volume=27&issue=3&aid=1354.

Kelsey:1998:OKR

[Kel98] Robert Lynn Kelsey. Objectbased knowledge representation implemented in SGML. Computer-produced typeface. thesis (ph.d.), New Mexico State University, Las Cruces, NM, USA, 1998. xiii + 173 pp.

Kennedy:1996:IVW

[Ken96]  Randall C. Kennedy.  In-

stant VRML Worlds. Que Corporation, Indianapolis, IN, USA, July 1996. ISBN 1-56276-421-7. xii + 128 pp. LCCN TK5105.888.K45 1996. US$34.99.

Keogh:1997:WMG

[Keo97] James Edward Keogh. The Web master’s guide to VB script. AP Professional, Boston, MA, USA, 1997. ISBN 0-12-703875-2. ix + 254 pp. LCCN QA76.73.V27K46 1997.

Kerr:1997:WPI

[Ker97] Mark Kerr. Web publishing: introduction to HTML. Library Information Technology Centre, London, UK, 1997. ISBN 1-900508-06-0 (spiral). 105 pp. LCCN ????

Kesseler:1995:SAH

[Kes95] M. Kesseler. A schema-based approach to HTML authoring. In O’Reilly and Associates and Web Consortium (W3C) [WWW95], pages 619– 632. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95. URL http:

//www.ora.com/gnn/bus/ora/
item/wj1.html. The World

Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Keyhani:1994:BEJ

[Key94] Andrea Keyhani. Building an electronic journal. In Williams [Wil94b], pages 257–261. ISBN ???? LCCN ????

Kobayashi:1998:DSS

[KFHU98] K. Kobayashi, M. Fujii, M. Hayamizu, and K. Ujihara. Data sharing of shape model with VRML environment. In Jacucci [Jac98], pages 187–194. ISBN 0-412-83540-1. LCCN TS155.6 .I569 1998. Includes CD-ROM.

Kerven:1996:HHD

[KFZ96] David Kerven, Jeff Foust, and John Zakour. HTML 3 How-To: The Definitive HTML Problem-Solver. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN 1-57169-050-6. xxx + 640 pp. LCCN QA76.76.H94 K47 1996. US$39.99. URL http:/

                                        /w3.waite.com/waite/waite/
                                        books.new/HTML_3_How_To/html/
                                        html3htcov.html.

Kerven:1997:HPH

[KFZ97] David Kerven, Jeff Foust, and John Zakour. HTML 3.2 Plus how-to. Waite Group Press, Corte Madera, CA, USA, 1997. ISBN 1-57169-091-3. xxiii + 939 pp. LCCN QA76.76.H94K4713 1997. US$49.99.

Knapp:1997:HJE

[KG97] Michael Knapp and Wolfram Gieseke. HTML and JAVA 100% efficace. Ed. Micro application, Paris, France, 1997. ISBN 2-7429-0924-9. 364 pp. LCCN ????

Kumar:1994:SWS

[KGK94] Vinay Kumar, Jay Glicksman, and Glenn A. Kramer. SHAREd Web to support design teams.

In Proceedings of the 3rd Workshop on Enabling Technologies: Infrastructure for Collaborative Enterprises (Apr 17–19 1994: Morgantown, WV, USA), pages 178–182. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. IEEE catalog number 94TH0632-0.

Kidder:1995:HPI

[KH95a] Gayle Kidder and Stuart (Stuart H.) Harris. HTML publishing with Internet Assistant: your guide to using Microsoft’s HTML add-on. Ventana Press, Chapel Hill, NC, USA, 1995. ISBN 1-56604-273-9. xx + 250 pp. LCCN QA76.76.H94H373 1995. US$29.95. Includes floppy.

Koopman:1995:SYO

[KH95b] Ann Koopman and Sharon Hay.

Swim at your own risk — no librarian on duty: Large-scale application of Mosaic in an academic library. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
LibApps/hay/WWWPap.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Kidder:1997:OHPb

[KH97] Gayle Kidder and Stuart (Stuart H.) Harris. Official HTML publishing for Netscape: your complete guide to Web page design and production. Ventana Press, Chapel Hill, NC, USA, second edition, 1997. ISBN 1-56604-650-5. xxviii + 710 pp. LCCN QA76.76.H94H372 1997.

Khozeimeh:1994:CCE

[Kho94] Khalil Khozeimeh, editor. Computing in civil engineering: proceedings of the First Congress held in conjunction with A/E/C Systems ’94 (Washington, DC, June 20–22, 1994). ASCE, New York, NY, USA, 1994. CODEN CCENEX. ISBN 0-7844-0026-1. LCCN TA345 .C657 1994.

Kelsey:1997:OMK

[KHW97] R. L. Kelsey, R. T. Hartley, and R. B. Webster. An objectbased methodology for knowledge representation in SGML. In IEEE [IEE97d], pages 304– 313. ISBN 0-8186-8203-5, 0-8186-8204-3 (case), 0-8186-8205-1 (microfiche). ISSN 1082-3409. LCCN Q334 .I54 1997.

Kidder:1997:OHPa

[Kid97] Gayle Kidder. Official HTML publishing for Netscape, 1997. Includes CD-ROM.

Kikuta:1996:WCO

[Kik96] M. Kikuta. What is CALS? 2. the organizational structure in the network age. Joho kanri = Information & documentation, 38(10):901–911, January 1996. CODEN JOKAAB. ISSN 0021-7298.

Kilpelainen:1999:SXC

[Kil99] Pekka Kilpeläinen. SGML & XML content models. Markup Languages: Theory & Practice, 1(2):53–76, Spring 1999. CO-DEN MLTPFG. ISSN 1099-6621.

Kim:1994:OOP

[Kim94] Won Kim. Observations on the ODMG-93 proposal for an object-oriented database language. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):4–9, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kim:1997:SPP

[Kim97a] Y.-D. Kim, editor. Selected papers from the proceedings of 1996 ICC&IC, volume 33(3/4) of Computers and Industrial Engineering. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1997. ISBN ???? ISSN 0360-8352 (print), 1879-0550 (electronic). LCCN ????

Krishnamurthy:1997:DMS

[KIM^{+}97b] Kasthurirangan Krishnamurthy, Michael Iuliano, Charles. McLean, National Institute of Standards, and Technology (U. S.). Development of manufacturing systems models using VRML. NIS-TIR; 6093. U.S. Dept. of Commerce, Technology Administration, National Institute of Standards and Technology, Gaithersburg, MD, USA, 1997. various pp.

Kircz:1991:RSS

[Kir91] J. G. Kircz. Rhetorical structure of scientific articles: the case for argumentational analysis in information retrieval. Journal of Documentation, 47(4):354–372, December 1991. CODEN JDO-CAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Kiser:1995:SGM

[Kis95] Betsy N. Kiser. Standard Generalized Markup Language: Why reference librarians should care. RSR: Reference Services Review, 18(3):37–??, 1995. ISSN 0090-7324.

Kandlur:1998:MCN

[KJR98] Dilip D. Kandlur, Kevin Jeffay, and Timothy Roscoe, editors. Multimedia computing and networking 1999: 25–27 January, 1999, San Jose, California, number 3654 in SPIE Proceedings. SPIE Optical Engineering Press, Bellingham, WA, USA, 1998. ISBN 0-8194-3125-7. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.3654.

Kieninger:1994:HHS

[KK94] T. Kieninger and N. Kuhn. Hyperbraille — a hypertext system for the blind. In ACM [ACM94a], pages 92–99. ISBN 0-89791-649-2. LCCN HV1569.5.A24 1994. ACM order number 444940.

Kummerfeld:1995:ICC

[KK95] Robert J. Kummerfeld and J. Kay. An individualised course for the C programming language. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/
kummerfeld/kummerfeld.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Kim:1998:NGC

[KK98] H. J. Kim and K. S. Kim.

A novel GUI for car navigation system based on VRML. In IEEE [IEE98d], pages 182– 183. ISBN 0-7803-4357-3 (softbound), 0-7803-4358-1 (casebound), 0-7803-4359-X (microfiche). LCCN TK7869 .I16 1998. IEEE Catalog Number 98CH36160.

Kim:1997:SMC

[KKK97] N. K. Kim, Y. Kim, and S.-H.

Kang. Subdivision methods of converting STEP into VRML on Web. In Kim [Kim97a], pages 497–500. ISBN ???? ISSN 0360-8352 (print), 1879-0550 (electronic). LCCN ????

Kappetijn:1991:EME

[KL91] J. Kappetijn and G. Lutz. Electronic manuals eliminate mountains of paper. Engineering and Automation, 13(1):4–7, January–February 1991. CO-DEN EAPPEY. ISSN 0931-6221.

Klein:1994:HMS

[Kle94] H.-J. Klein. How to modify SQL queries in order to guarantee sure answers. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(3): 14–20, September 1994. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Klee:1996:TSS

[Kle96] Karl J. Klee, editor. Twenty-Seventh SIGCSE Technical Symposium on Computer Science Education, volume 28(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY, USA, March 1996. CODEN SIGSD3. ISBN 0-89791-757-X. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA1 .A86. ACM order number: 457960.

Kling:1995:ECW

[Kli95]  Arnold Kling.  The eco-

nomic consequences of the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
ComEc/kling/intro.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Kirriemuir:1994:DAD

[KLM94] J. W. Kirriemuir, M. Lawes, and D. Miller. Development of an academic department’s World Wide Web pages. Axis, 1(4): 16–24, ???? 1994. CODEN AXISEN. ISSN 1352-8971.

Kikuchi:1996:UID

[KMA^{+}96] H. Kikuchi, Y. Mishina, M. Ashizawa, N. Yamazaki, and H. Fujisawa. User interface for a digital library to support construction of a ‘virtual personal library’. In IEEE [IEE96i], pages 429–432. ISBN 0-8186-7399-0. LCCN ???? IEEE Catalog No. 96CB35954.

Kirstein:1995:COP

[KMK95] Peter Kirstein and Goli Montasser-Kohsari. The C-ODA project: Experiences and tools. The Computer Journal, 38(8):670– 680, ???? 1995. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/
computer_journal/Volume_38/
Issue_08/Vol38_08.body.html#
AbstractKirstein.

Klaproth:1999:MDM

[KMS99] F. Klaproth, W. Muehl, and R. Schimmer. A meta data model for distributed library systems: Building a magic mirror (XML/RDF) for RDBs. In IEEE [IEE99a], pages 89– 91. ISBN 0-7695-0219-9, 0-7695-0221-0 (microfiche). ISSN 1092-9959. LCCN TK5103.5.F678 1999; ZA4080 .F67 1999. IEEE Computer Society Order Number PR00219.

Kent:1995:CWAb

[KN95a] R. E. Kent and C. Neuss. Creating a Web analysis and visualization environment. Computer Networks and ISDN Systems, 28(1–2):109–??, December 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kent:1995:CWAa

[KN95b] Robert E. Kent and Christian Neuss. Creating Web analysis and visualization environment. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Autools/kent/kent.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Kochtanek:1994:SFT

[Koc94] Thomas R. Kochtanek. Standards for full text document storage. In Williams [Wil94b], pages 301–307. ISBN ???? LCCN ????

Kohlhepp:1995:NGW

[Koh95] Robert J. Kohlhepp. Next generation Web browsing. Network Computing, 6(9):48–??, August 1, 1995. CODEN NCOMEV. ISSN 1046-4468. URL http:

                                        //techweb.cmp.com/techweb/
                                        nc/609/609sneak1.html.

Komachi:1991:ISO

[Kom91] Y. Komachi. International standardization overview on document description languages and fonts. Journal of the Information Processing Society of Japan = Joho Shori, 32(10): 1110–1117, ???? 1991. CODEN JOSHA4. ISSN 0447-8053.

Komachi:1993:ISF

[Kom93a] Y. Komachi. International standardization for font information interchange. Journal of the Information Processing Society of Japan = Joho Shori, 34(7): 915–921, July 1993. CODEN JOSHA4. ISSN 0447-8053.

Kommers:1993:SDE

[Kom93b] P. Kommers. Scenarios for the development of educational hypermedia. Educational and Training Technology International, 30(3):234–254, August 1993. CODEN ETTIEU. ISSN 0954-7304.

Komachi:1994:ISM

[Kom94]  Y. Komachi.  International

standardization for multimedia/ hypermedia information interchange. Journal of the Information Processing Society of Japan = Joho Shori, 35(7):632–641, July 1994. CODEN JOSHA4. ISSN 0447-8053.

Korpela:1998:IWL

[Kor98]  Jukka Korpela.  Internet

watch: Lurching toward Babel: HTML and CSS and XML. Computer, 31(7):103– 104, July 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1998/pdf/r7103.
pdf.

Kosek:1998:HTD

[Kos98] Jiri Kosek. HTML: tvorba dokonalych www stranek: podrobny pruvodce. Pruvodce. Grada, Praha, vyd. 1. edition, 1998. ISBN 80-7169-608-0. 291 pp. LCCN ????

Kennel:1996:WWW

[KPD96] A. Kennel, L. Perrochon, and A. Darvishi. WAB: World Wide Web access for blind and visually impaired computer users. SIGCAPH Newsletter, June 1996. CODEN SGNWD2. ISSN 0163-5727.

Kim:1999:NSV

[KPIE99] Shi B. Kim, Senior Project-Industrial, and Manufacturing Engineering. NT server-based visual basic application with HTML. Thesis (b.s.), California Polytechnic State University, San Luis Obispo, CA, USA, 1999. 99 pp.

Kuhlen:1995:HIR

[KR95] R. Kuhlen and M. Rittberger, editors. Hypertext — Information Retrieval — Multimedia Proceedings HIM ’95, volume 20 of Schriften zur Informationswissenschaft. Universitätsverlag Konstanz, Konstanz, Germany, 1995. ISBN 3-87940-509-3. LCCN ????

Kristensen:1998:TRX

[Kri98] Anders Kristensen. Template resolution in XML/HTML. Computer Networks and ISDN Systems, 30(1–7):239–249, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/
store/comnet/sub/1998/30/
1-7/1837.pdf.

Kristensen:1999:FXF

[Kri99] Anders Kristensen. Formsheets and the XML forms language. Computer Networks (Amsterdam, Netherlands: 1999), 31 (11–16):1189–1201, May 17, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.com/cas/tree/
store/comnet/sub/1999/31/
11-16/2155.pdf.

Kroeker:1999:NTN

[Kro99] Kirk L. Kroeker. New tools: Net development: Sun’s Java embedded server; MetaCreation’s Web-savvy graphics tool; WebCompiler’s HTML packaging tool. component technology: ProtoVIew Development’s ActiveX tools; Avilon Software’s load balancing component system. software development: Verilog’s test checker; Red Hat and Metrowerks’ development tools for Linux; the Object Factory’s optimization tool; Acumen Systems’s SDK for imaging; Aonix’s process-oriented lifecycle environment; Baan’s embedded software development suite. Computer, 32(5):103– 107, May 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co1999/pdf/r5103.
                                        pdf.

Koegel:1993:HHE

[KRRK93] John F. Koegel, Lloyd W. Rutledge, John L. Rutledge, and Can Keskin. HyOctane: A Hy-Time engine for an MMIS. In ACM [ACM93d], pages 129–136. ISBN ???? LCCN ????

Kruger:1990:EPS

[Kru90a] M. Kruger. Electronic publishing with Standard Generalized Markup Language (SGML). In Anonymous [Ano90d], pages 11/1–8. ISBN 92-835-0550-6. LCCN TL500.N6 C6 no.466.

Kruger:1990:TPS

[Kru90b] M. Kruger. Technical publication: strategies and systems. Nachrichten für Dokumentation, 41(5):285–289, October 1990. CODEN NADOAW. ISSN 0027-7436.

Kruse:1996:USI

[Kru96] Matt Kruse. Using server-side includes. Dr. Dobb’s Journal of Software Tools, 21(2):52, 54, 56, February 1, 1996. CODEN DDJOEB. ISSN 1044-789X.

Krueger:1998:XHF

[Kru98]  M. Krueger.  XML/HTML

“on the fly” aus SGML. In Anonymous [Ano98l], pages 2– ?? ISBN ???? LCCN ????

Kent:1995:APH

[KS95a] Paul Kent and Gary Stein.

Adobe PageMill Handbook: Web Authoring Without HTML: Macintosh Edition. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1995. ISBN 1-56830-272-X. xxii + 266 pp. LCCN TK5105.8885.P34K46 1995. US$35.00.

Kimball:1995:WDS

[KS95b] Ralph Kimball and Kevin Strehlo. Why decision support fails and how to fix it. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24(3):92–97, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Knappe:1996:CSA

[KS96] Heiko Knappe and Till S.

Schwalm. Chip Special/ Anwenderpraxis: HTML und Java. Vogel, Würzburg, Germany, 1996. ISBN 3-8259-1382-1 (??invalid checksum??). 138 pp. LCCN ???? 34

DM. URL http://medweb.
uni-muenster.de/zbm/liti/
design.html#473.  Includes

CD-ROM.

Kompatsiaris:1998:RVI

[KS98a] I. Kompatsiaris and M. G. Strintzis. 3D representation of videoconference image sequences using VRML 2.0. In Hutchison and Schafer [HS98], pages 1–12. ISBN 3-540-64594-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1425.

Kompatsiaris:1998:VVI

[KS98b] I. Kompatsiaris and M. G. Strintzis. Visualisation of videoconference image sequences using VRML 2.0. In Theodoridis and other [To98], pages 1089– 1092. ISBN 960-7620-05-4, 960-7620-07-0. LCCN TK5102.9.E97 1998. Four volumes.

Konopnicki:1998:IGW

[KS98c] David Konopnicki and Oded Shmueli. Information gathering in the World-Wide Web: the W3QL query language and the W3QS system. ACM Transactions on Database Systems, 23(4):369–410, December 1998. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tods/1998-23-4/p369-
konopnicki/p369-konopnicki.
pdf;   http://www.acm.org/
pubs/citations/journals/tods/
1998-23-4/p369-konopnicki/
;   http://www.acm.org:80/
pubs/citations/journals/tods/
1998-23-4/p369-konopnicki/.

Kim:1996:OOO

[KSC96] HyunKi Kim, HakGene Shin, and JaeWoo Chang. OOHS: an object-oriented hypermedia system. In IEEE [IEE96d], pages 496–501. ISBN 0-8186-7579-9. LCCN QA 76.6 C6295 1996. IEEE catalog number 96CB35986.

Katkere:1996:IVW

[KSGJ96] Arun Katkere, Jennifer Schlenzig, Amarnath Gupta, and Ramesh Jain. Interactive video on WWW: beyond VCR-like interfaces. Computer Networks and ISDN Systems, 28(7–11): 1559–1572, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=
1578.

Katkere:1995:VBW

[KSJ95] A. Katkere, J. Schlenzig, and R. Jain. VRML-Based WWW interface to MPI video. In Nadeau and Moreland [NM95b], pages 25–31, 137. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Krishnamurthy:1998:HMS

[KSM^{+}98] K. Rangan Krishnamurthy, J. P.

Shewchuk, Charles R. McLean, National Institute of Standards, and Technology (U. S.). A hybrid manufacturing system modeling environment using VRML. NISTIR; 6188. U.S. Dept. of Commerce, Technology Administration, National Institute of Standards and Technology, Gaithersburg, MD, USA, 1998. 11 pp.

Kikuta:1995:IPI

[KTK^{+}95] Y. Kikuta, K. Tazawa, N. Kumagai, N. Rokujyou, M. Hiraga, and T. Hashimoto. Intellectual productivity improvement tool for white-collar workers-multimedia information retrieval system FIND2. Joho kanri = Information & documentation, 37(12):1075– 1087, March 1995. CODEN JOKAAB. ISSN 0021-7298.

Kuchlin:1997:IPI

[Küc97] Wolfgang W. Küchlin, editor. ISSAC ’97. Proceedings of the 1997 International Symposium on Symbolic and Algebraic Computation, July 21–23, 1997, Maui, Hawaii. ACM Press, New York, NY, USA, 1997. ISBN ???? LCCN ????

Kuchling:1998:XEM

[Kuc98]  Andrew Kuchling.  XML,

the eXtensible Markup Language. Linux Journal, 55:64–65, November 1998. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Kulkarni:1994:OOE

[Kul94] Krishna G. Kulkarni. Objectoriented extensions in SQL3: A status report. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):478, June 1994. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Kumar:1996:ITC

[Kum96] B. Kumar, editor. Information technology in civil and structural engineering design: taking stock and future directions: Selection of papers presented at the International Conference on Information Technology in Civil & Engineering Design, held between August 14 and 16 at the University of Strathclyde, Glasgow, Scotland. Civil-Comp, Edinburgh, Scotland, 1996. ISBN 0-948749-39-3. LCCN ????

Kilpelaeinen:1997:SE

[KW97a] P. Kilpelaeinen and D. Wood.

SGML and exceptions. In Nicholas and Wood [NW97], pages 39–50. CODEN LNCSD9. ISBN 3-540-63620-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .P63 1996.

Kilpelainen:1997:SE

[KW97b] P. Kilpeläinen and D. Wood.

SGML and exceptions. Lecture Notes in Computer Science, 1293:39–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kimber:1997:AHH

[KW97c] W. Eliot Kimber and Julia A.

Woods. Application of Hy-Time hyperlinks and finite coordinate spaces to historical writing, analysis, and presentation. Journal of the American Society for Information Science, 48(7): 603–??, ???? 1997. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kim:1996:OOH

[KZSC96] HyunKi Kim, ZongCheol Zhoo, HakGene Shin, and JaeWoo Chang. An object-oriented hypermedia system for structured documents. In Du and Sheng [DS96], pages 286–295. ISBN 0-9641699-4-0. LCCN ????

Ladd:1996:UHJb

[Lad96]  Eric Ladd.  Using HTML,

Java and CGI, Platinum Edition. Macmillan Computer Publishing, Indianapolis, IN, USA, November 1996. ISBN 0-614-20289-2. ???? pp. LCCN ???? US$69.99.

Lam:1996:AHM

[Lam96] K. C. Lam. An application of hypertext markup language (HTML) for institute of textiles and clothing in the Internet. In Anonymous [Ano96-40], pages 313–314. ISBN 951-722-545-8. LCCN ????

Lampton:1997:HPI

[Lam97] Christopher Lampton. Home page: an introduction to Web page design. A first book. Franklin Watts, New York, NY, USA, 1997. ISBN 0-531-20262-3. ???? pp. LCCN QA76.76.H94 L34 1997.

Lankford:1991:PDS

[Lan91] Jeff Lankford. Public domain SGML tools wanted. TUGboat, 12(2):302, June 1991. ISSN 0896-3207.

Lantos:1996:HAL

[Lan96] I. Lantos. HTML applications in language learning and teaching. In Borchart [Bor96], pages 300– ?? (or 317–??). ISBN ???? ISSN 1085-3480. LCCN ????

Laplante:1998:HWW

[Lap98]  Bruce Laplante.  Hands-on

World Wide Web using HTML, 1998. 1 videocassette (VHS)(90 min.).

Lapeyre:1999:ATC

[Lap99] Deborah A. Lapeyre. Annotated table of contents: Developing SGML DTDs from text to model to markup. structuring XML documents. Markup Languages: Theory & Practice, 1(1):113–118, Winter 1999. CO-DEN MLTPFG. ISSN 1099-6621.

LaRoe:1995:CCW

[LaR95] R. John LaRoe. Connecting classrooms to the Web: An introduction to HTML. In Association of Small Computer Users in Education (ASCUE) Summer Conference. Proceedings (28th, North Myrtle Beach, South Carolina, June 18–22, 1995), pages 6–?? (or 120–124). ????, ????, June 1995.

LaTorra:1989:STM

[LaT89] Michael LaTorra. Standard text markup: What SGML means for technical writers. Technical Communication, 36(4):382– ??, November 1, 1989. CODEN TLCMBT. ISSN 0049-3155.

Lauritsen:1993:KDL

[Lau93] M. Lauritsen. Knowing documents (legal). In ACM [ACM93c], pages 184–191. ISBN 0-89791-606-9. LCCN ????

URL http://www.acm.org/
pubs/contents/proceedings/
ai/158976.

Lau:1994:BHI

[Lau94] T. Lau. Building a hypermedia information system on the Internet. In Wilkes [Wil94a], pages 192–197. ISBN 0-7803-1936-2. LCCN T10.5.I54 1994. IEEE Catalog No. 94CH3442-1.

Lavery:1997:IIC

[Lav97] William John Lavery, editor. ICT 97: International Conference on Telecommunications: bridging east & west through communications: 2–5 April 1997, Melbourne, Australia: proceedings, number 2 in International Conference on Telecommunications. Office of Continuing Education, Monash University, Clayton, Victoria, Australia, 1997. ISBN 0-7326-1257-8 (vol. 1), 0-7326-1258-6 (vol. 2), 0-7326-1259-4 (vol. 3), 0-7326-1256-X (set). LCCN ???? Three volumes.

Lawgon:1998:NBA

[Law98] George Lawgon. News briefs:

Alliance wants to give voice to the net; Intel pushes for Unix standards; groups seek to end browser wars; initiative hopes to drive basic chip research; HTML+TIME promises better multimedia; keyword browsing promises simple Web access. Computer, 31(12):21–23, December 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1998/pdf/rz021.
pdf.

Lie:1996:HWB

[LB96] Hakon Wium Lie and Bert Bos.

How to Write Better HTML with Style Sheets. Addison-Wesley Longman, Reading, MA, USA, October 1996. ISBN 0-201-41998-X. xv + 279 pp. LCCN TK5105.888.L54 1997. US$29.95.

Lie:1997:CSS

[LB97] Håkon Wium Lie and Bert Bos.

Cascading Style Sheets: Designing for the Web. Addison-Wes-ley, Reading, MA, USA, 1997. ISBN 0-201-41998-X. xv + 279 pp. LCCN TK5105.888.L54 1997. US$29.95.

Landauer:1998:PVW

[LB98a] Christopher Landauer and Kirstie L. Bellman, editors. Proceedings of the Virtual Worlds and Simulation Conference (VWSIM ’98): 1998 Western Multiconference, San Diego, California, January 11– 14, 1998, Catamaran Resort Hotel, volume 30(2) of Simulation series. Society for Computer Simulation International, San Diego, CA, USA, 1998. ISBN 1-56555-137-0. ISSN 0735-9276. LCCN QA76.9.V5 V578 1998.

Li:1998:DVU

[LB98b] L. Li and A. Barnes. Data visualization using Java and VRML. In Anonymous [Ano98t], pages 911–916. ISBN 1-58025-149-8. LCCN ????

Lie:1999:CSS

[LB99a] Håkon Wium Lie and Bert Bos. Cascading Style Sheets: designing for the Web. Addison-Wes-ley, Reading, MA, USA, second edition, 1999. ISBN 0-201-59625-3. xvii + 396 pp. LCCN QA76.76.H94 L54 1999.

Livingston:1999:ECD

[LB99b] Dan Livingston and Micah Brown. Essential CSS and DHTML for Web Professionals. The Prentice Hall essential Web professionals series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-012760-4. xviii + 220 pp. LCCN QA76.76.H94L58 1999.

                                        URL http://www.phptr.com/
                                        ptrbooks/ptr_0130127604.html.

Lysohir:1998:UWA

[LC98] Steve Lysohir and Computer Channel, Inc. Understanding Web-based applications HTML, dynamic HTML, SSL and more, 1998. 1 videocassette (47 min.).

Lu:1998:CSE

[LCLL98] T. Lu, C. Chiang, M. Lin, and C. Lee. A collaborative scene editor for VRML worlds. In Anonymous [Ano98m], pages C– 53–C–62. ISBN ???? ISSN 0167-7055 (print), 1467-8659 (electronic). LCCN ????

Lemay:1996:LLW

[LCM96] Laura Lemay, Justin Couch, and Kelly Murdock. Laura Lemay’s Web Workshop: VRML 2.0 and 3D Graphics. SAMS Publishing, Indianapolis, IN, USA, September 1996. ISBN 1-57521-143-2. xxxi + 462 pp. LCCN T385.L428 1996. US$39.99.

Lu:1993:SUS

[LCW93] Hongjun Lu, Hock Chuan Chan, and Kwok Kee Wei. A survey on usage of SQL. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(4): 60–65, December 1993. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

LeJeune:1995:NHE

[LD95a] Urban A. LeJeune and Jeff Duntemann. Netscape & HTML explorer. Coriolis Group, Scottsdale, AZ, USA, 1995. ISBN 1-883577-57-8. xxv + 723 pp. LCCN TK5105.883.N48 L45 1995. US$39.99, CDN$54.99.

Lieberman:1995:WWB

[LD95b] Jonathan Lieberman and Nadine O’Connor DiVito. WILT: A WWW based Interactive Language Teaching Tool. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ???? URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/
lieberman/lieberman.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Lemay:1997:TYWa

[LD97a] Laura Lemay and Arman Danesh. Teach yourself Web Publishing with HTML 4 in 14 days. SAMS Publishing, Indianapolis, IN, USA, 2nd professional reference ed. / with revisions for the 2nd ed. by arman danesh edition, 1997. ISBN 1-57521-305-2 (paperback). xxxiii + 1126 pp. LCCN QA76.76.H94 L4528 1997.

Lemay:1997:TYWb

[LD97b] Laura Lemay and Arman Danesh. Teach yourself Web publishing with HTML 4 in a week. Sams.net Pub., Indianapolis, IN, USA, 4th ed. / with revision for the fourth edition by arman danesh edition, 1997. ISBN 1-57521-336-2. xxxi + 621 pp. LCCN TK5105.875.I57 L45 1997. US$29.99.

Lemay:1998:H

[LD98] Laura Lemay and M. (Michel)

Dreyfus, trad. HTML 4. Le programmeur. Simon and Schuster MacMillan, Paris, France, 1998. ISBN 2-7440-0373-5. x + 567 pp. LCCN ????

Liu:1996:DWS

[LDW^{+}96] Yew-Huey Liu, P. Dantzig, C. E.

Wu, J. Challenger, and L. M. Ni. A distributed Web server and its performance analysis on multiple platforms. In IEEE [IEE96i], pages 665–672. ISBN 0-8186-7399-0. LCCN ???? IEEE Catalog No. 96CB35954.

Liu:1996:DCM

[LDWN96] Yew-Huey Liu, P. Dantzig, C. E.

Wu, and L. M. Ni. A distributed connection manager interface for Web services on IBM SP systems. In IEEE [IEE96i], pages 2–9. ISBN 0-8186-7399-0. LCCN ???? IEEE Catalog No. 96CB35954.

Lejeune:1996:NNH

[Lej96] Urban Lejeune. New Netscape 2 and HTML Explorer. Coriolis Group, Scottsdale, AZ, USA, May 1996. ISBN 1-883577-91-8. xx + 683 pp. LCCN TK5105.883.N48L46 1996. US$39.99.

Lemay:1995:TYM

[Lem95a] Laura Lemay. Teach yourself more Web publishing with HTML in a week. Sams.net Pub., Indianapolis, IN, USA, 1995. ISBN 1-57521-005-3. xxi + 451 pp. LCCN QA76.76.H94 L45 1995. Contents: New features in HTML and the Netscape extensions — Tables –Creating, using, and understanding images — More about sound and video – Gateway and CGI scripts — Useful forms and scripts – Testing, revising, and maintaining Web presentations – Managing larger presentations and sites — Web server hints, tricks, and tips — Web server security and access control — Real-life informational presentations — Real-life interactive presentations — HTML 3.0 – Future developments in HTML and the Web — Netscape extensions and HTML tables tag reference — Sources for further information — MIME types and file extensions.

Lemay:1995:TYWb

[Lem95b] Laura Lemay. Teach yourself Web publishing with HTML in 14 days. Sams.net Pub., Indianapolis, IN, USA, 1995. ISBN 1-57521-014-2. xxxv + 804 pp. LCCN QA76.76.H94L454 1995. US$39.99, CDN$53.99, UK£37.50. Includes CD-ROM.

Lemay:1995:TYWa

[Lem95c] Laura Lemay. Teach yourself Web publishing with HTML in a week. Howard W. Sams, Indianapolis, IN 46268, USA, 1995. ISBN 0-672-30667-0. xxi + 397 pp. LCCN QA76.76.H94 L4 1995.

Lemay:1996:TYWa

[Lem96a] Laura Lemay. Teach Yourself Web Publishing with HTML 3.0 in 14 Days: Premier. SAMS Publishing, Indianapolis, IN, USA, second edition, June 1996. ISBN 1-57521-096-7. xl + 1054 pp. LCCN QA76.76.H94L452 1996. US$59.99.

Lemay:1996:TYWb

[Lem96b] Laura Lemay. Teach yourself Web publishing with HTML 3.0 in a week. Sams.net Pub., Indianapolis, IN, USA, second edition, 1996. ISBN 1-57521-064-9. xxii + 518 pp. LCCN QA76.76.H94 L45 1996.

Lemay:1996:TYWc

[Lem96c] Laura Lemay. Teach Yourself Web Publishing with HTML 3.2 in a Week. SAMS Publishing, Indianapolis, IN, USA, third edition, September 1996. ISBN 1-57521-192-0. xvii + 582 pp. LCCN QA76.76.H94L45 1996b. US$29.99.

Lemay:1997:AYW

[Lem97] L. Lemay. Animating your Web page. Web Techniques, 2(2):18– ??, 20–22, February 1997. CO-DEN WETEFA. ISSN 1086-556X.

Lengauer:1993:AEF

[Len93] T. Lengauer, editor. Algorithms — ESA’93. First Annual European Symposium Proceedings: September 1993, Bad Honnef, Germany, volume 726 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. ISBN 3-540-57273-2, 0-387-57273-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A43E83 1993.

Leverette:1993:SCS

[Lev93a] T. Leverette. Status of the CALS SGML registry and library. In Anonymous [Ano93j], pages AP–14. ISBN ???? ISSN ???? LCCN ????

Levy:1993:DRD

[Lev93b] D. M. Levy. Document reuse and document systems. Electronic Publishing—Origination, Dissemination, and Design, 6 (4):339–348, December 1993. CODEN EPODEU. ISSN 0894-3982.

Levinson:1995:RSM

[Lev95]  E. Levinson.  RFC 1874:

SGML media types, December 26, 1995. URL ftp://ftp.

                                        internic.net/rfc/rfc1874.
                                        txt;   http://www.math.utah.
                                        edu/pub/rfc/rfc1874.txt.

Levin:1996:DMW

[Lev96a] A. Levin. Document management and the World Wide Web. Managing Office Technology Magazine, 41(5):47–49, May 1996. CODEN MOTEE3. ISSN 1070-4051.

Levinson:1996:ESD

[Lev96b] E. Levinson. Exchanging SGML documents using Internet mail and MIME. Computer Standards and Interfaces, 18(1): 93–102, January 1996. CO-DEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Levitt:1996:NWC

[Lev96c] J. Levitt. The new Web clients [browser comparison]. Information Week, 576(576):73–74, 76, 78, April 22, 1996. CODEN IN-FWE4. ISSN 8750-6874.

Lewis:1991:ECT

[Lew91] S. Lewis. European conformance testing service for SGML. In Anonymous [Ano91b], pages 189–191. ISBN ???? LCCN ????

Lewitt:1995:WAM

[Lew95] J. Lewitt. Web authoring made simple. Information Week, December 1995. CODEN IN-FWE4. ISSN 8750-6874.

Lewis:1997:BWI

[Lew97] S. Lewis. Building a Webbased installation guide using VRML. In IEEE [IEE97b], pages 211–222. ISBN 0-7803-4184-8 (softbound), 0-7803-4185-6 (casebound), 0-7803-4186-4 (microfiche). LCCN T 10.5 I54 1997. IEEE catalog number 97CH36123.

Llirbat:1997:ECR

[LFS97] François Llirbat, Françoise Fabret, and Eric Simon. Eliminating costly redundant computations from SQL trigger executions. SIGMOD Record (ACM Special Interest Group on Management of Data), 26(2):428–439, June 1997. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lieberman:1994:IAD

[LG94a] I. Lieberman and R. Geer. Interactive Authoring and Display System — a PC based document access/reference environment. In IEEE [IEE94a], pages 95–101. ISBN 0-7803-1910-9. LCCN TK 7895 A8 A98 1994. IEEE catalog no. 94CH3436-3.

Lysakowski:1994:CCD

[LG94b] Rich Lysakowski and Charles E. Gragg, editors. Computerized chemical data standards: databases, data interchange and information systems: Symposium — May 5–6, 1993, Atlanta, GA, USA, volume 1214 of ASTM Special Technical Publication 1994. ASTM, Philadelphia, PA, USA, 1994. ISBN 0-8031-1876-7. ISSN 0066-0558. LCCN QD8 .C66 1994.

Love:1995:DIC

[LG95] P. K. Love and M. V. Gosper. Developing interactive course materials: Using HTML to integrate conventional and Internet resources. In Pearce and Ellis [PE95], pages 359–367. ISBN 0-7325-1220-4. LCCN ????

Logan:1997:EPA

[LG97] Elisabeth Logan and Myke Gluck, editors. Electronic publishing: applications and implications: contributed papers to the ASIS midyear meeting, Minneapolis, MN, May 1995, ASIS monograph series. Information Today, Medford, NJ, USA, 1997. ISBN 1-57387-036-6. LCCN Z286.E43 L64 1997.

Lehner:1996:UHD

[LH96] K. Lehner and D. Hartmann.

Using HTML documents as an explanation facility tool in civil engineering software. In Kumar [Kum96], pages 127–132. ISBN 0-948749-39-3. LCCN ????

Lee:1996:DHE

[LHKC96] Taek-Kyung Lee, Jinseok Hong, Young-Kon Ko, and Yoon-Chul Choy. Development of Hy-Time engine and SGML searching engine for hypermedia systems. Journal of KISS(B) (Software and Applications), 23(8): 882–896, August 1996. CODEN CKNBFV.

Lee:1997:IVO

[LHL97] J.-R. Lee, F. C. Hadipriono, and R. E. Larew. Implementing VRML and object-oriented techniques in a pavement construction management system. In Davani and Guest [DG97], pages 165–170. ISBN 1-56555-104-4. ISSN 0735-9276. LCCN T65.5.C65 I583 1997.

Lea:1997:CPA

[LHMM97] R. Lea, Y. Honda, K. Matsuda, and S. Matsuda. Community place: Architecture and performance. In Spencer and Brutzman [SB97], pages 41–50. ISBN 0-89791-886-X. LCCN ???? URL http://

www.acm.org/pubs/contents/

                                        proceedings/graph/253437/.

ACM order number 434973.

Lin:1996:KHD

[LHW96] Zhangxi Lin, M. Hamalainen, and A. B. Whinston. Knowledgebased HTML document generation for automating Web publishing. Expert Systems with Applications, 10(3-4):381–391, ???? 1996. CODEN ESAPEH. ISSN 0957-4174.

LITA:1998:BHD

[Lib98] Library and Information Technology Association (U. S.). Internet and Resources Interest Group. Beyond HTML dynamic and interactive applications for libraries, July 1, 1998. 2 sound cassettes.

Lichnerowicz:1991:ITI

[Lic91] A. Lichnerowicz, editor. Intelligent Text and Image Handling. Proceedings of a Conference. RIAO ’91. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1991. ISBN 0-444-89361-X. LCCN QA76.9.T48C66 1991.

Lie:1995:WWT

[Lie95]  H. W. Lie.  World-Wide

Web technology. In Vandoni [Van95b], page 277. ISBN 92-9083-076-X. LCCN QA76 .C101 1995.

Light:1994:STU

[Lig94] Richard B. Light. The SGML tagger: user’s manual. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1994. ISBN 0-19-268244-X (paperback). 60 pp. LCCN ???? Includes 1 3.5 in. computer disk.

Light:1995:HMG

[Lig95] R. Light. HTML makes a great delivery vehicle for Web-based information. it just isn‘t a sensible place for much of that information to live in. Archives and Museum Informatics, 9(4): 381–387, ???? 1995. CODEN AMUIEA. ISSN 1042-1467.

Light:1997:PX

[Lig97] Richard B. Light. Presenting XML. Sams.net Pub., Indianapolis, IN, USA, 1997. ISBN 1-57521-334-6 (paperback). xxix + 414 pp. LCCN QA76.76.H94L54 1997. US$24.99. URL http://www.

amazon.com/exec/obidos/ISBN=
1-57521-334-6/r/8788-6902305-
556196.

Likavec:1995:MLN

[Lik95]  Jaromir Likavec.  Mosaic

launches new dimensions in the LAN management. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
InfOrg/likavec/lanmanagement.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Lilley:1995:AWP

[Lil95] C. Lilley. Active Web pages (internet). In Anonymous [Ano95j], pages 267–281. ISBN 0-9521097-2-7. LCCN T385.E975 1995.

Lim:1995:UCIa

[Lim95a] Jong-Gyun Lim. Using Coollists to index HTML documents in the Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Searching/lim/coollist.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Lim:1995:UCIb

[Lim95b] Jong-Gyun Lim. Using Coollists to index HTML documents in the Web. Computer Networks and ISDN Systems, 28(1– 2):147–154, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1995&
                                        volume=28&issue=1-2&aid=1538.

Lim:1996:UCI

[Lim96] J-G Lim. Using Coollists to index HTML documents in the Web. Computer Networks and ISDN Systems, 28(1):147–??, ???? 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Linde:1994:HMT

[Lin94] P. L. Linde. HTML and MO-SAIC: a taste for more: use of Hyper Text Mark-up Language (HTML) for the development of interactive multimedia through World Wide Web (WWW). In Plattner and Kiers [PK94b], pages 212/1– 212/7 (vol. 1). ISBN ???? LCCN ???? 2 vol.

Linder:1996:WBM

[Lin96] Josh Linder. Web browsers:

More than meets the eye. Network Computing, 7(7):59– ??, May 1, 1996. CODEN NCOMEV. ISSN 1046-4468.

Lin:1999:SCH

[Lin99] Forest Lin. Short course in HTML. Scott/Jones, Inc., Publishers, El Granada, CA, USA, 1999. ISBN 1-57676-024-3. vii + 236 + 82 pp. LCCN ????

Little:1996:ISE

[Lit96] J. C. Little, editor. Information systems education conference: 13th — October 1996, St Louis, MO, volume 13 of ISECON -Annual Conference-1996. DPMA, ????, 1996. ISBN ???? LCCN ????

Liu:1997:LWT

[Liu97a] Kim M. Liu. Lightweight webbrowsing through HTML validation: research project. Master of science, plan ii, Department of Electrical Engineering and Computer Sciences, University of California, Berkeley, Berkeley, CA, USA, 1997. 21 pp.

Liu:1997:EIH

[Liu97b] Trudy H. Liu. Embedding interactive haptic objects in HTML. Thesis (m. eng.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1997. 88 pp.

Liu:1998:STG

[Liu98] Zhuozhou Liu. A software tool for generating questionnaire surveys on the World Wide Web: a special HTML editor and a CGI program. Thesis (m.s.), University of South Carolina, Columbia, SC, USA, 1998. vi + 122 pp.

Li:1993:ESP

[LJ93] Shanping Li and N. P. Juster. An exchange strategy for product information. IFIP Transactions B [Applications in Technology], B-10:175–186, ???? 1993. CODEN ITBTEH. ISSN 0926-5481.

Lander:1996:SPC

[LJ96] R. Lander and J. E. James. The security professional in cyberspace. Security Management, 40(1):38–42, 44, January 1996. CODEN SECME6. ISSN 0145-9406.

Lavenant:1995:PMW

[LK95] M. G. Lavenant and John A. Kruper. Phoenicia: A model for World-Wide Web based campus information systems. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Campus.Infosys/kruper/lavenant.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Lambert:1997:HHT

[LL97] T. B. Lambert and J. S. Laurent. HTML help: Transition without fear. In Anonymous [Ano97b], pages 398–?? ISBN 0-914548-94-8, 0-914548-95-6. LCCN ????

Lister:1998:HWW

[LL98] Brad Lister and Bruce Laplante.

Hands-on World Wide Web, 1998. 2 videocassettes (180 min.) participant materials.

Leventhal:1998:DXI

[LLF98] Michael Leventhal, David Lewis, and Matthew Fuchs. Designing XML Internet applications. The Charles F. Goldfarb series on open information management. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-616822-1. xxxii + 582 pp. LCCN QA76.76.H94L46.

Lohman:1991:SIE

[LLL^{+}91] Guy M. Lohman, George Lapis, Tobin Lehman, Rakesh Agrawal, Roberta Cochrane, John McPherson, C. Mohan, Hamid Pirahesh, and Jennifer Widom. Starburst II: the extender strikes back! (relational database). SIGMOD Record (ACM Special Interest Group on Management of Data), 20 (2):447, June 1991. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Li:1997:CAD

[LLY97] H. Li, B. Li, and X. Yu, editors. Computer aided design and computer graphics 5th international conference on computer-aided design and computer graphics: International conference; 5th — December 1997, Shenzhen, China, volume 2 of International Conference on Computer Aided Design and Computer Graphics 1997. IAP, ????, 1997. ISBN 7-80003-411-9. LCCN ????

Lavagnino:1995:SMG

[LM95a] J. Lavagnino and E. Mylonas. The show must go on: problems of tagging performance texts. Computers and the humanities, 29(2):113–121, ???? 1995. CO-DEN COHUAD. ISSN 0010-4817.

Ling:1995:DSA

[LM95b] T. W. Ling and Y. Masunaga, editors. Database Systems for Advanced Applications ’95 Proceedings of the Fourth International Conference on Database Systems for Advanced Applications. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1995. ISBN 981-02-2220-3. LCCN QA76.9.D3I5565 1995.

Lawrence:1996:LHM

[LM96] David Lawrence and Dave Mark. Learn HTML 3.0 on the Macintosh. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-88793-2. ???? pp. LCCN QA76.76.H94 L39 1996.

Laurent:1998:HCT

[LM98] S. Laurent and C. McKee.

HTML conversion tools: The good, the bad, and the ugly. In Anonymous [Ano98u], pages 319–?? ISBN 1-891709-00-3, 1-891709-01-1. LCCN ????

Lea:1996:JVW

[LMM96] Rodger Lea, Kouichi Matsuda, and Ken Miyashita. Java for 3D and VRML Worlds. New Riders Publishing, Carmel, IN, USA, December 1, 1996. ISBN 1-56205-689-1. xiii + 399 pp. LCCN QA76.73.J38L43 1996. US$45.00.

Larson:1996:CID

[LMO^{+}96] R. R. Larson, J. McDonough, P. O’Leary, L. Kuntz, and R. Moon. Cheshire II: designing a next-generation online catalog. Journal of the American Society for Information Science, 47(7):555–567, July 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Laxminarayan:1998:IIC

[LMT98] Swamy Laxminarayan and Evangelia Micheli-Tzanakou, editors. 1998 IEEE International Conferenece on Information Technology Applications in Biomedicine: proceedings: ITAB ’98, 16–17 May 1998, Marriott Hotel, Washington DC. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-7803-4973-3, 0-7803-4974-1. LCCN R858.A2I157 1998. IEEE catalog number 98EX188.

Lim:1998:ESH

[LN98] S.-J. Lim and Y.-K. Ng. Extracting structures of HTML documents. In IEEE [IEE98f], pages 420–426. ISBN 0-8186-7225-0. LCCN TK5105.5 .I53 1998.

Lim:1999:WTR

[LN99] S.-J. Lim and Y.-K. Ng. Web-View: A tool for retrieving internal structures and extracting information from HTML documents. In Chen and Lochovsky [CL99], pages 71–80. ISBN 0-7695-0084-6, 0-7695-0085-4 (casebound), 0-7695-0086-2 (microfiche). LCCN QA76.9.D3 I5565 1999.

Lo:1995:ILV

[Lo95] C. Y. Lo. Integrating links and versioning in document management. Australian Computer Science Communications, 17(1): 339–346, ???? 1995. CODEN ACSCDD. ISSN 0157-3055.

Ladd:1996:PEU

[LO96] Eric Ladd and Jim O’Donnell. Platinum Edition Using HTML, Java and CGI. Que Corporation, Indianapolis, IN, USA, November 1996. ISBN 0-7897-0932-5. xlvi + 1471 pp. LCCN QA76.76.H94L33 1996. US$69.99.

Ladd:1997:UHJ

[LO^{+}97] Eric Ladd, Jim O’Donnell, et al., editors. Using HTML 4.0, Java 1.1, and JavaScript 1.2. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1477-9. xxix + 1395 pp. LCCN QA76.76.H94L333 1998. 138,00 DM. URL http:/

/www.mut.ch/profit.htx?&x_
pk=kd84mf7ms9dm.

Ladd:1998:UHJ

[LO98] Eric Ladd and Jim O’Donnell.

Using HTML 4.0, Java 1.1, and JavaScript 1.2. Que Corporation, Indianapolis, IN, USA, second edition, 1998. ISBN 0-7897-1477-9. xxix + 1395 pp. LCCN QA76.76.H94L333 1998.

Ladd:1999:PIH

[LO99a] Eric Ladd and Jim O’Donnell.

Programmation Internet, HTML 4, XML et Java R 2: ressources d’experts. CampusPress France, Paris, France, 1999. ISBN 2-7440-0628-9. xvii + 1175 pp. LCCN ????

Ladd:1999:UHX

[LO^{+}99b] Eric Ladd, Jim O’Donnell, et al.

Using HTML 4, XML, and Java 1.2: Platinum Edition. Que Corporation, Indianapolis, IN, USA, 1999. ISBN 0-7897-1759-X. xxiii + 1282 pp. LCCN QA76.76.H94L332. US$49.99.

Ladd:1996:UHJa

[LOA^{+}96] Eric Ladd, Jim O’Donnell, Jerry Ablan, et al. Using HTML 3.2, Java 1.1, and CGI: Platinum Edition. Que Corporation, Indianapolis, IN, USA, November 1, 1996. ISBN 0-7897-0932-5. xlvi + 1471 pp. LCCN QA76.76.H94L33 1996. US$69.99. Includes two CD-ROMs.

Lobin:1999:IXS

[Lob99] H. Lobin. Informationsmodellierung in XML und SGML. Springer-Verlag Inc., New York, NY, USA, 1999. ISBN 3-540-65356-2. 250 (est.) pp. LCCN ????

Locatis:1997:CMI

[Loc97] Craig Locatis. Creating multimedia for the Internet using HTML. Lister Hill monograph; LHNCBC 97-1. U.S. Dept. of Health and Human Services, Public Health Service National Institutes of Health, National Library of Medicine, Bethesda, MD, USA, 1997. ISBN ???? various pp. LCCN ????

Loeffen:1994:TDS

[Loe94] A. Loeffen. Text databases: a survey of text models and systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):97–106, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lotus:1995:LOP

[Lot95] Lotus Development Corporation. Lotus outlines plans to deliver powerful integration of Notes and World Wide Web— announces Lotus Notes server that includes native Notes support for HTTP, HTML and Java technology. press release, December 13, 1995. URL http:

//www.Lotus.com/corpcomm/
3406.htm.

LaLonde:1996:SWW

[LP96] W. LaLonde and J. Pugh.

Smalltalk on the World Wide Web. Journal of Object Oriented Programming, 9(4):82, 84– 85, July–August 1996. CODEN JOOPEC. ISSN 0896-8438.

Li:1997:MPT

[LQ97] S. Li and G. Qian. A method for product technical document integration with SGML. In Li et al. [LLY97], pages 695–698. ISBN 7-80003-411-9. LCCN ????

Litwin:1994:ADF

[LR94] Witold Litwin and Tore Risch, editors. Applications of databases: first international conference, ADB-94, Vadstena, Sweden, June 21–23, 1994: proceedings, number 819 in Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. ISBN 0-387-58183-9, 3-540-58183-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 A67 1994.

Lefevre:1995:OSD

[LR95] P. Lefevre and F. Reynaud. ODIL: An SGML description language of the layout structure of documents. In IEEE [IEE95f], pages 480–488. ISBN 0-8186-7128-9. LCCN TA1650 .I58 1995. Two volumes.

Lee:1996:MFS

[LRG96] Yuet C. Lee, Shekar Rao, and Arif Ghafoor, editors. Multimedia: Full-Service Impact on Business, Education, and the Home: 23–24 October 1995, Philadelphia, Pennsylvania, volume 2617 of Proceedings of the SPIE — The International Society for Optical Engineering. SPIE Optical Engineering Press, Bellingham, WA, USA, 1996. CODEN PSISDG. ISBN 0-8194-1981-8. ISSN 0277-786X (print), 1996-756X (electronic). LCCN QA76.575 .M848 1996.

Lee:1996:TAD

[LS96] J. C. Lee and C. E. Swietlik. Toward automated document reformatting: a SGML markup system. In Naraghi-Pour [NP96], pages 137–141. ISBN 0-8186-7352-4. LCCN TA 168 S727p 1996. IEEE Catalog No. 96TB100023.

Li:1998:UHO

[LS98] C. X. Li and J. Sun. Using hyperlink to organize SAS HTML output. In Anonymous [Ano98t], pages 986–990. ISBN 1-58025-149-8. LCCN ????

Lie:1999:MWP

[LS99] Håkon Wium Lie and Janne Saarela. Multipurpose Web publishing using HTML, XML, and CSS. Communications of the ACM, 42(10):95–101, October 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/cacm/1999-42-10/
p95-lie/.

Liu:1996:GPD

[LSC96] L. Y. Liu, A. L. Stumpf, and S. Chin. Global project documentation and communications using HTML on the World Wide Web. In Vanegas and Chinowsky [VC96], pages 15–20. ISBN 0-7844-0182-9. LCCN TA345 .C649 1996.

Liechti:1998:SGF

[LSI98] Olivier Liechti, Mark J. Sifer, and Tadao Ichikawa. Structured graph format: XML metadata for describing Web site structure. Computer Networks and ISDN Systems, 30(1–7): 11–21, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

URL http://www.elsevier.
com/cas/tree/store/comnet/
sub/1998/30/1-7/1850.pdf.

Lakshmanan:1996:DLQ

[LSS96] L. V. S. Lakshmanan, F. Sadri, and I. N. Subramanian. A declarative language for querying and restructuring the Web.

In IEEE [IEE96o], pages 12– 21. ISBN 0-8186-7289-7. LCCN QA76.9.D3 I593 1996. IEEE Catalog No. 96TB100021.

Ludeking:1995:MSB

[LSW95] L. D. Ludeking, D. N. Smithe, and G. Warren. The MMACE Source Book and Collector Tool. In IEEE [IEE95a], page 243. ISBN 0-7803-2669-5. LCCN TA 2020 I612c 1995. IEEE Catalog No. 95CH35796.

LeVan:1986:LDF

[LT86] H. Le Van and E. Terreni. A language to describe formatting directives for SGML documents. In Désarménien [Dés86], pages 98–119. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN Z253.4.T47 T49 1986.

Lemay:1998:TYW

[LT98] Laura Lemay and Denise Tyler. Teach yourself Web Publishing with HTML 4 in 21 days. SAMS Publishing, Indianapolis, IN, USA, new edition, 1998. ISBN 0-672-31345-6. xxii + 795 pp. LCCN QA76.76.H94L44 1998.

Lemay:1999:TYW

[LT99] Laura Lemay and Denise Tyler. Teach yourself Web Publishing with HTML 4 in 21 days. SAMS Publishing, Indianapolis, IN, USA, professional reference edition. edition, 1999. ISBN 0-672-31408-8. xxxv + 1231 pp. LCCN QA76.76.H94 L4529 1999.

Lunemann:1995:AUS

[Lun95a] Rhonda S. Lunemann. Advantages of using SGML (nuts and bolts of SGML). Technical Communication, 42(2):344–346, May 1995. CODEN TLCMBT. ISSN 0049-3155.

Lunemann:1995:CGF

[Lun95b] Rhonda S. Lunemann. Creating graphics files for SGML documents. Technical Communication, 42(1):160–163, February 1995. CODEN TLCMBT. ISSN 0049-3155.

Lunemann:1995:NBS

[Lun95c] Rhonda S. Lunemann. Nuts and bolts of SGML. Technical Communication, 42(2):344–??, May 1, 1995. CODEN TLCMBT. ISSN 0049-3155.

Lunt:1996:GGW

[Lun96] P. Lunt. Groupware gets Webified [intranets in banking organisations]. ABA Banking Journal, 88(6):70–72, June 1996. CO-DEN ABAJD5. ISSN 0194-5947.

Luoma:1997:ISM

[Luo97] R. N. Luoma. Implementing SGML in the mainstream: The first steps. In Anonymous [Ano97b], page 494. ISBN 0-914548-94-8, 0-914548-95-6. LCCN ????

Lutz:1998:NBM

[Lut98] Michael J. Lutz. New books:

Making wavelets in the real world; protecting the OB; Diffie on privacy; digital design explained; E-commerce du jour; seeing software; structures for XML; Smalltalk patterns. Computer, 31(6):97, June 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1998/pdf/r6097.
pdf.

Lynch:1996:HW

[Lyn96] K. Lynch. HTML: On and off the Web. In Debreceny and Ellis [DE96], pages 501–502. ISBN 1-875855-16-5. LCCN ????

Lewontin:1995:DWP

[LZ95] Steve Lewontin and Mary Ellen Zurko. The DCE Web Project: Providing authorization and other distributed services to the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Security/lewontin/Web_DCE_

Conf_94.html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Lopresti:1998:DRV

[LZ98] Daniel P. Lopresti and Jiangying Zhou, editors. Document recognition V: 28–29 January 1998, San Jose, California, number 3305 in SPIE Proceedings. SPIE Optical Engineering Press, Bellingham, WA, USA, 1998. ISBN 0-8194-2745-4. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.3305.

Lowe:1995:FMD

[LZSD95] B. Lowe, J. Zobel, and R. Sacks-Davis. A formal model for databases of structured text. In Ling and Masunaga [LM95b], pages 449–456. ISBN 981-02-2220-3. LCCN QA76.9.D3I5565 1995.

Morrison:1996:JU

[M^{+}96a] Michael Morrison et al. Java unleashed: Everything you need to master Java. Howard W. Sams, Indianapolis, IN 46268, USA, 1996. ISBN 1-57521-049-5 (paperback). xxxii + 971 pp. LCCN QA76.73.J38J384 1996. US$49.99, CDN$69.99, UK£46.95. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1-57521-049-5.html;
 http://www.mcp.com/cgi-bin/
bag?isbn=1--57521-049-5&last=

/. Includes CD-ROM. CD-ROM includes Symantec Café Lite, Sun’s Java 1.01 Developers Kit, plug-ins and helper apps, and HTML editors and utilities for Windows and Macintosh. Covers Version 1.0.

Morrow:1996:JU

[M^{+}96b] Cindy Morrow et al., editors.

Java Unleashed. Sams.net Pub., Indianapolis, IN, USA, 1996. ISBN 1-57521-049-5. xxxii + 971 pp. LCCN QA76.73.J38J384 1996. US$49.99.

                                        URL http://merchant.superlibrary.
                                        com:8000/catalog/hg/PRODUCT/
                                        PAGE/15752/bud/1-57521-049-

5.html. Includes CD-ROM with Symantec Café Lite, Sun’s Java 1.01 Developers Kit, plug-ins and helper apps, and HTML editors and utilities for Windows and Macintosh.

Mullery:1997:ISN

[M^{+}97] Al Mullery et al., editors. Intelligence in services and networks: technology for cooperative competition: Fourth International Conference on Intelligence in Services and Networks, IS&N ’97, Cernobbio, Italy, May 27–29, 1997: proceedings, volume 1238 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63135-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.A1 I47985 1997.

Maler:1996:DSD

[MA96] Eve Maler and Jeanne El Andaloussi. Developing SGML DTDs: from text to model to markup. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-309881-

8. xxiv + 532 pp. LCCN

QA76.76.H94 M25 1996.

Macleod:1990:ECL

[Mac90a] I. A. Macleod. Extending the command language interface to handle marked-up documents. In Henderson [Hen90], pages 192–196. ISBN 0-938734-48-2. LCCN Z699.A1 .A506 v.27.

Macleod:1990:SRS

[Mac90b] I. A. Macleod. Storage and retrieval of structured documents. Information Processing and Management, 26(2): 197–208, ???? 1990. CO-DEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Macleod:1991:QLR

[Mac91] I. A. Macleod. A query language for retrieving information from hierarchic text structures. The Computer Journal, 34(3):254– 264, June 1991. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Mackin:1993:WFC

[Mac93] J. Mackin. Work flow control using SGML and Hytime. In Anonymous [Ano93j], pages ST– 134–ST–143. ISBN ???? ISSN ???? LCCN ????

Macaulay:1995:WCG

[Mac95]  Tyson Macaulay.  WWW

clearinghouse for government information. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
NationWideSites/macaulay/
macaulay.html. A selected few

of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Mackie:1996:JVC

[Mac96a] Jamieson Mackie. Java and VRML combination may silence critics of both. Silicon Graphics World, 6(11):27, 29, November 1996. ISSN 1057-7041.

Macleod:1996:SN

[Mac96b] I. A. Macleod. SGML into the nineties. Computer Standards and Interfaces, 18(1):1–??, 1996. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Macmillan:1998:DHC

[Mac98] Macmillan Digital Publishing. Dynamic HTML construction kit, Windows 95/Windows NT 4.0 edition, 1998. ISBN 1-57595-082-0 (container). Includes CD-ROM.

Maguire:1994:SSP

[Mag94] Mark Maguire. Secure SGML: A proposal to the information community. Scholarly Publishing, 25(3):146–156, April 1994. ISSN 0036-634X.

Magel:1997:IWI

[Mag97] Ken Magel. Internet watch: Is it too late to put the user back into HTML? Computer, 30(12):131–132, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

                                        pdf.computer.org/co/books/
                                        co1997/pdf/rz131.pdf.

Malaika:1998:UXR

[Mal98] S. Malaika. Using XML in relational database applications. In Kitsuregawa et al. [K^{+}98], pages 167–?? ISBN 0-7695-0071-4 (softbound), 0-7803-5581-4 (casebound), 0-7695-0073-0 (microfiche). ISSN 1063-6382. LCCN QA76.9.D3 I5582 1999. IEEE Catalog Number 99CB36337.

Mann:1995:NGN

[Man95] K. Mann. Novell gets net-wise (Internet Publisher). Network Computing, 4(5):42–43, May 1995. CODEN NCOMEV. ISSN 0966-7873.

Mansfield:1997:CGV

[Man97] Richard Mansfield. The comprehensive guide to VBScript: the encyclopedic reference for VBScript, HTML and ActiveX. Ventana Press, Chapel Hill, NC, USA, 1997. ISBN 1-56604-470-7. lxvi + 790 pp. LCCN QA76.73.V27M36 1996.

Markey:1991:EHS

[Mar91] Brian D. Markey. Emerging hypermedia standards-hypermedia marketplace prepares for Hy-Time and MHEG. In Anonymous [Ano91f], pages 59–74. ISBN ???? LCCN ????

Markey:1992:HM

[Mar92] B. D. Markey. HyTime and MHEG. In IEEE [IEE92], pages 25–40. ISBN 0-8186-2655-0 (paperback), 0-8186-2656-9 (microfiche). LCCN QA75.5.C58 1992. IEEE catalog no. 92CH3098-1.

Marshall:1994:HBC

[Mar94] A. D. Marshall. Hypertext based computer vision teaching packages. Proceedings of the SPIE — The International Society for Optical Engineering, 2347:207–218, ???? 1994. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Mars:1995:TVL

[Mar95a] N. J. I. Mars, editor. Towards Very Large Knowledge Bases. Knowledge Building and Knowledge Sharing 1995. IOS Press, Amsterdam, The Netherlands, 1995. ISBN 90-5199-217-3. LCCN Q334 .T737 1995.

Marshall:1995:DHC

[Mar95b] A. D. Marshall. Developing hypertext courseware on the World Wide Web. In Maurer [Mau95], pages 418–423. ISBN 1-880094-15-0. LCCN LB1028.5.W67 1995.

Martin:1995:OFU

[Mar95c] Gerard Martin. Online fair use of copyrighted material: Issues and concerns. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Pub/
martin/WWW-Online-Essay.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Marrin:1997:EAI

[Mar97a] Chris Marrin. External Authoring Interface (EAI) Proposal. Silicon Graphics, Inc., Mountain View, CA, USA, 1997. URL

http://www.vrml.org/.

Marti:1997:VIF

[Mar97b] J. Marti. Viewing IGES files through VRML. In Yagel and Hagen [YH97], pages 471–474. ISBN 1-58113-011-2 (ACM), 0-8186-8262-0 (IEEE), 0-8186-8262-0, 0-8186-8263-9 (casebound), 0-8186-8264-7. ISSN 1070-2385. LCCN T385 .I192 1997.

Maranget:1998:HLH

[Mar98] Luc Maranget. HeVeA: L^{A}TEX to HTML translator. World-Wide Web document., 1998.

URL http://para.inria.fr/~
maranget/hevea/. The transla-

tor is written in Objective Caml.

Martin:1999:PCG

[Mar99] Teresa A. Martin. Project cool guide to XML for Web designers. Wiley, New York, NY, USA, 1999. ISBN 0-471-34401-X. xxi + 298 pp. LCCN QA76.76.H94 M277 1999.

Masys:1997:AMI

[Mas97] Daniel R. Masys, editor. The emergence of ‘internetable’ health care: systems that really work: proceedings, 1997 AMIA Annual Fall Symposium, formerly SCAMC, October 25–29, 1997, Opryland Hotel, Nashville, TN, AMIA Annual Symposium 1997. Hanley and Belfus, Philadelphia, PA, USA, 1997. ISBN ???? ISSN 1091-8280. LCCN W26.5 S978 1997.

Mason:1998:XRP

[Mas98] Jon Mason. XML and RDF: the promise and the reality of new Web architectures. Computer Networks and ISDN Systems, 30(1–7):763–765, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/
store/comnet/sub/1998/30/
1-7/1947.pdf.

Matzen:1993:FLM

[Mat93] Richard Walter Matzen. A formal language model for detecting ambiguity in SGML. Thesis (ph.d.), Oklahoma State University, Stillwater, OK, USA, 1993. xii + 144 pp.

Mateosian:1997:MRJ

[Mat97] Richard Mateosian. Micro review: Real SIPs: Java, C++, SGML: Destabilizing forces for companies and careers? IEEE Micro, 17(3):3–4, May/ June 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//pascal.computer.org/mi/
books/mi1997/pdf/m3003.pdf.

Matsumura:1998:MVX

[Mat98] Miko Matsumura. Miko’s vision: XML speeds along in standards land. JavaWorld: IDG’s magazine for the Java community, 3(2), February 1998. CO-DEN ???? ISSN 1091-8906.

URL http://www.javaworld.
com/javaworld/jw-02-1998/
jw-02-miko.htm.

Matzen:1999:NGT

[Mat99] R. W. Matzen. A new generation of tools for SGML. Markup Languages: Theory & Practice, 1(1):47–74, Winter 1999. CODEN MLTPFG. ISSN 1099-6621. URL http://

mitpress.mit.edu/journals/
MLANG/matzen.pdf.

Maurer:1991:H

[Mau91] H. Maurer, editor. Hypertext/Hypermedia ’91. Springer-Verlag Inc., New York, NY, USA, 1991. ISBN 3-540-54145-4. LCCN ????

Maurer:1993:EMH

[Mau93] H. Maurer, editor. Educational Multimedia and Hypermedia Annual, 1993 Proceedings of ED-MEDIA 93 - World Conference on Educational Multimedia and Hypermedia. Association for the Advancement of Computing in Education, Charlottesville, VA, USA, 1993. ISBN 1-880094-06-1. LCCN ????

Maurer:1995:EMH

[Mau95] Hermann A. Maurer, editor. Educational Multimedia and Hypermedia, 1995. Proceedings of ED-MEDIA 95 — World Conference on Educational Multimedia and Hypermedia, Graz, Austria, June 17–21, 1995. Association for the Advancement of Computing in Education, Charlottesville, VA, USA, 1995. ISBN 1-880094-15-0. LCCN LB1028.5.W67 1995.

Mauve:1998:TTA

[Mau98] M. Mauve. TeCo3D: a 3D telecooperation application based on VRML and Java [3654-22]. In Kandlur et al. [KJR98], pages 240–251. ISBN 0-8194-3125-7. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.3654.

Mamrak:1990:GPS

[MB90] Sandra A. Mamrak and Julie Barnes. Guidelines for the preparation of SGML document type definitions. Technical research report OSU-CISRC-7/90-TR18, Computer and Information Science Research Center, Ohio State University, Columbus, OH, USA, 1990. 21 pp.

Mamrak:1991:CPS

[MB91] S. A. Mamrak and J. A. Barnes. Considerations for the preparation of SGML document type definitions. Electronic Publishing—Origination, Dissemination, and Design, 4(1):27– 42, March 1991. CODEN EPODEU. ISSN 0894-3982.

Meyer:1995:ABV

[MB95a] T. Meyer and D. Brookshire Conner. Adding behavior to VRML. In Nadeau and Moreland [NM95b], pages 45–51.

ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Munzner:1995:VSW

[MB95b] T. Munzner and P. Burchard.

Visualizing the structure of the World Wide Web in 3D hyperbolic space. In Nadeau and Moreland [NM95b], pages 33– 38, 138. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Moin:1997:SRO

[MB97a] S. Moin and J.-J. Brault. Structure recognition of on-line handwritten documents based on SGML. In IEEE [IEE97f], pages 649–652. ISBN 0-8186-7899-2, 0-8186-7898-4, 0-8186-7900-X. LCCN TA1640 .I57 1997. Two volumes. IEEE Order catalog number 97TB100138. IEEE Computer Society order number PR07898.

Murshed:1997:ADI

[MB97b] Nabeel A. Murshed and Flavio Bortolozzi, editors. Advances in document image analysis: First Brazilian Symposium, BSDIA’97, Curitabia, Brazil, November 2–5, 1997: proceedings, volume 1339 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63791-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TA1640.B73 1997.

Moline:1990:PHS

[MBB90] J. Moline, D. Benigni, and J. Baronas, editors. Proceedings of the Hypertext Standardization Workshop (NIST SP 500-178). National Institute for Standards and Technology, Gaithersburg, MD, USA, 1990. ISBN ???? LCCN ????

Meyer:1995:WTD

[MBB95] T. Meyer, D. Blair, and D. Brookshire Conner. WAXweb: Toward dynamic MOO-based VRML. In Nadeau and Moreland [NM95b], pages 105–108. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Meyer:1995:WMBa

[MBH95a] Tom Meyer, David Blair, and Suzanne Hader. WAXweb: A MOO-based collaborative hypermedia system for WWW. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        VR/meyer.waxweb/meyer.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Meyer:1995:WMBb

[MBH95b] Tom Meyer, David Blair, and Suzanne Hader. WAXweb: A MOO-based collaborative hypermedia system for WWW. Computer Networks and ISDN Systems, 28(1–2):77–84, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:/

/www.elsevier.com/cgi-bin/
cas/tree/store/comnet/cas_
sub/browse/browse.cgi?year=
1995&volume=28&issue=1-2&
aid=1521.

Macleod:1991:SDN

[MBHL91] I. A. Macleod, D. T. Barnard, D. Hamilton, and M. Levison. SGML documents and nonlinear text retrieval. In Lichnerowicz [Lic91], pages 226–244. ISBN 0-444-89361-X. LCCN QA76.9.T48C66 1991.

Macnee:1995:PDE

[MBK^{+}95] C. A. Macnee, W. Behrendt, J. R. Kalmus, K. G. Jeffery, and M. D. Wilson. Presenting dynamically expandable hypermedia. Information and Software Technology, 37(7):339–350, July 1995. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

Morris:1993:CEL

[MBM93] R. A. Morris, E. M. Blachman, and C. Meyer. A constraint-based editor for linguistic scholars. Electronic Publishing—Origination, Dissemination, and Design, 6(4):349– 360, December 1993. CODEN EPODEU. ISSN 0894-3982.

Mamrak:1993:BAD

[MBO93] S. Mamrak, J. Barnes, and C. S.

O’Connell. Benefits of automating data translation. IEEE Software, 10(4):82–88, July 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Mea:1996:HGS

[MBRB96] Vincenzo Della Mea, Carlo Alberto Beltrami, Vito Roberto, and Davide Brunato. HTML generation and semantic markup for telepathology. Computer Networks and ISDN Systems, 28 (7–11):1085–1094, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1996&
                                        volume=28&issue=7-11&aid=
                                        1602.

Marrin:1997:TYV

[MC97a] Chris Marrin and Bruce Campbell. Teach Yourself VRML 2.0 in 21 Days. SAMS Publishing, Indianapolis, IN, USA, February 1997. ISBN 1-57521-193-9 (paperback). xviii + 479 pp. LCCN QA76.76.H94M26 1997. US$39.99. Includes CD ROM.

McLain:1997:IR

[MC97b] Tim McLain and Classroom Connect. Internet revealed, 1997. 4 videocassettes (107 min.).

McArthur:1994:WWW

[McA94] Douglas C. McArthur. World Wide Web and HTML. Dr. Dobb’s Journal of Software Tools, 19(15):18–20, 22, 24, 26, 86, December 1994. CODEN DDJOEB. ISSN 1044-789X.

McCarthy:1997:UWM

[McC97a] V. McCarthy. Use the Web to manage your network. Datamation, 43(1):128–133, January 1997. CODEN DTMNAT. ISSN 0011-6963.

McClees:1997:CWP

[McC97b] M. McClees. Creating Web pages with HTML. In Gerdin-Jelger et al. [GJTW97], pages 561–?? ISBN 90-5199-362-5, 4-274-90184-X. ISSN 0926-9630. LCCN ????

Meade:1998:MDH

[MCC98] James G. Meade, David Crowder, and Rhonda Crowder. Microsoft Dynamic HTML explorer. Ventana Press, Chapel Hill, NC, USA, 1998. ISBN 1-56604-798-6. xix + 440 pp. LCCN QA76.76.H94M43 1998.

McFedries:1996:CIG

[McF96] Paul McFedries. The complete idiot’s guide to creating an HTML Web page. Que Corporation, Indianapolis, IN, USA, 1996. ISBN 0-7897-0722-5. xviii + 312 pp. LCCN QA76.76 .H94M387 1996. US$24.99, CDN$33.99, UK£23.49.

McFedries:1997:CIG

[McF97a] Paul McFedries. The complete idiot’s guide to creating an HTML Web page. Que Corporation, Indianapolis, IN, USA, second edition, 1997. ISBN 0-7897-1146-X. xxiv + 328 pp. LCCN QA 76.76 H94 M387 1997.

McFedries:1997:CPW

[McF97b] Paul McFedries. Creando una pagina Web con HTML ‘!facil! Prentice-Hall Hispanoamericana, Mexico, DF, Mexico, second edition, 1997. ISBN 968-880-982-9. xxiv + 328 pp. LCCN ????

McFedries:1998:CIG

[McF98] Paul McFedries. The Complete idiot’s guide to creating an HTML 4 Web page. Que Corporation, Indianapolis, IN, USA, third edition, 1998. ISBN 0-7897-1490-6. xxv + 380 pp. LCCN QA76.76.H94M387 1998. Includes CD-ROM.

McGaffey:1991:SVT

[McG91] Robert W. McGaffey. SGML versus/and TEX. TUGboat, 12 (34):406–408, November 1991. ISSN 0896-3207.

McGaffey:1992:ATU

[McG92] Robert McGaffey. Automatic tables using SGML, C, and TEX. TUGboat, 13(3):291–294, October 1992. ISSN 0896-3207.

Madigan:1995:RHL

[MCG^{+}95] D. Madigan, C. R. Chapman, J. R. Gavrin, O. Villumsens, and J. H. Boose. Repertory hypergrids for large-scale hypermedia linking. International Journal of Human-Computer Studies, 43(3):465–481, September 1995. CODEN IHSTEI. ISSN 1071-5819.

McGrath:1998:PSS

[McG98a] Sean McGrath. ParseMe.1st:

SGML for Software Developers. The Charles F. Goldfarb series on open information management. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, January 1998. ISBN 0-13-488967-3 (hardcover). xxiii + 341 pp. LCCN QA76.76.H94M388 1998. US$33.75.

McGrath:1998:XEB

[McG98b] Sean McGrath. XML by Example: Building E-Commerce Applications. The Charles F. Goldfarb series on open information management. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-960162-7. xlviii + 470 pp. LCCN QA76.76.H94M3883 1998. US$39.95. URL http:

//www.phptr.com/ptrbooks/
ptr_0139601627.html.

McLachlan:1995:SLA

[McL95] Gordon McLachlan. Standards:

A look at SGML. HP professional, 9(6):48–??, June 1995. ISSN 0896-145X.

McLain:1998:NHE

[McL98] Tim McLain. News, HTML and email, 1998. 1 videocassette (34 min.).

McMurdo:1996:HL

[McM96] G. McMurdo. HTML for the lazy. Journal of Information Science, Principles and Practice, 22(3):198–212, ???? 1996.

CODEN JIOSED. ISSN 0165-5515 (print), 1741-6485 (electronic).

Maeng:1991:TPG

[MCP91] Seung-Rul Maeng, Zeung-Seong Cho, and Dong-Soon Park. Text processing with a glyph mapping model. Korea Information Science Society Review, 9(6):74– 78, ???? 1991. CODEN CHK-WEN. ISSN 1015-9908.

Mason:1995:CMB

[MCPD95] Joanna Mason, Marek Czernuszenko, Dana Plepys, and Thomas A. DeFanti. CAVEview: Mosaic-based virtual reality. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN

???? URL http://www.ncsa.
uiuc.edu/SDG/IT94/Proceedings/
VR/mason/CAVEview.paper.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Mattos:1994:RDT

[MD94] Nelson Mattos and Linda G. DeMichiel. Recent design tradeoffs in SQL3. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(4): 84–90, December 1994. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Martin:1997:PCG

[MD97] Teresa A. Martin and Glenn Davis. The Project Cool Guide to HTML. Wiley, New York, NY, USA, 1997. ISBN 0-471-17371-1. xix + 245 pp. LCCN QA76.76.H94M275 1997. US$19.95.

McCarthy:1998:BWJ

[MD98a] Martin McCarthy and Alligator Descartes. Building 3D worlds in Java and VRML. Prentice Hall Europe, New York, NY, USA, 1998. ISBN 0-13-748625-1 (paperback). 382 pp. LCCN QA76.73.J38M3527 1998.

McCarthy:1998:RAB

[MD98b] Martin McCarthy and Alligator Descartes. Reality architecture: building 3D worlds with Java and VRML. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-748625-1. xiv + 382 pp. LCCN QA76.73.J38M3527 1998.

Murthy:1998:SLI

[MDK98] T. A. V. Murthy, N. Datta, and R. P. (Ramesh Pal) Kumar, editors. Sustainable library and information services: seminar papers at the Fortythird All India Library Conference, Punjab University Library, Chandigarh, 5-8 November, 1997, volume 43 of All India Library Conference 1999. Indian Library Association, Delhi, India, 1998. ISBN 81-85216-31-2. LCCN Z845.I4 A55 1997.

Martin:1999:EKW

[ME99] P. Martin and P. Eklund. Embedding knowledge in Web documents: CGs versus XMLbased metadata languages. Lecture Notes in Computer Science, 1640:230–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meese:1996:OHJ

[Mee96] Philip David Meese. The one hour Java applet — everything you need to know to build your first Java applet. using this tutorial, at the end of one or two hours you will have a working Java bar-chart applet and HTML page. Datamation, 42 (5):51–61, March 1996. CODEN DTMNAT. ISSN 0011-6963.

Megginson:1998:SXD

[Meg98] David Megginson. Structuring XML documents. The Charles F. Goldfarb series on open information management. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-642299-3. xxxvii + 420 pp. LCCN QA76.76.H94 M44 1998. US$44.95. URL http://www.

prenhall.com/allbooks/ptr_
0136422993.html.

Meinke:1997:CCS

[Mei97] J. G. Meinke, editor. Consortium for Computing in Small Colleges: Southeastern conference; 10th — November 1996, Johnson City, TN, volume 12(3) of Journal of Computing in Small Colleges 1997. Consortium for Computing in Small Colleges, ????, 1997. ISBN ???? LCCN ????

Melby:1995:EET

[Mel95] A. Melby. E-Tif: an electronic terminology interchange format. Computers and the humanities, 29(2):159–165, ???? 1995. CO-DEN COHUAD. ISSN 0010-4817.

Melker:1999:IWN

[Mel99] Alexander I. Melker, editor. International Workshop on Nondestructive Testing and Computer Simulations in Science and Engineering: 8–12 June 1998, St. Petersburg, Russia, number 3687 in SPIE Proceedings. SPIE Optical Engineering Press, Bellingham, WA, USA, 1999. ISBN 0-8194-3161-3. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.3687.

Mendelson:1995:HA

[Men95a] Edward Mendelson. HTML addons. Australian Personal Computer, 16(7):46–??, July 1, 1995. ISSN 0725-4415.

Mendelson:1995:NER

[Men95b] Edward Mendelson. No experience required. PC Magazine, 14 (17):203–??, October 10, 1995. CODEN PCMGEP. ISSN 0888-8507.

Meyers:1999:HWC

[Mey99]  Paul Meyers.  The HTML

Web classroom. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-796111-1. xiv + 306 pp. LCCN QA76.76.H94M49 1999.

Maly:1995:WWA

[MFSF95] Kurt Maly, J. French, A. Selman, and E. Fox. WATERS — the Wide Area Technical Report Server. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Databases/maly/maly.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Marshall:1999:AIN

[MFVG99] I. Marshall, M. Fry, L. Velasco, and A. Ghosh. Active information networks and XML. Lecture Notes in Computer Science, 1653:60–72, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McAlpine:1994:NAC

[MG94a] K. McAlpine and P. Golder. A new architecture for a collaborative authoring systemcollaborwriter. Computer Supported Cooperative Work [CSCW], 2(3):159–174, ???? 1994. CODEN CSCWEQ. ISSN 0925-9724.

MoralesGerman:1994:SPE

[MG94b] D. Morales-German. An SGMLbased programming environment for literate programming. In Botsford et al. [BGG^{+}94], pages 42–49. ISBN ???? LCCN ????

Muller:1994:PIC

[MG94c] Hausi A. Muller and Mari Georges, editors. Proceedings: International Conference on Software Maintenance, September 19–23, 1994, Victoria, British Columbia, Canada, International Conference on Software Maintenance 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6332-4, 0-8186-6330-8, 0-8186-6331-6. LCCN QA76.76.S64 I58 1994. IEEE Catalog No. 94CH3485-0.

Maxwell:1996:BCV

[MGBW96] D. Maxwell, C. Graham, D. Blair, and J. Waldrop. Building compelling VRML worlds. In ACM [ACM96a], pages 511– 512. ISBN 0-89791-746-4. ISSN 1069-529X. LCCN T385 .S54 1996. ACM order number: 428960.

Matzen:1993:MSA

[MGH93] R. W. Matzen, K. M. George, and G. E. Hedrick. A model for studying ambiguity in SGML element declarations. In Deaton et al. [DGBH93], pages 668–676. ISBN 0-89791-567-4. LCCN QA76.76.A65 S95 1993.

Matzen:1997:FLM

[MGH97] R. W. Matzen, K. M. George, and G. E. Hedrick. A formal language model for parsing SGML. The Journal of Systems and Software, 36(2):147–166, February 1997. CODEN JSSODM.

ISSN 0164-1212 (print), 1873-1228 (electronic).

Mana-Gomez:1997:HEA

[MGVS97] A. Mana-Gomez and F. Villalba-Sanchez. HTML 2.0 extension for adaptative image transmission of synthetic images. In Arabnia and Ahmed [AA97], pages 422–428. ISBN 0-9648666-9-2. LCCN ????

Morris:1996:WPD

[MH96] Mary E. S. Morris and Randy J. Hinrichs. Web Page Design: A Different Multimedia. SunSoft Press, Englewood Cliffs, NJ, USA, 1996. ISBN 0-13-239880-X. 306 pp. LCCN QA76.76.H94M65 1996. US$24.95.

Marshall:1994:CIW

[MHMS^{+}94] A. D. Marshall, S. Hurley, S. N. McIntosh-Smith, R. R. Martin, and N. M. Stephens. Courseware on the Internet and the World Wide Web. In Raitt and Jeapes [RJ94], pages 341–355. ISBN 0-904933-88-1. LCCN ????

Mercuri:1998:UHJ

[MHP98] R. Mercuri, N. Herrmann, and J. Popyack. Using HTML and JavaScript in introductory programming courses. In Joyce [Joy98], pages 176–180. ISBN 0-89791-994-7. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA1 .A86 v.30, no.1.

Michie:1997:HTW

[Mic97] Donald Michie. HTML Techniques for Webmasters: Tables, Frames and Advanced Layout. Wiley, New York, NY, USA, January 1997. ISBN 0-471-17610-9. ???? pp. LCCN ???? US$29.95.

Michard:1999:XLA

[Mic99] Alain Michard. XML: langage et applications. Eyrolles, Paris, France, 2e tirage. edition, 1999. ISBN 2-212-09052-8. xi + 361 pp. LCCN ????

Miller:1993:TDO

[Mil93]  Rodger S. Miller.  Textual

databases: an object-based model. In Williama [Wil93a], pages 313–317. ISBN 0-938734-73-3. LCCN ????

Miller:1995:CNB

[Mil95] Michael J. Miller. Casting a net beyond HTML. PC Magazine, 14(20):75–??, November 21, 1995. CODEN PCMGEP. ISSN 0888-8507.

Mintz:1996:SER

[Min96] F. J. Mintz, editor. Safety engineering and risk analysis, 1996: presented at the 1996 ASME International Mechanical Engineering Congress and Exposition, November 17–22, 1996, Atlanta, Georgia, volume 6. American Society of Mechanical Engineers, 345 E. 47th St., New York, NY 10017, USA, 1996. ISBN 0-7918-1538-2. LCCN T55.A1 S226 1996.

Anonymous:1997:YMC

[Min97] Your Mining Co. guide to HTML, 1997. ISSN 1521-2068.

                                        URL http://html.miningco.
                                        com/;   mailto:html.guide@
                                        miningco.com?Subject=Subscribe.

The Mining Co., New York, NY, USA.

Mintert:1998:XHG

[Min98] S. Mintert. XML — herkunft, grundlagen, vorteile. In Anonymous [Ano98l], pages 1–?? ISBN ???? LCCN ????

Mione:1996:XOB

[Mio96] A. N. Mione. X opens on Broadway. Digital Systems Report, 18 (4):8–10, Fall 1996. CODEN DSREFP. ISSN 1086-9638.

Matsushima:1988:TSM

[MK88] T. Matsushima and H. Kohsaki. Trends in standardization of musical description. Journal of the Information Processing Society of Japan = Joho Shori, 29 (6):566–578, ???? 1988. CO-DEN JOSHA4. ISSN 0447-8053.

Mueller:1996:UAI

[MK96a] A. Mueller and S. Kutschekmanesch. Using abductive inference and dynamic indexing to retrieve multimedia SGML documents. In Ruthven [Rut96a], pages 8–?? ISBN 3-540-14582-

6. LCCN QA76.575 .M58 1995.

Musciano:1996:HDG

[MK96b] Chuck Musciano and Bill Kennedy. HTML: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1996. ISBN 1-56592-175-5. xx + 385 pp. LCCN QA76.76.H94 M87 199. US$29.95, CDN$42.95.

Musciano:1997:HDG

[MK97] Chuck Musciano and Bill Kennedy. HTML: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1997. ISBN 1-56592-235-2. xvii + 531 pp. LCCN QA76.76.H94M88 1997. US$32.95. URL http://www.

ora.com/catalog/html2/;
http://www.oreilly.com/catalog/
html2.

MacHale:1998:LDHd

[MK98a] Robert MacHale and KeyStone Learning Systems. Learning dynamic HTML, 1998. 3 videocassettes (277 min.).

MacHale:1998:LDHa

[MK98b] Robert MacHale and KeyStone Learning Systems. Learning dynamic HTML. level 1, 1998. 1 videocassette (115 min.) (3 1/2 in.) + 1 keynote booklet.

MacHale:1998:LDHb

[MK98c] Robert MacHale and KeyStone Learning Systems. Learning dynamic HTML. level 2, 1998. 1 videocassette (110 min.) booklet.

MacHale:1998:LH

[MK98d] Robert MacHale and KeyStone Learning Systems. Learning HTML 4.x, 1998. 6 videocassettes.

MacHale:1998:LHLa

[MK98e] Robert MacHale and KeyStone Learning Systems. Learning HTML 4.X. level 1, 1998. 1 videocassette (126 min.) disk (3 1/2 in.) in pocket attached to back of container.

MacHale:1998:LHLc

[MK98f] Robert MacHale and KeyStone Learning Systems. Learning HTML 4.X. level 3, 1998. 1 videocassette (117 min.).

MacHale:1998:LHLd

[MK98g] Robert MacHale and KeyStone Learning Systems. Learning HTML 4.X. level 4, 1998. 1 videocassette (106 min.) disk (3 1/2 in.) in pocket attached to back of container.

MacHale:1998:LHLe

[MK98h] Robert MacHale and KeyStone Learning Systems. Learning HTML 4.X. level 5, 1998. 1 videocassette (99 min.).

MacHale:1998:LHLf

[MK98i] Robert MacHale and KeyStone Learning Systems. Learning HTML 4.X. level 6, 1998. 1 videocassette (91 min.).

Musciano:1998:HDG

[MK98j] Chuck Musciano and Bill Kennedy. HTML: The Definitive Guide. A nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 1998. ISBN 1-56592-492-4. xviii + 587 pp. LCCN QA 76.76 H94 M88 1998. US$32.95. URL http://www.

oreilly.com/catalog/html3.

Mayer-Kress:1995:HMI

[MKBB95] Gottfried Mayer-Kress, William Bender, and John Bazik. Hyper-media on the Internet as a tool for approaching global problems: A teleconferencing experiment. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
CorInfSys/gmk/Mayer-Kress/
HungerConf.html. A selected

few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Moeckel:1998:MMI

[MKEB98] G. Moeckel, M. Keil, T. Exner, and J. Brickmann. Molecular modeling information transfer with VRML: From small molecules to large systems in bioscience. In Altman et al. [A^{+}98], pages 327–338. ISBN 981-02-3278-0. LCCN QH506.P324 1998.

Mamrak:1987:SAS

[MKNS87] S. A. Mamrak, M. J. Kaelbling, C. K. Nicholas, and M. Share. A software architecture for supporting the exchange of electronic manuscripts. Communications of the ACM, 30(5): 408–414, May 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org/
                                        pubs/toc/Abstracts/0001-0782/
                                        22905.html.

McGrath:1997:DRW

[ML97] F. M. McGrath and K. Latimer. Developing a robust WWW interface to scholarly databases: HTML (PHP + mSQL + Open Web) = increased functionality. In Masys [Mas97], pages 973– ?? ISBN ???? ISSN 1091-8280. LCCN W26.5 S978 1997.

Megginson:1999:XDR

[MLD^{+}99] David Megginson, Michael Leventhal, Robert Ducharme, David Lewis, and Mathew Fuchs. XML Developer’s Resource Library. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-020311-4. ???? pp. LCCN

                                        ???? URL http://www.phptr.
                                        com/bookseri/sgml.html;
                                        http://www.phptr.com/ptrbooks/
                                        ptr_0130203114.html.

Mattos:1990:AIO

[MM90a] N. M. Mattos and B. Mitschang. An approach to integrated office document processing and management. SIGOIS Bulletin, 11 (2-3):118–122, ???? 1990. CO-DEN SIGBEL. ISSN 0894-0819.

Maymard:1990:DSS

[MM90b] C. Maymard and E. Meynieux.

Document standards: the state of the art. Genie Logiciel and Systemes Experts, December 1990. CODEN GLSEED. ISSN 0295-6322.

Melton:1995:OET

[MM95a] Jim Melton and Nelson Mendoca Mattos. An overview of the emerging third-generation SQL standard. SIGMOD Record (ACM Special Interest Group on Management of Data), 24 (2):468, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Muller:1995:UWW

[MM95b] N. J. Muller and L. L. Muller.

Using the World Wide Web for network support. International Journal of Network Management, 5(6):326–341, November– December 1995. ISSN 1055-7148.

Meyer:1996:BIS

[MM96] Eric A. Meyer and Peter E. Murray. Borealis image server. Computer Networks and ISDN Systems, 28(7–11):1123–1137, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=

Moorthy:1997:ALE

[MM97] A. L. Moorthy and P. B. Mangla, editors. Information technology applications in academic libraries in India with emphasis on network services & information sharing: papers presented at the Fourth National Convention for Automation of Libraries in Education and Research (CALIBER-97) at Patiala, 6-8 March 1997. Information and Library Network Centre, Ahmedabad, India, 1997. ISBN 81-900825-0-7. LCCN ????

Morris:1998:C

[MM98] Mary E. S. Morris and Paul Massie. CyberCareers? Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-748872-6. xxxiii + 352 pp. LCCN QA76.25 .M677 1998. US$24.95. Includes chapter on SGML and HTML.

Marden:1999:IWT

[MM99a] Philip M. Marden, Jr. and Ethan V. Munson. Internet watch: Today’s style sheet standards: The great vision blinded. Computer, 32(11): 123–125, November 1999. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1999/pdf/ry123.

pdf. Contains strong criticism of CSS (Cascading Style Sheets) for HTML, and XSL (Extensible Style Language) for XML.

McBride:1999:HMS

[MM99b] Peter K. McBride and Nat McBride. HTML 4.0 made simple. Made simple books. Made Simple Books, Oxford, UK, 1999. ISBN 0-7506-4178-9 (paperback). ix + 198 pp. LCCN ????

McConville:1996:WWL

[MMJ96] D. McConville, J. Magid, and P. Jones. So wide a Web, so little time. Educom review, 31(3): 44–48, May–June 1996. CODEN EDREEW. ISSN 1045-9146.

Morris:1998:SGM

[MMRS98] J. A. Morris, J. E. Mattison, C. D. Riley, and R. Sokolowski. Standard generalized markup language (SGML) in health care. In Anonymous [Ano98o], pages 293–312. ISBN ???? ISSN 1048-8987. LCCN ????

MacKie-Mason:1995:SFA

[MMV95] J. K. MacKie-Mason and H. R.

Varian. Some FAQs about usage-based pricing. Computer Networks and ISDN Systems, 28(1–2):257–??, December 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Macleod:1995:DST

[MN95a] I. A. Macleod and D. Narine.

A depository for structured text objects. In Revell and Tjoa [RT95], pages 272–282. ISBN 3-540-60303-4. LCCN QA267.A1 L43 no.978.

Moreland:1995:VRB

[MN95b] J. L. Moreland and D. R. Nadeau. The Virtual Reality Behavior System (VRBS): Using Perl as a behavior language for VRML. In Nadeau and Moreland [NM95b], pages 63– 68. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Marin-Navarro:1991:AWS

[MNA91a] J. Marin-Navarro and P. E. Alevantis. Alice in the Wonderland of SGML: streamlining text entry in the CELEX databases. Electronic Library, 9(3):155–160, June 1991. CO-DEN ELLIDZ. ISSN 0264-0473.

MarinNavarro:1991:AWS

[MNA91b] J. Marin-Navarro and P. E. Alevantis. Alice in the Wonderland of SGML: streamlining text entry in the CELEX databases. Electronic Library, 9(3):155–160, June 1991. CO-DEN ELLIDZ. ISSN 0264-0473.

Munson:1998:PDD

[MNW98] Ethan V. Munson, Charles K. Nicholas, and Derick Wood, editors. Principles of digital document processing: 4th international workshop, PODDP ’98, Saint-Malo, France, March 29– 30, 1998: proceedings, volume 1481 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. ISBN 3-540-65086-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575.P63 1998.

Mohr:1999:DDA

[Moh99] Stephen T. Mohr. Designing distributed applications: with XML, ASP, IE5, LDAP, and MSMQ. Programmer to programmer. Wrox Press, Chicago, IL, USA, 1999. ISBN 1-86100-227-0. xiii + 458 pp. LCCN ????

Molesworth:1991:OIE

[Mol91] R. Molesworth. The open interchange of electronic documents (open document architecture (ODA)). In Anonymous [Ano91e], pages 4/1–6. ISBN ???? LCCN ????

Moline:1995:ISV

[Mol95a] Jodi Moline. Inline support for vector files: Engineering drawings on the WWW. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
DDay/moline/moline.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Moller:1995:CDE

[Mol95b] H. Moller. CODE-consistent document engineering: consistency and correctness. In Kuhlen and Rittberger [KR95], pages 145–162. ISBN 3-87940-509-3. LCCN ????

Monson:1998:WS

[Mon98] Lynn Monson. The WIDL specification. Dr. Dobb’s Journal of Software Tools, 23(11):92, 94–96, November 1998. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/1998/1998_11/WIDL.
TXT;   http://www.ddj.com/
ftp/1998/1998_11/WIDL.ZIP.

Morgan:1995:UWW

[Mor95a] E. Lease Morgan. Using World Wide Web, and WAIS technologies to create electronic information systems. In Anonymous [Ano95f], pages 3–11. CODEN JFOIEU. ISBN ???? ISSN 1049-6505. LCCN ????

Morris:1995:HFPa

[Mor95b] Mary E. S. Morris. HTML for Fun and Profit. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-359290-1. xxii + 264 pp. LCCN QA76.76.H94 M66 1995. US$35.95. See [Mor95c]. CD-ROM contains shareware and extensive examples of HTML pages and sample perl scripts. Contents: The basics — Hypertext-linking documents — Multimedia-going beyond text — Tables – Using the Common Gateway Interface (CGI)

— Server includes — Creating forms — Processing data from forms — Style guide — Worksaving tools — Putting data on the Internet — Future directions — Xmosaic – WinMosaic — MacMosaic.

Morris:1995:HFPb

[Mor95c] Mary E. S. Morris. HTML for Fun and Profit: Gold Signature Edition. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-242488-6. xxviii + 306 pp. LCCN QA76.76.H94M63 1996¡. US$39.95. Includes CD ROM. Extends the coverage of [Mor95b] with a chapter on Netscape, plus HTML examples on CD ROM, and a Microsoft Windows 3.1 web server on CD ROM.

Morris:1996:HHH

[Mor96] Bruce Morris. HTML in action: hot tools for cool Web sites. Microsoft Press, Bellevue, WA, USA, 1996. ISBN 1-55615-948-X. xiii + 265 pp. LCCN QA76.76.H94 M64 1996.

Moroney:1999:UVC

[Mor99a] N. Moroney. Using VRML to communicate device gamuts [3648-15]. In Beretta and Eschbach [BE99], pages 120–128. ISBN 0-8194-3119-2. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.3648.

Morrison:1999:XU

[Mor99b] Michael Morrison. XML unleashed. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-672-31514-9. 1000 pp. LCCN ????

Moscati:1996:ICI

[Mos96] Paola Moscati, editor. III Convegno internazionale di archeologia e informatica = III International Symposium on Computing and Archaeology. Rome, Italy, volume 7 of Archeologia e Calcolatori. All ’Insegna del Giglio, Firenze, Italia, 1996. ISBN 88-7814-109-7. ISSN 1120-6861. LCCN ???? Two volumes.

McDonald:1995:RMR

[MOTA95] C. J. McDonald, J. M. Overhage, W. M. Tierney, and G. Abernathy. The Regenstrief Medical Record System: Crossinstitutional usage, note writing, and MOSAIC/HTML. In Gardner [Gar95], pages 1029–?? ISBN 1-56053-123-1. ISSN 0195-4210. LCCN W 26.5 S978 1995.

Mumick:1994:IMS

[MP94] Inderpal Singh Mumick and Hamid Pirahesh. Implementation of magic-sets in a relational database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):103–114, June 1994. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Mark:1995:HHT

[MP95a] G. I. Mark and P. Pacher. HTML hypertext tutor for the study of time-dependent scattering. In Chapman [Cha95b], pages 301–308. ISBN 80-901974-0-X. LCCN ????

Mogul:1995:IWL

[MP95b] Jeffrey C. Mogul and Venkata N.

Padmanabhan. Improving WWW latency. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
mogul/HTTPLatency.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Molinari:1996:FRH

[MP96] A. Molinari and G. Pasi. A fuzzy representation of HTML documents for information retrieval systems. In IEEE [IEE96k], pages 107–112. ISBN 0-7803-3646-1 (casebound), 0-7803-3645-3 (softbound), 0-7803-3647-X (microfiche). LCCN TJ212.2.I3249 1996. IEEE catalog number 96CH35998.

Mack:1997:H

[MP97] E. Stephen Mack and Janan Platt. HTML 4.0. No experience required. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1997. ISBN 0-7821-2143-8. xxv + 677 pp. LCCN QA76.76.H94 M23 1997.

Mack:1998:H

[MP98a] E. Stephen Mack and Janan Platt. HTML 4.0. Biblioteca Profesional. Anaya Multimedia, Madrid, Spain, 1998. ISBN 84-415-0464-4. 768 pp. LCCN ????

Murray:1998:HUR

[MP98b] William H. Murray and Chris H. Pappas. HTML 4.0 User’s Resource. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-927104-X. xxiii + 387 pp. LCCN QA76.76.H94M87 1998. US$34.95.

Murray:1998:JHU

[MP98c] William H. Murray and Chris H. Pappas. Javascript and HTML 4.0 user’s resource. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-977422-X. ???? pp. LCCN QA76.73.J39M87 1998.

Murray:1999:JHU

[MP99] William H. Murray and Chris H. Pappas. JavaScript and HTML 4.0 user’s resource. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-977422-X. xxi + 521 pp. LCCN QA76.73.J39 M87 1999.

Min:1993:DMC

[MR93a] Zheng Min and R. Rada. Dexter-Groupware model for collaborative authoring. In Maurer [Mau93], pages 383–390. ISBN 1-880094-06-1. LCCN ????

Min:1993:MCS

[MR93b] Zheng Min and Roy Rada. A model for computer supported collaborative work and document re-use. Intelligent Tutoring Media, 4(1):3–14, February 1993. CODEN ITMEER. ISSN 0957-9133.

Mohr:1993:TOT

[MR93c] W. Mohr and L. Rostek. TEDI:

an object-oriented terminology editor. In Schmitz [Sch93], pages 363–374. ISBN ???? LCCN ????

McLaughlin:1995:JCM

[MR95] Margaret L. McLaughlin and Sheizaf Rafaeli. The Journal of Computer-Mediated Communication. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/
mclaughlin.journal/mclaughlin_

Jour.html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Morris:1996:HFP

[MR96] Mary E. S. Morris and Eric J.

Ray. HTML for fun and profit. Technical Communication, 43 (1):89–??, ???? 1996. CODEN TLCMBT. ISSN 0049-3155.

Mann:1999:HSF

[MR99a] M. Mann and R. Robischon.

HTML standards-history and future. In Bullington et al. [BCG99], pages 51–58. ISBN 0-7890-0768-1. ISSN 0361-526X. LCCN Z692.S5N67 1998.

Mylonas:1999:TEI

[MR99b] E. Mylonas and A. Renear, editors. Text Encoding Initiative: Anniversary conference; 10th — November 1997, Providence, RI, volume 33(1) of Computers and the Humanities 1999; /2. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN ???? ISSN 0010-4817. LCCN ????

Murray-Rust:1999:CMX

[MRR99] Peter Murray-Rust and Henry S. Rzepa. Chemical markup, XML, and the Worldwide Web. 1. Basic principles. Journal of Chemical Information and Computer Sciences, 39(6):928–942, ???? 1999. CODEN JCISD8. ISSN 0095-2338.

Morton:1995:SPS

[MS95] D. Morton and S. Silcot. Systems for providing searchable access to collections of HTML documents. In Debreceny and Ellis [DE95], pages 313–330. ISBN 0-646-23946-5. LCCN ????

MacHale:1998:LDHc

[MS98a] Robert MacHale and Key-Stone Learning Systems. Learning dynamic HTML. level 3, 1998. 1 videocassette (52 min.) booklet.

MacHale:1998:LHLb

[MS98b] Robert MacHale and Key-Stone Learning Systems. Learning HTML 4.X. level 2, 1998. 1 videocassette (117 min.).

Morris:1998:HFP

[MS98c] Mary E. S. Morris and John E. Simpson. HTML for fun and profit. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, third edition, 1998. ISBN 0-13-079672-7. xviii + 378 pp. LCCN QA76.76.H94M648 1998. URL

http://www.fdds.com/books/
books/Morris/Morris.html.

Meyer:1995:MHE

[MSIM95] T. Meyer, R. Suresh, D. Ilg, and B. Moxon. Mosaic, HDF and EOSDIS: Providing access to earth sciences data. Computer Networks and ISDN Systems, 28(1–2):221–??, December 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Miyamoto:1997:DCP

[MSMY97] Y. Miyamoto, T. Sawai, K. Matsui, and K. Yuasa. Development of CAI-SYSTEM for a plant by VRML. In Anonymous [Ano97-42], pages 719– 720. ISBN 0-7803-4208-9. LCCN TK7870.K37 1997. SICE catalog number 97 PR 0001-3. IEEE catalog number 97 TH 8323.

Madigan:1997:LLP

[MSW97] C. Madigan, M. K. Silber, and S. Wilson. Lessons learned prototyping an SGML-based computerized document management system. IEEE Transactions on Professional Communication, 40(2):139–143, June 1997. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Miyama:1999:NAP

[MTH99] Shoken M. Miyama, Kohji Tomisaka, and Tomoyuki Hanawa, editors. Numerical astrophysics: proceedings of the International Conference on Numerical Astrophysics 1998 (NAP98), held at the National Olympic Memorial Youth Center, Tokyo, Japan, March 10– 13, 1998, volume 240 of Astrophysics and space science library. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-5566-0. LCCN QB461.A1 A783 v.240.

Maruyama:1999:XJD

[MTU99] Hiroshi Maruyama, Kento Tamura, and Naohiko Uramoto. XML and Java: developing Web applications. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-48543-5, 0-201-61611-4 (CD-ROM). ix + 386 pp. LCCN QA76.76.H94M28 1999. US$39.95.

Mudry:1998:DC

[Mud98]  Robert Jon Mudry.  The

DHTML Companion. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-796046-8. xi + 333 pp. LCCN QA76.76.H94M74 1998. US$34.95.

Mueller:1995:UWI

[Mue95] Bernd Mueller. Using WWW as an information system to reduce the average period of study by better information providing and to relieve administration. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Educ/mueller/Paper.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Mullen:1996:HQR

[Mul96a] Robert Mullen. HTML Quick Reference. Que Corporation, Indianapolis, IN, USA, July 1996. ISBN 0-7897-0867-1. xvi + 222 pp. LCCN QA76.76.H94M7823 1996. US$19.99.

Muller:1996:WGH

[Mul96b] Nathan J. Muller. The Webmaster’s guide to HTML: for advanced Web developers. Mc-Graw-Hill, New York, NY, USA, 1996. ISBN 0-07-912273-6. xvii + 251 pp. LCCN QA76.76.H94 M79 1996.

Mullen:1997:HQR

[Mul97a] Robert Mullen. HTML 3.2 Quick Reference. Que Corporation, Indianapolis, IN, USA, second edition, 1997. ISBN 0-7897-1144-3. xxxii + 201 pp. LCCN QA76.76.H94H76 1997. US$19.99.

Mullen:1997:HPR

[Mul97b] Robert Mullen. The HTML programmer’s reference. Ventana Press, Chapel Hill, NC, USA, 1997. ISBN 1-56604-597-5. xvii + 348 pp. LCCN QA76.76.H94M78 1997.

Mulroy:1997:VGR

[Mul97c] P. Mulroy. VRML gets real the MPEG 4 way. In Anonymous [Ano97-50], pages 4–?? ISBN ???? ISSN 0963-3308. LCCN ????

Mullen:1998:HPR

[Mul98] Robert Mullen. The HTML 4 programmer’s reference: all platforms. Ventana Press, Chapel Hill, NC, USA, second edition, 1998. ISBN 1-56604-730-7. xl + 311 pp. LCCN QA 76.76 H94 M82 1998.

Munnecke:1995:WWW

[Mun95] Tom Munnecke. World Wide Web and the demise of the clockwork universe. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Overviews/munnecke/abstract.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Murayama:1990:DTF

[Mur90] N. Murayama. DTP today and future. Journal of the Information Processing Society of Japan = Joho Shori, 31(11): 1483–1494, ???? 1990. CODEN JOSHA4. ISSN 0447-8053.

Murphy:1992:PCT

[Mur92a] H. J. Murphy, editor. Proceedings of Conference. Technology and Persons with Disabilities. California State Univ., Northridge, Northridge, CA, USA, ???? 1992. ISBN ???? LCCN ????

Murphy:1992:PSA

[Mur92b] H. J. Murphy, editor. Proceedings of the Seventh Annual Conference ‘Technology and Persons with Disabilities’. California State Univ., Northridge, Northridge, CA, USA, 1992. ISBN ???? LCCN ????

Marovic:1999:IVF

[MVKJ99] B. S. Marovic, D. J. Valentino, W. J. Karplus, and Z. Jovanovic. Interactive visualization of 3D fields and images on inexpensive workstations using VRML. In Westwood [Wes99], pages 219– 220. ISBN 90-5199-445-1, 4-274-90273-0. ISSN 0926-9630. LCCN W 26.5 M489-9 1999.

Melby:1992:TSI

[MW92] A. K. Melby and S. E. Wright.

TEI-TERM: an SGML-based interchange format for terminology files. In Anonymous [Ano92h], pages 19–37. ISBN 0-85142-295-0. LCCN ????

Miller:1997:PTE

[MWEK97] James E. Miller, Curt White, Carl Erickson, and Bruce Klein, editors. Proceedings of the Twenty-eighth SIGCSE Technical Symposium on Computer Science Education, February 1997, San Jose, CA, volume 29(1). ACM Press, New York, NY, USA, 1997. ISBN 0-89791-889-4. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA 76.27 S53 v.29 no.1 1997.

Wu:1997:CCS

[mWfC97] Tsan ming Wu and Yuan feng Cheng. Ching ching sung sung tzu hsiu hsueh HTML tsui hsin pan. Sung kang tien nao tu shu tzu liao ku fen yu hsien kung ssu, Taipei, Taiwan, chu pan edition, 1997. ISBN 957-22-2418-2 (paperback). various pp. LCCN ????

Mitchell:1999:RVC

[MWVE99] S. C. Mitchell, A. Wahle, C. Von Birgelen, and R. Erbel. Realtime visualization of coronary interventions using VRML. In Chen and Clough [CC99], pages 279–287. ISBN 0-8194-3132-X. ISSN 0277-786X (print), 1996-756X (electronic). LCCN RC78.7.D53 M45 1999.

Mazumdar:1998:ASS

[MYBP98] S. Mazumdar, G. Yuan, W. Bao, and J. Price. Adding semantics to SGML databases. Lecture Notes in Computer Science, 1375:563–574, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myka:1995:PPD

[Myk95] Andreas Myka. Putting paper documents in the World-Wide Web. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
myka/myka.html. A selected few

of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Murata:1996:RMT

[MYN96] T. Murata, H. Yamamoto, and J. Nagata. Realtime machine translation system for World Wide Web. Transactions of the Institute of Electronics, Information and Communication Engineers B-I, J79B-I(5):364–372, May 1996. CODEN DJBTES. ISSN 0915-1877.

Magalhaes:1996:CGI

[MZZ96] L. P. Magalhaes, J. A. Zuffo, and M. K. Zuffo, editors. Computer graphics and image processing: Computer graphics in Brazil: Symposium — 1996, volume 20(3) of Computers and Graphics 1996. Pergamon, New York, NY, USA, 1996. ISBN ???? ISSN 0097-8493 (print), 1873-7684 (electronic). LCCN ????

Nadrchal:1990:MMI

[Nad90] J. Nadrchal, editor. Man-Machine Interface in the Scientific Environment. Proceedings of the 8th European Summer School on Computing Techniques in Physics. Skalský Dvůr, Czechoslovakia, 19–28 September 1989, volume 61 of Computer Physics Communications. European Summer Schools on Computing Techniques in Physics, North-Holland Publishing Co.; Elsevier Science Publishers B. V., Amsterdam, The Netherlands, 1990. In addition to the papers presented at the colloquium, each paper is followed by a summary of the discussion about that paper.

Nadeau:1996:UIM

[Nad96] D. R. Nadeau. User interface metaphor in virtual reality using VRML. Behavior research methods, instruments, and computers, 28(2):170–173, May 1996. CODEN BRMCEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Nagy:1995:DIA

[Nag95] G. Nagy. Document image analysis: What is missing? In Braccini et al. [BDV95], pages 577–587. ISBN 3-540-60298-4. LCCN TA1632.I552 1995.

Nagasawa:1999:BVE

[Nag99] M. Nagasawa. Beyond VRML: Efficient 3-D data transfer for remote visualization. In Miyama et al. [MTH99], pages 415–422. ISBN 0-7923-5566-0. LCCN QB461.A1 A783 v.240.

Nakamura:1994:DJC

[Nak94] Y. Nakamura. Development of Japanese-language CD-ROM manuals using SGML. In Anonymous [Ano94r], pages 212–214. ISBN 0-914548-79-4, 0-914548-80-8. LCCN T 10.5 S63 1994.

Nadeau:1996:OPV

[NAM96] David R. Nadeau, Andrea L.

Ames, and John L. Moreland. Optimizing the performance of VRML worlds. Dr. Dobb’s Journal of Software Tools, 21(7):16– 18, 20, 22, 24, July 1996. CO-DEN DDJOEB. ISSN 1044-789X.

Nakagawa:1996:DCS

[NANK96] Y. Nakagawa, T. Akasaka, K. Nishimura, and M. Kurihara. Development of a CAI system using HTML under a computer network environment. Memoirs of the Hokkaido Institute of Technology, ???? 1996. CODEN HODKDL. ISSN 0385-0862.

NIST:1988:SGM

[Nat88] National Institute for Standards and Technology, Gaithersburg, MD, USA. Standard Generalized Markup Language (SGML), September 26, 1988. 3 pp.

NETG:1998:HF

[Nat98] National Education Training Group, Inc. HTML 4.0. fundamentals, 1998. Includes CD-ROM.

NETG:1999:HAT

[Nat99a] National Education Training Group, Inc. HTML 4.0. advanced topics, 1999. Includes CD-ROM.

NIST:1999:HMS

[Nat99b] National Institute of Standards and Technology (U. S.). A Hybrid Manufacturing System Modeling Environment Using VRML. Gaithersburg, MD, USA, March 26, 1999. Shipping List no: 99-0334-M.

LeNoel:1991:SPC

[NB91] Bernard Le Noel and Michel Biezunski. SGML-ODA: présentation des concepts et comparaison fonctionnelle. AFNOR technique. AFNOR, Paris, France, 1991. ISBN 2-12-488011-X. ix + 87 pp. LCCN ????

Nelson:1995:WWW

[NB95] Michael L. Nelson and David J. Bianco. The World Wide Web and technology transfer at NASA Langley. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
NationWideSites/nelson/paper.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Nordin:1993:RSG

[NBM93] B. Nordin, D. T. Barnard, and I. A. Macleod. A review of the Standard Generalized Markup Language (SGML). Computer Standards and Interfaces, 15(1):5–19, May 1993. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Nixon:1998:ILP

[NDM98] Carol Nixon, M. Heide Dengler, and Mare L. McHenry, editors. Internet librarian ’98: Proceedings of the second Internet Librarian Conference, Monterey, CA, November 1–5, 1998. Information Today, Medford, NJ, USA, 1998. ISBN 1-57387-067-6. LCCN Z674.82.I59 I585 1998.

Negishi:1992:RDF

[Neg92] M. Negishi. Recent development in full text database applications. Journal of the Information Processing Society of Japan = Joho Shori, 33(4):413–420, ???? 1992. CODEN JOSHA4. ISSN 0447-8053.

Negishi:1995:SBE

[Neg95] M. Negishi. SGML based editing system for full text databases of scientific journals. International information, communication and education, 14(1):94–??, 1995. CODEN IICEE2. ISSN 0970-1850.

Neou:1996:HCJb

[Neo96a] Vivian Neou. HTML 3.0 CD with JavaScript. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, September 1996. ISBN 0-13-243353-2. ???? pp. LCCN ???? US$44.95.

Neou:1996:HCJc

[Neo96b] Vivian Neou. HTML 3.0 CD with JavaScript. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, September 1996. ISBN 0-13-243353-2. ???? pp. LCCN ???? US$44.95.

Neou:1997:HCJa

[Neo97] Vivian Neou. HTML 3.0 CD JavaScript for MacIntosh: CD-ROM With Guide. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, August 1997. ISBN 0-13-243353-2. ???? pp. LCCN ???? US$44.95.

Neou:1998:HCJ

[Neo98] Vivian Neou. HTML 4.0 CD with JavaScript for Windows 95, 98 and NT. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-095783-6. xi + 619 pp. LCCN QA76.76.H94N435 1999. US$49.95. Includes CD-ROM.

Neou:1999:HCJ

[Neo99] Vivian Neou. HTML 4.0 CD with JavaScript for Windows 95, 98 and NT. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-095783-6. xi + 619 pp. LCCN QA76.76.H94N435 1999.

Nerbonne:1997:LDP

[Ner97] John A. Nerbonne, editor. Linguistic databases: Papers presented at a conference held Mar. 23–24, 1995, University of Groningen, volume 77 of CSLI lecture notes. CSLI Publications, Stanford, CA, USA, 1997. ISBN 1-57586-093-7 (hardback), 1-57586-092-9 (paperback). LCCN P128.D37 L56 1997.

Neubauer:1994:QID

[Neu94] Wolfram Neubauer, editor.

Qualität und Information: Deutscher Dokumentartag 1993, Friedrich-Schiller-Universität, Jena, 28. bis 30. September 1993: Proceedings. Deutsche Gesellschaft für Dokumentationrs, Frankfurt am Main, Germany, 1994. ISBN 3-925474-23-4. ISSN 0721-1058. LCCN Z674.5.G28D48 1993.

Neubauer:1996:EPB

[Neu96] Karl Wilhelm Neubauer, editor.

Elektronisches Publizieren und Bibliotheken (Electronic Publishing and Libraries). Papers presented at the Bielefelder Kolloquium, sponsored by the Bibliothek der Universitat Bielefeld, the British Council, and the Buchhandler-Vereinigung GmbH, February 1996, volume 65 of Zeitschrift fur Bibliothekswesen und Bibliographie. Vittorio Klostermann, Frankfurt am Main, Germany, 1996. ISBN 3-465-02885-6. ISSN 0514-6364. LCCN Z678.93.E5 E423 1996.

Neugass:1997:EWP

[Neu97a] Jonah Neugass. Easy Web publishing with HTML 3.2. Que Corporation, Indianapolis, IN, USA, March 1997. ISBN 0-7897-1143-5. ix + 172 pp. LCCN QA76.76.H94N48 1997. US$24.99.

Neumann:1997:USC

[Neu97b] Andreas Neumann. Unambiguity of SGML content modelspushdown automata revisited. Forschungsbericht, Nr. 97-05; Forschungsbericht (Universitat Trier. Mathematik/Informatik); Nr. 97-05 ??, Universitát Trier, Trier, Germany, 1997. 22 + v pp.

Neugass:1998:SSH

[Neu98] Jonah Neugass. Step by step hsueh hui HTML chih tso wang yeh. Beginners’; 3. Tien nao jen wen hua shih yeh ku fen yu hsien kung ssu, Tai-pei shih, tseng ting pan. edition, 1998. ISBN 957-8309-09-0 (paperback). 194 pp. LCCN ????

Newcomb:1990:ECM

[New90] S. R. Newcomb. Explanatory cover material for section 7.2 of X3V1.8M/SD-7, fifth draft (music description language). In Moline et al. [MBB90], pages 179–188. ISBN ???? LCCN ????

Newmann:1992:MVI

[New92] Scott Newmann. Multi-vendor interoperability through SQL access. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(2):414, June 1992. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Newcomb:1995:SMIa

[New95a] Steven R. Newcomb. Standards: Multimedia interchange using SGML/HyTime. Part I: Structures. IEEE MultiMedia, 2(2):86–89, Summer 1995. CO-DEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

URL http://dlib.computer.
org/mu/books/mu1995/pdf/u20086.
pdf.

Newcomb:1995:SMIb

[New95b] Steven R. Newcomb. Standards: Multimedia interchange using SGML/HyTime. Part II: Principles and applications. IEEE MultiMedia, 2 (3):60–64, Fall 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

URL http://dlib.computer.
org/mu/books/mu1995/pdf/u30060.
pdf.

Newing:1996:WEY

[New96] R. Newing. Web enabling your database. Conspectus, pages 12–13, July 1996. CODEN CONSF8. ISSN 1351-0908.

Nebesh:1996:LUA

[NF96] B. Nebesh and M. Feldman.

Learning to use Ada 95 components using HTML linking. In ACM [ACM96c], pages 207–210. ISBN 0-89791-808-8. LCCN QA 76.73 A35 T75 1996.

Nebesh:1997:UHL

[NF97] B. Nebesh and M. B. Feldman. Using HTML linking to help novice programmers to reuse components. In Anonymous [Ano97z], pages 315–319. ISBN 0-262-51095-2. LCCN Q334 .N37 1997.

North:1999:STY

[NH99] Simon North and Paul Hermans.

SAMS teach yourself XML in 21 days. SAMS Publishing, Indianapolis, IN, USA, 1999. ISBN 1-57521-396-6. xiii + 580 pp. LCCN QA76.76.H94N67 1999.

Nickerson:1992:WWH

[Nic92] G. Nickerson. WorldWide Web: hypertext from CERN. Computers in Libraries, 12(11):75– 77, December 1992. CODEN CPLIE8. ISSN 1041-7915.

Nicol:1995:DWI

[Nic95] G. T. Nicol. Dyna Web: Integrating large SGML repositories and the WWW. In O’Reilly and Associates and Web Consortium (W3C) [WWW95], pages 549– 558. ISBN 1-56592-169-0. ISSN 1085-2301. LCCN TK5105.888 .I68 1995. US$39.95. URL http:

//www.ora.com/gnn/bus/ora/
item/wj1.html. The World

Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Nicchiotti:1997:SRV

[Nic97] G. Nicchiotti. A system for 3D reconstruction and VRML conversion for virtual museums applications. In IEEE [IEE97a], pages 421–424. ISBN 0-7803-4137-6, 0-7803-4138-4, 0-7803-4139-2. LCCN TK 5102.9 I58 1997.

Nielsen:1999:VPR

[Nie99] A. Nielsen. VRML programs for room ventilation applications. In Anonymous [Ano99b], page ?? ISBN ???? ISSN 1395-7953. LCCN ????

Nilsen:1996:VRI

[NJDF96] C. Nilsen, J. Joiner, J. Damico, and J. Friesen. Virtual realtime inspection of nuclear material via VRML and secure Web pages. In Anonymous [Ano96x], pages 1250–1254. ISBN ???? LCCN ????

Neith:1999:DSM

[NK99] M. Neith and O. E. Kiselyov.

Disseminating and storing of meteorological data with extensible markup language (XML). In Anonymous [Ano99a], pages 554–557. ISBN ???? LCCN ????

Newcomb:1991:HHD

[NKN91a] Steven R. Newcomb, Neill A.

Kipp, and Victoria T. Newcomb. The “HyTime”: hypermedia/time-based document structuring language. Communications of the ACM, 34(11): 67–83, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/
125495.html.

Newcomb:1991:HTB

[NKN91b] Steven R. Newcomb, Neill A.

Kipp, and Victoria T. Newcomb. The “HyTime”: hypermedia/time-based document structuring language. Communications of the ACM, 34(11): 67–83, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/
125495.html.

Nielsen:1995:TUL

[NL95] H. F. Nielsen and H. W. Lie. Towards a uniform library of common code. A presentation of the CERN World-Wide Web library. Computer Networks and ISDN Systems, 28(1–2):13–??, December 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Nadeau:1995:VRB

[NM95a] D. R. Nadeau and J. L. Moreland. The Virtual Reality Behavior System (VRBS): A behavior protocol for VRML. In Nadeau and Moreland [NM95b], pages 53–61. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Nadeau:1995:SVR

[NM95b] David R. Nadeau and John L. Moreland, editors. 1995 Symposium on the Virtual Reality Modeling Language, VRML ’95, San Diego, California, December 14–15, 1995. ACM Press, New York, NY, USA, 1995. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Nebel:1995:RFF

[NM95c] E. Nebel and L. Masinter.

RFC 1867: Form-based file upload in HTML, November

7, 1995. URL ftp://ftp.
internic.net/rfc/rfc1867.
txt;   http://www.math.utah.
edu/pub/rfc/rfc1867.txt.

Newcomb:1992:SBI

[NN92] S. R. Newcomb and V. T. Newcomb. Some background information about HyTime (hypermedia language). Journal of the Institute of Image Electronics Engineers of Japan, 21(5): 459–467, October 1992. ISSN 0285-9831.

Nishioka:1997:GEV

[NN97] D. Nishioka and M. Nagasawa.

GhostSpace: Efficient VRML browsing on ATM. In Arabnia and Ahmed [AA97], pages 16– 20. ISBN 0-9648666-9-2. LCCN ????

Noreault:1996:USD

[Nor96] T. R. Noreault. Use of SGML in digital libraries. In Anonymous [Ano96n], pages 81–87. ISBN 89-87028-00-3. LCCN ????

Normark:1999:PWW

[Nør99]  Kurt Nørmark.  Program-

ming World Wide Web pages in Scheme. ACM SIGPLAN Notices, 34(12):37–46, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic).  URL http://ftp.
informatik.rwth-aachen.de/

dblp/db/indices/a-tree/n/
N=oslash=rmark:Kurt.html;
 http://www.cs.auc.dk/~normark/

laml/. Describes a literate programming environment, LAML (Lisp Abstracted Markup Language), for HTML dynamically generated from Scheme programs.

Notess:1996:CLK

[Not96] G. R. Notess. Catch the lightning: keeping current with the Net. Online, 20(3):86–88, May– June 1996. CODEN ONLIDN. ISSN 0146-5422.

Naraghi-Pour:1996:PTS

[NP96] Mort Naraghi-Pour, editor. Proceedings of the Twenty-Eighth Southeastern Symposium on System Theory: March 31– April 2, 1996, Baton Rouge, Louisiana. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7352-4. LCCN TA 168 S727p 1996. IEEE Catalog No. 96TB100023.

Neches:1995:OPI

[NPPR95] Anna-Lena Neches, Cesar Pina, Paul Postel, and Craig Milo Rogers. Online procurement at ISI: Transitioning to the Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
CorInfSys/neches/broker.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Neou:1996:HCJa

[NR96a] Vivian Neou and Mimi Recker.

HTML 3.0 CD With JavaScript for the Mac and Power Mac. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, June 1, 1996. ISBN 0-13-439399-6. xix + 441 pp. LCCN QA76.76.H94 N43 1996. US$44.95.

Neou:1996:HCI

[NR96b] Vivian Neou and Mimi Recker.

HTML CD: an Internet publishing toolkit: for Windows. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-232331-1. xxxviii + 311 pp. LCCN QA76.76.H94 N46 1996. US$39.95, CDN$54.00.

Nica:1996:USD

[NR96c] A. Nica and E. A. Rundensteiner. Uniform structured document handling using a constraint-based object approach. In Adam et al. [ABHY96], pages 83–101. ISBN 3-540-61410-9. LCCN Z699 .D532 1996.

Neou:1997:HCJb

[NR97] Vivian Neou and Mimi Recker.

HTML 3.2 CD With JavaScript for Windows 95. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-270125-1. xxi + 439 pp. LCCN QA76.76.H94N434 1997. US$44.95.

Nguyen:1995:UDS

[NS95a] T. L. Nguyen and T. A. Sharkey. Using a database and SGML to build singlesource documents. In Anonymous [Ano95-37], pages 291–?? ISBN 0-914548-86-7, 0-914548-87-5. LCCN T 11 S63 1995.

Nielsen:1995:SUIb

[NS95b] J. Nielsen and D. Sano. Sun-Web: User interface design for Sun Microsystem’s internal Web. Computer Networks and ISDN Systems, 28(1–2): 179–??, December 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Nielsen:1995:SUIa

[NS95c] Jakob Nielsen and Darrell Sano. SunWeb: User interface design for Sun Microsystem’s internal Web. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

                                        http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/HCI/
                                        nielsen/sunweb.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Nguyen:1996:ARD

[NS96] T. Nguyen and V. Srinivasan. Accessing relational databases from the World Wide Web. In Jagadish and Mumick [JM96], pages 529–540. CODEN SRECD8. ISBN 0-89791-794-4.

ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.25 no.2 1996.

Nikolaou:1998:RAT

[NS98] Christos Nikolaou and Constantine Stephanidis, editors. Research and advanced technology for digital libraries: second European conference, ECDL ’98, Heraklion, Crete, Cyprus, September 21–23, 1998: proceedings, volume 1513 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. ISBN 3-540-65101-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ZA4082.E85E25 1998.

Nakano:1991:IBT

[NTI91] Y. Nakano, H. Tanigawa, and Y. Imai. An interface/application builder with transmission control facility for SGML document databases. In Anonymous [Ano91c], pages 41– 46. ISBN ???? LCCN ????

Noreault:1989:TFI

[NTR89] T. Noreault, G. K. M. Tobin, and R. M. Robin. Towards full information database display. In Salvendy and Smith [SS89], pages 180–186. ISBN 0-444-88078-X. LCCN QA76.9.H85 I59 1989 v.2.

Nicholas:1992:ISO

[NW92] C. K. Nicholas and L. A.

Welsch. On the interchangeability of SGML and ODA. Electronic Publishing—Origination, Dissemination, and Design, 5 (3):105–130, September 1992. CODEN EPODEU. ISSN 0894-3982.

Nicholas:1997:PDP

[NW97] Charles K. Nicholas and Derick Wood, editors. Principles of document processing: Third International Workshop, PODP’96, Palo Alto, California, USA, September 23, 1996: proceedings, volume 1293 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63620-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .P63 1996.

Oogane:1998:IMA

[OA98] T. Oogane and C. Asakawa. An interactive method for accessing tables in HTML. In ACM [ACM98], pages 126– 128. ISBN 1-58113-020-1. LCCN ???? URL http://

                                        www.acm.org/pubs/contents/
                                        proceedings/assets/274497/.

ACM order number 444980.

OrdieresMere:1997:IVW

[OBO97] J. Ordieres Mere, A. Bello Garcia, and F. Ortega Fernandez. Impact of VRML and WWW technology in a construction project life cycle. In Riitahuhta [Rii97], pages 261–264. ISBN 951-722-788-4. LCCN ????

OConnor:1992:D

[O’C92a] Mary Ann O’Connor. DataBasics. CD-ROM professional, 5(6):130–131, November 1992. CODEN CRPFEX. ISSN 1049-0833.

OConnor:1992:DMS

[O’C92b] Mary Ann O’Connor. DataBasics: Markup, SGML, and hypertext for full-text databases: Part I. CD-ROM professional, 5(3):112–??, May 1, 1992. CO-DEN CRPFEX. ISSN 1049-0833.

OConnor:1992:MSH

[O’C92c] Mary Ann O’Connor. Markup, SGML, and hypertext for fulltext databases, part II. CD-ROM professional, 5(4):123– 125, July 1992. CODEN CRPFEX. ISSN 1049-0833.

OConnor:1997:WSI

[O’C97] Hugh O’Connor. Web sites for information on intranets, HMTL/SGML/JAVA/VRML coding and web page/sites design. American Association of Retired Persons, Washington, DC, USA, version 2 edition, February 20, 1997. 10 pp.

OCLC:1988:SKG

[OCL88] OCLC. SGML keying guidelines. OCLC Reference Services, Dublin, OH, USA, revised 1.0. edition, 1988. 12–?? pp.

Oliver:1997:H

[OH97a] Dick Oliver and Molly E.

Holzschlag. HTML 4. Le Grand Poche. Simon & Schuster Macmillan (France), Paris, France, 1997. ISBN 2-7440-0366-2. ix + 309 pp. LCCN ????

Oliver:1997:STY

[OH97b] Dick Oliver and Molly E. Holzschlag. Sam’s Teach yourself HTML 4 in 24 hours. Sams.net Pub., Indianapolis, IN, USA, second edition, 1997. ISBN 1-57521-366-4 (paperback). xxviii + 405 pp. LCCN QA76.76.H94O4524 1997.

Oliver:1997:TYHb

[OH97c] Dick Oliver and Molly E. Holzschlag. Teach yourself HTML 4 in 24 hours. Sams.net Pub., Indianapolis, IN, USA, second edition, 1997. ISBN 1-57521-366-4 (paperback). xxvi + 405 pp. LCCN QA76.76.H94O4524 1997.

Oezsu:1997:OSH

[OISEM97] M. T. Oezsu, P. Iglinski, D. Szafron, and S. El-Medani. An object-oriented SGML/ HYTIME compliant multimedia database management system. In Hollan and Foley [HF97], pages 239–250. ISBN 0-201-32232-3 (Addison-Wesley), 0-89791-991-2 (ACM). LCCN QA76.575.A36 1997.

OKane:1995:IRA

[O’K95] K. C. O’Kane. Information retrieval by automatically generated HTML documents in large document collections. In Anonymous [Ano95v], pages 94–95. ISBN ???? LCCN ????

Okerson:1993:SPE

[Oke93] Ann Okerson, editor. Scholarly publishing on the electronic networks: the new generation: visions and opportunities in notfor-profit publishing: proceedings of the second symposium, December 5–8, 1992, the Washington Vista Hotel, Washington, DC, USA. Office of Scientific and Academic Publishing, Association of Research Libraries, Washington, DC, USA, 1993. ISBN 0-918006-61-9. LCCN Z286.E43 S36 1993.

Okerson:1995:FPP

[Oke95] A. Okerson, editor. Filling the pipeline and paying the piper: proceedings of the fourth symposium: November 5–7, 1994, the Washington Vista Hotel, Washington, DC, USA. Association of Research Libraries, Washington, DC 20036, USA, 1995. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Oklahoma:1997:HHM

[Okl97] Oklahoma Dept.of Libraries.

HTML, how to make library Web pages. Oklahoma Dept. of Libraries, Oklahoma City, OK, USA, May 2, 1997. 49 pp.

Ogawa:1992:RDF

[OKT92] R. Ogawa, Y. Kikuchi, and K. Takahashi. Recent developments in full text database technologies. Journal of the Information Processing Society of Japan = Joho Shori, 33(4):

404–412, ???? 1992. CODEN JOSHA4. ISSN 0447-8053.

Olson:1997:ODS

[OL97] M. R. Olson and B. S. Lee. Object databases for SGML document management. In El-Rewini et al. [ERPNS97a], pages 39– 56. ISBN 0-8186-7734-1, 0-8186-7749-X. ISSN 1060-3425. LCCN ????

Oliver:1997:TYHa

[Oli97a] Dick Oliver. Teach yourself HTML 3.2 in 24 hours. SAMS Publishing, Indianapolis, IN, USA, March 1997. ISBN 1-57521-235-8. xxii + 386 pp. LCCN QA76.76.H94O45 1997. US$19.99.

Oliver:1997:HTT

[Oli97b] William G. Oliver. HTML tutorial texts, forms, lists and tables, 1997.

Oliver:1998:H

[Oli98a] Dick Oliver. HTML 4. Le tout en poche. MacMillan Publishing Company, New York, NY, USA, 1998. ISBN 2-7440-0507-X. x + 349 pp. LCCN ????

Oliver:1998:STY

[Oli98b] Dick Oliver. Sams teach yourself HTML 4 in 24 hours. SAMS Publishing, Indianapolis, IN, USA, third edition, 1998. ISBN 0-672-31369-3. xxi + 471 pp. LCCN QA76.76.H94O4524 1998.

Oliver:1999:STY

[Oli99] Dick Oliver. Sams teach yourself HTML in 24 hours. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fourth edition, 1999. ISBN 0-672-31346-4. xv + 428 pp. LCCN TK5105.888.S624 1999.

Olympia:1996:CFA

[Oly96] P. L. Olympia. Cold Fusion 1.5, Autobahn. DBMS, 9(8):33– 34, 36, 38, July 1996. CODEN DBMSEO. ISSN 1041-5173.

Oldham:1996:TIR

[OM96] J. D. Oldham and V. W. Marek.

Toward intelligent representation of database content. In Ras and Michalewicz [RM96a], pages 274–284. ISBN 3-540-61286-6. LCCN QA76.76.E95I582 1996.

Osawa:1997:APD

[OMY97] N. Osawa, H. Morita, and T. Yuba. Animation for performance debugging of parallel computing systems. In Spencer and Brutzman [SB97], pages 101–108. ISBN 0-89791-886-X. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/253437/.

ACM order number 434973.

OnWordPressDevelopmentStaff:1995:MGS

[OnW95] OnWord Press Development Staff. Manager’s Guide to SGML. High Mountain Press, Incorporated, Santa Fe, NM, USA, 1995. ISBN 1-56690-049-2. 200 pp. LCCN ???? US$49.95.

OppoSiteSoftware:1997:HPT

[Opp97]  OppoSite Software.  HTML

power tools, 1997. Includes CD-ROM.

Oostendorp:1995:TIW

[OPW95] K. A. Oostendorp, W. F. Punch, and R. W. Wiggins. A tool for individualizing the Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Agents/oostendorp/oostendorp.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Onifer-Reino:1996:ISI

[OR96] J. Onifer-Reino. Implementing SGML: Issues and strategies for the beginner. In Anonymous [Ano96b], pages 397–398. ISBN 0-914548-89-1, 0-914548-90-5. LCCN ????

OReilly:1997:WDL

[O’R97] O’Reilly & Associates, Inc. Web Developer’s Library on the Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-363-4. ???? pp. LCCN ???? US$59.95.

URL http://www.ora.com/
catalog/webrlw/;   http:/
/www.oreilly.com/catalog/
webrlw.

Oliver:1995:DHN

[ORCK95] Katie S. Oliver, Christine Remenyik, Alexander Crosby, and Jim Kazmer. Dynamic hypertext navigation and display management. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
MedTrack/oliver/OliverRem.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Ore:1988:TDT

[Ore88] E. S. Ore. Text as data: text as end product. Literary and Linguistic Computing, 3(3):162– 165, ???? 1988. CODEN LL-COEI. ISSN 0268-1145.

NISO:1991:EMP

[Org91] National Information Standards Organization. Electronic Manuscript Preparation and Markup: American National Standard for Electronic Manuscript Preparation and Markup. Transaction Publishers, Rutgers University, New Brunswick, NJ 08903, USA, 1991. ISBN 0-88738-945-7. xv + 167 pp. LCCN Z286.E43 E428 1991. ANSI/NISO Z39.59-1988. Approved December 1, 1988, by American National Standards Institute. Developed by The National Information Standards Organization.

Orlando:1993:EDI

[Orl93] B. Orlando. Electronic data interchange (EDI) in a CALS/CE environment: a user’s view. In Anonymous [Ano93c], pages 1– 50. ISBN ???? LCCN ????

Ochiai:1997:SPH

[OSE97] N. Ochiai, Y. Sota, and H. Ezumi. Self-study program on HTML browser — application to clinical nursing general remarks course. In Gerdin-Jelger et al. [GJTW97], pages 360–363. ISBN 90-5199-362-5, 4-274-90184-X. ISSN 0926-9630. LCCN ????

Ozsu:1996:DMS

[OSEM^{+}96] M. T. Ozsu, D. Szafron, G. El-Medani, S. El-Medani, P. Iglinski, M. Schoene, and C. Vittal. Database management support for a news-on-demand application. Proceedings of the SPIE — The International Society for Optical Engineering, 2617:24–32, ???? 1996. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Ozsu:1995:OMD

[OSEMV95] M. T. Ozsu, D. Szafron, G. El-Medani, and C. Vittal. An object-oriented multimedia database system for a news-on-demand application. Multimedia Systems, 3(5-6):182–203, ???? 1995. CODEN MUSYEW. ISSN 0942-4962.

Ohta:1994:DQS

[OST94a] M. Ohta, O. Sato, and K. Tazawa. Development of Q and A system-technology for information sharing system: FIND2. Fujitsu, 45(5):436–442, ???? 1994. CODEN FUJTAR. ISSN 0016-2515.

Oster:1994:IPS

[Ost94b] J. W. Oster. An introduction to the principles of an SGML application. In Anonymous [Ano94r], pages 303–?? ISBN 0-914548-79-4, 0-914548-80-8. LCCN T 10.5 S63 1994.

Ohtsuki:1997:SDF

[OSYM97] M. Ohtsuki, J. Segawa, N. Yoshida, and A. Makinouchi. Structured document framework for design patterns based on SGML. In IEEE [IEE97g], pages 320–323. ISBN 0-8186-8106-3, 0-8186-8107-1, 0-8186-8105-5. ISSN 0730-3157. LCCN QA76.6 .C62 1997. IEEE Catalog Number 97CB36112. IEEE Computer Society Order Number PR08105.

Oliver:1997:CRW

[OuP97] Dick Oliver and Chin u Pak.

Choboja rul wihan HTML 3.2 24 sigan wansong. Inpo Buk, Soul, Korea, 1997. ISBN 89-8054-310-7 (paperback). 360 pp. LCCN ????

Oliver:1995:LLA

[OW95] R. Oliver and M. Wild, editors.

Learning without limits: Australian computers in education conference — July 1995, Hyatt, Australia, volume 2. ECAWA, ????, 1995. ISBN 1-875534-08-3, 1-875534-10-5. LCCN ????

Owen:1995:IWW

[Owe95]  G. Scott Owen.  Integrat-

ing World Wide Web technology into courses in computer graphics and scientific visualization. Computer Graphics, 29(3):12–14, August 1995. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Owen:1996:IWW

[Owe96] G. Scott Owen. Integrating World Wide Web technology into undergraduate education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (??):101–103, ???? 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ozer:1996:AW

[Oze96] Jan Ozer. Audio … Video … Live from the WEB! PC Magazine, 15(6):100–??, March 26, 1996. CODEN PCMGEP. ISSN 0888-8507.

Palme:1998:SFP

[Pal98] J. Palme. From start to finish — publishing in SGML. In Czerminski [Cze98], pages 108–117. ISBN 83-86230-35-5. LCCN ????

Paoli:1995:CWN

[Pao95] Jean Paoli. Cooperative work on the network: edit the WWW!

Computer Networks and ISDN Systems, 27(6):841–847, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/
store/comnet/sub/1995/27/
6/1458.pdf.

Paoli:1996:EWT

[Pao96] Jean Paoli. Extending the Web’s tag set using SGML: authoring new tags with Grif Symposia. Computer Networks and ISDN Systems, 28(7–11): 1095–1104, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=
1625.

Paquel:1994:DIW

[Paq94] N. Paquel. The document interface to a world of objects. Objects in Europe, 1(2):11–12, 14– 15, Spring 1994.

Parsons:1995:NWI

[Par95] Michael Parsons. News: Web innovation: SGI, Sun and Netscape boost VRML, Java. IDG International News Service, page ??, December 05, 1995.

Pardi:1999:XA

[Par99] William J. Pardi. XML in action. Microsoft Press, Bellevue, WA, USA, 1999. ISBN 0-7356-0562-9. xii + 329 pp. LCCN QA76.76.H94 P374 1999.

Patterson:1997:UH

[Pat97] Lois Patterson. Using HTML 3.2. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1450-7. xvii + 386 pp. LCCN QA76.76.H94 U85 1998. US$19.99.

Patterson:1998:IH

[Pat98] Lois R. Patterson. Ispolzovanie HTML 4. Viliams, Kiev, Ukraine, third edition, 1998. ISBN 966-7416-08-9. 382 pp. LCCN ????

Paulson:1995:BYO

[Pau95] Linus Paulson. Build Your Own Web Site for Almost Free: Create a World Wide Web Site with HTML, Perl, and Linux. Waite Group Press, Corte Madera, CA, USA, September 1995. ISBN 1-57169-053-0. 700 pp. LCCN ???? US$36.95.

Pelletier:1996:SPD

[PAV96] S.-J. Pelletier, J.-F. Arcand, and J. Velissarios. STEALTH: a personal digital assistant for information filtering. In Anonymous [Ano96-31], pages 455– 474. ISBN ???? LCCN ????

Polk:1988:EAM

[PB88] W. T. Polk and Lawrence E. Bassham. Expert assistance for manipulating of SGML document type definitions. In Solem [Sol88], pages 79–84. ISBN 0-89791-291-8. LCCN QA 76.9 T48 A25 1988.

Pirouz:1998:HKW

[PB98] Raymond Pirouz and Lucie Blazkova. HTML: kouzla na webu. Unis, Brno, 1998. ISBN 80-86097-17-X. 294 pp. LCCN ????

Preis:1996:VEL

[PBB^{+}96] K. Preis, I. Bardi, O. Biro, R. Hoschek, M. Mayr, and I. Ticar. A virtual electromagnetic laboratory for the classroom and the WWW. In IEEE [IEE96g], page 277. ISBN ???? LCCN ????

Pimentel:1997:TRM

[PBFT97] M. G. C. Pimentel, L. Baldochi, F. Fagundes, and C. A. C. Teixeira. Temporal relations in multimedia objects: WWW presentation from HyTime specification. In IEEE [IEE97c], pages 84–91. ISBN 0-8186-7916-6, 0-8186-7918-2 (microfiche). LCCN QA76.575 .I33 1997. IEEE Computer Society order number PR07916. IEEE order plan catalog number 97TB100116.

Pardue:1997:AME

[PC97] J. H. Pardue and D. Chatterjee.

An approach to measuring extent of organizational use of Web functionalities: A content analysis of HTML tags. In Gupta [Gup97], pages 233–234. ISBN ???? LCCN ????

Pajon:1995:BEL

[PCL^{+}95] J. L. Pajon, Y. Collenot, X. Lhomme, N. Tsingos, F. Sillion, P. Guilloteau, P. Vuyslteker, G. Grillon, and D. David. Building and exploiting levels of detail: An overview and some VRML experiments. In Nadeau and Moreland [NM95b], pages 117–122. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Prah:1995:MCD

[PD95] Chris Prah and Diane DiGiovanni. Mosaic as corporate data collector and dispenser. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
CorInfSys/prah/prah.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Perret:1996:MAP

[PD96] Stéphane Perret and Andrzej Duda. Mobile assistant programming for efficient information access on the WWW. Computer Networks and ISDN Systems, 28(7–11):1373–1383, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=
1609.

Pearce:1995:ASC

[PE95] J. M. Pearce and A. Ellis, editors. Australian Society for Computers in Learning in Tertiary Education: ASCILITE 95: Annual conference; 12th — 1995 Dec, ASCILITE -Conference Proceedings- 1995; 12th. The University of Melbourne, Faculty of Science, Melbourne, Victoria, Australia, 1995. ISBN 0-7325-1220-4. LCCN ????

Pelton:1997:EGU

[Pel97a] L. F. Pelton. Exploring geometry by using VRML to build 3-D worlds. In Anonymous [Ano97v], pages 144–?? ISBN 1-56484-112-X. LCCN ????

Pelton:1997:IVT

[Pel97b] T. Pelton. An introduction to VRML: A tool for building 3-D worlds. In Anonymous [Ano97v], pages 341–346. ISBN 1-56484-112-X. LCCN ????

Perron:1995:LWC

[Per95] Daniel Perron. Learning on the WWW: A case study. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Educ/perron/perron.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Perry:1996:ASH

[Per96]  L. A. Perry.  The ABC’s

of SGML/HTML: Understanding the concept of CALS and SGML. In Anonymous [Ano96b], pages 399–401. ISBN 0-914548-89-1, 0-914548-90-5. LCCN ????

Pesce:1995:VBB

[Pes95] Mark Pesce. VRML browsing and building cyberspace: the definitive resource for VRML technology. New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-498-8. xxxiii + 424 pp. LCCN TK5105.88 .P47 1995. Foreword by Tim Berners-Lee. Includes CD-ROM.

Pesce:1996:LV

[Pes96a] Mark Pesce. Learning VRML 2.0. New Riders Publishing, Carmel, IN, USA, December 1996. ISBN 1-56205-675-1. ???? pp. LCCN ???? US$39.99.

Pesce:1996:VNS

[Pes96b] Mark Pesce. VRML 2.0 the Next Step in Cyberspace. New Riders Publishing, Carmel, IN, USA, September 1996. ISBN 1-56205-675-1. ???? pp. LCCN ???? US$40.00.

Pesce:1996:VFT

[Pes96c] Mark Pesce. VRML flying through the Web. New Riders Publishing, Carmel, IN, USA, 1996. ISBN 1-56205-521-6. xii + 367 pp. LCCN T385.P476 1996. US$35.00, CDN$47.95, UK£32.99. Includes CD-ROM.

Petersen:1990:CIN

[Pet90] H. Petersen. Conception and implementation of a network service for the processing of documents. Informacne Systemy, 19 (4):321–337, August 1990. CO-DEN ISYSDJ. ISSN 0323-1984.

Petruzzelli:1995:CTI

[Pet95] Jim Petruzzelli. Critical thinking with the interactive text: Decentering textual authority with Mosaic. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/
petruzzelli/petruzzelli.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Petrovsky:1998:DHA

[Pet98a] Michele Petrovsky. Dynamic HTML in action. Osborne/ McGraw-Hill, Berkeley, CA, USA, 1998. ISBN 0-07-882437-0. xxvii + 441 pp. LCCN QA76.76.H94P485 1998.

Petrovsky:1998:MDH

[Pet98b] Michele Petrovsky. Manual de dynamic HTML. McGraw-Hill, New York, NY, USA, 1998. ISBN 84-481-1481-7. xxiv + 365 pp. LCCN ????

Petrovsky:1998:ICC

[Pet98c] Michele Jo Petrovsky. Implementing CDF channels. Mc-Graw-Hill, New York, NY, USA, 1998. ISBN 0-07-049887-3. xviii + 267 pp. LCCN TK5105.887 .P48 1998.

Pearlstein:1995:WCE

[PF95] Robert Pearlstein and Peter C. FitzGerald. The Web as a computational engine for chemistry and molecular biology. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        BioChem/pearlstein/fitzgerald.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Pfaffenberger:1997:DH

[Pfa97a] Bryan Pfaffenberger. Discovering HTML 4. AP Professional, Boston, MA, USA, May 1997. ISBN 0-12-553167-2. xx + 419 pp. LCCN QA76.76.H94P488 1998. US$34.95, CDN$48.95.

Pfaffenberger:1997:WPX

[Pfa97b] Bryan Pfaffenberger. Web Publishing with XML in Six Easy Steps. AP Professional, Boston, MA, USA, June 1997. ISBN 0-12-553166-4. xxviii + 364 pp. LCCN QA76.76.H94P52 1998. US$34.95, CDN$48.95.

Pfaffenberger:1998:DH

[Pfa98a] Bryan Pfaffenberger. Discovering HTML 4. AP Professional, Boston, MA, USA, 1998. ISBN 0-12-553167-2 (paperback). xx + 419 pp. LCCN QA76.76.H94P488 1998.

Pfaffenberger:1998:WPX

[Pfa98b] Bryan Pfaffenberger. Web publishing with XML in six easy steps. AP Professional, Boston, MA, USA, 1998. ISBN 0-12-553166-4 (paper ). xxviii + 364 pp. LCCN QA76.76.H94P52 1998.

Pfaffenberger:1998:XWP

[Pfa98c] Bryan Pfaffenberger. XML for Web publishing in six easy steps. AP Professional, Boston, MA, USA, 1998. ISBN 0-12-553166-4 (paperback). xxviii + 364 pp. LCCN QA76.76.H94P52 1998.

Pfaffenberger:1999:WPX

[Pfa99] Bryan Pfaffenberger. Web publishing with XML in six easy steps. AP Professional, Boston, MA, USA, 1999. ISBN 0-12-553166-4 (paper). xxviii + 364 pp. LCCN QA76.76.H94 P52 1999.

Pfeifer:1995:SSD

[PFH95] U. Pfeifer, N. Fuhr, and T. Huynh. Searching structured documents with the enhanced retrieval functionality of freeWAIS-sf and SFgate. Computer Networks and ISDN Systems, 27(6):1027–1036, April 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Pfaffenberger:1998:HB

[PG98] Bryan Pfaffenberger and Alexis D.

Gutzman. HTML 4 bible. I D G Books Worldwide, Indianapolis, IN, USA, 1998. ISBN 0-7645-3220-0 (paperback). li + 903 pp. LCCN QA76.76.H94P494 1998.

Pitty:1997:PSH

[PGRC97] D. Pitty, C. Gordon, P. Reeves, and A. Capey. The place of SGML and HTML in building electronic patient records. In Pappas et al. [PMS97], pages 329–333. ISBN 90-5199-343-9, 4-274-90169-6. ISSN 0926-9630. LCCN W 26.5 M489-7 1997. Two volumes.

Poppelier:1992:BR

[PH92] Nico Poppelier and Amy Hendrickson. Book reviews. Nederlandstalige TEX Gebruikersgroep (Dutch-speaking TEX Users Group), 92(1):135–138, May 1992.

Palowitch:1996:MSN

[PH96] C. Palowitch and L. Horowitz. Meta-information structures for networked information resources. Cataloging & classification quarterly, 21(3-4):109–130, ???? 1996. CODEN CCQUDB. ISSN 0163-9374.

Palme:1997:RMM

[PH97a] J. Palme and A. Hopmann. RFC 2110: MIME E-mail encapsulation of aggregate documents, such as HTML (MHTML), March 1997. URL ftp://ftp.

internic.net/rfc/rfc2110.
txt;   http://www.math.utah.
edu/pub/rfc/rfc2110.txt. Sta-

tus: PROPOSED STANDARD.

Park:1997:OVM

[PH97b] S. Park and T. Han. Objectoriented VRML for multi-user environments. In Spencer and Brutzman [SB97], pages 25–32. ISBN 0-89791-886-X. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/253437/.

ACM order number 434973.

Phillips:1998:PH

[Phi98a] Lee Ann Phillips. Practical HTML 4. Phillips, Lee Ann. Using HTML 4. Que Corporation, Indianapolis, IN, USA, 1998. ISBN 0-7897-2148-1 (paperback). xi + 700 pp. LCCN QA76.76.H94 P4942 1998.

Phillips:1998:UH

[Phi98b] Lee Anne Phillips. Using HTML 4. Que Corporation, Indianapolis, IN, USA, fourth edition, 1998. ISBN 0-7897-1562-7 (paperback). vii + 688 pp. LCCN QA76.76.H94P4942 1998.

Poulin:1997:OFL

[PHL97] D. Poulin, G. Huard, and A. Lavoie. The other formalization of law: SGML modelling and tagging. In ACM [ACM97b], pages 82–88. ISBN 0-89791-924-6. LCCN ????

URL http://www.acm.org/
pubs/contents/proceedings/

ai/261618/. ACM order number 604972.

Pirouz:1998:HWM

[PI98] Raymond Pirouz and Ardith Ibañez. HTML Web magic.

New Riders Publishing, Carmel, IN, USA, second edition, 1998. ISBN 1-56830-475-7 (paperback). xviii + 284 pp. LCCN QA76.76.H94P4943 1998.

Pietrowicz:1997:BRL

[Pie97] Stephen Pietrowicz. Book review: Late Night VRML 2.0 with Java. JavaWorld: IDG’s magazine for the Java community, 2(11), November 1997. CODEN ???? ISSN 1091-8906.

URL http://www.javaworld.
com/javaworld/jw-11-1997/
jw-11-bookreview-vrml.htm.

Piggott:1996:ITI

[Pig96] S. Piggott. International trade information. Business Information Review, 13(2):101–105, June 1996. CODEN BIREEY. ISSN 0266-3821.

Pinkerton:1995:FWP

[Pin95] Brian Pinkerton. Finding what people want: Experiences with the WebCrawler. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Searching/pinkerton/WebCrawler.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Pitti:1995:SGM

[Pit95] D. V. Pitti. Standard generalized markup language and the transformation of cataloging.

The Serials Librarian, 25(3): 243–??, 1995. CODEN SELID4. ISSN 0361-526X.

Pitty:1997:SSE

[Pit97] D. Pitty. SGML and the structure of the electronic medical record. In Teasdale [Tea97b], pages 106–112. ISBN ???? LCCN ????

Pitkow:1995:TIP

[PJ95] James E. Pitkow and R. K.

Jones. Towards an intelligent publishing environment. Computer Networks and ISDN Systems, 27(6):729–737, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://

www.elsevier.com/cas/tree/
store/comnet/sub/1995/27/
6/1454.pdf.

Plattner:1994:IAC

[PK94a] B. R. Plattner and J. P. A.

Kiers, editors. INET ’94: Annual conference — June 1994, Prague, Czech Republic, number 1 in Proceedings — INET 1994. Internet Society, Reston, VA, USA, 1994. ISBN ???? LCCN ????

Plattner:1994:PII

[PK94b] B. R. Plattner and J. P. A.

Kiers, editors. Proceedings of INET’94/JENC5. INET’94, the Annual Conference of the Internet Society held in conjuction with 5th Joint European Networking Conference (JENC5). Internet Soc, Reston, VA, USA, 1994. ISBN ???? LCCN ???? 2 vol.

Perrochon:1995:WWW

[PK95a] L. Perrochon and A. Kennel. World Wide Web access for blind people. In Anonymous [Ano95q], pages A43/1–27 (or A42/1–4??). ISBN ???? LCCN ????

Punin:1995:SHE

[PK95b] John R. Punin and Mukkai S. Krishnamoorthy. A simple HTML editor. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Autools/krishnamoorthy/user.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Paulson:1999:BBS

[PKM99] Olaf B. Paulson, Gitte Moos Knudson, and Torben Moos, editors. Brain barrier systems: proceedings of a symposium held at the Royal Danish Academy of Sciences and Letters, August 23–27, 1998, volume 45. Munksgaard, Copenhagen, Denmark, 1999. ISBN 87-16-12177-5. ISSN 0105-3639. LCCN WL 200 A392b 1998.

Pinckney:1995:DDE

[PKT95] Tom Pinckney, M. Frans Kaashoek, and Joshua A. Tauber. Dynamic documents: Extensibility and adaptability in the WWW. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
pinckney/dd.html. A selected

few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Plewe:1995:GPU

[Ple95] Brandon S. Plewe. The GeoWeb project: Using WAIS and the World Wide Web to aid location of distributed data sets. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Autools/plewe/plewe.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Padmanabhan:1995:IHL

[PM95] V. N. Padmanabhan and J. C.

Mogul. Improving HTTP latency. Computer Networks and ISDN Systems, 28(1–2): 25–??, December 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Pitts-Moultis:1999:XRT

[PM99] Natanya Pitts-Moultis. XML in record time. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1999. ISBN 0-7821-2340-6. xxv + 548 pp. LCCN QA76.76.H94P4945 1999.

Pitts-Moultis:1998:XBB

[PMK98] Natanya Pitts-Moultis and Cheryl Kirk. XML black book: Indispensable problem solver. Coriolis Group, Scottsdale, AZ, USA, 1998. ISBN 1-57610-284-X. xxii + 790 pp. LCCN QA76.76.H94 P4944 1999.

Pappas:1997:MIE

[PMS97] C. (Costas) Pappas, N. (Nicos) Maglaveras, and J. R. (Jean-Raoul) Scherrer, editors. Medical Informatics Europe ’97, volume 43 of Studies in Health Technology and Informatics. IOS Press, Amsterdam, The Netherlands, 1997. ISBN 90-5199-343-9, 4-274-90169-6. ISSN 0926-9630. LCCN W 26.5 M489-7 1997. Two volumes.

Pitts-Moultis:1998:DHB

[PMSC98] Natanya Pitts-Moultis, C. C. (Claire C.) Sanders, and Ramesh Chandak. Dynamic HTML black book. Coriolis Group, Scottsdale, AZ, USA, 1998. ISBN 1-57610-188-6. xxiii + 662 pp. LCCN QA76.76.H94P58 1998.

Pitts-Moultis:1998:HSS

[PMTJ98] Natanya Pitts-Moultis, Ed Tittel, and Stephen N. James. HTML style sheets design guide. Coriolis Group, Scottsdale, AZ, USA, 1998. ISBN 1-57610-211-4 (paperback). xv + 232 pp. LCCN QA76.76.H94P59 1998.

Pichler:1995:VMP

[POA^{+}95] M. Pichler, G. Orasche, K. Andrews, E. Grossman, and M. McCahill. VRweb: A multiprotocol VRML browser. In Nadeau and Moreland [NM95b], pages 77–85, 139. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Poler:1995:IWM

[Pol95] Ariel Poler. Improving WWW marketing through user information and non-intrusive communications. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Security/poler/poler.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Poppelier:1991:STS

[Pop91] N. A. F. M. Poppelier. SGML and TEX in scientific publishing. TUGboat, 12(1):105–109, March 1991. ISSN 0896-3207.

Poppelier:1992:BRJ

[Pop92] Nico Poppelier. Book reviews:

Jane Hahn, L^{A}TEX for Everyone; Eric van Herwijnen, Practical SGML. TUGboat, 13(2):182– 184, July 1992. ISSN 0896-3207.

Popham:1993:USHa

[Pop93a] M. G. Popham. Use of SGML and HyTime in UK universities.

Information services and use, 13 (2):103–109, ???? 1993. CO-DEN ISUDX8. ISSN 0167-5265.

Popham:1993:USHb

[Pop93b] M. G. Popham. Use of SGML and HyTime in UK universities. In Van Halm [Van93a], pages 103–110. ISBN ???? ISSN 0167-5265. LCCN ????

Popham:1994:US

[Pop94a] Mike Popham. An update on SGML. Information Management & Technology, 27(6):247– 250, 262, November 1, 1994. CODEN IMTHEM. ISSN 0266-6960.

Poppelier:1994:EHP

[Pop94b] Nico Poppelier. Eric van Herwijnen, Practical SGML. TUGboat, 15(1):24–25, March 1994. ISSN 0896-3207.

Pope:1997:USS

[Pop97] P. Pope. Using the SAS system for large volume HTML document production. In Anonymous [Ano97-31], pages 988– 993. ISBN 1-55544-961-1. LCCN ????

Powell:1994:AWW

[Pow94] J. Powell. Adventures with the World-Wide WEB: creating a hypertext library information system. Database, 17(1):59–60, 62–66, February 1994. CODEN DTBSDQ. ISSN 0162-4105.

Powell:1995:SWW

[Pow95] James Powell. Spinning The World-Wide Web: An HTML primer. Database, 18(1):54–59, February–March 1995. CODEN DTBSDQ. ISSN 0162-4105.

Powell:1997:HP

[Pow97]  James E. Powell.  HTML

Plus! Multimedia Ser. Wadsworth, Pacific Grove, CA, USA, July 1997. ISBN 0-534-51626-2. xiv + 418 pp. LCCN QA76.76.H94P68 1997. US$27.75.

Powell:1998:HCR

[Pow98a] Thomas A. Powell. HTML:

the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 1998. ISBN 0-07-882397-8 (paperback). xxix + 1073 pp. LCCN QA76.76.H94 P684 1998.

Powers:1998:DH

[Pow98b]  Shelley Powers.  Dynamic

HTML. I D G Books Worldwide, Indianapolis, IN, USA, 1998. ISBN 0-7645-8053-1. xxiii + 571 pp. LCCN QA76.76.H94P69 1998.

Powell:1999:HCR

[Pow99] Thomas A. Powell. HTML:

the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, second edition, 1999. ISBN 0-07-211977-2 (paperback). xxii + 1130 pp. LCCN QA76.76.H94 P684 1998. Includes HTML 4, CSS 1 and 2, DHTML, and XML.

Parr:1995:LCM

[PR95a] T. J. Parr and T. F. Rohaly. A language for creating and manipulating VRML. In Nadeau and Moreland [NM95b], pages 123–131. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Pitkow:1995:SYR

[PR95b] Jim Pitkow and Mimi Recker. A simple yet robust caching algorithm based on dynamic access patterns. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
pitkow/caching.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Paquet:1998:CAV

[PR98a] E. Paquet and M. Rioux. Content-based access of VRML libraries. In Ip and Smeulders [IS98], pages 20–32. ISBN 3-540-64826-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575 .I318 1998.

Paquet:1998:CSEb

[PR98b] E. Paquet and M. Rioux. A content-based search engine for CAD and VRML databases. In Roller [Rol98], pages 427–434. ISBN 0-9532576-0-6, 0-9532576-1-4, 0-9532576-2-2, 0-9532576-3-0, 0-9532576-4-9. LCCN ????

Paquet:1998:CSEa

[PR98c] E. Paquet and M. Rioux. A content-based search engine for VRML databases. In IEEE [IEE98a], pages 541–546. ISBN 0-8186-8497-6 (softbound), 0-7803-5063-4 (casebound), 0-8186-8499-2 (microfiche). ISSN 1063-6919. LCCN TA1632 .I36 1998. IEEE Catalog Number 98CB36231. IEEE Computer Society Order Number PR08497.

Prescod:1996:MMP

[Pre96] P. Prescod. Multiple media publishing in SGML. In ACM [ACM96d], pages 3–9. ISBN 0-89791-799-5. LCCN QA 76.9 D6 I56 1996.

Price:1987:ST

[Pri87] Lynne A. Price. SGML and TEX. TUGboat, 8(2):221–225, July 1987. ISSN 0896-3207.

Price:1993:MIF

[Pri93] Roger Price. MHEG: An introduction to the future international standard for hypermedia object interchange. In Proceedings of the 1st ACM International Conference on Multimedia Aug 1–6, 1993, Anaheim, CA, USA, pages 121–128. ACM Press, New York, NY, USA, 1993. ISBN ???? LCCN ???? US$.

Prince:1995:CWB

[Pri95] C. J. Prince. Caught in the web (banking on the Internet). Bank Systems + Technology, 32 (7):44–48, July 1995. CODEN BSYTEE. ISSN 1045-9472.

Price:1996:PSI

[Pri96a] Lynne A. Price. Practical SGML as an introduction to SGML: Finds the book’s strengths and weaknesses in its details. Asterisk: the journal of computer documentation, 20(2):36–??, May 1, 1996. CODEN ASTRF7. ISSN 0731-1001.

Price:1996:RCT

[Pri96b]  Lynne A. Price.  Ronald

C. Turner, Timothy A. Douglass, and Audrey J. Turner, README.1ST: SGML for Writers and Editors. TUGboat, 17(1): 35–37, March 1996. ISSN 0896-3207.

Prior:1996:DAM

[Pri96c] F. Prior. Database access methods for medical imaging: DI-COM, SQL, and HTML. In IEEE [IEE96f], pages 288–292. ISBN 0-8186-7560-8. LCCN R857.O6 I57 1995. IEEE Computer Society Press order number PR07560.

Priestley:1997:NIH

[Pri97] M. Priestley. Navigation issues in hypertext: Documenting complex hierarchies with HTML frames. In ACM [ACM97a], pages 223–236. ISBN 0-89791-861-4. LCCN QA 76.9 D6 I56 1997.

Proffer:1995:TTW

[Pro95] William J. Proffer. Taxing times: WWW access to federal and state tax forms, information and assistance. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Fin/proffer/proffer_paper.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Price:1988:ESA

[PS88] Lynne A. Price and Joe Schneider. Evolution of an SGML application generator. In Solem [Sol88], pages 51–60. ISBN 0-89791-291-8. LCCN QA 76.9 T48 A25 1988.

Poo:1995:PIS

[PS95] G. S. Poo and E. S. Seumahu, editors. Proceedings of IEEE Singapore International Conference on Networks/International Conference on Information Engineering 1995. Theme: Electrotechnology 2000: Communications and Networks: 3–7 July 1995, Singapore. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2579-6. LCCN TK5105.2 .S5 1995. IEEE Catalog No. 95TH8061.

Pokrovskii:1996:HEC

[PS96] S. B. Pokrovskii and G. G.

Stepanov. A hypertext environment: capabilities and perspectives. Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 22(3):150–155, May–June 1996. CODEN PROGD3. ISSN 0132-3474, 0361-7688.

Pardi:1998:DHA

[PS98a] William J. Pardi and Eric M. Schurman. Dynamic HTML in action. Microsoft Press, Bellevue, WA, USA, 1998. ISBN 1-57231-820-1. xiii + 377 pp. LCCN QA76.76.H94P37 1998.

Patterson:1998:UH

[PS98b] Lois R. Patterson and Todd Stauffer. Using HTML 4. Que Corporation, Indianapolis, IN, USA, third edition, 1998. ISBN 0-7897-1450-7. xvii + 386 pp. LCCN QA76.76.H94U85 1998.

Piron:1992:UEB

[PSG92] C. Piron, J. L. Sanson, and G. Giordano. Using electronic book metaphor for the viewing of R and D projects. Technical Report 93NO00009, EDF-Electricité de France, Clamart, France, October 1992. 87 pp.

Pimentel:1997:HOO

[PST97] M. G. C. Pimentel, L. P. Scott, and C. A. C. Teixeira. Hy-Time from the object-oriented point of view. In Bose [Bos97], pages 655–660. ISBN 0-9696869-2-7 (valid ISBN allocated to two books: the 7th and 8th workshop). LCCN QA76.9.S63 I56 1998.

Pimentel:1997:HEA

[PSTW97] M. G. C. Pimentel, L. P. Scott, C. A. C. Teixeira, and C. J.

Winter. Hydrogen energy: After 21/2 decades of research — it’s HYTIME for entrepreneurs, bankers, economists. In Bose [Bos97], pages 368–375. ISBN 0-9696869-2-7 (valid ISBN allocated to two books: the 7th and 8th workshop). LCCN QA76.9.S63 I56 1998.

Pitts:1998:HSS

[PTJ98] Natanya Pitts, Ed Tittel, and Stephen N. James. HTML style sheets design guide. Coriolis Group, Scottsdale, AZ, USA, 1998. ISBN 1-57610-211-4 (paperback). xv + 232 pp. LCCN QA76.76.H94P59 1998.

Pullinger:1995:LPE

[Pul95] D. J. Pullinger. Learning from putting electronic journals on SuperJANET: the SuperJournal Project. Interlending and Document Supply, 23(1):20–27, ???? 1995. CODEN IDSUDQ. ISSN 0264-1615.

Purcell:1998:WDC

[Pur98] Lee Purcell. Web developer.com guide to creating Web channels. Wiley, New York, NY, USA, 1998. ISBN 0-471-25168-2. ???? pp. LCCN QA76.625 .P87 1998.

Putz:1994:IIS

[Put94] Steve Putz. Interactive information services using World-Wide Web hypertext. Computer Networks and ISDN Systems, 27 (2):273–280, November 1, 1994. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1994&
                                        volume=27&issue=2&aid=1391.

Price-Wilkin:1994:GBW

[PW94] John Price-Wilkin. A gateway between the World-Wide Web and PAT: Exploiting SGML through the Web. Public-access Computer Systems Review, 5(7): 5–27, ???? 1994. ISSN 1063-164X.

Pfaffenberger:1996:SWS

[PW96] Bryan Pfaffenberger and Dave Wall. 10 secrets for Web success: what it takes to do your site right. Ventana Press, Chapel Hill, NC, USA, 1996. ISBN 1-56604-370-0. ???? pp. LCCN TK5105.888 .P4255 1996.

Powell:1998:HPR

[PW98] Thomas A. Powell and Dan Whitworth. HTML programmer’s reference. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 1998. ISBN 0-07-882559-8 (paperback). xiii + 396 pp. LCCN QA76.76.H94P495 1998.

Powell:1999:HRD

[PW99] Thomas A. Powell and Dan Whitworth. HTML 4: les references du programmeur. Pro-Micro. Osman Eyrolles Multimedia, Paris, France, 1999. ISBN 2-7464-0001-4. xii + 459 pp. LCCN ????

Pichumani:1998:DFI

[PWHK98] R. Pichumani, D. Walker, L. Heinrichs, and C. Karadi. The design of frog island: A VRML world for biology. In Landauer and Bellman [LB98a], pages 31–38. ISBN 1-56555-137-0. ISSN 0735-9276. LCCN QA76.9.V5 V578 1998.

Pacifici:1998:MSS

[PY98] Giovanni Pacifici and Alaa Youssef. A markup system for shared HTML documents. Research report RC. International Business Machines Corporation. Research Division ; 21338 International Business Machines Corporation. Research Division. Research report ; RC 21338. RC 21338 (96124), IBM T.J. Watson Research Center, Yorktown Heights, NY, USA, 1998. 19 pp.

Qayyum:1996:UIV

[Qay96] H. Qayyum. Using IVDS and VBI for interactive television. In IEEE [IEE96a], pages 657– 666. CODEN ITCEDA. ISBN 0-7803-3029-3 (softbound), 0-7803-3030-7 (casebound), 0-7803-3031-5 (microfiche). ISSN 0098-3063. LCCN TK7869 .I16 1996.

Quint:1995:SAE

[QRV95] Vincent Quint, Cécile Roisin, and Irène Vatton. A structured authoring environment for the World-Wide Web. Computer Networks and ISDN Systems, 27 (6):831–840, April 3, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/comnet/
                                        sub/1995/27/6/1462.pdf.

Quirchmayr:1998:DES

[QSBC98] Gerald Quirchmayr, Erich Schweighofer, and T. J. M. Bench-Capon, editors. Database and Expert Systems Applications: 9th International Conference, DEXA ’98, Vienna, Austria, August 24–28, 1998, proceedings, volume 1460 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. ISBN 3-540-64950-6 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 D15592 1998.

Quesenbery:1997:BGH

[Que97] W. Quesenbery. A beginner’s guide to HTML. In Anonymous [Ano97b], pages 342–?? ISBN 0-914548-94-8, 0-914548-95-6. LCCN ????

Quesenbery:1998:BGH

[Que98] W. Quesenbery. A beginner’s guide to HTML and Web design. In Anonymous [Ano98u], pages 275–?? ISBN 1-891709-00-3, 1-891709-01-1. LCCN ????

Quinn:1995:GRC

[Qui95] Christine A. Quinn. From grass roots to corporate image — the maturation of the Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/

Campus.Infosys/quinn/quinn.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Radev:1995:RWS

[Rad95] Dragomir R. Radev. REN-DEZVOUS: A WWW synchronization system. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
radev/radev.html. A selected

few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Radecki:1997:HMM

[Rad97] S. Radecki. HTML: Making the move. In Anonymous [Ano97b], pages 343–346. ISBN 0-914548-94-8, 0-914548-95-6. LCCN ????

Radoslavova:1998:STC

[Rad98] D. Radoslavova. SGML tagging of complex texts in old Slavic liturgical manuscripts. In Dobreva [Dob98], pages 165–167. ISBN 954-9650-02-2. LCCN ????

Raggett:1994:EWS

[Rag94a] D. Raggett. Extending WWW to support platform independent virtual reality. In Plattner and Kiers [PK94b], pages 242/1, 242/3–242/6 (vol. 1). ISBN ???? LCCN ???? 2 vol.

Raggett:1994:RHDa

[Rag94b] D. Raggett. A review of the HTML+ document format. In Cailliau [Cai94b], pages 135– 146. CODEN CNISE9. ISBN ???? ISSN 0169-7552 (print), 1879-2324 (electronic). LCCN ???? Published in Computer Networks and ISDN Systems, 27, No. 2.

Raggett:1994:RHDc

[Rag94c] D. Raggett. A review of the HTML+ document format. In Plattner and Kiers [PK94a], pages 213–220. ISBN ???? LCCN ????

Raggett:1994:RHDb

[Rag94d] David Raggett. A review of the HTML+ document format. Computer Networks and ISDN Systems, 27(2):135– 145, November 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1994&
                                        volume=27&issue=2&aid=1382.

Raggett:1994:RHDd

[Rag94e] David Raggett. A review of the HTML+ document format. In Plattner and Kiers [PK94b], pages 213/1–213/10 (vol. 1). ISBN ???? LCCN ???? 2 vol.

Raggett:1996:IOH

[Rag96a] D. Raggett. Inserting objects into HTML. In Anonymous [Ano96-46], pages 29–42. ISBN 1-56592-210-7. ISSN 1085-2301. LCCN ???? US$24.95.

Raggett:1996:RHT

[Rag96b]  D. Raggett.  RFC 1942:

HTML tables, May 15, 1996.

URL ftp://ftp.internic.
net/rfc/rfc1942.txt;   http:
//www.math.utah.edu/pub/rfc/
rfc1942.txt.

Raggett:1998:CYW

[Rag98] Dave Raggett. Clean up your Web pages with HP’s HTML Tidy. Computer Networks and ISDN Systems, 30(1–7): 730–732, April 1, 1998. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

URL http://www.elsevier.
com/cas/tree/store/comnet/
sub/1998/30/1-7/1933.pdf.

Rahtz:1995:ALL

[Rah95] Sebastian Rahtz. Another look at L^{A}TEX to SGML conversion. TUGboat, 16(3):315–324, September 1995. ISSN 0896-3207.

Raitt:1990:OII

[Rai90] D. I. Raitt, editor. Online Information 90. 14th International Online Information Meeting. Proceedings. Learned Information, Medford, NJ, USA, ???? 1990. ISBN 0-904933-75-X. LCCN ????

Raitt:1992:OII

[Rai92] D. I. Raitt, editor. Online Information 92. 16th International Online Information Meeting Proceedings. Learned Information, Medford, NJ, USA, 1992. ISBN 0-904933-83-0. LCCN ????

Rakov:1995:ISM

[Rak95] Simon A. Rakov. The Internet sampler: A Mosaic-based museum kiosk about the Internet. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Museum/rakov/rakov.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Raman:1997:DJP

[Ram97] T. V. Raman. Documents are not just for printing. Mathematical and computer modelling, 25(4):21–??, ???? 1997. CO-DEN MCMOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

Ramalho:1998:LAH

[Ram98] Jose Antonio Ramalho. Learn advanced HTML 4.0 with DHTML. Wordware Publishing, Plano, TX, USA, 1998. ISBN 1-55622-586-5 (paperback). xxii + 752 pp. LCCN QA76.76.H94R36 1998.

Ramalho:1999:LAH

[Ram99] Jose Antonio Ramalho. Learn advanced HTML 4.0 with DHTML. Wordware Publishing, Plano, TX, USA, 1999. ISBN 1-55622-586-5 (paperback). xxii + 752 pp. LCCN QA76.76.H94R36 1999.

Randolph:1998:CMN

[Ran98] G. B. Randolph. Cues, mental nodes, cognitive overhead, and HTML: A theoretical framework for Web communication. In Hoadley and Benbasat [HB98], pages 1079–1080. ISBN ???? LCCN ????

Ramalho:1999:SDW

[RARH99] José Carlos Ramalho, José João Almeida, Jorge Gustavo Rocha, and Pedro Henriques. SGML documents: Where does quality go? Markup Languages: Theory & Practice, 1(1):75–90, Winter 1999. CODEN MLTPFG. ISSN 1099-6621.

Robertson:1999:HH

[RAS99] Greg Robertson, Tim Altom, and Simply Written, Inc. Hands on HTML. Prima Tech, Rocklin, CA, USA, 1999. ISBN 0-7615-1885-1. xxv + 547 pp. LCCN ????

Rath:1995:SSH

[Rat95] H. Holger Rath. Spezifikation von strukturierten Hypermedia-Dokumenten auf der Basis von SGML und HyTime. In Kuhlen and Rittberger [KR95], pages 129–144. ISBN 3-87940-509-3. LCCN ????

Rath:1998:XCC

[Rat98] H. H. Rath. XML: chance and challenge for online information providers. In Raitt et al. [RMG98], pages 339–346. ISBN 1-900871-31-9. LCCN ????

Roussos:1995:FRI

[RB95a] Maria Roussos and Paul Brenner. The file room: An interactive archive of cultural censorship. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Arts/
brenner/roussos.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Rutledge:1995:HSH

[RB95b] L. Rutledge and J. F. Buford. HyTime: A Standard for Hypermedia Document Systems. Springer-Verlag Inc., New York, NY, USA, 1995. ISBN 3-540-58260-6. 350 pp. LCCN ???? US$59.00.

Rosch:1996:RTM

[RB96] P. Rosch and M. Baentsch. Reviewing two multimedia presentation (quasi-) standards. In IEEE [IEE96h], page ?? ISBN 0-8186-7511-X. LCCN QA76.575 .M84 1996.

Rasmussen:1995:GIM

[RBBN95] Mary Rasmussen, Brad Blumenthal, Steve Borysewicz, and Paul Neumann. The global interactive museum using Mosaic. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Museum/rasmussen/global_museum.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Rice:1995:UMU

[RBBW95] Judy Rice, Rosina Bignall, Dalinda Bond, and Phillip J. Windley. Uses of Mosaic in a university setting. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ???? URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/
rice.university/article.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Rutledge:1995:MTH

[RBR95] L. Rutledge, J. F. Buford, and J. L. Rutledge. Modeling techniques for HyTime. In Chua et al. [CPK95], pages 67–82. ISBN 981-02-2502-4. LCCN QA76.575.M834 1995.

Rada:1995:HIU

[RBZ95] R. Rada, G. Bird, and Min Zheng. Hypertext interchange using ICA. Journal of Documentation, 51(2):99–117, June 1995. CODEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Rada:1994:SSN

[RC94] Roy Rada and George S. Carson. Sharing standards: The new media. Communications of the ACM, 37(9):23– 25, September 1994. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-0782/

214782.html. Discusses the problem of the multiplicity of document markup formats, and the role that SGML and Open-Doc might play in providing a common intermediate format for document exchange.

Rathe:1995:BWH

[RC95] R. Rathe and G. Cornwall. Beyond the Web: HTML as an Abstract Document Specification. In Gardner [Gar95], pages 973– ?? ISBN 1-56053-123-1. ISSN 0195-4210. LCCN W 26.5 S978 1995.

Rathe:1996:BSE

[RC96] R. Rathe and G. Cornwall. Beyond the superhighway: Exploiting HTML in various instructional contexts. In Cimino [Cim96], pages 947–?? ISBN 1-56053-208-4. LCCN W 26.5 S978 1996.

Roehl:1997:LNV

[RCRBR97] Bernie Roehl, Justin Couch, Cindy Reed-Ballreich, and Tim Rohaly. Late Night VRML 2.0 With Java. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 1-56276-504-3. xxxvi + 710 pp. LCCN QA76.76.H94L36 1997. US$49.99. URL http:/

/ece.uwaterloo.ca/~broehl/
vrml/lnvj.

Rabarijoana:1999:EXC

[RDC99] A. Rabarijoana, R. Dieng, and O. Corby. Exploitation of XML for corporate knowledge management. Lecture Notes in Computer Science, 1621:373– 378, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rebelsky:1995:WRI

[Reb95] Samuel A. Rebelsky. A Web of resources for introductory computer science. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/
rebelsky/rebelsky.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Reding:1999:HIB

[Red99] Elizabeth Eisner Reding. HTML:

illustrated brief. Illustrated series. Course Technology, Cambridge, MA, USA, revised / use with any text editor or browser! edition, 1999. ISBN 0-7600-6077-0. various pp. LCCN ????

Regan:1998:TLW

[Reg98] T. Regan. Taking living worlds into peoples living rooms. In Spencer and Roehl [SR98], pages 71–76. ISBN 1-58113-022-8. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/271897/.

Rehn:1995:IWH

[Reh95] G. Rehn. An introduction to writing HTML and developing hypermedia for the WWW. In Oliver and Wild [OW95], pages 371–375. ISBN 1-875534-08-3, 1-875534-10-5. LCCN ????

Reinke:1993:TSI

[Rei93] U. Reinke. Towards a standard interchange format for terminographic data. In Schmitz [Sch93], pages 270–282. ISBN ???? LCCN ????

Remy:1997:DPJ

[Rem97] Martin Remy. Design patterns, Java and Web development. Dr. Dobb’s Journal of Software Tools, 22(6):28, 30, 32, 34, 98–100, June 1997. CODEN DDJOEB. ISSN 1044-789X.

Ressler:1995:AUV

[Res95]  Sandy Ressler.  Approaches

using virtual environments with Mosaic. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        VR/ressler/veMain.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Ressler:1997:AEP

[Res97] Sandy Ressler. The Art of Electronic Publishing: the Internet and beyond. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-488172-

9. xvii + 450 pp. LCCN Z286.E43R47 1997. US$39.95.

Retzlaff:1996:CCS

[Ret96] Don Retzlaff. C/C++ source to HTML converter: Use those ubiquitous Web browsers to advantage and show C or C++ code in a more readable format. C/C++ Users Journal, 14(10): 57–58, October 1996. CODEN CCUJEX. ISSN 1075-2838.

Rodriguez:1995:SSL

[RFM95] Arturo A. Rodriguez, Martin Fisher, and Brian Markey. Standards: Scripting languages emerge in standards bodies. IEEE MultiMedia, 2(4):88– 92, Winter 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

URL http://dlib.computer.
org/mu/books/mu1995/pdf/u40088.

pdf. Discusses the proposed Standard Multimedia/ Hypermedia Scripting Language (SMSL), a superset of Java.

Register:1995:IPI

[RG95] Joel D. Register and John L.

Gerone. Interfaces for public information and scientific research. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/HCI/
register/register.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Raman:1997:DMM

[RG97a] T. V. Raman and D. Gries. Documents mean more than just paper! Mathematical and computer modelling, 26(1):45– ??, ???? 1997. CODEN MC-MOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

Roberts:1997:SSP

[RG97b] A. Roberts and D. Garnet. An SGML solution for patient records. In Waegemann [Wae97], pages 273–278. ISBN 0-9640667-3-4. LCCN ????

Risse:1997:EIV

[RH97] Thomas Risse and Matthias Hemmje. Entwurf und Implementierung einer VRML Bibliothek fur ein objektrelationales Datenbanksystem auf der Basis des VRML 2.0 Standards. GMD-Studien; Nr. 324. GMD-Forschungszentrum Informationstechnik GmbH, Sankt Augustin, Germany, September 1997. ISBN 3-88457-324-1. 143 pp. LCCN ????

Ricciuti:1992:PSG

[Ric92] M. Ricciuti. Publishing software gets serious. Datamation, 38(24):65–66, 68–69, December 1992. CODEN DTMNAT. ISSN 0011-6963.

Richardson:1994:HLU

[Ric94] P. Richardson. HTML, the language used to create WWW pages, and a tool that makes life a lot easier. .EXE: the software developers’ magazine, 9(6): 92–??, 1994. CODEN EXEEE5. ISSN 0268-6872.

Rice:1995:HH

[Ric95a] Judy Cossel Rice. Hypermedia for HyperKids. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/
rice.hyperkids/Hyper.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Richardson:1995:HML

[Ric95b]  P. Richardson.  HyperText

Markup Language 3.0. .EXE: the software developers’ magazine, 10(2):29–??, 1995. CO-DEN EXEEE5. ISSN 0268-6872.

Richey:1995:CHS

[Ric95c] Jeff Richey. Condition handling in SQL persistent stored modules. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(3):98– 103, September 1995. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Richmond:1996:EOS

[Ric96] Alice Richmond. Enticing online shoppers to buy — a human behavior study. Computer Networks and ISDN Systems, 28 (7–11):1469–1480, May 1, 1996.

CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=
1611.

Richardson:1998:VBB

[Ric98] J. F. Richardson. VRML based behaviour database editor. In Heudin [Heu98], pages 49–62. ISBN 3-540-64780-5 (softcover). LCCN QA76.9.H85 V628 1998 Bar.

Ridley:1995:IIA

[Rid95] Michael Ridley. Innovation and implementation: Adopting and managing World Wide Web services in academic libraries. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
LibApps/ridley/ridley.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Rieger:1994:SPA

[Rie94] Wolfgang Rieger. SGML für die Praxis — Ansatz und Einsatz von ISO 8879. (German) [SGML for practice — Introduction and use of ISO 8879]. Springer-Verlag Inc., New York, NY, USA, 1994. ISBN 3-540-57534-0. ???? pp. LCCN ???? DM98.00. A copy of this bibliography, translated to SGML format, is expected to be included in a floppy disk with this book.

Riitahuhta:1997:IWC

[Rii97] A. Riitahuhta, editor. ICED 97: world class design by world class methods: proceedings of the 11th International Conference on Engineering Design in Tampere, August 1997, number 25 in Schriftenreihe WDK Workshop Design Konstruction 1997. Tampere University of Technology, Tampere, Finland, 1997. ISBN 951-722-788-4. LCCN ????

Raitt:1993:OII

[RJ93] D. I. Raitt and B. Jeapes, editors. Online Information 93. 17th International Online Information Meeting Proceedings. Learned Information, Medford, NJ, USA, 1993. ISBN 0-904933-85-7. LCCN ????

Raitt:1994:OII

[RJ94] D. I. Raitt and B. Jeapes, editors. Online Information 94. 18th International Online Information Meeting. Proceedings. Learned Information, Medford, NJ, USA, 1994. ISBN 0-904933-88-1. LCCN ????

Raitt:1995:OII

[RJ95] D. I. Raitt and B. Jeapes, editors. Online Information 95. 19th International Online Information Meeting Proceedings: London, 5–7 December 1995. Learned Inf. (Europe), Oxford, UK, 1995. ISBN 0-904933-94-6. LCCN ????

Randall:1996:UH

[RJKS96] Neil Randall, John Jung, Greg Knauss, and Tom Savola. Using HTML: The Fast and Easy Way to Learn. Que Corporation, Indianapolis, IN, USA, 1996. ISBN 0-7897-0622-9. xii + 374 pp. LCCN QA76.76.H94R37 1996. US$24.99, CDN$33.99, UK£22.99.

Robertson:1995:VFDb

[RJN95a] D. Robertson, W. Johnston, and W. Nip. Virtual frog dissection: Interactive 3D graphics via the Web. Computer Networks and ISDN Systems, 28(1– 2):155–??, December 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Robertson:1995:VFDa

[RJN95b] David W. Robertson, William E. Johnston, and Wing Nip. Virtual frog dissection: Interactive 3D graphics via the Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
BioChem/robertson/robertson.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Ryu:1998:DIS

[RKL^{+}98] Eunsook Ryu, Yonghun Kim, Kangchan Lee, Kyuchul Lee, Sangki Lee, Hyunki Kim, Haeran Lee, and Zongchul Zhoo. Design and implementation of SGML document repository system. In Callaos et al. [COW98], pages 129–133. ISBN 980-07-5078-9 (vol. 1), 980-07-5079-7 (vol. 2), 980-07-5080-0 (vol. 3), 980-07-5081-9 (vol. 4). LCCN ???? Four volumes.

Rossello:1996:QIM

[RKW96] R. Rossello, A. Karsh, and M. Walter. Quark Immedia: making pages interactive. Seybold Report on Desktop Publishing, 10(11):3–11, July 1996. CO-DEN SRDFED. ISSN 0889-9762.

Roden:1995:MCW

[RL95a] Peter Roden and Suzana Lisanti. MIT’s campus-wide information system: Scaling up for World Wide Web support. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Campus.Infosys/roden/roden.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Rzepa:1995:CWH

[RL95b] Henry S. Rzepa and Christopher Leach. Chemistry in the Web: Hyperactive molecules. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
BioChem/rzepa/rzepa.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Raggett:1996:DGH

[RLA96] Dave Raggett, Jenny Lam, and Ian Alexander. Definitive Guide to HTML 3.0: Electronic Publishing on the World Wide Web. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-87693-0. xv + 398 pp. LCCN QA76.76.H94 H765 1996. US$26.95.

Raggett:1998:RH

[RLAK98] Dave Raggett, Jenny Lam, Ian Alexander, and Michael Kmiec. Raggett on HTML 4. Addison-Wesley, Reading, MA, USA, second edition, 1998. ISBN 0-201-17805-2. xv + 437 pp. LCCN QA76.76.H94 R35 1998. US$29.95.

Rhodes-Marriott:1995:SSQ

[RM95a] A. Rhodes-Marriott. To SGML or not to SGML. that is the question? In Anonymous [Ano95-37], pages 269– 271. ISBN 0-914548-86-7, 0-914548-87-5. LCCN T 11 S63 1995.

Roth:1995:PAP

[RM95b] Al Roth and Andre Marien, editors. The practical application of Prolog: proceedings of the Third International Conference on the Practical Application of Prolog: 3rd–6th April 1995, Hotel Sofitel Paris St Jacques, Paris, France. Alinmead Software, Blackpool, UK, 1995. ISBN 0-9525554-0-9. LCCN QA76.73.P76 I58 1995.

Ras:1996:FIS

[RM96a] Zbigniew W. Ras and Maciek Michalewicz, editors. Foundations of intelligent systems: 9th international symposium, ISMIS ’96, Zakopane, Poland, June 9–13, 1996: proceedings. Springer-Verlag Inc., New York, NY, USA, 1996. ISBN 3-540-61286-6. LCCN QA76.76.E95I582 1996.

Rubinsky:1996:BHS

[RM96b] Yuri Rubinsky and Murray Maloney. Beyond HTML: SGML Publishing on the World Wide Web. Charles F. Goldfarb series on open information management. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, November 1996. ISBN 0-13-519984-0. ???? pp. LCCN QA76.76.H94R83 1997. US$39.95.

Ronquillo:1997:PIO

[RM97a] O. Ronquillo and J. Moorman.

A process for implementing online documentation and online help using SGML. In Anonymous [Ano97b], pages 383–386. ISBN 0-914548-94-8, 0-914548-95-6. LCCN ????

Rubinsky:1997:SWS

[RM97b] Yuri Rubinsky and Murray Maloney. SGML on the Web: small steps beyond HTML. Charles F. Goldfarb series on open information management. Ventana Communications Group, Incorporated, Research Triangle Park, NC, USA, 1997. ISBN 0-13-519984-0. xxvi + 499 pp. LCCN QA76.76.H94R83 1997. US$59.95.

Rostek:1993:WWS

[RMF93] L. Rostek, W. Mohr, and D. H. Fischer. Weaving a web: the structure and creation of an object network representing an electronic reference work. Electronic Publishing—Origination, Dissemination, and Design, 6 (4):495–505, December 1993. CODEN EPODEU. ISSN 0894-3982.

Raitt:1998:OII

[RMG98] David I. Raitt, Brian McKenna, and Catherine Graham, editors. Online Information 98: 22nd International Online Information Meeting: proceedings, London, 8–10 December 1998. Learned Information Europe Ltd, Oxford, UK, 1998. ISBN 1-900871-31-9. LCCN ????

Raposo:1997:WRV

[RMR97] A. Raposo, L. Magalhaes, and I. Ricarte. Working with remote VRML scenes through low-bandwidth connections. In de Figueiredo and Netto [dN97], pages 34–41. ISBN 0-8186-8102-0, 0-8186-8103-9, 0-8186-8104-7. LCCN T385 .S59 1997. IEEE Computer Society order number PR08102. IEEE order plan catalog number 97TB100164.

Roach:1995:MGI

[Roa95] J. Michael Roach. 10 minute guide to Internet Assistant for Word. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 1-56761-616-X. vi + 168 pp. LCCN QA76.76.H94 R63 1995. US$10.99, CDN$14.95.

Robinson:1994:EDHb

[Rob94a] B. Robinson. Electronic document handling using SGML: Management report. The British Library research and development reports, 6157(??):1– ??, 1994. CODEN BLRRDP. ISSN 0308-2385.

Robinson:1994:EDHa

[Rob94b] Brian Robinson. Electronic document handling using SGML: hypertext interchange and SGML. Technical Report British Library R and D report 6155, British Library Research and Development Department, London, UK, 1994. 19 pp.

Robinson:1994:TPT

[Rob94c] Peter Robinson. The transcription of primary textual sources using SGML, volume 6 of Office for Humanities Communication publications. Office for Humanities Communication, Oxford, UK, 1994. ISBN 1-897791-07-0. vii + 136 pp. LCCN QA76.9.T48 R63 1994.

Robin:1996:ESP

[Rob96] B. Robin. Electronic scholarly publishing: today’s technical alternatives. In Robin et al. [RPWW96], pages 822– 826. ISBN ???? LCCN ????

Robbins:1998:HIM

[Rob98] Curt Robbins. HTML 4.0 Intermediate: mastering frames and tables. DDC Publishing, Inc., ????, 1998. viii + 142 pp.

Rockley:1993:PLD

[Roc93a]  A. Rockley.  Putting large

documents online. In ACM [ACM93a], pages 273–281. ISBN 0-89791-630-1. LCCN QA76.9.D6I57 1993.

Rockley:1993:OHS

[Roc93b] Ann Rockley. Ontario Hydro and SGML. Technical Communication, 40(3):383–386, August 1993. CODEN TLCMBT. ISSN 0049-3155.

Roddick:1992:SSQ

[Rod92] John F. Roddick. SQL/SE: a query language extension for databases supporting schema evolution. SIGMOD Record (ACM Special Interest Group on Management of Data), 21(3): 10–16, September 1992. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roesser:1987:NIS

[Roe87]  J. R. Roesser.  A nota-

tion for the input and storage of mathematics:-an application of SGML. In Salvendy et al. [SSH87], pages 137–142. ISBN 0-444-42847-X. LCCN QA76.9.I58 I593 1987.

Roehl:1995:DVR

[Roe95] B. Roehl. Distributed virtual reality — an overview. In Nadeau and Moreland [NM95b], pages 39–43. ISBN 0-89791-818-5. LCCN QA76.76.H94 S95 1995. ACM order number 434953.

Roehl:1996:VTT

[Roe96] Bernie Roehl. VRML tech talk. Linux Journal, 21:??, January 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Role:1991:SSD

[Rol91] F. Role. The SGML standard:

describing the logical structure of documents. Documentaliste – Sciences de l’Information, 28(4-5):187–192, July–October 1991. CODEN DSINE6. ISSN 0012-4508.

Roller:1998:AMD

[Rol98] D. (Dieter) Roller, editor. Automotive mechatronics: design & engineering. 31st ISATA, Dusseldorf, Germany, 2nd–5th June 1998, volume 1 of ISATA - Proceedings- 1998. ISATA Dusseldorf Trade Fair, Croydon, UK, 1998. ISBN 0-9532576-0-6, 0-9532576-1-4, 0-9532576-2-2, 0-9532576-3-0, 0-9532576-4-9. LCCN ????

Ross:1988:SIC

[Ros88] E. Ross. SGML and in-house CD-ROM publishing software. In Anonymous [Ano88d], pages 619–621 (vol. 2). ISBN ???? LCCN ???? 2 vol.

Ross:1996:UGM

[Ros96] John Ross. The underground guide to Microsoft Internet assistant: slightly askew advice on mastering the Web with Win-Word. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-48944-9. xvii + 174 pp. LCCN QA76.76.H94 R67 1996.

Rosch:1997:DVS

[Ros97] H. Rosch. Does VRML supersedes all 3D-CAD-formats? In Anonymous [Ano97-48], pages 247–254. ISBN ???? ISSN 0943-7207. LCCN ????

Rotter:1997:AIR

[Rot97]  R. Rotter.  Access to in-

formation: a reality in the AGORA project effective dissemination requires effective information management: SGML a universal format? In Anonymous [Ano97y], pages 265–274. ISBN 92-828-0111-X. LCCN CD974.4 .D46 1996.

Rousseau:1995:PW

[Rou95] B. Rousseau. Publishing on the Web. In Vandoni [Van95b], pages 279–293. ISBN 92-9083-076-X. LCCN QA76 .C101 1995.

Rouyer:1998:DHW

[Rou98] Jeff Rouyer. Dynamic HTML web magic. New Riders Publishing, Carmel, IN, USA, 1998. ISBN 1-56830-421-8. xiv + 296 pp. LCCN QA76.76.H94R68 1998.

Reinwald:1998:SOH

[RP98] Berthold Reinwald and Hamid Pirahesh. SQL open heterogeneous data access. SIGMOD Record (ACM Special Interest Group on Management of Data), 27(2):506–507, June 1998. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Robin:1996:TTE

[RPWW96] B. Robin, J. D. Price, J. Willis, and D. A. Willis, editors. Technology and Teacher Education Annual, 1996 Proceedings of SITE 96 - Seventh International Conference of the Society for Information Technology and Teacher Education (SITE). Assoc. Advancement of Comput. Educ, Charlottesville, VA, USA, 1996. ISBN ???? LCCN ????

Rousseau:1994:WDP

[RR94] Bertrand Rousseau and Mario Ruggier. Writing documents for paper and WWW. A strategy based on FrameMaker and WebMaker. Computer Networks and ISDN Systems, 27(2):205– 214, November 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1994&
volume=27&issue=2&aid=1383.

Ray:1996:DH

[RR96a] Deborah Ray and Eric Ray.

Dummies 101: HTML. I D G Books Worldwide, Indianapolis, IN, USA, August 1996. ISBN 0-7645-0032-5. xviii + 222 pp. LCCN QA76.76.H94R388 1996. US$24.99.

Ray:1996:HDQ

[RR96b] Deborah S. Ray and Eric J. Ray. HTML for dummies, quick reference. I D G Books Worldwide, Indianapolis, IN, USA, 1996. ISBN 1-56884-990-7. xiv + 196 pp. LCCN QA76.76.H94R39 1996. US$12.99, CDN$17.99, UK£11.99.

Riva:1996:LSH

[RR96c] Alberto Riva and Marco Ramoni. LispWeb: a specialized HTTP server for distributed AI applications. Computer Networks and ISDN Systems, 28 (7–11):953–961, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1996&
                                        volume=28&issue=7-11&aid=
                                        1615.

Ray:1997:HDQ

[RR97a] Deborah S. Ray and Eric J. Ray. HTML for dummies quick reference. –For dummies. I D G Books Worldwide, Indianapolis, IN, USA, second edition, 1997. ISBN 0-7645-0248-4. x + 230 pp. LCCN QA76.76.H94 R39 1997.

Ray:1997:MH

[RR97b] Deborah S. Ray and Eric J. Ray. Mastering HTML 4.0.

Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1997. ISBN 0-7821-2102-0 (paperback). xxv + 1013 pp. LCCN QA76.76.H94 R39 1997. US$49.99. Includes CD-ROM.

Randolph:1998:CSD

[RR98a] Elaine Randolph and Jeff Randolph. Cascading stylesheets and dynamic HTML. In Anonymous [Ano98u], pages 293– ?? ISBN 1-891709-00-3, 1-891709-01-1. LCCN ????

URL http://www.ocstc.org/
ana_conf/pdf/tt4f.pdf.

Ray:1998:DH

[RR98b] Deborah S. Ray and Eric J.

Ray. Dummies 101. HTML 4. –For dummies. I D G Books Worldwide, Indianapolis, IN, USA, 1998. ISBN 0-7645-0205-0. xx + 266 pp. LCCN QA76.76.H94R3883

1998. URL http://lcweb.loc.
gov/catdir/toc/98-70133.html.

Ray:1998:HDQ

[RR98c] Deborah S. Ray and Eric J. Ray.

HTML 4 for dummies quick reference. –For dummies. I D G Books Worldwide, Indianapolis, IN, USA, 1998. ISBN 0-7645-0332-4. x + 228 pp. LCCN QA76.76.H94R389 1998.

Ray:1999:MHa

[RR99a] Deborah S. Ray and Eric J.

Ray. Mastering HTML 4. Financial Times Management, SYBEX, 1999. ISBN 0-7821-2523-9. xxi + 873 pp. LCCN QA76.76.H94R394 1999.

Ray:1999:MHb

[RR99b] Deborah S. Ray and Eric J. Ray. Mastering HTML 4. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, 1999. ISBN 0-7821-2523-9, 0-7821-2524-7. xxi + 873 pp. LCCN QA76.76.H94R394 1999.

Ray:1999:MHc

[RR99c] Deborah S. Ray and Eric J. Ray. Mastering HTML 4. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, premium edition, 1999. ISBN 0-7821-2524-7. 1188 pp. LCCN QA76.76.H94 R394 1999.

Ramakrishnan:1993:CDD

[RRS^{+}93] Raghu Ramakrishnan, William G. Roth, Praveen Seshadri, Divesh Srivastava, and S. Sudarshan. The CORAL deductive database system. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):544–545 (or 541–545??), June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Roth:1993:MFI

[RS93] G. I. Roth and K. Sullow. The Multimedia Forum — an integrated digital information service. In Raitt and Jeapes [RJ93], pages 263–272. ISBN 0-904933-85-7. LCCN ????

Rose:1996:QQS

[RS96] E. A. Rose and G. A. Schloss.

QSBA: a query, search and browse assistant for images on the Web. In Anonymous [Ano96-35], pages 417– 426. ISBN ???? LCCN ????

Ras:1997:FIS

[RS97] Zbigniew Ras and Andrzej Skowron, editors. Foundations of intelligent systems: 10th international symposium, ISMIS ’97, Charlotte, North Carolina, USA, October 15–18, 1997: proceedings, volume 1325 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63614-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.E95 I582 1997.

Rosen:1999:FMH

[RSV99] B. R. Rosen, R. J. Seitz, and J. Volkmann, editors. Functional mapping of the human brain: International conference; 5th — June 1999, Dusseldorf, Germany, volume 9(2) of NEU-ROIMAGE 1999; number 6. Academic Press, New York, USA, 1999. ISBN ???? ISSN 1053-8119. LCCN ????

Revell:1995:DES

[RT95] Norman Revell and A. Min Tjoa, editors. Database and expert systems applications: 6th international conference, DEXA ’95, London, United Kingdom, September 4–8, 1995: proceedings, volume 978 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. ISBN 3-540-60303-4. LCCN QA267.A1 L43 no.978.

Reed:1998:EHF

[RT98] Duncan Reed and Peter Thomas. Essential HTML fast. The Essential series. Springer-Verlag Inc., New York, NY, USA, 1998. ISBN 3-540-76199-3 (Berlin) (paperback). x + 133 pp. LCCN QA76.76.H94 R44 1998.

Raymond:1996:DRD

[RTW96] D. Raymond, F. Tompa, and D. Wood. From data representation to data model: Metasemantic issues in the evolution of SGML. Computer Standards and Interfaces, 18(1): 25–36, January 1996. CO-DEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Rubinsky:1992:SVT

[Rub92] Yuri Rubinsky. SGML: the vision, the technology, and the business. CALS Expo ’92 tutorials. ????, ????, 1992. ISBN ???? various pp. LCCN ????

Rubinski:1996:BH

[Rub96] Yuri Rubinski. Beyond HTML. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, May 1996. ISBN 0-614-14488-4. ???? pp. LCCN ???? US$29.95.

Rudich:1996:WDC

[Rud96] J. Rudich. Web development on the cheap [software tools]. Inc, 18(9):91–92, ???? 1996. CODEN INCCDU. ISSN 0162-8968.

Rule:1998:DHH

[Rul98] Jeff Rule. Dynamic HTML: the HTML developer’s guide. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-37961-9. xx + 265 pp. LCCN QA76.76.H94R85 1998.

Rule:1999:DHH

[Rul99] Jeff Rule. Dynamic HTML: the HTML developer’s guide. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-37961-9. xx + 265 pp. LCCN QA76.76.H94 R85 1999.

Rung:1995:EIM

[Run95] Steven F. Rung. Electronic imaging of a Macintosh page. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Autools/rung/rung.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Ruskin:1998:NIT

[Rus98] K. J. Ruskin. New Internet technology: HTML tools, Java, teleconferencing, and streaming multimedia. In Ikeda et al. [IDK98], pages 229–234. ISBN 0-444-50008-1. LCCN WO 200 I61-2 1998; RD80.95 .I68 1998.

Russell:1999:HES

[Rus99] Chris Russell. HTML in easy steps. Computer Step, Southam, third edition, 1999. ISBN 1-84078-048-7. 192 pp. LCCN ????

Ruthven:1996:MPF

[Rut96a] Ian Ruthven, editor. Miro ’95: proceedings of the Final Workshop on Multimedia Information Retrieval (MIRO ’95), Glasgow, Scotland, 18–20 September 1995. Springer-Verlag Inc., New York, NY, USA, 1996. ISBN 3-540-14582-6. LCCN QA76.575 .M58 1995.

Rutledge:1996:PHD

[Rut96b] L. Rutledge. Presenting HyTime documents with HTML. In Carlson and Makedon [CM96], pages 611–615. ISBN 1-880094-21-5. LCCN ????

Ruggiero:1994:OHN

[RV94] F. Ruggiero and R. Van Kleij. On-line hypermedia newspapers: an experiment with ‘L’Unione Sarda’. International Journal of Modern Physics C [Physics and Computers], 5(5): 899–906, October 1994. CO-DEN IJMPEO. ISSN 0129-1831 (print), 1793-6586 (electronic).

Reding:1999:HIC

[RV99] Elizabeth Eisner Reding and Sasha Vodnik. HTML: illustrated complete. Illustrated series. Course Technology, Cambridge, MA, USA, 1999. ISBN 0-7600-5842-3. various pp. LCCN ????

Robinson:1990:AS

[RW90] B. Robinson and G. Wu. Applications of SGML. In Raitt [Rai90], pages 75–82. ISBN 0-904933-75-X. LCCN ????

Robinson:1992:AS

[RW92] B. Robinson and G. Wu. Applications of SGML. University Computing, 14(2):53–57, ???? 1992. CODEN UNCOET. ISSN 0265-4385.

Rees:1995:SCT

[RW95a] Michael Rees and Tak Woo. A synchronous collaboration tool for WWW. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/
rees/SynColTol.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Richmond:1995:IAI

[RW95b] Alan Richmond and Nick White.

Implementation of an astrophysics information system on the World-Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Astronomy/richmond/richmond.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Rath:1996:MSW

[RW96a] H. H. Rath and H.-P. Wiedling. Making SGML work: Introducing SGML into an enterprise and using its possibilities in advanced applications. Computer Standards and Interfaces, 18(1):37–53, January 1996. CO-DEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Ready:1996:HHD

[RW96b] Kevin Ready and Janine Warner. Hybrid HTML Design: A Multi Browser HTML Reference. New Riders Publishing, Carmel, IN, USA, May 1996. ISBN 1-56205-617-4. xxvii + 412 pp. LCCN QA76.76.H94R43 1996. US$35.00.

Ressler:1998:MVA

[RW98] S. Ressler and Q. Wang. Making VRML accessible for people with disabilities. In ACM [ACM98], pages 144–148. ISBN 1-58113-020-1. LCCN ????

                                        URL http://www.acm.org/
                                        pubs/contents/proceedings/
                                        assets/274497/. ACM order

number 444980.

Ressler:1997:UVA

[RWBS97] S. Ressler, Q. Wang, S. Bodarky, and C. Sheppard. Using VRML to access manufacturing data. In Spencer and Brutzman [SB97], pages 109–116. ISBN 0-89791-886-X.

LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/253437/.

ACM order number 434973.

Snodgrass:1994:TLS

[SAA^{+}94] Richard T. Snodgrass, Ilsoo Ahn, Gadi Ariav, Don Batory, James Clifford, Curtis E. Dyreson, Ramez Elmasri, Fabio Grandi, Christian S. Jensen, Wolfgang Käfer, Nick Kline, Krishna Kulkarni, T. Y. Cliff Leung, Nikos Lorentzos, John F. Roddick, Arie Segev, Michael D. Soo, and Suryanarayana M. Sripada. TSQL2 language specification. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(1):65–86, March 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Sadun:1997:HTM

[Sad97] Erica Sadun. HTML, template master CD-ROM, 1997. ISBN 1-886801-06-1. Includes CD-ROM.

Sakui:1994:TD

[Sak94] M. Sakui. TEX and database.

Joho kanri = Information & documentation, 37(3):246–256, June 1994. CODEN JOKAAB. ISSN 0021-7298.

Salamone:1996:HPM

[Sal96] S. Salamone. How to put mainframes on the Web. BYTE Magazine, 21(6):53–54, June 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sammons:1995:MIR

[Sam95] N. Sammons. Multi-platform interrogation and reporting with Rscan. In USENIX Association [USE95], pages 75–87. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

Sandoval:1994:SOP

[San94] Victor Sandoval. SGML: un outil pour la gestion électronique de documents. (French) [SGML: A tool for electronic text processing]. Techniques de l’information. Hermes, Paris, France, 1994. ISBN 2-86601-440-5. 174 pp. LCCN ????

Sandewall:1996:TWW

[San96] Erik Sandewall. Towards a world-wide data base. Computer Networks and ISDN Systems, 28(7–11):1513–1522, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://www.

                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1996&
                                        volume=28&issue=7-11&aid=
                                        1628.

Sangin:1999:HCP

[San99] Jean-Pierre Sangin. HTML 4: créez des pages Web qui vendent! Collection Informatique Collection Informatique (Montreal, Quebec). Editions JCI, Sainte-Agathe, PQ, Canada, 1999. ISBN 2-921599-54-6. ???? pp. LCCN ????

Sapwater:1995:W

[Sap95] E. Sapwater. Webbed. Photo Electronic Imaging, 38(7):42–??, July 1, 1995. CODEN PELIE6. ISSN 0146-0153.

Savage:1995:SSU

[Sav95] Dean Savage. Social science uses of Mosaic at the City University of New York. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
SocialSci/savage/savage.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Singh:1995:AM

[SB95] Guy Singh and Roger Binns. An API to Mosaic. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
singh/ixiwww94.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Spencer:1997:PVS

[SB97] Stephen N. Spencer and Don Brutzman, editors. Proceedings VRML 97: second Symposium on the Virtual Reality Modeling Language, Monterey, California, February 24–26, 1997. ACM Press, New York, NY, USA, 1997. ISBN 0-89791-886-X. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/253437/.

ACM order number 434973. Spencer:1999:PVF

[SB99] Stephen N. Spencer and Christian Bauer, editors. Proceedings VRML 99: fourth Symposium on the Virtual Reality Modeling Language, Paderborn, Germany, February 16–19, 1999. ACM Press, New York, NY, USA, 1999. ISBN 1-58113-079-1. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/299246/.

Schwartz:1995:HID

[SBD^{+}95] Michael F. Schwartz, C. Mic Bowman, Peter B. Danzig, Darren R. Hardy, and Udi Manber. The Harvest information discovery and access system. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Searching/schwartz.harvest/
schwartz.harvest.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Shaya:1998:AXJ

[SBGK98] E. J. Shaya, J. H. Blackwell, J. E. Gass, and V. E. Kargatis. Application of XML to journal table archiving. In Anonymous [Ano98b], page 13.06. ISBN ???? ISSN 0002-7537. LCCN ????

Shaya:1999:XAA

[SBGO99] E. Shaya, J. Blackwell, J. Gass, and N. Oliversen. XML at the ADC: Steps to a next generation data archive. In Paulson et al. [PKM99], pages 83.04–?? ISBN 87-16-12177-5. ISSN 0105-3639. LCCN WL 200 A392b 1998.

Savola:1996:UHS

[SBJ96] Tom Savola, Mark Brown, and John Jung. Using HTML: Special Edition. Que Corporation, Indianapolis, IN, USA, second edition, April 1996. ISBN 0-7897-0758-6. xxxiii + 921 pp. LCCN QA76.76.H94U85 1996. US$49.99.

Salvail:1996:WGI

[SBM96] Leon Salvail, Steven K. Besler, and Sacha Michel Mallais. Webmaster’s guide to Internet connectivity. New Riders Publishing, Carmel, IN, USA, 1996. ISBN 1-56205-574-7. ???? pp. LCCN TK5105.875.I57 S355 1996.

Smith:1995:VVTa

[SBS95a] David Smith, Richard Boyd, and Alan Scott. Virtus VRML toolkit: Create 3-D worlds for the Web without programming. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1995. ISBN 1-56830-247-9. xvii + 283 pp. LCCN QA76.76.H94S625 1995. Includes CD-ROM.

Spasojevic:1995:UWA

[SBS95b] Mirjana Spasojevic, Mic Bowman, and Alfred Spector. Using a wide-area file system within the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ???? URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
spasojevic/paper.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Smyth:1998:ACB

[SC98] Barry Smyth and Padraig Cunningham, editors. Advances in case-based reasoning: 4th European workshop, EWCBR-98, Dublin, Ireland, September 23–25, 1998: proceedings, volume 1488 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag Inc., New York, NY, USA, 1998. ISBN 3-540-64990-5 (paperback). LCCN QA267.A1 L43 no.1488.

StLaurent:1999:BXA

[SC99] Simon St. Laurent and Ethan Cerami. Building XML applications. McGraw-Hill, New York, NY, USA, 1999. ISBN 0-07-134116-1, 0-07-134115-3. xx + 551 pp. LCCN QA76.76.H94 S69 1999.

Scheller:1988:DSA

[Sch88] Angela Scheller. Document standards: Availability and products. Computer Networks and ISDN Systems, 16(1–2): 138–142, September 1988. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

Scheller:1990:ESI

[Sch90a] A. Scheller. Einsatz von SGML im Deutschen Forschungsnetz (English: Usage of SGML within the German Research Network DFN). Informationstechnik — IT, 32(3):186–193, June 1990. CODEN ICSAES. ISSN 0179-9738.

Scheller:1990:USW

[Sch90b] A. Scheller. Usage of SGML within the German Research Network DFN. Informationstechnik — IT, 32(3):186–193, June 1990. CODEN ICSAES. ISSN 0179-9738.

Schmidt:1992:WDO

[Sch92] S. Schmidt. Where does open information processing stand? the ODA/ODIF and SGML document standards. Bulletin des Schweizerischen Elektrotechnischen Vereins and des Verbandes Schweizerischer Elektrizitaetswerke, 83(1):31–34, January 1992. CODEN BEVEDP. ISSN 0036-1321.

Schmitz:1993:TTK

[Sch93] K.-D Schmitz, editor. TKE ’93 Terminology and Knowledge Engineering Proceedings Third International Congress on Terminology and Knowledge Engineering. Indeks Verlag, Frankfurt/Main, Germany, 1993. ISBN ???? LCCN ????

Scharf:1995:HVQ

[Sch95a] Dean Scharf. HTML: Visual quick reference. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0411-0. 159 pp. LCCN QA76.76.H94 S255 1995. US$14.99, CDN$20.95.

Schohl:1995:PCP

[Sch95b] W. Schohl. Practical chances of PC-based fulltext databases in smaller publishing houses and press departments. Nachrichten für Dokumentation, 46(1):3–12, January–February 1995. CO-DEN NADOAW. ISSN 0027-7436.

Schuyler:1995:VTL

[Sch95c] Michael Schuyler. The view from the top left corner. at home on the Web with HTML. Computers in Libraries, 15(9):27– ??, October 1, 1995. CODEN CPLIE8. ISSN 1041-7915.

Schwartz:1995:FFL

[Sch95d] Christian Schwartz. Flexible formatting with Linuxdoc-SGML. Linux Journal, 18:??, October 1995. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Scharf:1996:HVQ

[Sch96a] Dean Scharf. HTML: Visual quick reference. Que Corporation, Indianapolis, IN, USA, second edition, 1996. ISBN 0-7897-0786-1. ix + 183 pp. LCCN QA76.76.H94S255 1996. US$16.99, CDN$23.95, UK£15.49. URL http://www.

mcp.com/que.

Schick:1996:SCP

[Sch96b] William Schick. Second commentary on Practical SGML: Explores the practicality of Practical SGML. Asterisk: the journal of computer documentation, 20(2):33–??, May 1, 1996. CODEN ASTRF7. ISSN 0731-1001.

Schulzrinne:1996:WWW

[Sch96c] H. Schulzrinne. World Wide Web: whence, whither, what next? IEEE network, 10(2):10– 17, March–April 1996. CODEN IENEET. ISSN 0890-8044.

Scharf:1997:HVQ

[Sch97a] Dean Scharf. HTML 3.2 Visual Quick Reference. Que Corporation, Indianapolis, IN, USA, third edition, March 1997. ISBN 0-7897-1103-6. ix + 211 pp. LCCN QA 76.76 H94S253 1997. US$19.99.

Schmalstieg:1997:LOL

[Sch97b] D. Schmalstieg. Lodestar: An octree-based level of detail generator for VRML. In Spencer and Brutzman [SB97], pages 125–132. ISBN 0-89791-886-X. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/253437/.

ACM order number 434973.

Schmuller:1998:DHM

[Sch98] Joseph Schmuller. Dynamic HTML: master the essentials. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1998. ISBN 0-7821-2277-9. xxiv + 581 pp. LCCN QA76.76.H94S5272 1998.

Schnoor:19xx:IHJa

[Schxxa] Peter Schnoor. Intranet, HTML og Java. KnowWare, Ordrupvej 63 C, 2920 Charlottenlund, Denmark, 19xx. ISBN ???? ???? pp. LCCN

????  URL http://www.
knowware.dk/dk/dk06.htm.

Schnoor:19xx:IHJb

[Schxxb] Peter Schnoor. Intranet, HTML und Java. KnowWare, Ordrupvej 63 C, 2920 Charlottenlund, Denmark, 19xx. ISBN ???? 48 pp. LCCN ????

URL http://www.knowware.
dk;   http://www.knowware.
dk/de/de06.htm. Nr. 133.

Scott:1996:EVW

[Sco96a] Adrian Scott. Enter the virtual world: Creating client-side VRML with JavaScript. Java Report: The Source for Java Development, 1(3):40–45, July and August 1996. CODEN JREPFI. ISSN 1086-4660. URL http://

www.sigs.com/publications/
docs/java/9607/java9607.toc.
html.

Scott:1996:MJV

[Sco96b]  Adrian Scott.  The mar-

riage of Java and VRML. Java Report: The Source for Java Development, 1(6): ??, November/December 1996. CODEN JREPFI. ISSN 1086-4660. URL http://

www.sigs.com/publications/
docs/java/9611/java9611.scott.
html;   http://www.sigs.com/
publications/docs/java/9611/
java9611.toc.html.

Shelly:1999:HICa

[SCW99] Gary B. Shelly, Thomas J. Cashman, and Denise M. Woods. HTML: introductory concepts and techniques. Course Technology, Cambridge, MA, USA, 1999. ISBN 0-7895-4724-4. various pp. LCCN ????

Sacks-Davis:1996:SIM

[SD96a] R. Sacks-Davis. The structured information manager: A database system for SGML documents. In Vijayaraman et al. [V^{+}96], pages 596–?? ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996.

Stanek:1996:WPU

[SD96b] William Robert Stanek and Steven J. DeRose. Web publishing unleashed: HTML, Java, CGI, VRML, SGML. Howard W. Sams, Indianapolis, IN 46268, USA, March 1, 1996. ISBN 1-57521-051-7. xxii + 916 pp. LCCN QA76.76.H94S73 1996. US$49.99. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1-57521-051-7.html.

Shah:1998:CHP

[SD98] A. Shah and T. Darugar. Creating high performance Web applications using Tcl, display templates, XML, and database content. In USENIX [USE98], pages 121–130. ISBN 1-880446-98-7. LCCN QA76.73.T44 T34 1998.

Sacks-Davis:1995:DSS

[SDAMZ95] R. Sacks-Davis, T. Arnold-Moore, and J. Zobel. Database systems for structured documents. IEICE Transactions on Information and Systems, E78-D(11):1335–1342, November 1995. CODEN ITISEF. ISSN 0916-8532.

Schilit:1996:TLC

[SDK^{+}96] Bill N. Schilit, Fred Douglis, David M. Kristol, Paul Krzyzanowski, James Sienicki, and John A. Trotter. TeleWeb: loosely connected access to the World Wide Web. Computer Networks and ISDN Systems, 28 (7–11):1431–1444, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1996&
                                        volume=28&issue=7-11&aid=

1575. http://garuda.imag.fr/www5cd/www172/overview.htm.

SDSC:1997:VR

[SDS97] VRML repository. San Diego Supercomputer Center, 1997.

                                        URL http://www.sdsc.edu/
                                        vrml.

Sullivan:1995:HME

[SDSS95] Brian Sullivan, Jonathan Ryan Day, Jeff Spitulnik, and Elliot Soloway. HTML made easy:

HTML Claris XTND translator. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Autools/sullivan/HTMLMadeEasy.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Seaman:1994:LAE

[Sea94] D. M. Seaman. ’A library and apparatus of every kind’: the Electronic Text Center at the University of Virginia. Information Technology and Libraries, 13(1):15–19, March 1994. CO-DEN ITLBDC. ISSN 0730-9295.

Seaman:1995:CPS

[Sea95] D. Seaman. Campus publishing in standardized electronic formats — HTML and TEI. In Okerson [Oke95], pages 93–102. ISBN 0-918006-25-2. LCCN Z286.E43 .F55 1995.

Seaman:1996:ETC

[Sea96a] D. Seaman. The Electronic Text Center and on-line archive of electronic texts. In Neubauer [Neu96], pages 55–57. ISBN 3-465-02885-6. ISSN 0514-6364. LCCN Z678.93.E5 E423 1996.

Seaman:1996:SAC

[Sea96b] D. M. Seaman. Selection, access, and control in a library of electronic texts. Cataloging & classification quarterly, 22(3-4):75– 84, ???? 1996. CODEN CC-QUDB. ISSN 0163-9374.

Sedayao:1995:MWK

[Sed95] Jeff Sedayao. “Mosaic will kill my network!” studying network traffic patterns of Mosaic use. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
sedayao/mos_traf_paper.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Seesing:1996:EPW

[See96] P. R. Seesing. Electronic publication on the World-Wide Web: the process and the pitfalls. In Anonymous [Ano96k], pages 145–155. ISBN 0-7803-3689-5 (softbound), 0-7803-3690-9 (casebound), 0-7803-3691-7 (microfiche). LCCN T10.5.I54 1996. IEEE catalog number 96CH36009.

Seidman:1995:HES

[Sei95] James L. Seidman. An HTML extension to support client-side image maps. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
seidman/seidman.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Seidman:1996:RAP

[Sei96] J. Seidman. RFC 1980: A proposed extension to HTML: Client-side image maps, August 14, 1996. URL ftp://ftp.

internic.net/rfc/rfc1980.
txt;   http://www.math.utah.
edu/pub/rfc/rfc1980.txt.

Scheller:1990:JED

[SF90] A. Scheller and C. Fuhrhop.

Joint editing with DAPHNE — a system integrating CGMs into SGML documents. In Vandoni and Duce [VD90], pages 113–124. ISBN 0-444-88683-4. LCCN ????

Sarna:1996:PSB

[SF96] D. Sarna and G. Febish.

Paradigm shift — the business reality of VRML — VRML will have a profound impact on business. Datamation, 42(10):27–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sasaki:1997:HCI

[SFO97] Nozo Sasaki, Mami Fujisaki, and Akihiro Ota. HTML and CGI immun. Yongjin Chulpansa, Soul, Korea, 1-pan. edition, 1997. ISBN 89-314-0732-7. 267 pp. LCCN ????

Silverman:1993:WIA

[SG93] D. Silverman and M. Gross.

What’s so important about data conversion? In Williama [Wil93a], pages 383–385. ISBN 0-938734-73-3. LCCN ????

Sullow:1993:MFI

[SGBO^{+}93] K. Sullow, I. Gabel-Becker, M. Ockenfeld, W. Putz, and G. Roth. MultiMedia Forum: an interactive online journal. Electronic Publishing—Origination, Dissemination, and Design, 6(4):413–422, December 1993. CODEN EPODEU. ISSN 0894-3982.

Srinivas:1995:AMH

[SGJ^{+}95] K. Srinivas, K. Gopinath, V. Jagannathan, R. Karinthi, Matthew Fuchs, Y. V. Reddy, George Almasi, and Tad Davis. Applications of Mosaic in health care delivery. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
MedTrack/srinivas/artemis.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

SGMLAssociates:1987:TSN

[SGM87] <TAG>: the SGML newsletter, 1987. ISSN 1067-9197. BDS, Inc., Sterling, VA, USA.

SGMLSighyper:1991:SSN

[SGM91] SGML SIGhyper newsletter, 1991. Techno Teacher, Inc., Tallahassee, FL.

Smith:1996:EAV

[SGP96] J. R. Smith, R. V. Grimes, and T. A. Plant. Engineering applications of virtual reality. In Bolas et al. [BFM96], pages 332– 339. CODEN PSISDG. ISBN 0-8194-2027-1. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.2653.

Shah:1995:SRS

[Sha95a] Rawn Shah. Show report: SIG-GRAPH 95/Java BOF. Sunworld Online, 09:??, 1995. ISSN 1091-8914. URL http://www.

sun.com/sunworldonline/swol-
09-1995/swol-09-siggraph.
html.

Shapiro:1995:VPF

[Sha95b] Ehud Shapiro. Virtual places — A foundation for human interaction. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
WebProd/shapiro/demo1.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Shaw:1997:ISS

[Sha97] N. Shaw. Integrating SGML, STEP/EXPRESS, HTML, and Java to make product data accessible across the Web. In Anonymous [Ano97h], pages 115–118. ISBN 1-901782-01-8. LCCN ????

Shahi:1998:CVM

[Sha98] Avantika Shahi. A cybertheatre a virtual movie zone designed using VRML and JavaScript. ????, ????, 1998. 44 pp.

Sherman:1988:CRH

[She88] Chris Sherman, editor. The CD ROM Handbook. Intertext Pubs./McGraw-Hill Book Co., New York, NY, USA, 1988. ISBN 0-07-056578-3. xxiii + 510 pp. LCCN TK7882.C56 C332

198. US$59.95.

Sherwood:1995:TSA

[She95] Kaitlin Duck Sherwood. Technical and sociological aspects of developing campus-wide webs: UIUC College of Engineering. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Campus.Infosys/sherwood/sherwood.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Shelly:1996:HI

[She96] Gary B. Shelly. HTML: An Introduction. Course Technology, Cambridge, MA, USA, 1996. ISBN 0-7895-1242-4. 312 pp. LCCN ???? Includes instructional resource package.

Shimojo:1993:MDR

[Shi93]  S. Shimojo.  Multimedia-its

dream and reality: research and products. Journal of the Information Processing Society of Japan = Joho Shori, 34(4):

389–399, April 1993. CODEN JOSHA4. ISSN 0447-8053.

Shih:1997:SUS

[Shi97] Yung-Ming Shih. A study of the usefulness of the structure of HTML documents on retrieval effectiveness. Thesis (m.s.), State University of New York at Binghamton, Thomas J. Watson School of Engineering and Applied Science, Binghamton, NY, USA, 1997. ix + 76 pp.

Shimazu:1998:TCR

[Shi98] H. Shimazu. A textual casebased reasoning system using XML on the World-Wide Web. In Smyth and Cunningham [SC98], pages 274– 285. ISBN 3-540-64990-5 (paperback). LCCN QA267.A1 L43 no.1488.

Stribling:1996:RWC

[SHO^{+}96] D. Stribling, T. Hunter, L. Olszewski, A. Corrigan, R. Mullis, and L. Allen. A real world conversion to SGML. In ACM [ACM96d], pages 75–86. ISBN 0-89791-799-5. LCCN QA 76.9 D6 I56 1996.

Shreve:1995:SRC

[Shr95] G. M. Shreve. SGML representation of concept systems — identifying, tagging and retrieving term-concept structures in textual context. In Wright and Strehlow [WS95b], pages 157– 168. ISBN 0-8031-1984-4. ISSN 0066-0558. LCCN T11 .S6765

Scharf:1995:UMR

[SHW95] Ronald Scharf, Stefan Hartmann, and Werner Wolz. Using Mosaic for remote test system control supports distributed engineering. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

                                        http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/CSCW/
                                        scharf/scharf.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Suzuki:1996:SDM

[SIF^{+}96] K. Suzuki, M. Imamura, Y. Fujii, O. Moriguchi, and Y. Maruta. SGML document management technologies for CALS. Mitsubishi Denki Giho, 70(10):76– 81, ???? 1996. CODEN MTD-NAF. ISSN 0369-2302.

Sigler:1996:HT

[Sig96] Douglas Sigler. HTML toolbox. Internet World, 7(4):51–??, April 1, 1996. CODEN IERNE8. ISSN 1064-3923.

Simpson:1999:JX

[Sim99] John E. Simpson. Just XML. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-943417-8 (paperback). xiv + 381 pp. LCCN QA76.76.H94S57 1998.

                                        URL http://www.phptr.com/
                                        ptrbooks/ptr_0139434178.html.

Singleton:1996:WW

[Sin96] A. Singleton. Wired on the web. BYTE Magazine, 21(1): 77–78, 80, January 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Singh:1997:HTI

[Sin97] D. V. Singh. HTML: A tool for information exchange on World Wide Web. In Moorthy and Mangla [MM97], pages 144–145. ISBN 81-900825-0-7. LCCN ????

Sjolin:1995:WFE

[Sjo95] Martin Sjolin. A WWW front end to an OODBMS. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Databases/sjolin/sjolin.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Simms:1998:THS

[SK98] E. Simms and J. Kagan. Taking an HTML snapshot of a data warehouse: Just say, cheese! In Anonymous [Ano98t], pages 551–556. ISBN 1-58025-149-8. LCCN ????

Skala:1998:WSI

[Ska98] Vaclav Skala, editor. WSCG ’98:

the Sixth International Conference in Central Europe on Computer Graphics and Visualization ’98: February 9–13, 1998, Plzen, Czech Republic: conference proceedings, program & posters, volume 3. University of West Bohemia, Plzen, Czech Republic, 1998. ISBN 80-7082-359-3 (vol. 1) 80-7082-360-7 (vol. 2) 80-7082-361-5 (vol. 3) 80-7082-362-3 (vol. 4). LCCN ???? Four volumes.

Skala:1999:WTI

[Ska99] Vaclav Skala, editor. WSCG ’99: the 7-th International Conference in Central Europe on Computer Graphics, Visualization and Interactive Digital Media ’99: February 8.–12., 1999, Plzen, Czech Republic, volume Four volumes. University of West Bohemia, Plzen, Czech Republic, 1999. ISBN 80-7082-490-5 (set), 80-7082-494-8 (vol. 1), 80-7082-495-6 (vol. 2), 80-7082-496-4 (vol. 3), 80-7082-525-1 (vol. 4). LCCN ????

Skinner:1994:CFP

[Ski94]  R. Skinner.  Cross-platform

formatting programs. Library Software Review, 13(2):152– 156, Summer 1994. CODEN LSREEA. ISSN 0742-5759.

Skok:1996:EII

[Sko96] D. Skok. Evolving imaging into the Internet revolution. Imaging Magazine, 5(1):104–106, 108, 110–112, 114, 118, 120–122, January 1996. CODEN IM-NYE9. ISSN 1063-4320.

Skovira:1998:HHP

[Sko98] Robert J. Skovira. HTML and home page design: an introduction. Easy Learning and Teaching, Bratislava, Slovakia, 1998. ISBN 80-88812-06-2. 112 pp. LCCN ????

Si:1997:PCD

[SL97] A. Si and Hong Va Leong.

On page coherence for dynamic HTML pages. In Allen and Rasmussen [AR97], pages 263– ?? ISBN 0-89791-868-1. LCCN Z 699 A1 A27 1997.

Slater:1994:CW

[Sla94] A. F. Slater. Controlled by the Web. Computer Networks and ISDN Systems, 27(2):289– 295, November 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1994&
volume=27&issue=2&aid=1395.

Slater:1995:EWCb

[Sla95a] A. F. Slater. Extending W3 clients. Computer Networks and ISDN Systems, 28(1–2): 61–??, December 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Slater:1995:EWCa

[Sla95b] Alan Falconer Slater. Extending W3 clients. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
slater/ew3c.html. A selected

few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Schultheiss:1993:OSB

[SM93a] P. M. Schultheiss and H. Messer. Optimal and suboptimal broadband source location estimation. IEEE Trans. Signal Processing, 41(9):2752–2763, September 1993. CODEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic).

Sperberg-McQueen:1993:SGM

[SM93b] C. M. Sperberg-McQueen. The standard generalized markup language (SGML): A brief introduction. Proceedings of the ASIS annual meeting, 30(??): 285, 1993. CODEN PAISDQ. ISSN 0044-7870.

Smith:1994:TTT

[SM94] C. Smith and T. Mayes. Telework training through World Wide Web. In Anonymous [Ano94h], pages 10/1–10/3. ISBN ???? LCCN ????

Salomon:1995:EJB

[SM95] Marc E. Salomon and David C. Martin. An electronic journal browser implemented in the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        MedTrack/salomon/salomon.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Steyer:1998:H

[SM98a] Ralph Steyer and Olivier Manesse. HTML 4.0. Guide de l’utilisateur; 2053. Micro Application, Paris, France, 1998. ISBN 2-7429-1053-0. 391 pp. LCCN ????

Stuhrmann:1998:HHD

[SM98b] Jan Christoph Stuhrmann and Andrea Misera. HTML 4 et HTML dynamique: grand livre. Grand Livre; 2050. Micro Application, Paris, France, 1998. ISBN 2-7429-1050-6. 575 pp. LCCN ????

Schatz:1996:FDC

[SMC^{+}96] B. Schatz, W. H. Mischo, T. W.

Cole, J. B. Hardin, A. P. Bishop, and Hsinchun Chen. Federating diverse collections of scientific literature. Computer, 29(5): 28–36, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sugimoto:1998:ESA

[SMDO98] S. Sugimoto, A. Maeda, M. Dartois, and J. Ohta. Experimental studies on an appletbased document viewer for multilingual WWW documents — functional extension of and lessons learned from multilingual HTML. In Nikolaou and Stephanidis [NS98], pages 199– 214. ISBN 3-540-65101-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ZA4082.E85E25 1998.

Sperberg-McQueen:1995:HMMa

[SMG95a] C. M. Sperberg-McQueen and R. F. Goldstein. HTML to the max: A manifesto for adding SGML intelligence to the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Autools/sperberg-mcqueen/
sperberg.html. A selected few

of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Sperberg-McQueen:1995:HMMb

[SMG95b] C. M. Sperberg-McQueen and Robert F. Goldstein. HTML to the max: A manifesto for adding SGML intelligence to the World-Wide Web. Computer Networks and ISDN Systems, 28 (1–2):3–11, December 1, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1995&
volume=28&issue=1-2&aid=1524.

Sperberg-McQueen:1996:HMM

[SMG96] C. M. Sperberg-McQueen and R. F. Goldstein. HTML to the max: A manifesto for adding SGML intelligence to the World-Wide Web. Computer Networks and ISDN Systems, 28(1):3–??, ???? 1996. CODEN CNISE9.

ISSN 0169-7552 (print), 1879-2324 (electronic).

Sperberg-McQueen:1999:CDH

[SMH99] C. M. Sperberg-McQueen and C. Huitfeldt. Concurrent document hierarchies in MECS and SGML. In Anonymous [Ano99e], pages 29–42. ISBN ???? ISSN 0268-1145. LCCN ????

Smith:1986:ISP

[Smi86a] Joan M. Smith. Implications of SGML for the preparation of scientific publications. The Computer Journal, 29(3):193– 200, June 1986. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Smith:1986:SGM

[Smi86b] Joan M. Smith. The Standard Generalized Markup Language and Related Issues. British National Bibliography Research Fund, London, UK, 1986. ISBN 0-7123-3082-8. 63 pp. LCCN Z286 .E43 S65 1986.

Smith:1987:SGM

[Smi87a]  J. M. Smith.  The stan-

dard generalized markup language (SGML): an application for OSI. In Gerner and Spaniol [GS87], pages 254–259. ISBN 3-540-17213-0. LCCN TK5105.5 .G561 1987.

Smith:1987:SGMc

[Smi87b]  J. M. Smith.  The stan-

dard generalized markup language (SGML): an application for OSI. In Gerner and Spaniol [GS87], pages 254–259. ISBN 3-540-17213-0. LCCN TK5105.5 .G561 1987.

Smith:SGM86

[Smi87c] Joan M. Smith. The Standard Generalized Markup Language and Related Issues. British National Bibliography Research Fund, London, UK, 1987. ISBN 0-7123-3082-8. 63 pp. LCCN Z286 .E43 S65 1986.

Smith:1987:SGMa

[Smi87d] Joan M. Smith. The standard generalized markup language (SGML): guidelines for authors. British national bibliography research fund report; No. 27. Longwood Publishing Group, Wolfeboro, NH, USA, 1987. ISBN 0-7123-3112-3. ix + 66 pp. LCCN QA76.73.S44 S65x 1987.

Smith:SGM87

[Smi87e] Joan M. Smith. The Standard Generalized Markup Language (SGML): guidelines for authors. British Library Publications Sales Unit, Wolfeboro, NH, USA, 1987. ISBN 0-7123-3112-3. ix + 66 pp. LCCN QA76.73.S44 S65x 1987.

Smith:1987:SGMb

[Smi87f] Joan M. Smith. The standard generalized markup language (SGML): guidelines for editors and publishers. British national bibliography research fund report; No. 26. Longwood Publishing Group, Wolfeboro, NH, USA, 1987. ISBN 0-7123-3111-5 (microfiche). ix + 51 pp. LCCN ????

Smith:1989:SGM

[Smi89] J. M. Smith. Standard generalized markup language and related standards. Computer Communications, 12(2):80–84, April 1989. CODEN COCOD7. ISSN 0140-3664 (print), 1873-703X (electronic).

Smith:1991:CEI

[Smi91] J. M. Smith. CALS in Europe:

an integrated information base. Logistics Information Management, 4(1):26–31, ???? 1991. CODEN LINMEC. ISSN 0957-6053.

Smith:1992:SRS

[Smi92a] Joan M. Smith. SGML and Related Standards: Document Description and Processing Languages. The Ellis Horwood series in computers and their applications. Ellis Horwood, New York, NY, USA, 1992. ISBN 0-13-441056-4 (paperback), 0-13-806506-3. xviii + 151 pp. LCCN QA76.73 .S42 1992. US$50.25.

Smith:SRS92

[Smi92b] Joan M. Smith. SGML and Related Standards: Document Description and Processing Languages. Ellis Horwood, New York, NY, USA, 1992. ISBN 0-13-806506-3. xviii + 151 pp. LCCN QA76.73 .S42 1992. US$50.25.

Smith:1993:DEP

[Smi93] M. Smith. DynaText: an electronic publishing system. Computers and the humanities, 27(5-6):415–420, 1993. CODEN CO-HUAD. ISSN 0010-4817.

Smith:1995:VVTb

[Smi95] David Smith. Virtus VRML Toolkit for Macintosh. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, December 1995. ISBN 1-56830-246-0. ???? pp. LCCN ???? US$40.00.

Smith:1996:FCP

[Smi96a] Norman E. Smith. First commentary on Practical SGML: Surveys the programming aspects that this edition omits. Asterisk: the journal of computer documentation, 20(2):30– ??, May 1, 1996. CODEN AS-TRF7. ISSN 0731-1001.

Smith:1996:PGS

[Smi96b]  Norman E. Smith.  Prac-

tical Guide to SGML Filters. Wordware Publishing, Plano, TX, USA, August 1996. ISBN 1-55622-511-3. 400 pp. LCCN QA76.76.H94S63 1996. US$49.95.

Smith:1997:GET

[Smi97a] Gary (Gary A.) Smith. The German electronic textbook, 1997.

                                        URL http://www.wm.edu/CAS/
                                        modlang/grammnu.html.

Smith:1997:PGS

[Smi97b] Norman E. (Norman Earl)

Smith. Practical guide to SGML filters. Wordware Publishing, Plano, TX, USA, 1997. ISBN 1-55622-511-3. xvi + 455 pp. LCCN QA76.76.H94S63 1997.

Smith:1998:PGS

[Smi98] Norman E. (Norman Earl)

Smith. Practical guide to SGML/XML filters. Wordware Publishing, Plano, TX, USA, 1998. ISBN 1-55622-587-3 (paperback). xxv + 842 pp. LCCN QA76.76.H94S635 1998.

Stuhrmann:1998:H

[SMiPM98] Christoph Stuhrmann, Andrea Misera, Joachim informaticien Paul, and Jean-Marc Moster. HTML 4. Grand Livre; 2050. Micro Application, Paris, France, 1998. ISBN 2-7429-1423-4. 888 pp. LCCN ????

Soreide:1995:MARa

[SMM95a] N. N. Soreide, L. McCarty, and D. C. McClurg. Mosaic access to realtime data from the TOGA-TAO array of moored buoys. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
EarSci/soreide/soreide.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Soreide:1995:MARb

[SMM95b] N. N. Soreide, L. C. McCarty, and D. C. McClurg. Mosaic access to real-time data from the TOGA-TAO array of moored buoys. Computer Networks and ISDN Systems, 28(1– 2):189–??, December 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Shatdal:1995:APA

[SN95] Ambuj Shatdal and Jeffrey F. Naughton. Adaptive parallel aggregation algorithms. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):104–114, May 1995. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Szwillus:1996:SBE

[SN96] Gerd Szwillus and Lisa Neal. Structure-Based Editors and Environments. Academic Press, New York, USA, 1996. ISBN 0-12-681890-8. viii + 374 pp. LCCN QA76.9.H85 S89 1996. US$75.00. Discusses editors for Java, HTML, Visual Basic, Visual C, ….

Sager:1996:MLP

[SNLT96] N. Sager, N. T. Nhan, M. Lyman, and L. J. Tick. Medical language processing with SGML display. In Cimino [Cim96], pages 547–551. ISBN 1-56053-208-4. LCCN W 26.5 S978 1996.

Spencer:1995:CIA

[SNS95] Stephan M. Spencer, Max L. Nibert, and Jean-Yves Sgro.

Communicating information about virus structure and biology via the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Educ/spencer/spencer.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Soares:1998:ITV

[Soa98] Nuno Soares. Integration of telepresence and VRML models across computer networks. University of Leeds, School of Computer Studies, Leeds, 1998.

SoftQuad:1991:SP

[Sof91a] SoftQuad, Inc. The SGML Primer. SoftQuad, Inc., Toronto, Ontario, Canada, third edition, 1991. ISBN ???? 36 pp. LCCN QA76.73.S44 S45 1991. SoftQuad’s quick reference guide to the essentials of the standard: the SGML needed for reading a DTD and marked-up documents and discussing them reasonably.

SoftQuad:SP91

[Sof91b] SoftQuad, Inc. The SGML Primer. SoftQuad, Inc., Toronto, Ontario, Canada, 1991. 36 pp. LCCN QA76.73.S44 S45 1991.

BareBonesSoftware:1997:B

[Sof97] Bare Bones Software. BBedit, 1997. Includes CD-ROM.

Soiffer:1997:MPR

[Soi97] Neil Soiffer. MathML: A proposal for representing mathematics in HTML. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 31(3): 44–45, September 1997. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). Poster abstract only. See http:/

/www.w3.org/pub/WWW/TR/WD-

math/ for the proposal details.

Solem:1988:PAC

[Sol88] Ann Solem, editor. Proceedings of the ACM Conference on Document Processing Systems, December 5–9, 1988, Santa Fe, NM, USA. ACM Press, New York, NY, USA, 1988. ISBN 0-89791-291-8. LCCN QA 76.9 T48 A25 1988.

Soo:1995:LMH

[Soo95] Jonathan Soo. Live multimedia over HTTP. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
soo/www94a.html. A selected

few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Sugano:1997:SMA

[SOUH97] H. Sugano, K. Otani, H. Ueda, and S. Hiraiwa. SpaceFusion: A multi-server architecture for shared virtual environments. In Spencer and Brutzman [SB97], pages 51–58. ISBN 0-89791-886-X. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/253437/.

ACM order number 434973.

Sosinsky:1996:WPR

[SP96] Barrie A. Sosinsky and Elisabeth Parker. The Web page recipe book. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-460296-X. ???? pp. LCCN QA76.76.H94 S67 1996.

Schurman:1999:DHA

[SP99a] Eric M. Schurman and William J.

Pardi. Dynamic HTML in action. Microsoft Press, Bellevue, WA, USA, second edition, 1999. ISBN 0-7356-0563-7 (paperback). xv + 497 pp. LCCN QA76.76.H94P37 1999.

Smith:1999:SOX

[SP99b] Howard Smith and Kevin Poulter. Share the ontology in XML-based trading architectures. Communications of the ACM, 42(3):110–111, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/cacm/1999-42-3/p110-
smith/.

Spetka:1995:TRB

[Spe95] Scott Spetka. The TkWWW robot: Beyond browsing. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Agents/spetka/spetka.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Spencer:1999:XDI

[Spe99] Paul Spencer. XML design and implementation: programming with XML, ASP, and IE5. Programmer to programmer. Wrox Press, Chicago, IL, USA, 1999. ISBN 1-86100-228-9. xi + 426 pp. LCCN QA76.76.H64 S63 1999.

Spicer:1996:QEA

[Spi96a] J. Spicer. Quick and easy AF applications using SAS/ ASSIST software, SAS data views, block menus and HTML. In Anonymous [Ano96-36], pages 764–769 (vol. 1). ISBN ???? LCCN ???? Two volumes.

Spivak:1996:SP

[Spi96b] Jeffrey Spivak. The SGML Primer. DF - Computer Applications Series. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, 1996. ISBN 0-7895-0194-5. LCCN ???? US$13.95.

Scott:1996:HII

[SPR^{+}96a] Richard Scott, Tim Poulsen, Sue Reber, Gail Sandler, and Logical Operations Staff. HTML 3.2 for the Internet and Intranets. Ziff-Davis Press, Emeryville, CA, USA, October 1996. ISBN 1-56276-493-4. xviii + 246 pp. LCCN QA76.76.H94 S3 1996. US$29.99.

Spring:1996:RMD

[Spr96b] M. B. Spring. Reference model for data interchange standards. Computer, 29(8):87–88, August 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Slone:1993:CAN

[SR93] B. J. Slone, III and C. E.

Richardson. A call for the adoption of nuclear utility information standards. Transactions of the American Nuclear Society, 68(pt.A):377–??, ???? 1993. CODEN TANSAO. ISSN 0003-018X.

Stauffer:1996:UH

[SR96] Todd Stauffer and Neil Randall.

Using HTML. Que Corporation, Indianapolis, IN, USA, second edition, December 1996. ISBN 0-7897-0985-6. xxii + 374 pp. LCCN QA76.76.H94R37 1996b. US$24.99.

Schengili-Roberts:1997:AHC

[SR97] Keith Schengili-Roberts. The advanced HTML companion. AP Professional, Boston, MA, USA, 1997. ISBN 0-12-623540-6. xvii + 387 pp. LCCN QA76.76.H94S257 1997. US$29.95. Tips and tricks for even the most experienced Web designers, extensive coverage of programming techniques involving HTML, including CGI, JavaScript and VB-Script, how to use HTML with style sheets, Java, ActiveX, plugins and other goodies.

Spencer:1998:PVT

[SR98] Stephen N. Spencer and Bernie Roehl, editors. Proceedings VRML 98: third Symposium on the Virtual Reality Modeling Language, Monterey, California, February 16–19, 1998. ACM Press, New York, NY, USA, 1998. ISBN 1-58113-022-8. LCCN ???? URL http://

                                        www.acm.org/pubs/contents/
                                        proceedings/graph/271897/.

Sraeel:1996:PCL

[Sra96] H. Sraeel. Page Call links CSRs with Web customers [customer service representatives]. Bank Systems + Technology, 33(4):27, April 1996. CODEN BSYTEE. ISSN 1045-9472.

Schwabe:1996:SHA

[SRB96] D. Schwabe, G. Rossi, and S. D. J. Barbosa. Systematic hypermedia application design with OOHDM. In ACM [ACM96b], pages 116–128. ISBN 0-89791-778-2. LCCN QA76.76.H94 A26 1996. ACM order number: 614960.

Schengili-Roberts:1998:AHC

[SRSC98] Keith Schengili-Roberts and Kim Silk-Copeland. The advanced HTML companion. AP Professional, Boston, MA, USA, second edition, 1998. ISBN 0-12-623542-2 (paperback). xvi + 646 pp. LCCN QA76.76.H94S257 1998.

Smith:1988:SUG

[SS88] Joan M. Smith and Robert Stutely. SGML: The User’s Guide to ISO 8879. Ellis Horwood series in computers and their applications. Ellis Horwood and Halsted Press, New York, NY, USA and New York, USA, 1988. ISBN 0-470-21126-1 (Halsted Press), 0-7458-0221-

4. 175 pp. LCCN QA76.73.S44 S441 1988. UK£22.50.

Salvendy:1989:DUH

[SS89] G. Salvendy and M. J. Smith, editors. Designing and Using Human-Computer Interfaces and Knowledge Based Systems. Proceedings of the Third International Conference on Human Computer Interaction. Vol. II. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1989. ISBN 0-444-88078-X. LCCN QA76.9.H85 I59 1989 v.2.

Slavney:1995:PSD

[SS95a] Susan Slavney and Thomas C.

Stein. Planetary science data distribution on the Web: The Magellan standard products catalog. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Astronomy/slavney/slavneystein.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Smeaton:1995:ISW

[SS95b] Calum Smeaton and Alan Slater. Integrating simulations and W3 courseware. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ???? URL

                                        http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Educ/
                                        smeaton/ise_www/ise_www94.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Smithson:1995:ICS

[SS95c] Brian Smithson and Barbara Singer. An information clearinghouse server for industry consortia. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        CorInfSys/smithson/smithson.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Summitt:1995:CCW

[SS95d] Paul M. Summitt and Mary J. Summitt. Creating Cool 3D Web Worlds with VRML. I D G Books Worldwide, Indianapolis, IN, USA, December 1995. ISBN 1-56884-796-3. xxiv + 520 pp. LCCN QA76.76.H94S85 1995. US$34.99.

Soergaard:1997:PSW

[SS97] P. Soergaard and T. I. Sandahl. Problems with styles in word processing: A weak foundation for electronic publishing with SGML. In El-Rewini et al. [ERPNS97b], pages 137– 147. ISBN 0-8186-7734-1, 0-8186-7862-3. ISSN 1060-3425. LCCN ????

Shewale:1998:HHT

[SS98] N. G. Shewale and M. N. Shewale. Hypertext and HTML tools and techniques to develop Web pages. In Murthy et al. [MDK98], pages 213–220. ISBN 81-85216-31-2. LCCN Z845.I4 A55 1997.

Segbert:1997:ICL

[SSB97] M. Segbert, Katarina Steinwachs, and Peter Burnett, editors. International Conference on Library Automation in Central and Eastern Europe. April 10– 13, 1996, Budapest, Hungary, volume 17551 EN of EUR. European Commission, Luxembourg, Luxembourg, 1997. ISBN 92-828-0305-8. LCCN ????

Steyer:1998:HHD

[SSB98] Ralph Steyer, Celine Stoll, and Hubert Bertrand. HTML 4 [ et HTML dynamique]. PC Poche; 2039. Micro Application, Paris, France, 1998. ISBN 2-7429-1039-5. 440 pp. LCCN ????

Salvendy:1987:SES

[SSH87] Gavriel Salvendy, Steven L.

Sauter, and Joseph J. Hurrell, Jr., editors. Social, Ergonomic and Stress Aspects of Work with Computers. Proceedings of the Second International Conference on Human-Computer Interaction. Vol.1. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1987. ISBN 0-444-42847-X. LCCN QA76.9.I58 I593 1987.

Salvendy:1997:DCS

[SSK97] Gavriel Salvendy, Michael J. Smith, and Richard J. Koubek, editors. Design of computing systems: proceedings of the Seventh International Conference on Human-Computer Interaction (HCI International ’97), San Francisco, California, USA, August 24–29 1997, volume 21 of Advances in human factors/ergonomics. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1997. ISBN 0-444-82183-X. ISSN 0921-2647. LCCN TA166 .A38 v.21A.

StLaurent:1997:DHP

[St.97a] Simon St. Laurent. Dynamic HTML: a primer. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1997. ISBN 1-55828-569-5. xv + 330 pp. LCCN QA76.76.H94 S7 1997.

StLaurent:1997:XP

[St.97b] Simon St. Laurent. XML: a primer. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1997.

ISBN 1-55828-592-X. ???? pp. LCCN QA76.76.H94 S72 1997.

StLaurent:1998:XP

[St.98] Simon St. Laurent. XML: a primer. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1998. ISBN 1-55828-592-X. xix + 348 pp. LCCN QA76.76.H94S72 1997.

StLaurent:1999:IXD

[St.99a] Simon St. Laurent. Inside XML DTDs: scientific and technical. Enterprise computing. McGraw-Hill, New York, NY, USA, 1999. ISBN 0-07-134621-X. xii + 466 pp. LCCN QA76.76.H94 S7 1999.

StLaurent:1999:XP

[St.99b] Simon St. Laurent. XML: a primer. IDG Books, San Mateo, CA, USA, second edition, 1999. ISBN 0-7645-3310-X. xxv + 429 pp. LCCN QA76.76.H94 S72 1999.

DSR:1989:SOS

[sta89] Dansk standardiseringsraad.

SGML and ODA: standardiseringens resultater, 2. del: standards for document processing and interchange. DS/INF; 48 DS-informationsblad; 48. Dansk standardiseringsraad, København, Danmark, 1989. ISBN 87-7310-055-2. 108 pp. LCCN ????

Stauffer:1996:HE

[Sta96] Todd Stauffer. HTML by Example. Que Corporation, Indianapolis, IN, USA, June 1996. ISBN 0-7897-0812-4. xxiv + 548 pp. LCCN QA76.76.H94S74 1996. US$34.99.

Stauffer:1997:HWP

[Sta97] Todd Stauffer. HTML Web publishing 6-in-1. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1407-8 (paperback). xxii + 438 pp. LCCN QA76.76.H94S743 1997.

Sakai:1998:LST

[STA98a] N. Sakai, A. Takasu, and J. Adachi. Logical structure transformation between SGML documents. In Wakayama et al. [W^{+}98a], pages 183–198. ISBN 0-7923-8059-2. LCCN T58.64 .I46 1998.

Stachowiak:1998:AES

[Sta98b] Maciej Stachowiak. Automated extraction of structured data from HTML documents. Thesis (m.eng.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, ????, 1998. 45 pp.

Stein:1995:IGR

[Ste95a] T. I. Stein, editor. 1995 International Geoscience and Remote Sensing Symposium, IGARSS ’95. Quantitative Remote Sensing for Science and Applications. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2568-0, 0-7803-2567-2 (paperback).

LCCN QE 33.2 R4 I58 1995. Three volumes. IEEE catalog no. 95CH35770.

Stern:1995:ESH

[Ste95b] David Stern. Expert systems:

HTML, the WWW, and the librarian. Computers in Libraries, 15(4):56–58, April 1995. CO-DEN CPLIE8. ISSN 1041-7915.

Stern:1995:SDB

[Ste95c] David Stern. SGML documents:

A better system for communicating knowledge. Special Libraries, 86(2):117–124, Spring 1995. CODEN SPLBAN. ISSN 0038-6723.

Stewart:1995:VWW

[Ste95d]  Joe Stewart.  VRML and

the World Wide Web. Dr. Dobb’s Developer Update, 2(6): 2–3, June 1995. CODEN ???? ISSN 1079-8595.

Stertzbach:1996:RPS

[Ste96a] L. A. Stertzbach. Researching and planning an SGML implementation. In Anonymous [Ano96b], pages 239–240. ISBN 0-914548-89-1, 0-914548-90-5. LCCN ????

Stewart:1996:CYO

[Ste96b] Doshia Stewart, editor. Creating Your Own VRML Web Pages. Que Corporation, Indianapolis, IN, USA, June 1996. ISBN 0-7897-0731-4. ???? pp. LCCN ???? US$29.99.

Stewart:1996:UV

[Ste96c] Doshia Stewart, editor. Using VRML. Que Corporation, Indianapolis, IN, USA, February 1996. ISBN 0-7897-0730-6. ???? pp. LCCN ???? US$24.99.

Stewman:1996:AIR

[Ste96d] J. H. Stewman, editor. Artificial intelligence: Research symposium; 9th — May 1996, Key West, FL, Proceedings of the Florida Artificial Intelligence Research Symposium 1996; Conf 9. Florida AI Research Society, ????, 1996. ISBN 0-9620173-8-8. LCCN ????

Stewman:1996:PNF

[Ste96e] J. H. Stewman, editor. Proceedings of the Ninth Florida Artificial Intelligence Research Symposium, FLAIRS-96. Eckerd College, St. Petersburg, FL, USA, 1996. ISBN ???? LCCN ????

Steyer:1998:TSH

[Ste98] Ralph Steyer. Todo sobre HTML 4. Marcombo, Barcelona, Spain, 1998. ISBN 84-267-1152-9. 392 pp. LCCN ????

Stickels:1994:SPP

[Sti94] T. C. Stickels. SGML — A pilot project for the Department of Energy. In Anonymous [Ano94r], pages 246–248. ISBN 0-914548-79-4, 0-914548-80-8. LCCN T 10.5 S63 1994.

Strehlow:1994:USR

[STM^{+}94] Richaerd A. Strehlow, Thomas O. Tallant, James D. Mason, Philip L. Kienlen, and Karen T. Barry. The use of SGML for retrieval of chemical data. In Lysakowski and Gragg [LG94b], pages 138–145. ISBN 0-8031-1876-7. ISSN 0066-0558. LCCN QD8 .C66 1994.

Stiles:1997:AVI

[STM97] R. Stiles, S. Tewari, and M. Mehta. Adapting VRML 2.0 for immersive use. In Spencer and Brutzman [SB97], pages 75–82. ISBN 0-89791-886-X. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/253437/.

ACM order number 434973.

Stiles:1998:AVF

[STMM98] R. Stiles, S. Tewari, M. Mehta, and L. McCarthy. Adapting VRML for free-form immersed manipulation. In Spencer and Roehl [SR98], pages 89–94. ISBN 1-58113-022-8. LCCN ???? URL http://

www.acm.org/pubs/contents/
proceedings/graph/271897/.

Shklar:1995:III

[STMS95] Leon Shklar, Satish Thatte, Howard Marcus, and Amit Sheth. The InfoHarness information integration platform. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Searching/shklar/shklar.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Stonebraker:1993:MD

[Sto93] Michael Stonebraker. The Miro DBMS. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):439, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Strom:1995:EPD

[Str95] Karen M. Strom. Electronic publication and data distribution for the star formation group in the five college astronomy department. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Astronomy/strom/strom.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Strom:1996:TTI

[Str96] D. Strom. Top ten interesting trends on and around the World-Wide Web. ConneXions, 10(2): 9–13, February 1996. CODEN CONNFR. ISSN 0894-5926.

Stutely:1987:SGM

[Stu87] Robert Stutely. The standard generalised markup language. In Earnshaw [Ear87], pages 176– 189. ISBN 0-387-96527-0, 3-540-96527-0. LCCN Z286.E43 W67 1987. US$29.50.

Su:1996:PTI

[Su96] Stanley Y. W. Su, editor. Proceedings of the twelfth International Conference on Data Engineering, February 26–March 1, 1996, New Orleans, Louisiana. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7240-4. LCCN QA76.9.D3 I5582 1996. IEEE Catalog No. 96CB35888.

Sugimoto:1996:CTE

[Sug96] S. Sugimoto. Component technologies and environment for digital libraries. Joho-Shori (J. Information Processing Soc. Japan), 37(9):820–825, September 1996. CODEN JOSHA4. ISSN 0447-8053.

Sullow:1996:AMD

[Sul96] K. Sullow. AMPHORE — a movie documentation workbench. Nachrichten für Dokumentation, 47(2):67–74, March– April 1996. CODEN NADOAW. ISSN 0027-7436.

Summers:1989:MCC

[Sum89] R. J. Summers. Macro CALS capabilities survey. In Anonymous [Ano89c], pages 449–461 (vol. 2). ISBN ???? LCCN ???? 3 vol.

Summerfield:1994:STL

[Sum94] H. Summerfield. SGML transformation in the land of the word processor. In Anonymous [Ano94e], pages 1215– 1259. ISBN 1-871802-32-6. LCCN ????

Sundsted:1997:HJC

[Sun97] Todd Sundsted. How-to Java: 3D computer graphics: Getting the hang of VRML. JavaWorld: IDG’s magazine for the Java community, 2(8), August 1997. CODEN ???? ISSN 1091-8906.

URL http://www.javaworld.
com/javaworld/jw-08-1997/
jw-08-howto.html.

Sun:1998:GMD

[Sun98a] J. F. Sun. A generalized macrobased data reporting system to produce both HTML and text files. In Anonymous [Ano98t], pages 827–832. ISBN 1-58025-149-8. LCCN ????

Sun:1998:UHF

[Sun98b] James Sun. Update HTML formatted SAS output within Web browser. In Anonymous [Ano98t], pages 899– ?? ISBN 1-58025-149-8. LCCN ???? URL http:/

/www2.sas.com/proceedings/
sugi23/Internet/p160.pdf.

Selvaraj:1998:IIC

[SV98] Henry Selvaraj and Brijesh Verma, editors. ICCIMA 1998, International Conference on Computational Intelligence and Multimedia Applications 1998, February 1998, Gippsland, Australia. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1998. ISBN 981-02-3352-3. LCCN QA76.9.S63 I56 1998.

Shelly:1999:HICb

[SVC99] Gary B. Shelly, Misty Vermaat, and Thomas J. Cashman. HTML: introductory concepts and techniques. Course Technology, Cambridge, MA, USA, 1999. ISBN 0-7895-4724-4. various pp. LCCN ????

Sieburg:1995:CWS

[SVOM95] Hans B. Sieburg, Stella Veretnik, Jay M. Otero, and Donald E. Mosier. Cybermouse WWW simulation server. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
MedTrack/sieburg/sieburg.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Swaine:1996:PPL

[Swa96] Michael Swaine. Programming paradigms: Looking for an HTML book. Dr. Dobb’s Journal of Software Tools, 21(2):117– 119, February 1996. CODEN DDJOEB. ISSN 1044-789X.

Schmitz:1995:DS

[SWC^{+}95] John Schmitz, Marsha Woodbury, Mary Connors, Aaron Buckley, and Chip Aubry. Discovery system. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/
schmitz.discover/conference.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Shelly:1999:HCC

[SWC99] Gary B. Shelly, Denise Woods, and Thomas J. Cashman. HTML complete concepts and techniques. Shelly Cashman series. Course Technology, Cambridge, MA, USA, 1999. ISBN 0-7895-4723-6. 448 pp. LCCN ????

Savola:1995:SEU

[SWH95] Tom Savola, Alan Westenbroek, and Joseph Heck. Special edition using HTML. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0236-3. xxi + 673 pp. LCCN QA76.76.H94 S25 1995. US$39.99. Includes CD-ROM.

AdobeSystems:1997:AFS

[Sys97]  Adobe Systems.  Adobe

FrameMaker + SGML, 1997. Includes CD-ROMs and quick reference card.

Szillat:1995:SPE

[Szi95] Horst Szillat. SGML — eine praktische Einfúhrung (English: SGML — a practical introduction). International Thomson Publishing, 5101 Madison Road, Cincinnati, OH 45227, USA; Bonn, Germany, 1995. ISBN 3-929821-75-3. xvi + 226 pp. LCCN ???? DM79.00; oeS 616.00; sFr 77.50.

Soreide:1995:ARD

[SZM^{+}95] N. N. Soreide, W. H. Zhu, D. C.

McClurg, M. J. McPhaden, D. W. Denbo, and M. W. Renton. Access to real-time data from the TAO array of moored buoys via Unix/X-Windows or the World Wide Web and Mosaic. In IEEE [IEE95b], pages 360–365 (vol. 1). ISBN 0-933957-14-9. LCCN TC 1505 O33197 1995. Three volumes. IEEE Catalog No. 95CH35870.

Tittel:1996:CB

[T^{+}96] Ed Tittel et al. CGI bible. I D G Books Worldwide, Indianapolis, IN, USA, second edition, November 1996. ISBN 0-7645-8016-7. xxxi + 618 pp. LCCN TK5105.565.C45 1996. US$39.99.

Tittel:1997:BVW

[T^{+}97] Ed Tittel et al. Building VRML worlds. Osborne/McGraw-Hill, Berkeley, CA, USA, 1997. ISBN 0-07-882233-5. xxiii + 381 pp. LCCN QA76.76.H94B85 1997.

Tung:19xx:HCJ

[T^{+}xx] Wu Shiu Tung et al. HTML, CGI, Java, and VB Script. Run PC, ??, Taiwan, 19xx. ISBN 957-99138-0-3 (??invalid checksum??). ???? pp. LCCN ???? 450 NT. Includes CD-ROM.

Tokunaga:1999:RTI

[TA99] H. Tokunaga and J.-I. Aoe. Research of the technical interpretive article database by XML.

In Anonymous [Ano99f], pages 281–284. ISBN ???? LCCN ????

Tabata:1996:WDL

[Tab96a] K. Tabata. What’s a digital library? Joho-Shori (J. Information Processing Soc. Japan), 37(9):814–819, September 1996. CODEN JOSHA4. ISSN 0447-8053.

Tatge:1996:QEH

[TAB96b] P. K. Tatge, K. Auten, and A. Balaban. Quick and easy HTML. In Anonymous [Ano96b], pages 268–?? ISBN 0-914548-89-1, 0-914548-90-5. LCCN ????

Theodoulidis:1994:OTD

[TABA^{+}94] Babis Theodoulidis, Aziz Ait-Braham, George Andrianopoulos, Jayant Chaudhary, George Karvelis, and Simon Sou. The ORES temporal database management system. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):511, June 1994. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Takada:1996:IMW

[Tak96] T. Takada. Internationalization and multilingualism of the World-Wide Web. NTT R and D, 45(2):167–172, ???? 1996. CODEN NTTDEC. ISSN 0915-2326.

Tanaka:1991:SS

[Tan91] Y. Tanaka. Standardization of SGML. Journal of the Information Processing Society of Japan = Joho Shori, 32(10): 1118–1125, ???? 1991. CODEN JOSHA4. ISSN 0447-8053.

Taniguchi:1998:EPC

[Tan98] M. Taniguchi. Event processing for complicated routes in VRML 2.0. In Spencer and Roehl [SR98], pages 83–88. ISBN 1-58113-022-8. LCCN ????

URL http://www.acm.org/
pubs/contents/proceedings/
graph/271897/.

Tatters:1996:TYN

[Tat96] Wes Tatters. Teach yourself Netscape Web publishing in a week. Sams.net Pub., Indianapolis, IN, USA, 1996. ISBN 1-57521-068-1. xxvi + 634 pp. LCCN TK5105.883.N48 T38 1996. URL http://

merchant.superlibrary.com:
8000/catalog/hg/PRODUCT/PAGE/
15752/bud/1-57521-068-1.html.

Tauber:1996:RFP

[Tau96] J. Tauber. Resolving formal public identifiers on the WWW: A proposal for delegating SGML open catalogues. In Debreceny and Ellis [DE96], pages 443–446. ISBN 1-875855-16-5. LCCN ????

Taubin:1997:GCV

[Tau97] Gabriel Taubin. Geometry coding and VRML. Research report RC. International Business Machines Corporation. Research Division ; 20925 International Business Machines Corporation. Research Division. Research report ; RC 20925. RC 20925 (92674), IBM T.J. Watson Research Center, Yorktown Heights, NY, USA, July 21, 1997. 28 pp.

Taylor:1990:DPY

[Tay90] M. Taylor. Desktop publishingwhat you need to know. In Anonymous [Ano90d], pages 10/1–25. ISBN 92-835-0550-6. LCCN TL500.N6 C6 no.466.

Taylor:1995:EKB

[Tay95a] A. Taylor. Extracting knowledge from biological descriptions. In Mars [Mar95a], pages 114–120. ISBN 90-5199-217-3. LCCN Q334 .T737 1995.

Taylor:1995:UPB

[Tay95b] A. Taylor. Using Prolog for biological descriptions. In Roth and Marien [RM95b], pages 587–597. ISBN 0-9525554-0-9. LCCN QA76.73.P76 I58 1995.

Taylor:1995:CCWa

[Tay95c] Dave Taylor. Creating cool Web pages with HTML. IDG Books, San Mateo, CA, USA, 1995. ISBN 1-56884-454-9. xix + 244 pp. LCCN QA76.76.H94 T39 1995. US$19.99.

Taylor:1995:CCWb

[Tay95d] Dave Taylor. Creating cool Web pages with HTML. IDG Books, San Mateo, CA, USA, Macworld authorized edition, 1995. ISBN 1-56884-705-X. xx + 280 pp. LCCN QA76.76.H94 T392 1995.

Taylor:1995:CCWc

[Tay95e] Dave Taylor. Creating Cool Web Pages with HTML. I D G Books Worldwide, Indianapolis, IN, USA, second edition, December 1995. ISBN 1-56884-822-6. xxiv + 294 pp. LCCN QA76.76.H94T392 1995b. US$24.99.

Taylor:1996:MCC

[Tay96] Dave Taylor. Macworld Creating Cool HTML 3 Web Pages. I D G Books Worldwide, Indianapolis, IN, USA, December 1996. ISBN 0-7645-4013-0. xxi + 357 pp. LCCN QA76.76.H94T393 1997. US$34.99.

Taylor:1997:CCW

[Tay97a] Dave Taylor. Creating Cool Web Pages with HTML 3.2. I D G Books Worldwide, Indianapolis, IN, USA, third edition, 1997. ISBN 0-7645-3040-2. xvi + 343 pp. LCCN QA76.76.H94T38 1997. US$34.99.

Taylor:1997:MCC

[Tay97b] Dave Taylor. Macworld creating cool HTML 3.2 Web pages. IDG Books, San Mateo, CA, USA, Macworld authorized edition, 1997. ISBN 0-7645-4013-0. xxi + 357 pp. LCCN QA76.76.H94T393 1997.

Taylor:1998:CCH

[Tay98] Dave Taylor. Creating cool HTML 4 web pages. I D G Books Worldwide, Indianapolis, IN, USA, 1998. ISBN 0-7645-3201-4. xxvi + 433 pp. LCCN QA76.76.H94T385 1998.

Tescher:1996:IIL

[TB96a] Andrew G. Tescher and V. Michael Bove, Jr., editors. Integration Issues in Large Commercial Media Delivery Systems: 23– 24 October 1995, Philadelphia, Pennsylvania, volume 2615 of Proceedings of the SPIE — The International Society for Optical Engineering. SPIE Optical Engineering Press, Bellingham, WA, USA, 1996. CODEN PSISDG. ISBN 0-8194-1979-6. ISSN 0277-786X (print), 1996-756X (electronic). LCCN QA76.575 .I576 1996.

Thistlewaite:1996:AFH

[TB96b] Paul Thistlewaite and Steve Ball. Active FORMs [in HTML]. Computer Networks and ISDN Systems, 28(7–11): 1355–1364, May 1, 1996. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1996&
volume=28&issue=7-11&aid=
1596.

Tayeb-Bey:1997:GFD

[TBS97] S. Tayeb-Bey and A. S. Saidi. Grammatical formalism for document understanding system: From documents towards HTML text. Lecture Notes in Computer Science, 1339:165– 175, 1997. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Thioulouse:1996:NWW

[TC96] J. Thioulouse and F. Chevenet.

NetMul, a World-Wide Web user interface for multivariate analysis software. Computational Statistics and Data Analysis, 21(3):369–372, March 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Thomas:1992:RAA

[TCLV92] R. Thomas, S. M. Cornick, D. A.

Leishman, and D. J. Vanier. Research at the Advanced Construction Technology Laboratory. In Grierson et al. [GRA92], pages 35–51. ISBN 1-85166-787-3. LCCN TA174.I469 1992.

Thurber:1995:DSK

[TD95] Bonnie Thurber and Bob Davis.

Designing a server with a K–8 school in mind. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/
thurber/thurber.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Turner:1995:RMF

[TDT95] Ronald C. Turner, Timothy A.

Douglass, and Audrey J. Turner. Read Me First: SGML for Writers and Editors. The Charles Goldfarb series on open information management. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, August 1995. ISBN 0-13-432717-9. xi + 241 pp. LCCN QA76.76.H94T87 1996. US$33.75.

Teague:1997:HPH

[Tea97a] Jason Cranford Teague. How to Program HTML Frames: Interface Design and JavaScript. Ziff-Davis Press, Emeryville, CA, USA, 1997. ISBN 1-56276-495-0. xii + 431 pp. LCCN QA76.76.H94T43 1997. US$39.99.

Teasdale:1997:BCS

[Tea97b] S. Teasdale, editor. British Computer Society: Primary Health Care Specialist Group: Annual conference — September 1997, Cambridge, Proceedings of the Annual Conference — Primary Health Care Specialist Group of the British Computer Society. Primary Health Care Specialist Group, British Computer Society; 1997, ????, 1997. ISBN ???? LCCN ????

Teague:1998:DWW

[Tea98a] Jason Cranford Teague. DHTML for the World Wide Web. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1998. ISBN 0-201-35341-5 (paperback). viii + 248 pp. LCCN QA76.76.H94 T428 1998. US$17.95.

Teasdale:1998:BCS

[Tea98b] S. Teasdale, editor. British Computer Society: Primary Health Care Specialists Group: Annual conference — September 1998, Cambridge, Proceedings of the Annual Conference — Primary Health Care Specialist Group of the British Computer Society. Primary Health Care Specialist Group, of the British Computer Society; 1998, ????, 1998. ISBN ???? LCCN ????

Tennant:1995:VLF

[Ten95] R. Tennant. The virtual library foundation: staff training and support. Information Technology and Libraries, 14(1):46–49, March 1995. CODEN ITLBDC. ISSN 0730-9295.

Tennant:1996:PHA

[Ten96] Roy Tennant. Practical HTML:

A Self-Paced Tutorial. Library Solutions Press, 2137 Oregon St., Berkeley, CA, USA, 1996. ISBN 1-882208-19-6. xii + 106 pp. LCCN QA76.76.H94T46 1996. US$54.00.

Tentner:1999:HPC

[Ten99] A. Tentner, editor. High performance computing: HPC ’99: Conference; 7th — April 1999, San Diego, CA, Proceedings of the High Performance Computing Symposium 1999; 7th. Society for Computer Simulation, ????, 1999. ISBN 1-56555-166-4. LCCN ????

Tessier:1996:SDB

[Tes96a] Tom Tessier. Sharing data between Web page frames using JavaScript: Implementing a hidden-frame technique. Dr. Dobb’s Journal of Software Tools, 21(5):72, 74–75, 86– 87, May 1, 1996. CODEN DDJOEB. ISSN 1044-789X.

Tessier:1996:UJC

[Tes96b] Tom Tessier. Using JavaScript to create interactive Web pages. Dr. Dobb’s Journal of Software Tools, 21(3):84, 86–89, 96–97, March 1996. CODEN DDJOEB. ISSN 1044-789X.

Tewari:1995:ICD

[Tew95] Rajiv Tewari. Innovative courseware development with Mosaic and WWW. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

                                        http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/Educ/
                                        tewari/tewari.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

TPWD:1999:NE

[Tex99] Texas Parks and Wildlife Dept. Nature and the environment,

                                        1999. URL http://www.tpwd.
                                        state.tx.us/nature/nature.
                                        htm.

Takahashi:1995:MDD

[TF95] Z. Takahashi and Y. Fukada. Manual development and development support system. Fujitsu, 46(5):493–503, ???? 1995. CODEN FUJTAR. ISSN 0016-2515.

Tittel:1996:WPS

[TGHE96] Ed Tittel, Mark Gaither, Sebastian Hassinger, and Mike Erwin. Web programming Secrets with HTML, CGI, and Perl. I D G Books Worldwide, Indianapolis, IN, USA, 1996. ISBN 1-56884-848-X. xxxiii + 828 pp. LCCN QA76.76.H94W44 1996. US$44.99, CDN$62.99. Includes CD-ROM.

Thimbleby:1991:LTT

[Thi91] H. Thimbleby. Low-tech TEX?

(typesetting program). In Proceedings of the Fourth Annual Conference on Computers and the Writing Process, pages 124– 130. Univ. Sussex, Brighton, UK, 1991.

Thomas:1993:STP

[Tho93] R. E. Thomas. SGML tables for the PHIGS slide set. Technical report, Rutherford Appleton Laboratory, Chilton, Didcot, Berks, UK, 1993. 27 pp.

Thomas:1997:BRB

[Tho97]  Lowell Thomas.  Book re-

view: Building VRML worlds: How to plan, design, and construct VRML sites. Java Report: The Source for Java Development, 2(6):25–??, June 1997. CODEN JREPFI. ISSN 1086-4660. URL http://

www.sigs.com/publications/

                                        docs/java/9706/java9706.toc.
                                        html.

Thralls:1995:BHA

[Thr95] Robert Keith Thralls, Sr. Building HTML application systems: Converting existing MS-Windows applications to HTML. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

                                        http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/WebProd/
                                        thralls/WhitePaper.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Tittel:1995:FWW

[Tit95a]  Ed Tittel.  Foundations of

World Wide Web programming with HTML and CGI. I D G Books Worldwide, Indianapolis, IN, USA, 1995. ISBN 1-56884-703-3. xliii + 648 pp. LCCN TK 5105.888 F68 1995.

Tittel:1995:HDa

[Tit95b] Ed Tittel. HTML for Dummies. I D G Books Worldwide, Indianapolis, IN, USA, 1995. ISBN 1-56884-330-5. xl + 452 pp. LCCN QA76.76.H94T57 1995.

Tittel:1995:HDb

[Tit95c] Ed Tittel. HTML for Dummies. American Printing House for the Blind, Louisville, lte edition, 1995. ISBN 0-614-20551-4. 560 pp. LCCN ???? US$140.00.

Tittel:1998:WPX

[Tit98] Ed Tittel. What’s the point of XML? find out what this much discussed, but little understood technology can do for you. Sunworld Online, 12(2):??, February 1998. ISSN 1091-8914.

URL http://www.sun.com/
sunworldonline/swol-02-1998/
swol-02-xml.html.

Tittel:1996:HD

[TJ96a] Ed Tittel and Steve James.

HTML for dummies. IDG Books, San Mateo, CA, USA, second edition, 1996. ISBN 1-56884-647-9. xx + 428 pp. LCCN QA76.76.H94T58 1996. US$29.99, CDN$42.99, UK£28.99. Includes floppy.

Tittel:1996:MHD

[TJ96b] Ed Tittel and Steve James.

More HTML for Dummies. I D G Books Worldwide, Indianapolis, IN, USA, May 1996. ISBN 1-56884-996-6. xxvi + 390 pp. LCCN QA76.76.H94T59 1996. US$22.99.

Tittel:1997:HD

[TJ97a] Ed Tittel and Steven N. James.

HTML for dummies. For dummies. IDG Books, San Mateo, CA, USA, third edition, 1997. ISBN 0-7645-0214-X. xxiv + 407 pp. LCCN QA76.76.H94 T58 1997. Includes CD-ROM.

Tittel:1997:MHD

[TJ97b] Ed Tittel and Steven N. James.

More HTML for dummies. –For dummies. IDG Books, San Mateo, CA, USA, second edition, 1997. ISBN 0-7645-0233-6. xxvi + 390 pp. LCCN QA76.76.H94 T59 1997.

Tittel:1998:HD

[TJ98] Ed Tittel and Stephen N. James. HTML 4 for dummies. –For dummies. IDG Books, San Mateo, CA, USA, 1998. ISBN 0-7645-0331-6 (paperback). xxiv + 373 pp. LCCN QA76.76.H94T568 1998.

URL http://lcweb.loc.gov/
catdir/toc/98-70136.html.

Tittel:1998:HPN

[TJD98] Ed Tittel, Stephen Nelson James, and Michel Dreyfus. HTML 4 pour les nuls. Pour les nuls. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1998. ISBN 2-7361-2962-8. 401 pp. LCCN ????

Toyoda:1990:SDS

[TKB90] Kiyoshi Toyoda, Eri Kumagai, and Tatsuo Bando. SGML document structure editor. National Technical Report, 36(5): 89–95 (or 597–603??), October 1990. CODEN NTROAV. ISSN 0028-0291.

Tittel:1998:XD

[TMC98] Ed Tittel, Norbert Mikula, and Ramesh Chandak. XML for dummies. –For dummies. IDG Books, San Mateo, CA, USA, 1998. ISBN 0-7645-0360-X. xxviii + 367 pp. LCCN QA76.76.H94T595 1998.

Theodoridis:1998:SPI

[To98] S. Theodoridis and other, editors. Signal processing IX: theories and applications: proceedings of EUSIPCO-98, Ninth European Signal Processing Conference, Rhodes, Greece, 8–11 September 1998, volume 2. Typorama Publications, Patras, Greece, 1998. ISBN 960-7620-05-4, 960-7620-07-0. LCCN TK5102.9.E97 1998. Four volumes.

Towheed:1995:NUW

[Tow95] Syed S. Towheed. NASA’s use of the World Wide Web to deliver Shoemaker-Levy 9 collision data in near-real time. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Astronomy/towheed/towheed.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Toyama:1998:SES

[Toy98] Motomichi Toyama. Super-SQL: An extended SQL for database publishing and presentation. SIGMOD Record (ACM Special Interest Group on Management of Data), 27 (2):584–586, June 1998. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Tittel:1999:HD

[TPM99] Ed Tittel and Natanya Pitts-Moultis. HTML 4 for dummies. –For dummies. IDG Books, San Mateo, CA, USA, second edition, 1999. ISBN 0-7645-0572-6 (paperback). xxii + 408 pp. LCCN ????

Tjoa:1992:DDE

[TR92] A. M. Tjoa and I. Ramos, editors. DEXA 92. Database and Expert Systems Applications Proceedings of the International Conference. Springer-Verlag Inc., New York, NY, USA, 1992. ISBN 3-211-82400-6. LCCN QA76.9.D32D363 1992.

Trasher:1997:BWA

[Tra97] Sloan Trasher. Building Web applications using VBScript. New Riders Publishing, Carmel, IN, USA, 1997. ISBN 1-56205-666-2. ???? pp. LCCN QA76.73.V27T73 1997.

Tamiosso:1997:BIA

[TRMR97] F. Tamiosso, A. Raposo, L. Magalhaes, and I. Ricarte. Building interactive animations using VRML and Java. In de Figueiredo and Netto [dN97], pages 42–48. ISBN 0-8186-8102-0, 0-8186-8103-9, 0-8186-8104-7. LCCN T385 .S59 1997. IEEE Computer Society order number PR08102. IEEE order plan catalog number 97TB100164.

Toren:1999:HBP

[TRMW99] B. J. Toren, M. Randall, S. Murphy, and M. J. Whitney. HTML for beginners (public services emphasis). In Bullington et al.

[BCG99], pages 23–26. ISBN 0-7890-0768-1. ISSN 0361-526X. LCCN Z692.S5N67 1998.

Thalman:1996:WFI

[TS96] N. M. Thalman and V. Skala, editors. WSCG 96. Fourth International Conference in Central Europe on Computer Graphics and Visualization 96, in cooperation with IFIP Working Group 5.10 on Computer Graphics and Virtual Worlds. Conference Proceedings. Univ. West Bohemia, Plzen, Czech Republic, 1996. ISBN 80-7082-238-4. LCCN ???? 2 vol.

Tittel:1997:HPG

[TS97] Ed Tittel and James Michael Stewart. The hip pocket guide to HTML 3.2. I D G Books Worldwide, Indianapolis, IN, USA, 1997. ISBN 0-7645-8017-5 (paperback). xx + 234 pp. LCCN QA76.76.H94T56 1997.

Tolstykh:1999:CLO

[TS99] V. N. Tolstykh and I. V. Shturtz.

Creating large optimized VRML worlds. In Melker [Mel99], pages 220–223. ISBN 0-8194-3161-3. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.3687.

Tittel:1998:HPG

[TSP98] Ed Tittel, James Michael Stewart, and Natanya Pitts. The hip pocket guide to HTML 4. I D G Books Worldwide, Indianapolis, IN, USA, 1998. ISBN 0-7645-3196-4. xx + 219 pp. LCCN QA76.76.H94T563 1998.

Tittel:1996:BVW

[TSWS96] Ed Tittel, Charlie Scott, Paul Wolfe, and Claire Sanders. Building VRML Worlds. Osborne/McGraw-Hill, Berkeley, CA, USA, 1996. ISBN 0-07-882233-5. xxiii + 381 pp. LCCN QA76.76.H94B85 1997. US$39.95.

Taniguchi:1995:DEL

[TSY^{+}95] T. Taniguchi, Y. Sawada, T. Yoshida, T. Kakimoto, M. Nagao, M. Harada, and T. Ishikawa. Development of the electronic library ‘Ariadne’. 2. the input and editing of information in Ariadne. Joho kanri = Information & documentation, 38(4):324–337, July 1995. CODEN JOKAAB. ISSN 0021-7298.

Testani:1998:CET

[TTG98] S. Testani, J. Tumasz, and S. Gnanamgari. CIMBLE: Experiential training in a VRML environment. In Landauer and Bellman [LB98a], pages 67–71. ISBN 1-56555-137-0. ISSN 0735-9276. LCCN QA76.9.V5 V578 1998.

Tanigawa:1993:GC

[TTN93] H. Tanigawa, S. Tsuji, and Y. Nakano. Groupware CASE. National Technical Report, 39 (1):72–79, February 1993. CO-DEN NTROAV. ISSN 0028-0291.

Tresorukov:1997:ASG

[TTS97] V. N. Tresorukov, A. A.

Terekhov, and V. A. Sobolev. Application of Standard General Markup Language (SGML) in a library environment. In Segbert et al. [SSB97], pages 153–156. ISBN 92-828-0305-8. LCCN ????

Tubbs:1995:PFD

[Tub95] Dave Tubbs. Pop, fizz, and dance — Java and VRML are new graphics technologies that will alter the state of the Net. Internet World, 6(11):72, ???? 1995. CODEN IERNE8. ISSN 1064-3923.

Tucker:1999:UIE

[Tuc99] Andrew Tucker. Using Internet Explorer’s HTMLParser. Dr. Dobb’s Journal of Software Tools, 24(8):82, 84–87, August 1999. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ftp/1999/1999_
08/htmlpars.txt;   http://
www.ddj.com/ftp/1999/1999_
08/htmlpars.zip.

Tusa:1997:UWI

[Tus97] S. Tusa. Untangle the Web:

Introduction to browsers and HTML. In Christiansen and Leathem [CL97], pages 251–254. ISBN 0-7890-0324-4. ISSN 0361-526X. LCCN Z692.S5N67 1996.

Travis:1995:SIG

[TW95] Brian E. Travis and Dale C.

Waldt. The SGML Implementation Guide: A Blueprint for SGML Migration. Springer-Verlag Inc., New York, NY, USA, 1995. ISBN 0-387-57730-0 (New York), 3-540-57730-0 (Berlin). xxvi + 522 pp. LCCN QA76.76.H94 T73 1995. US$49.50.

Tjoa:1998:PNI

[TW98] A Min Tjoa and Roland R. Wagner, editors. Proceedings: Ninth International Workshop on Database and Expert Systems Applications, August 26– 28, 1998, Vienna, Austria, volume 9. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8353-8, 0-8186-8355-4. LCCN QA76.9.D3 D15592 1998. IEEE Computer Society Order number PR08353.

Ubell:1994:MED

[Ube94] Michael Ubell. The Montage extensible DataBlade architecture. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):482, June 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ubois:1996:WMS

[Ubo96] Jeff Ubois. Wave Maker Starwave VP Patrick Naughton — who helped create Java at Sun — offers some strong opinions on Java, Visual Basic, VRML, Internet over cable, and more. Internet World, 7(3):96–??, ???? 1996. CODEN IERNE8. ISSN 1064-3923.

Udell:1995:PM

[Ude95a] Jon Udell. Perl magic. BYTE Magazine, 20(12):115–??, December 1, 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Udell:1995:PMO

[Ude95b] Jon Udell. Perl magic — our Webmeister explains the Perl and HTML programming tricks unde the hood of the new Virtual Press Room. BYTE Magazine, 20(12):115–??, ???? 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Uhler:1995:IRT

[Uhl95] Stephen Uhler. Incorporating real-time audio on the Web. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
DDay/uhler/uhler.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Ulmer:1998:GMV

[UL98] H. Ulmer and C. Lindenbeck.

Geology meets virtual reality: VRML visualization server applications (D). In Skala [Ska98], pages 402–408. ISBN 80-7082-359-3 (vol. 1) 80-7082-360-7 (vol. 2) 80-7082-361-5 (vol. 3) 80-7082-362-3 (vol. 4). LCCN ???? Four volumes.

UTKCAS:1999:IH

[Uni99] University of Tennessee, Knoxville, Computing and Administrative Systems. Introduction to HTML, 1999. URL http://

                                        www.cas.utk.edu/html/default.
                                        html.

Unsworth:1996:SHC

[Uns96] J. Unsworth. Software for humanities computing: Image annotation, Unicode, and SGML packages from the Institute for Advanced Technology in the Humanities. In Anonymous [Ano96f], pages 297–?? ISBN 82-7283-083-3. ISSN 0804-8800. LCCN ????

USENIX:1987:PSU

[USE87] USENIX Association, editor. Proceedings of the Summer 1987 USENIX Conference: June 8– 12, 1987, Phoenix, Arizona, USA. USENIX Association, Berkeley, CA, USA, 1987. ISBN ???? LCCN ????

USENIX:1989:PWU

[USE89] USENIX Association, editor. Proceedings of the Winter 1989 USENIX Conference: January 30–February 3, 1989, San Diego, California, USA. USENIX Association, Berkeley, CA, USA, 1989. ISBN ???? LCCN ????

USENIX:1995:PNS

[USE95] USENIX Association, editor. Proceedings of the Ninth Systems Administration Conference (LISA IX): September 17–22, 1995, Monterey, CA, USA. USENIX Association, Berkeley, CA, USA, 1995. ISBN 1-880446-73-1. LCCN QA 76.76 O63 S97 1995.

USENIX:1996:PUA

[USE96] USENIX Association, editor.

Proceedings of the USENIX 1996 annual technical conference: January 22–26, 1996, San Diego, California, USA, USENIX Conference Proceedings 1996. USENIX Association, Berkeley, CA, USA, 1996. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

USENIX:1997:ATT

[USE97a] USENIX, editor. 5th Annual Tcl/Tk Workshop ’97, July 14–17, 1997. Boston, MA. USENIX Association, Berkeley, CA, USA, July 14–17, 1997. ISBN 1-880446-87-1. LCCN QA76.73.T44 T44 1997.

USENIX:1997:PUS

[USE97b] USENIX, editor. Proceedings of the USENIX Symposium on Internet Technologies and Systems: December 8– 11, 1997, Monterey, California. USENIX Association, Berkeley, CA, USA, 1997. ISBN 1-880446-91-X. LCCN TK5105.875.I57 U96 1997.

USENIX:1998:PSA

[USE98] USENIX, editor. Proceedings of the sixth annual Tcl/Tk Conference: September 18–24 [i.e. 14–18], 1998, San Diego, California. USENIX Association, Berkeley, CA, USA, 1998. ISBN 1-880446-98-7. LCCN QA76.73.T44 T34 1998.

Ueno:1997:BBV

[USO97] K. Ueno, K. Suzuki, and H. Ozawa. Book based view and scroll based view: Which is suited for HTML viewers? In Salvendy et al. [SSK97], pages 795–798. ISBN 0-444-82183-X. ISSN 0921-2647. LCCN TA166 .A38 v.21A.

Uzawa:1992:DID

[Uza92] A. Uzawa. Document image database systems. Journal of the Information Processing Society of Japan = Joho Shori, 33(5): 497–504, ???? 1992. CODEN JOSHA4. ISSN 0447-8053.

Vijayaraman:1996:PTS

[V^{+}96] T. M. Vijayaraman et al., editors. Proceedings of the twentysecond international Conference on Very Large Data Bases, September 3–6, 1996, Mumbai (Bombay), India. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1996. ISBN 1-55860-382-4. LCCN QA76.9D3 I559 1996.

Vacca:1996:VBV

[Vac96a] John Vacca. VRML: Bringing Virtual Reality to the Internet. Academic Press, New York, USA, May 1996. ISBN 0-12-709910-7, 0-12-709911-5 (with CD-ROM). xxxv + 552 pp. LCCN T385.V322 1996. US$34.95.

Vacca:1996:WWT

[Vac96b] John R. Vacca. 3D worlds on the Web — thanks to the stabilization of the Virtual Reality Modeling Language (VRML), the tools needed to create and view 3D worlds on the Web are becoming more widely available. Computer Graphics World, 19 (5):43–??, ???? 1996. CODEN CGWODH. ISSN 0271-4159.

Vacca:1998:VCE

[Vac98] John R. Vacca. VRML Clearly Explained. AP Professional, Boston, MA, USA, second edition, 1998. ISBN 0-12-710008-3, 0-12-710009-1 (CD-ROM). xxvi + 822 pp. LCCN T385.V3223 1998. US$44.95. Includes CD-ROM.

Valauskas:1996:BCS

[Val96]  E. J. Valauskas.  Britan-

nica CD 2.0 sets new standards. Database, 19(1):70–73, February–March 1996. CODEN DTBSDQ. ISSN 0162-4105.

vanVliet:1988:DMT

[van88] J. C. van Vliet, editor. Document Manipulation and Typography. Proceedings of the International Conference on Electronic Publishing, Document Manipulation and Typography, Nice (France), April 20– 22, 1988. Cambridge University Press, New York, NY, USA, 1988. ISBN 0-521-36294-6. LCCN Z286.E43 I571 1988. US$47.50.

VanHerwijnen:1989:UTI

[Van89] Eric Van Herwijnen. The use of text interchange standards for submitting physics articles to journals. Computer Physics Communications, 57(1–3):244– 250, December 2, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/001046558990221X.

International Conference on Computing in High-Energy Physics.

vanderLaan:1990:ATR

[van90a] C. G. van der Laan. Announcing two reports: SGML-L^{A}TEX and journal style guidelines. TUGboat, 11(1):86, April 1990. ISSN 0896-3207.

vanHerwijnen:1990:PS

[van90b] Eric van Herwijnen. Practical SGML. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0635-X, 0-7923-9434-8. xviii + 307 pp. LCCN QA76.73.S44 V36 1990. UK£24.90, US$49.00.

vanderLaan:1991:STD

[van91a] Kees van der Laan. SGML (, TEX and …). TUGboat, 12(1): 90–104, March 1991. ISSN 0896-3207.

VanEss-Dykema:1991:PSE

[Van91b] Carol Van Ess-Dykema. Practical SGML: Eric van Herwijnen. Computational Linguistics, 17(1):110–??, March 1991. CO-DEN AJCLD9. ISSN 0891-2017.

VanHalm:1993:HHS

[Van93a] J. Van Halm, editor. Hypermedia and hypertext standards: Workshop — April 1993, Amsterdam, volume 13 (2) of Information services and Use 1993. IOS Press, Amsterdam, The Netherlands, 1993. ISBN ???? ISSN 0167-5265. LCCN ????

vanHerwijnen:1993:ST

[van93b] Eric van Herwijnen. SGML tutorial, 1993. 1 computer disk.

vanHerwijnen:1994:PS

[van94] Eric van Herwijnen. Practical SGML. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, second edition, 1994. ISBN 0-7923-9434-8. xx + 288 pp. LCCN QA76.73.S44 V36 1994.

vanBrakel:1995:EJP

[van95a] P. A. van Brakel. Electronic journals: publishing via Internet’s World Wide Web. Electronic Library, 13(4):389–396, August 1995. CODEN ELLIDZ. ISSN 0264-0473.

Vandoni:1995:CSC

[Van95b] C. E. Vandoni, editor. 1995 CERN School of Computing: Arles, France, 20 August–2 September 1995: proceedings, volume 95-05 of CERN. CERN, Geneva, Switzerland, 1995.

ISBN 92-9083-076-X. LCCN QA76 .C101 1995.

VanHeyningen:1995:UCS

[Van95c] Marc D. VanHeyningen. The unified computer science technical report index: Lessons in indexing diverse resources. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
DDay/vanheyningen/paper.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Vandergrift:1998:CPV

[Van98] David Vandergrift. Cal Poly’s virtual Performing Arts Center: an everyday use of 3D visualization with VRML 2.0. Thesis (b.s.), California Polytechnic State University, San Luis Obispo, CA, USA, 1998. iv + 24 pp.

Varian:1995:SFA

[Var95] Hal Varian. Some FAQs about usage-based pricing. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
ComEc/varian/useFAQs.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Vaughan:1994:ISD

[Vau94] G. Vaughan. Integrating SGML and data management technologies. In Anonymous [Ano94e], pages 1145–1214. ISBN 1-871802-32-6. LCCN ????

vonBultzingsloewen:1989:OSQ

[vB89] G. von Bultzingsloewen. Optimizing SQL queries for parallel execution. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(4): 17–22, December 1989. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Verkoulen:1995:SHV

[VB95a] Peter A. C. Verkoulen and Henk M. Blanken. SGML/ HyTime voor de ondersteuning van het gezamenlijke ontwikkelen van multimediaapplicaties. Informatie, 37(7):470–??, 1995. CODEN INFTCR. ISSN 0019-9907.

Voelker:1995:MIS

[VB95b] G. M. Voelker and B. N. Bershad. Mobisaic: an information system for a mobile wireless computing environment. In IEEE [IEE95g], pages 185–190. ISBN 0-8186-6345-6. LCCN QA76.59.W67 1994. IEEE catalog no. 94TH06734.

Vollhardt:1995:MGW

[VB95c] H. Vollhardt and J. Brickmann.

3D molecular graphics on the World Wide Web. In Hunter and Klein [HK95b], pages 663–673.

ISBN 981-02-2578-4. LCCN QH323.5.P33 1996.

Viegnes:1996:HJB

[VB96] Laurent Viegnes and Stéphane Boix. HTML & JAVA: bien débuter. Edition Micro Application, Paris, France, 1996. ISBN 2-7429-0733-5. 358 pp. LCCN ???? 59 FF. URL http://www.

microapp.com/. Includes CD-

ROM.

Vanegas:1996:CCE

[VC96] Jorge Vanegas and Paul Chinowsky, editors. Computing in civil engineering: proceedings of the third congress held in conjunction with A/E/C Systems ’96, Anaheim, California, June 17–19, 1996. American Society of Civil Engineers, New York, NY, USA, 1996. ISBN 0-7844-0182-9. LCCN TA345 .C649 1996.

Vitali:1997:EHP

[VCB97] Fabio Vitali, Chao-Min Chiu, and Michael Bieber. Extending HTML in a principled way with displets. Computer Networks and ISDN Systems, 29(8– 13):1115–1128, September 30, 1997. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http://www.

elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1997&
volume=29&issue=08-13&aid=
1686.

vanderLaan:1989:SLM

[vCL89] C. G. van der Laan, D. C. Coleman, and J. R. Luyten. SGML– L^{A}TEX 1. Mathematical formulas. RC-Rapport 24, Rekencentrum der Rijksuniversiteit te Groningen, Groningen, The Netherlands, 1989. (English version).

Vandoni:1990:EPE

[VD90] C. E. Vandoni and D. A. Duce, editors. Eurographics ’90. Proceedings of the European Computer Graphics Conference and Exhibition. North-Holland Publishing Co., Amsterdam, The Netherlands, 1990. ISBN 0-444-88683-4. LCCN ????

Verdejo:1998:VCT

[VD98] Felisa Verdejo and Gordon Davies, editors. The virtual campus: trends for higher education and training: IFIP TC3/WG3.3 & WG3.6 Joint Working Conference on the Virtual Campus: Trends for Higher Education and Training, 27–29 November 1997, Madrid, Spain. Chapman and Hall, Ltd., London, UK, 1998. ISBN 0-412-83550-9. LCCN LC5803.C65I353 1997.

Venditto:1996:DT

[Ven96a] Gus Venditto. Dueling tools. Internet World, 7(4):36–??, April 1, 1996. CODEN IERNE8. ISSN 1064-3923.

VentanaDevelopmentStaff:1996:VPP

[Ven96b] Ventana Development Staff. VRML Power Publishing, and Online Companion. Ventana Communications Group, Incorporated, Research Triangle Park, NC, USA, July 1996. ISBN 1-56604-450-2. ???? pp. LCCN ???? US$49.95.

Venditto:1997:MFH

[Ven97] Gus Venditto. Microsoft Frontpage 97: HTML and Beyond. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel:

(503) 282-5215, February 1997. ISBN 1-55828-527-X. xix + 411 pp. LCCN TK5105.888.V46 1996. US$29.95.

vanOssenbruggen:1995:WAS

[vES95] J. van Ossenbruggen, A. Eliëns, and B. Schönhage. Web applications and SGML. Electronic Publishing—Origination, Dissemination, and Design, 8(2/3): 51–62, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Veziroglu:1996:HEW

[Vez96] T. Nejat Veziroglu, editor. Hydrogen energy progress XI: proceedings of the 11th World Hydrogen Energy Conference, Stuttgart, Germany, 23–28 June 1996, volume 11 of World Hydrogen Energy Conference. International Association for Hydrogen Energy, ????, 1996. ISBN ???? LCCN ???? Three volumes.

Vizine-Goetz:1995:SWB

[VGGB95] Diane Vizine-Goetz, Jean Godby, and Mark Bendig. Spectrum: a Web-based tool for describing electronic resources. Computer Networks and ISDN Systems, 27 (6):985–1001, April 3, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

URL http://www.elsevier.
com/cas/tree/store/comnet/
sub/1995/27/6/1459.pdf.

Huu:1987:ESD

[vH87] Le van Huu. An environment for SGML document preparation. In USENIX Association [USE87], pages 43–52. ISBN ???? LCCN ????

Varela:1995:PDW

[VH95] Carlos A. Varela and Caroline C. Hayes. Providing data on the Web: From examples to programs. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/DDay/
varela/paper.html. A selected

few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

ViaGrafix:1997:PHIb

[Via97a] ViaGrafix. Programming in HTML. intermediate, 1997. Includes CD-ROM.

ViaGrafix:1997:PHIa

[Via97b] ViaGrafix. Programming in HTML. introduction, 1997. Includes CD-ROM.

ViaGrafix:1997:PHS

[Via97c] ViaGrafix. Programming in HTML standard 3.2. advanced, 1997. 1 videocassette (VHS).

ViaGrafix:1997:TYHb

[Via97d]  ViaGrafix.  Teach yourself

HTML 3.2. advanced, 1997. 1 videocassette (60 min.).

ViaGrafix:1997:TYHa

[Via97e]  ViaGrafix.  Teach yourself

HTML 3.2. introduction, 1997. 1 videocassette (55 min.).

ViaGrafix:1998:PH

[Via98a] ViaGrafix. Programming in HTML 4, 1998. Includes six CD-ROMs and six videocassettes.

ViaGrafix:1998:PHC

[Via98b] ViaGrafix. Programming in HTML 4 [CD-ROM tutorial for windows]. vol. 1-2, 1998. Includes two CD-ROMs.

Vignaud:1989:LSD

[Vig89] Dominique Vignaud, editor. L’Edition structurée des documents: SGML: applications a l’édition française. Editions du Cercle de la librairie, Paris, France, 1989. ISBN 2-7654-0420-8. LCCN ????

Vincent:1998:EHD

[Vin98] B. Vincent. Enhancing HTML documents with ActiveX[TM].

In Anonymous [Ano98u], pages 315–?? ISBN 1-891709-00-3, 1-891709-01-1. LCCN ????

Vint:1999:SAW

[Vin99] Danny R. Vint. SGML at work.

Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-636572-8. xvi + 845 pp. LCCN QA76.76.H94V56 1999.

Vogl:1998:VAW

[VMM98] S. Vogl, K. Manske, and M. Muehlhaeuser. VRML approach to Web video browsing [3654-25]. In Kandlur et al. [KJR98], pages 276–285. ISBN 0-8194-3125-7. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.3654.

Vanouplines:1995:OIS

[VN95] P. Vanouplines and P. Nieuwenhuysen. Online information sources about astronomy: the Chiron case. In Raitt and Jeapes [RJ95], pages 311–325. ISBN 0-904933-94-6. LCCN ????

Vodnik:1999:DHI

[Vod99] Sasha Vodnik. Dynamic HTML:

illustrated introductory. Course Technology, Cambridge, MA, USA, 1999. ISBN 0-7600-6079-7 (paperback). various pp. LCCN ????

Volk:1997:MTS

[Vol97]  M. Volk.  Markup of a

test suite with SGML. In Nerbonne [Ner97], pages 59– 76. ISBN 1-57586-093-7 (hardback), 1-57586-092-9 (paperback). LCCN P128.D37 L56 1997.

VonHagen:1997:SD

[Von97] Bill Von Hagen. SGML for Dummies. I D G Books Worldwide, Indianapolis, IN, USA, 1997. ISBN 0-7645-0175-5. xxiv + 383 pp. LCCN QA76.76.H94V66 1997. US$29.99. Includes CD-ROM.

Venkatrao:1995:SCN

[VP95] Murali Venkatrao and Michael Pizzo. SQL/CLI — A new binding style for SQL. SIG-MOD Record (ACM Special Interest Group on Management of Data), 24(4):72–77, December 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

VRML:1997:VIS

[VRM97] VRML 97, International Specification ISO/IEC IS 14772-1, December 1997. URL http://www/

                                        vrml.org/.

Anonymous:1998:VDJ

[VRM98] VRML developer’s journal, page various, 1998. ISSN 1086-8550. Sys-Con Publications, Jersey City, NJ, USA.

Vekar:1995:CVP

[VS95] John Vekar and Saul Schwartz. The Canadian virtual policy and administration community. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/

Campus.Infosys/vekar/vekar.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

vanOostendorp:1993:CPD

[vSSW93] H. van Oostendorp, B. Scholten, E. Schultheiss, and M. Walsteijn. Customised publishing: developments and an empirical study. Informatie, 35(4):280– 288, April 1993. CODEN IN-FTCR. ISSN 0019-9907.

Vetter:1994:MWW

[VSW94] Ronald J. Vetter, Chris Spell, and Charles Ward. Mosaic and the World Wide Web. Computer, 27(10):49–57, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vakaloudis:1998:SQO

[VT98] A. Vakaloudis and B. Theodoulidis.

The storage and querying of 3D objects for the dynamic composition of VRML worlds. In Banissi et al. [BKS98], pages 267–273. ISBN 0-8186-8509-3, 0-8186-8510-7, 0-8186-8511-5. ISSN 1093-9547. LCCN TA1634.I132 1998. IEEE catalog number 97TB1200165.

Vakaloudis:1999:MQD

[VT99] A. Vakaloudis and B. Theodoulidis.

Modelling, querying and developing spatiotemporal applications based on VRML. In Skala [Ska99], pages 118–125.

ISBN 80-7082-490-5 (set), 80-7082-494-8 (vol. 1), 80-7082-495-6 (vol. 2), 80-7082-496-4 (vol. 3), 80-7082-525-1 (vol. 4). LCCN ????

Vanier:1994:SP

[VTW94] Dana J. Vanier, J. Russell Thomas, and Jamie L. Worling. Standards processing 2000. In Khozeimeh [Kho94], pages 137– 144. CODEN CCENEX. ISBN 0-7844-0026-1. LCCN TA345 .C657 1994. Two volumes.

Vulpe:1997:SSM

[Vul97] M. Vulpe. SGML: Show me which road I’m on. In Bearman and Trant [BT97], pages 235–242. ISBN 1-885626-13-4. LCCN AM2.M876 1997.

Veraart:1996:SSE

[VW96] V. E. Veraart and S. L. Wright. Supporting software engineering education with a local Web site. In Klee [Kle96], pages 275–279. CODEN SIGSD3. ISBN 0-89791-757-X. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA1 .A86. ACM order number: 457960.

Vossen:1988:EDL

[VY88] Gottfried Vossen and Jim Yacabucci. An extension of the database language SQL to capture more relational concepts. SIGMOD Record (ACM Special Interest Group on Management of Data), 17(4):70–78, December 1988. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wang:1997:IFW

[W^{+}97] Yao Wang et al., editors.

1997 IEEE First Workshop on Multimedia Signal Processing: June 23–25, 1997, Princeton, New Jersey, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-3780-8, 0-7803-3781-6, 0-7803-3782-4. LCCN TK5102.9.I338 1997. IEEE catalog number 97TH8256.

Wakayama:1998:IPI

[W^{+}98a] Toshiro Wakayama et al., editors. Information and process integration in enterprises: rethinking documents. Based on papers presented at the International Working Conference on Information and Process Integration in Enterprises, MIT, Nov. 14–15, 1996, volume SECS 428 of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 0-7923-8059-2. LCCN T58.64 .I46 1998.

Westwood:1998:MMV

[W^{+}98b] James D. Westwood et al., editors. Medicine meets virtual reality: art, science, technology: healthcare (r)evolution. San Diego, California, January 28–31, 1998, volume 50 of Studies in health technology and informatics. IOS Press, Amsterdam, The Netherlands, 1998. ISBN 90-5199-386-2, 4-274-90210-2. ISSN 0926-9630. LCCN W26.5 M45 1998.

Woodruff:1996:IDW

[WAB^{+}96] Allison Woodruff, Paul M. Aoki, Eric Brewer, Paul Gauthier, and Lawrence A. Rowe. An investigation of documents from the World Wide Web. Computer Networks and ISDN Systems, 28 (7–11):963–980, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

                                        tronic).  URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/comnet/cas_sub/
                                        browse/browse.cgi?year=1996&
                                        volume=28&issue=7-11&aid=
                                        1617.

Waegemann:1997:TEH

[Wae97] C. P. Waegemann, editor. Toward an electronic health record Europe ’97: Conference — October 1997, London, Toward and Electronic Health Record Europe 1997. Medical Record Institute, Newton, MA, USA, 1997. ISBN 0-9640667-3-4. LCCN ????

Wagner:1997:SWT

[Wag97] Mitch Wagner. SourceCraft Web tool combines Java, HTML. ComputerWorld, 31(6):51, 53, February 10, 1997. CODEN CMPWAB. ISSN 0010-4841.

Walter:1990:OCA

[Wal90] M. Walter. OSU’s Chameleon architecture: a grammatical approach to translation and DTDs. Seybold Report on Publishing Systems, 20(7):17–23, December 1990. CODEN SRPSD9. ISSN 0736-7260.

Walter:1992:SJT

[Wal92] M. Walter. SGML for journals: toward electronic delivery. Seybold Report on Publishing Systems, 21(18):1, 3–13, 16–19, June 1992. CODEN SRPSD9. ISSN 0736-7260.

Wallmannsberger:1994:PPR

[Wal94] J. Wallmannsberger. PORIDGE:

postmodern rhizomatics in digitally generated environments-do we need a metatheory for W3? Electronic Library, 12(6):345– 351, December 1994. CODEN ELLIDZ. ISSN 0264-0473.

Walter:1995:MIN

[Wal95] M. Walter. Making an Internet newspaper: SGML, HTML and signposts from Wyoming. Seybold Report on Publishing Systems, 24(10):3–6, January 1995. CODEN SRPSD9. ISSN 0736-7260.

Wang:1997:IVM

[Wan97]  Tianxi Wang.  Interactive

VRML modeling and the user interface. Thesis (m.s.), Department of Computer Science, Southern Illinois University at Carbondale, Carbondale, IL, USA, 1997. iv + 50 pp.

Wang:1999:DIX

[Wan99] Yili Wang. Design and implementation of an XML based information retrieval system. ????, ????, ????, 1999. 60 pp.

Weymouth:1995:UAR

[WAO^{+}95] Terry E. Weymouth, Dan Atkins, Gary Olson, Bob Clauer, Atul Prakash, Thomas Finholt, and Craig E. Rasmussen. The Upper Atmospheric Research Collaboratory: UARC. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/CSCW/
weymouth/Weymouth.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Warren:1987:SSS

[War87] P. T. Warren. SGML and style sheets: the implications for electronic document preparation. University Computing, 9 (2):81–86, June 1987. CODEN UNCOET. ISSN 0265-4385.

Ward:1998:TSV

[War98] Chris Ward. Techniques for the simplification of a 3D VRML scene composed of textured triangles. University of Leeds, School of Computer Studies, Leeds, 1998.

Watt:1995:NSV

[Wat95] Peggy Watt. News: Several vendors announce plans for VRML. Network World, page ??, December 11, 1995. ISSN 0887-7661 (print), 1944-7655 (elec-

tronic).  URL http://www.
iquest.net/index.html.

Watson:1996:WSW

[Wat96] J. Watson. Weaving a stronger Web: not WWW but XXX. The Computer Bulletin, 8(pt.3): 20–21, June 1996. CODEN COBUAH. ISSN 0010-4531 (print), 1464-357X (electronic).

Watson:1999:JPW

[Wat99]  Mark Watson.  Java pro-

gramming for Windows: using Microsoft AFC, WFC, and XML. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-516-9. xxiii + 344 pp. LCCN QA76.73.J38W383 1999.

Williams:1996:SVS

[WB96] D. B. Williams and H. W.

Beck. SGML versus semantic data models in building digital agricultural libraries. In Stewman [Ste96e], pages 405–409. ISBN ???? LCCN ????

Wadge:1998:IH

[WBSY98] B. Wadge, G. Brown, M. C.

Schraefel, and T. Yildirim. Intensional HTML. In Munson et al. [MNW98], pages 128– 139. ISBN 3-540-65086-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.575.P63 1998.

Willcocks:1992:S

[WC92] M. Willcocks and F. J. Cave.

SGML. Desktop Publishing Commentary, 8(7):6–13, December 1992. CODEN DP-COEQ. ISSN 0957-3178.

Wilke:1999:HBT

[WCM99] M. I. Wilke, D. Cochenour, and S. Moran. HTML for beginners (technical services emphasis). In Bullington et al. [BCG99], pages 15–22. ISBN 0-7890-0768-1. ISSN 0361-526X. LCCN Z692.S5N67 1998.

Wray:1995:OIMb

[WCP^{+}95a] R. E. Wray III, R. Chong, J. Phillips, S. Rogers, W. Walsh, and J. Laird. Organizing information in Mosaic: A classroom experiment. Computer Networks and ISDN Systems, 28(1–2):167–??, December 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Wray:1995:OIMa

[WCP^{+}95b] Robert E. Wray, III, Ronald Chong, Joseph Phillips, Seth Rogers, William Walsh, and John Laird. Organizing information in Mosaic: A classroom experiment. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/
wray/wrayre-experiment-mosaic94.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Watson:1990:OSA

[WD90] B. C. Watson and R. J. Davis. ODA and SGML: an assessment of co-existence possibilities. Computer Standards and Interfaces, 11(3):169–176, 1990. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Watson:1991:OSA

[WD91] B. C. Watson and R. J. Davis.

ODA and SGML: An assessment of co-existence possibilities. Computer Standards and Interfaces, 11(3):169–176, 1991. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Ward:1995:ESI

[WD95] Pat Ward and Kristopher Davis.

Empowering students in the information age. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Educ/
ward/ward.html. A selected few

of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Wang:1997:UVI

[WDM97] Y. Wang, J. Dong, and H. L.

Marcus. The use of VRML to integrate design and solid freeform fabrication. In Bourell [Bou97b], pages 669–676. ISBN ???? ISSN 1053-2153. LCCN ????

Wesley:1992:SMO

[WE92] T. Wesley and J. Engelen.

SGML — a major opportunity for access to information. In Murphy [Mur92b], pages 593– 597. ISBN ???? LCCN ????

Weakley:1997:IAW

[Wea97]  William R. Weakley.  In-

teractive authoring on the Web: an HTML tutorial using JavaScript. Thesis (m.a.), Illinois State University, Springfield, IL, USA, 1997. iv + 93 pp.

Weber:1995:WSK

[Web95a] Jay Weber. A Webmaster’s starter kit. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

                                        http://www.ncsa.uiuc.edu/
                                        SDG/IT94/Proceedings/DDay/
                                        weber/WeberPaper.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Weber:1995:CWP

[Web95b] Kate Weber. Chapter 6, in which Pooh proposes improvements to Web authoring tools, having seen said tools for the Unix platform. Computer Networks and ISDN Systems, 27(6): 823–829, April 3, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/comnet/
                                        sub/1995/27/6/1465.pdf.

Wegner:1996:VEC

[Weg96] Tom Wegner. VRML Explorer with CD-ROM. Coriolis Group, Scottsdale, AZ, USA, January 1996. ISBN 1-883577-74-8. ???? pp. LCCN ???? US$39.99.

Weiss:1993:DDS

[Wei93] E. H. Weiss. Of document databases, SGML, and rhetorical neutrality. IEEE Transactions on Professional Communication, 36(2):58–61, June 1993. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Weinberger:1994:DIP

[Wei94] David Weinberger. Document imaging and publishing meet SGML: Management and creation lines blur. Advanced Imaging, 9(8):30–??, August 1994. CODEN ADIMEZ. ISSN 1042-0711.

Weibel:1995:WWW

[Wei95] S. L. Weibel. The World Wide Web and emerging Internet resource discovery standards for scholarly literature. Library Trends, 43(4):627–644, Spring 1995. CODEN LIBTA3. ISSN 0024-2594.

Weibel:1996:CLN

[Wei96a] S. Weibel. The changing landscape of network resource description. Library Hi Tech, 14 (1):7–10, ???? 1996. CODEN LIHTD2. ISSN 0737-8831.

Weibel:1996:IFC

[Wei96b] Stuart Weibel. The Internet front: The changing landscape of networked resource description. Library Hi Tech, 14(1):

7–??, 1996. CODEN LIHTD2. ISSN 0737-8831.

Weinman:1997:DWG

[Wei97] Lynda Weinman. Designing Web graphics 2. New Riders Publishing, Carmel, IN, USA, 1997. ISBN 1-56205-715-4. ???? pp. LCCN T385.W4545 1997.

Welz:1995:MBW

[Wel95a] Gary Welz. The media business on the WWW. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
ComEc/welz/welzwwwconf.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Welz:1995:VAF

[Wel95b] Gary Welz. VRML IN ACTION — the first wave of VRML viewers and applications are being demonstrated. Internet World, 6 (11):82–??, ???? 1995. CODEN IERNE8. ISSN 1064-3923.

Westwood:1999:MMV

[Wes99] James D. Westwood, editor. Medicine meets virtual reality: the convergence of physical & informational technologies: options for a new era in healthcare, number 62 in Studies in Health Technology and Informatics 1999. Ohmsha and IOS Press, Tokyo and Amsterdam, 1999. ISBN 90-5199-445-1, 4-274-90273-0. ISSN 0926-9630. LCCN W 26.5 M489-9 1999.

Widom:1989:SSS

[WF89] J. Widom and S. J. Finkelstein. A syntax and semantics for set-oriented production rules in relational database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 18(3):36–45, September 1989. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Widom:1990:SOP

[WF90] Jennifer Widom and Sheldon J. Finkelstein. Set-oriented production rules in relational database systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):259–270, June 1990. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Warrick:1998:VBA

[WF98a] P. A. Warrick and W. R. J.

Funnell. VRML based anatomical teaching (VAT): Work in Progress(Poster). In Laxminarayan and Micheli-Tzanakou [LMT98], pages 71–75. ISBN 0-7803-4973-3, 0-7803-4974-1. LCCN R858.A2I157 1998. IEEE catalog number 98EX188.

Wexler:1998:OMH

[WF98b] Steve Wexler and Brett Foster.

The official Microsoft HTML help authoring kit: understanding, creating, and migrating to Microsoft HTML Help for the Microsoft Windows 95 and Windows NT 4.0 operating systems. WexTech Systems, Redmond, WA, USA, 1998. ISBN 1-57231-603-9. xiii + 298 pp. LCCN QA76.76.H94W49 1998.

Wray:1998:DVE

[WH98] M. Wray and R. Hawkes. Distributed virtual environments and VRML: an event-based architecture. In Anonymous [Ano98s], pages 43–52. ISBN ???? ISSN 0169-7552 (print), 1879-2324 (electronic). LCCN ????

Whang:1992:MIS

[Wha92] Kyu-Yong Whang. Multimedia information system platforms. Korea Information Science Society Review, 10(5):5–9, ???? 1992. CODEN CHKWEN. ISSN 1015-9908.

Willard:1995:WBM

[WHC95] K. E. Willard, J. H. Hallgren, and D. P. Connelly. W3 based medical information systems vs. custom client server applications. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
MedTrack/willard/UMHC_www/
UMHC_www_paper.html. A se-

lected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Whetzel:1996:IWW

[Whe96] J. K. Whetzel. Integrating the World Wide Web and database technology. AT&T Technical Journal, 75(2):38–46, March– April 1996. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

White:1991:SD

[Whi91] M. S. White. Standards for DIP. In Anonymous [Ano91d], pages 13–20. ISBN 0-86353-256-X. LCCN ????

Whitehead:1995:AFEb

[Whi95a] S. D. Whitehead. Auto-FAQ:

An experiment in cyberspace leveraging. Computer Networks and ISDN Systems, 28(1–2): 137–??, December 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Whitehead:1995:AFEa

[Whi95b] Steven D. Whitehead. Auto-FAQ: An experiment in cyberspace leveraging. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

URL http://www.ncsa.uiuc.
edu/SDG/IT94/Proceedings/
Agents/whitehead/whitehead.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

White:1996:HAA

[Whi96a] Bebo White. HTML and the art of authoring for the World Wide Web. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-9691-X. xviii + 314 pp. LCCN QA76.76.H94 W58 1996.

White:1996:UOD

[Whi96b]  C. White.  Using object

databases on the World-Wide Web. Info DB, 10(1):11, February 1996. CODEN IFDBEB. ISSN 0891-6004.

White:1996:JCA

[Whi96c] Greg White. Java command-line arguments. Dr. Dobb’s Journal of Software Tools, 21(2):58–??, 60, 62, 96, 98–99, February 1, 1996. CODEN DDJOEB. ISSN 1044-789X.

White:1996:NLH

[Whi96d] Greg White. A new look for HTML. Online & CDROM review: the international journal of online & optical information systems, 20(2):97–??, April 1, 1996. CODEN ONCDEW. ISSN 0309-314X.

Whitehouse:1996:AVA

[Whi96e] K. Whitehouse. Applications: VRML adds a new dimension to Web browsing. IEEE Computer Graphics and Applications, 16 (4):7–9, July 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Whitney:1996:ASI

[Whi96f] G. Whitney, editor. American Society for Information Science:

The Digital revolution: Midyear meeting — May 1996, San Diego, CA. Information Today, Medford, NJ, USA, 1996. ISBN 1-57387-028-5. LCCN ????

Widom:1994:RIA

[Wid94] Jennifer Widom. Research issues in active database systems: Report from the closing panel at RIDE-ADS ’94. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(3):41–43, September 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wiedling:1994:EPN

[Wie94] H.-P. Wiedling. Erstellung, Pflege und Nutzung von technischen Dokumenten auf der Basis von SGML. In Neubauer [Neu94], pages 231–244. ISBN 3-925474-23-4. ISSN 0721-1058. LCCN Z674.5.G28D48 1993.

Williams:1990:NOM

[Wil90] M. E. Williams, editor. National Online Meeting. Proceedings - 1990. Learned Information, Medford, NJ, USA, 1990. ISBN 0-938734-44-X. LCCN ????

Wilson:1992:CIC

[Wil92]  E. Wilson.  A comparison

of interfaces: computer, designer, and user. In Tjoa and Ramos [TR92], pages 326–331. ISBN 3-211-82400-6. LCCN QA76.9.D32D363 1992.

Williama:1993:NOM

[Wil93a] M. E. Williama, editor. 14th National Online Meeting. Proceedings - 1993. Learned Information, Medford, NJ, USA, 1993. ISBN 0-938734-73-3. LCCN ????

Williams:1993:CD

[Wil93b] Lowell Williams. Compound documents. Dr. Dobb’s Journal of Software Tools, 18(3):32, 34, 38–39, 101, March 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Wilson:1993:CSL

[Wil93c] Eve Wilson. The case for SGML: a law database, hypertext and information retrieval. International Yearbook of Law, Computers, and Technology, 7:59–??, 1993. ISSN 0965-528X.

Wilkes:1994:IPS

[Wil94a] T. Wilkes, editor. IPCC 94 Proceedings. Scaling New Heights in Technical Communication. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, ???? 1994. ISBN 0-7803-1936-2. LCCN T10.5.I54 1994. IEEE Catalog No. 94CH3442-1.

Williams:1994:NOM

[Wil94b] M. E. Williams, editor. 15th National Online Meeting. Proceedings - 1994. Learned Information, Medford, NJ, USA, 1994. ISBN 0-938734-84-9. LCCN ????

Wilson:1994:HLL

[Wil94c] E. Wilson. Hypertext and law librarians. Law librarian (London), 25(1):30–42, March 1994. CODEN LALIE2. ISSN 0023-9275.

Wilder:1995:PWW

[Wil95] C. Wilder. Plugging in to the Web (World Wide Web consultants). Information Week, April 1995. CODEN INFWE4. ISSN 8750-6874.

Wilimek:1996:MHH

[Wil96a] Debbie Wilimek. Mastering HTML for Help Authors. Blue Sky Software Corporation, La Jolla, CA, second edition, April 1996. ISBN 0-9647236-3-8. 110 pp. LCCN ???? US$19.95.

Williamson:1996:IYT

[Wil96b]  S. Williamson.  If yester-

day is tomorrow, this must be Toronto [Internet software]. In Frieder and Wigglesworth [FW96], pages 116– 118. ISBN 0-8186-7389-3, 0-8186-7390-7 (softbound), 0-8186-7391-5 (microfiche). LCCN QA76.76.D47I595 1996. IEEE Catalog No. 96TB100054.

Wiley:1997:CAV

[Wil97a] Wiley Computer Publishing.

CD-ROM to accompany VRML 2.0 sourcebook, 2nd edition, 1997. Includes CD-ROM.

Williams:1997:NOM

[Wil97b] Martha E. Williams, editor. National Online Meeting proceedings–1997: proceedings of the 18th National Online Meeting, New York, May 13– 15, 1997, volume 18 of National Online Meeting proceedings. Information Today, Medford, NJ, USA, 1997. ISBN 1-57387-043-9. LCCN ????

WLBF:1998:BFC

[Wil98] William L.Bryant Foundation, University of Central Florida. The Bryant Foundation collection: an HTML gallery of Caribbean art, 1998.

URL http://www.oir.ucf.
edu/bryant/.

Williams:1999:NOM

[Wil99a] Martha E. Williams, editor. National Online Meeting proceedings–1999: proceedings of the 20th National Online Meeting, New York, May 18–20, 1999, volume 20. Information Today, Medford, NJ, USA, 1999. ISBN 1-57387-084-6. LCCN ????

Williamson:1999:HMR

[Wil99b] Heather Williamson. HTML master reference. I D G Books Worldwide, Indianapolis, IN, USA, 1999. ISBN 0-7645-3256-1 (hardbound ). xi + 1359 pp. LCCN QA76.76.H94W59 1999.

Williamson:1999:WCD

[Wil99c] Heather Williamson. Writing cross-browser Dynamic HTML. Apress, Berkeley, CA, USA, 1999. ISBN 1-893115-05-4. ???? pp. LCCN ????

Winter:1996:HIH

[Win96] C.-J. Winter. Hytime — it is high time for hydrogen! A plea for the introduction of hydrogen transportation: Courageously, prudently, step by step. In Veziroglu [Vez96], pages 3–12. ISBN ???? LCCN ???? Three volumes.

Winter:1997:HEA

[Win97a] C. J. Winter. HYDROGEN ENERGY: After 2 1/2 decades of research — it’s HYTIME for entrepreneurs, bankers, economists. In Bose [Bos97], pages 368–375. ISBN 0-9696869-2-7 (valid ISBN allocated to two books: the 7th and 8th workshop). LCCN QA76.9.S63 I56 1998.

Winter:1997:HIH

[Win97b] C.-J. Winter. HYTIME: it’s high time for a hydrogen energy business forum hydrogen ’96: closing remarks by the science program chairman. International journal of hydrogen energy, 22(9):847–??, ???? 1997. ISSN 0360-3199.

Wittbecker:1986:GTB

[Wit86] Alan E. Wittbecker. Genesis of a TEX-based markup metalanguage. In Conference proceedings: setting the direction / ACM SIGUCCS User Services Conference XIV, Le Centre Sheraton, Montreal, Quebec, September 27th – October 1st, 1986, pages 31–33. ACM Press, New York, NY, USA, 1986.

ISBN 0-89791-208-X. LCCN QA74.A15 1986. URL http://

                                        www.acm.org/pubs/contents/
                                        proceedings/userservices/

324239/. ACM order number 459860.

Withers:1996:SPI

[Wit96a] John Withers. SGML Publishing on the Internet. Ventana Communications Group, Incorporated, Research Triangle Park, NC, USA, February 1996. ISBN 1-56604-366-2. 500 pp. LCCN ???? US$39.95.

Witthaus:1996:MAV

[Wit96b] Nick Witthaus. Mastering Active VRML. Prima Tech Publishing, Rocklin, CA, USA, October 1996. ISBN 0-7615-0742-6. 300 pp. LCCN ???? US$35.00.

Wong:1996:SSM

[WKD96] J. Wong, S. Kini, and K. Doobagunta. Synchronization in specification-based multimedia presentations. Software—Practice and Experience, 26(1):71– 81, January 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wilberg:1996:DEE

[WKV^{+}96] J. Wilberg, A. Kuth, H.-T. Vierhaus, R. Camposano, and W. Rosenstiel. A design exploration environment. In IEEE [IEE96n], pages 77–80. ISBN 0-8186-7502-0. LCCN TK 7874.75 G74 1996. IEEE Catalog No. 96TB100041.

Wonneberger:1992:SQA

[WM92] Reinhard Wonneberger and Frank Mittelbach. SGML — questions and answers. TUGboat, 13(2):221–223, July 1992. ISSN 0896-3207.

Werle:1998:PHH

[WM98] Rainer Werle and Jean-Marc Moster, trad. Programmation HTML 4: en 18 Heures chrono ! Micro Application, Paris, France, 1998. ISBN 2-7429-1198-7. 537 pp. LCCN ????

Walsh:1999:DDG

[WM99] Norman Walsh and Leonard Muellner. DocBook: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-580-7. xii + 635 pp. LCCN QA76.76.H94 W35 1999. US$36.95. URL http:

//www.oreilly.com/catalog/
9781565925809;   http://www.
oreilly.com/catalog/docbook.

Includes CD-ROM.

Weibel:1995:ASPa

[WMGL95a] Stuart Weibel, Eric Miller, Jean Godby, and Ralph LeVan. An architecture for scholarly publishing on the World Wide Web. In Anonymous [Ano95-34], page ???? ISBN

????  LCCN ????  URL

http://www.ncsa.uiuc.edu/
SDG/IT94/Proceedings/Pub/

weibel/weibel_www_paper.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Weibel:1995:ASPb

[WMGL95b] Stuart Weibel, Eric Miller, Jean Godby, and Ralph LeVan. An architecture for scholarly publishing on the World Wide Web. Computer Networks and ISDN Systems, 28(1–2):239– 245, December 1, 1995. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic).  URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1995&
volume=28&issue=1-2&aid=1526.

Woods:1997:WSE

[WNB97] D. J. Woods, A. Norton, and G. Bell. Wired for speed: Efficient routes in VRML 2.0. In Spencer and Brutzman [SB97], pages 133–138. ISBN 0-89791-886-X. LCCN ????

URL http://www.acm.org/
pubs/contents/proceedings/
graph/253437/. ACM order

number 434973.

Wills:1995:PSW

[WNC95] Linda Wills, Philip Newcomb, and Elliot Chickofsky, editors. Proceedings: Second Working Conference on Reverse Engineering: July 14– 16, 1995, Toronto, Ontario, Canada, Working Conference on Reverse Engineering 1995, 2nd. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7111-4. LCCN TA 168.5 W67 1995. IEEE Catalog No. 95TB8101.

Wolfsthal:1991:SCQ

[Wol91] Yaron Wolfsthal. Style control in the Quill document editing system. Software—Practice and Experience, 21(6):625–638, June 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wolf:1995:IOL

[Wol95] K. D. Wolf. The implementation of an open learning environment under World Wide Web. In Maurer [Mau95], pages 689–694. ISBN 1-880094-15-0. LCCN LB1028.5.W67 1995.

Wonneberger:1990:SDP

[Won90] Reinhard Wonneberger. Structured document processing: the L^{A}TEX approach. In Nadrchal [Nad90], pages 177–189. CO-DEN CPHCBZ. Invited paper.

Wonneberger:1992:AST

[Won92] Reinhard Wonneberger. Approaching SGML from TEX. TUGboat, 13(2):226–227, July 1992. ISSN 0896-3207.

Wonneberger:1994:TIE

[Won94] R. Wonneberger. TEX in an industrial environment. Electronic Publishing—Origination, Dissemination, and Design, 7 (1):3–19, March 1994. CODEN EPODEU. ISSN 0894-3982.

Wood:1995:SGM

[Woo95a] D. Wood. Standard Generalized Markup Language: Mathematical and philosophical issues. Lecture Notes in Computer Science, ??(1000):344–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wood:1995:WS

[Woo95b] E. M. Wood. So what is SGML? Journal of systems management, 46(5):24–29, September–October 1995. CO-DEN JSYMA9. ISSN 0022-4839.

Wood:1995:WIS

[Woo95c] Eileen M. Wood. So what is SGML? Journal of systems management, 46(5):24–??, September 1, 1995. CODEN JSYMA9. ISSN 0022-4839.

WWW:1997:AHS

[Wor97] World Wide Web Consortium. Advancing HTML: style and substance, volume 2(1) of World Wide Web journal. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Winter 1997. ISBN 1-56592-264-6. ISSN 1085-2298. vii + 246 pp. LCCN ????

Werkman:1995:SRL

[WP95] Keith Werkman and Jeff Poulin. Software reuse libraries with Mosaic. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN

???? URL http://www.ncsa.
uiuc.edu/SDG/IT94/Proceedings/
DDay/werkman/www94.html. A

selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Wnek:1998:ACS

[WP98] J. Wnek and R. J. Price. Automated conversion of structured documents into SGML [3305-19]. In Lopresti and Zhou [LZ98], pages 141–150. ISBN 0-8194-2745-4. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510.S63 v.3305.

Wu:1994:SSS

[WR94] Gilbert S. K. Wu and Brian Robinson. SGML support for secure document systems. British Library Research and Development report 6158., British Library Research and Development Department, London, UK, 1994. 59 pp.

Whitaker:1995:CPI

[WR95] B. J. Whitaker and H. S. Rzepa.

Chemical publishing via the Internet. In Collier [Col95c], pages 62–71. ISBN ???? LCCN ????

Wallace:1999:HXG

[WR99] Malcolm Wallace and Colin Runciman. Haskell and XML: Generic combinators or typebased translation? ACM SIG-PLAN Notices, 34(9):148–159, September 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.

                                        acm.org/pubs/citations/proceedings/
                                        fp/317636/p148-wallace. Pro-

ceedings of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Wright:1996:IHP

[Wri96] Steve Wright. Instant HTML Programmer’s Reference. Wrox Press, Chicago, IL, USA, December 1996. ISBN 1-86100-076-6. xii + 285 pp. LCCN QA76.76.H94W75 1996. US$15.00.

Watson:1995:CCS

[WS95a] B. C. Watson and K. Shafer. Creating custom SGML DTDs for documentation products. In ACM [ACM95b], pages 189–196. ISBN 0-89791-713-8. LCCN QA 76.9 D6 I57 1995.

Wright:1995:SHT

[WS95b] Sue Ellen Wright and Richard A. Strehlow, editors. Standardizing and harmonizing terminology: theory and practice: Symposium — October 5–6, 1993, volume 1223 of ASTM Special Technical Publication. ASTM, Philadelphia, PA, USA, 1995. ISBN 0-8031-1984-4. ISSN 0066-0558. LCCN T11 .S6765 1995.

Wang:1999:IVM

[WS99] Qiming Wang and Bonita Saunders. Interactive 3D visualization of mathematical functions using VRML. U.S. Dept. of Commerce, Technology Administration, National Institute of Standards and Technology ; U.S. Dept. of Energy, Gaithersburg, MD, USA, 1999. 7 pp.

Wood:1993:LM

[WSM93] N. Wood, B. Schmidt, and J. Myers. Lexical metadatabases. In Raitt and Jeapes [RJ93], pages 485–493. ISBN 0-904933-85-7. LCCN ????

Wakatori:1996:OSI

[WSM96] R. Wakatori, T. Sakairi, and Y. Mano. ODA: standard interchange format for a variety of documents. Joho-Shori (J. Information Processing Soc. Japan), 37(3):199–206, March 1996. CODEN JOSHA4. ISSN 0447-8053.

Wang:1998:IVM

[WSoSS98] Qiming Wang, Bonita Saunders, National Institute of Standards, and Technology (U. S.). Interactive 3D visualization of mathematical functions using VRML. NISTIR; 6289. U.S. Department of Commerce, Technology Administration, National Institute of Standards and Technology, Gaithersburg, MD, USA, 1998. 7 pp.

Wesley:1993:CCC

[WT93] T. Wesley and C. Tobin. CAPS copyright communication and access to information for persons with special needs. SIG-CAPH Newsletter, September 1993. CODEN SGNWD2. ISSN 0163-5727.

Woelk:1995:MIP

[WT95] Darrell Woelk and Christine Tomlinson. The MCC InfoSleuth Project: Intelligent search management via semantic agents. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        Searching/woelk/woelk.html.

A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.

Wu:1989:STP

[Wu89] Gilbert S. K. Wu. SGML theory and practice. British Library research paper, 68 0269-9257. British Library Research and Development Department, London, UK, 1989. ISBN 0-7123-3211-1 (paperback). vii + 92 pp. LCCN ????

Welch:1996:TTH

[WU96] B. Welch and S. Uhler. Tcl/ Tk HTML tools. In Association [Ass96], pages 173–182. ISBN 1-880446-78-2. LCCN QA76.73.T44 T44 1996.

Wu:1998:SBP

[Wu98] Yihe Wu. Simulation of the biological processes of steroid hormones with VRML. ????, ????, 1998. ix + 89 pp.

Wusteman:1996:EJF

[Wus96] J. Wusteman. Electronic journal formats. Program, 30(4):319– 343, October 1996. CODEN PRGMBD. ISSN 0033-0337.

Warmer:1989:IAS

[WV89] J. Warmer and S. Van Egmond.

The implementation of the Amsterdam SGML Parser. Electronic Publishing—Origination, Dissemination, and Design, 2 (2):65–90, July 1989. CODEN EPODEU. ISSN 0894-3982.

Warmer:1991:PSD

[WV91] J. Warmer and H. Van Vliet.

Processing SGML documents. Electronic Publishing—Origination, Dissemination, and Design, 4(1):3–26, March 1991. CODEN EPODEU. ISSN 0894-3982.

Weinman:1998:CHD

[WW98] Lynda Weinman and William Weinman. Creative HTML design. New Riders Publishing, Carmel, IN, USA, 1998. ISBN 1-56205-704-9. xxx + 434 pp. LCCN QA76.76.H94W45 1998.

Weinman:1997:CHD

[WWEI97] Lynda Weinman, William Weinman, Jennifer Eberhardt, and Dayna Isley. Creative HTML Design. New Riders Publishing, Carmel, IN, USA, January 1997. ISBN 1-56205-704-9. xxx + 434 pp. LCCN QA76.76.H94W45 1998. US$39.99.

Walczowski:1996:VDT

[WWN96] L. T. Walczowski, W. A. J.

Waller, and D. Nalbantis. VLSI design training with the help of the World Wide Web. In Anonymous [Ano96u], pages 8/1–3. ISBN ???? LCCN ????

OReilly:1995:WWW

[WWW95] O’Reilly and Associates and Web Consortium (W3C), editors. World Wide Web Journal: The Fourth International WWW Conference Proceedings, pages xii + 735, 1995. ISBN 1-56592-169-0. ISSN 1085-2301. US$39.95. URL http://www.

ora.com/gnn/bus/ora/item/

wj1.html. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800. The World Wide Web Journal is a quarterly publication that provides timely, indepth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

Wendt:1991:CPS

[WZ91] H. Wendt and C. Zurn. Computer-assisted publishing with its standards and compatibility problems. In Anonymous [Ano91a], pages 2/1–3. ISBN ???? LCCN ????

Wagner:1995:QRS

[WZ95] W. P. Wagner and J. S. Zaveri. The quiet revolution of SGML.

In Anonymous [Ano95h], pages 729–?? ISBN ???? LCCN ????

Xu:1997:CSU

[Xu97] Linghua Xu. A comparative study of the use of HTML, Java, perl and C in the implementation of World Wide Web applications. Thesis (m.s.), Major in Engineering/Computer Science, South Dakota State University, Brookings, SD, USA, 1997. ix + 86 pp.

Yamada:1993:CMD

[Yam93] M. Yamada. Conversion method from document image to logically structured document based on ODA. Transactions of the Institute of Electronics, Information and Communication Engineers D-II, J76D-II(11):2274– 2284, November 1993. CODEN DTGDE7. ISSN 0915-1923.

Yamada:1994:CMD

[Yam94] M. Yamada. Conversion method from document image to logically structured document based on open-document architecture (ODA). Systems and computers in Japan, 25(13):47–61, November 1994. CODEN SCJAEP. ISSN 0882-1666 (print), 1520-684X (electronic).

Yee:1999:WFD

[Yee99] Michael Townsend Yee. Web forms and dynamic HTML: simplifying data transfer on the internet. Thesis (m.s.), California Polytechnic State University, San Luis Obispo, CA, USA, 1999. xi + 68 pp.

LiaoYang:1996:SSS

[YF96] Liao Yang and Xu Fupei. SGML supporting system and design of a DTD editor model. Mini-Micro Systems, 17(11):51– 55, November 1996. CODEN XWJXEH. ISSN 1000-1220.

Yagel:1997:VPO

[YH97] Roni Yagel and Hans Hagen, editors. Visualization ’97: proceedings, October 19–24, 1997, Phoenix, Arizona. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 1-58113-011-2 (ACM), 0-8186-8262-0 (IEEE), 0-8186-8262-0, 0-8186-8263-9 (casebound), 0-8186-8264-7. ISSN 1070-2385. LCCN T385 .I192 1997.

Yao:1996:SSS

[YHHYK96] Lin-Ju Yao, Hsiu-Hsen, and Chen Yuan-Kuo. SSQL: a semistructured query language for SGML document retrievals. In ACM [ACM96d], pages 221–228. ISBN 0-89791-799-5. LCCN QA 76.9 D6 I56 1996.

Yoshikawa:1996:ASD

[YIU96] M. Yoshikawa, O. Ichikawa, and S. Uemura. Amalgamating SGML documents and databases. In Apers et al. [ABG96], pages 259–274. CO-DEN LNCSD9. ISBN 3-540-61057-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D3 I55844 1996.

Yamakawa:1990:DPV

[YKT90] T. Yamakawa, Y. Kawabata, and H. Tamura. Desktop publishing viewed from the office automation industries. Journal of the Information Processing Society of Japan = Joho Shori, 31 (11):1508–1517, ???? 1990. CO-DEN JOSHA4. ISSN 0447-8053.

Yergeau:1997:RIH

[YNAD97] F. Yergeau, G. Nicol, G. Adams, and M. Duerst. RFC 2070: Internationalization of the hypertext markup language, January 6, 1997. URL ftp://ftp.

internic.net/rfc/rfc2070.
txt;   http://www.math.utah.
edu/pub/rfc/rfc2070.txt. This

Internet Request-for-Comment document describes an extension of the HTML 2.0 DTD to support the Unicode character set.

Yokoi:1993:VLK

[Yok93] T. Yokoi. A very large-scale knowledge base embodying an intelligence space-technologies for building and evolving knowledge infrastructures. Journal of the Information Processing Society of Japan = Joho Shori, 34 (12):1449–1457, December 1993. CODEN JOSHA4. ISSN 0447-8053.

Yourdon:1996:JWS

[You96] Edward Yourdon. Java, the Web, and software development. Computer, 29(8):25–30, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yamamoto:1996:WIA

[YT96] S. Yamamoto and K. Tokumaru. WebBASE: an Internet application development system. American Programmer, 9(8):27–35, August 1996. CODEN AM-PRFD. ISSN 1048-5600.

Yuzuguchi:1989:DTP

[Yuz89] A. Yuzuguchi. Desk top publishing now in Japan. Keisoku Jido Seigyo Gakkai = Journal of the Society of Instrument and Control Engineers, 28(3):199– 203, March 1989. CODEN KE-SEA4. ISSN 0453-4662.

Young:1996:MDI

[YWJ96] A. Young, D. Woolston, and J. Johnson. Managing divergent information. Object Magazine, 5 (9):50–55, February 1996. CO-DEN OBMAFO. ISSN 1055-3614.

Yamazaki:1997:SPE

[YYOO97] S. Yamazaki, H. Yoshihara, K. Ohe, and K. Ohashi. The standardization of procedures for the exchange of medical information using SGML. In Waegemann [Wae97], pages 263–266. ISBN 0-9640667-3-4. LCCN ????

Zacker:1997:TMG

[Zac97] Craig Zacker. Ten Minute Guide to HTML Style Sheets. Que Corporation, Indianapolis, IN, USA, January 1997. ISBN 0-7897-1034-X. xii + 162 pp.

LCCN QA76.76.H94Z33 1997. US$14.99.

Zagler:1994:CHP

[ZBW94] Wolfgang L. Zagler, Geoffrey Busby, and Roland R. Wagner, editors. Computers for Handicapped Persons. 4th International Conference, ICCHP ’94 Proceedings. Springer-Verlag Inc., New York, NY, USA, 1994. ISBN 3-540-58476-5. LCCN HV1569.5.I56 1994.

Zakour:1997:HHD

[ZFK97] John Zakour, Jeff Foust, and David Kerven. HTML 4 howto: the definitive HTML 4 problem-solver. Waite Group Press, Corte Madera, CA, USA, 1997. ISBN 1-57169-125-1 (paperback). xxi + 1067 pp. LCCN QA76.76.H94Z35 1997.

Zhang:1995:AOS

[Zha95] Jian Zhang. Application of OODB and SGML techniques in text database: An electronic dictionary system. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(1):3–8, March 1995. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ziff-DavisPressDevelopmentStaff:1996:SES

[Zif96] Ziff-Davis Press Development Staff. SGML Elements of Style. Que Corporation, Indianapolis, IN, USA, May 1996. ISBN 1-56276-414-4. ???? pp. LCCN ???? US$29.99.

Zimmermann:1995:ACT

[Zim95] P. Zimmermann. Application of CALS to technical publications for military aircraft. In Anonymous [Ano95e], pages 152–172. ISBN ???? LCCN ????

Zimmerman:1997:WPE

[Zim97] Paul H. Zimmerman. Web page essentials. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 1-57576-368-0. 173 pp. LCCN TK5105.888.Z55 1997.

Zuckerman:1996:JLP

[ZJ96] A. E. Zuckerman and L. B. Jacques. A Java language program for writing clinical cases using custom extensions to HTML. In Cimino [Cim96], pages 906–?? ISBN 1-56053-208-4. LCCN W 26.5 S978 1996.

Zachary:1995:ECC

[ZJEP95] J. L. Zachary, C. R. Johnson, E. N. Eide, and K. W. Parker. An entry-level course in computational engineering and science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(1): 209–213, March 1995. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zucker:1995:EPU

[ZKM95] J. Zucker, R. M. Kahn, and P. Molholt. An electronic PDR using fielded searches in HTML. In Gardner [Gar95], pages 1012– ?? ISBN 1-56053-123-1. ISSN 0195-4210. LCCN W 26.5 S978 1995.

Zimmerman:1996:CAH

[ZMP96] D. E. Zimmerman, M. L.

Muraski, and M. E. Palmquist. A comparative analysis of HTML and Windows-based multimedia software — an exploratory study. In Anonymous [Ano96k], pages 173–184. ISBN 0-7803-3689-5 (softbound), 0-7803-3690-9 (casebound), 0-7803-3691-7 (microfiche). LCCN T10.5.I54 1996. IEEE catalog number 96CH36009.

Zheng:1993:DMC

[ZR93a] Min Zheng and Roy Rada.

Dexter-Groupware model for collaborative authoring. In Maurer [Mau93], pages 383–390. ISBN 1-880094-06-1. LCCN ????

Zheng:1993:SMB

[ZR93b] Min Zheng and Roy Rada.

SHyD — A model for bridging text and hypermedia. In Proceedings — ACM Computer Science Conference Proceedings of the 21st Annual Computer Science Conference Feb 16–18 1993, Indianapolis, IN, USA, pages 418–424. ACM Press, New York, NY, USA, 1993. ISBN ???? LCCN ????

Zheng:1993:SAT

[ZR93c] Min Zheng and Roy Rada. A standard-based approach to text and hypertext mutual conversion and interchange. In Anonymous [Ano93k], pages 116–121 (or 116–122??). ISBN ????

LCCN ???? IEEE catalog no. 93CH3367-0.

Zheng:1993:TMC

[ZR93d] Min Zheng and Roy Rada. Text-hypertext mutual conversion and hypertext interchange through SGML. In Bhargava et al. [BFY93], pages 139–147. ISBN 0-89791-626-3. LCCN QA76.9.D3I5586 1993.

Zimmerman:1995:ICW

[ZY95] Frederick Zimmerman and Peter Young. Information for a changing world: CIESIN’s use of WWW to facilitate access to data and information relevant to human interactions in the environment. In Anonymous [Ano95-34], page ???? ISBN ???? LCCN ????

                                        URL http://www.ncsa.uiuc.
                                        edu/SDG/IT94/Proceedings/
                                        EarSci/zimmerman/zimmerman.

html. A selected few of the 199 papers in these proceedings were republished in the journal Computer Networks and ISDN Systems, 28(1–2), December 1995.